GNU Awk

GAWK: Effective AWK Programming

A User’s Guide for GNU Awk Edition 4.1 April, 2015

Arnold D. Robbins

“To boldly go where no man has gone before” is a Registered Trademark of Paramount Pictures Corporation.

Published by:

Free Software Foundation 51 Franklin Street, Fifth Floor Boston, MA 02110-1301 USA Phone: +1-617-542-5942 Fax: +1-617-542-2652

Email: gnu@gnu.org
URL: http://www.gnu.org/

ISBN 1-882114-28-0

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989, 1991, 1992, 1993, 1996–2005, 2007, 2009–2015 Free Software Foundation, Inc.

This is Edition 4.1 of GAWK: Effective AWK Programming: A User’s Guide for GNU Awk, for the 4.1.2 (or later) version of the GNU implementation of AWK. Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with the Invariant Sections being “GNU General Public License”, with the Front-Cover Texts being “A GNU Manual”, and with the Back-Cover Texts as in (a) below. A copy of the license is included in the section entitled “GNU Free Documentation License”. a. The FSF’s Back-Cover Text is: “You have the freedom to copy and modify this GNU

manual.”

To my parents, for their love, and for the wonderful example they set for me. To my wife, Miriam, for making me complete.Thank you for building your life together with me. To our children, Chana, Rivka, Nachum, and Malka, for enrichening our lives in innumerable ways.

Short Contents

Foreword to the Third Edition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Foreword to the Fourth Edition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Preface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

Part I: The awk Language 1 Getting Started with awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2 Running awk and gawk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 3 Regular Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 4 Reading Input Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 5 Printing Output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 6 Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 7 Patterns, Actions, and Variables . . . . . . . . . . . . . . . . . . . . . . . 131 8 Arrays in awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 9 Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175

Part II: Problem Solving with awk 10 A Library of awk Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 11 Practical awk Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251

Part III: Moving Beyond Standard awk with gawk 12 Advanced Features of gawk . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 13 Internationalization with gawk . . . . . . . . . . . . . . . . . . . . . . . . . 315 14 Debugging awk Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 15 Arithmetic and Arbitrary-Precision Arithmetic with gawk… 341 16 Writing Extensions for gawk. . . . . . . . . . . . . . . . . . . . . . . . . . . 355

Part IV: Appendices A The Evolution of the awk Language. . . . . . . . . . . . . . . . . . . . . 413 B Installing gawk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 C Implementation Notes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 447 D Basic Programming Concepts . . . . . . . . . . . . . . . . . . . . . . . . . 459 Glossary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 463 GNU General Public License . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475 GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . . . . 487 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495

Table of Contents

Foreword to the Third Edition . . . . . . . . . . . . . . . . . . . . 1

Foreword to the Fourth Edition . . . . . . . . . . . . . . . . . . . 3

Preface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

History of awk and gawk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 A Rose by Any Other Name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Using This Book . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Typographical Conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

Dark Corners . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 The GNU Project and This Book. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 How to Contribute . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Acknowledgments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

Part I: The awk Language

1 Getting Started with awk . . . . . . . . . . . . . . . . . . . . . 17

1.1 How to Run awk Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 1.1.1 One-Shot Throwaway awk Programs . . . . . . . . . . . . . . . . . . . . . . 17 1.1.2 Running awk Without Input Files . . . . . . . . . . . . . . . . . . . . . . . . 18 1.1.3 Running Long Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 1.1.4 Executable awk Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 1.1.5 Comments in awk Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 1.1.6 Shell Quoting Issues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

1.1.6.1 Quoting in MS-Windows Batch Files. . . . . . . . . . . . . . . . . 23 1.2 Data files for the Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 1.3 Some Simple Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 1.4 An Example with Two Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 1.5 A More Complex Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 1.6 awk Statements Versus Lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 1.7 Other Features of awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 1.8 When to Use awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 1.9 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

2 Running awk and gawk . . . . . . . . . . . . . . . . . . . . . . . . . 31

2.1 Invoking awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 2.2 Command-Line Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 2.3 Other Command-Line Arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 2.4 Naming Standard Input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 2.5 The Environment Variables gawk Uses . . . . . . . . . . . . . . . . . . . . . . . . . 39

2.5.1 The AWKPATH Environment Variable . . . . . . . . . . . . . . . . . . . . . . 39 iv GAWK: Effective AWK Programming

2.5.2 The AWKLIBPATH Environment Variable . . . . . . . . . . . . . . . . . . . 40 2.5.3 Other Environment Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

2.6 gawk’s Exit Status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 2.7 Including Other Files into Your Program . . . . . . . . . . . . . . . . . . . . . . 42 2.8 Loading Dynamic Extensions into Your Program . . . . . . . . . . . . . . 43 2.9 Obsolete Options and/or Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 2.10 Undocumented Options and Features. . . . . . . . . . . . . . . . . . . . . . . . . 44 2.11 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

3 Regular Expressions. . . . . . . . . . . . . . . . . . . . . . . . . . . 45

3.1 How to Use Regular Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 3.2 Escape Sequences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 3.3 Regular Expression Operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 3.4 Using Bracket Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 3.5 How Much Text Matches? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 3.6 Using Dynamic Regexps. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 3.7 gawk-Specific Regexp Operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 3.8 Case Sensitivity in Matching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 3.9 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

4 Reading Input Files . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

4.1 How Input Is Split into Records . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 4.1.1 Record Splitting with Standard awk . . . . . . . . . . . . . . . . . . . . . . 59 4.1.2 Record Splitting with gawk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

4.2 Examining Fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 4.3 Nonconstant Field Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 4.4 Changing the Contents of a Field. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 4.5 Specifying How Fields Are Separated . . . . . . . . . . . . . . . . . . . . . . . . . . 67

4.5.1 Whitespace Normally Separates Fields . . . . . . . . . . . . . . . . . . . . 68 4.5.2 Using Regular Expressions to Separate Fields . . . . . . . . . . . . . 68 4.5.3 Making Each Character a Separate Field . . . . . . . . . . . . . . . . . 69 4.5.4 Setting FS from the Command Line . . . . . . . . . . . . . . . . . . . . . . 69 4.5.5 Making the Full Line Be a Single Field . . . . . . . . . . . . . . . . . . . 71 4.5.6 Field-Splitting Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

4.6 Reading Fixed-Width Data. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 4.7 Defining Fields by Content . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 4.8 Multiple-Line Records. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 4.9 Explicit Input with getline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

4.9.1 Using getline with No Arguments . . . . . . . . . . . . . . . . . . . . . . . 78 4.9.2 Using getline into a Variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 4.9.3 Using getline from a File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 4.9.4 Using getline into a Variable from a File . . . . . . . . . . . . . . . . 80 4.9.5 Using getline from a Pipe. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 4.9.6 Using getline into a Variable from a Pipe . . . . . . . . . . . . . . . 82 4.9.7 Using getline from a Coprocess . . . . . . . . . . . . . . . . . . . . . . . . . 82 4.9.8 Using getline into a Variable from a Coprocess . . . . . . . . . . 83 4.9.9 Points to Remember About getline . . . . . . . . . . . . . . . . . . . . . 83

4.9.10 Summary of getline Variants. . . . . . . . . . . . . . . . . . . . . . . . . . . 84 4.10 Reading Input with a Timeout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 4.11 Directories on the Command Line . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 4.12 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 4.13 Exercises . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

5 Printing Output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

5.1 The print Statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 5.2 print Statement Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 5.3 Output Separators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 5.4 Controlling Numeric Output with print. . . . . . . . . . . . . . . . . . . . . . . 92 5.5 Using printf Statements for Fancier Printing . . . . . . . . . . . . . . . . . 92

5.5.1 Introduction to the printf Statement . . . . . . . . . . . . . . . . . . . . 92 5.5.2 Format-Control Letters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 5.5.3 Modifiers for printf Formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 5.5.4 Examples Using printf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

5.6 Redirecting Output of print and printf . . . . . . . . . . . . . . . . . . . . . . 98 5.7 Special Files for Standard Preopened Data Streams. . . . . . . . . . . 100 5.8 Special File names in gawk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

5.8.1 Accessing Other Open Files with gawk. . . . . . . . . . . . . . . . . . . 101 5.8.2 Special Files for Network Communications. . . . . . . . . . . . . . . 102 5.8.3 Special File name Caveats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

5.9 Closing Input and Output Redirections. . . . . . . . . . . . . . . . . . . . . . . 102 5.10 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 5.11 Exercises . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

6 Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107

6.1 Constants, Variables, and Conversions . . . . . . . . . . . . . . . . . . . . . . . . 107

6.1.1 Constant Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 6.1.1.1 Numeric and String Constants . . . . . . . . . . . . . . . . . . . . . . 107 6.1.1.2 Octal and Hexadecimal Numbers . . . . . . . . . . . . . . . . . . . 107 6.1.1.3 Regular Expression Constants . . . . . . . . . . . . . . . . . . . . . . 108

6.1.2 Using Regular Expression Constants. . . . . . . . . . . . . . . . . . . . . 109 6.1.3 Variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110

6.1.3.1 Using Variables in a Program. . . . . . . . . . . . . . . . . . . . . . . 110 6.1.3.2 Assigning Variables on the Command Line. . . . . . . . . . 111

6.1.4 Conversion of Strings and Numbers. . . . . . . . . . . . . . . . . . . . . . 111 6.1.4.1 How awk Converts Between Strings and Numbers … 111 6.1.4.2 Locales Can Influence Conversion. . . . . . . . . . . . . . . . . . . 112

6.2 Operators: Doing Something with Values . . . . . . . . . . . . . . . . . . . . . 113 6.2.1 Arithmetic Operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 6.2.2 String Concatenation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 6.2.3 Assignment Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 6.2.4 Increment and Decrement Operators . . . . . . . . . . . . . . . . . . . . 119

6.3 Truth Values and Conditions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 6.3.1 True and False in awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 6.3.2 Variable Typing and Comparison Expressions. . . . . . . . . . . . 120 vi GAWK: Effective AWK Programming

6.3.2.1 String Type versus Numeric Type . . . . . . . . . . . . . . . . . . 121 6.3.2.2 Comparison Operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 6.3.2.3 String Comparison with POSIX Rules . . . . . . . . . . . . . . 124

6.3.3 Boolean Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 6.3.4 Conditional Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125

6.4 Function Calls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 6.5 Operator Precedence (How Operators Nest) . . . . . . . . . . . . . . . . . . 128 6.6 Where You Are Makes a Difference. . . . . . . . . . . . . . . . . . . . . . . . . . . 129 6.7 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129

7 Patterns, Actions, and Variables. . . . . . . . . . . . 131

7.1 Pattern Elements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 7.1.1 Regular Expressions as Patterns . . . . . . . . . . . . . . . . . . . . . . . . . 131 7.1.2 Expressions as Patterns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 7.1.3 Specifying Record Ranges with Patterns . . . . . . . . . . . . . . . . . 133 7.1.4 The BEGIN and END Special Patterns. . . . . . . . . . . . . . . . . . . . . 134

7.1.4.1 Startup and Cleanup Actions . . . . . . . . . . . . . . . . . . . . . . . 134 7.1.4.2 Input/Output from BEGIN and END Rules . . . . . . . . . . . 135

7.1.5 The BEGINFILE and ENDFILE Special Patterns . . . . . . . . . . . 135 7.1.6 The Empty Pattern . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136

7.2 Using Shell Variables in Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 7.3 Actions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 7.4 Control Statements in Actions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138

7.4.1 The if-else Statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 7.4.2 The while Statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 7.4.3 The do-while Statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 7.4.4 The for Statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 7.4.5 The switch Statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 7.4.6 The break Statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 7.4.7 The continue Statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 7.4.8 The next Statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 7.4.9 The nextfile Statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 7.4.10 The exit Statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146

7.5 Predefined Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 7.5.1 Built-in Variables That Control awk . . . . . . . . . . . . . . . . . . . . . 147 7.5.2 Built-in Variables That Convey Information . . . . . . . . . . . . . 149 7.5.3 Using ARGC and ARGV. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155

7.6 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157

8 Arrays in awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159

8.1 The Basics of Arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 8.1.1 Introduction to Arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 8.1.2 Referring to an Array Element . . . . . . . . . . . . . . . . . . . . . . . . . . 161 8.1.3 Assigning Array Elements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 8.1.4 Basic Array Example. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 8.1.5 Scanning All Elements of an Array . . . . . . . . . . . . . . . . . . . . . . 163 8.1.6 Using Predefined Array Scanning Orders with gawk . . . . . . 164

8.2 Using Numbers to Subscript Arrays . . . . . . . . . . . . . . . . . . . . . . . . . . 167 8.3 Using Uninitialized Variables as Subscripts . . . . . . . . . . . . . . . . . . . 167 8.4 The delete Statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 8.5 Multidimensional Arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169

8.5.1 Scanning Multidimensional Arrays. . . . . . . . . . . . . . . . . . . . . . . 171 8.6 Arrays of Arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 8.7 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173

9 Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175

9.1 Built-in Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 9.1.1 Calling Built-in Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 9.1.2 Numeric Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 9.1.3 String-Manipulation Functions . . . . . . . . . . . . . . . . . . . . . . . . . . 177

9.1.3.1 More about ‘\’ and ‘&’ with sub(), gsub(), and

gensub() . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 9.1.4 Input/Output Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 9.1.5 Time Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 9.1.6 Bit-Manipulation Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 9.1.7 Getting Type Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 9.1.8 String-Translation Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198

9.2 User-Defined Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 9.2.1 Function Definition Syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 9.2.2 Function Definition Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 9.2.3 Calling User-Defined Functions . . . . . . . . . . . . . . . . . . . . . . . . . . 202

9.2.3.1 Writing a Function Call . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 9.2.3.2 Controlling Variable Scope . . . . . . . . . . . . . . . . . . . . . . . . . 203 9.2.3.3 Passing Function Arguments by Value Or by Reference

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 9.2.4 The return Statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 9.2.5 Functions and Their Effects on Variable Typing . . . . . . . . . 208

9.3 Indirect Function Calls. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 9.4 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214

Part II: Problem Solving with awk

viii GAWK: Effective AWK Programming

10 A Library of awk Functions . . . . . . . . . . . . . . . . . 217

10.1 Naming Library Function Global Variables . . . . . . . . . . . . . . . . . . 218 10.2 General Programming . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219

10.2.1 Converting Strings to Numbers. . . . . . . . . . . . . . . . . . . . . . . . . 219 10.2.2 Assertions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 10.2.3 Rounding Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 10.2.4 The Cliff Random Number Generator . . . . . . . . . . . . . . . . . . 223 10.2.5 Translating Between Characters and Numbers . . . . . . . . . . 223 10.2.6 Merging an Array into a String. . . . . . . . . . . . . . . . . . . . . . . . . 225 10.2.7 Managing the Time of Day . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 10.2.8 Reading a Whole File at Once. . . . . . . . . . . . . . . . . . . . . . . . . . 227 10.2.9 Quoting Strings to Pass to the Shell . . . . . . . . . . . . . . . . . . . . 228

10.3 Data file Management . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 10.3.1 Noting Data file Boundaries . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 10.3.2 Rereading the Current File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 10.3.3 Checking for Readable Data files . . . . . . . . . . . . . . . . . . . . . . . 231 10.3.4 Checking for Zero-Length Files . . . . . . . . . . . . . . . . . . . . . . . . . 232 10.3.5 Treating Assignments as File names . . . . . . . . . . . . . . . . . . . . 232

10.4 Processing Command-Line Options . . . . . . . . . . . . . . . . . . . . . . . . . 233 10.5 Reading the User Database . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 10.6 Reading the Group Database. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 10.7 Traversing Arrays of Arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 10.8 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 248 10.9 Exercises . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249

11 Practical awk Programs . . . . . . . . . . . . . . . . . . . . . 251

11.1 Running the Example Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 11.2 Reinventing Wheels for Fun and Profit . . . . . . . . . . . . . . . . . . . . . . 251

11.2.1 Cutting Out Fields and Columns . . . . . . . . . . . . . . . . . . . . . . . 251 11.2.2 Searching for Regular Expressions in Files . . . . . . . . . . . . . . 256 11.2.3 Printing Out User Information . . . . . . . . . . . . . . . . . . . . . . . . . 260 11.2.4 Splitting a Large File into Pieces . . . . . . . . . . . . . . . . . . . . . . . 262 11.2.5 Duplicating Output into Multiple Files . . . . . . . . . . . . . . . . . 263 11.2.6 Printing Nonduplicated Lines of Text. . . . . . . . . . . . . . . . . . . 265 11.2.7 Counting Things . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269

11.3 A Grab Bag of awk Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 11.3.1 Finding Duplicated Words in a Document . . . . . . . . . . . . . . 271 11.3.2 An Alarm Clock Program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 272 11.3.3 Transliterating Characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 11.3.4 Printing Mailing Labels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 11.3.5 Generating Word-Usage Counts . . . . . . . . . . . . . . . . . . . . . . . . 278 11.3.6 Removing Duplicates from Unsorted Text . . . . . . . . . . . . . . 280 11.3.7 Extracting Programs from Texinfo Source Files . . . . . . . . . 281 11.3.8 A Simple Stream Editor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 11.3.9 An Easy Way to Use Library Functions . . . . . . . . . . . . . . . . 285 11.3.10 Finding Anagrams from a Dictionary. . . . . . . . . . . . . . . . . . 292 11.3.11 And Now for Something Completely Different . . . . . . . . . 293

11.4 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294 11.5 Exercises . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294

Part III: Moving Beyond Standard awk with

  gawk

12 Advanced Features of gawk . . . . . . . . . . . . . . . . . 299

12.1 Allowing Nondecimal Input Data. . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 12.2 Controlling Array Traversal and Array Sorting . . . . . . . . . . . . . . 300

12.2.1 Controlling Array Traversal . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 12.2.2 Sorting Array Values and Indices with gawk . . . . . . . . . . . . 304

12.3 Two-Way Communications with Another Process. . . . . . . . . . . . 305 12.4 Using gawk for Network Programming. . . . . . . . . . . . . . . . . . . . . . . 307 12.5 Profiling Your awk Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 12.6 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312

13 Internationalization with gawk . . . . . . . . . . . . . 315

13.1 Internationalization and Localization . . . . . . . . . . . . . . . . . . . . . . . . 315 13.2 GNU gettext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 13.3 Internationalizing awk Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 13.4 Translating awk Programs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319

13.4.1 Extracting Marked Strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 13.4.2 Rearranging printf Arguments . . . . . . . . . . . . . . . . . . . . . . . . 319 13.4.3 awk Portability Issues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320

13.5 A Simple Internationalization Example . . . . . . . . . . . . . . . . . . . . . . 321 13.6 gawk Can Speak Your Language. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 13.7 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323

14 Debugging awk Programs . . . . . . . . . . . . . . . . . . . 325

14.1 Introduction to the gawk Debugger . . . . . . . . . . . . . . . . . . . . . . . . . . 325 14.1.1 Debugging in General . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 14.1.2 Debugging Concepts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 14.1.3 awk Debugging. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326

14.2 Sample gawk Debugging Session. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 14.2.1 How to Start the Debugger. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 327 14.2.2 Finding the Bug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 327

14.3 Main Debugger Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 330 14.3.1 Control of Breakpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 330 14.3.2 Control of Execution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 332 14.3.3 Viewing and Changing Data . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 14.3.4 Working with the Stack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 334 14.3.5 Obtaining Information About the Program and the

Debugger State . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 335 14.3.6 Miscellaneous Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 337

14.4 Readline Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 338 14.5 Limitations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 14.6 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 x GAWK: Effective AWK Programming

15 Arithmetic and Arbitrary-Precision

Arithmetic with gawk . . . . . . . . . . . . . . . . . . . . . . . 341

15.1 A General Description of Computer Arithmetic. . . . . . . . . . . . . . 341 15.2 Other Stuff to Know . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 342 15.3 Arbitrary-Precision Arithmetic Features in gawk. . . . . . . . . . . . . 344 15.4 Floating-Point Arithmetic: Caveat Emptor! . . . . . . . . . . . . . . . . . 344

15.4.1 Floating-Point Arithmetic Is Not Exact . . . . . . . . . . . . . . . . 345 15.4.1.1 Many Numbers Cannot Be Represented Exactly … 345 15.4.1.2 Be Careful Comparing Values . . . . . . . . . . . . . . . . . . . . . 345 15.4.1.3 Errors Accumulate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 346

15.4.2 Getting the Accuracy You Need . . . . . . . . . . . . . . . . . . . . . . . . 347 15.4.3 Try a Few Extra Bits of Precision and Rounding . . . . . . . 347 15.4.4 Setting the Precision . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 348 15.4.5 Setting the Rounding Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . 349

15.5 Arbitrary-Precision Integer Arithmetic with gawk. . . . . . . . . . . . 350 15.6 Standards Versus Existing Practice. . . . . . . . . . . . . . . . . . . . . . . . . . 351 15.7 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 353

16 Writing Extensions for gawk . . . . . . . . . . . . . . . . 355

16.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 355 16.2 Extension Licensing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 355 16.3 How It Works at a High Level. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 355 16.4 API Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 357

16.4.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 357 16.4.2 General-Purpose Data Types . . . . . . . . . . . . . . . . . . . . . . . . . . . 359 16.4.3 Memory Allocation Functions and Convenience Macros

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 16.4.4 Constructor Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 16.4.5 Registration Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363

16.4.5.1 Registering An Extension Function . . . . . . . . . . . . . . . . 363 16.4.5.2 Registering An Exit Callback Function . . . . . . . . . . . . 364 16.4.5.3 Registering An Extension Version String. . . . . . . . . . . 365 16.4.5.4 Customized Input Parsers . . . . . . . . . . . . . . . . . . . . . . . . . 365 16.4.5.5 Customized Output Wrappers . . . . . . . . . . . . . . . . . . . . . 368 16.4.5.6 Customized Two-way Processors . . . . . . . . . . . . . . . . . . 370

16.4.6 Printing Messages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 16.4.7 Updating ERRNO. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 16.4.8 Requesting Values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 372 16.4.9 Accessing and Updating Parameters . . . . . . . . . . . . . . . . . . . . 372 16.4.10 Symbol Table Access. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 373

16.4.10.1 Variable Access and Update by Name . . . . . . . . . . . . 373 16.4.10.2 Variable Access and Update by Cookie . . . . . . . . . . . 373 16.4.10.3 Creating and Using Cached Values . . . . . . . . . . . . . . . 375

16.4.11 Array Manipulation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 377 16.4.11.1 Array Data Types. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 377 16.4.11.2 Array Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 16.4.11.3 Working With All The Elements of an Array . . . . . 379

16.4.11.4 How To Create and Populate Arrays . . . . . . . . . . . . . 383 16.4.12 API Variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 385

16.4.12.1 API Version Constants and Variables . . . . . . . . . . . . . 385 16.4.12.2 Informational Variables . . . . . . . . . . . . . . . . . . . . . . . . . . 386

16.4.13 Boilerplate Code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 16.5 How gawk Finds Extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 16.6 Example: Some File Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388

16.6.1 Using chdir() and stat(). . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 16.6.2 C Code for chdir() and stat() . . . . . . . . . . . . . . . . . . . . . . . 390 16.6.3 Integrating the Extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 397

16.7 The Sample Extensions in the gawk Distribution . . . . . . . . . . . . 398 16.7.1 File-Related Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 398 16.7.2 Interface to fnmatch() . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 16.7.3 Interface to fork(), wait(), and waitpid() . . . . . . . . . . . 402 16.7.4 Enabling In-Place File Editing . . . . . . . . . . . . . . . . . . . . . . . . . 402 16.7.5 Character and Numeric values: ord() and chr() . . . . . . . 403 16.7.6 Reading Directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 16.7.7 Reversing Output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 16.7.8 Two-Way I/O Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 405 16.7.9 Dumping and Restoring an Array . . . . . . . . . . . . . . . . . . . . . . 405 16.7.10 Reading an Entire File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 16.7.11 Extension Time Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 16.7.12 API Tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407

16.8 The gawkextlib Project . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 16.9 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 408 16.10 Exercises . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 409

Part IV: Appendices

Appendix A The Evolution of the awk Language

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 413

A.1 Major Changes Between V7 and SVR3.1 . . . . . . . . . . . . . . . . . . . . . 413 A.2 Changes Between SVR3.1 and SVR4. . . . . . . . . . . . . . . . . . . . . . . . . 414 A.3 Changes Between SVR4 and POSIX awk. . . . . . . . . . . . . . . . . . . . . 414 A.4 Extensions in Brian Kernighan’s awk. . . . . . . . . . . . . . . . . . . . . . . . . 415 A.5 Extensions in gawk Not in POSIX awk . . . . . . . . . . . . . . . . . . . . . . . 415 A.6 History of gawk Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 417 A.7 Common Extensions Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 423 A.8 Regexp Ranges and Locales: A Long Sad Story . . . . . . . . . . . . . . 424 A.9 Major Contributors to gawk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 425 A.10 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 427 xii GAWK: Effective AWK Programming

Appendix B Installing gawk . . . . . . . . . . . . . . . . . . . 429

B.1 The gawk Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 B.1.1 Getting the gawk Distribution. . . . . . . . . . . . . . . . . . . . . . . . . . . 429 B.1.2 Extracting the Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 B.1.3 Contents of the gawk Distribution. . . . . . . . . . . . . . . . . . . . . . . 430

B.2 Compiling and Installing gawk on Unix-Like Systems . . . . . . . . . 433 B.2.1 Compiling gawk for Unix-Like Systems . . . . . . . . . . . . . . . . . . 433 B.2.2 Additional Configuration Options . . . . . . . . . . . . . . . . . . . . . . . 433 B.2.3 The Configuration Process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 434

B.3 Installation on Other Operating Systems . . . . . . . . . . . . . . . . . . . . . 435

B.3.1 Installation on PC Operating Systems . . . . . . . . . . . . . . . . . . 435 B.3.1.1 Installing a Prepared Distribution for PC Systems . . 435 B.3.1.2 Compiling gawk for PC Operating Systems . . . . . . . . . 435 B.3.1.3 Testing gawk on PC Operating Systems . . . . . . . . . . . . 437 B.3.1.4 Using gawk on PC Operating Systems . . . . . . . . . . . . . . 437 B.3.1.5 Using gawk In The Cygwin Environment . . . . . . . . . . . 438 B.3.1.6 Using gawk In The MSYS Environment . . . . . . . . . . . . 439

B.3.2 Compiling and Installing gawk on Vax/VMS and OpenVMS

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 439 B.3.2.1 Compiling gawk on VMS . . . . . . . . . . . . . . . . . . . . . . . . . . . 439 B.3.2.2 Compiling gawk Dynamic Extensions on VMS . . . . . . 439 B.3.2.3 Installing gawk on VMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 B.3.2.4 Running gawk on VMS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 441 B.3.2.5 The VMS GNV Project . . . . . . . . . . . . . . . . . . . . . . . . . . . . 442 B.3.2.6 Some VMS Systems Have An Old Version of gawk . . 442

B.4 Reporting Problems and Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 442 B.5 Other Freely Available awk Implementations . . . . . . . . . . . . . . . . . 443 B.6 Summary. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 446

Appendix C Implementation Notes . . . . . . . . . . 447

C.1 Downward Compatibility and Debugging. . . . . . . . . . . . . . . . . . . . . 447 C.2 Making Additions to gawk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 447

C.2.1 Accessing The gawk Git Repository . . . . . . . . . . . . . . . . . . . . . 447 C.2.2 Adding New Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 448 C.2.3 Porting gawk to a New Operating System . . . . . . . . . . . . . . . 449 C.2.4 Why Generated Files Are Kept In Git . . . . . . . . . . . . . . . . . . 451

C.3 Probable Future Extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 453 C.4 Some Limitations of the Implementation . . . . . . . . . . . . . . . . . . . . . 453 C.5 Extension API Design . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 453

C.5.1 Problems With The Old Mechanism . . . . . . . . . . . . . . . . . . . . 454 C.5.2 Goals For A New Mechanism . . . . . . . . . . . . . . . . . . . . . . . . . . . 454 C.5.3 Other Design Decisions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 C.5.4 Room For Future Growth. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 456

C.6 Compatibility For Old Extensions. . . . . . . . . . . . . . . . . . . . . . . . . . . . 456 C.7 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 457

Appendix D Basic Programming Concepts . . 459

D.1 What a Program Does. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 459 D.2 Data Values in a Computer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460

Glossary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 463

GNU General Public License . . . . . . . . . . . . . . . . . . . 475

GNU Free Documentation License . . . . . . . . . . . . . 487

ADDENDUM: How to use this License for your documents . . . . . . . . 493

Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495

Foreword to the Third Edition 1

Foreword to the Third Edition

Arnold Robbins and I are good friends. We were introduced in 1990 by circumstances—and our favorite programming language, AWK. The circumstances started a couple of years earlier. I was working at a new job and noticed an unplugged Unix computer sitting in the corner. No one knew how to use it, and neither did I. However, a couple of days later, it was running, and I was root and the one-and-only user. That day, I began the transition from statistician to Unix programmer.

On one of many trips to the library or bookstore in search of books on Unix, I found the gray AWK book, a.k.a. Alfred V. Aho, Brian W. Kernighan, and Peter J. Weinberger’s The AWK Programming Language (Addison-Wesley, 1988). awk’s simple programming paradigm—find a pattern in the input and then perform an action—often reduced complex or tedious data manipulations to a few lines of code. I was excited to try my hand at programming in AWK.

Alas, the awk on my computer was a limited version of the language described in the gray book. I discovered that my computer had “old awk” and the book described “new awk.” I learned that this was typical; the old version refused to step aside or relinquish its name. If a system had a new awk, it was invariably called nawk, and few systems had it. The best way to get a new awk was to ftp the source code for gawk from prep.ai.mit.edu. gawk was a version of new awk written by David Trueman and Arnold, and available under the GNU General Public License.

(Incidentally, it’s no longer difficult to find a new awk. gawk ships with GNU/Linux, and you can download binaries or source code for almost any system; my wife uses gawk on her VMS box.)

My Unix system started out unplugged from the wall; it certainly was not plugged into a network. So, oblivious to the existence of gawk and the Unix community in general, and desiring a new awk, I wrote my own, called mawk. Before I was finished, I knew about gawk, but it was too late to stop, so I eventually posted to a comp.sources newsgroup.

A few days after my posting, I got a friendly email from Arnold introducing himself. He suggested we share design and algorithms and attached a draft of the POSIX standard so that I could update mawk to support language extensions added after publication of The AWK Programming Language.

Frankly, if our roles had been reversed, I would not have been so open and we probably would have never met. I’m glad we did meet. He is an AWK expert’s AWK expert and a genuinely nice person. Arnold contributes significant amounts of his expertise and time to the Free Software Foundation.

This book is the gawk reference manual, but at its core it is a book about AWK programming that will appeal to a wide audience. It is a definitive reference to the AWK language as defined by the 1987 Bell Laboratories release and codified in the 1992 POSIX Utilities standard.

On the other hand, the novice AWK programmer can study a wealth of practical programs that emphasize the power of AWK’s basic idioms: data-driven control flow, pattern matching with regular expressions, and associative arrays. Those looking for something new can try out gawk’s interface to network protocols via special /inet files.

The programs in this book make clear that an AWK program is typically much smaller and faster to develop than a counterpart written in C. Consequently, there is often a payoff to prototyping an algorithm or design in AWK to get it running quickly and expose problems early. Often, the interpreted performance is adequate and the AWK prototype becomes the product.

The new pgawk (profiling gawk), produces program execution counts. I recently experimented with an algorithm that for n lines of input, exhibited ∼ Cn^{2} performance, while theory predicted ∼ Cn log n behavior. A few minutes poring over the awkprof.out profile pinpointed the problem to a single line of code. pgawk is a welcome addition to my programmer’s toolbox.

Arnold has distilled over a decade of experience writing and using AWK programs, and developing gawk, into this book. If you use AWK or want to learn how, then read this book.

Michael Brennan Author of mawk March 2001

Foreword to the Fourth Edition 3

Foreword to the Fourth Edition

Some things don’t change. Thirteen years ago I wrote: “If you use AWK or want to learn how, then read this book.” True then, and still true today.

Learning to use a programming language is about more than mastering the syntax. One needs to acquire an understanding of how to use the features of the language to solve practical programming problems. A focus of this book is many examples that show how to use AWK.

Some things do change. Our computers are much faster and have more memory. Consequently, speed and storage inefficiencies of a high-level language matter less. Prototyping in AWK and then rewriting in C for performance reasons happens less, because more often the prototype is fast enough.

Of course, there are computing operations that are best done in C or C++. With gawk 4.1 and later, you do not have to choose between writing your program in AWK or in C/C++. You can write most of your program in AWK and the aspects that require C/C++ capabilities can be written in C/C++, and then the pieces glued together when the gawk module loads the C/C++ module as a dynamic plug-in. Chapter 16 [Writing Extensions for gawk], page 355, has all the details, and, as expected, many examples to help you learn the ins and outs.

I enjoy programming in AWK and had fun (re)reading this book. I think you will too. Michael Brennan Author of mawk October 2014

Preface 5

Preface

Several kinds of tasks occur repeatedly when working with text files. You might want to extract certain lines and discard the rest. Or you may need to make changes wherever certain patterns appear, but leave the rest of the file alone. Such jobs are often easy with awk. The awk utility interprets a special-purpose programming language that makes it easy to handle simple data-reformatting jobs.

The GNU implementation of awk is called gawk; if you invoke it with the proper options or environment variables, it is fully compatible with the POSIX^{1} specification of the awk language and with the Unix version of awk maintained by Brian Kernighan. This means that all properly written awk programs should work with gawk. So most of the time, we don’t distinguish between gawk and other awk implementations.

Using awk you can:

• Manage small, personal databases

• Generate reports

• Validate data

• Produce indexes and perform other document-preparation tasks

• Experiment with algorithms that you can adapt later to other computer languages

In addition, gawk provides facilities that make it easy to:

• Extract bits and pieces of data for processing

• Sort data

• Perform simple network communications

• Profile and debug awk programs

• Extend the language with functions written in C or C++

This book teaches you about the awk language and how you can use it effectively. You should already be familiar with basic system commands, such as cat and ls,^{2} as well as basic shell facilities, such as input/output (I/O) redirection and pipes.

Implementations of the awk language are available for many different computing environments. This book, while describing the awk language in general, also describes the particular implementation of awk called gawk (which stands for “GNU awk”). gawk runs on a broad range of Unix systems, ranging from Intel-architecture PC-based computers up through large-scale systems. gawk has also been ported to Mac OS X, Microsoft Windows (all versions) and OS/2 PCs, and OpenVMS.^{3}

1 The 2008 POSIX standard is accessible online at http://www.opengroup.org/onlinepubs/9699919799/.

2 These utilities are available on POSIX-compliant systems, as well as on traditional Unix-based systems.

If you are using some other operating system, you still need to be familiar with the ideas of I/O redirection

and pipes.

3 Some other, obsolete systems to which gawk was once ported are no longer supported and the code for

those systems has been removed.

History of awk and gawk

Recipe for a Programming Language

         1 part egrep 1 part snobol

2 parts ed  3 parts C

Blend all parts well using lex and yacc. Document minimally and release.

After eight years, add another part egrep and two more parts C. Document very well and release.

The name awk comes from the initials of its designers: Alfred V. Aho, Peter J. Weinberger, and Brian W. Kernighan. The original version of awk was written in 1977 at AT&T Bell Laboratories. In 1985, a new version made the programming language more powerful, introducing user-defined functions, multiple input streams, and computed regular expressions. This new version became widely available with Unix System V Release 3.1 (1987). The version in System V Release 4 (1989) added some new features and cleaned up the behavior in some of the “dark corners” of the language. The specification for awk in the POSIX Command Language and Utilities standard further clarified the language. Both the gawk designers and the original awk designers at Bell Laboratories provided feedback for the POSIX specification.

Paul Rubin wrote gawk in 1986. Jay Fenlason completed it, with advice from Richard Stallman. John Woods contributed parts of the code as well. In 1988 and 1989, David Trueman, with help from me, thoroughly reworked gawk for compatibility with the newer awk. Circa 1994, I became the primary maintainer. Current development focuses on bug fixes, performance improvements, standards compliance, and, occasionally, new features.

In May 1997, Jürgen Kahrs felt the need for network access from awk, and with a little help from me, set about adding features to do this for gawk. At that time, he also wrote the bulk of TCP/IP Internetworking with gawk (a separate document, available as part of the gawk distribution). His code finally became part of the main gawk distribution with gawk version 3.1.

John Haque rewrote the gawk internals, in the process providing an awk-level debugger. This version became available as gawk version 4.0 in 2011.

See Section A.9 [Major Contributors to gawk], page 425, for a full list of those who have made important contributions to gawk.

A Rose by Any Other Name

The awk language has evolved over the years. Full details are provided in Appendix A [The Evolution of the awk Language], page 413. The language described in this book is often referred to as “new awk.” By analogy, the original version of awk is referred to as “old awk.”

On most current systems, when you run the awk utility you get some version of new awk.^{4} If your system’s standard awk is the old one, you will see something like this if you try the test program:

4 Only Solaris systems still use an old awk for the default awk utility. A more modern awk lives in

/usr/xpg6/bin on these systems.

     $ awk 1 /dev/null
      error awk: syntax error near line 1
      error awk: bailing out near line 1

In this case, you should find a version of new awk, or just install gawk!

Throughout this book, whenever we refer to a language feature that should be available in any complete implementation of POSIX awk, we simply use the term awk. When referring to a feature that is specific to the GNU implementation, we use the term gawk.

Using This Book

The term awk refers to a particular program as well as to the language you use to tell this program what to do. When we need to be careful, we call the language “the awk language,” and the program “the awk utility.” This book explains both how to write programs in the awk language and how to run the awk utility. The term “awk program” refers to a program written by you in the awk programming language.

Primarily, this book explains the features of awk as defined in the POSIX standard. It does so in the context of the gawk implementation. While doing so, it also attempts to describe important differences between gawk and other awk implementations.^{5} Finally, it notes any gawk features that are not in the POSIX standard for awk.

This book has the difficult task of being both a tutorial and a reference. If you are a novice, feel free to skip over details that seem too complex. You should also ignore the many cross-references; they are for the expert user and for the Info and HTML (http://

www.gnu.org/software/gawk/manual/) versions of the book.

There are sidebars scattered throughout the book. They add a more complete explanation of points that are relevant, but not likely to be of interest on first reading. All appear in the index, under the heading “sidebar.”

Most of the time, the examples use complete awk programs. Some of the more advanced sections show only the part of the awk program that illustrates the concept being described.

Although this book is aimed principally at people who have not been exposed to awk, there is a lot of information here that even the awk expert should find useful. In particular, the description of POSIX awk and the example programs in Chapter 10 [A Library of awk Functions], page 217, and in Chapter 11 [Practical awk Programs], page 251, should be of interest.

This book is split into several parts, as follows:

• Part I describes the awk language and the gawk program in detail. It starts with the basics, and continues through all of the features of awk. It contains the following chapters: − Chapter 1 [Getting Started with awk], page 17, provides the essentials you need to

know to begin using awk. − Chapter 2 [Running awk and gawk], page 31, describes how to run gawk, the meaning of its command-line options, and how it finds awk program source files. − Chapter 3 [Regular Expressions], page 45, introduces regular expressions in general,

and in particular the flavors supported by POSIX awk and gawk.

5 All such differences appear in the index under the entry “differences in awk and gawk.”

− Chapter 4 [Reading Input Files], page 59, describes how awk reads your data. It introduces the concepts of records and fields, as well as the getline command. I/O redirection is first described here. Network I/O is also briefly introduced here.

− Chapter 5 [Printing Output], page 89, describes how awk programs can produce

output with print and printf. − Chapter 6 [Expressions], page 107, describes expressions, which are the basic build-

ing blocks for getting most things done in a program. − Chapter 7 [Patterns, Actions, and Variables], page 131, describes how to write pat-

terns for matching records, actions for doing something when a record is matched, and the predefined variables awk and gawk use.

− Chapter 8 [Arrays in awk], page 159, covers awk’s one-and-only data structure: the associative array. Deleting array elements and whole arrays is described, as well as sorting arrays in gawk. The chapter also describes how gawk provides arrays of arrays.

− Chapter 9 [Functions], page 175, describes the built-in functions awk and gawk provide, as well as how to define your own functions. It also discusses how gawk lets you call functions indirectly.

• Part II shows how to use awk and gawk for problem solving. There is lots of code here for you to read and learn from. This part contains the following chapters: − Chapter 10 [A Library of awk Functions], page 217, provides a number of functions

meant to be used from main awk programs. − Chapter 11 [Practical awk Programs], page 251, provides many sample awk pro-

grams.

Reading these two chapters allows you to see awk solving real problems.

• Part III focuses on features specific to gawk. It contains the following chapters:

− Chapter 12 [Advanced Features of gawk], page 299, describes a number of advanced features. Of particular note are the abilities to control the order of array traversal, have two-way communications with another process, perform TCP/IP networking, and profile your awk programs.

− Chapter 13 [Internationalization with gawk], page 315, describes special features

for translating program messages into different languages at runtime. − Chapter 14 [Debugging awk Programs], page 325, describes the gawk debugger. − Chapter 15 [Arithmetic and Arbitrary-Precision Arithmetic with gawk], page 341,

describes advanced arithmetic facilities. − Chapter 16 [Writing Extensions for gawk], page 355, describes how to add new

variables and functions to gawk by writing extensions in C or C++.

• Part IV provides the appendices, the Glossary, and two licenses that cover the gawk source code and this book, respectively. It contains the following appendices: − Appendix A [The Evolution of the awk Language], page 413, describes how the

awk language has evolved since its first release to the present. It also describes how gawk has acquired features over time.

− Appendix B [Installing gawk], page 429, describes how to get gawk, how to compile it on POSIX-compatible systems, and how to compile and use it on different non-POSIX systems. It also describes how to report bugs in gawk and where to get other freely available awk implementations.

− Appendix C [Implementation Notes], page 447, describes how to disable gawk’s extensions, as well as how to contribute new code to gawk, and some possible future directions for gawk development.

− Appendix D [Basic Programming Concepts], page 459, provides some very cursory background material for those who are completely unfamiliar with computer programming. The [Glossary], page 463, defines most, if not all, of the significant terms used throughout the book. If you find terms that you aren’t familiar with, try looking them up here.

− [GNU General Public License], page 475, and [GNU Free Documentation License], page 487, present the licenses that cover the gawk source code and this book, respectively.

Typographical Conventions

This book is written in Texinfo (http://www.gnu.org/software/texinfo/), the GNU

documentation formatting language. A single Texinfo source file is used to produce both the printed and online versions of the documentation. Because of this, the typographical conventions are slightly different than in other books you may have read.

Examples you would type at the command line are preceded by the common shell primary and secondary prompts, ‘$’ and ‘>’. Input that you type is shown like this. Output from the command is preceded by the glyph “ ⊣ ”. This typically represents the command’s standard output. Error messages and other output on the command’s standard error are preceded by the glyph “ error ”. For example:

     $ echo hi on stdout
     ⊣ hi on stdout
     $ echo hello on stderr 1>&2
      error hello on stderr

In the text, almost anything related to programming, such as command names, variable and function names, and string, numeric and regexp constants appear in this font. Code fragments appear in the same font and quoted, ‘like this’. Things that are replaced by the user or programmer appear in this font. Options look like this: -f. File names are indicated like this: /path/to/ourfile. Some things are emphasized like this, and if a point needs to be made strongly, it is done like this. The first occurrence of a new term is usually its definition and appears in the same font as the previous occurrence of “definition” in this sentence.

Characters that you type at the keyboard look like this. In particular, there are special characters called “control characters.” These are characters that you type by holding down both the CONTROL key and another key, at the same time. For example, a Ctrl-d is typed by first pressing and holding the CONTROL key, next pressing the d key, and finally releasing both keys.

For the sake of brevity, throughout this book, we refer to Brian Kernighan’s version of awk as “BWK awk.” (See Section B.5 [Other Freely Available awk Implementations], page 443, for information on his and other versions.)

Dark Corners

Dark corners are basically fractal—no matter how much you illuminate, there’s always a smaller but darker one. —Brian Kernighan

Until the POSIX standard (and GAWK: Effective AWK Programming), many features of awk were either poorly documented or not documented at all. Descriptions of such features (often called “dark corners”) are noted in this book with the picture of a flashlight in the margin, as shown here. They also appear in the index under the heading “dark corner.”

But, as noted by the opening quote, any coverage of dark corners is by definition incomplete.

Extensions to the standard awk language that are supported by more than one awk implementation are marked “(c.e.),” and listed in the index under “common extensions” and “extensions, common.”

The GNU Project and This Book

The Free Software Foundation (FSF) is a nonprofit organization dedicated to the production and distribution of freely distributable software. It was founded by Richard M. Stallman, the author of the original Emacs editor. GNU Emacs is the most widely used version of Emacs today.

The GNU^{6} Project is an ongoing effort on the part of the Free Software Foundation to create a complete, freely distributable, POSIX-compliant computing environment. The FSF uses the GNU General Public License (GPL) to ensure that its software’s source code is always available to the end user. A copy of the GPL is included in this book for your reference (see [GNU General Public License], page 475). The GPL applies to the C language source code for gawk. To find out more about the FSF and the GNU Project online, see the GNU Project’s home page (http://www.gnu.org). This book may also be read from

GNU’s website (http://www.gnu.org/software/gawk/manual/).

A shell, an editor (Emacs), highly portable optimizing C, C++, and Objective-C compilers, a symbolic debugger and dozens of large and small utilities (such as gawk), have all been completed and are freely available. The GNU operating system kernel (the HURD), has been released but remains in an early stage of development.

Until the GNU operating system is more fully developed, you should consider using GNU/Linux, a freely distributable, Unix-like operating system for Intel, Power Architecture, Sun SPARC, IBM S/390, and other systems.^{7} Many GNU/Linux distributions are available for download from the Internet.

The book you are reading is actually free—at least, the information in it is free to anyone. The machine-readable source code for the book comes with gawk. (Take a moment to check the Free Documentation License in [GNU Free Documentation License], page 487.)

The book itself has gone through multiple previous editions. Paul Rubin wrote the very first draft of The GAWK Manual; it was around 40 pages long. Diane Close and Richard Stallman improved it, yielding a version that was around 90 pages and barely described the original, “old” version of awk.

6 GNU stands for “GNU’s Not Unix.”

7 The terminology “GNU/Linux” is explained in the [Glossary], page 463.

I started working with that version in the fall of 1988. As work on it progressed, the FSF published several preliminary versions (numbered 0.x). In 1996, edition 1.0 was released with gawk 3.0.0. The FSF published the first two editions under the title The GNU Awk User’s Guide.

This edition maintains the basic structure of the previous editions. For FSF edition 4.0, the content was thoroughly reviewed and updated. All references to gawk versions prior to 4.0 were removed. Of significant note for that edition was the addition of Chapter 14 [Debugging awk Programs], page 325.

For FSF edition 4.1, the content has been reorganized into parts, and the major new additions are Chapter 15 [Arithmetic and Arbitrary-Precision Arithmetic with gawk], page 341, and Chapter 16 [Writing Extensions for gawk], page 355.

This book will undoubtedly continue to evolve. If you find an error in the book, please report it! See Section B.4 [Reporting Problems and Bugs], page 442, for information on submitting problem reports electronically.

How to Contribute

As the maintainer of GNU awk, I once thought that I would be able to manage a collection of publicly available awk programs and I even solicited contributions. Making things available on the Internet helps keep the gawk distribution down to manageable size.

The initial collection of material, such as it is, is still available at ftp://ftp. freefriends.org/arnold/Awkstuff. In the hopes of doing something more broad, I acquired the awk.info domain.

However, I found that I could not dedicate enough time to managing contributed code: the archive did not grow and the domain went unused for several years.

Late in 2008, a volunteer took on the task of setting up an awk-related website—http:// awk.info—and did a very nice job.

If you have written an interesting awk program, or have written a gawk extension that you would like to share with the rest of the world, please see http://awk.info/?contribute for how to contribute it to the website.

Acknowledgments

The initial draft of The GAWK Manual had the following acknowledgments:

Many people need to be thanked for their assistance in producing this manual. Jay Fenlason contributed many ideas and sample programs. Richard Mlynarik and Robert Chassell gave helpful comments on drafts of this manual. The paper A Supplemental Document for AWK by John W. Pierce of the Chemistry Department at UC San Diego, pinpointed several issues relevant both to awk implementation and to this manual, that would otherwise have escaped us.

I would like to acknowledge Richard M. Stallman, for his vision of a better world and for his courage in founding the FSF and starting the GNU Project.

Earlier editions of this book had the following acknowledgements:

The following people (in alphabetical order) provided helpful comments on various versions of this book: Rick Adams, Dr. Nelson H.F. Beebe, Karl Berry, Dr. Michael Brennan, Rich Burridge, Claire Cloutier, Diane Close, Scott Deifik, Christopher (“Topher”) Eliot, Jeffrey Friedl, Dr. Darrel Hankerson, Michal Jaegermann, Dr. Richard J. LeBlanc, Michael Lijewski, Pat Rankin, Miriam Robbins, Mary Sheehan, and Chuck Toporek.

Robert J. Chassell provided much valuable advice on the use of Texinfo. He also deserves special thanks for convincing me not to title this book How to Gawk Politely. Karl Berry helped significantly with the TEX part of Texinfo.

I would like to thank Marshall and Elaine Hartholz of Seattle and Dr. Bert and Rita Schreiber of Detroit for large amounts of quiet vacation time in their homes, which allowed me to make significant progress on this book and on gawk itself.

Phil Hughes of SSC contributed in a very important way by loaning me his laptop GNU/Linux system, not once, but twice, which allowed me to do a lot of work while away from home.

David Trueman deserves special credit; he has done a yeoman job of evolving gawk so that it performs well and without bugs. Although he is no longer involved with gawk, working with him on this project was a significant pleasure.

The intrepid members of the GNITS mailing list, and most notably Ulrich Drepper, provided invaluable help and feedback for the design of the internationalization features.

Chuck Toporek, Mary Sheehan, and Claire Cloutier of O’Reilly & Associates contributed significant editorial help for this book for the 3.1 release of gawk.

Dr. Nelson Beebe, Andreas Buening, Dr. Manuel Collado, Antonio Colombo, Stephen Davies, Scott Deifik, Akim Demaille, Darrel Hankerson, Michal Jaegermann, Jürgen Kahrs, Stepan Kasal, John Malmberg, Dave Pitts, Chet Ramey, Pat Rankin, Andrew Schorr, Corinna Vinschen, and Eli Zaretskii (in alphabetical order) make up the current gawk “crack portability team.” Without their hard work and help, gawk would not be nearly the robust, portable program it is today. It has been and continues to be a pleasure working with this team of fine people.

Notable code and documentation contributions were made by a number of people. See Section A.9 [Major Contributors to gawk], page 425, for the full list.

Thanks to Michael Brennan for the Forewords.

Thanks to Patrice Dumas for the new makeinfo program. Thanks to Karl Berry, who continues to work to keep the Texinfo markup language sane.

Robert P.J. Day, Michael Brennan, and Brian Kernighan kindly acted as reviewers for the 2015 edition of this book. Their feedback helped improve the final work.

I would also like to thank Brian Kernighan for his invaluable assistance during the testing and debugging of gawk, and for his ongoing help and advice in clarifying numerous points about the language. We could not have done nearly as good a job on either gawk or its documentation without his help.

Brian is in a class by himself as a programmer and technical author. I have to thank him (yet again) for his ongoing friendship and for being a role model to me for close to 30 years! Having him as a reviewer is an exciting privilege. It has also been extremely humbling…

I must thank my wonderful wife, Miriam, for her patience through the many versions of this project, for her proofreading, and for sharing me with the computer. I would like to thank my parents for their love, and for the grace with which they raised and educated me. Finally, I also must acknowledge my gratitude to G-d, for the many opportunities He has sent my way, as well as for the gifts He has given me with which to take advantage of those opportunities.

Arnold Robbins Nof Ayalon Israel February 2015

Part I: The awk Language

Chapter 1: Getting Started with awk 17

1 Getting Started with awk

The basic function of awk is to search files for lines (or other units of text) that contain certain patterns. When a line matches one of the patterns, awk performs specified actions on that line. awk continues to process input lines in this way until it reaches the end of the input files.

Programs in awk are different from programs in most other languages, because awk programs are data driven (i.e., you describe the data you want to work with and then what to do when you find it). Most other languages are procedural; you have to describe, in great detail, every step the program should take. When working with procedural languages, it is usually much harder to clearly describe the data your program will process. For this reason, awk programs are often refreshingly easy to read and write.

When you run awk, you specify an awk program that tells awk what to do. The program consists of a series of rules (it may also contain function definitions, an advanced feature that we will ignore for now; see Section 9.2 [User-Defined Functions], page 199). Each rule specifies one pattern to search for and one action to perform upon finding the pattern.

Syntactically, a rule consists of a pattern followed by an action. The action is enclosed in braces to separate it from the pattern. Newlines usually separate rules. Therefore, an awk program looks like this:

     pattern { action }
     pattern { action }
     ...

1.1 How to Run awk Programs

There are several ways to run an awk program. If the program is short, it is easiest to include it in the command that runs awk, like this:

     awk 'program' input-file1 input-file2 ...

When the program is long, it is usually more convenient to put it in a file and run it with a command like this:

     awk -f program-file input-file1 input-file2 ...

This section discusses both mechanisms, along with several variations of each.

1.1.1 One-Shot Throwaway awk Programs

Once you are familiar with awk, you will often type in simple programs the moment you want to use them. Then you can write the program as the first argument of the awk command, like this:

     awk 'program' input-file1 input-file2 ...

where program consists of a series of patterns and actions, as described earlier.

This command format instructs the shell, or command interpreter, to start awk and use the program to process records in the input file(s). There are single quotes around program so the shell won’t interpret any awk characters as special shell characters. The quotes also cause the shell to treat all of program as a single argument for awk, and allow program to be more than one line long.

This format is also useful for running short or medium-sized awk programs from shell scripts, because it avoids the need for a separate file for the awk program. A self-contained shell script is more reliable because there are no other files to misplace.

Later in this chapter, in Section 1.3 [Some Simple Examples], page 24, we’ll see examples of several short, self-contained programs.

1.1.2 Running awk Without Input Files

You can also run awk without any input files. If you type the following command line:

     awk 'program'

awk applies the program to the standard input, which usually means whatever you type on the keyboard. This continues until you indicate end-of-file by typing Ctrl-d. (On non-POSIX operating systems, the end-of-file character may be different. For example, on OS/2,

it is Ctrl-z.)

As an example, the following program prints a friendly piece of advice (from Douglas Adams’s The Hitchhiker’s Guide to the Galaxy), to keep you from worrying about the complexities of computer programming:

     $ awk 'BEGIN { print "Don\47t Panic!" }'
     ⊣ Don't Panic!

awk executes statements associated with BEGIN before reading any input. If there are no other statements in your program, as is the case here, awk just stops, instead of trying to read input it doesn’t know how to process. The ‘\47’ is a magic way (explained later) of getting a single quote into the program, without having to engage in ugly shell quoting tricks.

NOTE: If you use Bash as your shell, you should execute the command ‘set +H’ before running this program interactively, to disable the C shell-style command history, which treats ‘!’ as a special character. We recommend putting this command into your personal startup file.

This next simple awk program emulates the cat utility; it copies whatever you type on the keyboard to its standard output (why this works is explained shortly):

     $ awk '{ print }'
     Now is the time for all good men
     ⊣ Now is the time for all good men
     to come to the aid of their country.
     ⊣ to come to the aid of their country.
     Four score and seven years ago, ...
     ⊣ Four score and seven years ago, ...
     What, me worry?
     ⊣ What, me worry?
     Ctrl-d

1.1.3 Running Long Programs

Sometimes awk programs are very long. In these cases, it is more convenient to put the program into a separate file. In order to tell awk to use that file for its program, you type:

     awk -f source-file input-file1 input-file2 ...

The -f instructs the awk utility to get the awk program from the file source-file (see Section 2.2 [Command-Line Options], page 31). Any file name can be used for source-file. For example, you could put the program:

     BEGIN { print "Don't Panic!" }

into the file advice. Then this command:

     awk -f advice

does the same thing as this one:

     awk 'BEGIN { print "Don\47t Panic!" }'

This was explained earlier (see Section 1.1.2 [Running awk Without Input Files], page 18). Note that you don’t usually need single quotes around the file name that you specify with -f, because most file names don’t contain any of the shell’s special characters. Notice that in advice, the awk program did not have single quotes around it. The quotes are only needed for programs that are provided on the awk command line. (Also, placing the program in a file allows us to use a literal single quote in the program text, instead of the magic ‘\47’.)

If you want to clearly identify an awk program file as such, you can add the extension .awk to the file name. This doesn’t affect the execution of the awk program but it does make “housekeeping” easier.

1.1.4 Executable awk Programs

Once you have learned awk, you may want to write self-contained awk scripts, using the ‘#!’ script mechanism. You can do this on many systems.^{1} For example, you could update the file advice to look like this:

     #! /bin/awk -f
     BEGIN { print "Don't Panic!" }

After making this file executable (with the chmod utility), simply type ‘advice’ at the shell and the system arranges to run awk as if you had typed ‘awk -f advice’:

     $ chmod +x advice
     $ advice
     ⊣ Don't Panic!

(We assume you have the current directory in your shell’s search path variable [typically $PATH]. If not, you may need to type ‘./advice’ at the shell.)

Self-contained awk scripts are useful when you want to write a program that users can invoke without their having to know that the program is written in awk.

1 The ‘#!’ mechanism works on GNU/Linux systems, BSD-based systems, and commercial Unix systems.

Understanding ‘#!’

awk is an interpreted language. This means that the awk utility reads your program and then processes your data according to the instructions in your program. (This is different from a compiled language such as C, where your program is first compiled into machine code that is executed directly by your system’s processor.) The awk utility is thus termed an interpreter. Many modern languages are interpreted.

The line beginning with ‘#!’ lists the full file name of an interpreter to run and a single optional initial command-line argument to pass to that interpreter. The operating system then runs the interpreter with the given argument and the full argument list of the executed program. The first argument in the list is the full file name of the awk program. The rest of the argument list contains either options to awk, or data files, or both. (Note that on many systems awk may be found in /usr/bin instead of in /bin.)

Some systems limit the length of the interpreter name to 32 characters. Often, this can be dealt with by using a symbolic link.

You should not put more than one argument on the ‘#!’ line after the path to awk. It does not work. The operating system treats the rest of the line as a single argument and passes it to awk. Doing this leads to confusing behavior—most likely a usage diagnostic of some sort from awk.

Finally, the value of ARGV[0] (see Section 7.5 [Predefined Variables], page 147) varies depending upon your operating system. Some systems put ‘awk’ there, some put the full pathname of awk (such as /bin/awk), and some put the name of your script (‘advice’). Don’t rely on the value of ARGV[0] to provide your script name.

1.1.5 Comments in awk Programs

A comment is some text that is included in a program for the sake of human readers; it is not really an executable part of the program. Comments can explain what the program does and how it works. Nearly all programming languages have provisions for comments, as programs are typically hard to understand without them.

In the awk language, a comment starts with the number sign character (‘#’) and continues to the end of the line. The ‘#’ does not have to be the first character on the line. The awk language ignores the rest of a line following a number sign. For example, we could have put the following into advice:

    # This program prints a nice, friendly message.  It helps
    # keep novice users from being afraid of the computer.
    BEGIN    { print "Don't Panic!" }

You can put comment lines into keyboard-composed throwaway awk programs, but this usually isn’t very useful; the purpose of a comment is to help you or another person understand the program when reading it at a later time.

CAUTION: As mentioned in Section 1.1.1 [One-Shot Throwaway awk Programs], page 17, you can enclose short to medium-sized programs in single quotes, in order to keep your shell scripts self-contained. When doing so, don’t put an apostrophe (i.e., a single quote) into a comment (or anywhere else in your program). The shell interprets the quote as the closing quote for the entire program. As a result, usually the shell prints a message about mismatched quotes, and if awk actually runs, it will probably print strange messages about syntax errors. For example, look at the following:

          $ awk 'BEGIN { print "hello" } # let's be cute'
          >

The shell sees that the first two quotes match, and that a new quoted object begins at the end of the command line. It therefore prompts with the secondary prompt, waiting for more input. With Unix awk, closing the quoted string produces this result:

          $ awk '{ print "hello" } # let's be cute'
          > '
           error awk: can't open file be
           error source line number 1

Putting a backslash before the single quote in ‘let's’ wouldn’t help, because backslashes are not special inside single quotes. The next subsection describes the shell’s quoting rules.

1.1.6 Shell Quoting Issues

For short to medium-length awk programs, it is most convenient to enter the program on the awk command line. This is best done by enclosing the entire program in single quotes. This is true whether you are entering the program interactively at the shell prompt, or writing it as part of a larger shell script:

     awk 'program text' input-file1 input-file2 ...

Once you are working with the shell, it is helpful to have a basic knowledge of shell quoting rules. The following rules apply only to POSIX-compliant, Bourne-style shells (such as Bash, the GNU Bourne-Again Shell). If you use the C shell, you’re on your own.

Before diving into the rules, we introduce a concept that appears throughout this book, which is that of the null, or empty, string.

The null string is character data that has no value. In other words, it is empty. It is written in awk programs like this: "". In the shell, it can be written using single or double quotes: "" or ''. Although the null string has no characters in it, it does exist. For example, consider this command:

     $ echo ""

Here, the echo utility receives a single argument, even though that argument has no characters in it. In the rest of this book, we use the terms null string and empty string interchangeably. Now, on to the quoting rules:

• Quoted items can be concatenated with nonquoted items as well as with other quoted items. The shell turns everything into one argument for the command.

• Preceding any single character with a backslash (‘\’) quotes that character. The shell removes the backslash and passes the quoted character on to the command.

• Single quotes protect everything between the opening and closing quotes. The shell does no interpretation of the quoted text, passing it on verbatim to the command. It is impossible to embed a single quote inside single-quoted text. Refer back to Section 1.1.5 [Comments in awk Programs], page 20, for an example of what happens if you try.

• Double quotes protect most things between the opening and closing quotes. The shell does at least variable and command substitution on the quoted text. Different shells may do additional kinds of processing on double-quoted text. Because certain characters within double-quoted text are processed by the shell, they must be escaped within the text. Of note are the characters ‘$’, ‘`’, ‘\’, and ‘"’, all of which must be preceded by a backslash within double-quoted text if they are to be passed on literally to the program. (The leading backslash is stripped first.) Thus, the example seen previously in Section 1.1.2 [Running awk Without Input Files], page 18:

     awk 'BEGIN { print "Don\47t Panic!" }'

could instead be written this way:

     $ awk "BEGIN { print \"Don't Panic!\" }"
     ⊣ Don't Panic!

Note that the single quote is not special within double quotes.

• Null strings are removed when they occur as part of a non-null command-line argument, while explicit null objects are kept. For example, to specify that the field separator FS should be set to the null string, use:

     awk -F "" 'program' files # correct

Don’t use this:

     awk -F"" 'program' files  # wrong!

In the second case, awk attempts to use the text of the program as the value of FS, and the first file name as the text of the program! This results in syntax errors at best, and confusing behavior at worst.

Mixing single and double quotes is difficult. You have to resort to shell quoting tricks, like this:

 $ awk 'BEGIN { print "Here is a single quote <'"'"'>" }'
  ⊣ Here is a single quote <'>

This program consists of three concatenated quoted strings. The first and the third are single-quoted, and the second is double-quoted.

This can be “simplified” to:

 $ awk 'BEGIN { print "Here is a single quote <'\''>" }'
  ⊣ Here is a single quote <'>

Judge for yourself which of these two is the more readable.

Another option is to use double quotes, escaping the embedded, awk-level double quotes:

 $ awk "BEGIN { print \"Here is a single quote <'>\" }"
  ⊣ Here is a single quote <'>

This option is also painful, because double quotes, backslashes, and dollar signs are very common in more advanced awk programs.

A third option is to use the octal escape sequence equivalents (see Section 3.2 [Escape Sequences], page 46) for the single- and double-quote characters, like so:

 $ awk 'BEGIN { print "Here is a single quote <\47>" }'
  ⊣ Here is a single quote <'>
 $ awk 'BEGIN { print "Here is a double quote <\42>" }'

     ⊣ Here is a double quote <">

This works nicely, but you should comment clearly what the escapes mean.

A fourth option is to use command-line variable assignment, like this:

     $ awk -v sq="'" 'BEGIN { print "Here is a single quote <" sq ">" }'
     ⊣ Here is a single quote <'>

(Here, the two string constants and the value of sq are concatenated into a single string that is printed by print.)

If you really need both single and double quotes in your awk program, it is probably best to move it into a separate file, where the shell won’t be part of the picture and you can say what you mean.

1.1.6.1 Quoting in MS-Windows Batch Files

Although this book generally only worries about POSIX systems and the POSIX shell, the following issue arises often enough for many users that it is worth addressing.

The “shells” on Microsoft Windows systems use the double-quote character for quoting, and make it difficult or impossible to include an escaped double-quote character in a command-line script. The following example, courtesy of Jeroen Brink, shows how to print all lines in a file surrounded by double quotes:

     gawk "{ print \"\042\" $0 \"\042\" }" file

1.2 Data files for the Examples

Many of the examples in this book take their input from two sample data files. The first, mail-list, represents a list of peoples’ names together with their email addresses and information about those people. The second data file, called inventory-shipped, contains information about monthly shipments. In both files, each line is considered to be one record.

In mail-list, each record contains the name of a person, his/her phone number, his/her email address, and a code for his/her relationship with the author of the list. The columns are aligned using spaces. An ‘A’ in the last column means that the person is an acquaintance. An ‘F’ in the last column means that the person is a friend. An ‘R’ means that the person is a relative:

     Amelia       555-5553     amelia.zodiacusque@gmail.com    F
     Anthony      555-3412     anthony.asserturo@hotmail.com   A
     Becky        555-7685     becky.algebrarum@gmail.com      A
     Bill         555-1675     bill.drowning@hotmail.com       A
     Broderick    555-0542     broderick.aliquotiens@yahoo.com R
     Camilla      555-2912     camilla.infusarum@skynet.be     R
     Fabius       555-1234     fabius.undevicesimus@ucb.edu    F
     Julie        555-6699     julie.perscrutabor@skeeve.com   F
     Martin       555-6480     martin.codicibus@hotmail.com    A
     Samuel       555-3430     samuel.lanceolis@shu.edu        A
     Jean-Paul    555-2127     jeanpaul.campanorum@nyu.edu     R

The data file inventory-shipped represents information about shipments during the year. Each record contains the month, the number of green crates shipped, the number of red boxes shipped, the number of orange bags shipped, and the number of blue packages shipped, respectively. There are 16 entries, covering the 12 months of last year and the first four months of the current year. An empty line separates the data for the two years:

Jan  13  25  15 115
Feb  15  32  24 226
Mar  15  24  34 228
Apr  31  52  63 420
May  16  34  29 208
Jun  31  42  75 492
Jul  24  34  67 436
Aug  15  34  47 316
Sep  13  55  37 277
Oct  29  54  68 525
Nov  20  87  82 577
Dec  17  35  61 401
Jan  21  36  64 620
Feb  26  58  80 652
Mar  24  75  70 495
Apr  21  70  74 514

The sample files are included in the gawk distribution, in the directory awklib/eg/data.

1.3 Some Simple Examples

The following command runs a simple awk program that searches the input file mail-list for the character string ‘li’ (a grouping of characters is usually called a string; the term string is based on similar usage in English, such as “a string of pearls” or “a string of cars in a train”):

awk '/li/ { print $0 }' mail-list

When lines containing ‘li’ are found, they are printed because ‘print $0’ means print the current line. (Just ‘print’ by itself means the same thing, so we could have written that instead.)

You will notice that slashes (‘/’) surround the string ‘li’ in the awk program. The slashes indicate that ‘li’ is the pattern to search for. This type of pattern is called a regular expression, which is covered in more detail later (see Chapter 3 [Regular Expressions], page 45). The pattern is allowed to match parts of words. There are single quotes around the awk program so that the shell won’t interpret any of it as special shell characters.

Here is what this program prints:

$ awk '/li/ { print $0 }' mail-list
⊣ Amelia       555-5553     amelia.zodiacusque@gmail.com    F
⊣ Broderick    555-0542     broderick.aliquotiens@yahoo.com R
⊣ Julie        555-6699     julie.perscrutabor@skeeve.com   F
⊣ Samuel       555-3430     samuel.lanceolis@shu.edu        A

In an awk rule, either the pattern or the action can be omitted, but not both. If the pattern is omitted, then the action is performed for every input line. If the action is omitted, the default action is to print all lines that match the pattern.

Thus, we could leave out the action (the print statement and the braces) in the previous example and the result would be the same: awk prints all lines matching the pattern ‘li’. By comparison, omitting the print statement but retaining the braces makes an empty action that does nothing (i.e., no lines are printed).

Many practical awk programs are just a line or two long. Following is a collection of useful, short programs to get you started. Some of these programs contain constructs that haven’t been covered yet. (The description of the program will give you a good idea of what is going on, but you’ll need to read the rest of the book to become an awk expert!) Most of the examples use a data file named data. This is just a placeholder; if you use these programs yourself, substitute your own file names for data. For future reference, note that there is often more than one way to do things in awk. At some point, you may want to look back at these examples and see if you can come up with different ways to do the same things shown here:

• Print every line that is longer than 80 characters:

         awk 'length($0) > 80' data

The sole rule has a relational expression as its pattern and has no action—so it uses the default action, printing the record.

• Print the length of the longest input line:

         awk '{ if (length($0) > max) max = length($0) }
              END { print max }' data

The code associated with END executes after all input has been read; it’s the other side of the coin to BEGIN.

• Print the length of the longest line in data:

         expand data | awk '{ if (x < length($0)) x = length($0) }
                            END { print "maximum line length is " x }'

This example differs slightly from the previous one: the input is processed by the expand utility to change TABs into spaces, so the widths compared are actually the right-margin columns, as opposed to the number of input characters on each line.

• Print every line that has at least one field:

         awk 'NF > 0' data

This is an easy way to delete blank lines from a file (or rather, to create a new file similar to the old file but from which the blank lines have been removed).

• Print seven random numbers from 0 to 100, inclusive:

         awk 'BEGIN { for (i = 1; i <= 7; i++)
                          print int(101 * rand()) }'

• Print the total number of bytes used by files:

         ls -l files | awk '{ x += $5 }
                            END { print "total bytes: " x }'

• Print the total number of kilobytes used by files:

         ls -l files | awk '{ x += $5 }
            END { print "total K-bytes:", x / 1024 }'

• Print a sorted list of the login names of all users:

         awk -F: '{ print $1 }' /etc/passwd | sort

• Count the lines in a file:

         awk 'END { print NR }' data

• Print the even-numbered lines in the data file:

         awk 'NR % 2 == 0' data

If you used the expression ‘NR % 2 == 1’ instead, the program would print the oddnumbered lines.

1.4 An Example with Two Rules

The awk utility reads the input files one line at a time. For each line, awk tries the patterns of each rule. If several patterns match, then several actions execute in the order in which they appear in the awk program. If no patterns match, then no actions run.

After processing all the rules that match the line (and perhaps there are none), awk reads the next line. (However, see Section 7.4.8 [The next Statement], page 144, and also see Section 7.4.9 [The nextfile Statement], page 145.) This continues until the program reaches the end of the file. For example, the following awk program contains two rules:

     /12/  { print $0 }
     /21/  { print $0 }

The first rule has the string ‘12’ as the pattern and ‘print $0’ as the action. The second rule has the string ‘21’ as the pattern and also has ‘print $0’ as the action. Each rule’s action is enclosed in its own pair of braces.

This program prints every line that contains the string ‘12’ or the string ‘21’. If a line contains both strings, it is printed twice, once by each rule.

This is what happens if we run this program on our two sample data files, mail-list

and inventory-shipped:
     $ awk '/12/ { print $0 }
     >      /21/ { print $0 }' mail-list inventory-shipped
     ⊣ Anthony      555-3412     anthony.asserturo@hotmail.com   A
     ⊣ Camilla      555-2912     camilla.infusarum@skynet.be     R
     ⊣ Fabius       555-1234     fabius.undevicesimus@ucb.edu    F
     ⊣ Jean-Paul    555-2127     jeanpaul.campanorum@nyu.edu     R
     ⊣ Jean-Paul    555-2127     jeanpaul.campanorum@nyu.edu     R
     ⊣ Jan  21  36  64 620
     ⊣ Apr  21  70  74 514

Note how the line beginning with ‘Jean-Paul’ in mail-list was printed twice, once for each rule.

1.5 A More Complex Example

Now that we’ve mastered some simple tasks, let’s look at what typical awk programs do. This example shows how awk can be used to summarize, select, and rearrange the output of another utility. It uses features that haven’t been covered yet, so don’t worry if you don’t understand all the details:

     ls -l | awk '$6 == "Nov" { sum += $5 }
                  END { print sum }'

This command prints the total number of bytes in all the files in the current directory that were last modified in November (of any year). The ‘ls -l’ part of this example is a system command that gives you a listing of the files in a directory, including each file’s size and the date the file was last modified. Its output looks like this:

     -rw-r--r--  1 arnold   user   1933 Nov  7 13:05 Makefile
     -rw-r--r--  1 arnold   user  10809 Nov  7 13:03 awk.h
     -rw-r--r--  1 arnold   user    983 Apr 13 12:14 awk.tab.h
     -rw-r--r--  1 arnold   user  31869 Jun 15 12:20 awkgram.y
     -rw-r--r--  1 arnold   user  22414 Nov  7 13:03 awk1.c
     -rw-r--r--  1 arnold   user  37455 Nov  7 13:03 awk2.c
     -rw-r--r--  1 arnold   user  27511 Dec  9 13:07 awk3.c
     -rw-r--r--  1 arnold   user   7989 Nov  7 13:03 awk4.c

The first field contains read-write permissions, the second field contains the number of links to the file, and the third field identifies the file’s owner. The fourth field identifies the file’s group. The fifth field contains the file’s size in bytes. The sixth, seventh, and eighth fields contain the month, day, and time, respectively, that the file was last modified. Finally, the ninth field contains the file name.

The ‘$6 == "Nov"’ in our awk program is an expression that tests whether the sixth field of the output from ‘ls -l’ matches the string ‘Nov’. Each time a line has the string ‘Nov’ for its sixth field, awk performs the action ‘sum += $5’. This adds the fifth field (the file’s size) to the variable sum. As a result, when awk has finished reading all the input lines, sum is the total of the sizes of the files whose lines matched the pattern. (This works because awk variables are automatically initialized to zero.)

After the last line of output from ls has been processed, the END rule executes and prints the value of sum. In this example, the value of sum is 80600.

These more advanced awk techniques are covered in later sections (see Section 7.3 [Actions], page 137). Before you can move on to more advanced awk programming, you have to know how awk interprets your input and displays your output. By manipulating fields and using print statements, you can produce some very useful and impressive-looking reports.

1.6 awk Statements Versus Lines

Most often, each line in an awk program is a separate statement or separate rule, like this:

     awk '/12/  { print $0 }
          /21/  { print $0 }' mail-list inventory-shipped

However, gawk ignores newlines after any of the following symbols and keywords:

     ,    {    ?    :    ||    &&    do    else

A newline at any other point is considered the end of the statement.^{2}

If you would like to split a single statement into two lines at a point where a newline would terminate it, you can continue it by ending the first line with a backslash character (‘\’). The backslash must be the final character on the line in order to be recognized as a continuation character. A backslash is allowed anywhere in the statement, even in the middle of a string or regular expression. For example:

2 The ‘?’ and ‘:’ referred to here is the three-operand conditional expression described in Section 6.3.4

[Conditional Expressions], page 125. Splitting lines after ‘?’ and ‘:’ is a minor gawk extension; if --posix

is specified (see Section 2.2 [Command-Line Options], page 31), then this extension is disabled.

     awk '/This regular expression is too long, so continue it\
      on the next line/ { print $1 }'

We have generally not used backslash continuation in our sample programs. gawk places no limit on the length of a line, so backslash continuation is never strictly necessary; it just makes programs more readable. For this same reason, as well as for clarity, we have kept most statements short in the programs presented throughout the book. Backslash continuation is most useful when your awk program is in a separate source file instead of entered from the command line. You should also note that many awk implementations are more particular about where you may use backslash continuation. For example, they may not allow you to split a string constant using backslash continuation. Thus, for maximum portability of your awk programs, it is best not to split your lines in the middle of a regular expression or a string.

CAUTION: Backslash continuation does not work as described with the C shell. It works for awk programs in files and for one-shot programs, provided you are using a POSIX-compliant shell, such as the Unix Bourne shell or Bash. But the C shell behaves differently! There you must use two backslashes in a row, followed by a newline. Note also that when using the C shell, every newline in your awk program must be escaped with a backslash. To illustrate:

          % awk 'BEGIN { \
          ?     print \\
          ?             "hello, world" \
          ? }'
          ⊣ hello, world

Here, the ‘%’ and ‘?’ are the C shell’s primary and secondary prompts, analogous to the standard shell’s ‘$’ and ‘>’. Compare the previous example to how it is done with a POSIX-compliant shell:

          $ awk 'BEGIN {
          >   print \
          >       "hello, world"
          > }'
          ⊣ hello, world

awk is a line-oriented language. Each rule’s action has to begin on the same line as the pattern. To have the pattern and action on separate lines, you must use backslash continuation; there is no other option.

Another thing to keep in mind is that backslash continuation and comments do not mix. As soon as awk sees the ‘#’ that starts a comment, it ignores everything on the rest of the line. For example:

     $ gawk 'BEGIN { print "dont panic" # a friendly \
     >                                    BEGIN rule
     > }'
      error gawk: cmd. line:2:                BEGIN rule
      error gawk: cmd. line:2:                ^syntax error

In this case, it looks like the backslash would continue the comment onto the next line. However, the backslash-newline combination is never even noticed because it is “hidden” inside the comment. Thus, the BEGIN is noted as a syntax error.

When awk statements within one rule are short, you might want to put more than one of them on a line. This is accomplished by separating the statements with a semicolon (‘;’). This also applies to the rules themselves. Thus, the program shown at the start of this section could also be written this way:

     /12/ { print $0 } ; /21/ { print $0 }

NOTE: The requirement that states that rules on the same line must be separated with a semicolon was not in the original awk language; it was added for consistency with the treatment of statements within an action.

1.7 Other Features of awk

The awk language provides a number of predefined, or built-in, variables that your programs can use to get information from awk. There are other variables your program can set as well to control how awk processes your data.

In addition, awk provides a number of built-in functions for doing common computational and string-related operations. gawk provides built-in functions for working with timestamps, performing bit manipulation, for runtime string translation (internationalization), determining the type of a variable, and array sorting.

As we develop our presentation of the awk language, we will introduce most of the variables and many of the functions. They are described systematically in Section 7.5 [Predefined Variables], page 147, and in Section 9.1 [Built-in Functions], page 175.

1.8 When to Use awk

Now that you’ve seen some of what awk can do, you might wonder how awk could be useful for you. By using utility programs, advanced patterns, field separators, arithmetic statements, and other selection criteria, you can produce much more complex output. The awk language is very useful for producing reports from large amounts of raw data, such as summarizing information from the output of other utility programs like ls. (See Section 1.5 [A More Complex Example], page 26.)

Programs written with awk are usually much smaller than they would be in other languages. This makes awk programs easy to compose and use. Often, awk programs can be quickly composed at your keyboard, used once, and thrown away. Because awk programs are interpreted, you can avoid the (usually lengthy) compilation part of the typical edit-compile-test-debug cycle of software development.

Complex programs have been written in awk, including a complete retargetable assembler for eight-bit microprocessors (see [Glossary], page 463, for more information), and a microcode assembler for a special-purpose Prolog computer. The original awk’s capabilities were strained by tasks of such complexity, but modern versions are more capable.

If you find yourself writing awk scripts of more than, say, a few hundred lines, you might consider using a different programming language. The shell is good at string and pattern matching; in addition, it allows powerful use of the system utilities. Python offers a nice balance between high-level ease of programming and access to system facilities.^{3}

1.9 Summary

• Programs in awk consist of pattern–action pairs.

• An action without a pattern always runs. The default action for a pattern without one

  is ‘{ print $0 }’.
• Use either ‘awk 'program' files’ or ‘awk -f program-file files’ to run awk.

• You may use the special ‘#!’ header line to create awk programs that are directly executable.

• Comments in awk programs start with ‘#’ and continue to the end of the same line.

• Be aware of quoting issues when writing awk programs as part of a larger shell script (or MS-Windows batch file).

• You may use backslash continuation to continue a source line. Lines are automatically continued after a comma, open brace, question mark, colon, ‘||’, ‘&&’, do, and else.

3 Other popular scripting languages include Ruby and Perl.

Chapter 2: Running awk and gawk 31

2 Running awk and gawk

This chapter covers how to run awk, both POSIX-standard and gawk-specific command-line options, and what awk and gawk do with nonoption arguments. It then proceeds to cover how gawk searches for source files, reading standard input along with other files, gawk’s environment variables, gawk’s exit status, using include files, and obsolete and undocumented options and/or features.

Many of the options and features described here are discussed in more detail later in the book; feel free to skip over things in this chapter that don’t interest you right now.

2.1 Invoking awk

There are two ways to run awk—with an explicit program or with one or more program files. Here are templates for both of them; items enclosed in […] in these templates are optional:

awk [options] -f progfile [--] file … awk [options] [--] 'program' file …

In addition to traditional one-letter POSIX-style options, gawk also supports GNU long options.

It is possible to invoke awk with an empty program:

     awk '' datafile1 datafile2

Doing so makes little sense, though; awk exits silently when given an empty program. If --lint has been specified on the command line, gawk issues a warning that the program is empty.

2.2 Command-Line Options

Options begin with a dash and consist of a single character. GNU-style long options consist of two dashes and a keyword. The keyword can be abbreviated, as long as the abbreviation allows the option to be uniquely identified. If the option takes an argument, either the keyword is immediately followed by an equals sign (‘=’) and the argument’s value, or the keyword and the argument’s value are separated by whitespace. If a particular option with a value is given more than once, it is the last value that counts.

Each long option for gawk has a corresponding POSIX-style short option. The long and short options are interchangeable in all contexts. The following list describes options mandated by the POSIX standard:

-F fs
--field-separator fs

Set the FS variable to fs (see Section 4.5 [Specifying How Fields Are Separated], page 67).

-f source-file
--file source-file

Read the awk program source from source-file instead of in the first nonoption argument. This option may be given multiple times; the awk program consists of the concatenation of the contents of each specified source-file.

-v var=val
--assign var=val

Set the variable var to the value val before execution of the program begins. Such variable values are available inside the BEGIN rule (see Section 2.3 [Other Command-Line Arguments], page 37). The -v option can only set one variable, but it can be used more than once, setting another variable each time, like this: ‘awk -v foo=1 -v bar=2 ...’.

CAUTION: Using -v to set the values of the built-in variables may lead to surprising results. awk will reset the values of those variables as it needs to, possibly ignoring any initial value you may have given.

-W gawk-opt

Provide an implementation-specific option. This is the POSIX convention for providing implementation-specific options. These options also have corresponding GNU-style long options. Note that the long options may be abbreviated, as long as the abbreviations remain unique. The full list of gawk-specific options is provided next.

-- Signal the end of the command-line options. The following arguments are not treated as options even if they begin with ‘-’. This interpretation of -- follows the POSIX argument parsing conventions. This is useful if you have file names that start with ‘-’, or in shell scripts, if you have file names that will be specified by the user that could start with ‘-’. It is also useful for passing options on to the awk program; see Section 10.4 [Processing Command-Line Options], page 233.

The following list describes gawk-specific options:

-b
--characters-as-bytes

Cause gawk to treat all input data as single-byte characters. In addition, all output written with print or printf is treated as single-byte characters. Normally, gawk follows the POSIX standard and attempts to process its input data according to the current locale (see Section 6.6 [Where You Are Makes a Difference], page 129). This can often involve converting multibyte characters into wide characters (internally), and can lead to problems or confusion if the input data does not contain valid multibyte characters. This option is an easy way to tell gawk, “Hands off my data!”

-c
--traditional

Specify compatibility mode, in which the GNU extensions to the awk language are disabled, so that gawk behaves just like BWK awk. See Section A.5 [Extensions in gawk Not in POSIX awk], page 415, which summarizes the extensions. Also see Section C.1 [Downward Compatibility and Debugging], page 447.

-C
--copyright

Print the short version of the General Public License and then exit.

-d[file]

--dump-variables[=file]

Print a sorted list of global variables, their types, and final values to file. If no file is provided, print this list to a file named awkvars.out in the current directory. No space is allowed between the -d and file, if file is supplied. Having a list of all global variables is a good way to look for typographical errors in your programs. You would also use this option if you have a large program with a lot of functions, and you want to be sure that your functions don’t inadvertently use global variables that you meant to be local. (This is a particularly easy mistake to make with simple variable names like i, j, etc.)

-D[file]

--debug[=file]

Enable debugging of awk programs (see Section 14.1 [Introduction to the gawk Debugger], page 325). By default, the debugger reads commands interactively from the keyboard (standard input). The optional file argument allows you to specify a file with a list of commands for the debugger to execute noninteractively. No space is allowed between the -D and file, if file is supplied.

-e program-text --source program-text

Provide program source code in the program-text. This option allows you to mix source code in files with source code that you enter on the command line. This is particularly useful when you have library functions that you want to use from your command-line programs (see Section 2.5.1 [The AWKPATH Environment Variable], page 39).

-E file

--exec file

Similar to -f, read awk program text from file. There are two differences from

-f:

• This option terminates option processing; anything else on the command line is passed on directly to the awk program.

• Command-line variable assignments of the form ‘var=value’ are disallowed.

This option is particularly necessary for World Wide Web CGI applications that pass arguments through the URL; using this option prevents a malicious (or other) user from passing in options, assignments, or awk source code (via -e) to the CGI application.^{1} This option should be used with ‘#!’ scripts (see Section 1.1.4 [Executable awk Programs], page 19), like so:

     #! /usr/local/bin/gawk -E
     awk program here ...

1 For more detail, please see Section 4.4 of RFC 3875 (http://www.ietf.org/rfc/rfc3875). Also see the

explanatory note sent to the gawk bug mailing list (http://lists.gnu.org/archive/html/bug-gawk/ 2014-11/msg00022.html).

-g
--gen-pot

Analyze the source program and generate a GNU gettext portable object template file on standard output for all string constants that have been marked for translation. See Chapter 13 [Internationalization with gawk], page 315, for information about this option.

-h

--help Print a “usage” message summarizing the short- and long-style options that gawk accepts and then exit.

-i source-file --include source-file

Read an awk source library from source-file. This option is completely equivalent to using the @include directive inside your program. It is very similar to the -f option, but there are two important differences. First, when -i is used, the program source is not loaded if it has been previously loaded, whereas with -f, gawk always loads the file. Second, because this option is intended to be used with code libraries, gawk does not recognize such files as constituting main program input. Thus, after processing an -i argument, gawk still expects to find the main source code via the -f option or on the command line.

-l ext

--load ext

Load a dynamic extension named ext. Extensions are stored as system shared libraries. This option searches for the library using the AWKLIBPATH environment variable. The correct library suffix for your platform will be supplied by default, so it need not be specified in the extension name. The extension initialization routine should be named dl_load(). An alternative is to use the @load keyword inside the program to load a shared library. This advanced feature is described in detail in Chapter 16 [Writing Extensions for gawk], page 355.

-L[value]

--lint[=value]

Warn about constructs that are dubious or nonportable to other awk implementations. No space is allowed between the -L and value, if value is supplied. Some warnings are issued when gawk first reads your program. Others are issued at runtime, as your program executes. With an optional argument of ‘fatal’, lint warnings become fatal errors. This may be drastic, but its use will certainly encourage the development of cleaner awk programs. With an optional argument of ‘invalid’, only warnings about things that are actually invalid are issued. (This is not fully implemented yet.)

Some warnings are only printed once, even if the dubious constructs they warn about occur multiple times in your awk program. Thus, when eliminating problems pointed out by --lint, you should take care to search for all occurrences of each inappropriate construct. As awk programs are usually short, doing so is not burdensome.

-M

--bignum Force arbitrary-precision arithmetic on numbers. This option has no effect if gawk is not compiled to use the GNU MPFR and MP libraries (see Chapter 15 [Arithmetic and Arbitrary-Precision Arithmetic with gawk], page 341).

-n
--non-decimal-data

Enable automatic interpretation of octal and hexadecimal values in input data (see Section 12.1 [Allowing Nondecimal Input Data], page 299).

CAUTION: This option can severely break old programs. Use with care. Also note that this option may disappear in a future version

     of gawk.
-N
--use-lc-numeric

Force the use of the locale’s decimal point character when parsing numeric input data (see Section 6.6 [Where You Are Makes a Difference], page 129).

-o[file]

--pretty-print[=file]

Enable pretty-printing of awk programs. By default, the output program is created in a file named awkprof.out (see Section 12.5 [Profiling Your awk Programs], page 308). The optional file argument allows you to specify a different file name for the output. No space is allowed between the -o and file, if file is supplied.

NOTE: Due to the way gawk has evolved, with this option your program still executes. This will change in the next major release, such that gawk will only pretty-print the program and not run it.

-O
--optimize

Enable some optimizations on the internal representation of the program. At the moment, this includes just simple constant folding.

-p[file]

--profile[=file]

Enable profiling of awk programs (see Section 12.5 [Profiling Your awk Programs], page 308). By default, profiles are created in a file named awkprof.out. The optional file argument allows you to specify a different file name for the profile file. No space is allowed between the -p and file, if file is supplied. The profile contains execution counts for each statement in the program in the left margin, and function call counts for each function.

-P

--posix Operate in strict POSIX mode. This disables all gawk extensions (just like --traditional) and disables all extensions not allowed by POSIX. See Section A.7 [Common Extensions Summary], page 423, for a summary of the extensions in gawk that are disabled by this option. Also, the following additional restrictions apply:

• Newlines do not act as whitespace to separate fields when FS is equal to a single space (see Section 4.2 [Examining Fields], page 63).

• Newlines are not allowed after ‘?’ or ‘:’ (see Section 6.3.4 [Conditional Expressions], page 125).

• Specifying ‘-Ft’ on the command line does not set the value of FS to be a single TAB character (see Section 4.5 [Specifying How Fields Are Separated], page 67).

• The locale’s decimal point character is used for parsing input data (see Section 6.6 [Where You Are Makes a Difference], page 129).

If you supply both --traditional and --posix on the command line, --posix takes precedence. gawk issues a warning if both options are supplied.

-r
--re-interval

Allow interval expressions (see Section 3.3 [Regular Expression Operators], page 48) in regexps. This is now gawk’s default behavior. Nevertheless, this option remains (both for backward compatibility and for use in combination

with --traditional).
-S
--sandbox

Disable the system() function, input redirections with getline, output redirections with print and printf, and dynamic extensions. This is particularly useful when you want to run awk scripts from questionable sources and need to make sure the scripts can’t access your system (other than the specified input data file).

-t
--lint-old

Warn about constructs that are not available in the original version of awk from Version 7 Unix (see Section A.1 [Major Changes Between V7 and SVR3.1], page 413).

-V
--version

Print version information for this particular copy of gawk. This allows you to determine if your copy of gawk is up to date with respect to whatever the Free Software Foundation is currently distributing. It is also useful for bug reports (see Section B.4 [Reporting Problems and Bugs], page 442).

As long as program text has been supplied, any other options are flagged as invalid with a warning message but are otherwise ignored.

In compatibility mode, as a special case, if the value of fs supplied to the -F option is ‘t’, then FS is set to the TAB character ("\t"). This is true only for --traditional and not for --posix (see Section 4.5 [Specifying How Fields Are Separated], page 67).

The -f option may be used more than once on the command line. If it is, awk reads its program source from all of the named files, as if they had been concatenated together into one big file. This is useful for creating libraries of awk functions. These functions can be written once and then retrieved from a standard place, instead of having to be included in each individual program. The -i option is similar in this regard. (As mentioned in Section 9.2.1 [Function Definition Syntax], page 199, function names must be unique.)

With standard awk, library functions can still be used, even if the program is entered at the keyboard, by specifying ‘-f /dev/tty’. After typing your program, type Ctrl-d (the end-of-file character) to terminate it. (You may also use ‘-f -’ to read program source from the standard input, but then you will not be able to also use the standard input as a source of data.)

Because it is clumsy using the standard awk mechanisms to mix source file and commandline awk programs, gawk provides the -e option. This does not require you to preempt the standard input for your source code; it allows you to easily mix command-line and library source code (see Section 2.5.1 [The AWKPATH Environment Variable], page 39). As with -f, the -e and -i options may also be used multiple times on the command line.

If no -f or -e option is specified, then gawk uses the first nonoption command-line argument as the text of the program source code.

If the environment variable POSIXLY_CORRECT exists, then gawk behaves in strict POSIX mode, exactly as if you had supplied --posix. Many GNU programs look for this environment variable to suppress extensions that conflict with POSIX, but gawk behaves differently: it suppresses all extensions, even those that do not conflict with POSIX, and behaves in strict POSIX mode. If --lint is supplied on the command line and gawk turns on POSIX mode because of POSIXLY_CORRECT, then it issues a warning message indicating that POSIX mode is in effect. You would typically set this variable in your shell’s startup file. For a Bourne-compatible shell (such as Bash), you would add these lines to the .profile file in your home directory:

     POSIXLY_CORRECT=true
     export POSIXLY_CORRECT

For a C shell-compatible shell,^{2} you would add this line to the .login file in your home directory:

     setenv POSIXLY_CORRECT true

Having POSIXLY_CORRECT set is not recommended for daily use, but it is good for testing the portability of your programs to other environments.

2.3 Other Command-Line Arguments

Any additional arguments on the command line are normally treated as input files to be processed in the order specified. However, an argument that has the form var=value, assigns the value value to the variable var—it does not specify a file at all. (See Section 6.1.3.2 [Assigning Variables on the Command Line], page 111.) In the following example, count=1 is a variable assignment, not a file name:

     awk -f program.awk file1 count=1 file2

All the command-line arguments are made available to your awk program in the ARGV array (see Section 7.5 [Predefined Variables], page 147). Command-line options and the program text (if present) are omitted from ARGV. All other arguments, including variable assignments, are included. As each element of ARGV is processed, gawk sets ARGIND to the index in ARGV of the current element.

2 Not recommended.

Changing ARGC and ARGV in your awk program lets you control how awk processes the input files; this is described in more detail in Section 7.5.3 [Using ARGC and ARGV], page 155.

The distinction between file name arguments and variable-assignment arguments is made when awk is about to open the next input file. At that point in execution, it checks the file name to see whether it is really a variable assignment; if so, awk sets the variable instead of reading a file.

Therefore, the variables actually receive the given values after all previously specified files have been read. In particular, the values of variables assigned in this fashion are not available inside a BEGIN rule (see Section 7.1.4 [The BEGIN and END Special Patterns], page 134), because such rules are run before awk begins scanning the argument list.

The variable values given on the command line are processed for escape sequences (see Section 3.2 [Escape Sequences], page 46).

In some very early implementations of awk, when a variable assignment occurred before any file names, the assignment would happen before the BEGIN rule was executed. awk’s behavior was thus inconsistent; some command-line assignments were available inside the BEGIN rule, while others were not. Unfortunately, some applications came to depend upon this “feature.” When awk was changed to be more consistent, the -v option was added to accommodate applications that depended upon the old behavior.

The variable assignment feature is most useful for assigning to variables such as RS, OFS, and ORS, which control input and output formats, before scanning the data files. It is also useful for controlling state if multiple passes are needed over a data file. For example:

     awk 'pass == 1  { pass 1 stuff }
          pass == 2  { pass 2 stuff }' pass=1 mydata pass=2 mydata

Given the variable assignment feature, the -F option for setting the value of FS is not strictly necessary. It remains for historical compatibility.

2.4 Naming Standard Input

Often, you may wish to read standard input together with other files. For example, you may wish to read one file, read standard input coming from a pipe, and then read another file.

The way to name the standard input, with all versions of awk, is to use a single, standalone minus sign or dash, ‘-’. For example:

     some_command | awk -f myprog.awk file1 - file2

Here, awk first reads file1, then it reads the output of some command, and finally it reads

file2.

You may also use "-" to name standard input when reading files with getline (see Section 4.9.3 [Using getline from a File], page 80).

In addition, gawk allows you to specify the special file name /dev/stdin, both on the command line and with getline. Some other versions of awk also support this, but it is not standard. (Some operating systems provide a /dev/stdin file in the filesystem; however, gawk always processes this file name itself.)

2.5 The Environment Variables gawk Uses

A number of environment variables influence how gawk behaves.

2.5.1 The AWKPATH Environment Variable

In most awk implementations, you must supply a precise pathname for each program file, unless the file is in the current directory. But with gawk, if the file name supplied to the -f or -i options does not contain a directory separator ‘/’, then gawk searches a list of directories (called the search path) one by one, looking for a file with the specified name.

The search path is a string consisting of directory names separated by colons.^{3} gawk gets its search path from the AWKPATH environment variable. If that variable does not exist, or if it has an empty value, gawk uses a default path (described shortly).

The search path feature is particularly helpful for building libraries of useful awk functions. The library files can be placed in a standard directory in the default path and then specified on the command line with a short file name. Otherwise, you would have to type the full file name for each file.

By using the -i or -f options, your command-line awk programs can use facilities in awk library files (see Chapter 10 [A Library of awk Functions], page 217). Path searching is not done if gawk is in compatibility mode. This is true for both --traditional and --posix. See Section 2.2 [Command-Line Options], page 31.

If the source code file is not found after the initial search, the path is searched again after adding the suffix ‘.awk’ to the file name.

gawk’s path search mechanism is similar to the shell’s. (See The Bourne-Again SHell

manual (http://www.gnu.org/software/bash/manual/).) It treats a null entry in the

path as indicating the current directory. (A null entry is indicated by starting or ending the path with a colon or by placing two colons next to each other [‘::’].)

NOTE: To include the current directory in the path, either place . as an entry in the path or write a null entry in the path. Different past versions of gawk would also look explicitly in the current directory, either before or after the path search. As of version 4.1.2, this no longer happens; if you wish to look in the current directory, you must include . either as a separate entry or as a null entry in the search path.

The default value for AWKPATH is ‘.:/usr/local/share/awk’.^{4} Since . is included at the beginning, gawk searches first in the current directory and then in /usr/local/share/awk. In practice, this means that you will rarely need to change the value of AWKPATH.

gawk places the value of the search path that it used into ENVIRON["AWKPATH"]. This provides access to the actual search path value from within an awk program.

Although you can change ENVIRON["AWKPATH"] within your awk program, this has no effect on the running program’s behavior. This makes sense: the AWKPATH environment variable is used to find the program source files. Once your program is running, all the files have been found, and gawk no longer needs to use AWKPATH.

3 Semicolons on MS-Windows and MS-DOS.

4 Your version of gawk may use a different directory; it will depend upon how gawk was built and installed.

The actual directory is the value of $(datadir) generated when gawk was configured. You probably

don’t need to worry about this, though.

2.5.2 The AWKLIBPATH Environment Variable

The AWKLIBPATH environment variable is similar to the AWKPATH variable, but it is used to search for loadable extensions (stored as system shared libraries) specified with the -l option rather than for source files. If the extension is not found, the path is searched again after adding the appropriate shared library suffix for the platform. For example, on GNU/Linux systems, the suffix ‘.so’ is used. The search path specified is also used for extensions loaded via the @load keyword (see Section 2.8 [Loading Dynamic Extensions into Your Program], page 43).

If AWKLIBPATH does not exist in the environment, or if it has an empty value, gawk uses a default path; this is typically ‘/usr/local/lib/gawk’, although it can vary depending upon how gawk was built.

gawk places the value of the search path that it used into ENVIRON["AWKLIBPATH"]. This provides access to the actual search path value from within an awk program.

2.5.3 Other Environment Variables

A number of other environment variables affect gawk’s behavior, but they are more specialized. Those in the following list are meant to be used by regular users:

GAWK_MSEC_SLEEP

Specifies the interval between connection retries, in milliseconds. On systems that do not support the usleep() system call, the value is rounded up to an integral number of seconds.

GAWK_READ_TIMEOUT

Specifies the time, in milliseconds, for gawk to wait for input before returning with an error. See Section 4.10 [Reading Input with a Timeout], page 84.

GAWK_SOCK_RETRIES

Controls the number of times gawk attempts to retry a two-way TCP/IP (socket) connection before giving up. See Section 12.4 [Using gawk for Network Programming], page 307.

POSIXLY_CORRECT

Causes gawk to switch to POSIX-compatibility mode, disabling all traditional and GNU extensions. See Section 2.2 [Command-Line Options], page 31.

The environment variables in the following list are meant for use by the gawk developers for testing and tuning. They are subject to change. The variables are:

AWKBUFSIZE

This variable only affects gawk on POSIX-compliant systems. With a value of ‘exact’, gawk uses the size of each input file as the size of the memory buffer to allocate for I/O. Otherwise, the value should be a number, and gawk uses that number as the size of the buffer to allocate. (When this variable is not set, gawk uses the smaller of the file’s size and the “default” blocksize, which is usually the filesystem’s I/O blocksize.)

AWK_HASH If this variable exists with a value of ‘gst’, gawk switches to using the hash function from GNU Smalltalk for managing arrays. This function may be marginally faster than the standard function.

AWKREADFUNC

If this variable exists, gawk switches to reading source files one line at a time, instead of reading in blocks. This exists for debugging problems on filesystems on non-POSIX operating systems where I/O is performed in records, not in blocks.

GAWK_MSG_SRC

If this variable exists, gawk includes the file name and line number within the gawk source code from which warning and/or fatal messages are generated. Its purpose is to help isolate the source of a message, as there are multiple places that produce the same warning or error message.

GAWK_NO_DFA

If this variable exists, gawk does not use the DFA regexp matcher for “does it match” kinds of tests. This can cause gawk to be slower. Its purpose is to help isolate differences between the two regexp matchers that gawk uses internally. (There aren’t supposed to be differences, but occasionally theory and practice don’t coordinate with each other.)

GAWK_NO_PP_RUN

When gawk is invoked with the --pretty-print option, it will not run the program if this environment variable exists.

CAUTION: This variable will not survive into the next major release.

GAWK_STACKSIZE

This specifies the amount by which gawk should grow its internal evaluation stack, when needed.

INT_CHAIN_MAX

This specifies intended maximum number of items gawk will maintain on a hash chain for managing arrays indexed by integers.

STR_CHAIN_MAX

This specifies intended maximum number of items gawk will maintain on a hash chain for managing arrays indexed by strings.

TIDYMEM If this variable exists, gawk uses the mtrace() library calls from the GNU C library to help track down possible memory leaks.

2.6 gawk’s Exit Status

If the exit statement is used with a value (see Section 7.4.10 [The exit Statement], page 146), then gawk exits with the numeric value given to it.

Otherwise, if there were no problems during execution, gawk exits with the value of the C constant EXIT_SUCCESS. This is usually zero.

If an error occurs, gawk exits with the value of the C constant EXIT_FAILURE. This is usually one.

If gawk exits because of a fatal error, the exit status is two. On non-POSIX systems, this value may be mapped to EXIT_FAILURE.

2.7 Including Other Files into Your Program

This section describes a feature that is specific to gawk.

The @include keyword can be used to read external awk source files. This gives you the ability to split large awk source files into smaller, more manageable pieces, and also lets you reuse common awk code from various awk scripts. In other words, you can group together awk functions used to carry out specific tasks into external files. These files can be used just like function libraries, using the @include keyword in conjunction with the AWKPATH environment variable. Note that source files may also be included using the -i option.

Let’s see an example. We’ll start with two (trivial) awk scripts, namely test1 and test2. Here is the test1 script:

BEGIN {
    print "This is script test1."
}

and here is test2:

@include "test1"
BEGIN {
    print "This is script test2."
}

Running gawk with test2 produces the following result:

$ gawk -f test2
⊣ This is script test1.
⊣ This is script test2.

gawk runs the test2 script, which includes test1 using the @include keyword. So, to include external awk source files, you just use @include followed by the name of the file to be included, enclosed in double quotes.

NOTE: Keep in mind that this is a language construct and the file name cannot be a string variable, but rather just a literal string constant in double quotes.

The files to be included may be nested; e.g., given a third script, namely test3:

@include "test2"
BEGIN {
    print "This is script test3."
}

Running gawk with the test3 script produces the following results:

$ gawk -f test3
⊣ This is script test1.
⊣ This is script test2.
⊣ This is script test3.

The file name can, of course, be a pathname. For example:

@include "../io_funcs"

and:

@include "/usr/awklib/network"

are both valid. The AWKPATH environment variable can be of great value when using @include. The same rules for the use of the AWKPATH variable in command-line file searches (see Section 2.5.1 [The AWKPATH Environment Variable], page 39) apply to @include also.

This is very helpful in constructing gawk function libraries. If you have a large script with useful, general-purpose awk functions, you can break it down into library files and put those files in a special directory. You can then include those “libraries,” either by using the full pathnames of the files, or by setting the AWKPATH environment variable accordingly and then using @include with just the file part of the full pathname. Of course, you can keep library files in more than one directory; the more complex the working environment is, the more directories you may need to organize the files to be included.

Given the ability to specify multiple -f options, the @include mechanism is not strictly necessary. However, the @include keyword can help you in constructing self-contained gawk programs, thus reducing the need for writing complex and tedious command lines. In particular, @include is very useful for writing CGI scripts to be run from web pages.

As mentioned in Section 2.5.1 [The AWKPATH Environment Variable], page 39, the current directory is always searched first for source files, before searching in AWKPATH; this also applies to files named with @include.

2.8 Loading Dynamic Extensions into Your Program

This section describes a feature that is specific to gawk.

The @load keyword can be used to read external awk extensions (stored as system shared libraries). This allows you to link in compiled code that may offer superior performance and/or give you access to extended capabilities not supported by the awk language. The AWKLIBPATH variable is used to search for the extension. Using @load is completely equivalent to using the -l command-line option.

If the extension is not initially found in AWKLIBPATH, another search is conducted after appending the platform’s default shared library suffix to the file name. For example, on GNU/Linux systems, the suffix ‘.so’ is used:

     $ gawk '@load "ordchr"; BEGIN {print chr(65)}'
     ⊣ A

This is equivalent to the following example:

     $ gawk -lordchr 'BEGIN {print chr(65)}'
     ⊣ A

For command-line usage, the -l option is more convenient, but @load is useful for embedding inside an awk source file that requires access to an extension.

Chapter 16 [Writing Extensions for gawk], page 355, describes how to write extensions (in C or C++) that can be loaded with either @load or the -l option. It also describes the ordchr extension.

2.9 Obsolete Options and/or Features

This section describes features and/or command-line options from previous releases of gawk that either are not available in the current version or are still supported but deprecated (meaning that they will not be in the next release).

The process-related special files /dev/pid, /dev/ppid, /dev/pgrpid, and /dev/user

were deprecated in gawk 3.1, but still worked. As of version 4.0, they are no longer interpreted specially by gawk. (Use PROCINFO instead; see Section 7.5.2 [Built-in Variables That Convey Information], page 149.)

2.10 Undocumented Options and Features

Use the Source, Luke!

—Obi-Wan This section intentionally left blank.

2.11 Summary

• Use either ‘awk 'program' files’ or ‘awk -f program-file files’ to run awk.

• The three standard options for all versions of awk are -f, -F, and -v. gawk supplies these and many others, as well as corresponding GNU-style long options.

• Nonoption command-line arguments are usually treated as file names, unless they have the form ‘var=value’, in which case they are taken as variable assignments to be performed at that point in processing the input.

• All nonoption command-line arguments, excluding the program text, are placed in the ARGV array. Adjusting ARGC and ARGV affects how awk processes input.

• You can use a single minus sign (‘-’) to refer to standard input on the command line. gawk also lets you use the special file name /dev/stdin.

• gawk pays attention to a number of environment variables. AWKPATH, AWKLIBPATH, and POSIXLY_CORRECT are the most important ones.

• gawk’s exit status conveys information to the program that invoked it. Use the exit statement from within an awk program to set the exit status.

• gawk allows you to include other awk source files into your program using the @include statement and/or the -i and -f command-line options.

• gawk allows you to load additional functions written in C or C++ using the @load statement and/or the -l option. (This advanced feature is described later, in Chapter 16 [Writing Extensions for gawk], page 355.)

Chapter 3: Regular Expressions 45

3 Regular Expressions

A regular expression, or regexp, is a way of describing a set of strings. Because regular expressions are such a fundamental part of awk programming, their format and use deserve a separate chapter.

A regular expression enclosed in slashes (‘/’) is an awk pattern that matches every input record whose text belongs to that set. The simplest regular expression is a sequence of letters, numbers, or both. Such a regexp matches any string that contains that sequence. Thus, the regexp ‘foo’ matches any string containing ‘foo’. Thus, the pattern /foo/ matches any input record containing the three adjacent characters ‘foo’ anywhere in the record. Other kinds of regexps let you specify more complicated classes of strings.

Initially, the examples in this chapter are simple. As we explain more about how regular expressions work, we present more complicated instances.

3.1 How to Use Regular Expressions

A regular expression can be used as a pattern by enclosing it in slashes. Then the regular expression is tested against the entire text of each record. (Normally, it only needs to match some part of the text in order to succeed.) For example, the following prints the second field of each record where the string ‘li’ appears anywhere in the record:

     $ awk '/li/ { print $2 }' mail-list
     ⊣ 555-5553
     ⊣ 555-0542
     ⊣ 555-6699
     ⊣ 555-3430

Regular expressions can also be used in matching expressions. These expressions allow you to specify the string to match against; it need not be the entire current input record. The two operators ‘~’ and ‘!~’ perform regular expression comparisons. Expressions using these operators can be used as patterns, or in if, while, for, and do statements. (See Section 7.4 [Control Statements in Actions], page 138.) For example, the following is true if the expression exp (taken as a string) matches regexp:

     exp ~ /regexp/

This example matches, or selects, all input records with the uppercase letter ‘J’ somewhere in the first field:

     $ awk '$1 ~ /J/' inventory-shipped
     ⊣ Jan  13  25  15 115
     ⊣ Jun  31  42  75 492
     ⊣ Jul  24  34  67 436
     ⊣ Jan  21  36  64 620

So does this:

     awk '{ if ($1 ~ /J/) print }' inventory-shipped

This next example is true if the expression exp (taken as a character string) does not match regexp:

     exp !~ /regexp/

The following example matches, or selects, all input records whose first field does not contain the uppercase letter ‘J’:

     $ awk '$1 !~ /J/' inventory-shipped
     ⊣ Feb  15  32  24 226
     ⊣ Mar  15  24  34 228
     ⊣ Apr  31  52  63 420
     ⊣ May  16  34  29 208
     ...

When a regexp is enclosed in slashes, such as /foo/, we call it a regexp constant, much like 5.27 is a numeric constant and "foo" is a string constant.

3.2 Escape Sequences

Some characters cannot be included literally in string constants ("foo") or regexp constants (/foo/). Instead, they should be represented with escape sequences, which are character sequences beginning with a backslash (‘\’). One use of an escape sequence is to include a double-quote character in a string constant. Because a plain double quote ends the string, you must use ‘\"’ to represent an actual double-quote character as a part of the string. For example:

     $ awk 'BEGIN { print "He said \"hi!\" to her." }'
     ⊣ He said "hi!" to her.

The backslash character itself is another character that cannot be included normally; you must write ‘\\’ to put one backslash in the string or regexp. Thus, the string whose contents are the two characters ‘"’ and ‘\’ must be written "\"\\".

Other escape sequences represent unprintable characters such as TAB or newline. There is nothing to stop you from entering most unprintable characters directly in a string constant or regexp constant, but they may look ugly.

The following list presents all the escape sequences used in awk and what they represent. Unless noted otherwise, all these escape sequences apply to both string constants and regexp constants:

\\ A literal backslash, ‘\’.

\a The “alert” character, Ctrl-g, ASCII code 7 (BEL). (This often makes some sort of audible noise.)

\b Backspace, Ctrl-h, ASCII code 8 (BS).

\f Formfeed, Ctrl-l, ASCII code 12 (FF).

\n Newline, Ctrl-j, ASCII code 10 (LF).

\r Carriage return, Ctrl-m, ASCII code 13 (CR).

\t Horizontal TAB, Ctrl-i, ASCII code 9 (HT).

\v Vertical TAB, Ctrl-k, ASCII code 11 (VT).

\nnn The octal value nnn, where nnn stands for 1 to 3 digits between ‘0’ and ‘7’. For example, the code for the ASCII ESC (escape) character is ‘\033’.

\xhh... The hexadecimal value hh, where hh stands for a sequence of hexadecimal digits (‘0’–‘9’, and either ‘A’–‘F’ or ‘a’–‘f’). Like the same construct in ISO C, the escape sequence continues until the first nonhexadecimal digit is seen. (c.e.) However, using more than two hexadecimal digits produces undefined results. (The ‘\x’ escape sequence is not allowed in POSIX awk.)

CAUTION: The next major release of gawk will change, such that a maximum of two hexadecimal digits following the ‘\x’ will be used.

\/ A literal slash (necessary for regexp constants only). This sequence is used when you want to write a regexp constant that contains a slash (such as

/.*:\/home\/[[:alnum:]]+:.*/; the ‘[[:alnum:]]’ notation is discussed in

Section 3.4 [Using Bracket Expressions], page 50). Because the regexp is delimited by slashes, you need to escape any slash that is part of the pattern, in order to tell awk to keep processing the rest of the regexp.

\" A literal double quote (necessary for string constants only). This sequence is used when you want to write a string constant that contains a double quote (such as "He said \"hi!\" to her."). Because the string is delimited by double quotes, you need to escape any quote that is part of the string, in order to tell awk to keep processing the rest of the string.

In gawk, a number of additional two-character sequences that begin with a backslash have special meaning in regexps. See Section 3.7 [gawk-Specific Regexp Operators], page 54.

In a regexp, a backslash before any character that is not in the previous list and not listed in Section 3.7 [gawk-Specific Regexp Operators], page 54, means that the next character should be taken literally, even if it would normally be a regexp operator. For example, /a\+b/ matches the three characters ‘a+b’.

For complete portability, do not use a backslash before any character not shown in the previous list or that is not an operator.

Backslash Before Regular Characters

If you place a backslash in a string constant before something that is not one of the characters previously listed, POSIX awk purposely leaves what happens as undefined. There are two choices:

Strip the backslash out This is what BWK awk and gawk both do. For example, "a\qc" is the same as "aqc". (Because this is such an easy bug both to introduce and to miss, gawk warns you about it.) Consider ‘FS = "[ \t]+\|[ \t]+"’ to use vertical bars surrounded by whitespace as the field separator. There should be two backslashes in the string: ‘FS = "[ \t]+\\|[ \t]+"’.)

Leave the backslash alone Some other awk implementations do this. In such implementations, typing "a\qc" is the same as typing "a\\qc".

To summarize:

• The escape sequences in the preceding list are always processed first, for both string constants and regexp constants. This happens very early, as soon as awk reads your program.

• gawk processes both regexp constants and dynamic regexps (see Section 3.6 [Using Dynamic Regexps], page 52), for the special operators listed in Section 3.7 [gawk-Specific Regexp Operators], page 54.

• A backslash before any other character means to treat that character literally.

Escape Sequences for Metacharacters

Suppose you use an octal or hexadecimal escape to represent a regexp metacharacter. (See Section 3.3 [Regular Expression Operators], page 48.) Does awk treat the character as a literal character or as a regexp operator?

Historically, such characters were taken literally. However, the POSIX standard indicates that they should be treated as real metacharacters, which is what gawk does. In compatibility mode (see Section 2.2 [Command-Line Options], page 31), gawk treats the characters represented by octal and hexadecimal escape sequences literally when used in regexp constants. Thus, /a\52b/ is equivalent to /a\*b/.

3.3 Regular Expression Operators

You can combine regular expressions with special characters, called regular expression operators or metacharacters, to increase the power and versatility of regular expressions.

The escape sequences described earlier in Section 3.2 [Escape Sequences], page 46, are valid inside a regexp. They are introduced by a ‘\’ and are recognized and converted into corresponding real characters as the very first step in processing regexps.

Here is a list of metacharacters. All characters that are not escape sequences and that are not listed here stand for themselves:

\ This suppresses the special meaning of a character when matching. For example, ‘\$’ matches the character ‘$’.

^This matches the beginning of a string. ‘^@chapter’ matches ‘@chapter’ at the beginning of a string, for example, and can be used to identify chapter beginnings in Texinfo source files. The ‘^’ is known as an anchor, because it anchors the pattern to match only at the beginning of the string. It is important to realize that ‘^’ does not match the beginning of a line (the point right after a ‘\n’ newline character) embedded in a string. The condition is not true in the following example:

     if ("line1\nLINE 2" ~ /^L/) ...

$ This is similar to ‘^’, but it matches only at the end of a string. For example, ‘p$’ matches a record that ends with a ‘p’. The ‘$’ is an anchor and does not match the end of a line (the point right before a ‘\n’ newline character) embedded in a string. The condition in the following example is not true:

     if ("line1\nLINE 2" ~ /1$/) ...

. (period) This matches any single character, including the newline character. For example, ‘.P’ matches any single character followed by a ‘P’ in a string. Using concatenation, we can make a regular expression such as ‘U.A’, which matches any three-character sequence that begins with ‘U’ and ends with ‘A’. In strict POSIX mode (see Section 2.2 [Command-Line Options], page 31), ‘.’ does not match the nul character, which is a character with all bits equal to zero. Otherwise, nul is just another character. Other versions of awk may not be able to match the nul character.

[…] This is called a bracket expression.^{1} It matches any one of the characters that are enclosed in the square brackets. For example, ‘[MVX]’ matches any one of the characters ‘M’, ‘V’, or ‘X’ in a string. A full discussion of what can be inside the square brackets of a bracket expression is given in Section 3.4 [Using Bracket Expressions], page 50.

[^…] This is a complemented bracket expression. The first character after the ‘[’ must be a ‘^’. It matches any characters except those in the square brackets. For example, ‘[^awk]’ matches any character that is not an ‘a’, ‘w’, or ‘k’.

| This is the alternation operator and it is used to specify alternatives. The ‘|’ has the lowest precedence of all the regular expression operators. For example, ‘^P|[aeiouy]’ matches any string that matches either ‘^P’ or ‘[aeiouy]’. This means it matches any string that starts with ‘P’ or contains (anywhere within it) a lowercase English vowel. The alternation applies to the largest possible regexps on either side.

(…) Parentheses are used for grouping in regular expressions, as in arithmetic. They can be used to concatenate regular expressions containing the alternation operator, ‘|’. For example, ‘@(samp|code)\{[^}]+\}’ matches both ‘@code{foo}’ and ‘@samp{bar}’. (These are Texinfo formatting control sequences. The ‘+’ is explained further on in this list.)

* This symbol means that the preceding regular expression should be repeated as many times as necessary to find a match. For example, ‘ph*’ applies the ‘*’ symbol to the preceding ‘h’ and looks for matches of one ‘p’ followed by any number of ‘h’s. This also matches just ‘p’ if no ‘h’s are present. There are two subtle points to understand about how ‘*’ works. First, the ‘*’ applies only to the single preceding regular expression component (e.g., in ‘ph*’, it applies just to the ‘h’). To cause ‘*’ to apply to a larger subexpression, use parentheses: ‘(ph)*’ matches ‘ph’, ‘phph’, ‘phphph’, and so on. Second, ‘*’ finds as many repetitions as possible. If the text to be matched is ‘phhhhhhhhhhhhhhooey’, ‘ph*’ matches all of the ‘h’s.

+ This symbol is similar to ‘*’, except that the preceding expression must be matched at least once. This means that ‘wh+y’ would match ‘why’ and ‘whhy’, but not ‘wy’, whereas ‘wh*y’ would match all three.

1 In other literature, you may see a bracket expression referred to as either a character set, a character

class, or a character list.

? This symbol is similar to ‘*’, except that the preceding expression can be matched either once or not at all. For example, ‘fe?d’ matches ‘fed’ and ‘fd’, but nothing else.

{n}
{n,}

{n,m} One or two numbers inside braces denote an interval expression. If there is one number in the braces, the preceding regexp is repeated n times. If there are two numbers separated by a comma, the preceding regexp is repeated n to m times. If there is one number followed by a comma, then the preceding regexp is repeated at least n times:

wh{3}y Matches ‘whhhy’, but not ‘why’ or ‘whhhhy’.

wh{3,5}y Matches ‘whhhy’, ‘whhhhy’, or ‘whhhhhy’ only.

wh{2,}y Matches ‘whhy’, ‘whhhy’, and so on.

Interval expressions were not traditionally available in awk. They were added as part of the POSIX standard to make awk and egrep consistent with each other. Initially, because old programs may use ‘{’ and ‘}’ in regexp constants, gawk did not match interval expressions in regexps. However, beginning with version 4.0, gawk does match interval expressions by default. This is because compatibility with POSIX has become more important to most gawk users than compatibility with old programs. For programs that use ‘{’ and ‘}’ in regexp constants, it is good practice to always escape them with a backslash. Then the regexp constants are valid and work the way you want them to, using any version of awk.^{2} Finally, when ‘{’ and ‘}’ appear in regexp constants in a way that cannot be interpreted as an interval expression (such as /q{a}/), then they stand for themselves.

In regular expressions, the ‘*’, ‘+’, and ‘?’ operators, as well as the braces ‘{’ and ‘}’, have the highest precedence, followed by concatenation, and finally by ‘|’. As in arithmetic, parentheses can change how operators are grouped.

In POSIX awk and gawk, the ‘*’, ‘+’, and ‘?’ operators stand for themselves when there is nothing in the regexp that precedes them. For example, /+/ matches a literal plus sign. However, many other versions of awk treat such a usage as a syntax error.

If gawk is in compatibility mode (see Section 2.2 [Command-Line Options], page 31), interval expressions are not available in regular expressions.

3.4 Using Bracket Expressions

As mentioned earlier, a bracket expression matches any character among those listed between the opening and closing square brackets.

Within a bracket expression, a range expression consists of two characters separated by a hyphen. It matches any single character that sorts between the two characters, based upon the system’s native character set. For example, ‘[0-9]’ is equivalent to ‘[0123456789]’. (See Section A.8 [Regexp Ranges and Locales: A Long Sad Story], page 424, for an explanation of how the POSIX standard and gawk have changed over time. This is mainly of historical interest.)

2 Use two backslashes if you’re using a string constant with a regexp operator or function.

To include one of the characters ‘\’, ‘]’, ‘-’, or ‘^’ in a bracket expression, put a ‘\’ in front of it. For example:

     [d\]]

matches either ‘d’ or ‘]’. Additionally, if you place ‘]’ right after the opening ‘[’, the closing bracket is treated as one of the characters to be matched.

The treatment of ‘\’ in bracket expressions is compatible with other awk implementations and is also mandated by POSIX. The regular expressions in awk are a superset of the POSIX specification for Extended Regular Expressions (EREs). POSIX EREs are based on the regular expressions accepted by the traditional egrep utility.

Character classes are a feature introduced in the POSIX standard. A character class is a special notation for describing lists of characters that have a specific attribute, but the actual characters can vary from country to country and/or from character set to character set. For example, the notion of what is an alphabetic character differs between the United States and France.

A character class is only valid in a regexp inside the brackets of a bracket expression. Character classes consist of ‘[:’, a keyword denoting the class, and ‘:]’. Table 3.1 lists the character classes defined by the POSIX standard.

Class  Meaning
[:alnum:]  Alphanumeric characters
[:alpha:]  Alphabetic characters
[:blank:]  Space and TAB characters
[:cntrl:]  Control characters
[:digit:]  Numeric characters

[:graph:]

Characters that are both printable and visible (a space is printable but not visible, whereas an ‘a’ is both)

[:lower:]  Lowercase alphabetic characters
[:print:]  Printable characters (characters that are not control characters)

[:punct:]

Punctuation characters (characters that are not letters, digits, control characters, or space characters)

[:space:]  Space characters (such as space, TAB, and formfeed, to name a few)
[:upper:]  Uppercase alphabetic characters
[:xdigit:] Characters that are hexadecimal digits

Table 3.1: POSIX character classes

For example, before the POSIX standard, you had to write /[A-Za-z0-9]/ to match alphanumeric characters. If your character set had other alphabetic characters in it, this would not match them. With the POSIX character classes, you can write /[[:alnum:]]/ to match the alphabetic and numeric characters in your character set.

Some utilities that match regular expressions provide a nonstandard ‘[:ascii:]’ character class; awk does not. However, you can simulate such a construct using ‘[\x00-\x7F]’.

This matches all values numerically between zero and 127, which is the defined range of the ASCII character set. Use a complemented character list (‘[^\x00-\x7F]’) to match any single-byte characters that are not in the ASCII range.

Two additional special sequences can appear in bracket expressions. These apply to non-ASCII character sets, which can have single symbols (called collating elements) that are represented with more than one character. They can also have several characters that are equivalent for collating, or sorting, purposes. (For example, in French, a plain “e” and a grave-accented “è” are equivalent.) These sequences are:

Collating symbols Multicharacter collating elements enclosed between ‘[.’ and ‘.]’. For example, if ‘ch’ is a collating element, then ‘[[.ch.]]’ is a regexp that matches this collating element, whereas ‘[ch]’ is a regexp that matches either ‘c’ or ‘h’.

Equivalence classes Locale-specific names for a list of characters that are equal. The name is enclosed between ‘[=’ and ‘=]’. For example, the name ‘e’ might be used to represent all of “e,” “ê,” “è,” and “é.” In this case, ‘[[=e=]]’ is a regexp that matches any of ‘e’, ‘^e’, ‘é’, or ‘è’.

These features are very valuable in non-English-speaking locales.

CAUTION: The library functions that gawk uses for regular expression matching currently recognize only POSIX character classes; they do not recognize collating symbols or equivalence classes.

3.5 How Much Text Matches?

Consider the following:

     echo aaaabcd | awk '{ sub(/a+/, "<A>"); print }'

This example uses the sub() function to make a change to the input record. (sub() replaces the first instance of any text matched by the first argument with the string provided as the second argument; see Section 9.1.3 [String-Manipulation Functions], page 177.) Here, the regexp /a+/ indicates “one or more ‘a’ characters,” and the replacement text is ‘<A>’.

The input contains four ‘a’ characters. awk (and POSIX) regular expressions always match the leftmost, longest sequence of input characters that can match. Thus, all four ‘a’ characters are replaced with ‘<A>’ in this example:

     $ echo aaaabcd | awk '{ sub(/a+/, "<A>"); print }'
     ⊣ <A>bcd

For simple match/no-match tests, this is not so important. But when doing text matching and substitutions with the match(), sub(), gsub(), and gensub() functions, it is very important. Understanding this principle is also important for regexp-based record and field splitting (see Section 4.1 [How Input Is Split into Records], page 59, and also see Section 4.5 [Specifying How Fields Are Separated], page 67).

3.6 Using Dynamic Regexps

The righthand side of a ‘~’ or ‘!~’ operator need not be a regexp constant (i.e., a string of characters between slashes). It may be any expression. The expression is evaluated and converted to a string if necessary; the contents of the string are then used as the regexp. A regexp computed in this way is called a dynamic regexp or a computed regexp:

BEGIN { digits_regexp = "[[:digit:]]+" }
$0 ~ digits_regexp    { print }

This sets digits_regexp to a regexp that describes one or more digits, and tests whether the input record matches this regexp.

NOTE: When using the ‘~’ and ‘!~’ operators, be aware that there is a difference between a regexp constant enclosed in slashes and a string constant enclosed in double quotes. If you are going to use a string constant, you have to understand that the string is, in essence, scanned twice: the first time when awk reads your program, and the second time when it goes to match the string on the lefthand side of the operator with the pattern on the right. This is true of any stringvalued expression (such as digits_regexp, shown in the previous example), not just string constants.

What difference does it make if the string is scanned twice? The answer has to do with escape sequences, and particularly with backslashes. To get a backslash into a regular expression inside a string, you have to type two backslashes.

For example, /\*/ is a regexp constant for a literal ‘*’. Only one backslash is needed. To do the same thing with a string, you have to type "\\*". The first backslash escapes the second one so that the string actually contains the two characters ‘\’ and ‘*’.

Given that you can use both regexp and string constants to describe regular expressions, which should you use? The answer is “regexp constants,” for several reasons:

• String constants are more complicated to write and more difficult to read. Using regexp constants makes your programs less error-prone. Not understanding the difference between the two kinds of constants is a common source of errors.

• It is more efficient to use regexp constants. awk can note that you have supplied a regexp and store it internally in a form that makes pattern matching more efficient. When using a string constant, awk must first convert the string into this internal form and then perform the pattern matching.

• Using regexp constants is better form; it shows clearly that you intend a regexp match.

Using \n in Bracket Expressions of Dynamic Regexps Some older versions of awk do not allow the newline character to be used inside a bracket

expression for a dynamic regexp:

      $ awk '$0 ~ "[ \t\n]"'
       error awk: newline in character class [

error ]...

       error source line number 1
       error context is
       error  $0 ~ "[ >>>  \t\n]" <<<

But a newline in a regexp constant works with no problem:

      $ awk '$0 ~ /[ \t\n]/'
      here is a sample line
      ⊣ here is a sample line
      Ctrl-d

gawk does not have this problem, and it isn’t likely to occur often in practice, but it’s worth noting for future reference.

3.7 gawk-Specific Regexp Operators

GNU software that deals with regular expressions provides a number of additional regexp operators. These operators are described in this section and are specific to gawk; they are not available in other awk implementations. Most of the additional operators deal with word matching. For our purposes, a word is a sequence of one or more letters, digits, or underscores (‘_’):

\s Matches any whitespace character. Think of it as shorthand for ‘[[:space:]]’.

\S Matches any character that is not whitespace. Think of it as shorthand for

          ‘[^[:space:]]’.

\w Matches any word-constituent character—that is, it matches any letter, digit, or underscore. Think of it as shorthand for ‘[[:alnum:]_]’.

\W Matches any character that is not word-constituent. Think of it as shorthand

          for ‘[^[:alnum:]_]’.

\< Matches the empty string at the beginning of a word. For example, /\<away/ matches ‘away’ but not ‘stowaway’.

\> Matches the empty string at the end of a word. For example, /stow\>/ matches

          ‘stow’ but not ‘stowaway’.

\y Matches the empty string at either the beginning or the end of a word (i.e., the word boundary). For example, ‘\yballs?\y’ matches either ‘ball’ or ‘balls’, as a separate word.

\B Matches the empty string that occurs between two word-constituent characters. For example, /\Brat\B/ matches ‘crate’, but it does not match ‘dirty rat’. ‘\B’ is essentially the opposite of ‘\y’.

There are two other operators that work on buffers. In Emacs, a buffer is, naturally, an Emacs buffer. Other GNU programs, including gawk, consider the entire string to match as the buffer. The operators are:

\` Matches the empty string at the beginning of a buffer (string)

\' Matches the empty string at the end of a buffer (string)

Because ‘^’ and ‘$’ always work in terms of the beginning and end of strings, these operators don’t add any new capabilities for awk. They are provided for compatibility with other GNU software.

In other GNU software, the word-boundary operator is ‘\b’. However, that conflicts with the awk language’s definition of ‘\b’ as backspace, so gawk uses a different letter. An alternative method would have been to require two backslashes in the GNU operators, but this was deemed too confusing. The current method of using ‘\y’ for the GNU ‘\b’ appears to be the lesser of two evils.

The various command-line options (see Section 2.2 [Command-Line Options], page 31) control how gawk interprets characters in regexps:

No options In the default case, gawk provides all the facilities of POSIX regexps and the previously described GNU regexp operators.

--posix Match only POSIX regexps; the GNU operators are not special (e.g., ‘\w’ matches a literal ‘w’). Interval expressions are allowed.

--traditional

Match traditional Unix awk regexps. The GNU operators are not special, and interval expressions are not available. Because BWK awk supports them, the POSIX character classes (‘[[:alnum:]]’, etc.) are available. Characters described by octal and hexadecimal escape sequences are treated literally, even if they represent regexp metacharacters.

--re-interval

Allow interval expressions in regexps, if --traditional has been provided. Otherwise, interval expressions are available by default.

3.8 Case Sensitivity in Matching

Case is normally significant in regular expressions, both when matching ordinary characters (i.e., not metacharacters) and inside bracket expressions. Thus, a ‘w’ in a regular expression matches only a lowercase ‘w’ and not an uppercase ‘W’.

The simplest way to do a case-independent match is to use a bracket expression—for example, ‘[Ww]’. However, this can be cumbersome if you need to use it often, and it can make the regular expressions harder to read. There are two alternatives that you might prefer.

One way to perform a case-insensitive match at a particular point in the program is to convert the data to a single case, using the tolower() or toupper() built-in string functions (which we haven’t discussed yet; see Section 9.1.3 [String-Manipulation Functions], page 177). For example:

     tolower($1) ~ /foo/  { ... }

converts the first field to lowercase before matching against it. This works in any POSIXcompliant awk.

Another method, specific to gawk, is to set the variable IGNORECASE to a nonzero value (see Section 7.5 [Predefined Variables], page 147). When IGNORECASE is not zero, all regexp and string operations ignore case.

Changing the value of IGNORECASE dynamically controls the case sensitivity of the program as it runs. Case is significant by default because IGNORECASE (like most variables) is initialized to zero:

     x = "aB"
     if (x ~ /ab/) ...   # this test will fail
     IGNORECASE = 1
     if (x ~ /ab/) ...   # now it will succeed

In general, you cannot use IGNORECASE to make certain rules case insensitive and other rules case sensitive, as there is no straightforward way to set IGNORECASE just for the pattern of a particular rule.^{3} To do this, use either bracket expressions or tolower(). However, one thing you can do with IGNORECASE only is dynamically turn case sensitivity on or off for all the rules at once.

IGNORECASE can be set on the command line or in a BEGIN rule (see Section 2.3 [Other Command-Line Arguments], page 37; also see Section 7.1.4.1 [Startup and Cleanup Actions], page 134). Setting IGNORECASE from the command line is a way to make a program case insensitive without having to edit it.

In multibyte locales, the equivalences between upper- and lowercase characters are tested based on the wide-character values of the locale’s character set. Otherwise, the characters are tested based on the ISO-8859-1 (ISO Latin-1) character set. This character set is a superset of the traditional 128 ASCII characters, which also provides a number of characters suitable for use with European languages.^{4}

The value of IGNORECASE has no effect if gawk is in compatibility mode (see Section 2.2 [Command-Line Options], page 31). Case is always significant in compatibility mode.

3.9 Summary

• Regular expressions describe sets of strings to be matched. In awk, regular expression constants are written enclosed between slashes: /…/.

• Regexp constants may be used standalone in patterns and in conditional expressions, or as part of matching expressions using the ‘~’ and ‘!~’ operators.

• Escape sequences let you represent nonprintable characters and also let you represent regexp metacharacters as literal characters to be matched.

• Regexp operators provide grouping, alternation, and repetition.

• Bracket expressions give you a shorthand for specifying sets of characters that can match at a particular point in a regexp. Within bracket expressions, POSIX character classes let you specify certain groups of characters in a locale-independent fashion.

3 Experienced C and C++ programmers will note that it is possible, using something like ‘IGNORECASE =

1 && /foObAr/ { ... }’ and ‘IGNORECASE = 0 || /foobar/ { ... }’. However, this is somewhat obscure

and we don’t recommend it.

4 If you don’t understand this, don’t worry about it; it just means that gawk does the right thing.

• Regular expressions match the leftmost longest text in the string being matched. This matters for cases where you need to know the extent of the match, such as for text substitution and when the record separator is a regexp.

• Matching expressions may use dynamic regexps (i.e., string values treated as regular expressions).

• gawk’s IGNORECASE variable lets you control the case sensitivity of regexp matching. In

other awk versions, use tolower() or toupper().

Chapter 4: Reading Input Files 59

4 Reading Input Files

In the typical awk program, awk reads all input either from the standard input (by default, this is the keyboard, but often it is a pipe from another command) or from files whose names you specify on the awk command line. If you specify input files, awk reads them in order, processing all the data from one before going on to the next. The name of the current input file can be found in the predefined variable FILENAME (see Section 7.5 [Predefined Variables], page 147).

The input is read in units called records, and is processed by the rules of your program one record at a time. By default, each record is one line. Each record is automatically split into chunks called fields. This makes it more convenient for programs to work on the parts of a record.

On rare occasions, you may need to use the getline command. The getline command is valuable both because it can do explicit input from any number of files, and because the files used with it do not have to be named on the awk command line (see Section 4.9 [Explicit Input with getline], page 78).

4.1 How Input Is Split into Records

awk divides the input for your program into records and fields. It keeps track of the number of records that have been read so far from the current input file. This value is stored in a predefined variable called FNR, which is reset to zero every time a new file is started. Another predefined variable, NR, records the total number of input records read so far from all data files. It starts at zero, but is never automatically reset to zero.

4.1.1 Record Splitting with Standard awk

Records are separated by a character called the record separator. By default, the record separator is the newline character. This is why records are, by default, single lines. To use a different character for the record separator, simply assign that character to the predefined variable RS.

Like any other variable, the value of RS can be changed in the awk program with the assignment operator, ‘=’ (see Section 6.2.3 [Assignment Expressions], page 116). The new record-separator character should be enclosed in quotation marks, which indicate a string constant. Often, the right time to do this is at the beginning of execution, before any input is processed, so that the very first record is read with the proper separator. To do this, use the special BEGIN pattern (see Section 7.1.4 [The BEGIN and END Special Patterns], page 134). For example:

     awk 'BEGIN { RS = "u" }
          { print $0 }' mail-list

changes the value of RS to ‘u’, before reading any input. The new value is a string whose first character is the letter “u”; as a result, records are separated by the letter “u”. Then the input file is read, and the second rule in the awk program (the action with no pattern) prints each record. Because each print statement adds a newline at the end of its output, this awk program copies the input with each ‘u’ changed to a newline. Here are the results of running the program on mail-list:

$ awk 'BEGIN { RS = "u" }
>      { print $0 }' mail-list
⊣ Amelia       555-5553     amelia.zodiac
⊣ sq
⊣ e@gmail.com    F
⊣ Anthony      555-3412     anthony.assert
⊣ ro@hotmail.com   A
⊣ Becky        555-7685     becky.algebrar
⊣ m@gmail.com      A
⊣ Bill         555-1675     bill.drowning@hotmail.com       A
⊣ Broderick    555-0542     broderick.aliq
⊣ otiens@yahoo.com R
⊣ Camilla      555-2912     camilla.inf
⊣ sar
⊣ m@skynet.be     R
⊣ Fabi
⊣ s       555-1234     fabi
⊣ s.
⊣ ndevicesim
⊣ s@
⊣ cb.ed
⊣  F
⊣ J
⊣ lie        555-6699     j
⊣ lie.perscr
⊣ tabor@skeeve.com   F
⊣ Martin       555-6480     martin.codicib
⊣ s@hotmail.com    A
⊣ Sam
⊣ el       555-3430     sam
⊣ el.lanceolis@sh
⊣ .ed
⊣  A
⊣ Jean-Pa
⊣ l    555-2127     jeanpa
⊣ l.campanor
⊣ m@ny
⊣ .ed
⊣  R

⊣

Note that the entry for the name ‘Bill’ is not split. In the original data file (see Section 1.2 [Data files for the Examples], page 23), the line looks like this:

Bill         555-1675     bill.drowning@hotmail.com       A

It contains no ‘u’, so there is no reason to split the record, unlike the others, which each have one or more occurrences of the ‘u’. In fact, this record is treated as part of the previous record; the newline separating them in the output is the original newline in the data file, not the one added by awk when it printed the record!

Another way to change the record separator is on the command line, using the variableassignment feature (see Section 2.3 [Other Command-Line Arguments], page 37):

     awk '{ print $0 }' RS="u" mail-list

This sets RS to ‘u’ before processing mail-list.

Using an alphabetic character such as ‘u’ for the record separator is highly likely to produce strange results. Using an unusual character such as ‘/’ is more likely to produce correct behavior in the majority of cases, but there are no guarantees. The moral is: Know Your Data.

When using regular characters as the record separator, there is one unusual case that occurs when gawk is being fully POSIX-compliant (see Section 2.2 [Command-Line Options], page 31). Then, the following (extreme) pipeline prints a surprising ‘1’:

     $ echo | gawk --posix 'BEGIN { RS = "a" } ; { print NF }'
     ⊣ 1

There is one field, consisting of a newline. The value of the built-in variable NF is the number of fields in the current record. (In the normal case, gawk treats the newline as whitespace, printing ‘0’ as the result. Most other versions of awk also act this way.)

Reaching the end of an input file terminates the current input record, even if the last character in the file is not the character in RS.

The empty string "" (a string without any characters) has a special meaning as the value of RS. It means that records are separated by one or more blank lines and nothing else. See Section 4.8 [Multiple-Line Records], page 76, for more details.

If you change the value of RS in the middle of an awk run, the new value is used to delimit subsequent records, but the record currently being processed, as well as records already processed, are not affected.

After the end of the record has been determined, gawk sets the variable RT to the text in the input that matched RS.

4.1.2 Record Splitting with gawk

When using gawk, the value of RS is not limited to a one-character string. It can be any regular expression (see Chapter 3 [Regular Expressions], page 45). (c.e.) In general, each record ends at the next string that matches the regular expression; the next record starts at the end of the matching string. This general rule is actually at work in the usual case, where RS contains just a newline: a record ends at the beginning of the next matching string (the next newline in the input), and the following record starts just after the end of this string (at the first character of the following line). The newline, because it matches RS, is not part of either record.

When RS is a single character, RT contains the same single character. However, when RS is a regular expression, RT contains the actual input text that matched the regular expression.

If the input file ends without any text matching RS, gawk sets RT to the null string. The following example illustrates both of these features. It sets RS equal to a regular

expression that matches either a newline or a series of one or more uppercase letters with optional leading and/or trailing whitespace:

     $ echo record 1 AAAA record 2 BBBB record 3 |
     > gawk 'BEGIN { RS = "\n|( *[[:upper:]]+ *)" }
     >             { print "Record =", $0,"and RT = [" RT "]" }'
     ⊣ Record = record 1 and RT = [ AAAA ]
     ⊣ Record = record 2 and RT = [ BBBB ]
     ⊣ Record = record 3 and RT = [
     ⊣ ]

The square brackets delineate the contents of RT, letting you see the leading and trailing whitespace. The final value of RT is a newline. See Section 11.3.8 [A Simple Stream Editor], page 284, for a more useful example of RS as a regexp and RT.

If you set RS to a regular expression that allows optional trailing text, such as ‘RS = "abc(XYZ)?"’, it is possible, due to implementation constraints, that gawk may match the leading part of the regular expression, but not the trailing part, particularly if the input text that could match the trailing part is fairly long. gawk attempts to avoid this problem, but currently, there’s no guarantee that this will never happen.

NOTE: Remember that in awk, the ‘^’ and ‘$’ anchor metacharacters match the beginning and end of a string, and not the beginning and end of a line. As a result, something like ‘RS = "^[[:upper:]]"’ can only match at the beginning of a file. This is because gawk views the input file as one long string that happens to contain newline characters. It is thus best to avoid anchor metacharacters in the value of RS.

The use of RS as a regular expression and the RT variable are gawk extensions; they are not available in compatibility mode (see Section 2.2 [Command-Line Options], page 31). In compatibility mode, only the first character of the value of RS determines the end of the record.

RS = "\0" Is Not Portable There are times when you might want to treat an entire data file as a single record. The

only way to make this happen is to give RS a value that you know doesn’t occur in the input file. This is hard to do in a general way, such that a program always works for arbitrary input files.

You might think that for text files, the nul character, which consists of a character with all bits equal to zero, is a good value to use for RS in this case:

      BEGIN { RS = "\0" }  # whole file becomes one record?

gawk in fact accepts this, and uses the nul character for the record separator. This works for certain special files, such as /proc/environ on GNU/Linux systems, where the nul character is in fact the record separator. However, this usage is not portable to most other awk implementations.

Almost all other awk implementations^{1} store strings internally as C-style strings. C strings use the nul character as the string terminator. In effect, this means that ‘RS = "\0"’ is the same as ‘RS = ""’.

It happens that recent versions of mawk can use the nul character as a record separator. However, this is a special case: mawk does not allow embedded nul characters in strings. (This may change in a future version of mawk.)

See Section 10.2.8 [Reading a Whole File at Once], page 227, for an interesting way to read whole files. If you are using gawk, see Section 16.7.10 [Reading an Entire File], page 406, for another option.

4.2 Examining Fields

When awk reads an input record, the record is automatically parsed or separated by the awk utility into chunks called fields. By default, fields are separated by whitespace, like words in a line. Whitespace in awk means any string of one or more spaces, TABs, or newlines;^{2} other characters that are considered whitespace by other languages (such as formfeed, vertical tab, etc.) are not considered whitespace by awk.

The purpose of fields is to make it more convenient for you to refer to these pieces of the record. You don’t have to use them—you can operate on the whole record if you want—but fields are what make simple awk programs so powerful.

You use a dollar sign (‘$’) to refer to a field in an awk program, followed by the number of the field you want. Thus, $1 refers to the first field, $2 to the second, and so on. (Unlike in the Unix shells, the field numbers are not limited to single digits. $127 is the 127th field in the record.) For example, suppose the following is a line of input:

     This seems like a pretty nice example.

Here the first field, or $1, is ‘This’, the second field, or $2, is ‘seems’, and so on. Note that the last field, $7, is ‘example.’. Because there is no space between the ‘e’ and the ‘.’, the period is considered part of the seventh field.

NF is a predefined variable whose value is the number of fields in the current record. awk automatically updates the value of NF each time it reads a record. No matter how many fields there are, the last field in a record can be represented by $NF. So, $NF is the same as $7, which is ‘example.’. If you try to reference a field beyond the last one (such as $8 when the record has only seven fields), you get the empty string. (If used in a numeric operation, you get zero.)

1 At least that we know about.

2 In POSIX awk, newlines are not considered whitespace for separating fields.

The use of $0, which looks like a reference to the “zeroth” field, is a special case: it represents the whole input record. Use it when you are not interested in specific fields. Here are some more examples:

     $ awk '$1 ~ /li/ { print $0 }' mail-list
     ⊣ Amelia       555-5553     amelia.zodiacusque@gmail.com    F
     ⊣ Julie        555-6699     julie.perscrutabor@skeeve.com   F

This example prints each record in the file mail-list whose first field contains the string ‘li’.

By contrast, the following example looks for ‘li’ in the entire record and prints the first and last fields for each matching input record:

     $ awk '/li/ { print $1, $NF }' mail-list
     ⊣ Amelia F
     ⊣ Broderick R
     ⊣ Julie F
     ⊣ Samuel A

4.3 Nonconstant Field Numbers

A field number need not be a constant. Any expression in the awk language can be used after a ‘$’ to refer to a field. The value of the expression specifies the field number. If the value is a string, rather than a number, it is converted to a number. Consider this example:

     awk '{ print $NR }'

Recall that NR is the number of records read so far: one in the first record, two in the second, and so on. So this example prints the first field of the first record, the second field of the second record, and so on. For the twentieth record, field number 20 is printed; most likely, the record has fewer than 20 fields, so this prints a blank line. Here is another example of using expressions as field numbers:

     awk '{ print $(2*2) }' mail-list

awk evaluates the expression ‘(2*2)’ and uses its value as the number of the field to print. The ‘*’ represents multiplication, so the expression ‘2*2’ evaluates to four. The parentheses are used so that the multiplication is done before the ‘$’ operation; they are necessary whenever there is a binary operator^{3} in the field-number expression. This example, then, prints the type of relationship (the fourth field) for every line of the file mail-list. (All of the awk operators are listed, in order of decreasing precedence, in Section 6.5 [Operator Precedence (How Operators Nest)], page 128.)

If the field number you compute is zero, you get the entire record. Thus, ‘$(2-2)’ has the same value as $0. Negative field numbers are not allowed; trying to reference one usually terminates the program. (The POSIX standard does not define what happens when you reference a negative field number. gawk notices this and terminates your program. Other awk implementations may behave differently.)

3 A binary operator, such as ‘*’ for multiplication, is one that takes two operands. The distinction is

required because awk also has unary (one-operand) and ternary (three-operand) operators.

As mentioned in Section 4.2 [Examining Fields], page 63, awk stores the current record’s number of fields in the built-in variable NF (also see Section 7.5 [Predefined Variables], page 147). Thus, the expression $NF is not a special feature—it is the direct consequence of evaluating NF and using its value as a field number.

4.4 Changing the Contents of a Field

The contents of a field, as seen by awk, can be changed within an awk program; this changes what awk perceives as the current input record. (The actual input is untouched; awk never modifies the input file.) Consider the following example and its output:

     $ awk '{ nboxes = $3 ; $3 = $3 - 10
     >        print nboxes, $3 }' inventory-shipped
     ⊣ 25 15
     ⊣ 32 22
     ⊣ 24 14
     ...

The program first saves the original value of field three in the variable nboxes. The ‘-’ sign represents subtraction, so this program reassigns field three, $3, as the original value of field three minus ten: ‘$3 - 10’. (See Section 6.2.1 [Arithmetic Operators], page 114.) Then it prints the original and new values for field three. (Someone in the warehouse made a consistent mistake while inventorying the red boxes.)

For this to work, the text in $3 must make sense as a number; the string of characters must be converted to a number for the computer to do arithmetic on it. The number resulting from the subtraction is converted back to a string of characters that then becomes field three. See Section 6.1.4 [Conversion of Strings and Numbers], page 111.

When the value of a field is changed (as perceived by awk), the text of the input record is recalculated to contain the new field where the old one was. In other words, $0 changes to reflect the altered field. Thus, this program prints a copy of the input file, with 10 subtracted from the second field of each line:

     $ awk '{ $2 = $2 - 10; print $0 }' inventory-shipped
     ⊣ Jan 3 25 15 115
     ⊣ Feb 5 32 24 226
     ⊣ Mar 5 24 34 228
     ...

It is also possible to assign contents to fields that are out of range. For example:

     $ awk '{ $6 = ($5 + $4 + $3 + $2)
     >        print $6 }' inventory-shipped
     ⊣ 168
     ⊣ 297
     ⊣ 301
     ...

We’ve just created $6, whose value is the sum of fields $2, $3, $4, and $5. The ‘+’ sign represents addition. For the file inventory-shipped, $6 represents the total number of parcels shipped for a particular month.

Creating a new field changes awk’s internal copy of the current input record, which is the value of $0. Thus, if you do ‘print $0’ after adding a field, the record printed includes the new field, with the appropriate number of field separators between it and the previously existing fields.

This recomputation affects and is affected by NF (the number of fields; see Section 4.2 [Examining Fields], page 63). For example, the value of NF is set to the number of the highest field you create. The exact format of $0 is also affected by a feature that has not been discussed yet: the output field separator, OFS, used to separate the fields (see Section 5.3 [Output Separators], page 91).

Note, however, that merely referencing an out-of-range field does not change the value of either $0 or NF. Referencing an out-of-range field only produces an empty string. For example:

     if ($(NF+1) != "")
         print "can't happen"
     else
         print "everything is normal"

should print ‘everything is normal’, because NF+1 is certain to be out of range. (See Section 7.4.1 [The if-else Statement], page 138, for more information about awk’s if-else statements. See Section 6.3.2 [Variable Typing and Comparison Expressions], page 120, for more information about the ‘!=’ operator.)

It is important to note that making an assignment to an existing field changes the value of $0 but does not change the value of NF, even when you assign the empty string to a field. For example:

     $ echo a b c d | awk '{ OFS = ":"; $2 = ""
     >                       print $0; print NF }'
     ⊣ a::c:d
     ⊣ 4

The field is still there; it just has an empty value, delimited by the two colons between ‘a’ and ‘c’. This example shows what happens if you create a new field:

     $ echo a b c d | awk '{ OFS = ":"; $2 = ""; $6 = "new"
     >                       print $0; print NF }'
     ⊣ a::c:d::new
     ⊣ 6

The intervening field, $5, is created with an empty value (indicated by the second pair of adjacent colons), and NF is updated with the value six.

Decrementing NF throws away the values of the fields after the new value of NF and recomputes $0. Here is an example:

     $ echo a b c d e f | awk '{ print "NF =", NF;
     >                           NF = 3; print $0 }'
     ⊣ NF = 6
     ⊣ a b c

CAUTION: Some versions of awk don’t rebuild $0 when NF is decremented. Finally, there are times when it is convenient to force awk to rebuild the entire record,

using the current values of the fields and OFS. To do this, use the seemingly innocuous assignment:

     $1 = $1   # force record to be reconstituted
     print $0  # or whatever else with $0

This forces awk to rebuild the record. It does help to add a comment, as we’ve shown here.

There is a flip side to the relationship between $0 and the fields. Any assignment to $0 causes the record to be reparsed into fields using the current value of FS. This also applies to any built-in function that updates $0, such as sub() and gsub() (see Section 9.1.3 [String-Manipulation Functions], page 177).

Understanding $0 It is important to remember that $0 is the full record, exactly as it was read from the

input. This includes any leading or trailing whitespace, and the exact whitespace (or other characters) that separates the fields.

It is a common error to try to change the field separators in a record simply by setting FS and OFS, and then expecting a plain ‘print’ or ‘print $0’ to print the modified record.

But this does not work, because nothing was done to change the record itself. Instead, you must force the record to be rebuilt, typically with a statement such as ‘$1 = $1’, as described earlier.

4.5 Specifying How Fields Are Separated

The field separator, which is either a single character or a regular expression, controls the way awk splits an input record into fields. awk scans the input record for character sequences that match the separator; the fields themselves are the text between the matches.

In the examples that follow, we use the bullet symbol (•) to represent spaces in the output. If the field separator is ‘oo’, then the following line:

     moo goo gai pan

is split into three fields: ‘m’, ‘•g’, and ‘•gai•pan’. Note the leading spaces in the values of the second and third fields.

The field separator is represented by the predefined variable FS. Shell programmers take note: awk does not use the name IFS that is used by the POSIX-compliant shells (such as the Unix Bourne shell, sh, or Bash).

The value of FS can be changed in the awk program with the assignment operator, ‘=’ (see Section 6.2.3 [Assignment Expressions], page 116). Often, the right time to do this is at the beginning of execution before any input has been processed, so that the very first record is read with the proper separator. To do this, use the special BEGIN pattern (see Section 7.1.4 [The BEGIN and END Special Patterns], page 134). For example, here we set the value of FS to the string ",":

     awk 'BEGIN { FS = "," } ; { print $2 }'

Given the input line:

     John Q. Smith, 29 Oak St., Walamazoo, MI 42139

this awk program extracts and prints the string ‘•29•Oak•St.’.

Sometimes the input data contains separator characters that don’t separate fields the way you thought they would. For instance, the person’s name in the example we just used might have a title or suffix attached, such as:

     John Q. Smith, LXIX, 29 Oak St., Walamazoo, MI 42139

The same program would extract ‘•LXIX’ instead of ‘•29•Oak•St.’. If you were expecting the program to print the address, you would be surprised. The moral is to choose your data layout and separator characters carefully to prevent such problems. (If the data is not in a form that is easy to process, perhaps you can massage it first with a separate awk program.)

4.5.1 Whitespace Normally Separates Fields

Fields are normally separated by whitespace sequences (spaces, TABs, and newlines), not by single spaces. Two spaces in a row do not delimit an empty field. The default value of the field separator FS is a string containing a single space, " ". If awk interpreted this value in the usual way, each space character would separate fields, so two spaces in a row would make an empty field between them. The reason this does not happen is that a single space as the value of FS is a special case—it is taken to specify the default manner of delimiting fields.

If FS is any other single character, such as ",", then each occurrence of that character separates two fields. Two consecutive occurrences delimit an empty field. If the character occurs at the beginning or the end of the line, that too delimits an empty field. The space character is the only single character that does not follow these rules.

4.5.2 Using Regular Expressions to Separate Fields

The previous subsection discussed the use of single characters or simple strings as the value of FS. More generally, the value of FS may be a string containing any regular expression. In this case, each match in the record for the regular expression separates fields. For example, the assignment:

     FS = ", \t"

makes every area of an input line that consists of a comma followed by a space and a TAB into a field separator.

For a less trivial example of a regular expression, try using single spaces to separate fields the way single commas are used. FS can be set to "[ ]" (left bracket, space, right bracket). This regular expression matches a single space and nothing else (see Chapter 3 [Regular Expressions], page 45).

There is an important difference between the two cases of ‘FS = " "’ (a single space) and ‘FS = "[ \t\n]+"’ (a regular expression matching one or more spaces, TABs, or newlines). For both values of FS, fields are separated by runs (multiple adjacent occurrences) of spaces, TABs, and/or newlines. However, when the value of FS is " ", awk first strips leading and trailing whitespace from the record and then decides where the fields are. For example, the following pipeline prints ‘b’:

     $ echo ' a b c d ' | awk '{ print $2 }'
     ⊣ b

However, this pipeline prints ‘a’ (note the extra spaces around each letter):

     $ echo ' a  b  c  d ' | awk 'BEGIN { FS = "[ \t\n]+" }
     >                                  { print $2 }'
     ⊣ a

In this case, the first field is null, or empty.

The stripping of leading and trailing whitespace also comes into play whenever $0 is recomputed. For instance, study this pipeline:

     $ echo '   a b c d' | awk '{ print; $2 = $2; print }'
     ⊣  a b c d
     ⊣ a b c d

The first print statement prints the record as it was read, with leading whitespace intact. The assignment to $2 rebuilds $0 by concatenating $1 through $NF together, separated by the value of OFS (which is a space by default). Because the leading whitespace was ignored when finding $1, it is not part of the new $0. Finally, the last print statement prints the new $0.

There is an additional subtlety to be aware of when using regular expressions for field splitting. It is not well specified in the POSIX standard, or anywhere else, what ‘^’ means when splitting fields. Does the ‘^’ match only at the beginning of the entire record? Or is each field separator a new string? It turns out that different awk versions answer this question differently, and you should not rely on any specific behavior in your programs.

As a point of information, BWK awk allows ‘^’ to match only at the beginning of the record. gawk also works this way. For example:

     $ echo 'xxAA  xxBxx  C' |
     > gawk -F '(^x+)|( +)' '{ for (i = 1; i <= NF; i++)
     >                             printf "-->%s<--\n", $i }'
     ⊣ --><--
     ⊣ -->AA<--
     ⊣ -->xxBxx<--
     ⊣ -->C<--

4.5.3 Making Each Character a Separate Field

There are times when you may want to examine each character of a record separately. This can be done in gawk by simply assigning the null string ("") to FS. (c.e.) In this case, each individual character in the record becomes a separate field. For example:

     $ echo a b | gawk 'BEGIN { FS = "" }
     >                  {
     >                      for (i = 1; i <= NF; i = i + 1)
     >                          print "Field", i, "is", $i
     >                  }'
     ⊣ Field 1 is a
     ⊣ Field 2 is
     ⊣ Field 3 is b

Traditionally, the behavior of FS equal to "" was not defined. In this case, most versions of Unix awk simply treat the entire record as only having one field. In compatibility mode (see Section 2.2 [Command-Line Options], page 31), if FS is the null string, then gawk also behaves this way.

4.5.4 Setting FS from the Command Line

FS can be set on the command line. Use the -F option to do so. For example:

     awk -F, 'program' input-files

sets FS to the ‘,’ character. Notice that the option uses an uppercase ‘F’ instead of a lowercase ‘f’. The latter option (-f) specifies a file containing an awk program.

The value used for the argument to -F is processed in exactly the same way as assignments to the predefined variable FS. Any special characters in the field separator must be escaped appropriately. For example, to use a ‘\’ as the field separator on the command line, you would have to type:

     # same as FS = "\\"
     awk -F\\\\ '...' files ...

Because ‘\’ is used for quoting in the shell, awk sees ‘-F\\’. Then awk processes the ‘\\’ for escape characters (see Section 3.2 [Escape Sequences], page 46), finally yielding a single ‘\’ to use for the field separator.

As a special case, in compatibility mode (see Section 2.2 [Command-Line Options], page 31), if the argument to -F is ‘t’, then FS is set to the TAB character. If you type ‘-F\t’ at the shell, without any quotes, the ‘\’ gets deleted, so awk figures that you really want your fields to be separated with TABs and not ‘t’s. Use ‘-v FS="t"’ or ‘-F"[t]"’ on the command line if you really do want to separate your fields with ‘t’s. Use ‘-F '\t'’ when not in compatibility mode to specify that TABs separate fields.

As an example, let’s use an awk program file called edu.awk that contains the pattern /edu/ and the action ‘print $1’:

     /edu/   { print $1 }

Let’s also set FS to be the ‘-’ character and run the program on the file mail-list. The following command prints a list of the names of the people that work at or attend a university, and the first three digits of their phone numbers:

     $ awk -F- -f edu.awk mail-list
     ⊣ Fabius       555
     ⊣ Samuel       555
     ⊣ Jean

Note the third line of output. The third line in the original file looked like this:

     Jean-Paul    555-2127     jeanpaul.campanorum@nyu.edu     R

The ‘-’ as part of the person’s name was used as the field separator, instead of the ‘-’ in the phone number that was originally intended. This demonstrates why you have to be careful in choosing your field and record separators.

Perhaps the most common use of a single character as the field separator occurs when processing the Unix system password file. On many Unix systems, each user has a separate entry in the system password file, with one line per user. The information in these lines is separated by colons. The first field is the user’s login name and the second is the user’s encrypted or shadow password. (A shadow password is indicated by the presence of a single ‘x’ in the second field.) A password file entry might look like this:

     arnold:x:2076:10:Arnold Robbins:/home/arnold:/bin/bash

The following program searches the system password file and prints the entries for users whose full name is not indicated:

     awk -F: '$5 == ""' /etc/passwd

4.5.5 Making the Full Line Be a Single Field

Occasionally, it’s useful to treat the whole input line as a single field. This can be done easily and portably simply by setting FS to "\n" (a newline):^{4}

     awk -F'\n' 'program' files ...

When you do this, $1 is the same as $0.

Changing FS Does Not Affect the Fields According to the POSIX standard, awk is supposed to behave as if each record is split

into fields at the time it is read. In particular, this means that if you change the value of FS after a record is read, the values of the fields (i.e., how they were split) should reflect the old value of FS, not the new one.

However, many older implementations of awk do not work this way. Instead, they defer splitting the fields until a field is actually referenced. The fields are split using the current value of FS! This behavior can be difficult to diagnose. The following example illustrates the difference between the two methods:

      sed 1q /etc/passwd | awk '{ FS = ":" ; print $1 }'

which usually prints:

      root

on an incorrect implementation of awk, while gawk prints the full first line of the file, something like:

      root:x:0:0:Root:/:

(The sed^{5} command prints just the first line of /etc/passwd.)

4.5.6 Field-Splitting Summary

It is important to remember that when you assign a string constant as the value of FS, it undergoes normal awk string processing. For example, with Unix awk and gawk, the assignment ‘FS = "\.."’ assigns the character string ".." to FS (the backslash is stripped). This creates a regexp meaning “fields are separated by occurrences of any two characters.” If instead you want fields to be separated by a literal period followed by any single character,

use ‘FS = "\\.."’.

The following list summarizes how fields are split, based on the value of FS (‘==’ means “is equal to”):

FS == " " Fields are separated by runs of whitespace. Leading and trailing whitespace are ignored. This is the default.

FS == any other single character

Fields are separated by each occurrence of the character. Multiple successive occurrences delimit empty fields, as do leading and trailing occurrences. The character can even be a regexp metacharacter; it does not need to be escaped.

4 Thanks to Andrew Schorr for this tip.

5 The sed utility is a “stream editor.” Its behavior is also defined by the POSIX standard.

FS == regexp

Fields are separated by occurrences of characters that match regexp. Leading and trailing matches of regexp delimit empty fields.

FS == "" Each individual character in the record becomes a separate field. (This is a common extension; it is not specified by the POSIX standard.)

                              FS and IGNORECASE

The IGNORECASE variable (see Section 7.5.1 [Built-in Variables That Control awk], page 147) affects field splitting only when the value of FS is a regexp. It has no effect when FS is a single character, even if that character is a letter. Thus, in the following code:

      FS = "c"
      IGNORECASE = 1
      $0 = "aCa"
      print $1

The output is ‘aCa’. If you really want to split fields on an alphabetic character while ignoring case, use a regexp that will do it for you (e.g., ‘FS = "[c]"’). In this case, IGNORECASE will take effect.

4.6 Reading Fixed-Width Data

This section discusses an advanced feature of gawk. If you are a novice awk user, you might want to skip it on the first reading.

gawk provides a facility for dealing with fixed-width fields with no distinctive field separator. For example, data of this nature arises in the input for old Fortran programs where numbers are run together, or in the output of programs that did not anticipate the use of their output as input for other programs.

An example of the latter is a table where all the columns are lined up by the use of a variable number of spaces and empty fields are just spaces. Clearly, awk’s normal field splitting based on FS does not work well in this case. Although a portable awk program can use a series of substr() calls on $0 (see Section 9.1.3 [String-Manipulation Functions], page 177), this is awkward and inefficient for a large number of fields.

The splitting of an input record into fixed-width fields is specified by assigning a string containing space-separated numbers to the built-in variable FIELDWIDTHS. Each number specifies the width of the field, including columns between fields. If you want to ignore the columns between fields, you can specify the width as a separate field that is subsequently ignored. It is a fatal error to supply a field width that has a negative value. The following data is the output of the Unix w utility. It is useful to illustrate the use of FIELDWIDTHS:

 10:06pm  up 21 days, 14:04,  23 users
User     tty       login  idle   JCPU   PCPU  what
hzuo     ttyV0     8:58pm            9      5  vi p24.tex
hzang    ttyV3     6:37pm    50                -csh
eklye    ttyV5     9:53pm            7      1  em thes.tex
dportein ttyV6     8:17pm  1:47                -csh
gierd    ttyD3    10:00pm     1                elm
dave     ttyD4     9:47pm            4      4  w
brent    ttyp0    26Jun91  4:46  26:46   4:41  bash
dave     ttyq4    26Jun9115days     46     46  wnewmail

The following program takes this input, converts the idle time to number of seconds, and prints out the first two fields and the calculated idle time:

BEGIN  { FIELDWIDTHS = "9 6 10 6 7 7 35" }
NR > 2 {
    idle = $4
    sub(/^+/, "", idle)   # strip leading spaces
    if (idle == "")
        idle = 0
    if (idle ~ /:/) {
        split(idle, t, ":")
        idle = t[1] * 60 + t[2]
    }
    if (idle ~ /days/)
        idle *= 24 * 60 * 60
    print $1, $2, idle
}

NOTE: The preceding program uses a number of awk features that haven’t been introduced yet.

Running the program on the data produces the following results:

hzuo      ttyV0  0
hzang     ttyV3  50
eklye     ttyV5  0
dportein  ttyV6  107
gierd     ttyD3  1
dave      ttyD4  0
brent     ttyp0  286
dave      ttyq4  1296000

Another (possibly more practical) example of fixed-width input data is the input from a deck of balloting cards. In some parts of the United States, voters mark their choices by punching holes in computer cards. These cards are then processed to count the votes for any particular candidate or on any particular issue. Because a voter may choose not to vote on some issue, any column on the card may be empty. An awk program for processing such data could use the FIELDWIDTHS feature to simplify reading the data. (Of course, getting gawk to run on a system with card readers is another story!)

Assigning a value to FS causes gawk to use FS for field splitting again. Use ‘FS = FS’ to make this happen, without having to know the current value of FS. In order to tell which kind of field splitting is in effect, use PROCINFO["FS"] (see Section 7.5.2 [Built-in Variables That Convey Information], page 149). The value is "FS" if regular field splitting is being used, or "FIELDWIDTHS" if fixed-width field splitting is being used:

     if (PROCINFO["FS"] == "FS")
         regular field splitting ...
     else if  (PROCINFO["FS"] == "FIELDWIDTHS")
         fixed-width field splitting ...
     else
         content-based field splitting ... (see next section)

This information is useful when writing a function that needs to temporarily change FS or FIELDWIDTHS, read some records, and then restore the original settings (see Section 10.5 [Reading the User Database], page 238, for an example of such a function).

4.7 Defining Fields by Content

This section discusses an advanced feature of gawk. If you are a novice awk user, you might want to skip it on the first reading.

Normally, when using FS, gawk defines the fields as the parts of the record that occur in between each field separator. In other words, FS defines what a field is not, instead of what a field is. However, there are times when you really want to define the fields by what they are, and not by what they are not.

The most notorious such case is so-called comma-separated values (CSV) data. Many spreadsheet programs, for example, can export their data into text files, where each record is terminated with a newline, and fields are separated by commas. If commas only separated the data, there wouldn’t be an issue. The problem comes when one of the fields contains an embedded comma. In such cases, most programs embed the field in double quotes.^{6} So, we might have data like this:

     Robbins,Arnold,"1234 A Pretty Street, NE",MyTown,MyState,12345-6789,USA

The FPAT variable offers a solution for cases like this. The value of FPAT should be a string that provides a regular expression. This regular expression describes the contents of each field.

In the case of CSV data as presented here, each field is either “anything that is not a comma,” or “a double quote, anything that is not a double quote, and a closing double quote.” If written as a regular expression constant (see Chapter 3 [Regular Expressions], page 45), we would have /([^,]+)|("[^"]+")/. Writing this as a string requires us to escape the double quotes, leading to:

     FPAT = "([^,]+)|(\"[^\"]+\")"

Putting this to use, here is a simple program to parse the data:

     BEGIN {
         FPAT = "([^,]+)|(\"[^\"]+\")"

6 The CSV format lacked a formal standard definition for many years. RFC 4180 (http://www.ietf.org/

rfc/rfc4180.txt) standardizes the most common practices.

}
{
    print "NF = ", NF
    for (i = 1; i <= NF; i++) {
        printf("$%d = <%s>\n", i, $i)
    }
}

When run, we get the following:

$ gawk -f simple-csv.awk addresses.csv
NF =  7
$1 = <Robbins>
$2 = <Arnold>
$3 = <"1234 A Pretty Street, NE">
$4 = <MyTown>
$5 = <MyState>
$6 = <12345-6789>
$7 = <USA>

Note the embedded comma in the value of $3.

A straightforward improvement when processing CSV data of this sort would be to remove the quotes when they occur, with something like this:

if (substr($i, 1, 1) == "\"") {
    len = length($i)
    $i = substr($i, 2, len - 2)    # Get text within the two quotes
}

As with FS, the IGNORECASE variable (see Section 7.5.1 [Built-in Variables That Control awk], page 147) affects field splitting with FPAT.

Assigning a value to FPAT overrides field splitting with FS and with FIELDWIDTHS. Similar to FIELDWIDTHS, the value of PROCINFO["FS"] will be "FPAT" if content-based field splitting is being used.

NOTE: Some programs export CSV data that contains embedded newlines between the double quotes. gawk provides no way to deal with this. Even though a formal specification for CSV data exists, there isn’t much more to be done; the FPAT mechanism provides an elegant solution for the majority of cases, and the gawk developers are satisfied with that.

As written, the regexp used for FPAT requires that each field contain at least one character. A straightforward modification (changing the first ‘+’ to ‘*’) allows fields to be empty:

FPAT = "([^,]*)|(\"[^\"]+\")"

Finally, the patsplit() function makes the same functionality available for splitting regular strings (see Section 9.1.3 [String-Manipulation Functions], page 177).

To recap, gawk provides three independent methods to split input records into fields. The mechanism used is based on which of the three variables—FS, FIELDWIDTHS, or FPAT—was last assigned to.

4.8 Multiple-Line Records

In some databases, a single line cannot conveniently hold all the information in one entry. In such cases, you can use multiline records. The first step in doing this is to choose your data format.

One technique is to use an unusual character or string to separate records. For example, you could use the formfeed character (written ‘\f’ in awk, as in C) to separate them, making each record a page of the file. To do this, just set the variable RS to "\f" (a string containing the formfeed character). Any other character could equally well be used, as long as it won’t be part of the data in a record.

Another technique is to have blank lines separate records. By a special dispensation, an empty string as the value of RS indicates that records are separated by one or more blank lines. When RS is set to the empty string, each record always ends at the first blank line encountered. The next record doesn’t start until the first nonblank line that follows. No matter how many blank lines appear in a row, they all act as one record separator. (Blank lines must be completely empty; lines that contain only whitespace do not count.)

You can achieve the same effect as ‘RS = ""’ by assigning the string "\n\n+" to RS. This regexp matches the newline at the end of the record and one or more blank lines after the record. In addition, a regular expression always matches the longest possible sequence when there is a choice (see Section 3.5 [How Much Text Matches?], page 52). So, the next record doesn’t start until the first nonblank line that follows—no matter how many blank lines appear in a row, they are considered one record separator.

However, there is an important difference between ‘RS = ""’ and ‘RS = "\n\n+"’. In the first case, leading newlines in the input data file are ignored, and if a file ends without extra blank lines after the last record, the final newline is removed from the record. In the second case, this special processing is not done.

Now that the input is separated into records, the second step is to separate the fields in the records. One way to do this is to divide each of the lines into fields in the normal manner. This happens by default as the result of a special feature. When RS is set to the empty string and FS is set to a single character, the newline character always acts as a field separator. This is in addition to whatever field separations result from FS.^{7}

The original motivation for this special exception was probably to provide useful behavior in the default case (i.e., FS is equal to " "). This feature can be a problem if you really don’t want the newline character to separate fields, because there is no way to prevent it. However, you can work around this by using the split() function to break up the record manually (see Section 9.1.3 [String-Manipulation Functions], page 177). If you have a single-character field separator, you can work around the special feature in a different way, by making FS into a regexp for that single character. For example, if the field separator is a percent character, instead of ‘FS = "%"’, use ‘FS = "[%]"’.

Another way to separate fields is to put each field on a separate line: to do this, just set the variable FS to the string "\n". (This single-character separator matches a single newline.) A practical example of a data file organized this way might be a mailing list, where blank lines separate the entries. Consider a mailing list in a file named addresses, which looks like this:

7 When FS is the null string ("") or a regexp, this special feature of RS does not apply. It does apply to

the default field separator of a single space: ‘FS = " "’.

Jane Doe
123 Main Street
Anywhere, SE 12345-6789
John Smith
456 Tree-lined Avenue
Smallville, MW 98765-4321
...

A simple program to process this file is as follows:

# addrs.awk --- simple mailing list program
# Records are separated by blank lines.
# Each line is one field.
BEGIN { RS = "" ; FS = "\n" }
{
      print "Name is:", $1
      print "Address is:", $2
      print "City and State are:", $3
      print ""
}

Running the program produces the following output:

$ awk -f addrs.awk addresses
⊣ Name is: Jane Doe
⊣ Address is: 123 Main Street
⊣ City and State are: Anywhere, SE 12345-6789

⊣

⊣ Name is: John Smith
⊣ Address is: 456 Tree-lined Avenue
⊣ City and State are: Smallville, MW 98765-4321

⊣

...

See Section 11.3.4 [Printing Mailing Labels], page 276, for a more realistic program dealing with address lists. The following list summarizes how records are split, based on the value of RS:

RS == "\n"

Records are separated by the newline character (‘\n’). In effect, every line in the data file is a separate record, including blank lines. This is the default.

RS == any single character

Records are separated by each occurrence of the character. Multiple successive occurrences delimit empty records.

RS == "" Records are separated by runs of blank lines. When FS is a single character, then the newline character always serves as a field separator, in addition to whatever value FS may have. Leading and trailing newlines in a file are ignored.

RS == regexp

Records are separated by occurrences of characters that match regexp. Leading and trailing matches of regexp delimit empty records. (This is a gawk extension; it is not specified by the POSIX standard.)

If not in compatibility mode (see Section 2.2 [Command-Line Options], page 31), gawk sets RT to the input text that matched the value specified by RS. But if the input file ended without any text that matches RS, then gawk sets RT to the null string.

4.9 Explicit Input with getline

So far we have been getting our input data from awk’s main input stream—either the standard input (usually your keyboard, sometimes the output from another program) or the files specified on the command line. The awk language has a special built-in command called getline that can be used to read input under your explicit control.

The getline command is used in several different ways and should not be used by beginners. The examples that follow the explanation of the getline command include material that has not been covered yet. Therefore, come back and study the getline command after you have reviewed the rest of Parts I and II and have a good knowledge of how awk works.

The getline command returns 1 if it finds a record and 0 if it encounters the end of the file. If there is some error in getting a record, such as a file that cannot be opened, then getline returns −1. In this case, gawk sets the variable ERRNO to a string describing the error that occurred.

In the following examples, command stands for a string value that represents a shell command.

NOTE: When --sandbox is specified (see Section 2.2 [Command-Line Options], page 31), reading lines from files, pipes, and coprocesses is disabled.

4.9.1 Using getline with No Arguments

The getline command can be used without arguments to read input from the current input file. All it does in this case is read the next input record and split it up into fields. This is useful if you’ve finished processing the current record, but want to do some special processing on the next record right now. For example:

     # Remove text between /* and */, inclusive
     {
         if ((i = index($0, "/*")) != 0) {
             out = substr($0, 1, i - 1)  # leading part of the string
             rest = substr($0, i + 2)    # ... */ ...
             j = index(rest, "*/")       # is */ in trailing part?
             if (j > 0) {
                 rest = substr(rest, j + 2)  # remove comment
             } else {
                 while (j == 0) {
                     # get more text
                     if (getline <= 0) {

                         print("unexpected EOF or error:", ERRNO) > "/dev/stderr"
                         exit
                     }
                     # build up the line using string concatenation
                     rest = rest $0
                     j = index(rest, "*/")   # is */ in trailing part?
                     if (j != 0) {
                         rest = substr(rest, j + 2)
                         break
                     }
                 }
             }
             # build up the output line using string concatenation
             $0 = out rest
         }
         print $0
     }

This awk program deletes C-style comments (‘/* ... */’) from the input. It uses a number of features we haven’t covered yet, including string concatenation (see Section 6.2.2 [String Concatenation], page 115) and the index() and substr() built-in functions (see Section 9.1.3 [String-Manipulation Functions], page 177). By replacing the ‘print $0’ with other statements, you could perform more complicated processing on the decommented input, such as searching for matches of a regular expression. (This program has a subtle problem—it does not work if one comment ends and another begins on the same line.)

This form of the getline command sets NF, NR, FNR, RT, and the value of $0.

NOTE: The new value of $0 is used to test the patterns of any subsequent rules. The original value of $0 that triggered the rule that executed getline is lost. By contrast, the next statement reads a new record but immediately begins processing it normally, starting with the first rule in the program. See Section 7.4.8 [The next Statement], page 144.

4.9.2 Using getline into a Variable

You can use ‘getline var’ to read the next record from awk’s input into the variable var. No other processing is done. For example, suppose the next line is a comment or a special string, and you want to read it without triggering any rules. This form of getline allows you to read that line and store it in a variable so that the main read-a-line-and-check-each-rule loop of awk never sees it. The following example swaps every two lines of input:

     {
          if ((getline tmp) > 0) {
               print tmp
               print $0
          } else
               print $0
     }

It takes the following list:

     wan
     tew
     free
     phore

and produces these results:

     tew
     wan
     phore
     free

The getline command used in this way sets only the variables NR, FNR, and RT (and, of course, var). The record is not split into fields, so the values of the fields (including $0) and the value of NF do not change.

4.9.3 Using getline from a File

Use ‘getline < file’ to read the next record from file. Here, file is a string-valued expression that specifies the file name. ‘< file’ is called a redirection because it directs input to come from a different place. For example, the following program reads its input record from the file secondary.input when it encounters a first field with a value equal to 10 in the current input file:

     {
         if ($1 == 10) {
              getline < "secondary.input"
              print
         } else
              print
     }

Because the main input stream is not used, the values of NR and FNR are not changed. However, the record it reads is split into fields in the normal manner, so the values of $0 and the other fields are changed, resulting in a new value of NF. RT is also set.

According to POSIX, ‘getline < expression’ is ambiguous if expression contains unparenthesized operators other than ‘$’; for example, ‘getline < dir "/" file’ is ambiguous because the concatenation operator (not discussed yet; see Section 6.2.2 [String Concatenation], page 115) is not parenthesized. You should write it as ‘getline < (dir "/" file)’ if you want your program to be portable to all awk implementations.

4.9.4 Using getline into a Variable from a File

Use ‘getline var < file’ to read input from the file file, and put it in the variable var. As earlier, file is a string-valued expression that specifies the file from which to read.

In this version of getline, none of the predefined variables are changed and the record is not split into fields. The only variable changed is var.^{8} For example, the following program copies all the input files to the output, except for records that say ‘@include filename’. Such a record is replaced by the contents of the file filename:

8 This is not quite true. RT could be changed if RS is a regular expression.

     {
          if (NF == 2 && $1 == "@include") {
               while ((getline line < $2) > 0)
                    print line
               close($2)
          } else
               print
     }

Note here how the name of the extra input file is not built into the program; it is taken directly from the data, specifically from the second field on the @include line.

The close() function is called to ensure that if two identical @include lines appear in the input, the entire specified file is included twice. See Section 5.9 [Closing Input and Output Redirections], page 102.

One deficiency of this program is that it does not process nested @include statements (i.e., @include statements in included files) the way a true macro preprocessor would. See Section 11.3.9 [An Easy Way to Use Library Functions], page 285, for a program that does handle nested @include statements.

4.9.5 Using getline from a Pipe

Omniscience has much to recommend it. Failing that, attention to details would be useful.

—Brian Kernighan The output of a command can also be piped into getline, using ‘command | getline’.

In this case, the string command is run as a shell command and its output is piped into awk to be used as input. This form of getline reads one record at a time from the pipe. For example, the following program copies its input to its output, except for lines that begin with ‘@execute’, which are replaced by the output produced by running the rest of the line as a shell command:

     {
          if ($1 == "@execute") {
               tmp = substr($0, 10)        # Remove "@execute"
               while ((tmp | getline) > 0)
                    print
               close(tmp)
          } else
               print
     }

The close() function is called to ensure that if two identical ‘@execute’ lines appear in the input, the command is run for each one. Given the input:

     foo
     bar
     baz
     @execute who
     bletch

the program might produce:

     foo
     bar
     baz
     arnold     ttyv0   Jul 13 14:22
     miriam     ttyp0   Jul 13 14:23     (murphy:0)
     bill       ttyp1   Jul 13 14:23     (murphy:0)
     bletch

Notice that this program ran the command who and printed the result. (If you try this program yourself, you will of course get different results, depending upon who is logged in on your system.)

This variation of getline splits the record into fields, sets the value of NF, and recomputes the value of $0. The values of NR and FNR are not changed. RT is set.

According to POSIX, ‘expression | getline’ is ambiguous if expression contains unparenthesized operators other than ‘$’—for example, ‘"echo " "date" | getline’ is ambiguous because the concatenation operator is not parenthesized. You should write it as ‘("echo " "date") | getline’ if you want your program to be portable to all awk implementations.

NOTE: Unfortunately, gawk has not been consistent in its treatment of a construct like ‘"echo " "date" | getline’. Most versions, including the current version, treat it at as ‘("echo " "date") | getline’. (This is also how BWK awk behaves.) Some versions instead treat it as ‘"echo " ("date" | getline)’. (This is how mawk behaves.) In short, always use explicit parentheses, and then you won’t have to worry.

4.9.6 Using getline into a Variable from a Pipe

When you use ‘command | getline var’, the output of command is sent through a pipe to getline and into the variable var. For example, the following program reads the current date and time into the variable current_time, using the date utility, and then prints it:

     BEGIN {
          "date" | getline current_time
          close("date")
          print "Report printed on " current_time
     }

In this version of getline, none of the predefined variables are changed and the record is not split into fields. However, RT is set.

4.9.7 Using getline from a Coprocess

Reading input into getline from a pipe is a one-way operation. The command that is started with ‘command | getline’ only sends data to your awk program.

On occasion, you might want to send data to another program for processing and then read the results back. gawk allows you to start a coprocess, with which two-way communications are possible. This is done with the ‘|&’ operator. Typically, you write data to the coprocess first and then read the results back, as shown in the following:

     print "some query" |& "db_server"
     "db_server" |& getline

which sends a query to db_server and then reads the results.

The values of NR and FNR are not changed, because the main input stream is not used. However, the record is split into fields in the normal manner, thus changing the values of $0, of the other fields, and of NF and RT.

Coprocesses are an advanced feature. They are discussed here only because this is the section on getline. See Section 12.3 [Two-Way Communications with Another Process], page 305, where coprocesses are discussed in more detail.

4.9.8 Using getline into a Variable from a Coprocess

When you use ‘command |& getline var’, the output from the coprocess command is sent through a two-way pipe to getline and into the variable var.

In this version of getline, none of the predefined variables are changed and the record is not split into fields. The only variable changed is var. However, RT is set.

4.9.9 Points to Remember About getline

Here are some miscellaneous points about getline that you should bear in mind:

• When getline changes the value of $0 and NF, awk does not automatically jump to the start of the program and start testing the new record against every pattern. However, the new record is tested against any subsequent rules.

• Some very old awk implementations limit the number of pipelines that an awk program may have open to just one. In gawk, there is no such limit. You can open as many pipelines (and coprocesses) as the underlying operating system permits.

• An interesting side effect occurs if you use getline without a redirection inside a BEGIN rule. Because an unredirected getline reads from the command-line data files, the first getline command causes awk to set the value of FILENAME. Normally, FILENAME does not have a value inside BEGIN rules, because you have not yet started to process the command-line data files. (See Section 7.1.4 [The BEGIN and END Special Patterns], page 134; also see Section 7.5.2 [Built-in Variables That Convey Information], page 149.)

• Using FILENAME with getline (‘getline < FILENAME’) is likely to be a source of confusion. awk opens a separate input stream from the current input file. However, by not using a variable, $0 and NF are still updated. If you’re doing this, it’s probably by accident, and you should reconsider what it is you’re trying to accomplish.

• Section 4.9.10 [Summary of getline Variants], page 84, presents a table summarizing the getline variants and which variables they can affect. It is worth noting that those variants that do not use redirection can cause FILENAME to be updated if they cause awk to start reading a new input file.

• If the variable being assigned is an expression with side effects, different versions of awk behave differently upon encountering end-of-file. Some versions don’t evaluate the expression; many versions (including gawk) do. Here is an example, courtesy of Duncan Moore:

     BEGIN {
         system("echo 1 > f")
         while ((getline a[++c] < "f") > 0) { }

             print c
         }

Here, the side effect is the ‘++c’. Is c incremented if end-of-file is encountered before the element in a is assigned? gawk treats getline like a function call, and evaluates the expression ‘a[++c]’ before attempting to read from f. However, some versions of awk only evaluate the expression once they know that there is a string value to be assigned.

4.9.10 Summary of getline Variants

Table 4.1 summarizes the eight variants of getline, listing which predefined variables are set by each one, and whether the variant is standard or a gawk extension. Note: for each variant, gawk sets the RT predefined variable.

Variant  Effect  awk / gawk

getline  Sets $0, NF, FNR, NR, and RT     awk
getline var  Sets var, FNR, NR, and RT        awk
getline < file  Sets $0, NF, and RT             awk
getline var < file  Sets var and RT                awk
command | getline  Sets $0, NF, and RT             awk

command | getline var  Sets var and RT                awk

command |& getline  Sets $0, NF, and RT             gawk

command |& getline var  Sets var and RT                gawk

Table 4.1: getline variants and what they set

4.10 Reading Input with a Timeout

This section describes a feature that is specific to gawk.

You may specify a timeout in milliseconds for reading input from the keyboard, a pipe, or two-way communication, including TCP/IP sockets. This can be done on a per-input, per-command, or per-connection basis, by setting a special element in the PROCINFO array (see Section 7.5.2 [Built-in Variables That Convey Information], page 149):

     PROCINFO["input_name", "READ_TIMEOUT"] = timeout in milliseconds

When set, this causes gawk to time out and return failure if no data is available to read within the specified timeout period. For example, a TCP client can decide to give up on receiving any response from the server after a certain amount of time:

     Service = "/inet/tcp/0/localhost/daytime"
     PROCINFO[Service, "READ_TIMEOUT"] = 100
     if ((Service |& getline) > 0)
         print $0
     else if (ERRNO != "")
         print ERRNO

Here is how to read interactively from the user^{9} without waiting for more than five seconds:

9 This assumes that standard input is the keyboard.

     PROCINFO["/dev/stdin", "READ_TIMEOUT"] = 5000
     while ((getline < "/dev/stdin") > 0)
         print $0

gawk terminates the read operation if input does not arrive after waiting for the timeout period, returns failure, and sets ERRNO to an appropriate string value. A negative or zero value for the timeout is the same as specifying no timeout at all.

A timeout can also be set for reading from the keyboard in the implicit loop that reads input records and matches them against patterns, like so:

     $ gawk 'BEGIN { PROCINFO["-", "READ_TIMEOUT"] = 5000 }
     > { print "You entered: " $0 }'
     gawk
     ⊣ You entered: gawk

In this case, failure to respond within five seconds results in the following error message:

      error gawk: cmd. line:2: (FILENAME=- FNR=1) fatal: error reading input file `-
     ': Connection timed out

The timeout can be set or changed at any time, and will take effect on the next attempt to read from the input device. In the following example, we start with a timeout value of one second, and progressively reduce it by one-tenth of a second until we wait indefinitely for the input to arrive:

     PROCINFO[Service, "READ_TIMEOUT"] = 1000
     while ((Service |& getline) > 0) {
         print $0
         PROCINFO[Service, "READ_TIMEOUT"] -= 100
     }

NOTE: You should not assume that the read operation will block exactly after the tenth record has been printed. It is possible that gawk will read and buffer more than one record’s worth of data the first time. Because of this, changing the value of timeout like in the preceding example is not very useful.

If the PROCINFO element is not present and the GAWK_READ_TIMEOUT environment variable exists, gawk uses its value to initialize the timeout value. The exclusive use of the environment variable to specify timeout has the disadvantage of not being able to control it on a per-command or per-connection basis.

gawk considers a timeout event to be an error even though the attempt to read from the underlying device may succeed in a later attempt. This is a limitation, and it also means that you cannot use this to multiplex input from two or more sources.

Assigning a timeout value prevents read operations from blocking indefinitely. But bear in mind that there are other ways gawk can stall waiting for an input device to be ready. A network client can sometimes take a long time to establish a connection before it can start reading any data, or the attempt to open a FIFO special file for reading can block indefinitely until some other process opens it for writing.

4.11 Directories on the Command Line

According to the POSIX standard, files named on the awk command line must be text files; it is a fatal error if they are not. Most versions of awk treat a directory on the command line as a fatal error.

By default, gawk produces a warning for a directory on the command line, but otherwise ignores it. This makes it easier to use shell wildcards with your awk program:

$ gawk -f whizprog.awk * Directories could kill this program If either of the --posix or --traditional options is given, then gawk reverts to treating

a directory on the command line as a fatal error.

See Section 16.7.6 [Reading Directories], page 404, for a way to treat directories as usable data from an awk program.

4.12 Summary

• Input is split into records based on the value of RS. The possibilities are as follows:

Value of RS  Records are split on …  awk / gawk
Any single character  That character  awk
The empty string ("") Runs of two or more newlines awk
A regexp  Text that matches the regexp gawk

• FNR indicates how many records have been read from the current input file; NR indicates how many records have been read in total.

• gawk sets RT to the text matched by RS.

• After splitting the input into records, awk further splits the records into individual fields, named $1, $2, and so on. $0 is the whole record, and NF indicates how many fields there are. The default way to split fields is between whitespace characters.

• Fields may be referenced using a variable, as in $NF. Fields may also be assigned values, which causes the value of $0 to be recomputed when it is later referenced. Assigning to a field with a number greater than NF creates the field and rebuilds the record, using OFS to separate the fields. Incrementing NF does the same thing. Decrementing NF throws away fields and rebuilds the record.

• Field splitting is more complicated than record splitting:

Field separator value  Fields are split …  awk / gawk
FS == " "  On runs of whitespace  awk

FS == any single character  On that character  awk

FS == regexp  On text matching the regexp  awk

FS == ""

Such that each individual character is a separate field

                                                              gawk

FIELDWIDTHS == list of columns Based on character position  gawk

FPAT == regexp

On the text surrounding text matching the regexp

                                                              gawk

• Using ‘FS = "\n"’ causes the entire record to be a single field (assuming that newlines separate records).

• FS may be set from the command line using the -F option. This can also be done using command-line variable assignment.

• Use PROCINFO["FS"] to see how fields are being split.

• Use getline in its various forms to read additional records from the default input stream, from a file, or from a pipe or coprocess.

• Use PROCINFO[file, "READ_TIMEOUT"] to cause reads to time out for file.

• Directories on the command line are fatal for standard awk; gawk ignores them if not in POSIX mode.

4.13 Exercises

1. Using the FIELDWIDTHS variable (see Section 4.6 [Reading Fixed-Width Data], page 72), write a program to read election data, where each record represents one voter’s votes. Come up with a way to define which columns are associated with each ballot item, and print the total votes, including abstentions, for each item.

2. Section 4.9.1 [Using getline with No Arguments], page 78, presented a program to remove C-style comments (‘/* ... */’) from the input. That program does not work if one comment ends on one line and another one starts later on the same line. That can be fixed by making one simple change. What is it?

Chapter 5: Printing Output 89

5 Printing Output

One of the most common programming actions is to print, or output, some or all of the input. Use the print statement for simple output, and the printf statement for fancier formatting. The print statement is not limited when computing which values to print. However, with two exceptions, you cannot specify how to print them—how many columns, whether to use exponential notation or not, and so on. (For the exceptions, see Section 5.3 [Output Separators], page 91, and Section 5.4 [Controlling Numeric Output with print], page 92.) For printing with specifications, you need the printf statement (see Section 5.5 [Using printf Statements for Fancier Printing], page 92).

Besides basic and formatted printing, this chapter also covers I/O redirections to files and pipes, introduces the special file names that gawk processes internally, and discusses the close() built-in function.

5.1 The print Statement

Use the print statement to produce output with simple, standardized formatting. You specify only the strings or numbers to print, in a list separated by commas. They are output, separated by single spaces, followed by a newline. The statement looks like this:

     print item1, item2, ...

The entire list of items may be optionally enclosed in parentheses. The parentheses are necessary if any of the item expressions uses the ‘>’ relational operator; otherwise it could be confused with an output redirection (see Section 5.6 [Redirecting Output of print and printf], page 98).

The items to print can be constant strings or numbers, fields of the current record (such as $1), variables, or any awk expression. Numeric values are converted to strings and then printed.

The simple statement ‘print’ with no items is equivalent to ‘print $0’: it prints the entire current record. To print a blank line, use ‘print ""’. To print a fixed piece of text, use a string constant, such as "Don't Panic", as one item. If you forget to use the doublequote characters, your text is taken as an awk expression, and you will probably get an error. Keep in mind that a space is printed between any two items.

Note that the print statement is a statement and not an expression—you can’t use it in the pattern part of a pattern–action statement, for example.

5.2 print Statement Examples

Each print statement makes at least one line of output. However, it isn’t limited to only one line. If an item value is a string containing a newline, the newline is output along with the rest of the string. A single print statement can make any number of lines this way.

The following is an example of printing a string that contains embedded newlines:

     $ awk 'BEGIN { print "line one\nline two\nline three" }'
     ⊣ line one
     ⊣ line two
     ⊣ line three

The next example, which is run on the inventory-shipped file, prints the first two fields of each input record, with a space between them:

     $ awk '{ print $1, $2 }' inventory-shipped
     ⊣ Jan 13
     ⊣ Feb 15
     ⊣ Mar 15
     ...

A common mistake in using the print statement is to omit the comma between two items. This often has the effect of making the items run together in the output, with no space. The reason for this is that juxtaposing two string expressions in awk means to concatenate them. Here is the same program, without the comma:

     $ awk '{ print $1 $2 }' inventory-shipped
     ⊣ Jan13
     ⊣ Feb15
     ⊣ Mar15
     ...

To someone unfamiliar with the inventory-shipped file, neither example’s output makes much sense. A heading line at the beginning would make it clearer. Let’s add some headings to our table of months ($1) and green crates shipped ($2). We do this using a BEGIN rule (see Section 7.1.4 [The BEGIN and END Special Patterns], page 134) so that the headings are only printed once:

     awk 'BEGIN {  print "Month Crates"
                   print "----- ------" }
                {  print $1, $2 }' inventory-shipped

When run, the program prints the following:

     Month Crates
     ----- ------
     Jan 13
     Feb 15
     Mar 15
     ...

The only problem, however, is that the headings and the table data don’t line up! We can fix this by printing some spaces between the two fields:

     awk 'BEGIN { print "Month Crates"
                  print "----- ------" }
                { print $1, "     ", $2 }' inventory-shipped

Lining up columns this way can get pretty complicated when there are many columns to fix. Counting spaces for two or three columns is simple, but any more than this can take up a lot of time. This is why the printf statement was created (see Section 5.5 [Using printf Statements for Fancier Printing], page 92); one of its specialties is lining up columns of data.

NOTE: You can continue either a print or printf statement simply by putting a newline after any comma (see Section 1.6 [awk Statements Versus Lines], page 27).

5.3 Output Separators

As mentioned previously, a print statement contains a list of items separated by commas. In the output, the items are normally separated by single spaces. However, this doesn’t need to be the case; a single space is simply the default. Any string of characters may be used as the output field separator by setting the predefined variable OFS. The initial value of this variable is the string " " (i.e., a single space).

The output from an entire print statement is called an output record. Each print statement outputs one output record, and then outputs a string called the output record separator (or ORS). The initial value of ORS is the string "\n" (i.e., a newline character). Thus, each print statement normally makes a separate line.

In order to change how output fields and records are separated, assign new values to the variables OFS and ORS. The usual place to do this is in the BEGIN rule (see Section 7.1.4 [The BEGIN and END Special Patterns], page 134), so that it happens before any input is processed. It can also be done with assignments on the command line, before the names of the input files, or using the -v command-line option (see Section 2.2 [Command-Line Options], page 31). The following example prints the first and second fields of each input record, separated by a semicolon, with a blank line added after each newline:

$ awk 'BEGIN { OFS = ";"; ORS = "\n\n" }
>            { print $1, $2 }' mail-list
⊣ Amelia;555-5553

⊣

⊣ Anthony;555-3412

⊣

⊣ Becky;555-7685

⊣

⊣ Bill;555-1675

⊣

⊣ Broderick;555-0542

⊣

⊣ Camilla;555-2912

⊣

⊣ Fabius;555-1234

⊣

⊣ Julie;555-6699

⊣

⊣ Martin;555-6480

⊣

⊣ Samuel;555-3430

⊣

⊣ Jean-Paul;555-2127

⊣

If the value of ORS does not contain a newline, the program’s output runs together on a single line.

5.4 Controlling Numeric Output with print

When printing numeric values with the print statement, awk internally converts each number to a string of characters and prints that string. awk uses the sprintf() function to do this conversion (see Section 9.1.3 [String-Manipulation Functions], page 177). For now, it suffices to say that the sprintf() function accepts a format specification that tells it how to format numbers (or strings), and that there are a number of different ways in which numbers can be formatted. The different format specifications are discussed more fully in Section 5.5.2 [Format-Control Letters], page 93.

The predefined variable OFMT contains the format specification that print uses with sprintf() when it wants to convert a number to a string for printing. The default value of OFMT is "%.6g". The way print prints numbers can be changed by supplying a different format specification for the value of OFMT, as shown in the following example:

     $ awk 'BEGIN {
     >   OFMT = "%.0f"  # print numbers as integers (rounds)
     >   print 17.23, 17.54 }'
     ⊣ 17 18

According to the POSIX standard, awk’s behavior is undefined if OFMT contains anything but a floating-point conversion specification.

5.5 Using printf Statements for Fancier Printing

For more precise control over the output format than what is provided by print, use printf. With printf you can specify the width to use for each item, as well as various formatting choices for numbers (such as what output base to use, whether to print an exponent, whether to print a sign, and how many digits to print after the decimal point).

5.5.1 Introduction to the printf Statement

A simple printf statement looks like this:

     printf format, item1, item2, ...

As for print, the entire list of arguments may optionally be enclosed in parentheses. Here too, the parentheses are necessary if any of the item expressions uses the ‘>’ relational operator; otherwise, it can be confused with an output redirection (see Section 5.6 [Redirecting Output of print and printf], page 98).

The difference between printf and print is the format argument. This is an expression whose value is taken as a string; it specifies how to output each of the other arguments. It is called the format string.

The format string is very similar to that in the ISO C library function printf(). Most of format is text to output verbatim. Scattered among this text are format specifiers—one per item. Each format specifier says to output the next item in the argument list at that place in the format.

The printf statement does not automatically append a newline to its output. It outputs only what the format string specifies. So if a newline is needed, you must include one in the format string. The output separator variables OFS and ORS have no effect on printf statements. For example:

$ awk 'BEGIN {
>    ORS = "\nOUCH!\n"; OFS = "+"
>    msg = "Don\47t Panic!"
>    printf "%s\n", msg
> }'
⊣ Don't Panic!

Here, neither the ‘+’ nor the ‘OUCH!’ appears in the output message.

5.5.2 Format-Control Letters

A format specifier starts with the character ‘%’ and ends with a format-control letter—it tells the printf statement how to output one item. The format-control letter specifies what kind of value to print. The rest of the format specifier is made up of optional modifiers that control how to print the value, such as the field width. Here is a list of the format-control letters:

%c Print a number as a character; thus, ‘printf "%c", 65’ outputs the letter ‘A’.

The output for a string value is the first character of the string.

NOTE: The POSIX standard says the first character of a string is printed. In locales with multibyte characters, gawk attempts to convert the leading bytes of the string into a valid wide character and then to print the multibyte encoding of that character. Similarly, when printing a numeric value, gawk allows the value to be within the numeric range of values that can be held in a wide character. If the conversion to multibyte encoding fails, gawk uses the low eight bits of the value as the character to print. Other awk versions generally restrict themselves to printing the first byte of a string or to numeric values within the range of a single byte (0–255).

%d, %i Print a decimal integer. The two control letters are equivalent. (The ‘%i’ specification is for compatibility with ISO C.)

%e, %E Print a number in scientific (exponential) notation. For example:

printf "%4.3e\n", 1950

prints ‘1.950e+03’, with a total of four significant figures, three of which follow the decimal point. (The ‘4.3’ represents two modifiers, discussed in the next subsection.) ‘%E’ uses ‘E’ instead of ‘e’ in the output.

%f Print a number in floating-point notation. For example:

printf "%4.3f", 1950

prints ‘1950.000’, with a total of four significant figures, three of which follow the decimal point. (The ‘4.3’ represents two modifiers, discussed in the next subsection.) On systems supporting IEEE 754 floating-point format, values representing negative infinity are formatted as ‘-inf’ or ‘-infinity’, and positive infinity as ‘inf’ or ‘infinity’. The special “not a number” value formats as ‘-nan’ or ‘nan’ (see Section 15.2 [Other Stuff to Know], page 342).

%F Like ‘%f’, but the infinity and “not a number” values are spelled using uppercase letters. The ‘%F’ format is a POSIX extension to ISO C; not all systems support it. On those that don’t, gawk uses ‘%f’ instead.

%g, %G Print a number in either scientific notation or in floating-point notation, whichever uses fewer characters; if the result is printed in scientific notation, ‘%G’ uses ‘E’ instead of ‘e’.

%o Print an unsigned octal integer (see Section 6.1.1.2 [Octal and Hexadecimal Numbers], page 107).

%s Print a string.

%u Print an unsigned decimal integer. (This format is of marginal use, because all numbers in awk are floating point; it is provided primarily for compatibility with C.)

%x, %X Print an unsigned hexadecimal integer; ‘%X’ uses the letters ‘A’ through ‘F’ instead of ‘a’ through ‘f’ (see Section 6.1.1.2 [Octal and Hexadecimal Numbers], page 107).

%% Print a single ‘%’. This does not consume an argument and it ignores any modifiers.

NOTE: When using the integer format-control letters for values that are outside the range of the widest C integer type, gawk switches to the ‘%g’ format specifier. If --lint is provided on the command line (see Section 2.2 [Command-Line Options], page 31), gawk warns about this. Other versions of awk may print invalid values or do something else entirely.

5.5.3 Modifiers for printf Formats

A format specification can also include modifiers that can control how much of the item’s value is printed, as well as how much space it gets. The modifiers come between the ‘%’ and the format-control letter. We use the bullet symbol “•” in the following examples to represent spaces in the output. Here are the possible modifiers, in the order in which they may appear:

N$ An integer constant followed by a ‘$’ is a positional specifier. Normally, format specifications are applied to arguments in the order given in the format string. With a positional specifier, the format specification is applied to a specific argument, instead of what would be the next argument in the list. Positional specifiers begin counting with one. Thus:

     printf "%s %s\n", "don't", "panic"
     printf "%2$s %1$s\n", "panic", "don't"

prints the famous friendly message twice. At first glance, this feature doesn’t seem to be of much use. It is in fact a gawk extension, intended for use in translating messages at runtime. See Section 13.4.2 [Rearranging printf Arguments], page 319, which describes how and why to use positional specifiers. For now, we ignore them.

- (Minus) The minus sign, used before the width modifier (see later on in this list), says to left-justify the argument within its specified width. Normally, the argument is printed right-justified in the specified width. Thus:

     printf "%-4s", "foo"

prints ‘foo•’.

space For numeric conversions, prefix positive values with a space and negative values with a minus sign.

+ The plus sign, used before the width modifier (see later on in this list), says to always supply a sign for numeric conversions, even if the data to format is positive. The ‘+’ overrides the space modifier.

# Use an “alternative form” for certain control letters. For ‘%o’, supply a leading zero. For ‘%x’ and ‘%X’, supply a leading ‘0x’ or ‘0X’ for a nonzero result. For ‘%e’, ‘%E’, ‘%f’, and ‘%F’, the result always contains a decimal point. For ‘%g’ and ‘%G’, trailing zeros are not removed from the result.

0 A leading ‘0’ (zero) acts as a flag indicating that output should be padded with zeros instead of spaces. This applies only to the numeric output formats. This flag only has an effect when the field width is wider than the value to print.

' A single quote or apostrophe character is a POSIX extension to ISO C. It indicates that the integer part of a floating-point value, or the entire part of an integer decimal value, should have a thousands-separator character in it. This only works in locales that support such characters. For example:

     $ cat thousands.awk  Show source program
     ⊣ BEGIN { printf "%'d\n", 1234567 }
     $ LC_ALL=C gawk -f thousands.awk

⊣ 1234567 Results in "C" locale

     $ LC_ALL=en_US.UTF-8 gawk -f thousands.awk

⊣ 1,234,567 Results in US English UTF locale For more information about locales and internationalization issues, see Section 6.6 [Where You Are Makes a Difference], page 129.

NOTE: The ‘'’ flag is a nice feature, but its use complicates things: it becomes difficult to use it in command-line programs. For information on appropriate quoting tricks, see Section 1.1.6 [Shell Quoting Issues], page 21.

width This is a number specifying the desired minimum width of a field. Inserting any number between the ‘%’ sign and the format-control character forces the field to expand to this width. The default way to do this is to pad with spaces on the left. For example:

     printf "%4s", "foo"

prints ‘•foo’. The value of width is a minimum width, not a maximum. If the item value requires more than width characters, it can be as wide as necessary. Thus, the following:

               printf "%4s", "foobar"

prints ‘foobar’.

Preceding the width with a minus sign causes the output to be padded with spaces on the right, instead of on the left.

.prec A period followed by an integer constant specifies the precision to use when printing. The meaning of the precision varies by control letter:

          %d, %i, %o, %u, %x, %X

Minimum number of digits to print.

          %e, %E, %f, %F

Number of digits to the right of the decimal point.

%g, %G Maximum number of significant digits.

%s Maximum number of characters from the string that should print.

Thus, the following:

               printf "%.4s", "foobar"

prints ‘foob’.

The C library printf’s dynamic width and prec capability (e.g., "%*.*s") is supported. Instead of supplying explicit width and/or prec values in the format string, they are passed in the argument list. For example:

     w = 5
     p = 3
     s = "abcdefg"
     printf "%*.*s\n", w, p, s

is exactly equivalent to:

     s = "abcdefg"
     printf "%5.3s\n", s

Both programs output ‘••abc’. Earlier versions of awk did not support this capability. If you must use such a version, you may simulate this feature by using concatenation to build up the format string, like so:

     w = 5
     p = 3
     s = "abcdefg"
     printf "%" w "." p "s\n", s

This is not particularly easy to read, but it does work.

C programmers may be used to supplying additional modifiers (‘h’, ‘j’, ‘l’, ‘L’, ‘t’, and ‘z’) in printf format strings. These are not valid in awk. Most awk implementations silently ignore them. If --lint is provided on the command line (see Section 2.2 [Command-Line Options], page 31), gawk warns about their use. If --posix is supplied, their use is a fatal error.

5.5.4 Examples Using printf

The following simple example shows how to use printf to make an aligned table:

     awk '{ printf "%-10s %s\n", $1, $2 }' mail-list

This command prints the names of the people ($1) in the file mail-list as a string of 10 characters that are left-justified. It also prints the phone numbers ($2) next on the line. This produces an aligned two-column table of names and phone numbers, as shown here:

     $ awk '{ printf "%-10s %s\n", $1, $2 }' mail-list
     ⊣ Amelia     555-5553
     ⊣ Anthony    555-3412
     ⊣ Becky      555-7685
     ⊣ Bill       555-1675
     ⊣ Broderick  555-0542
     ⊣ Camilla    555-2912
     ⊣ Fabius     555-1234
     ⊣ Julie      555-6699
     ⊣ Martin     555-6480
     ⊣ Samuel     555-3430
     ⊣ Jean-Paul  555-2127

In this case, the phone numbers had to be printed as strings because the numbers are separated by dashes. Printing the phone numbers as numbers would have produced just the first three digits: ‘555’. This would have been pretty confusing.

It wasn’t necessary to specify a width for the phone numbers because they are last on their lines. They don’t need to have spaces after them.

The table could be made to look even nicer by adding headings to the tops of the columns. This is done using a BEGIN rule (see Section 7.1.4 [The BEGIN and END Special Patterns], page 134) so that the headers are only printed once, at the beginning of the awk program:

     awk 'BEGIN { print "Name      Number"
                  print "----      ------" }
                { printf "%-10s %s\n", $1, $2 }' mail-list

The preceding example mixes print and printf statements in the same program. Using just printf statements can produce the same results:

     awk 'BEGIN { printf "%-10s %s\n", "Name", "Number"
                  printf "%-10s %s\n", "----", "------" }
                { printf "%-10s %s\n", $1, $2 }' mail-list

Printing each column heading with the same format specification used for the column elements ensures that the headings are aligned just like the columns.

The fact that the same format specification is used three times can be emphasized by storing it in a variable, like this:

     awk 'BEGIN { format = "%-10s %s\n"
                  printf format, "Name", "Number"
                  printf format, "----", "------" }
                { printf format, $1, $2 }' mail-list

5.6 Redirecting Output of print and printf

So far, the output from print and printf has gone to the standard output, usually the screen. Both print and printf can also send their output to other places. This is called redirection.

NOTE: When --sandbox is specified (see Section 2.2 [Command-Line Options], page 31), redirecting output to files, pipes, and coprocesses is disabled.

A redirection appears after the print or printf statement. Redirections in awk are written just like redirections in shell commands, except that they are written inside the awk program.

There are four forms of output redirection: output to a file, output appended to a file, output through a pipe to another command, and output to a coprocess. We show them all for the print statement, but they work identically for printf:

print items > output-file

This redirection prints the items into the output file named output-file. The file name output-file can be any expression. Its value is changed to a string and then used as a file name (see Chapter 6 [Expressions], page 107). When this type of redirection is used, the output-file is erased before the first output is written to it. Subsequent writes to the same output-file do not erase output-file, but append to it. (This is different from how you use redirections in shell scripts.) If output-file does not exist, it is created. For example, here is how an awk program can write a list of peoples’ names to one file named name-list, and a list of phone numbers to another file named phone-list:

     $ awk '{ print $2 > "phone-list"
     >        print $1 > "name-list" }' mail-list
     $ cat phone-list
     ⊣ 555-5553
     ⊣ 555-3412
     ...
     $ cat name-list
     ⊣ Amelia
     ⊣ Anthony
     ...

Each output file contains one name or number per line.

print items >> output-file

This redirection prints the items into the preexisting output file named outputfile. The difference between this and the single-‘>’ redirection is that the old contents (if any) of output-file are not erased. Instead, the awk output is appended to the file. If output-file does not exist, then it is created.

print items | command

It is possible to send output to another program through a pipe instead of into a file. This redirection opens a pipe to command, and writes the values of items through this pipe to another process created to execute command. The redirection argument command is actually an awk expression. Its value is converted to a string whose contents give the shell command to be run. For example, the following produces two files, one unsorted list of peoples’ names, and one list sorted in reverse alphabetical order:

     awk '{ print $1 > "names.unsorted"
            command = "sort -r > names.sorted"
            print $1 | command }' mail-list

The unsorted list is written with an ordinary redirection, while the sorted list is written by piping through the sort utility. The next example uses redirection to mail a message to the mailing list bug-system. This might be useful when trouble is encountered in an awk script run periodically for system maintenance:

     report = "mail bug-system"
     print("Awk script failed:", $0) | report
     print("at record number", FNR, "of", FILENAME) | report
     close(report)

The close() function is called here because it’s a good idea to close the pipe as soon as all the intended output has been sent to it. See Section 5.9 [Closing Input and Output Redirections], page 102, for more information. This example also illustrates the use of a variable to represent a file or command—it is not necessary to always use a string constant. Using a variable is generally a good idea, because (if you mean to refer to that same file or command) awk requires that the string value be written identically every time.

print items |& command

This redirection prints the items to the input of command. The difference between this and the single-‘|’ redirection is that the output from command can be read with getline. Thus, command is a coprocess, which works together with but is subsidiary to the awk program. This feature is a gawk extension, and is not available in POSIX awk. See Section 4.9.7 [Using getline from a Coprocess], page 82, for a brief discussion. See Section 12.3 [Two-Way Communications with Another Process], page 305, for a more complete discussion.

Redirecting output using ‘>’, ‘>>’, ‘|’, or ‘|&’ asks the system to open a file, pipe, or coprocess only if the particular file or command you specify has not already been written to by your program or if it has been closed since it was last written to.

It is a common error to use ‘>’ redirection for the first print to a file, and then to use ‘>>’ for subsequent output:

# clear the file
print "Don't panic" > "guide.txt"
...
# append
print "Avoid improbability generators" >> "guide.txt"

This is indeed how redirections must be used from the shell. But in awk, it isn’t necessary. In this kind of case, a program should use ‘>’ for all the print statements, because the output file is only opened once. (It happens that if you mix ‘>’ and ‘>>’ output is produced in the expected order. However, mixing the operators for the same file is definitely poor style, and is confusing to readers of your program.)

As mentioned earlier (see Section 4.9.9 [Points to Remember About getline], page 83), many older awk implementations limit the number of pipelines that an awk program may have open to just one! In gawk, there is no such limit. gawk allows a program to open as many pipelines as the underlying operating system permits.

Piping into sh A particularly powerful way to use redirection is to build command lines and pipe them

into the shell, sh. For example, suppose you have a list of files brought over from a system where all the file names are stored in uppercase, and you wish to rename them to have names in all lowercase. The following program is both simple and efficient:

      { printf("mv %s %s\n", $0, tolower($0)) | "sh" }
      END { close("sh") }

The tolower() function returns its argument string with all uppercase characters converted to lowercase (see Section 9.1.3 [String-Manipulation Functions], page 177). The program builds up a list of command lines, using the mv utility to rename the files. It then sends the list to the shell for execution.

See Section 10.2.9 [Quoting Strings to Pass to the Shell], page 228, for a function that can help in generating command lines to be fed to the shell.

5.7 Special Files for Standard Preopened Data Streams

Running programs conventionally have three input and output streams already available to them for reading and writing. These are known as the standard input, standard output, and standard error output. These open streams (and any other open files or pipes) are often referred to by the technical term file descriptors.

These streams are, by default, connected to your keyboard and screen, but they are often redirected with the shell, via the ‘<’, ‘<<’, ‘>’, ‘>>’, ‘>&’, and ‘|’ operators. Standard error is typically used for writing error messages; the reason there are two separate streams, standard output and standard error, is so that they can be redirected separately.

In traditional implementations of awk, the only way to write an error message to standard error in an awk program is as follows:

     print "Serious error detected!" | "cat 1>&2"

This works by opening a pipeline to a shell command that can access the standard error stream that it inherits from the awk process. This is far from elegant, and it also requires a separate process. So people writing awk programs often don’t do this. Instead, they send the error messages to the screen, like this:

     print "Serious error detected!" > "/dev/tty"

(/dev/tty is a special file supplied by the operating system that is connected to your keyboard and screen. It represents the “terminal,”^{1} which on modern systems is a keyboard and screen, not a serial console.) This generally has the same effect, but not always: although the standard error stream is usually the screen, it can be redirected; when that happens, writing to the screen is not correct. In fact, if awk is run from a background job, it may not have a terminal at all. Then opening /dev/tty fails.

1 The “tty” in /dev/tty stands for “Teletype,” a serial terminal.

gawk, BWK awk, and mawk provide special file names for accessing the three standard streams. If the file name matches one of these special names when gawk (or one of the others) redirects input or output, then it directly uses the descriptor that the file name stands for. These special file names work for all operating systems that gawk has been ported to, not just those that are POSIX-compliant:

/dev/stdin

The standard input (file descriptor 0).

/dev/stdout

The standard output (file descriptor 1).

/dev/stderr

The standard error output (file descriptor 2).

With these facilities, the proper way to write an error message then becomes:

     print "Serious error detected!" > "/dev/stderr"

Note the use of quotes around the file name. Like with any other redirection, the value must be a string. It is a common error to omit the quotes, which leads to confusing results.

gawk does not treat these file names as special when in POSIX-compatibility mode. However, because BWK awk supports them, gawk does support them even when invoked with the --traditional option (see Section 2.2 [Command-Line Options], page 31).

5.8 Special File names in gawk

Besides access to standard input, standard output, and standard error, gawk provides access to any open file descriptor. Additionally, there are special file names reserved for TCP/IP networking.

5.8.1 Accessing Other Open Files with gawk

Besides the /dev/stdin, /dev/stdout, and /dev/stderr special file names mentioned earlier, gawk provides syntax for accessing any other inherited open file:

/dev/fd/N

The file associated with file descriptor N. Such a file must be opened by the program initiating the awk execution (typically the shell). Unless special pains are taken in the shell from which gawk is invoked, only descriptors 0, 1, and 2 are available.

The file names /dev/stdin, /dev/stdout, and /dev/stderr are essentially aliases for /dev/fd/0, /dev/fd/1, and /dev/fd/2, respectively. However, those names are more selfexplanatory.

Note that using close() on a file name of the form "/dev/fd/N", for file descriptor numbers above two, does actually close the given file descriptor.

5.8.2 Special Files for Network Communications

gawk programs can open a two-way TCP/IP connection, acting as either a client or a server. This is done using a special file name of the form:

     /net-type/protocol/local-port/remote-host/remote-port

The net-type is one of ‘inet’, ‘inet4’, or ‘inet6’. The protocol is one of ‘tcp’ or ‘udp’, and the other fields represent the other essential pieces of information for making a networking connection. These file names are used with the ‘|&’ operator for communicating with a coprocess (see Section 12.3 [Two-Way Communications with Another Process], page 305). This is an advanced feature, mentioned here only for completeness. Full discussion is delayed until Section 12.4 [Using gawk for Network Programming], page 307.

5.8.3 Special File name Caveats

Here are some things to bear in mind when using the special file names that gawk provides:

• Recognition of the file names for the three standard preopened files is disabled only in POSIX mode.

• Recognition of the other special file names is disabled if gawk is in compatibility mode (either --traditional or --posix; see Section 2.2 [Command-Line Options], page 31).

• gawk always interprets these special file names. For example, using ‘/dev/fd/4’ for output actually writes on file descriptor 4, and not on a new file descriptor that is dup()ed from file descriptor 4. Most of the time this does not matter; however, it is important to not close any of the files related to file descriptors 0, 1, and 2. Doing so results in unpredictable behavior.

5.9 Closing Input and Output Redirections

If the same file name or the same shell command is used with getline more than once during the execution of an awk program (see Section 4.9 [Explicit Input with getline], page 78), the file is opened (or the command is executed) the first time only. At that time, the first record of input is read from that file or command. The next time the same file or command is used with getline, another record is read from it, and so on.

Similarly, when a file or pipe is opened for output, awk remembers the file name or command associated with it, and subsequent writes to the same file or command are appended to the previous writes. The file or pipe stays open until awk exits.

This implies that special steps are necessary in order to read the same file again from the beginning, or to rerun a shell command (rather than reading more output from the same command). The close() function makes these things possible:

     close(filename)

or:

     close(command)

The argument filename or command can be any expression. Its value must exactly match the string that was used to open the file or start the command (spaces and other “irrelevant” characters included). For example, if you open a pipe with this:

     "sort -r names" | getline foo

then you must close it with this:

     close("sort -r names")

Once this function call is executed, the next getline from that file or command, or the next print or printf to that file or command, reopens the file or reruns the command. Because the expression that you use to close a file or pipeline must exactly match the expression used to open the file or run the command, it is good practice to use a variable to store the file name or command. The previous example becomes the following:

     sortcom = "sort -r names"
     sortcom | getline foo
     ...
     close(sortcom)

This helps avoid hard-to-find typographical errors in your awk programs. Here are some of the reasons for closing an output file:

• To write a file and read it back later on in the same awk program. Close the file after writing it, then begin reading it with getline.

• To write numerous files, successively, in the same awk program. If the files aren’t closed, eventually awk may exceed a system limit on the number of open files in one process. It is best to close each one when the program has finished writing it.

• To make a command finish. When output is redirected through a pipe, the command reading the pipe normally continues to try to read input as long as the pipe is open. Often this means the command cannot really do its work until the pipe is closed. For example, if output is redirected to the mail program, the message is not actually sent until the pipe is closed.

• To run the same program a second time, with the same arguments. This is not the same thing as giving more input to the first run! For example, suppose a program pipes output to the mail program. If it outputs several lines redirected to this pipe without closing it, they make a single message of several lines. By contrast, if the program closes the pipe after each line of output, then each line makes a separate message.

If you use more files than the system allows you to have open, gawk attempts to multiplex the available open files among your data files. gawk’s ability to do this depends upon the facilities of your operating system, so it may not always work. It is therefore both good practice and good portability advice to always use close() on your files when you are done with them. In fact, if you are using a lot of pipes, it is essential that you close commands when done. For example, consider something like this:

     {
         ...
         command = ("grep " $1 " /some/file | my_prog -q " $3)
         while ((command | getline) > 0) {
             process output of command
         }
         # need close(command) here
     }

This example creates a new pipeline based on data in each record. Without the call to close() indicated in the comment, awk creates child processes to run the commands, until it eventually runs out of file descriptors for more pipelines.

Even though each command has finished (as indicated by the end-of-file return status from getline), the child process is not terminated;^{2} more importantly, the file descriptor for the pipe is not closed and released until close() is called or awk exits.

close() silently does nothing if given an argument that does not represent a file, pipe, or coprocess that was opened with a redirection. In such a case, it returns a negative value, indicating an error. In addition, gawk sets ERRNO to a string indicating the error.

Note also that ‘close(FILENAME)’ has no “magic” effects on the implicit loop that reads through the files named on the command line. It is, more likely, a close of a file that was never opened with a redirection, so awk silently does nothing, except return a negative value.

When using the ‘|&’ operator to communicate with a coprocess, it is occasionally useful to be able to close one end of the two-way pipe without closing the other. This is done by supplying a second argument to close(). As in any other call to close(), the first argument is the name of the command or special file used to start the coprocess. The second argument should be a string, with either of the values "to" or "from". Case does not matter. As this is an advanced feature, discussion is delayed until Section 12.3 [Two-Way Communications with Another Process], page 305, which describes it in more detail and gives an example.

Using close()’s Return Value In many older versions of Unix awk, the close() function is actually a statement. It is

a syntax error to try and use the return value from close():

      command = "..."
      command | getline info
      retval = close(command)  # syntax error in many Unix awks

gawk treats close() as a function. The return value is −1 if the argument names something that was never opened with a redirection, or if there is a system problem closing the file or process. In these cases, gawk sets the predefined variable ERRNO to a string describing the problem.

In gawk, when closing a pipe or coprocess (input or output), the return value is the exit status of the command.^{3} Otherwise, it is the return value from the system’s close() or fclose() C functions when closing input or output files, respectively. This value is zero if the close succeeds, or −1 if it fails.

The POSIX standard is very vague; it says that close() returns zero on success and a nonzero value otherwise. In general, different implementations vary in what they report when closing pipes; thus, the return value cannot be used portably. In POSIX mode (see Section 2.2 [Command-Line Options], page 31), gawk just returns zero when closing a pipe.

5.10 Summary

• The print statement prints comma-separated expressions. Each expression is separated by the value of OFS and terminated by the value of ORS. OFMT provides the conversion format for numeric values for the print statement.

2 The technical terminology is rather morbid. The finished child is called a “zombie,” and cleaning up

after it is referred to as “reaping.”

3 This is a full 16-bit value as returned by the wait() system call. See the system manual pages for

information on how to decode this value.

• The printf statement provides finer-grained control over output, with format-control letters for different data types and various flags that modify the behavior of the formatcontrol letters.

• Output from both print and printf may be redirected to files, pipes, and coprocesses.

• gawk provides special file names for access to standard input, output, and error, and for network communications.

• Use close() to close open file, pipe, and coprocess redirections. For coprocesses, it is possible to close only one direction of the communications.

5.11 Exercises

1. Rewrite the program:

     awk 'BEGIN { print "Month Crates"
                  print "----- ------" }
                { print $1, "     ", $2 }' inventory-shipped

from Section 5.3 [Output Separators], page 91, by using a new value of OFS.

2. Use the printf statement to line up the headings and table data for the inventory-shipped example that was covered in Section 5.1 [The print Statement], page 89.

3. What happens if you forget the double quotes when redirecting output, as follows:

     BEGIN { print "Serious error detected!" > /dev/stderr }

Chapter 6: Expressions 107

6 Expressions

Expressions are the basic building blocks of awk patterns and actions. An expression evaluates to a value that you can print, test, or pass to a function. Additionally, an expression can assign a new value to a variable or a field by using an assignment operator.

An expression can serve as a pattern or action statement on its own. Most other kinds of statements contain one or more expressions that specify the data on which to operate. As in other languages, expressions in awk can include variables, array references, constants, and function calls, as well as combinations of these with various operators.

6.1 Constants, Variables, and Conversions

Expressions are built up from values and the operations performed upon them. This section describes the elementary objects that provide the values used in expressions.

6.1.1 Constant Expressions

The simplest type of expression is the constant, which always has the same value. There are three types of constants: numeric, string, and regular expression.

Each is used in the appropriate context when you need a data value that isn’t going to change. Numeric constants can have different forms, but are internally stored in an identical manner.

6.1.1.1 Numeric and String Constants

A numeric constant stands for a number. This number can be an integer, a decimal fraction, or a number in scientific (exponential) notation.^{1} Here are some examples of numeric constants that all have the same value:

     105
     1.05e+2
     1050e-1

A string constant consists of a sequence of characters enclosed in double quotation marks. For example:

     "parrot"

represents the string whose contents are ‘parrot’. Strings in gawk can be of any length, and they can contain any of the possible eight-bit ASCII characters, including ASCII nul (character code zero). Other awk implementations may have difficulty with some character codes.

6.1.1.2 Octal and Hexadecimal Numbers

In awk, all numbers are in decimal (i.e., base 10). Many other programming languages allow you to specify numbers in other bases, often octal (base 8) and hexadecimal (base 16). In octal, the numbers go 0, 1, 2, 3, 4, 5, 6, 7, 10, 11, 12, and so on. Just as ‘11’ in decimal is 1 times 10 plus 1, so ‘11’ in octal is 1 times 8 plus 1. This equals 9 in decimal. In hexadecimal, there are 16 digits. Because the everyday decimal number system only has ten digits (‘0’–‘9’), the letters ‘a’ through ‘f’ are used to represent the rest. (Case in the letters is usually irrelevant; hexadecimal ‘a’ and ‘A’ have the same value.) Thus, ‘11’ in hexadecimal is 1 times 16 plus 1, which equals 17 in decimal.

1 The internal representation of all numbers, including integers, uses double-precision floating-point num-

bers. On most modern systems, these are in IEEE 754 standard format. See Chapter 15 [Arithmetic

and Arbitrary-Precision Arithmetic with gawk], page 341, for much more information.

Just by looking at plain ‘11’, you can’t tell what base it’s in. So, in C, C++, and other languages derived from C, there is a special notation to signify the base. Octal numbers start with a leading ‘0’, and hexadecimal numbers start with a leading ‘0x’ or ‘0X’:

11 Decimal value 11

011 Octal 11, decimal value 9

0x11 Hexadecimal 11, decimal value 17

This example shows the difference:

     $ gawk 'BEGIN { printf "%d, %d, %d\n", 011, 11, 0x11 }'
     ⊣ 9, 11, 17

Being able to use octal and hexadecimal constants in your programs is most useful when working with data that cannot be represented conveniently as characters or as regular numbers, such as binary data of various sorts.

gawk allows the use of octal and hexadecimal constants in your program text. However, such numbers in the input data are not treated differently; doing so by default would break old programs. (If you really need to do this, use the --non-decimal-data commandline option; see Section 12.1 [Allowing Nondecimal Input Data], page 299.) If you have octal or hexadecimal data, you can use the strtonum() function (see Section 9.1.3 [String-Manipulation Functions], page 177) to convert the data into a number. Most of the time, you will want to use octal or hexadecimal constants when working with the built-in bitmanipulation functions; see Section 9.1.6 [Bit-Manipulation Functions], page 196, for more information.

Unlike in some early C implementations, ‘8’ and ‘9’ are not valid in octal constants. For example, gawk treats ‘018’ as decimal 18:

     $ gawk 'BEGIN { print "021 is", 021 ; print 018 }'
     ⊣ 021 is 17
     ⊣ 18

Octal and hexadecimal source code constants are a gawk extension. If gawk is in compatibility mode (see Section 2.2 [Command-Line Options], page 31), they are not available.

A Constant’s Base Does Not Affect Its Value Once a numeric constant has been converted internally into a number, gawk no longer

remembers what the original form of the constant was; the internal value is always used. This has particular consequences for conversion of numbers to strings:

      $ gawk 'BEGIN { printf "0x11 is <%s>\n", 0x11 }'
      ⊣ 0x11 is <17>

6.1.1.3 Regular Expression Constants

A regexp constant is a regular expression description enclosed in slashes, such as /^beginning and end$/. Most regexps used in awk programs are constant, but the ‘~’ and ‘!~’ matching operators can also match computed or dynamic regexps (which are typically just ordinary strings or variables that contain a regexp, but could be more complex expressions).

6.1.2 Using Regular Expression Constants

When used on the righthand side of the ‘~’ or ‘!~’ operators, a regexp constant merely stands for the regexp that is to be matched. However, regexp constants (such as /foo/) may be used like simple expressions. When a regexp constant appears by itself, it has the same meaning as if it appeared in a pattern (i.e., ‘($0 ~ /foo/)’). See Section 7.1.2 [Expressions as Patterns], page 131. This means that the following two code segments:

     if ($0 ~ /barfly/ || $0 ~ /camelot/)
         print "found"

and:

     if (/barfly/ || /camelot/)
         print "found"

are exactly equivalent. One rather bizarre consequence of this rule is that the following Boolean expression is valid, but does not do what its author probably intended:

     # Note that /foo/ is on the left of the ~
     if (/foo/ ~ $1) print "found foo"

This code is “obviously” testing $1 for a match against the regexp /foo/. But in fact, the expression ‘/foo/ ~ $1’ really means ‘($0 ~ /foo/) ~ $1’. In other words, first match the input record against the regexp /foo/. The result is either zero or one, depending upon the success or failure of the match. That result is then matched against the first field in the record. Because it is unlikely that you would ever really want to make this kind of test, gawk issues a warning when it sees this construct in a program. Another consequence of this rule is that the assignment statement:

     matches = /foo/

assigns either zero or one to the variable matches, depending upon the contents of the current input record.

Constant regular expressions are also used as the first argument for the gensub(), sub(), and gsub() functions, as the second argument of the match() function, and as the third argument of the split() and patsplit() functions (see Section 9.1.3 [String-Manipulation Functions], page 177). Modern implementations of awk, including gawk, allow the third argument of split() to be a regexp constant, but some older implementations do not.

Because some built-in functions accept regexp constants as arguments, confusion can arise when attempting to use regexp constants as arguments to user-defined functions (see Section 9.2 [User-Defined Functions], page 199). For example:

     function mysub(pat, repl, str, global)
     {
         if (global)
             gsub(pat, repl, str)
         else
             sub(pat, repl, str)
         return str

     }
     {
         ...
         text = "hi! hi yourself!"
         mysub(/hi/, "howdy", text, 1)
         ...
     }

In this example, the programmer wants to pass a regexp constant to the user-defined function mysub(), which in turn passes it on to either sub() or gsub(). However, what really happens is that the pat parameter is assigned a value of either one or zero, depending upon whether or not $0 matches /hi/. gawk issues a warning when it sees a regexp constant used as a parameter to a user-defined function, because passing a truth value in this way is probably not what was intended.

6.1.3 Variables

Variables are ways of storing values at one point in your program for use later in another part of your program. They can be manipulated entirely within the program text, and they can also be assigned values on the awk command line.

6.1.3.1 Using Variables in a Program

Variables let you give names to values and refer to them later. Variables have already been used in many of the examples. The name of a variable must be a sequence of letters, digits, or underscores, and it may not begin with a digit. Here, a letter is any one of the 52 upper- and lowercase English letters. Other characters that may be defined as letters in non-English locales are not valid in variable names. Case is significant in variable names; a and A are distinct variables.

A variable name is a valid expression by itself; it represents the variable’s current value. Variables are given new values with assignment operators, increment operators, and decrement operators (see Section 6.2.3 [Assignment Expressions], page 116). In addition, the sub() and gsub() functions can change a variable’s value, and the match(), split(), and patsplit() functions can change the contents of their array parameters (see Section 9.1.3 [String-Manipulation Functions], page 177).

A few variables have special built-in meanings, such as FS (the field separator) and NF (the number of fields in the current input record). See Section 7.5 [Predefined Variables], page 147, for a list of the predefined variables. These predefined variables can be used and assigned just like all other variables, but their values are also used or changed automatically by awk. All predefined variables’ names are entirely uppercase.

Variables in awk can be assigned either numeric or string values. The kind of value a variable holds can change over the life of a program. By default, variables are initialized to the empty string, which is zero if converted to a number. There is no need to explicitly initialize a variable in awk, which is what you would do in C and in most other traditional languages.

6.1.3.2 Assigning Variables on the Command Line

Any awk variable can be set by including a variable assignment among the arguments on the command line when awk is invoked (see Section 2.3 [Other Command-Line Arguments], page 37). Such an assignment has the following form:

     variable=text

With it, a variable is set either at the beginning of the awk run or in between input files. When the assignment is preceded with the -v option, as in the following:

     -v variable=text

the variable is set at the very beginning, even before the BEGIN rules execute. The -v option and its assignment must precede all the file name arguments, as well as the program text. (See Section 2.2 [Command-Line Options], page 31, for more information about the -v option.) Otherwise, the variable assignment is performed at a time determined by its position among the input file arguments—after the processing of the preceding input file argument. For example:

     awk '{ print $n }' n=4 inventory-shipped n=2 mail-list

prints the value of field number n for all input records. Before the first file is read, the command line sets the variable n equal to four. This causes the fourth field to be printed in lines from inventory-shipped. After the first file has finished, but before the second file is started, n is set to two, so that the second field is printed in lines from mail-list:

     $ awk '{ print $n }' n=4 inventory-shipped n=2 mail-list
     ⊣ 15
     ⊣ 24
     ...
     ⊣ 555-5553
     ⊣ 555-3412
     ...

Command-line arguments are made available for explicit examination by the awk program in the ARGV array (see Section 7.5.3 [Using ARGC and ARGV], page 155). awk processes the values of command-line assignments for escape sequences (see Section 3.2 [Escape Sequences], page 46).

6.1.4 Conversion of Strings and Numbers

Number-to-string and string-to-number conversion are generally straightforward. There can be subtleties to be aware of; this section discusses this important facet of awk.

6.1.4.1 How awk Converts Between Strings and Numbers

Strings are converted to numbers and numbers are converted to strings, if the context of the awk program demands it. For example, if the value of either foo or bar in the expression ‘foo + bar’ happens to be a string, it is converted to a number before the addition is performed. If numeric values appear in string concatenation, they are converted to strings. Consider the following:

     two = 2; three = 3
     print (two three) + 4

This prints the (numeric) value 27. The numeric values of the variables two and three are converted to strings and concatenated together. The resulting string is converted back to the number 23, to which 4 is then added.

If, for some reason, you need to force a number to be converted to a string, concatenate that number with the empty string, "". To force a string to be converted to a number, add zero to that string. A string is converted to a number by interpreting any numeric prefix of the string as numerals: "2.5" converts to 2.5, "1e3" converts to 1,000, and "25fix" has a numeric value of 25. Strings that can’t be interpreted as valid numbers convert to zero.

The exact manner in which numbers are converted into strings is controlled by the awk predefined variable CONVFMT (see Section 7.5 [Predefined Variables], page 147). Numbers are converted using the sprintf() function with CONVFMT as the format specifier (see Section 9.1.3 [String-Manipulation Functions], page 177).

CONVFMT’s default value is "%.6g", which creates a value with at most six significant digits. For some applications, you might want to change it to specify more precision. On most modern machines, 17 digits is usually enough to capture a floating-point number’s value exactly.^{2}

Strange results can occur if you set CONVFMT to a string that doesn’t tell sprintf() how to format floating-point numbers in a useful way. For example, if you forget the ‘%’ in the format, awk converts all numbers to the same constant string.

As a special case, if a number is an integer, then the result of converting it to a string is always an integer, no matter what the value of CONVFMT may be. Given the following code fragment:

     CONVFMT = "%2.2f"
     a = 12
     b = a ""

b has the value "12", not "12.00".

Pre-POSIX awk Used OFMT for String Conversion Prior to the POSIX standard, awk used the value of OFMT for converting numbers to

strings. OFMT specifies the output format to use when printing numbers with print. CONVFMT was introduced in order to separate the semantics of conversion from the semantics of printing. Both CONVFMT and OFMT have the same default value: "%.6g". In the vast majority of cases, old awk programs do not change their behavior. See Section 5.1 [The print Statement], page 89, for more information on the print statement.

6.1.4.2 Locales Can Influence Conversion

Where you are can matter when it comes to converting between numbers and strings. The local character set and language—the locale—can affect numeric formats. In particular, for awk programs, it affects the decimal point character and the thousands-separator character. The "C" locale, and most English-language locales, use the period character (‘.’) as the decimal point and don’t have a thousands separator. However, many (if not most) European and non-English locales use the comma (‘,’) as the decimal point character. European locales often use either a space or a period as the thousands separator, if they have one.

2 Pathological cases can require up to 752 digits (!), but we doubt that you need to worry about this.

The POSIX standard says that awk always uses the period as the decimal point when reading the awk program source code, and for command-line variable assignments (see Section 2.3 [Other Command-Line Arguments], page 37). However, when interpreting input data, for print and printf output, and for number-to-string conversion, the local decimal point character is used. In all cases, numbers in source code and in input data cannot have a thousands separator. Here are some examples indicating the difference in behavior, on a GNU/Linux system:

     $ export POSIXLY_CORRECT=1  Force POSIX behavior
     $ gawk 'BEGIN { printf "%g\n", 3.1415927 }'
     ⊣ 3.14159
     $ LC_ALL=en_DK.utf-8 gawk 'BEGIN { printf "%g\n", 3.1415927 }'
     ⊣ 3,14159
     $ echo 4,321 | gawk '{ print $1 + 1 }'
     ⊣ 5
     $ echo 4,321 | LC_ALL=en_DK.utf-8 gawk '{ print $1 + 1 }'
     ⊣ 5,321

The en_DK.utf-8 locale is for English in Denmark, where the comma acts as the decimal point separator. In the normal "C" locale, gawk treats ‘4,321’ as 4, while in the Danish locale, it’s treated as the full number including the fractional part, 4.321.

Some earlier versions of gawk fully complied with this aspect of the standard. However, many users in non-English locales complained about this behavior, because their data used a period as the decimal point, so the default behavior was restored to use a period as the decimal point character. You can use the --use-lc-numeric option (see Section 2.2 [Command-Line Options], page 31) to force gawk to use the locale’s decimal point character. (gawk also uses the locale’s decimal point character when in POSIX mode, either via --posix or the POSIXLY_CORRECT environment variable, as shown previously.)

Table 6.1 describes the cases in which the locale’s decimal point character is used and when a period is used. Some of these features have not been described yet.

Feature  Default  --posix or --use-lc-numeric

%'g  Use locale  Use locale
%g  Use period  Use locale
Input  Use period  Use locale
strtonum() Use period  Use locale

Table 6.1: Locale decimal point versus a period

Finally, modern-day formal standards and the IEEE standard floating-point representation can have an unusual but important effect on the way gawk converts some special string values to numbers. The details are presented in Section 15.6 [Standards Versus Existing Practice], page 351.

6.2 Operators: Doing Something with Values

This section introduces the operators that make use of the values provided by constants and variables.

6.2.1 Arithmetic Operators

The awk language uses the common arithmetic operators when evaluating expressions. All of these arithmetic operators follow normal precedence rules and work as you would expect them to.

The following example uses a file named grades, which contains a list of student names as well as three test scores per student (it’s a small class):

     Pat   100 97 58
     Sandy  84 72 93
     Chris  72 92 89

This program takes the file grades and prints the average of the scores:

     $ awk '{ sum = $2 + $3 + $4 ; avg = sum / 3
     >        print $1, avg }' grades
     ⊣ Pat 85
     ⊣ Sandy 83
     ⊣ Chris 84.3333

The following list provides the arithmetic operators in awk, in order from the highest precedence to the lowest:

x ^y

x ** y Exponentiation; x raised to the y power. ‘2 ^3’ has the value eight; the character sequence ‘**’ is equivalent to ‘^’. (c.e.)

- x Negation.

+ x Unary plus; the expression is converted to a number.

x * y Multiplication.

x / y Division; because all numbers in awk are floating-point numbers, the result is not rounded to an integer—‘3 / 4’ has the value 0.75. (It is a common mistake, especially for C programmers, to forget that all numbers in awk are floating point, and that division of integer-looking constants produces a real number, not an integer.)

x % y Remainder; further discussion is provided in the text, just after this list.

x + y Addition.

x - y Subtraction.

Unary plus and minus have the same precedence, the multiplication operators all have the same precedence, and addition and subtraction have the same precedence.

When computing the remainder of ‘x % y’, the quotient is rounded toward zero to an integer and multiplied by y. This result is subtracted from x; this operation is sometimes known as “trunc-mod.” The following relation always holds:

     b * int(a / b) + (a % b) == a

One possibly undesirable effect of this definition of remainder is that ‘x % y’ is negative if x is negative. Thus:

     -17 % 8 = -1

In other awk implementations, the signedness of the remainder may be machinedependent.

NOTE: The POSIX standard only specifies the use of ‘^’ for exponentiation.

For maximum portability, do not use the ‘**’ operator.

6.2.2 String Concatenation

It seemed like a good idea at the time.

—Brian Kernighan There is only one string operation: concatenation. It does not have a specific operator

to represent it. Instead, concatenation is performed by writing expressions next to one another, with no operator. For example:

     $ awk '{ print "Field number one: " $1 }' mail-list
     ⊣ Field number one: Amelia
     ⊣ Field number one: Anthony
     ...

Without the space in the string constant after the ‘:’, the line runs together. For example:

     $ awk '{ print "Field number one:" $1 }' mail-list
     ⊣ Field number one:Amelia
     ⊣ Field number one:Anthony
     ...

Because string concatenation does not have an explicit operator, it is often necessary to ensure that it happens at the right time by using parentheses to enclose the items to concatenate. For example, you might expect that the following code fragment concatenates

file and name:
     file = "file"
     name = "name"
     print "something meaningful" > file name

This produces a syntax error with some versions of Unix awk.^{3} It is necessary to use the following:

     print "something meaningful" > (file name)

Parentheses should be used around concatenation in all but the most common contexts, such as on the righthand side of ‘=’. Be careful about the kinds of expressions used in string concatenation. In particular, the order of evaluation of expressions used for concatenation is undefined in the awk language. Consider this example:

     BEGIN {
         a = "don't"
         print (a " " (a = "panic"))
     }

It is not defined whether the second assignment to a happens before or after the value of a is retrieved for producing the concatenated value. The result could be either ‘don't panic’,

or ‘panic panic’.

3 It happens that BWK awk, gawk, and mawk all “get it right,” but you should not rely on this.

The precedence of concatenation, when mixed with other operators, is often counterintuitive. Consider this example:

     $ awk 'BEGIN { print -12 " " -24 }'
     ⊣ -12-24

This “obviously” is concatenating −12, a space, and −24. But where did the space disappear to? The answer lies in the combination of operator precedences and awk’s automatic conversion rules. To get the desired result, write the program this way:

     $ awk 'BEGIN { print -12 " " (-24) }'
     ⊣ -12 -24

This forces awk to treat the ‘-’ on the ‘-24’ as unary. Otherwise, it’s parsed as follows:

−12 (" " − 24) ⇒ −12 (0 − 24) ⇒ −12 (−24) ⇒ −12−24

As mentioned earlier, when mixing concatenation with other operators, parenthesize. Otherwise, you’re never quite sure what you’ll get.

6.2.3 Assignment Expressions

An assignment is an expression that stores a (usually different) value into a variable. For example, let’s assign the value one to the variable z:

     z = 1

After this expression is executed, the variable z has the value one. Whatever old value z had before the assignment is forgotten.

Assignments can also store string values. For example, the following stores the value

"this food is good" in the variable message:
     thing = "food"
     predicate = "good"
     message = "this " thing " is " predicate

This also illustrates string concatenation. The ‘=’ sign is called an assignment operator. It is the simplest assignment operator because the value of the righthand operand is stored unchanged. Most operators (addition, concatenation, and so on) have no effect except to compute a value. If the value isn’t used, there’s no reason to use the operator. An assignment operator is different; it does produce a value, but even if you ignore it, the assignment still makes itself felt through the alteration of the variable. We call this a side effect.

The lefthand operand of an assignment need not be a variable (see Section 6.1.3 [Variables], page 110); it can also be a field (see Section 4.4 [Changing the Contents of a Field], page 65) or an array element (see Chapter 8 [Arrays in awk], page 159). These are all called lvalues, which means they can appear on the lefthand side of an assignment operator. The righthand operand may be any expression; it produces the new value that the assignment stores in the specified variable, field, or array element. (Such values are called rvalues.)

It is important to note that variables do not have permanent types. A variable’s type is simply the type of whatever value was last assigned to it. In the following program fragment, the variable foo has a numeric value at first, and a string value later on:

     foo = 1
     print foo
     foo = "bar"
     print foo

When the second assignment gives foo a string value, the fact that it previously had a numeric value is forgotten.

String values that do not begin with a digit have a numeric value of zero. After executing the following code, the value of foo is five:

     foo = "a string"
     foo = foo + 5

NOTE: Using a variable as a number and then later as a string can be confusing and is poor programming style. The previous two examples illustrate how awk works, not how you should write your programs!

An assignment is an expression, so it has a value—the same value that is assigned. Thus, ‘z = 1’ is an expression with the value one. One consequence of this is that you can write multiple assignments together, such as:

     x = y = z = 5

This example stores the value five in all three variables (x, y, and z). It does so because the value of ‘z = 5’, which is five, is stored into y and then the value of ‘y = z = 5’, which is five, is stored into x.

Assignments may be used anywhere an expression is called for. For example, it is valid to write ‘x != (y = 1)’ to set y to one, and then test whether x equals one. But this style tends to make programs hard to read; such nesting of assignments should be avoided, except perhaps in a one-shot program.

Aside from ‘=’, there are several other assignment operators that do arithmetic with the old value of the variable. For example, the operator ‘+=’ computes a new value by adding the righthand value to the old value of the variable. Thus, the following assignment adds five to the value of foo:

     foo += 5

This is equivalent to the following:

     foo = foo + 5

Use whichever makes the meaning of your program clearer.

There are situations where using ‘+=’ (or any assignment operator) is not the same as simply repeating the lefthand operand in the righthand expression. For example:

     # Thanks to Pat Rankin for this example
     BEGIN  {
         foo[rand()] += 5
         for (x in foo)
            print x, foo[x]
         bar[rand()] = bar[rand()] + 5
         for (x in bar)
            print x, bar[x]

     }

The indices of bar are practically guaranteed to be different, because rand() returns different values each time it is called. (Arrays and the rand() function haven’t been covered yet. See Chapter 8 [Arrays in awk], page 159, and see Section 9.1.2 [Numeric Functions], page 176, for more information.) This example illustrates an important fact about assignment operators: the lefthand expression is only evaluated once.

It is up to the implementation as to which expression is evaluated first, the lefthand or the righthand. Consider this example:

     i = 1
     a[i += 2] = i + 1

The value of a[3] could be either two or four.

Table 6.2 lists the arithmetic assignment operators. In each case, the righthand operand is an expression whose value is converted to a number.

Operator  Effect
lvalue += increment  Add increment to the value of lvalue.
lvalue -= decrement  Subtract decrement from the value of lvalue.
lvalue *= coefficient  Multiply the value of lvalue by coefficient.
lvalue /= divisor  Divide the value of lvalue by divisor.
lvalue %= modulus  Set lvalue to its remainder by modulus.
lvalue ^= power  Raise lvalue to the power power.
lvalue **= power  Raise lvalue to the power power. (c.e.)

Table 6.2: Arithmetic assignment operators

NOTE: Only the ‘^=’ operator is specified by POSIX. For maximum portability, do not use the ‘**=’ operator.

Syntactic Ambiguities Between ‘/=’ and Regular Expressions There is a syntactic ambiguity between the /= assignment operator and regexp constants

whose first character is an ‘=’. This is most notable in some commercial awk versions. For example:

      $ awk /==/ /dev/null
       error awk: syntax error at source line 1
       error context is
       error  >>> /= <<<
       error awk: bailing out at source line 1

A workaround is:

      awk '/[=]=/' /dev/null

gawk does not have this problem; BWK awk and mawk also do not.

6.2.4 Increment and Decrement Operators

Increment and decrement operators increase or decrease the value of a variable by one. An assignment operator can do the same thing, so the increment operators add no power to the awk language; however, they are convenient abbreviations for very common operations.

The operator used for adding one is written ‘++’. It can be used to increment a variable either before or after taking its value. To pre-increment a variable v, write ‘++v’. This adds one to the value of v—that new value is also the value of the expression. (The assignment expression ‘v += 1’ is completely equivalent.) Writing the ‘++’ after the variable specifies post-increment. This increments the variable value just the same; the difference is that the value of the increment expression itself is the variable’s old value. Thus, if foo has the value four, then the expression ‘foo++’ has the value four, but it changes the value of foo to five. In other words, the operator returns the old value of the variable, but with the side effect of incrementing it.

The post-increment ‘foo++’ is nearly the same as writing ‘(foo += 1) - 1’. It is not perfectly equivalent because all numbers in awk are floating point—in floating point, ‘foo + 1 - 1’ does not necessarily equal foo. But the difference is minute as long as you stick to numbers that are fairly small (less than 10^{12}).

Fields and array elements are incremented just like variables. (Use ‘$(i++)’ when you want to do a field reference and a variable increment at the same time. The parentheses are necessary because of the precedence of the field reference operator ‘$’.)

The decrement operator ‘--’ works just like ‘++’, except that it subtracts one instead of adding it. As with ‘++’, it can be used before the lvalue to pre-decrement or after it to post-decrement. Following is a summary of increment and decrement expressions:

++lvalue Increment lvalue, returning the new value as the value of the expression.

lvalue++ Increment lvalue, returning the old value of lvalue as the value of the expression.

--lvalue Decrement lvalue, returning the new value as the value of the expression. (This expression is like ‘++lvalue’, but instead of adding, it subtracts.)

lvalue-- Decrement lvalue, returning the old value of lvalue as the value of the expression. (This expression is like ‘lvalue++’, but instead of adding, it subtracts.)

Operator Evaluation Order

Doctor, it hurts when I do this!

Then don’t do that!

—Groucho Marx

What happens for something like the following?

      b = 6
      print b += b++

Or something even stranger?

      b = 6
      b += ++b + b++
      print b

In other words, when do the various side effects prescribed by the postfix operators (‘b++’) take effect? When side effects happen is implementation-defined. In other words, it is up to the particular version of awk. The result for the first example may be 12 or 13, and for the second, it may be 22 or 23.

In short, doing things like this is not recommended and definitely not anything that you can rely upon for portability. You should avoid such things in your own programs.

6.3 Truth Values and Conditions

In certain contexts, expression values also serve as “truth values”; i.e., they determine what should happen next as the program runs. This section describes how awk defines “true” and “false” and how values are compared.

6.3.1 True and False in awk

Many programming languages have a special representation for the concepts of “true” and “false.” Such languages usually use the special constants true and false, or perhaps their uppercase equivalents. However, awk is different. It borrows a very simple concept of true and false from C. In awk, any nonzero numeric value or any nonempty string value is true. Any other value (zero or the null string, "") is false. The following program prints ‘A

strange truth value’ three times:
     BEGIN {
        if (3.1415927)
            print "A strange truth value"
        if ("Four Score And Seven Years Ago")
            print "A strange truth value"
        if (j = 57)
            print "A strange truth value"
     }

There is a surprising consequence of the “nonzero or non-null” rule: the string constant "0" is actually true, because it is non-null.

6.3.2 Variable Typing and Comparison Expressions

The Guide is definitive. Reality is frequently inaccurate. —Douglas Adams, The Hitchhiker’s Guide to the Galaxy

Unlike in other programming languages, in awk variables do not have a fixed type. Instead, they can be either a number or a string, depending upon the value that is assigned to them. We look now at how variables are typed, and how awk compares variables.

6.3.2.1 String Type versus Numeric Type

The POSIX standard introduced the concept of a numeric string, which is simply a string that looks like a number—for example, " +2". This concept is used for determining the type of a variable. The type of the variable is important because the types of two variables determine how they are compared. Variable typing follows these rules:

• A numeric constant or the result of a numeric operation has the numeric attribute.

• A string constant or the result of a string operation has the string attribute.

• Fields, getline input, FILENAME, ARGV elements, ENVIRON elements, and the elements of an array created by match(), split(), and patsplit() that are numeric strings have the strnum attribute. Otherwise, they have the string attribute. Uninitialized variables also have the strnum attribute.

• Attributes propagate across assignments but are not changed by any use.

The last rule is particularly important. In the following program, a has numeric type, even though it is later used in a string operation:

     BEGIN {
          a = 12.345
          b = a " is a cute number"
          print b
     }

When two operands are compared, either string comparison or numeric comparison may be used. This depends upon the attributes of the operands, according to the following symmetric matrix:

STRING NUMERIC STRNUM
STRING  string  string  string
NUMERIC  string  numeric  numeric
STRNUM  string  numeric  numeric

The basic idea is that user input that looks numeric—and only user input—should be treated as numeric, even though it is actually made of characters and is therefore also a string. Thus, for example, the string constant " +3.14", when it appears in program source code, is a string—even though it looks numeric—and is never treated as a number for comparison purposes.

In short, when one operand is a “pure” string, such as a string constant, then a string comparison is performed. Otherwise, a numeric comparison is performed.

This point bears additional emphasis: All user input is made of characters, and so is first and foremost of string type; input strings that look numeric are additionally given the strnum attribute. Thus, the six-character input string ‘ +3.14’ receives the strnum attribute. In contrast, the eight characters " +3.14" appearing in program text comprise a string constant. The following examples print ‘1’ when the comparison between the two different constants is true, and ‘0’ otherwise:

     $ echo ' +3.14' | awk '{ print($0 == " +3.14") }'  True
     ⊣ 1
     $ echo ' +3.14' | awk '{ print($0 == "+3.14") }'  False
     ⊣ 0
     $ echo ' +3.14' | awk '{ print($0 == "3.14") }'  False
     ⊣ 0
     $ echo ' +3.14' | awk '{ print($0 == 3.14) }'  True
     ⊣ 1
     $ echo ' +3.14' | awk '{ print($1 == " +3.14") }'  False
     ⊣ 0
     $ echo ' +3.14' | awk '{ print($1 == "+3.14") }'  True
     ⊣ 1
     $ echo ' +3.14' | awk '{ print($1 == "3.14") }'  False
     ⊣ 0
     $ echo ' +3.14' | awk '{ print($1 == 3.14) }'  True
     ⊣ 1

6.3.2.2 Comparison Operators

Comparison expressions compare strings or numbers for relationships such as equality. They are written using relational operators, which are a superset of those in C. Table 6.3 describes them.

Expression  Result
x < y  True if x is less than y
x <= y  True if x is less than or equal to y
x > y  True if x is greater than y
x >= y  True if x is greater than or equal to y
x == y  True if x is equal to y
x != y  True if x is not equal to y
x ~ y  True if the string x matches the regexp denoted by y
x !~ y  True if the string x does not match the regexp denoted by y
subscript in array  True if the array array has an element with the subscript subscript

Table 6.3: Relational operators

Comparison expressions have the value one if true and zero if false. When comparing operands of mixed types, numeric operands are converted to strings using the value of CONVFMT (see Section 6.1.4 [Conversion of Strings and Numbers], page 111).

Strings are compared by comparing the first character of each, then the second character of each, and so on. Thus, "10" is less than "9". If there are two strings where one is a prefix of the other, the shorter string is less than the longer one. Thus, "abc" is less than

"abcd".

It is very easy to accidentally mistype the ‘==’ operator and leave off one of the ‘=’ characters. The result is still valid awk code, but the program does not do what is intended:

     if (a = b)   # oops! should be a == b
        ...

     else
        ...

Unless b happens to be zero or the null string, the if part of the test always succeeds. Because the operators are so similar, this kind of error is very difficult to spot when scanning the source code.

The following list of expressions illustrates the kinds of comparisons awk performs, as well as what the result of each comparison is:

1.5 <= 2.0

Numeric comparison (true)

"abc" >= "xyz"

String comparison (false)

1.5 != " +2"

String comparison (true)

"1e2" < "3"

String comparison (true)

a = 2; b = "2"

a == b String comparison (true)

a = 2; b = " +2"

a == b String comparison (false)

In this example:

     $ echo 1e2 3 | awk '{ print ($1 < $2) ? "true" : "false" }'
     ⊣ false

the result is ‘false’ because both $1 and $2 are user input. They are numeric strings— therefore both have the strnum attribute, dictating a numeric comparison. The purpose of the comparison rules and the use of numeric strings is to attempt to produce the behavior that is “least surprising,” while still “doing the right thing.”

String comparisons and regular expression comparisons are very different. For example:

     x == "foo"

has the value one, or is true if the variable x is precisely ‘foo’. By contrast:

     x ~ /foo/

has the value one if x contains ‘foo’, such as "Oh, what a fool am I!".

The righthand operand of the ‘~’ and ‘!~’ operators may be either a regexp constant (/…/) or an ordinary expression. In the latter case, the value of the expression as a string is used as a dynamic regexp (see Section 3.1 [How to Use Regular Expressions], page 45; also see Section 3.6 [Using Dynamic Regexps], page 52).

A constant regular expression in slashes by itself is also an expression. /regexp/ is an abbreviation for the following comparison expression:

     $0 ~ /regexp/

One special place where /foo/ is not an abbreviation for ‘$0 ~ /foo/’ is when it is the righthand operand of ‘~’ or ‘!~’. See Section 6.1.2 [Using Regular Expression Constants], page 109, where this is discussed in more detail.

6.3.2.3 String Comparison with POSIX Rules

The POSIX standard says that string comparison is performed based on the locale’s collating order. This is the order in which characters sort, as defined by the locale (for more discussion, see Section 6.6 [Where You Are Makes a Difference], page 129). This order is usually very different from the results obtained when doing straight character-by-character comparison.^{4}

Because this behavior differs considerably from existing practice, gawk only implements it when in POSIX mode (see Section 2.2 [Command-Line Options], page 31). Here is an example to illustrate the difference, in an en_US.UTF-8 locale:

     $ gawk 'BEGIN { printf("ABC < abc = %s\n",
     >                     ("ABC" < "abc" ? "TRUE" : "FALSE")) }'
     ⊣ ABC < abc = TRUE
     $ gawk --posix 'BEGIN { printf("ABC < abc = %s\n",
     >                             ("ABC" < "abc" ? "TRUE" : "FALSE")) }'
     ⊣ ABC < abc = FALSE

6.3.3 Boolean Expressions

A Boolean expression is a combination of comparison expressions or matching expressions, using the Boolean operators “or” (‘||’), “and” (‘&&’), and “not” (‘!’), along with parentheses to control nesting. The truth value of the Boolean expression is computed by combining the truth values of the component expressions. Boolean expressions are also referred to as logical expressions. The terms are equivalent.

Boolean expressions can be used wherever comparison and matching expressions can be used. They can be used in if, while, do, and for statements (see Section 7.4 [Control Statements in Actions], page 138). They have numeric values (one if true, zero if false) that come into play if the result of the Boolean expression is stored in a variable or used in arithmetic.

In addition, every Boolean expression is also a valid pattern, so you can use one as a pattern to control the execution of rules. The Boolean operators are:

boolean1 && boolean2

True if both boolean1 and boolean2 are true. For example, the following statement prints the current input record if it contains both ‘edu’ and ‘li’:

               if ($0 ~ /edu/ && $0 ~ /li/) print

The subexpression boolean2 is evaluated only if boolean1 is true. This can make a difference when boolean2 contains expressions that have side effects. In the case of ‘$0 ~ /foo/ && ($2 == bar++)’, the variable bar is not incremented if there is no substring ‘foo’ in the record.

boolean1 || boolean2

True if at least one of boolean1 or boolean2 is true. For example, the following statement prints all records in the input that contain either ‘edu’ or ‘li’:

               if ($0 ~ /edu/ || $0 ~ /li/) print

4 Technically, string comparison is supposed to behave the same way as if the strings were compared with

the C strcoll() function.

The subexpression boolean2 is evaluated only if boolean1 is false. This can make a difference when boolean2 contains expressions that have side effects. (Thus, this test never really distinguishes records that contain both ‘edu’ and ‘li’—as soon as ‘edu’ is matched, the full test succeeds.)

! boolean True if boolean is false. For example, the following program prints ‘no home!’

in the unusual event that the HOME environment variable is not defined:

               BEGIN { if (! ("HOME" in ENVIRON))
                           print "no home!" }

(The in operator is described in Section 8.1.2 [Referring to an Array Element], page 161.)

The ‘&&’ and ‘||’ operators are called short-circuit operators because of the way they work. Evaluation of the full expression is “short-circuited” if the result can be determined partway through its evaluation.

Statements that end with ‘&&’ or ‘||’ can be continued simply by putting a newline after them. But you cannot put a newline in front of either of these operators without using backslash continuation (see Section 1.6 [awk Statements Versus Lines], page 27).

The actual value of an expression using the ‘!’ operator is either one or zero, depending upon the truth value of the expression it is applied to. The ‘!’ operator is often useful for changing the sense of a flag variable from false to true and back again. For example, the following program is one way to print lines in between special bracketing lines:

     $1 == "START"   { interested = ! interested; next }
     interested      { print }
     $1 == "END"     { interested = ! interested; next }

The variable interested, as with all awk variables, starts out initialized to zero, which is also false. When a line is seen whose first field is ‘START’, the value of interested is toggled to true, using ‘!’. The next rule prints lines as long as interested is true. When a line is seen whose first field is ‘END’, interested is toggled back to false.^{5}

Most commonly, the ‘!’ operator is used in the conditions of if and while statements, where it often makes more sense to phrase the logic in the negative:

     if (! some condition || some other condition) {
         ... do whatever processing ...
     }

NOTE: The next statement is discussed in Section 7.4.8 [The next Statement], page 144. next tells awk to skip the rest of the rules, get the next record, and start processing the rules over again at the top. The reason it’s there is to avoid printing the bracketing ‘START’ and ‘END’ lines.

6.3.4 Conditional Expressions

A conditional expression is a special kind of expression that has three operands. It allows you to use one expression’s value to select one of two other expressions. The conditional expression in awk is the same as in the C language, as shown here:

5 This program has a bug; it prints lines starting with ‘END’. How would you fix it?

     selector ? if-true-exp : if-false-exp

There are three subexpressions. The first, selector, is always computed first. If it is “true” (not zero or not null), then if-true-exp is computed next, and its value becomes the value of the whole expression. Otherwise, if-false-exp is computed next, and its value becomes the value of the whole expression. For example, the following expression produces the absolute value of x:

     x >= 0 ? x : -x

Each time the conditional expression is computed, only one of if-true-exp and if-false-exp is used; the other is ignored. This is important when the expressions have side effects. For example, this conditional expression examines element i of either array a or array b, and increments i:

     x == y ? a[i++] : b[i++]

This is guaranteed to increment i exactly once, because each time only one of the two increment expressions is executed and the other is not. See Chapter 8 [Arrays in awk], page 159, for more information about arrays.

As a minor gawk extension, a statement that uses ‘?:’ can be continued simply by putting a newline after either character. However, putting a newline in front of either character does not work without using backslash continuation (see Section 1.6 [awk Statements Versus Lines], page 27). If --posix is specified (see Section 2.2 [Command-Line Options], page 31), this extension is disabled.

6.4 Function Calls

A function is a name for a particular calculation. This enables you to ask for it by name at any point in the program. For example, the function sqrt() computes the square root of a number.

A fixed set of functions are built in, which means they are available in every awk program. The sqrt() function is one of these. See Section 9.1 [Built-in Functions], page 175, for a list of built-in functions and their descriptions. In addition, you can define functions for use in your program. See Section 9.2 [User-Defined Functions], page 199, for instructions on how to do this. Finally, gawk lets you write functions in C or C++ that may be called from your program (see Chapter 16 [Writing Extensions for gawk], page 355).

The way to use a function is with a function call expression, which consists of the function name followed immediately by a list of arguments in parentheses. The arguments are expressions that provide the raw materials for the function’s calculations. When there is more than one argument, they are separated by commas. If there are no arguments, just write ‘()’ after the function name. The following examples show function calls with and without arguments:

     sqrt(x^2 + y^2)  one argument

atan2(y, x) two arguments rand() no arguments CAUTION: Do not put any space between the function name and the opening parenthesis! A user-defined function name looks just like the name of a variable—a space would make the expression look like concatenation of a variable with an expression inside parentheses. With built-in functions, space before the parenthesis is harmless, but it is best not to get into the habit of using space to avoid mistakes with user-defined functions.

Each function expects a particular number of arguments. For example, the sqrt() function must be called with a single argument, the number of which to take the square root:

     sqrt(argument)

Some of the built-in functions have one or more optional arguments. If those arguments are not supplied, the functions use a reasonable default value. See Section 9.1 [Built-in Functions], page 175, for full details. If arguments are omitted in calls to user-defined functions, then those arguments are treated as local variables. Such local variables act like the empty string if referenced where a string value is required, and like zero if referenced where a numeric value is required (see Section 9.2 [User-Defined Functions], page 199).

As an advanced feature, gawk provides indirect function calls, which is a way to choose the function to call at runtime, instead of when you write the source code to your program. We defer discussion of this feature until later; see Section 9.3 [Indirect Function Calls], page 208.

Like every other expression, the function call has a value, often called the return value, which is computed by the function based on the arguments you give it. In this example, the return value of ‘sqrt(argument)’ is the square root of argument. The following program reads numbers, one number per line, and prints the square root of each one:

     $ awk '{ print "The square root of", $1, "is", sqrt($1) }'
     1
     ⊣ The square root of 1 is 1
     3
     ⊣ The square root of 3 is 1.73205
     5
     ⊣ The square root of 5 is 2.23607
     Ctrl-d

A function can also have side effects, such as assigning values to certain variables or doing I/O. This program shows how the match() function (see Section 9.1.3 [String-Manipulation Functions], page 177) changes the variables RSTART and RLENGTH:

     {
         if (match($1, $2))
             print RSTART, RLENGTH
         else
             print "no match"
     }

Here is a sample run:

     $ awk -f matchit.awk
     aaccdd  c+
     ⊣ 3 2
     foo     bar
     ⊣ no match
     abcdefg e
     ⊣ 5 1

6.5 Operator Precedence (How Operators Nest)

Operator precedence determines how operators are grouped when different operators appear close by in one expression. For example, ‘*’ has higher precedence than ‘+’; thus, ‘a + b * c’ means to multiply b and c, and then add a to the product (i.e., ‘a + (b * c)’).

The normal precedence of the operators can be overruled by using parentheses. Think of the precedence rules as saying where the parentheses are assumed to be. In fact, it is wise to always use parentheses whenever there is an unusual combination of operators, because other people who read the program may not remember what the precedence is in this case. Even experienced programmers occasionally forget the exact rules, which leads to mistakes. Explicit parentheses help prevent any such mistakes.

When operators of equal precedence are used together, the leftmost operator groups first, except for the assignment, conditional, and exponentiation operators, which group in the opposite order. Thus, ‘a - b + c’ groups as ‘(a - b) + c’ and ‘a = b = c’ groups as ‘a =

(b = c)’.

Normally the precedence of prefix unary operators does not matter, because there is only one way to interpret them: innermost first. Thus, ‘$++i’ means ‘$(++i)’ and ‘++$x’ means ‘++($x)’. However, when another operator follows the operand, then the precedence of the unary operators can matter. ‘$x^2’ means ‘($x)^2’, but ‘-x^2’ means ‘-(x^2)’, because ‘-’ has lower precedence than ‘^’, whereas ‘$’ has higher precedence. Also, operators cannot be combined in a way that violates the precedence rules; for example, ‘$$0++--’ is not a valid expression because the first ‘$’ has higher precedence than the ‘++’; to avoid the problem the expression can be rewritten as ‘$($0++)--’.

This list presents awk’s operators, in order of highest to lowest precedence:

(…) Grouping.

$ Field reference.

++ -- Increment, decrement.

^** Exponentiation. These operators group right to left.

+ - ! Unary plus, minus, logical “not.”

* / % Multiplication, division, remainder.

+ - Addition, subtraction.

String concatenation There is no special symbol for concatenation. The operands are simply written side by side (see Section 6.2.2 [String Concatenation], page 115).

< <= == != > >= >> | |&

Relational and redirection. The relational operators and the redirections have the same precedence level. Characters such as ‘>’ serve both as relationals and as redirections; the context distinguishes between the two meanings. Note that the I/O redirection operators in print and printf statements belong to the statement level, not to expressions. The redirection does not produce an expression that could be the operand of another operator. As a result, it does not make sense to use a redirection operator near another operator of lower precedence without parentheses. Such combinations (e.g., ‘print foo > a ? b : c’) result in syntax errors. The correct way to write this statement is ‘print

          foo > (a ? b : c)’.

~ !~ Matching, nonmatching.

in Array membership.

&& Logical “and.”

|| Logical “or.”

?: Conditional. This operator groups right to left.

= += -= *= /= %= ^= **=

Assignment. These operators group right to left.

NOTE: The ‘|&’, ‘**’, and ‘**=’ operators are not specified by POSIX. For maximum portability, do not use them.

6.6 Where You Are Makes a Difference

Modern systems support the notion of locales: a way to tell the system about the local character set and language. The ISO C standard defines a default "C" locale, which is an environment that is typical of what many C programmers are used to.

Once upon a time, the locale setting used to affect regexp matching, but this is no longer true (see Section A.8 [Regexp Ranges and Locales: A Long Sad Story], page 424).

Locales can affect record splitting. For the normal case of ‘RS = "\n"’, the locale is largely irrelevant. For other single-character record separators, setting ‘LC_ALL=C’ in the environment will give you much better performance when reading records. Otherwise, gawk has to make several function calls, per input character, to find the record terminator.

Locales can affect how dates and times are formatted (see Section 9.1.5 [Time Functions], page 192). For example, a common way to abbreviate the date September 4, 2015, in the United States is “9/4/15.” In many countries in Europe, however, it is abbreviated “4.9.15.” Thus, the ‘%x’ specification in a "US" locale might produce ‘9/4/15’, while in a "EUROPE" locale, it might produce ‘4.9.15’.

According to POSIX, string comparison is also affected by locales (similar to regular expressions). The details are presented in Section 6.3.2.3 [String Comparison with POSIX Rules], page 124.

Finally, the locale affects the value of the decimal point character used when gawk parses input data. This is discussed in detail in Section 6.1.4 [Conversion of Strings and Numbers], page 111.

6.7 Summary

• Expressions are the basic elements of computation in programs. They are built from constants, variables, function calls, and combinations of the various kinds of values with operators.

• awk supplies three kinds of constants: numeric, string, and regexp. gawk lets you specify numeric constants in octal and hexadecimal (bases 8 and 16) as well as decimal (base 10). In certain contexts, a standalone regexp constant such as /foo/ has the same meaning as ‘$0 ~ /foo/’.

• Variables hold values between uses in computations. A number of built-in variables provide information to your awk program, and a number of others let you control how awk behaves.

• Numbers are automatically converted to strings, and strings to numbers, as needed by awk. Numeric values are converted as if they were formatted with sprintf() using the format in CONVFMT. Locales can influence the conversions.

• awk provides the usual arithmetic operators (addition, subtraction, multiplication, division, modulus), and unary plus and minus. It also provides comparison operators, Boolean operators, an array membership testing operator, and regexp matching operators. String concatenation is accomplished by placing two expressions next to each other; there is no explicit operator. The three-operand ‘?:’ operator provides an “ifelse” test within expressions.

• Assignment operators provide convenient shorthands for common arithmetic operations.

• In awk, a value is considered to be true if it is nonzero or non-null. Otherwise, the value is false.

• A variable’s type is set upon each assignment and may change over its lifetime. The type determines how it behaves in comparisons (string or numeric).

• Function calls return a value that may be used as part of a larger expression. Expressions used to pass parameter values are fully evaluated before the function is called. awk provides built-in and user-defined functions; this is described in Chapter 9 [Functions], page 175.

• Operator precedence specifies the order in which operations are performed, unless explicitly overridden by parentheses. awk’s operator precedence is compatible with that of C.

• Locales can affect the format of data as output by an awk program, and occasionally the format for data read as input.

Chapter 7: Patterns, Actions, and Variables 131

7 Patterns, Actions, and Variables

As you have already seen, each awk statement consists of a pattern with an associated action. This chapter describes how you build patterns and actions, what kinds of things you can do within actions, and awk’s predefined variables.

The pattern–action rules and the statements available for use within actions form the core of awk programming. In a sense, everything covered up to here has been the foundation that programs are built on top of. Now it’s time to start building something useful.

7.1 Pattern Elements

Patterns in awk control the execution of rules—a rule is executed when its pattern matches the current input record. The following is a summary of the types of awk patterns:

/regular expression/

A regular expression. It matches when the text of the input record fits the regular expression. (See Chapter 3 [Regular Expressions], page 45.)

expression

A single expression. It matches when its value is nonzero (if a number) or non-null (if a string). (See Section 7.1.2 [Expressions as Patterns], page 131.)

begpat, endpat

A pair of patterns separated by a comma, specifying a range of records. The range includes both the initial record that matches begpat and the final record that matches endpat. (See Section 7.1.3 [Specifying Record Ranges with Patterns], page 133.)

BEGIN

END Special patterns for you to supply startup or cleanup actions for your awk program. (See Section 7.1.4 [The BEGIN and END Special Patterns], page 134.)

BEGINFILE

ENDFILE Special patterns for you to supply startup or cleanup actions to be done on a per-file basis. (See Section 7.1.5 [The BEGINFILE and ENDFILE Special Patterns], page 135.)

empty The empty pattern matches every input record. (See Section 7.1.6 [The Empty Pattern], page 136.)

7.1.1 Regular Expressions as Patterns

Regular expressions are one of the first kinds of patterns presented in this book. This kind of pattern is simply a regexp constant in the pattern part of a rule. Its meaning is ‘$0 ~ /pattern/’. The pattern matches when the input record matches the regexp. For example:

     /foo|bar|baz/  { buzzwords++ }
     END            { print buzzwords, "buzzwords seen" }

7.1.2 Expressions as Patterns

Any awk expression is valid as an awk pattern. The pattern matches if the expression’s value is nonzero (if a number) or non-null (if a string). The expression is reevaluated each time the rule is tested against a new input record. If the expression uses fields such as $1, the value depends directly on the new input record’s text; otherwise, it depends on only what has happened so far in the execution of the awk program.

Comparison expressions, using the comparison operators described in Section 6.3.2 [Variable Typing and Comparison Expressions], page 120, are a very common kind of pattern. Regexp matching and nonmatching are also very common expressions. The left operand of the ‘~’ and ‘!~’ operators is a string. The right operand is either a constant regular expression enclosed in slashes (/regexp/), or any expression whose string value is used as a dynamic regular expression (see Section 3.6 [Using Dynamic Regexps], page 52). The following example prints the second field of each input record whose first field is precisely ‘li’:

     $ awk '$1 == "li" { print $2 }' mail-list

(There is no output, because there is no person with the exact name ‘li’.) Contrast this with the following regular expression match, which accepts any record with a first field that contains ‘li’:

     $ awk '$1 ~ /li/ { print $2 }' mail-list
     ⊣ 555-5553
     ⊣ 555-6699

A regexp constant as a pattern is also a special case of an expression pattern. The expression /li/ has the value one if ‘li’ appears in the current input record. Thus, as a pattern, /li/ matches any record containing ‘li’.

Boolean expressions are also commonly used as patterns. Whether the pattern matches an input record depends on whether its subexpressions match. For example, the following command prints all the records in mail-list that contain both ‘edu’ and ‘li’:

     $ awk '/edu/ && /li/' mail-list
     ⊣ Samuel       555-3430     samuel.lanceolis@shu.edu        A

The following command prints all records in mail-list that contain either ‘edu’ or ‘li’ (or both, of course):

     $ awk '/edu/ || /li/' mail-list
     ⊣ Amelia       555-5553     amelia.zodiacusque@gmail.com    F
     ⊣ Broderick    555-0542     broderick.aliquotiens@yahoo.com R
     ⊣ Fabius       555-1234     fabius.undevicesimus@ucb.edu    F
     ⊣ Julie        555-6699     julie.perscrutabor@skeeve.com   F
     ⊣ Samuel       555-3430     samuel.lanceolis@shu.edu        A
     ⊣ Jean-Paul    555-2127     jeanpaul.campanorum@nyu.edu     R

The following command prints all records in mail-list that do not contain the string ‘li’:

     $ awk '! /li/' mail-list
     ⊣ Anthony      555-3412     anthony.asserturo@hotmail.com   A
     ⊣ Becky        555-7685     becky.algebrarum@gmail.com      A
     ⊣ Bill         555-1675     bill.drowning@hotmail.com       A
     ⊣ Camilla      555-2912     camilla.infusarum@skynet.be     R
     ⊣ Fabius       555-1234     fabius.undevicesimus@ucb.edu    F
     ⊣ Martin       555-6480     martin.codicibus@hotmail.com    A

     ⊣ Jean-Paul    555-2127     jeanpaul.campanorum@nyu.edu     R

The subexpressions of a Boolean operator in a pattern can be constant regular expressions, comparisons, or any other awk expressions. Range patterns are not expressions, so they cannot appear inside Boolean patterns. Likewise, the special patterns BEGIN, END, BEGINFILE, and ENDFILE, which never match any input record, are not expressions and cannot appear inside Boolean patterns.

The precedence of the different operators that can appear in patterns is described in Section 6.5 [Operator Precedence (How Operators Nest)], page 128.

7.1.3 Specifying Record Ranges with Patterns

A range pattern is made of two patterns separated by a comma, in the form ‘begpat, endpat’. It is used to match ranges of consecutive input records. The first pattern, begpat, controls where the range begins, while endpat controls where the pattern ends. For example, the following:

     awk '$1 == "on", $1 == "off"' myfile

prints every record in myfile between ‘on’/‘off’ pairs, inclusive.

A range pattern starts out by matching begpat against every input record. When a record matches begpat, the range pattern is turned on, and the range pattern matches this record as well. As long as the range pattern stays turned on, it automatically matches every input record read. The range pattern also matches endpat against every input record; when this succeeds, the range pattern is turned off again for the following record. Then the range pattern goes back to checking begpat against each record.

The record that turns on the range pattern and the one that turns it off both match the range pattern. If you don’t want to operate on these records, you can write if statements in the rule’s action to distinguish them from the records you are interested in.

It is possible for a pattern to be turned on and off by the same record. If the record satisfies both conditions, then the action is executed for just that record. For example, suppose there is text between two identical markers (e.g., the ‘%’ symbol), each on its own line, that should be ignored. A first attempt would be to combine a range pattern that describes the delimited text with the next statement (not discussed yet, see Section 7.4.8 [The next Statement], page 144). This causes awk to skip any further processing of the current record and start over again with the next input record. Such a program looks like this:

     /^%$/,/^%$/    { next }
                    { print }

This program fails because the range pattern is both turned on and turned off by the first line, which just has a ‘%’ on it. To accomplish this task, write the program in the following manner, using a flag:

     /^%$/     { skip = ! skip; next }
     skip == 1 { next } # skip lines with `skip' set

In a range pattern, the comma (‘,’) has the lowest precedence of all the operators (i.e., it is evaluated last). Thus, the following program attempts to combine a range pattern with another, simpler test:

     echo Yes | awk '/1/,/2/ || /Yes/'

The intent of this program is ‘(/1/,/2/) || /Yes/’. However, awk interprets this as ‘/1/, (/2/ || /Yes/)’. This cannot be changed or worked around; range patterns do not combine with other patterns:

     $ echo Yes | gawk '(/1/,/2/) || /Yes/'
      error gawk: cmd. line:1: (/1/,/2/) || /Yes/
      error gawk: cmd. line:1:           ^syntax error

As a minor point of interest, although it is poor style, POSIX allows you to put a newline after the comma in a range pattern.

7.1.4 The BEGIN and END Special Patterns

All the patterns described so far are for matching input records. The BEGIN and END special patterns are different. They supply startup and cleanup actions for awk programs. BEGIN and END rules must have actions; there is no default action for these rules because there is no current record when they run. BEGIN and END rules are often referred to as “BEGIN and END blocks” by longtime awk programmers.

7.1.4.1 Startup and Cleanup Actions

A BEGIN rule is executed once only, before the first input record is read. Likewise, an END rule is executed once only, after all the input is read. For example:

     $ awk '
     > BEGIN { print "Analysis of \"li\"" }
     > /li/  { ++n }
     > END   { print "\"li\" appears in", n, "records." }' mail-list
     ⊣ Analysis of "li"
     ⊣ "li" appears in 4 records.

This program finds the number of records in the input file mail-list that contain the string ‘li’. The BEGIN rule prints a title for the report. There is no need to use the BEGIN rule to initialize the counter n to zero, as awk does this automatically (see Section 6.1.3 [Variables], page 110). The second rule increments the variable n every time a record containing the pattern ‘li’ is read. The END rule prints the value of n at the end of the run.

The special patterns BEGIN and END cannot be used in ranges or with Boolean operators (indeed, they cannot be used with any operators). An awk program may have multiple BEGIN and/or END rules. They are executed in the order in which they appear: all the BEGIN rules at startup and all the END rules at termination. BEGIN and END rules may be intermixed with other rules. This feature was added in the 1987 version of awk and is included in the POSIX standard. The original (1978) version of awk required the BEGIN rule to be placed at the beginning of the program, the END rule to be placed at the end, and only allowed one of each. This is no longer required, but it is a good idea to follow this template in terms of program organization and readability.

Multiple BEGIN and END rules are useful for writing library functions, because each library file can have its own BEGIN and/or END rule to do its own initialization and/or cleanup. The order in which library functions are named on the command line controls the order in which their BEGIN and END rules are executed. Therefore, you have to be careful when writing such rules in library files so that the order in which they are executed doesn’t matter.

See Section 2.2 [Command-Line Options], page 31, for more information on using library functions. See Chapter 10 [A Library of awk Functions], page 217, for a number of useful library functions.

If an awk program has only BEGIN rules and no other rules, then the program exits after the BEGIN rules are run.^{1} However, if an END rule exists, then the input is read, even if there are no other rules in the program. This is necessary in case the END rule checks the FNR and NR variables.

7.1.4.2 Input/Output from BEGIN and END Rules

There are several (sometimes subtle) points to be aware of when doing I/O from a BEGIN or END rule. The first has to do with the value of $0 in a BEGIN rule. Because BEGIN rules are executed before any input is read, there simply is no input record, and therefore no fields, when executing BEGIN rules. References to $0 and the fields yield a null string or zero, depending upon the context. One way to give $0 a real value is to execute a getline command without a variable (see Section 4.9 [Explicit Input with getline], page 78). Another way is simply to assign a value to $0.

The second point is similar to the first, but from the other direction. Traditionally, due largely to implementation issues, $0 and NF were undefined inside an END rule. The POSIX standard specifies that NF is available in an END rule. It contains the number of fields from the last input record. Most probably due to an oversight, the standard does not say that $0 is also preserved, although logically one would think that it should be. In fact, all of BWK awk, mawk, and gawk preserve the value of $0 for use in END rules. Be aware, however, that some other implementations and many older versions of Unix awk do not.

The third point follows from the first two. The meaning of ‘print’ inside a BEGIN or END rule is the same as always: ‘print $0’. If $0 is the null string, then this prints an empty record. Many longtime awk programmers use an unadorned ‘print’ in BEGIN and END rules, to mean ‘print ""’, relying on $0 being null. Although one might generally get away with this in BEGIN rules, it is a very bad idea in END rules, at least in gawk. It is also poor style, because if an empty line is needed in the output, the program should print one explicitly.

Finally, the next and nextfile statements are not allowed in a BEGIN rule, because the implicit read-a-record-and-match-against-the-rules loop has not started yet. Similarly, those statements are not valid in an END rule, because all the input has been read. (See Section 7.4.8 [The next Statement], page 144, and see Section 7.4.9 [The nextfile Statement], page 145,.)

7.1.5 The BEGINFILE and ENDFILE Special Patterns

This section describes a gawk-specific feature.

Two special kinds of rule, BEGINFILE and ENDFILE, give you “hooks” into gawk’s command-line file processing loop. As with the BEGIN and END rules (see the previous section), all BEGINFILE rules in a program are merged, in the order they are read by gawk, and all ENDFILE rules are merged as well.

The body of the BEGINFILE rules is executed just before gawk reads the first record from a file. FILENAME is set to the name of the current file, and FNR is set to zero.

1 The original version of awk kept reading and ignoring input until the end of the file was seen.

The BEGINFILE rule provides you the opportunity to accomplish two tasks that would otherwise be difficult or impossible to perform:

• You can test if the file is readable. Normally, it is a fatal error if a file named on the command line cannot be opened for reading. However, you can bypass the fatal error and move on to the next file on the command line. You do this by checking if the ERRNO variable is not the empty string; if so, then gawk was not able to open the file. In this case, your program can execute the nextfile statement (see Section 7.4.9 [The nextfile Statement], page 145). This causes gawk to skip the file entirely. Otherwise, gawk exits with the usual fatal error.

• If you have written extensions that modify the record handling (by inserting an “input parser”; see Section 16.4.5.4 [Customized Input Parsers], page 365), you can invoke them at this point, before gawk has started processing the file. (This is a very advanced feature, currently used only by the gawkextlib project (http://sourceforge.net/

    projects/gawkextlib).)

The ENDFILE rule is called when gawk has finished processing the last record in an input file. For the last input file, it will be called before any END rules. The ENDFILE rule is executed even for empty input files.

Normally, when an error occurs when reading input in the normal input-processing loop, the error is fatal. However, if an ENDFILE rule is present, the error becomes non-fatal, and instead ERRNO is set. This makes it possible to catch and process I/O errors at the level of the awk program.

The next statement (see Section 7.4.8 [The next Statement], page 144) is not allowed inside either a BEGINFILE or an ENDFILE rule. The nextfile statement is allowed only inside a BEGINFILE rule, not inside an ENDFILE rule.

The getline statement (see Section 4.9 [Explicit Input with getline], page 78) is restricted inside both BEGINFILE and ENDFILE: only redirected forms of getline are allowed.

BEGINFILE and ENDFILE are gawk extensions. In most other awk implementations, or if gawk is in compatibility mode (see Section 2.2 [Command-Line Options], page 31), they are not special.

7.1.6 The Empty Pattern

An empty (i.e., nonexistent) pattern is considered to match every input record. For example, the program:

     awk '{ print $1 }' mail-list

prints the first field of every record.

7.2 Using Shell Variables in Programs

awk programs are often used as components in larger programs written in shell. For example, it is very common to use a shell variable to hold a pattern that the awk program searches for. There are two ways to get the value of the shell variable into the body of the awk program.

A common method is to use shell quoting to substitute the variable’s value into the program inside the script. For example, consider the following program:

     printf "Enter search pattern: "
     read pattern
     awk "/$pattern/ "'{ nmatches++ }
          END { print nmatches, "found" }' /path/to/data

The awk program consists of two pieces of quoted text that are concatenated together to form the program. The first part is double-quoted, which allows substitution of the pattern shell variable inside the quotes. The second part is single-quoted.

Variable substitution via quoting works, but can potentially be messy. It requires a good understanding of the shell’s quoting rules (see Section 1.1.6 [Shell Quoting Issues], page 21), and it’s often difficult to correctly match up the quotes when reading the program.

A better method is to use awk’s variable assignment feature (see Section 6.1.3.2 [Assigning Variables on the Command Line], page 111) to assign the shell variable’s value to an awk variable. Then use dynamic regexps to match the pattern (see Section 3.6 [Using Dynamic Regexps], page 52). The following shows how to redo the previous example using this technique:

     printf "Enter search pattern: "
     read pattern
     awk -v pat="$pattern" '$0 ~ pat { nmatches++ }
            END { print nmatches, "found" }' /path/to/data

Now, the awk program is just one single-quoted string. The assignment ‘-v pat="$pattern"’ still requires double quotes, in case there is whitespace in the value of $pattern. The awk variable pat could be named pattern too, but that would be more confusing. Using a variable also provides more flexibility, as the variable can be used anywhere inside the program—for printing, as an array subscript, or for any other use—without requiring the quoting tricks at every point in the program.

7.3 Actions

An awk program or script consists of a series of rules and function definitions interspersed. (Functions are described later. See Section 9.2 [User-Defined Functions], page 199.) A rule contains a pattern and an action, either of which (but not both) may be omitted. The purpose of the action is to tell awk what to do once a match for the pattern is found. Thus, in outline, an awk program generally looks like this:

[pattern] { action } pattern [{ action }] …

     function name(args) { ... }

… An action consists of one or more awk statements, enclosed in braces (‘{…}’). Each

statement specifies one thing to do. The statements are separated by newlines or semicolons. The braces around an action must be used even if the action contains only one statement, or if it contains no statements at all. However, if you omit the action entirely, omit the braces as well. An omitted action is equivalent to ‘{ print $0 }’:

/foo/ { } match foo, do nothing — empty action /foo/ match foo, print the record — omitted action

The following types of statements are supported in awk:

Expressions Call functions or assign values to variables (see Chapter 6 [Expressions], page 107). Executing this kind of statement simply computes the value of the expression. This is useful when the expression has side effects (see Section 6.2.3 [Assignment Expressions], page 116).

Control statements Specify the control flow of awk programs. The awk language gives you C-like constructs (if, for, while, and do) as well as a few special ones (see Section 7.4 [Control Statements in Actions], page 138).

Compound statements Enclose one or more statements in braces. A compound statement is used in order to put several statements together in the body of an if, while, do, or for statement.

Input statements Use the getline command (see Section 4.9 [Explicit Input with getline], page 78). Also supplied in awk are the next statement (see Section 7.4.8 [The next Statement], page 144) and the nextfile statement (see Section 7.4.9 [The nextfile Statement], page 145).

Output statements Such as print and printf. See Chapter 5 [Printing Output], page 89.

Deletion statements For deleting array elements. See Section 8.4 [The delete Statement], page 168.

7.4 Control Statements in Actions

Control statements, such as if, while, and so on, control the flow of execution in awk programs. Most of awk’s control statements are patterned after similar statements in C.

All the control statements start with special keywords, such as if and while, to distinguish them from simple expressions. Many control statements contain other statements. For example, the if statement contains another statement that may or may not be executed. The contained statement is called the body. To include more than one statement in the body, group them into a single compound statement with braces, separating them with newlines or semicolons.

7.4.1 The if-else Statement

The if-else statement is awk’s decision-making statement. It looks like this:

     if (condition) then-body [else else-body]

The condition is an expression that controls what the rest of the statement does. If the condition is true, then-body is executed; otherwise, else-body is executed. The else part of the statement is optional. The condition is considered false if its value is zero or the null string; otherwise, the condition is true. Refer to the following:

     if (x % 2 == 0)

         print "x is even"
     else
         print "x is odd"

In this example, if the expression ‘x % 2 == 0’ is true (i.e., if the value of x is evenly divisible by two), then the first print statement is executed; otherwise, the second print statement is executed. If the else keyword appears on the same line as then-body and then-body is not a compound statement (i.e., not surrounded by braces), then a semicolon must separate then-body from the else. To illustrate this, the previous example can be rewritten as:

     if (x % 2 == 0) print "x is even"; else
             print "x is odd"

If the ‘;’ is left out, awk can’t interpret the statement and it produces a syntax error. Don’t actually write programs this way, because a human reader might fail to see the else if it is not the first thing on its line.

7.4.2 The while Statement

In programming, a loop is a part of a program that can be executed two or more times in succession. The while statement is the simplest looping statement in awk. It repeatedly executes a statement as long as a condition is true. For example:

     while (condition)
       body

body is a statement called the body of the loop, and condition is an expression that controls how long the loop keeps running. The first thing the while statement does is test the condition. If the condition is true, it executes the statement body. After body has been executed, condition is tested again, and if it is still true, body executes again. This process repeats until the condition is no longer true. If the condition is initially false, the body of the loop never executes and awk continues with the statement following the loop. This example prints the first three fields of each record, one per line:

     awk '
     {
         i = 1
         while (i <= 3) {
             print $i
             i++
         }
     }' inventory-shipped

The body of this loop is a compound statement enclosed in braces, containing two statements. The loop works in the following manner: first, the value of i is set to one. Then, the while statement tests whether i is less than or equal to three. This is true when i equals one, so the ith field is printed. Then the ‘i++’ increments the value of i and the loop repeats. The loop terminates when i reaches four.

A newline is not required between the condition and the body; however, using one makes the program clearer unless the body is a compound statement or else is very simple. The newline after the open brace that begins the compound statement is not required either, but the program is harder to read without it.

7.4.3 The do-while Statement

The do loop is a variation of the while looping statement. The do loop executes the body once and then repeats the body as long as the condition is true. It looks like this:

     do
       body
     while (condition)

Even if the condition is false at the start, the body executes at least once (and only once, unless executing body makes condition true). Contrast this with the corresponding while statement:

     while (condition)
         body

This statement does not execute the body even once if the condition is false to begin with. The following is an example of a do statement:

     {
         i = 1
         do {
             print $0
             i++
         } while (i <= 10)
     }

This program prints each input record 10 times. However, it isn’t a very realistic example, because in this case an ordinary while would do just as well. This situation reflects actual experience; only occasionally is there a real use for a do statement.

7.4.4 The for Statement

The for statement makes it more convenient to count iterations of a loop. The general form of the for statement looks like this:

     for (initialization; condition; increment)
       body

The initialization, condition, and increment parts are arbitrary awk expressions, and body stands for any awk statement.

The for statement starts by executing initialization. Then, as long as the condition is true, it repeatedly executes body and then increment. Typically, initialization sets a variable to either zero or one, increment adds one to it, and condition compares it against the desired number of iterations. For example:

     awk '
     {
         for (i = 1; i <= 3; i++)
             print $i
     }' inventory-shipped

This prints the first three fields of each input record, with one field per line.

It isn’t possible to set more than one variable in the initialization part without using a multiple assignment statement such as ‘x = y = 0’. This makes sense only if all the initial values are equal. (But it is possible to initialize additional variables by writing their assignments as separate statements preceding the for loop.)

The same is true of the increment part. Incrementing additional variables requires separate statements at the end of the loop. The C compound expression, using C’s comma operator, is useful in this context, but it is not supported in awk.

Most often, increment is an increment expression, as in the previous example. But this is not required; it can be any expression whatsoever. For example, the following statement prints all the powers of two between 1 and 100:

     for (i = 1; i <= 100; i *= 2)
         print i

If there is nothing to be done, any of the three expressions in the parentheses following the for keyword may be omitted. Thus, ‘for (; x > 0;)’ is equivalent to ‘while (x > 0)’. If the condition is omitted, it is treated as true, effectively yielding an infinite loop (i.e., a loop that never terminates).

In most cases, a for loop is an abbreviation for a while loop, as shown here:

     initialization
     while (condition) {
       body
       increment
     }

The only exception is when the continue statement (see Section 7.4.7 [The continue Statement], page 143) is used inside the loop. Changing a for statement to a while statement in this way can change the effect of the continue statement inside the loop.

The awk language has a for statement in addition to a while statement because a for loop is often both less work to type and more natural to think of. Counting the number of iterations is very common in loops. It can be easier to think of this counting as part of looping rather than as something to do inside the loop.

There is an alternative version of the for loop, for iterating over all the indices of an array:

     for (i in array)
         do something with array[i]

See Section 8.1.5 [Scanning All Elements of an Array], page 163, for more information on this version of the for loop.

7.4.5 The switch Statement

This section describes a gawk-specific feature. If gawk is in compatibility mode (see Section 2.2 [Command-Line Options], page 31), it is not available.

The switch statement allows the evaluation of an expression and the execution of statements based on a case match. Case statements are checked for a match in the order they are defined. If no suitable case is found, the default section is executed, if supplied.

Each case contains a single constant, be it numeric, string, or regexp. The switch expression is evaluated, and then each case’s constant is compared against the result in turn. The type of constant determines the comparison: numeric or string do the usual comparisons. A regexp constant does a regular expression match against the string value of the original expression. The general form of the switch statement looks like this:

switch (expression) {
case value or regular expression:
case-body
default:
default-body
}

Control flow in the switch statement works as it does in C. Once a match to a given case is made, the case statement bodies execute until a break, continue, next, nextfile, or exit is encountered, or the end of the switch statement itself. For example:

while ((c = getopt(ARGC, ARGV, "aksx")) != -1) {
switch (c) {
case "a":
# report size of all files
all_files = TRUE;
break
case "k":
BLOCK_SIZE = 1024       # 1K block size
break
case "s":
# do sums only
sum_only = TRUE
break
case "x":
# don't cross filesystems
fts_flags = or(fts_flags, FTS_XDEV)
break
case "?":
default:
usage()
break
}
}

Note that if none of the statements specified here halt execution of a matched case statement, execution falls through to the next case until execution halts. In this example, the case for "?" falls through to the default case, which is to call a function named usage(). (The getopt() function being called here is described in Section 10.4 [Processing Command-Line Options], page 233.)

7.4.6 The break Statement

The break statement jumps out of the innermost for, while, or do loop that encloses it. The following example finds the smallest divisor of any integer, and also identifies prime numbers:

# find smallest divisor of num
{

         num = $1
         for (divisor = 2; divisor * divisor <= num; divisor++) {
             if (num % divisor == 0)
                 break
         }
         if (num % divisor == 0)
             printf "Smallest divisor of %d is %d\n", num, divisor
         else
             printf "%d is prime\n", num
     }

When the remainder is zero in the first if statement, awk immediately breaks out of the containing for loop. This means that awk proceeds immediately to the statement following the loop and continues processing. (This is very different from the exit statement, which stops the entire awk program. See Section 7.4.10 [The exit Statement], page 146.)

The following program illustrates how the condition of a for or while statement could be replaced with a break inside an if:

     # find smallest divisor of num
     {
         num = $1
         for (divisor = 2; ; divisor++) {
             if (num % divisor == 0) {
                 printf "Smallest divisor of %d is %d\n", num, divisor
                 break
             }
             if (divisor * divisor > num) {
                 printf "%d is prime\n", num
                 break
             }
         }
     }

The break statement is also used to break out of the switch statement. This is discussed in Section 7.4.5 [The switch Statement], page 141.

The break statement has no meaning when used outside the body of a loop or switch. However, although it was never documented, historical implementations of awk treated the break statement outside of a loop as if it were a next statement (see Section 7.4.8 [The next Statement], page 144). Recent versions of BWK awk no longer allow this usage, nor does gawk.

7.4.7 The continue Statement

Similar to break, the continue statement is used only inside for, while, and do loops. It skips over the rest of the loop body, causing the next cycle around the loop to begin immediately. Contrast this with break, which jumps out of the loop altogether.

The continue statement in a for loop directs awk to skip the rest of the body of the loop and resume execution with the increment-expression of the for statement. The following program illustrates this fact:

     BEGIN {
          for (x = 0; x <= 20; x++) {
              if (x == 5)
                  continue
              printf "%d ", x
          }
          print ""
     }

This program prints all the numbers from 0 to 20—except for 5, for which the printf is skipped. Because the increment ‘x++’ is not skipped, x does not remain stuck at 5. Contrast the for loop from the previous example with the following while loop:

     BEGIN {
          x = 0
          while (x <= 20) {
              if (x == 5)
                  continue
              printf "%d ", x
              x++
          }
          print ""
     }

This program loops forever once x reaches 5, because the increment (‘x++’) is never reached.

The continue statement has no special meaning with respect to the switch statement, nor does it have any meaning when used outside the body of a loop. Historical versions of awk treated a continue statement outside a loop the same way they treated a break statement outside a loop: as if it were a next statement (see Section 7.4.8 [The next Statement], page 144). Recent versions of BWK awk no longer work this way, nor does

gawk.

7.4.8 The next Statement

The next statement forces awk to immediately stop processing the current record and go on to the next record. This means that no further rules are executed for the current record, and the rest of the current rule’s action isn’t executed.

Contrast this with the effect of the getline function (see Section 4.9 [Explicit Input with getline], page 78). That also causes awk to read the next record immediately, but it does not alter the flow of control in any way (i.e., the rest of the current action executes with a new input record).

At the highest level, awk program execution is a loop that reads an input record and then tests each rule’s pattern against it. If you think of this loop as a for statement whose body contains the rules, then the next statement is analogous to a continue statement. It skips to the end of the body of this implicit loop and executes the increment (which reads another record).

For example, suppose an awk program works only on records with four fields, and it shouldn’t fail when given bad input. To avoid complicating the rest of the program, write a “weed out” rule near the beginning, in the following manner:

     NF != 4 {
         printf("%s:%d: skipped: NF != 4\n", FILENAME, FNR) > "/dev/stderr"
         next
     }

Because of the next statement, the program’s subsequent rules won’t see the bad record. The error message is redirected to the standard error output stream, as error messages should be. For more detail, see Section 5.8 [Special File names in gawk], page 101.

If the next statement causes the end of the input to be reached, then the code in any END rules is executed. See Section 7.1.4 [The BEGIN and END Special Patterns], page 134.

The next statement is not allowed inside BEGINFILE and ENDFILE rules. See Section 7.1.5 [The BEGINFILE and ENDFILE Special Patterns], page 135.

According to the POSIX standard, the behavior is undefined if the next statement is used in a BEGIN or END rule. gawk treats it as a syntax error. Although POSIX does not disallow it, most other awk implementations don’t allow the next statement inside function bodies (see Section 9.2 [User-Defined Functions], page 199). Just as with any other next statement, a next statement inside a function body reads the next record and starts processing it with the first rule in the program.

7.4.9 The nextfile Statement

The nextfile statement is similar to the next statement. However, instead of abandoning processing of the current record, the nextfile statement instructs awk to stop processing the current data file.

Upon execution of the nextfile statement, FILENAME is updated to the name of the next data file listed on the command line, FNR is reset to one, and processing starts over with the first rule in the program. If the nextfile statement causes the end of the input to be reached, then the code in any END rules is executed. An exception to this is when nextfile is invoked during execution of any statement in an END rule; in this case, it causes the program to stop immediately. See Section 7.1.4 [The BEGIN and END Special Patterns], page 134.

The nextfile statement is useful when there are many data files to process but it isn’t necessary to process every record in every file. Without nextfile, in order to move on to the next data file, a program would have to continue scanning the unwanted records. The nextfile statement accomplishes this much more efficiently.

In gawk, execution of nextfile causes additional things to happen: any ENDFILE rules are executed if gawk is not currently in an END or BEGINFILE rule, ARGIND is incremented, and any BEGINFILE rules are executed. (ARGIND hasn’t been introduced yet. See Section 7.5 [Predefined Variables], page 147.)

With gawk, nextfile is useful inside a BEGINFILE rule to skip over a file that would otherwise cause gawk to exit with a fatal error. In this case, ENDFILE rules are not executed. See Section 7.1.5 [The BEGINFILE and ENDFILE Special Patterns], page 135.

Although it might seem that ‘close(FILENAME)’ would accomplish the same as nextfile, this isn’t true. close() is reserved for closing files, pipes, and coprocesses that are opened with redirections. It is not related to the main processing that awk does with the files listed in ARGV.

NOTE: For many years, nextfile was a common extension. In September 2012, it was accepted for inclusion into the POSIX standard. See the Austin

     Group website (http://austingroupbugs.net/view.php?id=607).

The current version of BWK awk and mawk also support nextfile. However, they don’t allow the nextfile statement inside function bodies (see Section 9.2 [User-Defined Functions], page 199). gawk does; a nextfile inside a function body reads the next record and starts processing it with the first rule in the program, just as any other nextfile statement.

7.4.10 The exit Statement

The exit statement causes awk to immediately stop executing the current rule and to stop processing input; any remaining input is ignored. The exit statement is written as follows:

exit [return code] When an exit statement is executed from a BEGIN rule, the program stops processing

everything immediately. No input records are read. However, if an END rule is present, as part of executing the exit statement, the END rule is executed (see Section 7.1.4 [The BEGIN and END Special Patterns], page 134). If exit is used in the body of an END rule, it causes the program to stop immediately.

An exit statement that is not part of a BEGIN or END rule stops the execution of any further automatic rules for the current record, skips reading any remaining input records, and executes the END rule if there is one. gawk also skips any ENDFILE rules; they do not execute.

In such a case, if you don’t want the END rule to do its job, set a variable to a nonzero value before the exit statement and check that variable in the END rule. See Section 10.2.2 [Assertions], page 220, for an example that does this.

If an argument is supplied to exit, its value is used as the exit status code for the awk process. If no argument is supplied, exit causes awk to return a “success” status. In the case where an argument is supplied to a first exit statement, and then exit is called a second time from an END rule with no argument, awk uses the previously supplied exit value. See Section 2.6 [gawk’s Exit Status], page 41, for more information.

For example, suppose an error condition occurs that is difficult or impossible to handle. Conventionally, programs report this by exiting with a nonzero status. An awk program can do this using an exit statement with a nonzero argument, as shown in the following example:

     BEGIN {
         if (("date" | getline date_now) <= 0) {
             print "Can't get system date" > "/dev/stderr"
             exit 1
         }
         print "current date is", date_now
         close("date")
     }

NOTE: For full portability, exit values should be between zero and 126, inclusive. Negative values, and values of 127 or greater, may not produce consistent results across different operating systems.

7.5 Predefined Variables

Most awk variables are available to use for your own purposes; they never change unless your program assigns values to them, and they never affect anything unless your program examines them. However, a few variables in awk have special built-in meanings. awk examines some of these automatically, so that they enable you to tell awk how to do certain things. Others are set automatically by awk, so that they carry information from the internal workings of awk to your program.

This section documents all of gawk’s predefined variables, most of which are also documented in the chapters describing their areas of activity.

7.5.1 Built-in Variables That Control awk

The following is an alphabetical list of variables that you can change to control how awk does certain things.

The variables that are specific to gawk are marked with a pound sign (‘#’). These variables are gawk extensions. In other awk implementations or if gawk is in compatibility mode (see Section 2.2 [Command-Line Options], page 31), they are not special. (Any exceptions are noted in the description of each variable.)

BINMODE # On non-POSIX systems, this variable specifies use of binary mode for all I/O. Numeric values of one, two, or three specify that input files, output files, or all files, respectively, should use binary I/O. A numeric value less than zero is treated as zero, and a numeric value greater than three is treated as three. Alternatively, string values of "r" or "w" specify that input files and output files, respectively, should use binary I/O. A string value of "rw" or "wr" indicates that all files should use binary I/O. Any other string value is treated the same as "rw", but causes gawk to generate a warning message. BINMODE is described in more detail in Section B.3.1.4 [Using gawk on PC Operating Systems], page 437. mawk (see Section B.5 [Other Freely Available awk Implementations], page 443) also supports this variable, but only using numeric values.

CONVFMT A string that controls the conversion of numbers to strings (see Section 6.1.4 [Conversion of Strings and Numbers], page 111). It works by being passed, in effect, as the first argument to the sprintf() function (see Section 9.1.3 [String-Manipulation Functions], page 177). Its default value is "%.6g". CONVFMT was introduced by the POSIX standard.

FIELDWIDTHS #

A space-separated list of columns that tells gawk how to split input with fixed columnar boundaries. Assigning a value to FIELDWIDTHS overrides the use of FS and FPAT for field splitting. See Section 4.6 [Reading Fixed-Width Data], page 72, for more information.

FPAT # A regular expression (as a string) that tells gawk to create the fields based on text that matches the regular expression. Assigning a value to FPAT overrides the use of FS and FIELDWIDTHS for field splitting. See Section 4.7 [Defining Fields by Content], page 74, for more information.

FS The input field separator (see Section 4.5 [Specifying How Fields Are Separated], page 67). The value is a single-character string or a multicharacter regular expression that matches the separations between fields in an input record. If the value is the null string (""), then each character in the record becomes a separate field. (This behavior is a gawk extension. POSIX awk does not specify the behavior when FS is the null string. Nonetheless, some other versions of awk also treat "" specially.) The default value is " ", a string consisting of a single space. As a special exception, this value means that any sequence of spaces, TABs, and/or newlines is a single separator.^{2} It also causes spaces, TABs, and newlines at the beginning and end of a record to be ignored. You can set the value of FS on the command line using the -F option:

     awk -F, 'program' input-files

If gawk is using FIELDWIDTHS or FPAT for field splitting, assigning a value to FS causes gawk to return to the normal, FS-based field splitting. An easy way to do this is to simply say ‘FS = FS’, perhaps with an explanatory comment.

IGNORECASE #

If IGNORECASE is nonzero or non-null, then all string comparisons and all regular expression matching are case-independent. This applies to regexp match-

ing with ‘~’ and ‘!~’, the gensub(), gsub(), index(), match(), patsplit(),

split(), and sub() functions, record termination with RS, and field splitting with FS and FPAT. However, the value of IGNORECASE does not affect array subscripting and it does not affect field splitting when using a single-character field separator. See Section 3.8 [Case Sensitivity in Matching], page 55.

LINT # When this variable is true (nonzero or non-null), gawk behaves as if the --lint command-line option is in effect (see Section 2.2 [Command-Line Options], page 31). With a value of "fatal", lint warnings become fatal errors. With a value of "invalid", only warnings about things that are actually invalid are issued. (This is not fully implemented yet.) Any other true value prints nonfatal warnings. Assigning a false value to LINT turns off the lint warnings. This variable is a gawk extension. It is not special in other awk implementations. Unlike with the other special variables, changing LINT does affect the production of lint warnings, even if gawk is in compatibility mode. Much as the --lint and --traditional options independently control different aspects of gawk’s behavior, the control of lint warnings during program execution is independent of the flavor of awk being executed.

OFMT A string that controls conversion of numbers to strings (see Section 6.1.4 [Conversion of Strings and Numbers], page 111) for printing with the print statement. It works by being passed as the first argument to the sprintf() function (see Section 9.1.3 [String-Manipulation Functions], page 177). Its default value is "%.6g". Earlier versions of awk used OFMT to specify the format for converting numbers to strings in general expressions; this is now done by CONVFMT.

OFS The output field separator (see Section 5.3 [Output Separators], page 91). It is output between the fields printed by a print statement. Its default value is " ", a string consisting of a single space.

2 In POSIX awk, newline does not count as whitespace.

ORS The output record separator. It is output at the end of every print statement. Its default value is "\n", the newline character. (See Section 5.3 [Output Separators], page 91.)

PREC # The working precision of arbitrary-precision floating-point numbers, 53 bits by default (see Section 15.4.4 [Setting the Precision], page 348).

ROUNDMODE #

The rounding mode to use for arbitrary-precision arithmetic on numbers, by default "N" (roundTiesToEven in the IEEE 754 standard; see Section 15.4.5 [Setting the Rounding Mode], page 349).

RS The input record separator. Its default value is a string containing a single newline character, which means that an input record consists of a single line of text. It can also be the null string, in which case records are separated by runs of blank lines. If it is a regexp, records are separated by matches of the regexp in the input text. (See Section 4.1 [How Input Is Split into Records], page 59.) The ability for RS to be a regular expression is a gawk extension. In most other awk implementations, or if gawk is in compatibility mode (see Section 2.2 [Command-Line Options], page 31), just the first character of RS’s value is used.

SUBSEP The subscript separator. It has the default value of "\034" and is used to separate the parts of the indices of a multidimensional array. Thus, the expression ‘foo["A", "B"]’ really accesses foo["A\034B"] (see Section 8.5 [Multidimensional Arrays], page 169).

TEXTDOMAIN #

Used for internationalization of programs at the awk level. It sets the default text domain for specially marked string constants in the source text, as well

as for the dcgettext(), dcngettext(), and bindtextdomain() functions (see

Chapter 13 [Internationalization with gawk], page 315). The default value of

TEXTDOMAIN is "messages".

7.5.2 Built-in Variables That Convey Information

The following is an alphabetical list of variables that awk sets automatically on certain occasions in order to provide information to your program.

The variables that are specific to gawk are marked with a pound sign (‘#’). These variables are gawk extensions. In other awk implementations or if gawk is in compatibility mode (see Section 2.2 [Command-Line Options], page 31), they are not special:

ARGC, ARGV

The command-line arguments available to awk programs are stored in an array called ARGV. ARGC is the number of command-line arguments present. See Section 2.3 [Other Command-Line Arguments], page 37. Unlike most awk arrays, ARGV is indexed from 0 to ARGC − 1. In the following example:

     $ awk 'BEGIN {
     >         for (i = 0; i < ARGC; i++)
     >             print ARGV[i]
     >      }' inventory-shipped mail-list

     ⊣ awk
     ⊣ inventory-shipped
     ⊣ mail-list
ARGV[0] contains ‘awk’, ARGV[1] contains ‘inventory-shipped’, and ARGV[2]

contains ‘mail-list’. The value of ARGC is three, one more than the index of the last element in ARGV, because the elements are numbered from zero. The names ARGC and ARGV, as well as the convention of indexing the array from 0 to ARGC − 1, are derived from the C language’s method of accessing command-line arguments. The value of ARGV[0] can vary from system to system. Also, you should note that the program text is not included in ARGV, nor are any of awk’s commandline options. See Section 7.5.3 [Using ARGC and ARGV], page 155, for information about how awk uses these variables.

ARGIND # The index in ARGV of the current file being processed. Every time gawk opens a new data file for processing, it sets ARGIND to the index in ARGV of the file name. When gawk is processing the input files, ‘FILENAME == ARGV[ARGIND]’ is always true. This variable is useful in file processing; it allows you to tell how far along you are in the list of data files as well as to distinguish between successive instances of the same file name on the command line. While you can change the value of ARGIND within your awk program, gawk automatically sets it to a new value when it opens the next file.

ENVIRON An associative array containing the values of the environment. The array indices are the environment variable names; the elements are the values of the particular environment variables. For example, ENVIRON["HOME"] might be "/home/arnold". Changing this array does not affect the environment passed on to any programs that awk may spawn via redirection or the system() function. (In a future version of gawk, it may do so.) Some operating systems may not have environment variables. On such systems, the ENVIRON array is empty (except for ENVIRON["AWKPATH"] and ENVIRON["AWKLIBPATH"]; see Section 2.5.1 [The AWKPATH Environment Variable], page 39, and see Section 2.5.2 [The AWKLIBPATH Environment Variable], page 40).

ERRNO # If a system error occurs during a redirection for getline, during a read for getline, or during a close() operation, then ERRNO contains a string describing the error. In addition, gawk clears ERRNO before opening each command-line input file. This enables checking if the file is readable inside a BEGINFILE pattern (see Section 7.1.5 [The BEGINFILE and ENDFILE Special Patterns], page 135). Otherwise, ERRNO works similarly to the C variable errno. Except for the case just mentioned, gawk never clears it (sets it to zero or ""). Thus, you should only expect its value to be meaningful when an I/O operation returns a failure value, such as getline returning −1. You are, of course, free to clear it yourself before doing an I/O operation.

FILENAME The name of the current input file. When no data files are listed on the command line, awk reads from the standard input and FILENAME is set to "-". FILENAME changes each time a new file is read (see Chapter 4 [Reading Input Files], page 59). Inside a BEGIN rule, the value of FILENAME is "", because there are no input files being processed yet.^{3} Note, though, that using getline (see Section 4.9 [Explicit Input with getline], page 78) inside a BEGIN rule can give

FILENAME a value.

FNR The current record number in the current file. awk increments FNR each time it reads a new record (see Section 4.1 [How Input Is Split into Records], page 59). awk resets FNR to zero each time it starts a new input file.

NF The number of fields in the current input record. NF is set each time a new record is read, when a new field is created, or when $0 changes (see Section 4.2 [Examining Fields], page 63). Unlike most of the variables described in this subsection, assigning a value to NF has the potential to affect awk’s internal workings. In particular, assignments to NF can be used to create fields in or remove fields from the current record. See Section 4.4 [Changing the Contents of a Field], page 65.

FUNCTAB # An array whose indices and corresponding values are the names of all the built-

in, user-defined, and extension functions in the program.

NOTE: Attempting to use the delete statement with the FUNCTAB array causes a fatal error. Any attempt to assign to an element of FUNCTAB also causes a fatal error.

NR The number of input records awk has processed since the beginning of the program’s execution (see Section 4.1 [How Input Is Split into Records], page 59). awk increments NR each time it reads a new record.

PROCINFO #

The elements of this array provide access to information about the running awk program. The following elements (listed alphabetically) are guaranteed to be available:

PROCINFO["egid"]

The value of the getegid() system call.

PROCINFO["euid"]

The value of the geteuid() system call.

PROCINFO["FS"]

This is "FS" if field splitting with FS is in effect, "FIELDWIDTHS" if field splitting with FIELDWIDTHS is in effect, or "FPAT" if field matching with FPAT is in effect.

PROCINFO["identifiers"]

A subarray, indexed by the names of all identifiers used in the text of the awk program. An identifier is simply the name of a variable (be it scalar or array), built-in function, user-defined function, or extension function. For each identifier, the value of the element is one of the following:

3 Some early implementations of Unix awk initialized FILENAME to "-", even if there were data files to be

processed. This behavior was incorrect and should not be relied upon in your programs.

"array" The identifier is an array.

"builtin"

The identifier is a built-in function.

"extension"

The identifier is an extension function loaded via @load or -l.

"scalar" The identifier is a scalar.

"untyped"

The identifier is untyped (could be used as a scalar or an array; gawk doesn’t know yet).

"user" The identifier is a user-defined function.

The values indicate what gawk knows about the identifiers after it has finished parsing the program; they are not updated while the program runs.

PROCINFO["gid"]

The value of the getgid() system call.

PROCINFO["pgrpid"]

The process group ID of the current process.

PROCINFO["pid"]

The process ID of the current process.

PROCINFO["ppid"]

The parent process ID of the current process.

PROCINFO["sorted_in"]

If this element exists in PROCINFO, its value controls the order in which array indices will be processed by ‘for (indx in array)’ loops. This is an advanced feature, so we defer the full description until later; see Section 8.1.5 [Scanning All Elements of an Array], page 163.

PROCINFO["strftime"]

The default time format string for strftime(). Assigning a new value to this element changes the default. See Section 9.1.5 [Time Functions], page 192.

PROCINFO["uid"]

The value of the getuid() system call.

PROCINFO["version"]

The version of gawk.

The following additional elements in the array are available to provide information about the MPFR and GMP libraries if your version of gawk supports arbitrary-precision arithmetic (see Chapter 15 [Arithmetic and Arbitrary-Precision Arithmetic with gawk], page 341):

PROCINFO["mpfr_version"]

The version of the GNU MPFR library.

PROCINFO["gmp_version"]

The version of the GNU MP library.

PROCINFO["prec_max"]

The maximum precision supported by MPFR.

PROCINFO["prec_min"]

The minimum precision required by MPFR.

The following additional elements in the array are available to provide information about the version of the extension API, if your version of gawk supports dynamic loading of extension functions (see Chapter 16 [Writing Extensions for gawk], page 355):

PROCINFO["api_major"]

The major version of the extension API.

PROCINFO["api_minor"]

The minor version of the extension API.

On some systems, there may be elements in the array, "group1" through "groupN" for some N. N is the number of supplementary groups that the process has. Use the in operator to test for these elements (see Section 8.1.2 [Referring to an Array Element], page 161). The PROCINFO array has the following additional uses:

• It may be used to provide a timeout when reading from any open input file, pipe, or coprocess. See Section 4.10 [Reading Input with a Timeout], page 84, for more information.

• It may be used to cause coprocesses to communicate over pseudo-ttys instead of through two-way pipes; this is discussed further in Section 12.3 [Two-Way Communications with Another Process], page 305.

RLENGTH The length of the substring matched by the match() function (see Section 9.1.3 [String-Manipulation Functions], page 177). RLENGTH is set by invoking the match() function. Its value is the length of the matched string, or −1 if no match is found.

RSTART The start index in characters of the substring that is matched by the match() function (see Section 9.1.3 [String-Manipulation Functions], page 177). RSTART is set by invoking the match() function. Its value is the position of the string where the matched substring starts, or zero if no match was found.

RT # The input text that matched the text denoted by RS, the record separator. It is set every time a record is read.

SYMTAB # An array whose indices are the names of all defined global variables and arrays in the program. SYMTAB makes gawk’s symbol table visible to the awk programmer. It is built as gawk parses the program and is complete before the program starts to run.

The array may be used for indirect access to read or write the value of a variable:

     foo = 5
     SYMTAB["foo"] = 4
     print foo    # prints 4

The isarray() function (see Section 9.1.7 [Getting Type Information], page 198) may be used to test if an element in SYMTAB is an array. Also, you may not use the delete statement with the SYMTAB array.

You may use an index for SYMTAB that is not a predefined identifier:

     SYMTAB["xxx"] = 5
     print SYMTAB["xxx"]

This works as expected: in this case SYMTAB acts just like a regular array. The only difference is that you can’t then delete SYMTAB["xxx"].

The SYMTAB array is more interesting than it looks. Andrew Schorr points out that it effectively gives awk data pointers. Consider his example:

     # Indirect multiply of any variable by amount, return result
     function multiply(variable, amount)
     {
         return SYMTAB[variable] *= amount
     }

NOTE: In order to avoid severe time-travel paradoxes,^{4} neither FUNCTAB nor SYMTAB is available as an element within the SYMTAB array.

4 Not to mention difficult implementation issues.

Changing NR and FNR awk increments NR and FNR each time it reads a record, instead of setting them to the

absolute value of the number of records read. This means that a program can change these variables and their new values are incremented for each record. The following example shows this:

      $ echo '1
      > 2
      > 3
      > 4' | awk 'NR == 2 { NR = 17 }
      > { print NR }'
      ⊣ 1
      ⊣ 17
      ⊣ 18
      ⊣ 19

Before FNR was added to the awk language (see Section A.1 [Major Changes Between V7 and SVR3.1], page 413), many awk programs used this feature to track the number of records in a file by resetting NR to zero when FILENAME changed.

7.5.3 Using ARGC and ARGV

Section 7.5.2 [Built-in Variables That Convey Information], page 149, presented the following program describing the information contained in ARGC and ARGV:

     $ awk 'BEGIN {
     >        for (i = 0; i < ARGC; i++)
     >            print ARGV[i]
     >      }' inventory-shipped mail-list
     ⊣ awk
     ⊣ inventory-shipped
     ⊣ mail-list

In this example, ARGV[0] contains ‘awk’, ARGV[1] contains ‘inventory-shipped’, and ARGV[2] contains ‘mail-list’. Notice that the awk program is not entered in ARGV. The other command-line options, with their arguments, are also not entered. This includes variable assignments done with the -v option (see Section 2.2 [Command-Line Options], page 31). Normal variable assignments on the command line are treated as arguments and do show up in the ARGV array. Given the following program in a file named showargs.awk:

     BEGIN {
         printf "A=%d, B=%d\n", A, B
         for (i = 0; i < ARGC; i++)
             printf "\tARGV[%d] = %s\n", i, ARGV[i]
     }
     END   { printf "A=%d, B=%d\n", A, B }

Running it produces the following:

     $ awk -v A=1 -f showargs.awk B=2 /dev/null
     ⊣ A=1, B=0
     ⊣  ARGV[0] = awk
     ⊣  ARGV[1] = B=2

     ⊣  ARGV[2] = /dev/null
     ⊣ A=1, B=2

A program can alter ARGC and the elements of ARGV. Each time awk reaches the end of an input file, it uses the next element of ARGV as the name of the next input file. By storing a different string there, a program can change which files are read. Use "-" to represent the standard input. Storing additional elements and incrementing ARGC causes additional files to be read.

If the value of ARGC is decreased, that eliminates input files from the end of the list. By recording the old value of ARGC elsewhere, a program can treat the eliminated arguments as something other than file names.

To eliminate a file from the middle of the list, store the null string ("") into ARGV in place of the file’s name. As a special feature, awk ignores file names that have been replaced with the null string. Another option is to use the delete statement to remove elements from ARGV (see Section 8.4 [The delete Statement], page 168).

All of these actions are typically done in the BEGIN rule, before actual processing of the input begins. See Section 11.2.4 [Splitting a Large File into Pieces], page 262, and see Section 11.2.5 [Duplicating Output into Multiple Files], page 263, for examples of each way of removing elements from ARGV.

To actually get options into an awk program, end the awk options with -- and then supply the awk program’s options, in the following manner:

     awk -f myprog.awk -- -v -q file1 file2 ...

The following fragment processes ARGV in order to examine, and then remove, the previously mentioned command-line options:

     BEGIN {
         for (i = 1; i < ARGC; i++) {
             if (ARGV[i] == "-v")
                 verbose = 1
             else if (ARGV[i] == "-q")
                 debug = 1
             else if (ARGV[i] ~ /^-./) {
                 e = sprintf("%s: unrecognized option -- %c",
                         ARGV[0], substr(ARGV[i], 2, 1))
                 print e > "/dev/stderr"
             } else
                 break
             delete ARGV[i]
         }
     }

Ending the awk options with -- isn’t necessary in gawk. Unless --posix has been specified, gawk silently puts any unrecognized options into ARGV for the awk program to deal with. As soon as it sees an unknown option, gawk stops looking for other options that it might otherwise recognize. The previous command line with gawk would be:

     gawk -f myprog.awk -q -v file1 file2 ...

Because -q is not a valid gawk option, it and the following -v are passed on to the awk program. (See Section 10.4 [Processing Command-Line Options], page 233, for an awk library function that parses command-line options.)

When designing your program, you should choose options that don’t conflict with gawk’s, because it will process any options that it accepts before passing the rest of the command line on to your program. Using ‘#!’ with the -E option may help (see Section 1.1.4 [Executable awk Programs], page 19, and see Section 2.2 [Command-Line Options], page 31,).

7.6 Summary

• Pattern–action pairs make up the basic elements of an awk program. Patterns are either normal expressions, range expressions, or regexp constants; one of the special keywords BEGIN, END, BEGINFILE, or ENDFILE; or empty. The action executes if the current record matches the pattern. Empty (missing) patterns match all records.

• I/O from BEGIN and END rules has certain constraints. This is also true, only more so, for BEGINFILE and ENDFILE rules. The latter two give you “hooks” into gawk’s file processing, allowing you to recover from a file that otherwise would cause a fatal error (such as a file that cannot be opened).

• Shell variables can be used in awk programs by careful use of shell quoting. It is easier to pass a shell variable into awk by using the -v option and an awk variable.

• Actions consist of statements enclosed in curly braces. Statements are built up from expressions, control statements, compound statements, input and output statements, and deletion statements.

• The control statements in awk are if-else, while, for, and do-while. gawk adds the switch statement. There are two flavors of for statement: one for performing general looping, and the other for iterating through an array.

• break and continue let you exit early or start the next iteration of a loop (or get out

of a switch).

• next and nextfile let you read the next record and start over at the top of your program or skip to the next input file and start over, respectively.

• The exit statement terminates your program. When executed from an action (or function body), it transfers control to the END statements. From an END statement body, it exits immediately. You may pass an optional numeric value to be used as awk’s exit status.

• Some predefined variables provide control over awk, mainly for I/O. Other variables convey information from awk to your program.

• ARGC and ARGV make the command-line arguments available to your program. Manipulating them from a BEGIN rule lets you control how awk will process the provided data files.

Chapter 8: Arrays in awk 159

8 Arrays in awk

An array is a table of values called elements. The elements of an array are distinguished by their indices. Indices may be either numbers or strings.

This chapter describes how arrays work in awk, how to use array elements, how to scan through every element in an array, and how to remove array elements. It also describes how awk simulates multidimensional arrays, as well as some of the less obvious points about array usage. The chapter moves on to discuss gawk’s facility for sorting arrays, and ends with a brief description of gawk’s ability to support true arrays of arrays.

8.1 The Basics of Arrays

This section presents the basics: working with elements in arrays one at a time, and traversing all of the elements in an array.

8.1.1 Introduction to Arrays

Doing linear scans over an associative array is like trying to club someone to death with a loaded Uzi. —Larry Wall

The awk language provides one-dimensional arrays for storing groups of related strings or numbers. Every awk array must have a name. Array names have the same syntax as variable names; any valid variable name would also be a valid array name. But one name cannot be used in both ways (as an array and as a variable) in the same awk program.

Arrays in awk superficially resemble arrays in other programming languages, but there are fundamental differences. In awk, it isn’t necessary to specify the size of an array before starting to use it. Additionally, any number or string, not just consecutive integers, may be used as an array index.

In most other languages, arrays must be declared before use, including a specification of how many elements or components they contain. In such languages, the declaration causes a contiguous block of memory to be allocated for that many elements. Usually, an index in the array must be a nonnegative integer. For example, the index zero specifies the first element in the array, which is actually stored at the beginning of the block of memory. Index one specifies the second element, which is stored in memory right after the first element, and so on. It is impossible to add more elements to the array, because it has room only for as many elements as given in the declaration. (Some languages allow arbitrary starting and ending indices—e.g., ‘15 .. 27’—but the size of the array is still fixed when the array is declared.)

A contiguous array of four elements might look like Figure 8.1, conceptually, if the element values are eight, "foo", "", and 30.

Figure 8.1: A contiguous array

Only the values are stored; the indices are implicit from the order of the values. Here, eight is the value at index zero, because eight appears in the position with zero elements before it.

Arrays in awk are different—they are associative. This means that each array is a collection of pairs—an index and its corresponding array element value:

Index Value

3       30
1       "foo"
0       8
2       ""

The pairs are shown in jumbled order because their order is irrelevant.^{1}

One advantage of associative arrays is that new pairs can be added at any time. For example, suppose a tenth element is added to the array whose value is "number ten". The result is:

Index Value

10      "number ten"
3       30
1       "foo"
0       8
2       ""

Now the array is sparse, which just means some indices are missing. It has elements 0–3 and 10, but doesn’t have elements 4, 5, 6, 7, 8, or 9.

Another consequence of associative arrays is that the indices don’t have to be nonnegative integers. Any number, or even a string, can be an index. For example, the following is an array that translates words from English to French:

Index Value

"dog"   "chien"
"cat"   "chat"
"one"   "un"

1 The ordering will vary among awk implementations, which typically use hash tables to store array

elements and values.

          1       "un"

Here we decided to translate the number one in both spelled-out and numeric form—thus illustrating that a single array can have both numbers and strings as indices. (In fact, array subscripts are always strings. There are some subtleties to how numbers work when used as array subscripts; this is discussed in more detail in Section 8.2 [Using Numbers to Subscript Arrays], page 167.) Here, the number 1 isn’t double-quoted, because awk automatically converts it to a string.

The value of IGNORECASE has no effect upon array subscripting. The identical string value used to store an array element must be used to retrieve it. When awk creates an array (e.g., with the split() built-in function), that array’s indices are consecutive integers starting at one. (See Section 9.1.3 [String-Manipulation Functions], page 177.)

awk’s arrays are efficient—the time to access an element is independent of the number of elements in the array.

8.1.2 Referring to an Array Element

The principal way to use an array is to refer to one of its elements. An array reference is an expression as follows:

     array[index-expression]

Here, array is the name of an array. The expression index-expression is the index of the desired element of the array.

The value of the array reference is the current value of that array element. For example, foo[4.3] is an expression referencing the element of array foo at index ‘4.3’.

A reference to an array element that has no recorded value yields a value of "", the null string. This includes elements that have not been assigned any value as well as elements that have been deleted (see Section 8.4 [The delete Statement], page 168).

NOTE: A reference to an element that does not exist automatically creates that array element, with the null string as its value. (In some cases, this is unfortunate, because it might waste memory inside awk.) Novice awk programmers often make the mistake of checking if an element exists by checking if the value is empty:

          # Check if "foo" exists in a:  Incorrect!
          if (a["foo"] != "") ...

This is incorrect for two reasons. First, it creates a["foo"] if it didn’t exist before! Second, it is valid (if a bit unusual) to set an array element equal to the empty string.

To determine whether an element exists in an array at a certain index, use the following expression:

     indx in array

This expression tests whether the particular index indx exists, without the side effect of creating that element if it is not present. The expression has the value one (true) if array[indx] exists and zero (false) if it does not exist. (We use indx here, because ‘index’ is the name of a built-in function.) For example, this statement tests whether the array frequencies contains the index ‘2’:

     if (2 in frequencies)
         print "Subscript 2 is present."

Note that this is not a test of whether the array frequencies contains an element whose value is two. There is no way to do that except to scan all the elements. Also, this does not create frequencies[2], while the following (incorrect) alternative does:

     if (frequencies[2] != "")
         print "Subscript 2 is present."

8.1.3 Assigning Array Elements

Array elements can be assigned values just like awk variables:

     array[index-expression] = value

array is the name of an array. The expression index-expression is the index of the element of the array that is assigned a value. The expression value is the value to assign to that element of the array.

8.1.4 Basic Array Example

The following program takes a list of lines, each beginning with a line number, and prints them out in order of line number. The line numbers are not in order when they are first read—instead, they are scrambled. This program sorts the lines by making an array using the line numbers as subscripts. The program then prints out the lines in sorted order of their numbers. It is a very simple program and gets confused upon encountering repeated numbers, gaps, or lines that don’t begin with a number:

     {
         if ($1 > max)
             max = $1
         arr[$1] = $0
     }
     END {
         for (x = 1; x <= max; x++)
             print arr[x]
     }

The first rule keeps track of the largest line number seen so far; it also stores each line into the array arr, at an index that is the line’s number. The second rule runs after all the input has been read, to print out all the lines. When this program is run with the following input:

     5  I am the Five man
     2  Who are you?  The new number two!
     4  . . . And four on the floor
     1  Who is number one?
     3  I three you.

Its output is:

     1  Who is number one?
     2  Who are you?  The new number two!
     3  I three you.

     4  . . . And four on the floor
     5  I am the Five man

If a line number is repeated, the last line with a given number overrides the others. Gaps in the line numbers can be handled with an easy improvement to the program’s END rule, as follows:

     END {
         for (x = 1; x <= max; x++)
             if (x in arr)
                 print arr[x]
     }

8.1.5 Scanning All Elements of an Array

In programs that use arrays, it is often necessary to use a loop that executes once for each element of an array. In other languages, where arrays are contiguous and indices are limited to nonnegative integers, this is easy: all the valid indices can be found by counting from the lowest index up to the highest. This technique won’t do the job in awk, because any number or string can be an array index. So awk has a special kind of for statement for scanning an array:

     for (var in array)
         body

This loop executes body once for each index in array that the program has previously used, with the variable var set to that index.

The following program uses this form of the for statement. The first rule scans the input records and notes which words appear (at least once) in the input, by storing a one into the array used with the word as the index. The second rule scans the elements of used to find all the distinct words that appear in the input. It prints each word that is more than 10 characters long and also prints the number of such words. See Section 9.1.3 [String-Manipulation Functions], page 177, for more information on the built-in function

length().
     # Record a 1 for each word that is used at least once
     {
         for (i = 1; i <= NF; i++)
             used[$i] = 1
     }
     # Find number of distinct words more than 10 characters long
     END {
         for (x in used) {
             if (length(x) > 10) {
                 ++num_long_words
                 print x
             }
         }
         print num_long_words, "words longer than 10 characters"
     }

See Section 11.3.5 [Generating Word-Usage Counts], page 278, for a more detailed example of this type.

The order in which elements of the array are accessed by this statement is determined by the internal arrangement of the array elements within awk and in standard awk cannot be controlled or changed. This can lead to problems if new elements are added to array by statements in the loop body; it is not predictable whether the for loop will reach them. Similarly, changing var inside the loop may produce strange results. It is best to avoid such things.

As a point of information, gawk sets up the list of elements to be iterated over before the loop starts, and does not change it. But not all awk versions do so. Consider this program,

named loopcheck.awk:
     BEGIN {
         a["here"] = "here"
         a["is"] = "is"
         a["a"] = "a"
         a["loop"] = "loop"
         for (i in a) {
             j++
             a[j] = j
             print i
         }
     }

Here is what happens when run with gawk (and mawk):

     $ gawk -f loopcheck.awk
     ⊣ here
     ⊣ loop
     ⊣ a
     ⊣ is

Contrast this to BWK awk:

     $ nawk -f loopcheck.awk
     ⊣ loop
     ⊣ here
     ⊣ is
     ⊣ a
     ⊣ 1

8.1.6 Using Predefined Array Scanning Orders with gawk

This subsection describes a feature that is specific to gawk.

By default, when a for loop traverses an array, the order is undefined, meaning that the awk implementation determines the order in which the array is traversed. This order is usually based on the internal implementation of arrays and will vary from one version of awk to the next.

Often, though, you may wish to do something simple, such as “traverse the array by comparing the indices in ascending order,” or “traverse the array by comparing the values in descending order.” gawk provides two mechanisms that give you this control:

• Set PROCINFO["sorted_in"] to one of a set of predefined values. We describe this now.

• Set PROCINFO["sorted_in"] to the name of a user-defined function to use for comparison of array elements. This advanced feature is described later in Section 12.2 [Controlling Array Traversal and Array Sorting], page 300.

The following special values for PROCINFO["sorted_in"] are available:

"@unsorted"

Array elements are processed in arbitrary order, which is the default awk behavior.

"@ind_str_asc"

Order by indices in ascending order compared as strings; this is the most basic sort. (Internally, array indices are always strings, so with ‘a[2*5] = 1’ the index is "10" rather than numeric 10.)

"@ind_num_asc"

Order by indices in ascending order but force them to be treated as numbers in the process. Any index with a non-numeric value will end up positioned as if it were zero.

"@val_type_asc"

Order by element values in ascending order (rather than by indices). Ordering is by the type assigned to the element (see Section 6.3.2 [Variable Typing and Comparison Expressions], page 120). All numeric values come before all string values, which in turn come before all subarrays. (Subarrays have not been described yet; see Section 8.6 [Arrays of Arrays], page 171.)

"@val_str_asc"

Order by element values in ascending order (rather than by indices). Scalar values are compared as strings. Subarrays, if present, come out last.

"@val_num_asc"

Order by element values in ascending order (rather than by indices). Scalar values are compared as numbers. Subarrays, if present, come out last. When numeric values are equal, the string values are used to provide an ordering: this guarantees consistent results across different versions of the C qsort() function,^{2} which gawk uses internally to perform the sorting.

"@ind_str_desc"

Like "@ind_str_asc", but the string indices are ordered from high to low.

"@ind_num_desc"

Like "@ind_num_asc", but the numeric indices are ordered from high to low.

"@val_type_desc"

Like "@val_type_asc", but the element values, based on type, are ordered from high to low. Subarrays, if present, come out first.

2 When two elements compare as equal, the C qsort() function does not guarantee that they will maintain

their original relative order after sorting. Using the string value to provide a unique ordering when the numeric values are equal ensures that gawk behaves consistently across different environments.

"@val_str_desc"

Like "@val_str_asc", but the element values, treated as strings, are ordered from high to low. Subarrays, if present, come out first.

"@val_num_desc"

Like "@val_num_asc", but the element values, treated as numbers, are ordered from high to low. Subarrays, if present, come out first.

The array traversal order is determined before the for loop starts to run. Changing PROCINFO["sorted_in"] in the loop body does not affect the loop. For example:

$ gawk '
> BEGIN {
>    a[4] = 4
>    a[3] = 3
>    for (i in a)
>        print i, a[i]
> }'
⊣ 4 4
⊣ 3 3
$ gawk '
> BEGIN {
>    PROCINFO["sorted_in"] = "@ind_str_asc"
>    a[4] = 4
>    a[3] = 3
>    for (i in a)
>        print i, a[i]
> }'
⊣ 3 3
⊣ 4 4

When sorting an array by element values, if a value happens to be a subarray then it is considered to be greater than any string or numeric value, regardless of what the subarray itself contains, and all subarrays are treated as being equal to each other. Their order relative to each other is determined by their index strings.

Here are some additional things to bear in mind about sorted array traversal:

• The value of PROCINFO["sorted_in"] is global. That is, it affects all array traversal for loops. If you need to change it within your own code, you should see if it’s defined and save and restore the value:

    ...
    if ("sorted_in" in PROCINFO) {
        save_sorted = PROCINFO["sorted_in"]
        PROCINFO["sorted_in"] = "@val_str_desc" # or whatever
    }
    ...
    if (save_sorted)
        PROCINFO["sorted_in"] = save_sorted

• As already mentioned, the default array traversal order is represented by "@unsorted". You can also get the default behavior by assigning the null string to PROCINFO["sorted_in"] or by just deleting the "sorted_in" element from the PROCINFO array with the delete statement. (The delete statement hasn’t been described yet; see Section 8.4 [The delete Statement], page 168.)

In addition, gawk provides built-in functions for sorting arrays; see Section 12.2.2 [Sorting Array Values and Indices with gawk], page 304.

8.2 Using Numbers to Subscript Arrays

An important aspect to remember about arrays is that array subscripts are always strings. When a numeric value is used as a subscript, it is converted to a string value before being used for subscripting (see Section 6.1.4 [Conversion of Strings and Numbers], page 111). This means that the value of the predefined variable CONVFMT can affect how your program accesses elements of an array. For example:

     xyz = 12.153
     data[xyz] = 1
     CONVFMT = "%2.2f"
     if (xyz in data)
         printf "%s is in data\n", xyz
     else
         printf "%s is not in data\n", xyz

This prints ‘12.15 is not in data’. The first statement gives xyz a numeric value. Assigning to data[xyz] subscripts data with the string value "12.153" (using the default conversion value of CONVFMT, "%.6g"). Thus, the array element data["12.153"] is assigned the value one. The program then changes the value of CONVFMT. The test ‘(xyz in data)’ generates a new string value from xyz—this time "12.15"—because the value of CONVFMT only allows two significant digits. This test fails, because "12.15" is different from

"12.153".

According to the rules for conversions (see Section 6.1.4 [Conversion of Strings and Numbers], page 111), integer values always convert to strings as integers, no matter what the value of CONVFMT may happen to be. So the usual case of the following works:

     for (i = 1; i <= maxsub; i++)

do something with array[i]

The “integer values always convert to strings as integers” rule has an additional consequence for array indexing. Octal and hexadecimal constants (see Section 6.1.1.2 [Octal and Hexadecimal Numbers], page 107) are converted internally into numbers, and their original form is forgotten. This means, for example, that array[17], array[021], and array[0x11] all refer to the same element!

As with many things in awk, the majority of the time things work as you would expect them to. But it is useful to have a precise knowledge of the actual rules, as they can sometimes have a subtle effect on your programs.

8.3 Using Uninitialized Variables as Subscripts

Suppose it’s necessary to write a program to print the input data in reverse order. A reasonable attempt to do so (with some test data) might look like this:

     $ echo 'line 1
     > line 2
     > line 3' | awk '{ l[lines] = $0; ++lines }
     > END {
     >     for (i = lines - 1; i >= 0; i--)
     >        print l[i]
     > }'
     ⊣ line 3
     ⊣ line 2

Unfortunately, the very first line of input data did not appear in the output!

Upon first glance, we would think that this program should have worked. The variable lines is uninitialized, and uninitialized variables have the numeric value zero. So, awk should have printed the value of l[0].

The issue here is that subscripts for awk arrays are always strings. Uninitialized variables, when used as strings, have the value "", not zero. Thus, ‘line 1’ ends up stored in l[""]. The following version of the program works correctly:

     { l[lines++] = $0 }
     END {
         for (i = lines - 1; i >= 0; i--)
            print l[i]
     }

Here, the ‘++’ forces lines to be numeric, thus making the “old value” numeric zero. This is then converted to "0" as the array subscript.

Even though it is somewhat unusual, the null string ("") is a valid array subscript. gawk warns about the use of the null string as a subscript if --lint is provided on the command line (see Section 2.2 [Command-Line Options], page 31).

8.4 The delete Statement

To remove an individual element of an array, use the delete statement:

     delete array[index-expression]

Once an array element has been deleted, any value the element once had is no longer available. It is as if the element had never been referred to or been given a value. The following is an example of deleting elements in an array:

     for (i in frequencies)
         delete frequencies[i]

This example removes all the elements from the array frequencies. Once an element is deleted, a subsequent for statement to scan the array does not report that element and using the in operator to check for the presence of that element returns zero (i.e., false):

     delete foo[4]
     if (4 in foo)
         print "This will never be printed"

It is important to note that deleting an element is not the same as assigning it a null value (the empty string, ""). For example:

     foo[4] = ""
     if (4 in foo)
       print "This is printed, even though foo[4] is empty"

It is not an error to delete an element that does not exist. However, if --lint is provided on the command line (see Section 2.2 [Command-Line Options], page 31), gawk issues a warning message when an element that is not in the array is deleted.

All the elements of an array may be deleted with a single statement by leaving off the subscript in the delete statement, as follows:

     delete array

Using this version of the delete statement is about three times more efficient than the equivalent loop that deletes each element one at a time.

This form of the delete statement is also supported by BWK awk and mawk, as well as by a number of other implementations.

NOTE: For many years, using delete without a subscript was a common extension. In September 2012, it was accepted for inclusion into the POSIX standard. See the Austin Group website (http://austingroupbugs.net/

     view.php?id=544).

The following statement provides a portable but nonobvious way to clear out an array:^{3}

     split("", array)

The split() function (see Section 9.1.3 [String-Manipulation Functions], page 177) clears out the target array first. This call asks it to split apart the null string. Because there is no data to split out, the function simply clears the array and then returns.

CAUTION: Deleting all the elements from an array does not change its type; you cannot clear an array and then use the array’s name as a scalar (i.e., a regular variable). For example, the following does not work:

          a[1] = 3
          delete a
          a = 3

8.5 Multidimensional Arrays

A multidimensional array is an array in which an element is identified by a sequence of indices instead of a single index. For example, a two-dimensional array requires two indices. The usual way (in many languages, including awk) to refer to an element of a two-dimensional array named grid is with grid[x,y].

Multidimensional arrays are supported in awk through concatenation of indices into one string. awk converts the indices into strings (see Section 6.1.4 [Conversion of Strings and Numbers], page 111) and concatenates them together, with a separator between them. This creates a single string that describes the values of the separate indices. The combined string is used as a single index into an ordinary, one-dimensional array. The separator used is the value of the built-in variable SUBSEP.

3 Thanks to Michael Brennan for pointing this out.

For example, suppose we evaluate the expression ‘foo[5,12] = "value"’ when the value of SUBSEP is "@". The numbers 5 and 12 are converted to strings and concatenated with an ‘@’ between them, yielding "5@12"; thus, the array element foo["5@12"] is set to "value".

Once the element’s value is stored, awk has no record of whether it was stored with a single index or a sequence of indices. The two expressions ‘foo[5,12]’ and ‘foo[5 SUBSEP 12]’ are always equivalent.

The default value of SUBSEP is the string "\034", which contains a nonprinting character that is unlikely to appear in an awk program or in most input data. The usefulness of choosing an unlikely character comes from the fact that index values that contain a string matching SUBSEP can lead to combined strings that are ambiguous. Suppose that SUBSEP is "@"; then ‘foo["a@b", "c"]’ and ‘foo["a", "b@c"]’ are indistinguishable because both are actually stored as ‘foo["a@b@c"]’.

To test whether a particular index sequence exists in a multidimensional array, use the same operator (in) that is used for single-dimensional arrays. Write the whole sequence of indices in parentheses, separated by commas, as the left operand:

     if ((subscript1, subscript2, ...) in array)
         ...

Here is an example that treats its input as a two-dimensional array of fields; it rotates this array 90 degrees clockwise and prints the result. It assumes that all lines have the same number of elements:

     {
          if (max_nf < NF)
               max_nf = NF
          max_nr = NR
          for (x = 1; x <= NF; x++)
               vector[x, NR] = $x
     }
     END {
          for (x = 1; x <= max_nf; x++) {
               for (y = max_nr; y >= 1; --y)
                    printf("%s ", vector[x, y])
               printf("\n")
          }
     }

When given the input:

     1 2 3 4 5 6
     2 3 4 5 6 1
     3 4 5 6 1 2
     4 5 6 1 2 3

the program produces the following output:

     4 3 2 1
     5 4 3 2
     6 5 4 3
     1 6 5 4

     2 1 6 5
     3 2 1 6

8.5.1 Scanning Multidimensional Arrays

There is no special for statement for scanning a “multidimensional” array. There cannot be one, because, in truth, awk does not have multidimensional arrays or elements—there is only a multidimensional way of accessing an array.

However, if your program has an array that is always accessed as multidimensional, you can get the effect of scanning it by combining the scanning for statement (see Section 8.1.5 [Scanning All Elements of an Array], page 163) with the built-in split() function (see Section 9.1.3 [String-Manipulation Functions], page 177). It works in the following manner:

     for (combined in array) {
         split(combined, separate, SUBSEP)
         ...
     }

This sets the variable combined to each concatenated combined index in the array, and splits it into the individual indices by breaking it apart where the value of SUBSEP appears. The individual indices then become the elements of the array separate.

Thus, if a value is previously stored in array[1, "foo"], then an element with index "1\034foo" exists in array. (Recall that the default value of SUBSEP is the character with code 034.) Sooner or later, the for statement finds that index and does an iteration with the variable combined set to "1\034foo". Then the split() function is called as follows:

     split("1\034foo", separate, "\034")

The result is to set separate[1] to "1" and separate[2] to "foo". Presto! The original sequence of separate indices is recovered.

8.6 Arrays of Arrays

gawk goes beyond standard awk’s multidimensional array access and provides true arrays of arrays. Elements of a subarray are referred to by their own indices enclosed in square brackets, just like the elements of the main array. For example, the following creates a two-element subarray at index 1 of the main array a:

     a[1][1] = 1
     a[1][2] = 2

This simulates a true two-dimensional array. Each subarray element can contain another subarray as a value, which in turn can hold other arrays as well. In this way, you can create arrays of three or more dimensions. The indices can be any awk expressions, including scalars separated by commas (i.e., a regular awk simulated multidimensional subscript). So the following is valid in gawk:

     a[1][3][1, "name"] = "barney"

Each subarray and the main array can be of different length. In fact, the elements of an array or its subarray do not all have to have the same type. This means that the main array and any of its subarrays can be nonrectangular, or jagged in structure. You can assign a scalar value to the index 4 of the main array a, even though a[1] is itself an array and not a scalar:

     a[4] = "An element in a jagged array"

The terms dimension, row, and column are meaningless when applied to such an array, but we will use “dimension” henceforth to imply the maximum number of indices needed to refer to an existing element. The type of any element that has already been assigned cannot be changed by assigning a value of a different type. You have to first delete the current element, which effectively makes gawk forget about the element at that index:

     delete a[4]
     a[4][5][6][7] = "An element in a four-dimensional array"

This removes the scalar value from index 4 and then inserts a three-level nested subarray containing a scalar. You can also delete an entire subarray or subarray of subarrays:

     delete a[4][5]
     a[4][5] = "An element in subarray a[4]"

But recall that you can not delete the main array a and then use it as a scalar.

The built-in functions that take array arguments can also be used with subarrays. For example, the following code fragment uses length() (see Section 9.1.3 [String-Manipulation Functions], page 177) to determine the number of elements in the main array a and its subarrays:

     print length(a), length(a[1]), length(a[1][3])

This results in the following output for our main array a:

     2, 3, 1

The ‘subscript in array’ expression (see Section 8.1.2 [Referring to an Array Element], page 161) works similarly for both regular awk-style arrays and arrays of arrays. For example, the tests ‘1 in a’, ‘3 in a[1]’, and ‘(1, "name") in a[1][3]’ all evaluate to one (true) for our array a.

The ‘for (item in array)’ statement (see Section 8.1.5 [Scanning All Elements of an Array], page 163) can be nested to scan all the elements of an array of arrays if it is rectangular in structure. In order to print the contents (scalar values) of a two-dimensional array of arrays (i.e., in which each first-level element is itself an array, not necessarily of the same length), you could use the following code:

     for (i in array)
         for (j in array[i])
             print array[i][j]

The isarray() function (see Section 9.1.7 [Getting Type Information], page 198) lets you test if an array element is itself an array:

     for (i in array) {
         if (isarray(array[i]) {
             for (j in array[i]) {
                 print array[i][j]
             }
         }
         else
             print array[i]
     }

If the structure of a jagged array of arrays is known in advance, you can often devise workarounds using control statements. For example, the following code prints the elements of our main array a:

 for (i in a) {
     for (j in a[i]) {
         if (j == 3) {
             for (k in a[i][j])
                 print a[i][j][k]
         } else
             print a[i][j]
     }
 }

See Section 10.7 [Traversing Arrays of Arrays], page 246, for a user-defined function that “walks” an arbitrarily dimensioned array of arrays.

Recall that a reference to an uninitialized array element yields a value of "", the null string. This has one important implication when you intend to use a subarray as an argument to a function, as illustrated by the following example:

 $ gawk 'BEGIN { split("a b c d", b[1]); print b[1][1] }'
  error gawk: cmd. line:1: fatal: split: second argument is not an array

The way to work around this is to first force b[1] to be an array by creating an arbitrary index:

 $ gawk 'BEGIN { b[1][1] = ""; split("a b c d", b[1]); print b[1][1] }'
  ⊣ a

8.7 Summary

• Standard awk provides one-dimensional associative arrays (arrays indexed by string values). All arrays are associative; numeric indices are converted automatically to strings.

• Array elements are referenced as array[indx]. Referencing an element creates it if it did not exist previously.

• The proper way to see if an array has an element with a given index is to use the in operator: ‘indx in array’.

• Use ‘for (indx in array) ...’ to scan through all the individual elements of an array. In the body of the loop, indx takes on the value of each element’s index in turn.

• The order in which a ‘for (indx in array)’ loop traverses an array is undefined in POSIX awk and varies among implementations. gawk lets you control the order by assigning special predefined values to PROCINFO["sorted_in"].

• Use ‘delete array[indx]’ to delete an individual element. To delete all of the elements in an array, use ‘delete array’. This latter feature has been a common extension for many years and is now standard, but may not be supported by all commercial versions

of awk.

• Standard awk simulates multidimensional arrays by separating subscript values with commas. The values are concatenated into a single string, separated by the value of SUBSEP. The fact that such a subscript was created in this way is not retained; thus, changing SUBSEP may have unexpected consequences. You can use ‘(sub1, sub2, ...) in array’ to see if such a multidimensional subscript exists in array.

• gawk provides true arrays of arrays. You use a separate set of square brackets for each dimension in such an array: data[row][col], for example. Array elements may thus be either scalar values (number or string) or other arrays.

• Use the isarray() built-in function to determine if an array element is itself a subarray.

Chapter 9: Functions 175

9 Functions

This chapter describes awk’s built-in functions, which fall into three categories: numeric, string, and I/O. gawk provides additional groups of functions to work with values that represent time, do bit manipulation, sort arrays, provide type information, and internationalize and localize programs.

Besides the built-in functions, awk has provisions for writing new functions that the rest of a program can use. The second half of this chapter describes these user-defined functions. Finally, we explore indirect function calls, a gawk-specific extension that lets you determine at runtime what function is to be called.

9.1 Built-in Functions

Built-in functions are always available for your awk program to call. This section defines all the built-in functions in awk; some of these are mentioned in other sections but are summarized here for your convenience.

9.1.1 Calling Built-in Functions

To call one of awk’s built-in functions, write the name of the function followed by arguments in parentheses. For example, ‘atan2(y + z, 1)’ is a call to the function atan2() and has two arguments.

Whitespace is ignored between the built-in function name and the opening parenthesis, but nonetheless it is good practice to avoid using whitespace there. User-defined functions do not permit whitespace in this way, and it is easier to avoid mistakes by following a simple convention that always works—no whitespace after a function name.

Each built-in function accepts a certain number of arguments. In some cases, arguments can be omitted. The defaults for omitted arguments vary from function to function and are described under the individual functions. In some awk implementations, extra arguments given to built-in functions are ignored. However, in gawk, it is a fatal error to give extra arguments to a built-in function.

When a function is called, expressions that create the function’s actual parameters are evaluated completely before the call is performed. For example, in the following code fragment:

     i = 4
     j = sqrt(i++)

the variable i is incremented to the value five before sqrt() is called with a value of four for its actual parameter. The order of evaluation of the expressions used for the function’s parameters is undefined. Thus, avoid writing programs that assume that parameters are evaluated from left to right or from right to left. For example:

     i = 5
     j = atan2(++i, i *= 2)

If the order of evaluation is left to right, then i first becomes six, and then 12, and atan2() is called with the two arguments six and 12. But if the order of evaluation is right to left, i first becomes 10, then 11, and atan2() is called with the two arguments 11 and

9.1.2 Numeric Functions

The following list describes all of the built-in functions that work with numbers. Optional parameters are enclosed in square brackets ([ ]):

atan2(y, x)

Return the arctangent of y / x in radians. You can use ‘pi = atan2(0, -1)’ to retrieve the value of π.

cos(x) Return the cosine of x, with x in radians.

exp(x) Return the exponential of x (e ^x) or report an error if x is out of range. The range of values x can have depends on your machine’s floating-point representation.

int(x) Return the nearest integer to x, located between x and zero and truncated toward zero. For example, int(3) is 3, int(3.9) is 3, int(-3.9) is −3, and int(-3) is −3 as well.

log(x) Return the natural logarithm of x, if x is positive; otherwise, return NaN (“not a number”) on IEEE 754 systems. Additionally, gawk prints a warning message when x is negative.

rand() Return a random number. The values of rand() are uniformly distributed between zero and one. The value could be zero but is never one.^{1} Often random integers are needed instead. Following is a user-defined function that can be used to obtain a random nonnegative integer less than n:

     function randint(n)
     {
         return int(n * rand())
     }

The multiplication produces a random number greater than or equal to zero and less than n. Using int(), this result is made into an integer between zero and n − 1, inclusive. The following example uses a similar function to produce random integers between one and n. This program prints a new random number for each input record:

     # Function to roll a simulated die.
     function roll(n) { return 1 + int(rand() * n) }
     # Roll 3 six-sided dice and
     # print total number of points.
     {
         printf("%d points\n", roll(6) + roll(6) + roll(6))
     }

1 The C version of rand() on many Unix systems is known to produce fairly poor sequences of random

numbers. However, nothing requires that an awk implementation use the C rand() to implement the awk version of rand(). In fact, gawk uses the BSD random() function, which is considerably better than rand(), to produce random numbers.

CAUTION: In most awk implementations, including gawk, rand() starts generating numbers from the same starting number, or seed, each time you run awk.^{2} Thus, a program generates the same results each time you run it. The numbers are random within one awk run but predictable from run to run. This is convenient for debugging, but if you want a program to do different things each time it is used, you must change the seed to a value that is different in each run. To do this, use srand().

sin(x) Return the sine of x, with x in radians.

sqrt(x) Return the positive square root of x. gawk prints a warning message if x is negative. Thus, sqrt(4) is 2.

srand([x])

Set the starting point, or seed, for generating random numbers to the value x. Each seed value leads to a particular sequence of random numbers.^{3} Thus, if the seed is set to the same value a second time, the same sequence of random numbers is produced again.

CAUTION: Different awk implementations use different randomnumber generators internally. Don’t expect the same awk program to produce the same series of random numbers when executed by different versions of awk.

If the argument x is omitted, as in ‘srand()’, then the current date and time of day are used for a seed. This is the way to get random numbers that are truly unpredictable. The return value of srand() is the previous seed. This makes it easy to keep track of the seeds in case you need to consistently reproduce sequences of random numbers. POSIX does not specify the initial seed; it differs among awk implementations.

9.1.3 String-Manipulation Functions

The functions in this section look at or change the text of one or more strings.

gawk understands locales (see Section 6.6 [Where You Are Makes a Difference], page 129) and does all string processing in terms of characters, not bytes. This distinction is particularly important to understand for locales where one character may be represented by multiple bytes. Thus, for example, length() returns the number of characters in a string, and not the number of bytes used to represent those characters. Similarly, index() works with character indices, and not byte indices.

CAUTION: A number of functions deal with indices into strings. For these functions, the first character of a string is at position (index) one. This is different from C and the languages descended from it, where the first character is at position zero. You need to remember this when doing index calculations, particularly if you are used to C.

2 mawk uses a different seed each time.

3 Computer-generated random numbers really are not truly random. They are technically known as

pseudorandom. This means that although the numbers in a sequence appear to be random, you can in fact generate the same sequence of random numbers over and over again.

In the following list, optional parameters are enclosed in square brackets ([ ]). Several functions perform string substitution; the full discussion is provided in the description of the sub() function, which comes toward the end, because the list is presented alphabetically.

Those functions that are specific to gawk are marked with a pound sign (‘#’). They are not available in compatibility mode (see Section 2.2 [Command-Line Options], page 31):

asort(source [, dest [, how ] ]) # asorti(source [, dest [, how ] ]) #

These two functions are similar in behavior, so they are described together. NOTE: The following description ignores the third argument, how, as it requires understanding features that we have not discussed yet. Thus, the discussion here is a deliberate simplification. (We do provide all the details later on; see Section 12.2.2 [Sorting Array Values and Indices with gawk], page 304, for the full story.)

Both functions return the number of elements in the array source. For asort(), gawk sorts the values of source and replaces the indices of the sorted values of source with sequential integers starting with one. If the optional array dest is specified, then source is duplicated into dest. dest is then sorted, leaving the indices of source unchanged. When comparing strings, IGNORECASE affects the sorting (see Section 12.2.2 [Sorting Array Values and Indices with gawk], page 304). If the source array contains subarrays as values (see Section 8.6 [Arrays of Arrays], page 171), they will come last, after all scalar values. Subarrays are not recursively sorted. For example, if the contents of a are as follows:

     a["last"] = "de"
     a["first"] = "sac"
     a["middle"] = "cul"

A call to asort():

     asort(a)

results in the following contents of a:

     a[1] = "cul"
     a[2] = "de"
     a[3] = "sac"

The asorti() function works similarly to asort(); however, the indices are sorted, instead of the values. Thus, in the previous example, starting with the same initial set of indices and values in a, calling ‘asorti(a)’ would yield:

     a[1] = "first"
     a[2] = "last"
     a[3] = "middle"
gensub(regexp, replacement, how [, target]) #

Search the target string target for matches of the regular expression regexp. If how is a string beginning with ‘g’ or ‘G’ (short for “global”), then replace all matches of regexp with replacement. Otherwise, how is treated as a number indicating which match of regexp to replace. If no target is supplied, use $0. It returns the modified string as the result of the function and the original target string is not changed. gensub() is a general substitution function. Its purpose is to provide more features than the standard sub() and gsub() functions. gensub() provides an additional feature that is not available in sub() or gsub(): the ability to specify components of a regexp in the replacement text. This is done by using parentheses in the regexp to mark the components and then specifying ‘\N’ in the replacement text, where N is a digit from 1 to 9. For example:

     $ gawk '
     > BEGIN {
     >      a = "abc def"
     >      b = gensub(/(.+) (.+)/, "\\2 \\1", "g", a)
     >      print b
     > }'
     ⊣ def abc

As with sub(), you must type two backslashes in order to get one into the string. In the replacement text, the sequence ‘\0’ represents the entire matched text, as does the character ‘&’. The following example shows how you can use the third argument to control which match of the regexp should be changed:

     $ echo a b c a b c |
     > gawk '{ print gensub(/a/, "AA", 2) }'
     ⊣ a b c AA b c

In this case, $0 is the default target string. gensub() returns the new string as its result, which is passed directly to print for printing. If the how argument is a string that does not begin with ‘g’ or ‘G’, or if it is a number that is less than or equal to zero, only one substitution is performed. If how is zero, gawk issues a warning message. If regexp does not match target, gensub()’s return value is the original unchanged value of target.

gsub(regexp, replacement [, target])

Search target for all of the longest, leftmost, nonoverlapping matching substrings it can find and replace them with replacement. The ‘g’ in gsub() stands for “global,” which means replace everywhere. For example:

     { gsub(/Britain/, "United Kingdom"); print }

replaces all occurrences of the string ‘Britain’ with ‘United Kingdom’ for all input records. The gsub() function returns the number of substitutions made. If the variable to search and alter (target) is omitted, then the entire input record ($0) is used. As in sub(), the characters ‘&’ and ‘\’ are special, and the third argument must be assignable.

index(in, find)

Search the string in for the first occurrence of the string find, and return the position in characters where that occurrence begins in the string in. Consider the following example:

     $ awk 'BEGIN { print index("peanut", "an") }'
     ⊣ 3

If find is not found, index() returns zero. With BWK awk and gawk, it is a fatal error to use a regexp constant for find. Other implementations allow it, simply treating the regexp constant as an expression meaning ‘$0 ~ /regexp/’.

length([string])

Return the number of characters in string. If string is a number, the length of the digit string representing that number is returned. For example, length("abcde") is five. By contrast, length(15 * 35) works out to three. In this example, 15 · 35 = 525, and 525 is then converted to the string "525", which has three characters. If no argument is supplied, length() returns the length of $0.

NOTE: In older versions of awk, the length() function could be called without any parentheses. Doing so is considered poor practice, although the 2008 POSIX standard explicitly allows it, to support historical practice. For programs to be maximally portable, always supply the parentheses.

If length() is called with a variable that has not been used, gawk forces the variable to be a scalar. Other implementations of awk leave the variable without a type. Consider:

     $ gawk 'BEGIN { print length(x) ; x[1] = 1 }'
     ⊣ 0
      error gawk: fatal: attempt to use scalar `x' as array
     $ nawk 'BEGIN { print length(x) ; x[1] = 1 }'
     ⊣ 0

If --lint has been specified on the command line, gawk issues a warning about this. With gawk and several other awk implementations, when given an array argument, the length() function returns the number of elements in the array. (c.e.) This is less useful than it might seem at first, as the array is not guaranteed to be indexed from one to the number of elements in it. If --lint is provided on the command line (see Section 2.2 [Command-Line Options], page 31), gawk warns that passing an array argument is not portable. If --posix is supplied, using an array argument is a fatal error (see Chapter 8 [Arrays in awk], page 159).

match(string, regexp [, array])

Search string for the longest, leftmost substring matched by the regular expression regexp and return the character position (index) at which that substring begins (one, if it starts at the beginning of string). If no match is found, return zero.

The regexp argument may be either a regexp constant (/…/) or a string constant ("…"). In the latter case, the string is treated as a regexp to be matched. See Section 3.6 [Using Dynamic Regexps], page 52, for a discussion of the difference between the two forms, and the implications for writing your program correctly. The order of the first two arguments is the opposite of most other string functions that work with regular expressions, such as sub() and gsub(). It might help to remember that for match(), the order is the same as for the ‘~’ operator:

‘string ~ regexp’.

The match() function sets the predefined variable RSTART to the index. It also sets the predefined variable RLENGTH to the length in characters of the matched substring. If no match is found, RSTART is set to zero, and RLENGTH to −1. For example:

     {
         if ($1 == "FIND")
             regex = $2
         else {
             where = match($0, regex)
             if (where != 0)
                 print "Match of", regex, "found at", where, "in", $0
            }
     }

This program looks for lines that match the regular expression stored in the variable regex. This regular expression can be changed. If the first word on a line is ‘FIND’, regex is changed to be the second word on that line. Therefore, if given:

     FIND ru+n
     My program runs
     but not very quickly
     FIND Melvin
     JF+KM
     This line is property of Reality Engineering Co.
     Melvin was here.

awk prints:

     Match of ru+n found at 12 in My program runs
     Match of Melvin found at 1 in Melvin was here.

If array is present, it is cleared, and then the zeroth element of array is set to the entire portion of string matched by regexp. If regexp contains parentheses, the integer-indexed elements of array are set to contain the portion of string matching the corresponding parenthesized subexpression. For example:

     $ echo foooobazbarrrrr |
     > gawk '{ match($0, /(fo+).+(bar*)/, arr)
     >         print arr[1], arr[2] }'
     ⊣ foooo barrrrr

In addition, multidimensional subscripts are available providing the start index and length of each matched subexpression:

     $ echo foooobazbarrrrr |
     > gawk '{ match($0, /(fo+).+(bar*)/, arr)
     >           print arr[1], arr[2]
     >           print arr[1, "start"], arr[1, "length"]
     >           print arr[2, "start"], arr[2, "length"]
     > }'
     ⊣ foooo barrrrr
     ⊣ 1 5
     ⊣ 9 7

There may not be subscripts for the start and index for every parenthesized subexpression, because they may not all have matched text; thus, they should be tested for with the in operator (see Section 8.1.2 [Referring to an Array Element], page 161). The array argument to match() is a gawk extension. In compatibility mode (see Section 2.2 [Command-Line Options], page 31), using a third argument is a fatal error.

patsplit(string, array [, fieldpat [, seps ] ]) #

Divide string into pieces defined by fieldpat and store the pieces in array and the separator strings in the seps array. The first piece is stored in array[1], the second piece in array[2], and so forth. The third argument, fieldpat, is a regexp describing the fields in string (just as FPAT is a regexp describing the fields in input records). It may be either a regexp constant or a string. If fieldpat is omitted, the value of FPAT is used. patsplit() returns the number of elements created. seps[i] is the separator string between array[i] and array[i+1]. Any leading separator will be in seps[0]. The patsplit() function splits strings into pieces in a manner similar to the way input lines are split into fields using FPAT (see Section 4.7 [Defining Fields by Content], page 74). Before splitting the string, patsplit() deletes any previously existing elements in the arrays array and seps.

split(string, array [, fieldsep [, seps ] ])

Divide string into pieces separated by fieldsep and store the pieces in array and the separator strings in the seps array. The first piece is stored in array[1], the second piece in array[2], and so forth. The string value of the third argument, fieldsep, is a regexp describing where to split string (much as FS can be a regexp describing where to split input records). If fieldsep is omitted, the value of FS is used. split() returns the number of elements created. seps is a gawk extension, with seps[i] being the separator string between array[i] and array[i+1]. If fieldsep is a single space, then any leading whitespace goes into seps[0] and any trailing whitespace goes into seps[n], where n is the return value of split() (i.e., the number of elements in array). The split() function splits strings into pieces in a manner similar to the way input lines are split into fields. For example:

     split("cul-de-sac", a, "-", seps)

splits the string "cul-de-sac" into three fields using ‘-’ as the separator. It sets the contents of the array a as follows:

     a[1] = "cul"
     a[2] = "de"
     a[3] = "sac"

and sets the contents of the array seps as follows:

     seps[1] = "-"
     seps[2] = "-"

The value returned by this call to split() is three. As with input field-splitting, when the value of fieldsep is " ", leading and trailing whitespace is ignored in values assigned to the elements of array but not in seps, and the elements are separated by runs of whitespace. Also, as with input field splitting, if fieldsep is the null string, each individual character in the string is split into its own array element. (c.e.) Note, however, that RS has no effect on the way split() works. Even though ‘RS = ""’ causes the newline character to also be an input field separator, this does not affect how split() splits strings. Modern implementations of awk, including gawk, allow the third argument to be a regexp constant (/…/) as well as a string. The POSIX standard allows this as well. See Section 3.6 [Using Dynamic Regexps], page 52, for a discussion of the difference between using a string constant or a regexp constant, and the implications for writing your program correctly. Before splitting the string, split() deletes any previously existing elements in the arrays array and seps. If string is null, the array has no elements. (So this is a portable way to delete an entire array with one statement. See Section 8.4 [The delete Statement], page 168.) If string does not match fieldsep at all (but is not null), array has one element only. The value of that element is the original string. In POSIX mode (see Section 2.2 [Command-Line Options], page 31), the fourth argument is not allowed.

sprintf(format, expression1, ...)

Return (without printing) the string that printf would have printed out with the same arguments (see Section 5.5 [Using printf Statements for Fancier Printing], page 92). For example:

     pival = sprintf("pi = %.2f (approx.)", 22/7)

assigns the string ‘pi = 3.14 (approx.)’ to the variable pival.

strtonum(str) #

Examine str and return its numeric value. If str begins with a leading ‘0’, strtonum() assumes that str is an octal number. If str begins with a leading ‘0x’ or ‘0X’, strtonum() assumes that str is a hexadecimal number. For example:

     $ echo 0x11 |
     > gawk '{ printf "%d\n", strtonum($1) }'
     ⊣ 17

Using the strtonum() function is not the same as adding zero to a string value; the automatic coercion of strings to numbers works only for decimal data, not for octal or hexadecimal.^{4}

Note also that strtonum() uses the current locale’s decimal point for recognizing numbers (see Section 6.6 [Where You Are Makes a Difference], page 129).

sub(regexp, replacement [, target])

Search target, which is treated as a string, for the leftmost, longest substring matched by the regular expression regexp. Modify the entire string by replacing the matched text with replacement. The modified string becomes the new value of target. Return the number of substitutions made (zero or one).

The regexp argument may be either a regexp constant (/…/) or a string constant ("…"). In the latter case, the string is treated as a regexp to be matched. See Section 3.6 [Using Dynamic Regexps], page 52, for a discussion of the difference between the two forms, and the implications for writing your program correctly.

This function is peculiar because target is not simply used to compute a value, and not just any expression will do—it must be a variable, field, or array element so that sub() can store a modified value there. If this argument is omitted, then the default is to use and alter $0.^{5} For example:

     str = "water, water, everywhere"
     sub(/at/, "ith", str)

sets str to ‘wither, water, everywhere’, by replacing the leftmost longest occurrence of ‘at’ with ‘ith’.

If the special character ‘&’ appears in replacement, it stands for the precise substring that was matched by regexp. (If the regexp can match more than one string, then this precise substring may vary.) For example:

     { sub(/candidate/, "& and his wife"); print }

changes the first occurrence of ‘candidate’ to ‘candidate and his wife’ on each input line. Here is another example:

     $ awk 'BEGIN {
     >         str = "daabaaa"
     >         sub(/a+/, "C&C", str)
     >         print str
     > }'
     ⊣ dCaaCbaaa

4 Unless you use the --non-decimal-data option, which isn’t recommended. See Section 12.1 [Allowing

Nondecimal Input Data], page 299, for more information.

5 Note that this means that the record will first be regenerated using the value of OFS if any fields have

been changed, and that the fields will be updated after the substitution, even if the operation is a “no-op”

such as ‘sub(/^/, "")’.

This shows how ‘&’ can represent a nonconstant string and also illustrates the “leftmost, longest” rule in regexp matching (see Section 3.5 [How Much Text Matches?], page 52). The effect of this special character (‘&’) can be turned off by putting a backslash before it in the string. As usual, to insert one backslash in the string, you must write two backslashes. Therefore, write ‘\\&’ in a string constant to include a literal ‘&’ in the replacement. For example, the following shows how to replace the first ‘|’ on each line with an ‘&’:

     { sub(/\|/, "\\&"); print }

As mentioned, the third argument to sub() must be a variable, field, or array element. Some versions of awk allow the third argument to be an expression that is not an lvalue. In such a case, sub() still searches for the pattern and returns zero or one, but the result of the substitution (if any) is thrown away because there is no place to put it. Such versions of awk accept expressions like the following:

     sub(/USA/, "United States", "the USA and Canada")

For historical compatibility, gawk accepts such erroneous code. However, using any other nonchangeable object as the third parameter causes a fatal error and your program will not run. Finally, if the regexp is not a regexp constant, it is converted into a string, and then the value of that string is treated as the regexp to match.

substr(string, start [, length ])

Return a length-character-long substring of string, starting at character number start. The first character of a string is character number one.^{6} For example,

substr("washington", 5, 3) returns "ing".

If length is not present, substr() returns the whole suffix of string that begins at character number start. For example, substr("washington", 5) returns "ington". The whole suffix is also returned if length is greater than the number of characters remaining in the string, counting from character start. If start is less than one, substr() treats it as if it was one. (POSIX doesn’t specify what to do in this case: BWK awk acts this way, and therefore gawk does too.) If start is greater than the number of characters in the string, substr() returns the null string. Similarly, if length is present but less than or equal to zero, the null string is returned. The string returned by substr() cannot be assigned. Thus, it is a mistake to attempt to change a portion of a string, as shown in the following example:

     string = "abcdef"
     # try to get "abCDEf", won't work
     substr(string, 3, 3) = "CDE"

It is also a mistake to use substr() as the third argument of sub() or gsub():

     gsub(/xyz/, "pdq", substr($0, 5, 20))  # WRONG

6 This is different from C and C++, in which the first character is number zero.

(Some commercial versions of awk treat substr() as assignable, but doing so is not portable.) If you need to replace bits and pieces of a string, combine substr() with string concatenation, in the following manner:

               string = "abcdef"
               ...
               string = substr(string, 1, 2) "CDE" substr(string, 6)
tolower(string)

Return a copy of string, with each uppercase character in the string replaced with its corresponding lowercase character. Nonalphabetic characters are left

          unchanged. For example, tolower("MiXeD cAsE 123") returns "mixed case
          123".
toupper(string)

Return a copy of string, with each lowercase character in the string replaced with its corresponding uppercase character. Nonalphabetic characters are left

          unchanged. For example, toupper("MiXeD cAsE 123") returns "MIXED CASE
          123".

Matching the Null String In awk, the ‘*’ operator can match the null string. This is particularly important for the

sub(), gsub(), and gensub() functions. For example:

      $ echo abc | awk '{ gsub(/m*/, "X"); print }'
      ⊣ XaXbXcX

Although this makes a certain amount of sense, it can be surprising.

9.1.3.1 More about ‘\’ and ‘&’ with sub(), gsub(), and gensub()

CAUTION: This subsubsection has been reported to cause headaches. You might want to skip it upon first reading.

When using sub(), gsub(), or gensub(), and trying to get literal backslashes and ampersands into the replacement text, you need to remember that there are several levels of escape processing going on.

First, there is the lexical level, which is when awk reads your program and builds an internal copy of it to execute. Then there is the runtime level, which is when awk actually scans the replacement string to determine what to generate.

At both levels, awk looks for a defined set of characters that can come after a backslash. At the lexical level, it looks for the escape sequences listed in Section 3.2 [Escape Sequences], page 46. Thus, for every ‘\’ that awk processes at the runtime level, you must type two backslashes at the lexical level. When a character that is not valid for an escape sequence follows the ‘\’, BWK awk and gawk both simply remove the initial ‘\’ and put the next character into the string. Thus, for example, "a\qb" is treated as "aqb".

At the runtime level, the various functions handle sequences of ‘\’ and ‘&’ differently. The situation is (sadly) somewhat complex. Historically, the sub() and gsub() functions treated the two-character sequence ‘\&’ specially; this sequence was replaced in the generated text with a single ‘&’. Any other ‘\’ within the replacement string that did not precede an ‘&’ was passed through unchanged. This is illustrated in Table 9.1.

You type  sub() sees  sub() generates
\&            &  The matched text
\\&           \&  A literal ‘&’
\\\&           \&  A literal ‘&’

   \\\\&          \\&  A literal ‘\&’
  \\\\\&          \\&  A literal ‘\&’
 \\\\\\&         \\\&  A literal ‘\\&’

\\q           \q  A literal ‘\q’

Table 9.1: Historical escape sequence processing for sub() and gsub()

This table shows the lexical-level processing, where an odd number of backslashes becomes an even number at the runtime level, as well as the runtime processing done by sub(). (For the sake of simplicity, the rest of the following tables only show the case of even numbers of backslashes entered at the lexical level.)

The problem with the historical approach is that there is no way to get a literal ‘\’ followed by the matched text.

Several editions of the POSIX standard attempted to fix this problem but weren’t successful. The details are irrelevant at this point in time.

At one point, the gawk maintainer submitted proposed text for a revised standard that reverts to rules that correspond more closely to the original existing practice. The proposed rules have special cases that make it possible to produce a ‘\’ preceding the matched text. This is shown in Table 9.2.

You type  sub() sees  sub() generates

 \\\\\\&         \\\&  A literal ‘\&’

\\\\&          \\&  A literal ‘\’, followed by the matched text
\\&           \&  A literal ‘&’
\\q           \q  A literal ‘\q’

    \\\\           \\    \\

Table 9.2: gawk rules for sub() and backslash

In a nutshell, at the runtime level, there are now three special sequences of characters (‘\\\&’, ‘\\&’, and ‘\&’) whereas historically there was only one. However, as in the historical case, any ‘\’ that is not part of one of these three sequences is not special and appears in the output literally.

gawk 3.0 and 3.1 follow these rules for sub() and gsub(). The POSIX standard took much longer to be revised than was expected. In addition, the gawk maintainer’s proposal was lost during the standardization process. The final rules are somewhat simpler. The results are similar except for one case.

The POSIX rules state that ‘\&’ in the replacement string produces a literal ‘&’, ‘\\’ produces a literal ‘\’, and ‘\’ followed by anything else is not special; the ‘\’ is placed straight into the output. These rules are presented in Table 9.3.

You type  sub() sees  sub() generates

 \\\\\\&         \\\&  A literal ‘\&’

\\\\&          \\&  A literal ‘\’, followed by the matched text
\\&           \&  A literal ‘&’
\\q           \q  A literal ‘\q’

    \\\\           \\    \

Table 9.3: POSIX rules for sub() and gsub()

The only case where the difference is noticeable is the last one: ‘\\\\’ is seen as ‘\\’ and produces ‘\’ instead of ‘\\’.

Starting with version 3.1.4, gawk followed the POSIX rules when --posix was specified (see Section 2.2 [Command-Line Options], page 31). Otherwise, it continued to follow the proposed rules, as that had been its behavior for many years.

When version 4.0.0 was released, the gawk maintainer made the POSIX rules the default, breaking well over a decade’s worth of backward compatibility.^{7} Needless to say, this was a bad idea, and as of version 4.0.1, gawk resumed its historical behavior, and only follows the POSIX rules when --posix is given.

The rules for gensub() are considerably simpler. At the runtime level, whenever gawk sees a ‘\’, if the following character is a digit, then the text that matched the corresponding parenthesized subexpression is placed in the generated output. Otherwise, no matter what character follows the ‘\’, it appears in the generated text and the ‘\’ does not, as shown in Table 9.4.

You type  gensub() sees  gensub() generates
&               &  The matched text
\\&              \&  A literal ‘&’
\\\\              \\  A literal ‘\’
\\\\&             \\&  A literal ‘\’, then the matched text

 \\\\\\&            \\\&  A literal ‘\&’

\\q              \q  A literal ‘q’

Table 9.4: Escape sequence processing for gensub()

7 This was rather naive of him, despite there being a note in this section indicating that the next major

version would move to the POSIX rules.

Because of the complexity of the lexical- and runtime-level processing and the special cases for sub() and gsub(), we recommend the use of gawk and gensub() when you have to do substitutions.

9.1.4 Input/Output Functions

The following functions relate to input/output (I/O). Optional parameters are enclosed in square brackets ([ ]):

close(filename [, how])

Close the file filename for input or output. Alternatively, the argument may be a shell command that was used for creating a coprocess, or for redirecting to or from a pipe; then the coprocess or pipe is closed. See Section 5.9 [Closing Input and Output Redirections], page 102, for more information. When closing a coprocess, it is occasionally useful to first close one end of the two-way pipe and then to close the other. This is done by providing a second argument to close(). This second argument (how) should be one of the two string values "to" or "from", indicating which end of the pipe to close. Case in the string does not matter. See Section 12.3 [Two-Way Communications with Another Process], page 305, which discusses this feature in more detail and gives an example. Note that the second argument to close() is a gawk extension; it is not available in compatibility mode (see Section 2.2 [Command-Line Options], page 31).

fflush([filename])

Flush any buffered output associated with filename, which is either a file opened for writing or a shell command for redirecting output to a pipe or coprocess. Many utility programs buffer their output (i.e., they save information to write to a disk file or the screen in memory until there is enough for it to be worthwhile to send the data to the output device). This is often more efficient than writing every little bit of information as soon as it is ready. However, sometimes it is necessary to force a program to flush its buffers (i.e., write the information to its destination, even if a buffer is not full). This is the purpose of the fflush() function—gawk also buffers its output, and the fflush() function forces gawk to flush its buffers. Brian Kernighan added fflush() to his awk in April 1992. For two decades, it was a common extension. In December 2012, it was accepted for inclusion into the POSIX standard. See the Austin Group website (http://

austingroupbugs.net/view.php?id=634).

POSIX standardizes fflush() as follows: if there is no argument, or if the argument is the null string (""), then awk flushes the buffers for all open output files and pipes.

NOTE: Prior to version 4.0.2, gawk would flush only the standard output if there was no argument, and flush all output files and pipes if the argument was the null string. This was changed in order to be compatible with Brian Kernighan’s awk, in the hope that standardizing this feature in POSIX would then be easier (which indeed proved to be the case).

With gawk, you can use ‘fflush("/dev/stdout")’ if you wish to flush only the standard output.

fflush() returns zero if the buffer is successfully flushed; otherwise, it returns a nonzero value. (gawk returns −1.) In the case where all buffers are flushed, the return value is zero only if all buffers were flushed successfully. Otherwise, it is −1, and gawk warns about the problem filename. gawk also issues a warning message if you attempt to flush a file or pipe that was opened for reading (such as with getline), or if filename is not an open file, pipe, or coprocess. In such a case, fflush() returns −1, as well.

Interactive Versus Noninteractive Buffering

As a side point, buffering issues can be even more confusing if your program is interactive (i.e., communicating with a user sitting at a keyboard).^{8} Interactive programs generally line buffer their output (i.e., they write out every line). Noninteractive programs wait until they have a full buffer, which may be many lines of output. Here is an example of the difference:

      $ awk '{ print $1 + $2 }'
      1 1
      ⊣ 2
      2 3
      ⊣ 5
      Ctrl-d

Each line of output is printed immediately. Compare that behavior with this example:

      $ awk '{ print $1 + $2 }' | cat
      1 1
      2 3
      Ctrl-d
      ⊣ 2
      ⊣ 5

Here, no output is printed until after the Ctrl-d is typed, because it is all buffered and sent down the pipe to cat in one shot.

system(command)

Execute the operating system command command and then return to the awk program. Return command’s exit status. For example, if the following fragment of code is put in your awk program:

     END {
          system("date | mail -s 'awk run done' root")
     }

the system administrator is sent mail when the awk program finishes processing input and begins its end-of-input processing.

8 A program is interactive if the standard output is connected to a terminal device. On modern systems,

this means your keyboard and screen.

Note that redirecting print or printf into a pipe is often enough to accomplish your task. If you need to run many commands, it is more efficient to simply print them down a pipeline to the shell:

              while (more stuff to do)
                  print command | "/bin/sh"
              close("/bin/sh")

However, if your awk program is interactive, system() is useful for running large self-contained programs, such as a shell or an editor. Some operating systems cannot implement the system() function. system() causes a fatal error if it is not supported.

NOTE: When --sandbox is specified, the system() function is disabled (see Section 2.2 [Command-Line Options], page 31).

Controlling Output Buffering with system()

The fflush() function provides explicit control over output buffering for individual files and pipes. However, its use is not portable to many older awk implementations. An alternative method to flush output buffers is to call system() with a null string as its argument:

     system("")   # flush output

gawk treats this use of the system() function as a special case and is smart enough not to run a shell (or other command interpreter) with the empty command. Therefore, with gawk, this idiom is not only useful, it is also efficient. Although this method should work with other awk implementations, it does not necessarily avoid starting an unnecessary shell. (Other implementations may only flush the buffer associated with the standard output and not necessarily all buffered output.)

If you think about what a programmer expects, it makes sense that system() should flush any pending output. The following program:

     BEGIN {
          print "first print"
          system("echo system echo")
          print "second print"
     }

must print:

     first print
     system echo
     second print

and not:

     system echo
     first print
     second print

If awk did not flush its buffers before calling system(), you would see the latter (undesirable) output.

9.1.5 Time Functions

awk programs are commonly used to process log files containing timestamp information, indicating when a particular log record was written. Many programs log their timestamps in the form returned by the time() system call, which is the number of seconds since a particular epoch. On POSIX-compliant systems, it is the number of seconds since 1970-01-01 00:00:00 UTC, not counting leap seconds.^{9} All known POSIX-compliant systems support timestamps from 0 through 2^{31} − 1, which is sufficient to represent times through 2038-01-19 03:14:07 UTC. Many systems support a wider range of timestamps, including negative timestamps that represent times before the epoch.

In order to make it easier to process such log files and to produce useful reports, gawk provides the following functions for working with timestamps. They are gawk extensions; they are not specified in the POSIX standard.^{10} However, recent versions of mawk (see Section B.5 [Other Freely Available awk Implementations], page 443) also support these functions. Optional parameters are enclosed in square brackets ([ ]):

mktime(datespec)

Turn datespec into a timestamp in the same form as is returned by systime(). It is similar to the function of the same name in ISO C. The argument, datespec, is a string of the form "YYYY MM DD HH MM SS [DST]". The string consists of six or seven numbers representing, respectively, the full year including century, the month from 1 to 12, the day of the month from 1 to 31, the hour of the day from 0 to 23, the minute from 0 to 59, the second from 0 to 60,^{11} and an optional daylight-savings flag. The values of these numbers need not be within the ranges specified; for example, an hour of −1 means 1 hour before midnight. The origin-zero Gregorian calendar is assumed, with year 0 preceding year 1 and year −1 preceding year 0. The time is assumed to be in the local time zone. If the daylight-savings flag is positive, the time is assumed to be daylight savings time; if zero, the time is assumed to be standard time; and if negative (the default), mktime() attempts to determine whether daylight savings time is in effect for the specified time. If datespec does not contain enough elements or if the resulting time is out of range, mktime() returns −1.

strftime([format [, timestamp [, utc-flag] ] ])

Format the time specified by timestamp based on the contents of the format string and return the result. It is similar to the function of the same name in ISO C. If utc-flag is present and is either nonzero or non-null, the value is formatted as UTC (Coordinated Universal Time, formerly GMT or Greenwich Mean Time). Otherwise, the value is formatted for the local time zone. The timestamp is in the same format as the value returned by the systime() function. If no timestamp argument is supplied, gawk uses the current time of day as the timestamp. Without a format argument, strftime() uses the value of PROCINFO["strftime"] as the format string (see Section 7.5 [Predefined Variables], page 147). The default string value is "%a %b %e %H:%M:%S %Z %Y". This format string produces output that is equivalent to that of the date utility. You can assign a new value to PROCINFO["strftime"] to change the default format; see the following list for the various format directives.

9 See [Glossary], page 463, especially the entries “Epoch” and “UTC.”

10 The GNU date utility can also do many of the things described here. Its use may be preferable for

simple time-related operations in shell scripts.

11 Occasionally there are minutes in a year with a leap second, which is why the seconds can go up to 60.

systime()

Return the current time as the number of seconds since the system epoch. On POSIX systems, this is the number of seconds since 1970-01-01 00:00:00 UTC, not counting leap seconds. It may be a different number on other systems.

The systime() function allows you to compare a timestamp from a log file with the current time of day. In particular, it is easy to determine how long ago a particular record was logged. It also allows you to produce log records using the “seconds since the epoch” format.

The mktime() function allows you to convert a textual representation of a date and time into a timestamp. This makes it easy to do before/after comparisons of dates and times, particularly when dealing with date and time data coming from an external source, such as a log file.

The strftime() function allows you to easily turn a timestamp into human-readable information. It is similar in nature to the sprintf() function (see Section 9.1.3 [String-Manipulation Functions], page 177), in that it copies nonformat specification characters verbatim to the returned string, while substituting date and time values for format specifications in the format string.

strftime() is guaranteed by the 1999 ISO C standard^{12} to support the following date format specifications:

%a The locale’s abbreviated weekday name.

%A The locale’s full weekday name.

%b The locale’s abbreviated month name.

%B The locale’s full month name.

%c The locale’s “appropriate” date and time representation. (This is ‘%A %B %d %T %Y’ in the "C" locale.)

%C The century part of the current year. This is the year divided by 100 and truncated to the next lower integer.

%d The day of the month as a decimal number (01–31).

%D Equivalent to specifying ‘%m/%d/%y’.

%e The day of the month, padded with a space if it is only one digit.

%F Equivalent to specifying ‘%Y-%m-%d’. This is the ISO 8601 date format.

%g The year modulo 100 of the ISO 8601 week number, as a decimal number (00– 99). For example, January 1, 2012, is in week 53 of 2011. Thus, the year of its ISO 8601 week number is 2011, even though its year is 2012. Similarly, December 31, 2012, is in week 1 of 2013. Thus, the year of its ISO week number is 2013, even though its year is 2012.

12 Unfortunately, not every system’s strftime() necessarily supports all of the conversions listed here.

%G The full year of the ISO week number, as a decimal number. %h Equivalent to ‘%b’. %H The hour (24-hour clock) as a decimal number (00–23). %I The hour (12-hour clock) as a decimal number (01–12). %j The day of the year as a decimal number (001–366). %m The month as a decimal number (01–12). %M The minute as a decimal number (00–59). %n A newline character (ASCII LF). %p The locale’s equivalent of the AM/PM designations associated with a 12-hour

clock.

%r The locale’s 12-hour clock time. (This is ‘%I:%M:%S %p’ in the "C" locale.) %R Equivalent to specifying ‘%H:%M’. %S The second as a decimal number (00–60). %t A TAB character. %T Equivalent to specifying ‘%H:%M:%S’. %u The weekday as a decimal number (1–7). Monday is day one. %U The week number of the year (with the first Sunday as the first day of week

one) as a decimal number (00–53).

%V The week number of the year (with the first Monday as the first day of week one) as a decimal number (01–53). The method for determining the week number is as specified by ISO 8601. (To wit: if the week containing January 1 has four or more days in the new year, then it is week one; otherwise it is week 53 of the previous year and the next week is week one.)

%w The weekday as a decimal number (0–6). Sunday is day zero. %W The week number of the year (with the first Monday as the first day of week

one) as a decimal number (00–53).

%x The locale’s “appropriate” date representation. (This is ‘%A %B %d %Y’ in the

"C" locale.)

%X The locale’s “appropriate” time representation. (This is ‘%T’ in the "C" locale.) %y The year modulo 100 as a decimal number (00–99). %Y The full year as a decimal number (e.g., 2015). %z The time zone offset in a ‘+HHMM’ format (e.g., the format necessary to produce

RFC 822/RFC 1036 date headers).

%Z The time zone name or abbreviation; no characters if no time zone is determinable.

%Ec %EC %Ex %EX %Ey %EY %Od %Oe %OH
%OI %Om %OM %OS %Ou %OU %OV %Ow %OW %Oy

“Alternative representations” for the specifications that use only the second letter (‘%c’, ‘%C’, and so on).^{13} (These facilitate compliance with the POSIX date utility.)

%% A literal ‘%’.

If a conversion specifier is not one of those just listed, the behavior is undefined.^{14} For systems that are not yet fully standards-compliant, gawk supplies a copy

of strftime() from the GNU C Library. It supports all of the just-listed format specifications. If that version is used to compile gawk (see Appendix B [Installing gawk], page 429), then the following additional format specifications are available:

%k The hour (24-hour clock) as a decimal number (0–23). Single-digit numbers are padded with a space.

%l The hour (12-hour clock) as a decimal number (1–12). Single-digit numbers are padded with a space.

%s The time as a decimal timestamp in seconds since the epoch.

Additionally, the alternative representations are recognized but their normal representations are used.

The following example is an awk implementation of the POSIX date utility. Normally, the date utility prints the current date and time of day in a well-known format. However, if you provide an argument to it that begins with a ‘+’, date copies nonformat specifier characters to the standard output and interprets the current time according to the format specifiers in the string. For example:

     $ date '+Today is %A, %B %d, %Y.'
     ⊣ Today is Monday, September 22, 2014.

Here is the gawk version of the date utility. It has a shell “wrapper” to handle the -u option, which requires that date run as if the time zone is set to UTC:

     #! /bin/sh
     #
     # date --- approximate the POSIX 'date' command
     case $1 in
     -u)  TZ=UTC0     # use UTC
          export TZ
          shift ;;

13 If you don’t understand any of this, don’t worry about it; these facilities are meant to make it easier to

“internationalize” programs. Other internationalization features are described in Chapter 13 [Interna-

tionalization with gawk], page 315.

14 This is because ISO C leaves the behavior of the C version of strftime() undefined and gawk uses the

system’s version of strftime() if it’s there. Typically, the conversion specifier either does not appear in

the returned string or appears literally.

     esac
     gawk 'BEGIN  {
         format = PROCINFO["strftime"]
         exitval = 0
         if (ARGC > 2)
             exitval = 1
         else if (ARGC == 2) {
             format = ARGV[1]
             if (format ~ /^\+/)
                 format = substr(format, 2)   # remove leading +
         }
         print strftime(format)
         exit exitval
     }' "$@"

9.1.6 Bit-Manipulation Functions

I can explain it for you, but I can’t understand it for you.

—Anonymous Many languages provide the ability to perform bitwise operations on two integer num-

bers. In other words, the operation is performed on each successive pair of bits in the operands. Three common operations are bitwise AND, OR, and XOR. The operations are described in Table 9.5.

Bit operator

AND  OR  XOR
Operands  0  1  0  1  0  1
0  0  0  0  1  0  1
1  0  1  1  1  1  0

Table 9.5: Bitwise operations

As you can see, the result of an AND operation is 1 only when both bits are 1. The result of an OR operation is 1 if either bit is 1. The result of an XOR operation is 1 if either bit is 1, but not both. The next operation is the complement; the complement of 1 is 0 and the complement of 0 is 1. Thus, this operation “flips” all the bits of a given value.

Finally, two other common operations are to shift the bits left or right. For example, if you have a bit string ‘10111001’ and you shift it right by three bits, you end up with ‘00010111’.^{15} If you start over again with ‘10111001’ and shift it left by three bits, you end up with ‘11001000’. The following list describes gawk’s built-in functions that implement the bitwise operations. Optional parameters are enclosed in square brackets ([ ]):

and(v1, v2 [, ...])

Return the bitwise AND of the arguments. There must be at least two.

15 This example shows that zeros come in on the left side. For gawk, this is always true, but in some

languages, it’s possible to have the left side fill with ones.

compl(val)

Return the bitwise complement of val.

lshift(val, count)

Return the value of val, shifted left by count bits.

or(v1, v2 [, ...])

Return the bitwise OR of the arguments. There must be at least two.

rshift(val, count)

Return the value of val, shifted right by count bits.

xor(v1, v2 [, ...])

Return the bitwise XOR of the arguments. There must be at least two.

For all of these functions, first the double-precision floating-point value is converted to the widest C unsigned integer type, then the bitwise operation is performed. If the result cannot be represented exactly as a C double, leading nonzero bits are removed one by one until it can be represented exactly. The result is then converted back into a C double. (If you don’t understand this paragraph, don’t worry about it.)

Here is a user-defined function (see Section 9.2 [User-Defined Functions], page 199) that illustrates the use of these functions:

# bits2str --- turn a byte into readable ones and zeros
function bits2str(bits,        data, mask)
{
if (bits == 0)
    return "0"
mask = 1
for (; bits != 0; bits = rshift(bits, 1))
    data = (and(bits, mask) ? "1" : "0") data
while ((length(data) % 8) != 0)
    data = "0" data
return data
}
BEGIN {
printf "123 = %s\n", bits2str(123)
printf "0123 = %s\n", bits2str(0123)
printf "0x99 = %s\n", bits2str(0x99)
comp = compl(0x99)
printf "compl(0x99) = %#x = %s\n", comp, bits2str(comp)
shift = lshift(0x99, 2)
printf "lshift(0x99, 2) = %#x = %s\n", shift, bits2str(shift)
shift = rshift(0x99, 2)
printf "rshift(0x99, 2) = %#x = %s\n", shift, bits2str(shift)

     }

This program produces the following output when run:

     $ gawk -f testbits.awk
     ⊣ 123 = 01111011
     ⊣ 0123 = 01010011
     ⊣ 0x99 = 10011001
     ⊣ compl(0x99) = 0xffffff66 = 11111111111111111111111101100110
     ⊣ lshift(0x99, 2) = 0x264 = 0000001001100100
     ⊣ rshift(0x99, 2) = 0x26 = 00100110

The bits2str() function turns a binary number into a string. Initializing mask to one creates a binary value where the rightmost bit is set to one. Using this mask, the function repeatedly checks the rightmost bit. ANDing the mask with the value indicates whether the rightmost bit is one or not. If so, a "1" is concatenated onto the front of the string. Otherwise, a "0" is added. The value is then shifted right by one bit and the loop continues until there are no more one bits.

If the initial value is zero, it returns a simple "0". Otherwise, at the end, it pads the value with zeros to represent multiples of 8-bit quantities. This is typical in modern computers.

The main code in the BEGIN rule shows the difference between the decimal and octal values for the same numbers (see Section 6.1.1.2 [Octal and Hexadecimal Numbers], page 107), and then demonstrates the results of the compl(), lshift(), and rshift() functions.

9.1.7 Getting Type Information

gawk provides a single function that lets you distinguish an array from a scalar variable. This is necessary for writing code that traverses every element of an array of arrays (see Section 8.6 [Arrays of Arrays], page 171).

isarray(x)

Return a true value if x is an array. Otherwise, return false.

isarray() is meant for use in two circumstances. The first is when traversing a multidimensional array: you can test if an element is itself an array or not. The second is inside the body of a user-defined function (not discussed yet; see Section 9.2 [User-Defined Functions], page 199), to test if a parameter is an array or not.

NOTE: Using isarray() at the global level to test variables makes no sense. Because you are the one writing the program, you are supposed to know if your variables are arrays or not. And in fact, due to the way gawk works, if you pass the name of a variable that has not been previously used to isarray(), gawk ends up turning it into a scalar.

9.1.8 String-Translation Functions

gawk provides facilities for internationalizing awk programs. These include the functions described in the following list. The descriptions here are purposely brief. See Chapter 13 [Internationalization with gawk], page 315, for the full story. Optional parameters are enclosed in square brackets ([ ]):

bindtextdomain(directory [, domain])

Set the directory in which gawk will look for message translation files, in case they will not or cannot be placed in the “standard” locations (e.g., during testing). It returns the directory in which domain is “bound.” The default domain is the value of TEXTDOMAIN. If directory is the null string (""), then bindtextdomain() returns the current binding for the given domain.

dcgettext(string [, domain [, category] ])

Return the translation of string in text domain domain for locale category category. The default value for domain is the current value of TEXTDOMAIN. The default value for category is "LC_MESSAGES".

dcngettext(string1, string2, number [, domain [, category] ])

Return the plural form used for number of the translation of string1 and string2 in text domain domain for locale category category. string1 is the English singular variant of a message, and string2 is the English plural variant of the same message. The default value for domain is the current value of TEXTDOMAIN. The default value for category is "LC_MESSAGES".

9.2 User-Defined Functions

Complicated awk programs can often be simplified by defining your own functions. Userdefined functions can be called just like built-in ones (see Section 6.4 [Function Calls], page 126), but it is up to you to define them (i.e., to tell awk what they should do).

9.2.1 Function Definition Syntax

It’s entirely fair to say that the awk syntax for local variable definitions is appallingly awful.

—Brian Kernighan Definitions of functions can appear anywhere between the rules of an awk program.

Thus, the general form of an awk program is extended to include sequences of rules and user-defined function definitions. There is no need to put the definition of a function before all uses of the function. This is because awk reads the entire program before starting to execute any of it.

The definition of a function named name looks like this:

function name([parameter-list])

     {

body-of-function

     }

Here, name is the name of the function to define. A valid function name is like a valid variable name: a sequence of letters, digits, and underscores that doesn’t start with a digit. Here too, only the 52 upper- and lowercase English letters may be used in a function name. Within a single awk program, any particular name can only be used as a variable, array, or function.

parameter-list is an optional list of the function’s arguments and local variable names, separated by commas. When the function is called, the argument names are used to hold the argument values given in the call.

A function cannot have two parameters with the same name, nor may it have a parameter with the same name as the function itself.

CAUTION: According to the POSIX standard, function parameters cannot have the same name as one of the special predefined variables (see Section 7.5 [Predefined Variables], page 147), nor may a function parameter have the same name as another function. Not all versions of awk enforce these restrictions. gawk always enforces the first restriction. With --posix (see Section 2.2 [Command-Line Options], page 31), it also enforces the second restriction.

Local variables act like the empty string if referenced where a string value is required, and like zero if referenced where a numeric value is required. This is the same as the behavior of regular variables that have never been assigned a value. (There is more to understand about local variables; see Section 9.2.5 [Functions and Their Effects on Variable Typing], page 208.)

The body-of-function consists of awk statements. It is the most important part of the definition, because it says what the function should actually do. The argument names exist to give the body a way to talk about the arguments; local variables exist to give the body places to keep temporary values.

Argument names are not distinguished syntactically from local variable names. Instead, the number of arguments supplied when the function is called determines how many argument variables there are. Thus, if three argument values are given, the first three names in parameter-list are arguments and the rest are local variables.

It follows that if the number of arguments is not the same in all calls to the function, some of the names in parameter-list may be arguments on some occasions and local variables on others. Another way to think of this is that omitted arguments default to the null string.

Usually when you write a function, you know how many names you intend to use for arguments and how many you intend to use as local variables. It is conventional to place some extra space between the arguments and the local variables, in order to document how your function is supposed to be used.

During execution of the function body, the arguments and local variable values hide, or shadow, any variables of the same names used in the rest of the program. The shadowed variables are not accessible in the function definition, because there is no way to name them while their names have been taken away for the arguments and local variables. All other variables used in the awk program can be referenced or set normally in the function’s body.

The arguments and local variables last only as long as the function body is executing. Once the body finishes, you can once again access the variables that were shadowed while the function was running.

The function body can contain expressions that call functions. They can even call this function, either directly or by way of another function. When this happens, we say the function is recursive. The act of a function calling itself is called recursion.

All the built-in functions return a value to their caller. User-defined functions can do so also, using the return statement, which is described in detail in Section 9.2.4 [The return Statement], page 206. Many of the subsequent examples in this section use the return statement.

In many awk implementations, including gawk, the keyword function may be abbreviated func. (c.e.) However, POSIX only specifies the use of the keyword function. This actually has some practical implications. If gawk is in POSIX-compatibility mode (see Section 2.2 [Command-Line Options], page 31), then the following statement does not define a function:

     func foo() { a = sqrt($1) ; print a }

Instead, it defines a rule that, for each record, concatenates the value of the variable ‘func’ with the return value of the function ‘foo’. If the resulting string is non-null, the action is executed. This is probably not what is desired. (awk accepts this input as syntactically valid, because functions may be used before they are defined in awk programs.^{16})

To ensure that your awk programs are portable, always use the keyword function when defining a function.

9.2.2 Function Definition Examples

Here is an example of a user-defined function, called myprint(), that takes a number and prints it in a specific format:

     function myprint(num)
     {
          printf "%6.3g\n", num
     }

To illustrate, here is an awk rule that uses our myprint() function:

     $3 > 0     { myprint($3) }

This program prints, in our special format, all the third fields that contain a positive number in our input. Therefore, when given the following input:

      1.2   3.4    5.6   7.8
      9.10 11.12 -13.14 15.16
     17.18 19.20  21.22 23.24

this program, using our function to format the results, prints:

        5.6
       21.2

This function deletes all the elements in an array (recall that the extra whitespace signifies the start of the local variable list):

     function delarray(a,    i)
     {
         for (i in a)
             delete a[i]
     }

When working with arrays, it is often necessary to delete all the elements in an array and start over with a new list of elements (see Section 8.4 [The delete Statement], page 168). Instead of having to repeat this loop everywhere that you need to clear out an array, your program can just call delarray(). (This guarantees portability. The use of ‘delete array’ to delete the contents of an entire array is a relatively recent^{17} addition to the POSIX standard.)

16 This program won’t actually run, because foo() is undefined.

The following is an example of a recursive function. It takes a string as an input parameter and returns the string in reverse order. Recursive functions must always have a test that stops the recursion. In this case, the recursion terminates when the input string is already empty:

     function rev(str)
     {
         if (str == "")
             return ""
         return (rev(substr(str, 2)) substr(str, 1, 1))
     }

If this function is in a file named rev.awk, it can be tested this way:

     $ echo "Don't Panic!" |
     > gawk -e '{ print rev($0) }' -f rev.awk
     ⊣ !cinaP t'noD

The C ctime() function takes a timestamp and returns it as a string, formatted in a well-known fashion. The following example uses the built-in strftime() function (see Section 9.1.5 [Time Functions], page 192) to create an awk version of ctime():

     # ctime.awk
     #
     # awk version of C ctime(3) function
     function ctime(ts,    format)
     {
         format = "%a %b %e %H:%M:%S %Z %Y"
         if (ts == 0)
             ts = systime()       # use current time as default
         return strftime(format, ts)
     }

You might think that ctime() could use PROCINFO["strftime"] for its format string. That would be a mistake, because ctime() is supposed to return the time formatted in a standard fashion, and user-level code could have changed PROCINFO["strftime"].

9.2.3 Calling User-Defined Functions

Calling a function means causing the function to run and do its job. A function call is an expression and its value is the value returned by the function.

9.2.3.1 Writing a Function Call

A function call consists of the function name followed by the arguments in parentheses. awk expressions are what you write in the call for the arguments. Each time the call is executed, these expressions are evaluated, and the values become the actual arguments. For example, here is a call to foo() with three arguments (the first being a string concatenation):

17 Late in 2012.

foo(x y, "lose", 4 * z)

CAUTION: Whitespace characters (spaces and TABs) are not allowed between the function name and the opening parenthesis of the argument list. If you write whitespace by mistake, awk might think that you mean to concatenate a variable with an expression in parentheses. However, it notices that you used a function name and not a variable name, and reports an error.

9.2.3.2 Controlling Variable Scope

Unlike in many languages, there is no way to make a variable local to a { … } block in awk, but you can make a variable local to a function. It is good practice to do so whenever a variable is needed only in that function.

To make a variable local to a function, simply declare the variable as an argument after the actual function arguments (see Section 9.2.1 [Function Definition Syntax], page 199). Look at the following example, where variable i is a global variable used by both functions

foo() and bar():
function bar()
{
    for (i = 0; i < 3; i++)
        print "bar's i=" i
}
function foo(j)
{
    i = j + 1
    print "foo's i=" i
    bar()
    print "foo's i=" i
}
BEGIN {
      i = 10
      print "top's i=" i
      foo(0)
      print "top's i=" i
}

Running this script produces the following, because the i in functions foo() and bar() and at the top level refer to the same variable instance:

top's i=10
foo's i=1
bar's i=0
bar's i=1
bar's i=2
foo's i=3

top's i=3

If you want i to be local to both foo() and bar(), do as follows (the extra space before i is a coding convention to indicate that i is a local variable, not an argument):

function bar(    i)
{
    for (i = 0; i < 3; i++)
        print "bar's i=" i
}
function foo(j,    i)
{
    i = j + 1
    print "foo's i=" i
    bar()
    print "foo's i=" i
}
BEGIN {
      i = 10
      print "top's i=" i
      foo(0)
      print "top's i=" i
}

Running the corrected script produces the following:

top's i=10
foo's i=1
bar's i=0
bar's i=1
bar's i=2
foo's i=1
top's i=10

Besides scalar values (strings and numbers), you may also have local arrays. By using a parameter name as an array, awk treats it as an array, and it is local to the function. In addition, recursive calls create new arrays. Consider this example:

function some_func(p1,      a)
{
    if (p1++ > 3)
        return
    a[p1] = p1
    some_func(p1)
    printf("At level %d, index %d %s found in a\n",
         p1, (p1 - 1), (p1 - 1) in a ? "is" : "is not")

         printf("At level %d, index %d %s found in a\n",
              p1, p1, p1 in a ? "is" : "is not")
         print ""
     }
     BEGIN {
         some_func(1)
     }

When run, this program produces the following output:

     At level 4, index 3 is not found in a
     At level 4, index 4 is found in a
     At level 3, index 2 is not found in a
     At level 3, index 3 is found in a
     At level 2, index 1 is not found in a
     At level 2, index 2 is found in a

9.2.3.3 Passing Function Arguments by Value Or by Reference

In awk, when you declare a function, there is no way to declare explicitly whether the arguments are passed by value or by reference.

Instead, the passing convention is determined at runtime when the function is called, according to the following rule: if the argument is an array variable, then it is passed by reference. Otherwise, the argument is passed by value.

Passing an argument by value means that when a function is called, it is given a copy of the value of this argument. The caller may use a variable as the expression for the argument, but the called function does not know this—it only knows what value the argument had. For example, if you write the following code:

     foo = "bar"
     z = myfunc(foo)

then you should not think of the argument to myfunc() as being “the variable foo.” Instead, think of the argument as the string value "bar". If the function myfunc() alters the values of its local variables, this has no effect on any other variables. Thus, if myfunc() does this:

     function myfunc(str)
     {
        print str
        str = "zzz"
        print str
     }

to change its first argument variable str, it does not change the value of foo in the caller. The role of foo in calling myfunc() ended when its value ("bar") was computed. If str also exists outside of myfunc(), the function body cannot alter this outer value, because it is shadowed during the execution of myfunc() and cannot be seen or changed from there.

However, when arrays are the parameters to functions, they are not copied. Instead, the array itself is made available for direct manipulation by the function. This is usually termed call by reference. Changes made to an array parameter inside the body of a function are visible outside that function.

NOTE: Changing an array parameter inside a function can be very dangerous if you do not watch what you are doing. For example:

          function changeit(array, ind, nvalue)
          {
               array[ind] = nvalue
          }
          BEGIN {
              a[1] = 1; a[2] = 2; a[3] = 3
              changeit(a, 2, "two")
              printf "a[1] = %s, a[2] = %s, a[3] = %s\n",
                      a[1], a[2], a[3]
          }
     prints ‘a[1] = 1, a[2] = two, a[3] = 3’, because changeit() stores "two" in

the second element of a. Some awk implementations allow you to call a function that has not been defined. They

only report a problem at runtime, when the program actually tries to call the function. For example:

     BEGIN {
         if (0)
             foo()
         else
             bar()
     }
     function bar() { ... }
     # note that `foo' is not defined

Because the ‘if’ statement will never be true, it is not really a problem that foo() has not been defined. Usually, though, it is a problem if a program calls an undefined function.

If --lint is specified (see Section 2.2 [Command-Line Options], page 31), gawk reports calls to undefined functions.

Some awk implementations generate a runtime error if you use either the next statement or the nextfile statement (see Section 7.4.8 [The next Statement], page 144, and see Section 7.4.9 [The nextfile Statement], page 145) inside a user-defined function. gawk does not have this limitation.

9.2.4 The return Statement

As seen in several earlier examples, the body of a user-defined function can contain a return statement. This statement returns control to the calling part of the awk program. It can also be used to return a value for use in the rest of the awk program. It looks like this:

return [expression] The expression part is optional. Due most likely to an oversight, POSIX does not define

what the return value is if you omit the expression. Technically speaking, this makes the returned value undefined, and therefore, unpredictable. In practice, though, all versions of awk simply return the null string, which acts like zero if used in a numeric context.

A return statement without an expression is assumed at the end of every function definition. So, if control reaches the end of the function body, then technically the function returns an unpredictable value. In practice, it returns the empty string. awk does not warn you if you use the return value of such a function.

Sometimes, you want to write a function for what it does, not for what it returns. Such a function corresponds to a void function in C, C++, or Java, or to a procedure in Ada. Thus, it may be appropriate to not return any value; simply bear in mind that you should not be using the return value of such a function.

The following is an example of a user-defined function that returns a value for the largest number among the elements of an array:

     function maxelt(vec,   i, ret)
     {
          for (i in vec) {
               if (ret == "" || vec[i] > ret)
                    ret = vec[i]
          }
          return ret
     }

You call maxelt() with one argument, which is an array name. The local variables i and ret are not intended to be arguments; there is nothing to stop you from passing more than one argument to maxelt() but the results would be strange. The extra space before i in the function parameter list indicates that i and ret are local variables. You should follow this convention when defining functions.

The following program uses the maxelt() function. It loads an array, calls maxelt(), and then reports the maximum number in that array:

     function maxelt(vec,   i, ret)
     {
          for (i in vec) {
               if (ret == "" || vec[i] > ret)
                    ret = vec[i]
          }
          return ret
     }
     # Load all fields of each record into nums.
     {
          for(i = 1; i <= NF; i++)
               nums[NR, i] = $i
     }
     END {
          print maxelt(nums)
     }

Given the following input:

      1 5 23 8 16
     44 3 5 2 8 26
     256 291 1396 2962 100
     -6 467 998 1101
     99385 11 0 225

the program reports (predictably) that 99,385 is the largest value in the array.

9.2.5 Functions and Their Effects on Variable Typing

awk is a very fluid language. It is possible that awk can’t tell if an identifier represents a scalar variable or an array until runtime. Here is an annotated sample program:

     function foo(a)
     {
         a[1] = 1   # parameter is an array
     }
     BEGIN {
         b = 1
         foo(b)  # invalid: fatal type mismatch
         foo(x)  # x uninitialized, becomes an array dynamically
         x = 1   # now not allowed, runtime error
     }

In this example, the first call to foo() generates a fatal error, so awk will not report the second error. If you comment out that call, though, then awk does report the second error.

Usually, such things aren’t a big issue, but it’s worth being aware of them.

9.3 Indirect Function Calls

This section describes an advanced, gawk-specific extension.

Often, you may wish to defer the choice of function to call until runtime. For example, you may have different kinds of records, each of which should be processed differently.

Normally, you would have to use a series of if-else statements to decide which function to call. By using indirect function calls, you can specify the name of the function to call as a string variable, and then call the function. Let’s look at an example.

Suppose you have a file with your test scores for the classes you are taking, and you wish to get the sum and the average of your test scores. The first field is the class name. The following fields are the functions to call to process the data, up to a “marker” field ‘data:’. Following the marker, to the end of the record, are the various numeric test scores.

Here is the initial file:

     Biology_101 sum average data: 87.0 92.4 78.5 94.9
     Chemistry_305 sum average data: 75.2 98.3 94.7 88.2
     English_401 sum average data: 100.0 95.6 87.1 93.4

To process the data, you might write initially:

{
    class = $1
    for (i = 2; $i != "data:"; i++) {
        if ($i == "sum")
            sum()   # processes the whole record
        else if ($i == "average")
            average()
        ...           # and so on
    }
}

This style of programming works, but can be awkward. With indirect function calls, you tell gawk to use the value of a variable as the name of the function to call.

The syntax is similar to that of a regular function call: an identifier immediately followed by an opening parenthesis, any arguments, and then a closing parenthesis, with the addition of a leading ‘@’ character:

the_func = "sum"
result = @the_func()   # calls the sum() function

Here is a full program that processes the previously shown data, using indirect function calls:

# indirectcall.awk --- Demonstrate indirect function calls
# average --- return the average of the values in fields $first - $last
function average(first, last,   sum, i)
{
    sum = 0;
    for (i = first; i <= last; i++)
        sum += $i
    return sum / (last - first + 1)
}
# sum --- return the sum of the values in fields $first - $last
function sum(first, last,   ret, i)
{
    ret = 0;
    for (i = first; i <= last; i++)
        ret += $i
    return ret
}

These two functions expect to work on fields; thus, the parameters first and last indicate where in the fields to start and end. Otherwise, they perform the expected computations and are not unusual:

     # For each record, print the class name and the requested statistics
     {
         class_name = $1
         gsub(/_/, " ", class_name)  # Replace _with spaces
         # find start
         for (i = 1; i <= NF; i++) {
             if ($i == "data:") {
                 start = i + 1
                 break
             }
         }
         printf("%s:\n", class_name)
         for (i = 2; $i != "data:"; i++) {
             the_function = $i
             printf("\t%s: <%s>\n", $i, @the_function(start, NF) "")
         }
         print ""
     }

This is the main processing for each record. It prints the class name (with underscores replaced with spaces). It then finds the start of the actual data, saving it in start. The last part of the code loops through each function name (from $2 up to the marker, ‘data:’), calling the function named by the field. The indirect function call itself occurs as a parameter in the call to printf. (The printf format string uses ‘%s’ as the format specifier so that we can use functions that return strings, as well as numbers. Note that the result from the indirect call is concatenated with the empty string, in order to force it to be a string value.)

Here is the result of running the program:

     $ gawk -f indirectcall.awk class_data1
     ⊣ Biology 101:
     ⊣  sum: <352.8>
     ⊣  average: <88.2>

⊣

     ⊣ Chemistry 305:
     ⊣  sum: <356.4>
     ⊣  average: <89.1>

⊣

     ⊣ English 401:
     ⊣  sum: <376.1>
     ⊣  average: <94.025>

The ability to use indirect function calls is more powerful than you may think at first. The C and C++ languages provide “function pointers,” which are a mechanism for calling a function chosen at runtime. One of the most well-known uses of this ability is the C qsort() function, which sorts an array using the famous “quicksort” algorithm (see the

Wikipedia article (http://en.wikipedia.org/wiki/Quicksort) for more information).

To use this function, you supply a pointer to a comparison function. This mechanism allows you to sort arbitrary data in an arbitrary fashion.

We can do something similar using gawk, like this:

# quicksort.awk --- Quicksort algorithm, with user-supplied
#                   comparison function
# quicksort --- C.A.R. Hoare's quicksort algorithm. See Wikipedia
#               or almost any algorithms or computer science text.
function quicksort(data, left, right, less_than,    i, last)
{
    if (left >= right)  # do nothing if array contains fewer
        return          # than two elements
    quicksort_swap(data, left, int((left + right) / 2))
    last = left
    for (i = left + 1; i <= right; i++)
        if (@less_than(data[i], data[left]))
            quicksort_swap(data, ++last, i)
    quicksort_swap(data, left, last)
    quicksort(data, left, last - 1, less_than)
    quicksort(data, last + 1, right, less_than)
}
# quicksort_swap --- helper function for quicksort, should really be inline
function quicksort_swap(data, i, j,      temp)
{
    temp = data[i]
    data[i] = data[j]
    data[j] = temp
}

The quicksort() function receives the data array, the starting and ending indices to sort (left and right), and the name of a function that performs a “less than” comparison. It then implements the quicksort algorithm.

To make use of the sorting function, we return to our previous example. The first thing to do is write some comparison functions:

# num_lt --- do a numeric less than comparison
function num_lt(left, right)
{
    return ((left + 0) < (right + 0))
}
# num_ge --- do a numeric greater than or equal to comparison

function num_ge(left, right)
{
    return ((left + 0) >= (right + 0))
}

The num_ge() function is needed to perform a descending sort; when used to perform a “less than” test, it actually does the opposite (greater than or equal to), which yields data sorted in descending order.

Next comes a sorting function. It is parameterized with the starting and ending field numbers and the comparison function. It builds an array with the data and calls quicksort() appropriately, and then formats the results as a single string:

# do_sort --- sort the data according to `compare'
#             and return it as a string
function do_sort(first, last, compare,      data, i, retval)
{
    delete data
    for (i = 1; first <= last; first++) {
        data[i] = $first
        i++
    }
    quicksort(data, 1, i-1, compare)
    retval = data[1]
    for (i = 2; i in data; i++)
        retval = retval " " data[i]
    return retval
}

Finally, the two sorting functions call do_sort(), passing in the names of the two comparison functions:

# sort --- sort the data in ascending order and return it as a string
function sort(first, last)
{
    return do_sort(first, last, "num_lt")
}
# rsort --- sort the data in descending order and return it as a string
function rsort(first, last)
{
    return do_sort(first, last, "num_ge")
}

Here is an extended version of the data file:

Biology_101 sum average sort rsort data: 87.0 92.4 78.5 94.9
Chemistry_305 sum average sort rsort data: 75.2 98.3 94.7 88.2
English_401 sum average sort rsort data: 100.0 95.6 87.1 93.4

Finally, here are the results when the enhanced program is run:

$ gawk -f quicksort.awk -f indirectcall.awk class_data2
⊣ Biology 101:
⊣  sum: <352.8>
⊣  average: <88.2>
⊣  sort: <78.5 87.0 92.4 94.9>
⊣  rsort: <94.9 92.4 87.0 78.5>

⊣

⊣ Chemistry 305:
⊣  sum: <356.4>
⊣  average: <89.1>
⊣  sort: <75.2 88.2 94.7 98.3>
⊣  rsort: <98.3 94.7 88.2 75.2>

⊣

⊣ English 401:
⊣  sum: <376.1>
⊣  average: <94.025>
⊣  sort: <87.1 93.4 95.6 100.0>
⊣  rsort: <100.0 95.6 93.4 87.1>

Another example where indirect functions calls are useful can be found in processing arrays. This is described in Section 10.7 [Traversing Arrays of Arrays], page 246.

Remember that you must supply a leading ‘@’ in front of an indirect function call.

Starting with version 4.1.2 of gawk, indirect function calls may also be used with builtin functions and with extension functions (see Chapter 16 [Writing Extensions for gawk], page 355). There are some limitations when calling built-in functions indirectly, as follows.

• You cannot pass a regular expression constant to a built-in function through an indirect function call.^{18} This applies to the sub(), gsub(), gensub(), match(), split() and

patsplit() functions.

• If calling sub() or gsub(), you may only pass two arguments, since those functions are unusual in that they update their third argument. This means that $0 will be updated.

gawk does its best to make indirect function calls efficient. For example, in the following case:

for (i = 1; i <= n; i++)
    @the_func()

gawk looks up the actual function to call only once.

18 This may change in a future version; recheck the documentation that comes with your version of gawk

to see if it has.

9.4 Summary

• awk provides built-in functions and lets you define your own functions.

• POSIX awk provides three kinds of built-in functions: numeric, string, and I/O. gawk provides functions that sort arrays, work with values representing time, do bit manipulation, determine variable type (array versus scalar), and internationalize and localize programs. gawk also provides several extensions to some of standard functions, typically in the form of additional arguments.

• Functions accept zero or more arguments and return a value. The expressions that provide the argument values are completely evaluated before the function is called. Order of evaluation is not defined. The return value can be ignored.

• The handling of backslash in sub() and gsub() is not simple. It is more straightforward in gawk’s gensub() function, but that function still requires care in its use.

• User-defined functions provide important capabilities but come with some syntactic inelegancies. In a function call, there cannot be any space between the function name and the opening left parenthesis of the argument list. Also, there is no provision for local variables, so the convention is to add extra parameters, and to separate them visually from the real parameters by extra whitespace.

• User-defined functions may call other user-defined (and built-in) functions and may call themselves recursively. Function parameters “hide” any global variables of the same names. You cannot use the name of a reserved variable (such as ARGC) as the name of a parameter in user-defined functions.

• Scalar values are passed to user-defined functions by value. Array parameters are passed by reference; any changes made by the function to array parameters are thus visible after the function has returned.

• Use the return statement to return from a user-defined function. An optional expression becomes the function’s return value. Only scalar values may be returned by a function.

• If a variable that has never been used is passed to a user-defined function, how that function treats the variable can set its nature: either scalar or array.

• gawk provides indirect function calls using a special syntax. By setting a variable to the name of a function, you can determine at runtime what function will be called at that point in the program. This is equivalent to function pointers in C and C++.

Part II: Problem Solving with awk

Chapter 10: A Library of awk Functions 217

10 A Library of awk Functions

Section 9.2 [User-Defined Functions], page 199, describes how to write your own awk functions. Writing functions is important, because it allows you to encapsulate algorithms and program tasks in a single place. It simplifies programming, making program development more manageable and making programs more readable.

In their seminal 1976 book, Software Tools,^{1} Brian Kernighan and P.J. Plauger wrote: Good Programming is not learned from generalities, but by seeing how significant programs can be made clean, easy to read, easy to maintain and modify, human-engineered, efficient and reliable, by the application of common sense and good programming practices. Careful study and imitation of good programs leads to better writing.

In fact, they felt this idea was so important that they placed this statement on the cover of their book. Because we believe strongly that their statement is correct, this chapter and Chapter 11 [Practical awk Programs], page 251, provide a good-sized body of code for you to read and, we hope, to learn from.

This chapter presents a library of useful awk functions. Many of the sample programs presented later in this book use these functions. The functions are presented here in a progression from simple to complex.

Section 11.3.7 [Extracting Programs from Texinfo Source Files], page 281, presents a program that you can use to extract the source code for these example library functions and programs from the Texinfo source for this book. (This has already been done as part of the gawk distribution.)

If you have written one or more useful, general-purpose awk functions and would like to contribute them to the awk user community, see [How to Contribute], page 11, for more information.

The programs in this chapter and in Chapter 11 [Practical awk Programs], page 251, freely use gawk-specific features. Rewriting these programs for different implementations of awk is pretty straightforward:

• Diagnostic error messages are sent to /dev/stderr. Use ‘| "cat 1>&2"’ instead of ‘> "/dev/stderr"’ if your system does not have a /dev/stderr, or if you cannot use

    gawk.

• A number of programs use nextfile (see Section 7.4.9 [The nextfile Statement], page 145) to skip any remaining input in the input file.

• Finally, some of the programs choose to ignore upper- and lowercase distinctions in their input. They do so by assigning one to IGNORECASE. You can achieve almost the same effect^{2} by adding the following rule to the beginning of the program:

         # ignore case
         { $0 = tolower($0) }

Also, verify that all regexp and string constants used in comparisons use only lowercase letters.

1 Sadly, over 35 years later, many of the lessons taught by this book have yet to be learned by a vast

number of practicing programmers.

2 The effects are not identical. Output of the transformed record will be in all lowercase, while IGNORECASE

preserves the original contents of the input record.

10.1 Naming Library Function Global Variables

Due to the way the awk language evolved, variables are either global (usable by the entire program) or local (usable just by a specific function). There is no intermediate state analogous to static variables in C.

Library functions often need to have global variables that they can use to preserve state information between calls to the function—for example, getopt()’s variable _opti (see Section 10.4 [Processing Command-Line Options], page 233). Such variables are called private, as the only functions that need to use them are the ones in the library.

When writing a library function, you should try to choose names for your private variables that will not conflict with any variables used by either another library function or a user’s main program. For example, a name like i or j is not a good choice, because user programs often use variable names like these for their own purposes.

The example programs shown in this chapter all start the names of their private variables with an underscore (‘_’). Users generally don’t use leading underscores in their variable names, so this convention immediately decreases the chances that the variable names will be accidentally shared with the user’s program.

In addition, several of the library functions use a prefix that helps indicate what function or set of functions use the variables—for example, _pw_byname() in the user database routines (see Section 10.5 [Reading the User Database], page 238). This convention is recommended, as it even further decreases the chance of inadvertent conflict among variable names. Note that this convention is used equally well for variable names and for private function names.^{3}

As a final note on variable naming, if a function makes global variables available for use by a main program, it is a good convention to start those variables’ names with a capital letter—for example, getopt()’s Opterr and Optind variables (see Section 10.4 [Processing Command-Line Options], page 233). The leading capital letter indicates that it is global, while the fact that the variable name is not all capital letters indicates that the variable is not one of awk’s predefined variables, such as FS.

It is also important that all variables in library functions that do not need to save state are, in fact, declared local.^{4} If this is not done, the variables could accidentally be used in the user’s program, leading to bugs that are very difficult to track down:

     function lib_func(x, y,    l1, l2)
     {
         ...
         # some_var should be local but by oversight is not
         use variable some_var
         ...
     }

A different convention, common in the Tcl community, is to use a single associative array to hold the values needed by the library function(s), or “package.” This significantly decreases the number of actual global names in use. For example, the functions described in Section 10.5 [Reading the User Database], page 238, might have used array elements

3 Although all the library routines could have been rewritten to use this convention, this was not done,

in order to show how our own awk programming style has evolved and to provide some basis for this

discussion.

4 gawk’s --dump-variables command-line option is useful for verifying this.

PW_data["inited"], PW_data["total"], PW_data["count"], and PW_data["awklib"],
instead of _pw_inited, _pw_awklib, _pw_total, and _pw_count.

The conventions presented in this section are exactly that: conventions. You are not required to write your programs this way—we merely recommend that you do so.

10.2 General Programming

This section presents a number of functions that are of general programming use.

10.2.1 Converting Strings to Numbers

The strtonum() function (see Section 9.1.3 [String-Manipulation Functions], page 177) is a gawk extension. The following function provides an implementation for other versions of

awk:
# mystrtonum --- convert string to number
function mystrtonum(str,        ret, n, i, k, c)
{
if (str ~ /^0[0-7]*$/) {
# octal
n = length(str)
ret = 0
for (i = 1; i <= n; i++) {
    c = substr(str, i, 1)
    # index() returns 0 if c not in string,
    # includes c == "0"
    k = index("1234567", c)
    ret = ret * 8 + k
}
} else if (str ~ /^0[xX][[:xdigit:]]+$/) {
# hexadecimal
str = substr(str, 3)    # lop off leading 0x
n = length(str)
ret = 0
for (i = 1; i <= n; i++) {
    c = substr(str, i, 1)
    c = tolower(c)
    # index() returns 0 if c not in string,
    # includes c == "0"
    k = index("123456789abcdef", c)
    ret = ret * 16 + k
}
} else if (str ~ \
/^[-+]?([0-9]+([.][0-9]*([Ee][0-9]+)?)?|([.][0-9]+([Ee][-+]?[0-9]+)?))$/) {

             # decimal number, possibly floating point
             ret = str + 0
         } else
             ret = "NOT-A-NUMBER"
         return ret
     }
     # BEGIN {     # gawk test harness
     #     a[1] = "25"
     #     a[2] = ".31"
     #     a[3] = "0123"
     #     a[4] = "0xdeadBEEF"
     #     a[5] = "123.45"
     #     a[6] = "1.e3"
     #     a[7] = "1.32"
     #     a[8] = "1.32E2"
     #
     #     for (i = 1; i in a; i++)
     #         print a[i], strtonum(a[i]), mystrtonum(a[i])
     # }

The function first looks for C-style octal numbers (base 8). If the input string matches a regular expression describing octal numbers, then mystrtonum() loops through each character in the string. It sets k to the index in "1234567" of the current octal digit. The return value will either be the same number as the digit, or zero if the character is not there, which will be true for a ‘0’. This is safe, because the regexp test in the if ensures that only octal values are converted.

Similar logic applies to the code that checks for and converts a hexadecimal value, which starts with ‘0x’ or ‘0X’. The use of tolower() simplifies the computation for finding the correct numeric value for each hexadecimal digit.

Finally, if the string matches the (rather complicated) regexp for a regular decimal integer or floating-point number, the computation ‘ret = str + 0’ lets awk convert the value to a number.

A commented-out test program is included, so that the function can be tested with gawk and the results compared to the built-in strtonum() function.

10.2.2 Assertions

When writing large programs, it is often useful to know that a condition or set of conditions is true. Before proceeding with a particular computation, you make a statement about what you believe to be the case. Such a statement is known as an assertion. The C language provides an <assert.h> header file and corresponding assert() macro that a programmer can use to make assertions. If an assertion fails, the assert() macro arranges to print a diagnostic message describing the condition that should have been true but was not, and then it kills the program. In C, using assert() looks this:

     #include <assert.h>

int myfunc(int a, double b)
{
     assert(a <= 5 && b >= 17.1);
     ...
}

If the assertion fails, the program prints a message similar to this:

prog.c:5: assertion failed: a <= 5 && b >= 17.1

The C language makes it possible to turn the condition into a string for use in printing the diagnostic message. This is not possible in awk, so this assert() function also requires a string version of the condition that is being tested. Following is the function:

# assert --- assert that a condition is true. Otherwise, exit.
function assert(condition, string)
{
    if (! condition) {
        printf("%s:%d: assertion failed: %s\n",
            FILENAME, FNR, string) > "/dev/stderr"
        _assert_exit = 1
        exit 1
    }
}
END {
    if (_assert_exit)
        exit 1
}

The assert() function tests the condition parameter. If it is false, it prints a message to standard error, using the string parameter to describe the failed condition. It then sets the variable _assert_exit to one and executes the exit statement. The exit statement jumps to the END rule. If the END rule finds _assert_exit to be true, it exits immediately.

The purpose of the test in the END rule is to keep any other END rules from running. When an assertion fails, the program should exit immediately. If no assertions fail, then _assert_exit is still false when the END rule is run normally, and the rest of the program’s END rules execute. For all of this to work correctly, assert.awk must be the first source file read by awk. The function can be used in a program in the following way:

function myfunc(a, b)
{
     assert(a <= 5 && b >= 17.1, "a <= 5 && b >= 17.1")
     ...
}

If the assertion fails, you see a message similar to the following:

mydata:1357: assertion failed: a <= 5 && b >= 17.1

There is a small problem with this version of assert(). An END rule is automatically added to the program calling assert(). Normally, if a program consists of just a BEGIN rule, the input files and/or standard input are not read. However, now that the program has an END rule, awk attempts to read the input data files or standard input (see Section 7.1.4.1 [Startup and Cleanup Actions], page 134), most likely causing the program to hang as it waits for input.

There is a simple workaround to this: make sure that such a BEGIN rule always ends with an exit statement.

10.2.3 Rounding Numbers

The way printf and sprintf() (see Section 5.5 [Using printf Statements for Fancier Printing], page 92) perform rounding often depends upon the system’s C sprintf() subroutine. On many machines, sprintf() rounding is unbiased, which means it doesn’t always round a trailing .5 up, contrary to naive expectations. In unbiased rounding, .5 rounds to even, rather than always up, so 1.5 rounds to 2 but 4.5 rounds to 4. This means that if you are using a format that does rounding (e.g., "%.0f"), you should check what your system does. The following function does traditional rounding; it might be useful if your awk’s printf does unbiased rounding:

     # round.awk --- do normal rounding
     function round(x,   ival, aval, fraction)
     {
        ival = int(x)    # integer part, int() truncates
        # see if fractional part
        if (ival == x)   # no fraction
           return ival   # ensure no decimals
        if (x < 0) {
           aval = -x     # absolute value
           ival = int(aval)
           fraction = aval - ival
           if (fraction >= .5)
              return int(x) - 1   # -2.5 --> -3
           else
              return int(x)       # -2.3 --> -2
        } else {
           fraction = x - ival
           if (fraction >= .5)
              return ival + 1
           else
              return ival
        }
     }
     # test harness
     # { print $0, round($0) }

10.2.4 The Cliff Random Number Generator

The Cliff random number generator (http: / / mathworld . wolfram . com /

CliffRandomNumberGenerator . html) is a very simple random number generator that “passes the noise sphere test for randomness by showing no structure.” It is easily programmed, in less than 10 lines of awk code:

# cliff_rand.awk --- generate Cliff random numbers
BEGIN { _cliff_seed = 0.1 }
function cliff_rand()
{
    _cliff_seed = (100 * log(_cliff_seed)) % 1
    if (_cliff_seed < 0)
        _cliff_seed = - _cliff_seed
    return _cliff_seed
}

This algorithm requires an initial “seed” of 0.1. Each new value uses the current seed as input for the calculation. If the built-in rand() function (see Section 9.1.2 [Numeric Functions], page 176) isn’t random enough, you might try using this function instead.

10.2.5 Translating Between Characters and Numbers

One commercial implementation of awk supplies a built-in function, ord(), which takes a character and returns the numeric value for that character in the machine’s character set. If the string passed to ord() has more than one character, only the first one is used.

The inverse of this function is chr() (from the function of the same name in Pascal), which takes a number and returns the corresponding character. Both functions are written very nicely in awk; there is no real reason to build them into the awk interpreter:

# ord.awk --- do ord and chr
# Global identifiers:
#    _ord_:        numerical values indexed by characters
#    _ord_init:    function to initialize _ord_
BEGIN    { _ord_init() }
function _ord_init(    low, high, i, t)
{
    low = sprintf("%c", 7) # BEL is ascii 7
    if (low == "\a") {    # regular ascii
        low = 0
        high = 127
    } else if (sprintf("%c", 128 + 7) == "\a") {
        # ascii, mark parity
        low = 128
        high = 255

    } else {        # ebcdic(!)
        low = 0
        high = 255
    }
    for (i = low; i <= high; i++) {
        t = sprintf("%c", i)
        _ord_[t] = i
    }
}

Some explanation of the numbers used by _ord_init() is worthwhile. The most prominent character set in use today is ASCII.^{5} Although an 8-bit byte can hold 256 distinct values (from 0 to 255), ASCII only defines characters that use the values from 0 to 127.^{6} In the now distant past, at least one minicomputer manufacturer used ASCII, but with mark parity, meaning that the leftmost bit in the byte is always 1. This means that on those systems, characters have numeric values from 128 to 255. Finally, large mainframe systems use the EBCDIC character set, which uses all 256 values. There are other character sets in use on some older systems, but they are not really worth worrying about:

function ord(str,    c)
{
    # only first character is of interest
    c = substr(str, 1, 1)
    return _ord_[c]
}
function chr(c)
{
    # force c to be numeric by adding 0
    return sprintf("%c", c + 0)
}
#### test code ####
# BEGIN {
#    for (;;) {
#        printf("enter a character: ")
#        if (getline var <= 0)
#            break
#        printf("ord(%s) = %d\n", var, ord(var))
#    }
# }

5 This is changing; many systems use Unicode, a very large character set that includes ASCII as a subset.

On systems with full Unicode support, a character can occupy up to 32 bits, making simple tests such as used here prohibitively expensive.

6 ASCII has been extended in many countries to use the values from 128 to 255 for country-specific

characters. If your system uses these extensions, you can simplify _ord_init() to loop from 0 to 255.

An obvious improvement to these functions is to move the code for the _ord_init function into the body of the BEGIN rule. It was written this way initially for ease of development. There is a “test program” in a BEGIN rule, to test the function. It is commented out for production use.

10.2.6 Merging an Array into a String

When doing string processing, it is often useful to be able to join all the strings in an array into one long string. The following function, join(), accomplishes this task. It is used later in several of the application programs (see Chapter 11 [Practical awk Programs], page 251).

Good function design is important; this function needs to be general, but it should also have a reasonable default behavior. It is called with an array as well as the beginning and ending indices of the elements in the array to be merged. This assumes that the array indices are numeric—a reasonable assumption, as the array was likely created with split() (see Section 9.1.3 [String-Manipulation Functions], page 177):

     # join.awk --- join an array into a string
     function join(array, start, end, sep,    result, i)
     {
         if (sep == "")
            sep = " "
         else if (sep == SUBSEP) # magic value
            sep = ""
         result = array[start]
         for (i = start + 1; i <= end; i++)
             result = result sep array[i]
         return result
     }

An optional additional argument is the separator to use when joining the strings back together. If the caller supplies a nonempty value, join() uses it; if it is not supplied, it has a null value. In this case, join() uses a single space as a default separator for the strings. If the value is equal to SUBSEP, then join() joins the strings with no separator between them. SUBSEP serves as a “magic” value to indicate that there should be no separation between the component strings.^{7}

10.2.7 Managing the Time of Day

The systime() and strftime() functions described in Section 9.1.5 [Time Functions], page 192, provide the minimum functionality necessary for dealing with the time of day in human-readable form. Although strftime() is extensive, the control formats are not necessarily easy to remember or intuitively obvious when reading a program.

The following function, getlocaltime(), populates a user-supplied array with preformatted time information. It returns a string with the current time formatted in the same way as the date utility:

     # getlocaltime.awk --- get the time of day in a usable format

7 It would be nice if awk had an assignment operator for concatenation. The lack of an explicit operator

for concatenation makes string operations more difficult than they really need to be.

# Returns a string in the format of output of date(1)
# Populates the array argument time with individual values:
#    time["second"]       -- seconds (0 - 59)
#    time["minute"]       -- minutes (0 - 59)
#    time["hour"]         -- hours (0 - 23)
#    time["althour"]      -- hours (0 - 12)
#    time["monthday"]     -- day of month (1 - 31)
#    time["month"]        -- month of year (1 - 12)
#    time["monthname"]    -- name of the month
#    time["shortmonth"]   -- short name of the month
#    time["year"]         -- year modulo 100 (0 - 99)
#    time["fullyear"]     -- full year
#    time["weekday"]      -- day of week (Sunday = 0)
#    time["altweekday"]   -- day of week (Monday = 0)
#    time["dayname"]      -- name of weekday
#    time["shortdayname"] -- short name of weekday
#    time["yearday"]      -- day of year (0 - 365)
#    time["timezone"]     -- abbreviation of timezone name
#    time["ampm"]         -- AM or PM designation
#    time["weeknum"]      -- week number, Sunday first day
#    time["altweeknum"]   -- week number, Monday first day
function getlocaltime(time,    ret, now, i)
{
    # get time once, avoids unnecessary system calls
    now = systime()
    # return date(1)-style output
    ret = strftime("%a %b %e %H:%M:%S %Z %Y", now)
    # clear out target array
    delete time
    # fill in values, force numeric values to be
    # numeric by adding 0
    time["second"]       = strftime("%S", now) + 0
    time["minute"]       = strftime("%M", now) + 0
    time["hour"]         = strftime("%H", now) + 0
    time["althour"]      = strftime("%I", now) + 0
    time["monthday"]     = strftime("%d", now) + 0
    time["month"]        = strftime("%m", now) + 0
    time["monthname"]    = strftime("%B", now)
    time["shortmonth"]   = strftime("%b", now)
    time["year"]         = strftime("%y", now) + 0
    time["fullyear"]     = strftime("%Y", now) + 0
    time["weekday"]      = strftime("%w", now) + 0

time["altweekday"]   = strftime("%u", now) + 0
time["dayname"]      = strftime("%A", now)
time["shortdayname"] = strftime("%a", now)
time["yearday"]      = strftime("%j", now) + 0
time["timezone"]     = strftime("%Z", now)
time["ampm"]         = strftime("%p", now)
time["weeknum"]      = strftime("%U", now) + 0
time["altweeknum"]   = strftime("%W", now) + 0
return ret
}

The string indices are easier to use and read than the various formats required by strftime(). The alarm program presented in Section 11.3.2 [An Alarm Clock Program], page 272, uses this function. A more general design for the getlocaltime() function would have allowed the user to supply an optional timestamp value to use instead of the current time.

10.2.8 Reading a Whole File at Once

Often, it is convenient to have the entire contents of a file available in memory as a single string. A straightforward but naive way to do that might be as follows:

function readfile(file,    tmp, contents)
{
if ((getline tmp < file) < 0)
    return
contents = tmp
while (getline tmp < file) > 0)
    contents = contents RT tmp
close(file)
return contents
}

This function reads from file one record at a time, building up the full contents of the file in the local variable contents. It works, but is not necessarily efficient.

The following function, based on a suggestion by Denis Shirokov, reads the entire contents of the named file in one shot:

# readfile.awk --- read an entire file at once
function readfile(file,     tmp, save_rs)
{
save_rs = RS
RS = "^$"
getline tmp < file
close(file)
RS = save_rs

    return tmp
}

It works by setting RS to ‘^$’, a regular expression that will never match if the file has contents. gawk reads data from the file into tmp, attempting to match RS. The match fails after each read, but fails quickly, such that gawk fills tmp with the entire contents of the file. (See Section 4.1 [How Input Is Split into Records], page 59, for information on RT and

RS.)

In the case that file is empty, the return value is the null string. Thus, calling code may use something like:

contents = readfile("/some/path")
if (length(contents) == 0)
    # file was empty ...

This tests the result to see if it is empty or not. An equivalent test would be ‘contents

== ""’.

See Section 16.7.10 [Reading an Entire File], page 406, for an extension function that also reads an entire file into memory.

10.2.9 Quoting Strings to Pass to the Shell

Michael Brennan offers the following programming pattern, which he uses frequently:

#! /bin/sh
awkp='
   ...
   '
input_program | awk "$awkp" | /bin/sh

For example, a program of his named flac-edit has this form:

$ flac-edit -song="Whoope! That's Great" file.flac

It generates the following output, which is to be piped to the shell (/bin/sh):

chmod +w file.flac
metaflac --remove-tag=TITLE file.flac
LANG=en_US.88591 metaflac --set-tag=TITLE='Whoope! That'"'"'s Great' file.flac
chmod -w file.flac

Note the need for shell quoting. The function shell_quote() does it. SINGLE is the one-character string "'" and QSINGLE is the three-character string "\"'\"":

# shell_quote --- quote an argument for passing to the shell
function shell_quote(s,             # parameter
    SINGLE, QSINGLE, i, X, n, ret)  # locals
{
    if (s == "")
        return "\"\""
    SINGLE = "\x27"  # single quote

         QSINGLE = "\"\x27\""
         n = split(s, X, SINGLE)
         ret = SINGLE X[1] SINGLE
         for (i = 2; i <= n; i++)
             ret = ret QSINGLE SINGLE X[i] SINGLE
         return ret
     }

10.3 Data file Management

This section presents functions that are useful for managing command-line data files.

10.3.1 Noting Data file Boundaries

The BEGIN and END rules are each executed exactly once, at the beginning and end of your awk program, respectively (see Section 7.1.4 [The BEGIN and END Special Patterns], page 134). We (the gawk authors) once had a user who mistakenly thought that the BEGIN rules were executed at the beginning of each data file and the END rules were executed at the end of each data file.

When informed that this was not the case, the user requested that we add new special patterns to gawk, named BEGIN_FILE and END_FILE, that would have the desired behavior. He even supplied us the code to do so.

Adding these special patterns to gawk wasn’t necessary; the job can be done cleanly in awk itself, as illustrated by the following library program. It arranges to call two usersupplied functions, beginfile() and endfile(), at the beginning and end of each data file. Besides solving the problem in only nine(!) lines of code, it does so portably; this works with any implementation of awk:

     # transfile.awk
     #
     # Give the user a hook for filename transitions
     #
     # The user must supply functions beginfile() and endfile()
     # that each take the name of the file being started or
     # finished, respectively.
     FILENAME != _oldfilename {
         if (_oldfilename != "")
             endfile(_oldfilename)
         _oldfilename = FILENAME
         beginfile(FILENAME)
     }
     END { endfile(FILENAME) }

This file must be loaded before the user’s “main” program, so that the rule it supplies is executed first.

This rule relies on awk’s FILENAME variable, which automatically changes for each new data file. The current file name is saved in a private variable, _oldfilename. If FILENAME does not equal _oldfilename, then a new data file is being processed and it is necessary to call endfile() for the old file. Because endfile() should only be called if a file has been processed, the program first checks to make sure that _oldfilename is not the null string. The program then assigns the current file name to _oldfilename and calls beginfile() for the file. Because, like all awk variables, _oldfilename is initialized to the null string, this rule executes correctly even for the first data file.

The program also supplies an END rule to do the final processing for the last file. Because this END rule comes before any END rules supplied in the “main” program, endfile() is called first. Once again, the value of multiple BEGIN and END rules should be clear.

If the same data file occurs twice in a row on the command line, then endfile() and beginfile() are not executed at the end of the first pass and at the beginning of the second pass. The following version solves the problem:

     # ftrans.awk --- handle datafile transitions
     #
     # user supplies beginfile() and endfile() functions
     FNR == 1 {
         if (_filename_!= "")
             endfile(_filename_)
         _filename_= FILENAME
         beginfile(FILENAME)
     }
     END { endfile(_filename_) }

Section 11.2.7 [Counting Things], page 269, shows how this library function can be used and how it simplifies writing the main program.

                 So Why Does gawk Have BEGINFILE and ENDFILE?

You are probably wondering, if beginfile() and endfile() functions can do the job, why does gawk have BEGINFILE and ENDFILE patterns?

Good question. Normally, if awk cannot open a file, this causes an immediate fatal error. In this case, there is no way for a user-defined function to deal with the problem, as the mechanism for calling it relies on the file being open and at the first record. Thus, the main reason for BEGINFILE is to give you a “hook” to catch files that cannot be processed. ENDFILE exists for symmetry, and because it provides an easy way to do per-file cleanup processing. For more information, refer to Section 7.1.5 [The BEGINFILE and ENDFILE Special Patterns], page 135.

10.3.2 Rereading the Current File

Another request for a new built-in function was for a function that would make it possible to reread the current file. The requesting user didn’t want to have to use getline (see Section 4.9 [Explicit Input with getline], page 78) inside a loop.

However, as long as you are not in the END rule, it is quite easy to arrange to immediately close the current input file and then start over with it from the top. For lack of a better name, we’ll call the function rewind():

# rewind.awk --- rewind the current file and start over
function rewind(    i)
{
# shift remaining arguments up
for (i = ARGC; i > ARGIND; i--)
    ARGV[i] = ARGV[i-1]
# make sure gawk knows to keep going
ARGC++
# make current file next to get done
ARGV[ARGIND+1] = FILENAME
# do it
nextfile
}

The rewind() function relies on the ARGIND variable (see Section 7.5.2 [Built-in Variables That Convey Information], page 149), which is specific to gawk. It also relies on the nextfile keyword (see Section 7.4.9 [The nextfile Statement], page 145). Because of this, you should not call it from an ENDFILE rule. (This isn’t necessary anyway, because gawk goes to the next file as soon as an ENDFILE rule finishes!)

10.3.3 Checking for Readable Data files

Normally, if you give awk a data file that isn’t readable, it stops with a fatal error. There are times when you might want to just ignore such files and keep going.^{8} You can do this by prepending the following program to your awk program:

# readable.awk --- library file to skip over unreadable files
BEGIN {
for (i = 1; i < ARGC; i++) {
    if (ARGV[i] ~ /^[a-zA-Z_][a-zA-Z0-9_]*=.*/ \
        || ARGV[i] == "-" || ARGV[i] == "/dev/stdin")
        continue    # assignment or standard input
    else if ((getline junk < ARGV[i]) < 0) # unreadable
        delete ARGV[i]
    else
        close(ARGV[i])
}
}

8 The BEGINFILE special pattern (see Section 7.1.5 [The BEGINFILE and ENDFILE Special Patterns],

page 135) provides an alternative mechanism for dealing with files that can’t be opened. However, the code here provides a portable solution.

This works, because the getline won’t be fatal. Removing the element from ARGV with delete skips the file (because it’s no longer in the list). See also Section 7.5.3 [Using ARGC and ARGV], page 155.

Because awk variable names only allow the English letters, the regular expression check purposely does not use character classes such as ‘[:alpha:]’ and ‘[:alnum:]’ (see Section 3.4 [Using Bracket Expressions], page 50).

10.3.4 Checking for Zero-Length Files

All known awk implementations silently skip over zero-length files. This is a by-product of awk’s implicit read-a-record-and-match-against-the-rules loop: when awk tries to read a record from an empty file, it immediately receives an end-of-file indication, closes the file, and proceeds on to the next command-line data file, without executing any user-level awk program code.

Using gawk’s ARGIND variable (see Section 7.5 [Predefined Variables], page 147), it is possible to detect when an empty data file has been skipped. Similar to the library file presented in Section 10.3.1 [Noting Data file Boundaries], page 229, the following library file calls a function named zerofile() that the user must provide. The arguments passed are the file name and the position in ARGV where it was found:

     # zerofile.awk --- library file to process empty input files
     BEGIN { Argind = 0 }
     ARGIND > Argind + 1 {
         for (Argind++; Argind < ARGIND; Argind++)
             zerofile(ARGV[Argind], Argind)
     }
     ARGIND != Argind { Argind = ARGIND }
     END {
         if (ARGIND > Argind)
             for (Argind++; Argind <= ARGIND; Argind++)
                 zerofile(ARGV[Argind], Argind)
     }

The user-level variable Argind allows the awk program to track its progress through ARGV. Whenever the program detects that ARGIND is greater than ‘Argind + 1’, it means that one or more empty files were skipped. The action then calls zerofile() for each such file, incrementing Argind along the way.

The ‘Argind != ARGIND’ rule simply keeps Argind up to date in the normal case.

Finally, the END rule catches the case of any empty files at the end of the command-line arguments. Note that the test in the condition of the for loop uses the ‘<=’ operator, not ‘<’.

10.3.5 Treating Assignments as File names

Occasionally, you might not want awk to process command-line variable assignments (see Section 6.1.3.2 [Assigning Variables on the Command Line], page 111). In particular, if you have a file name that contains an ‘=’ character, awk treats the file name as an assignment and does not process it.

Some users have suggested an additional command-line option for gawk to disable command-line assignments. However, some simple programming with a library file does the trick:

     # noassign.awk --- library file to avoid the need for a
     # special option that disables command-line assignments
     function disable_assigns(argc, argv,    i)
     {
         for (i = 1; i < argc; i++)
             if (argv[i] ~ /^[a-zA-Z_][a-zA-Z0-9_]*=.*/)
                 argv[i] = ("./" argv[i])
     }
     BEGIN {
         if (No_command_assign)
             disable_assigns(ARGC, ARGV)
     }

You then run your program this way:

     awk -v No_command_assign=1 -f noassign.awk -f yourprog.awk *

The function works by looping through the arguments. It prepends ‘./’ to any argument that matches the form of a variable assignment, turning that argument into a file name.

The use of No_command_assign allows you to disable command-line assignments at invocation time, by giving the variable a true value. When not set, it is initially zero (i.e., false), so the command-line arguments are left alone.

10.4 Processing Command-Line Options

Most utilities on POSIX-compatible systems take options on the command line that can be used to change the way a program behaves. awk is an example of such a program (see Section 2.2 [Command-Line Options], page 31). Often, options take arguments (i.e., data that the program needs to correctly obey the command-line option). For example, awk’s -F option requires a string to use as the field separator. The first occurrence on the command line of either -- or a string that does not begin with ‘-’ ends the options.

Modern Unix systems provide a C function named getopt() for processing commandline arguments. The programmer provides a string describing the one-letter options. If an option requires an argument, it is followed in the string with a colon. getopt() is also passed the count and values of the command-line arguments and is called in a loop. getopt() processes the command-line arguments for option letters. Each time around the loop, it returns a single character representing the next option letter that it finds, or ‘?’ if it finds an invalid option. When it returns −1, there are no options left on the command line.

When using getopt(), options that do not take arguments can be grouped together. Furthermore, options that take arguments require that the argument be present. The argument can immediately follow the option letter, or it can be a separate command-line argument.

Given a hypothetical program that takes three command-line options, -a, -b, and -c, where -b requires an argument, all of the following are valid ways of invoking the program:

     prog -a -b foo -c data1 data2 data3
     prog -ac -bfoo -- data1 data2 data3
     prog -acbfoo data1 data2 data3

Notice that when the argument is grouped with its option, the rest of the argument is considered to be the option’s argument. In this example, -acbfoo indicates that all of the -a, -b, and -c options were supplied, and that ‘foo’ is the argument to the -b option.

getopt() provides four external variables that the programmer can use:

optind The index in the argument value array (argv) where the first nonoption command-line argument can be found.

optarg The string value of the argument to an option.

opterr Usually getopt() prints an error message when it finds an invalid option. Setting opterr to zero disables this feature. (An application might want to print its own error message.)

optopt The letter representing the command-line option.

The following C fragment shows how getopt() might process command-line arguments for awk:

     int
     main(int argc, char *argv[])
     {
         ...
         /* print our own message */
         opterr = 0;
         while ((c = getopt(argc, argv, "v:f:F:W:")) != -1) {
             switch (c) {
             case 'f':    /* file */
                 ...
                 break;
             case 'F':    /* field separator */
                 ...
                 break;
             case 'v':    /* variable assignment */
                 ...
                 break;
             case 'W':    /* extension */
                 ...
                 break;
             case '?':
             default:
                 usage();

            break;
        }
    }
    ...
}

As a side point, gawk actually uses the GNU getopt_long() function to process both normal and GNU-style long options (see Section 2.2 [Command-Line Options], page 31).

The abstraction provided by getopt() is very useful and is quite handy in awk programs as well. Following is an awk version of getopt(). This function highlights one of the greatest weaknesses in awk, which is that it is very poor at manipulating single characters. Repeated calls to substr() are necessary for accessing individual characters (see Section 9.1.3 [String-Manipulation Functions], page 177).^{9}

The discussion that follows walks through the code a bit at a time:

# getopt.awk --- Do C library getopt(3) function in awk
# External variables:
#    Optind -- index in ARGV of first nonoption argument
#    Optarg -- string value of argument to current option
#    Opterr -- if nonzero, print our own diagnostic
#    Optopt -- current option letter
# Returns:
#    -1     at end of options
#    "?"    for unrecognized option
#    <c>    a character representing the current option
# Private Data:
#    _opti  -- index in multiflag option, e.g., -abc

The function starts out with comments presenting a list of the global variables it uses, what the return values are, what they mean, and any global variables that are “private” to this library function. Such documentation is essential for any program, and particularly for library functions.

The getopt() function first checks that it was indeed called with a string of options (the options parameter). If options has a zero length, getopt() immediately returns −1:

function getopt(argc, argv, options,    thisopt, i)
{
    if (length(options) == 0)    # no options given
        return -1
    if (argv[Optind] == "--") {  # all done
        Optind++
        _opti = 0
        return -1

9 This function was written before gawk acquired the ability to split strings into single characters using ""

as the separator. We have left it alone, as using substr() is more portable.

         } else if (argv[Optind] !~ /^-[^:[:space:]]/) {
             _opti = 0
             return -1
         }

The next thing to check for is the end of the options. A -- ends the command-line options, as does any command-line argument that does not begin with a ‘-’. Optind is used to step through the array of command-line arguments; it retains its value across calls to getopt(), because it is a global variable.

The regular expression that is used, /^-[^:[:space:]/, checks for a ‘-’ followed by anything that is not whitespace and not a colon. If the current command-line argument does not match this pattern, it is not an option, and it ends option processing. Continuing on:

         if (_opti == 0)
             _opti = 2
         thisopt = substr(argv[Optind], _opti, 1)
         Optopt = thisopt
         i = index(options, thisopt)
         if (i == 0) {
             if (Opterr)
                 printf("%c -- invalid option\n", thisopt) > "/dev/stderr"
             if (_opti >= length(argv[Optind])) {
                 Optind++
                 _opti = 0
             } else
                 _opti++
             return "?"
         }

The _opti variable tracks the position in the current command-line argument (argv[Optind]). If multiple options are grouped together with one ‘-’ (e.g., -abx), it is necessary to return them to the user one at a time.

If _opti is equal to zero, it is set to two, which is the index in the string of the next character to look at (we skip the ‘-’, which is at position one). The variable thisopt holds the character, obtained with substr(). It is saved in Optopt for the main program to use.

If thisopt is not in the options string, then it is an invalid option. If Opterr is nonzero, getopt() prints an error message on the standard error that is similar to the message from the C version of getopt().

Because the option is invalid, it is necessary to skip it and move on to the next option character. If _opti is greater than or equal to the length of the current command-line argument, it is necessary to move on to the next argument, so Optind is incremented and _opti is reset to zero. Otherwise, Optind is left alone and _opti is merely incremented.

In any case, because the option is invalid, getopt() returns "?". The main program can examine Optopt if it needs to know what the invalid option letter actually is. Continuing on:

         if (substr(options, i + 1, 1) == ":") {
             # get option argument

             if (length(substr(argv[Optind], _opti + 1)) > 0)
                 Optarg = substr(argv[Optind], _opti + 1)
             else
                 Optarg = argv[++Optind]
             _opti = 0
         } else
             Optarg = ""

If the option requires an argument, the option letter is followed by a colon in the options string. If there are remaining characters in the current command-line argument (argv[Optind]), then the rest of that string is assigned to Optarg. Otherwise, the next command-line argument is used (‘-xFOO’ versus ‘-x FOO’). In either case, _opti is reset to zero, because there are no more characters left to examine in the current command-line argument. Continuing:

         if (_opti == 0 || _opti >= length(argv[Optind])) {
             Optind++
             _opti = 0
         } else
             _opti++
         return thisopt
     }

Finally, if _opti is either zero or greater than the length of the current commandline argument, it means this element in argv is through being processed, so Optind is incremented to point to the next element in argv. If neither condition is true, then only _opti is incremented, so that the next option letter can be processed on the next call to

getopt().

The BEGIN rule initializes both Opterr and Optind to one. Opterr is set to one, because the default behavior is for getopt() to print a diagnostic message upon seeing an invalid option. Optind is set to one, because there’s no reason to look at the program name, which

is in ARGV[0]:
     BEGIN {
         Opterr = 1    # default is to diagnose
         Optind = 1    # skip ARGV[0]
         # test program
         if (_getopt_test) {
             while ((_go_c = getopt(ARGC, ARGV, "ab:cd")) != -1)
                 printf("c = <%c>, Optarg = <%s>\n",
                                            _go_c, Optarg)
             printf("non-option arguments:\n")
             for (; Optind < ARGC; Optind++)
                 printf("\tARGV[%d] = <%s>\n",
                                         Optind, ARGV[Optind])
         }
     }

The rest of the BEGIN rule is a simple test program. Here are the results of two sample runs of the test program:

     $ awk -f getopt.awk -v _getopt_test=1 -- -a -cbARG bax -x
     ⊣ c = <a>, Optarg = <>
     ⊣ c = <c>, Optarg = <>
     ⊣ c = <b>, Optarg = <ARG>
     ⊣ non-option arguments:
     ⊣  ARGV[3] = <bax>
     ⊣  ARGV[4] = <-x>
     $ awk -f getopt.awk -v _getopt_test=1 -- -a -x -- xyz abc
     ⊣ c = <a>, Optarg = <>
      error x -- invalid option
     ⊣ c = <?>, Optarg = <>
     ⊣ non-option arguments:
     ⊣  ARGV[4] = <xyz>
     ⊣  ARGV[5] = <abc>

In both runs, the first -- terminates the arguments to awk, so that it does not try to interpret the -a, etc., as its own options.

NOTE: After getopt() is through, user-level code must clear out all the elements of ARGV from 1 to Optind, so that awk does not try to process the command-line options as file names.

Using ‘#!’ with the -E option may help avoid conflicts between your program’s options and gawk’s options, as -E causes gawk to abandon processing of further options (see Section 1.1.4 [Executable awk Programs], page 19, and see Section 2.2 [Command-Line Options], page 31).

Several of the sample programs presented in Chapter 11 [Practical awk Programs], page 251, use getopt() to process their arguments.

10.5 Reading the User Database

The PROCINFO array (see Section 7.5 [Predefined Variables], page 147) provides access to the current user’s real and effective user and group ID numbers, and, if available, the user’s supplementary group set. However, because these are numbers, they do not provide very useful information to the average user. There needs to be some way to find the user information associated with the user and group ID numbers. This section presents a suite of functions for retrieving information from the user database. See Section 10.6 [Reading the Group Database], page 242, for a similar suite that retrieves information from the group database.

The POSIX standard does not define the file where user information is kept. Instead, it provides the <pwd.h> header file and several C language subroutines for obtaining user information. The primary function is getpwent(), for “get password entry.” The “password” comes from the original user database file, /etc/passwd, which stores user information along with the encrypted passwords (hence the name).

Although an awk program could simply read /etc/passwd directly, this file may not contain complete information about the system’s set of users.^{10} To be sure you are able to produce a readable and complete version of the user database, it is necessary to write a small C program that calls getpwent(). getpwent() is defined as returning a pointer to a struct passwd. Each time it is called, it returns the next entry in the database. When there are no more entries, it returns NULL, the null pointer. When this happens, the C program should call endpwent() to close the database. Following is pwcat, a C program that “cats” the password database:

10 It is often the case that password information is stored in a network database.

     /*
      * pwcat.c
      *
      * Generate a printable version of the password database.
      */
     #include <stdio.h>
     #include <pwd.h>
     int
     main(int argc, char **argv)
     {
         struct passwd *p;
         while ((p = getpwent()) != NULL)
             printf("%s:%s:%ld:%ld:%s:%s:%s\n",
                 p->pw_name, p->pw_passwd, (long) p->pw_uid,
                 (long) p->pw_gid, p->pw_gecos, p->pw_dir, p->pw_shell);
         endpwent();
         return 0;
     }

If you don’t understand C, don’t worry about it. The output from pwcat is the user database, in the traditional /etc/passwd format of colon-separated fields. The fields are:

Login name The user’s login name.

Encrypted password The user’s encrypted password. This may not be available on some systems.

User-ID The user’s numeric user ID number. (On some systems, it’s a C long, and not an int. Thus, we cast it to long for all cases.)

Group-ID The user’s numeric group ID number. (Similar comments about long versus int apply here.)

Full name The user’s full name, and perhaps other information associated with the user.

Home directory The user’s login (or “home”) directory (familiar to shell programmers as $HOME).

Login shell The program that is run when the user logs in. This is usually a shell, such as Bash.

A few lines representative of pwcat’s output are as follows:

$ pwcat
⊣ root:x:0:1:Operator:/:/bin/sh
⊣ nobody:*:65534:65534::/:
⊣ daemon:*:1:1::/:
⊣ sys:*:2:2::/:/bin/csh
⊣ bin:*:3:3::/bin:
⊣ arnold:xyzzy:2076:10:Arnold Robbins:/home/arnold:/bin/sh
⊣ miriam:yxaay:112:10:Miriam Robbins:/home/miriam:/bin/sh
⊣ andy:abcca2:113:10:Andy Jacobs:/home/andy:/bin/sh
...

With that introduction, following is a group of functions for getting user information. There are several functions here, corresponding to the C functions of the same names:

# passwd.awk --- access password file information
BEGIN {
# tailor this to suit your system
_pw_awklib = "/usr/local/libexec/awk/"
}
function _pw_init(    oldfs, oldrs, olddol0, pwcat, using_fw, using_fpat)
{
if (_pw_inited)
    return
oldfs = FS
oldrs = RS
olddol0 = $0
using_fw = (PROCINFO["FS"] == "FIELDWIDTHS")
using_fpat = (PROCINFO["FS"] == "FPAT")
FS = ":"
RS = "\n"
pwcat = _pw_awklib "pwcat"
while ((pwcat | getline) > 0) {
    _pw_byname[$1] = $0
    _pw_byuid[$3] = $0
    _pw_bycount[++_pw_total] = $0
}
close(pwcat)
_pw_count = 0
_pw_inited = 1
FS = oldfs
if (using_fw)
    FIELDWIDTHS = FIELDWIDTHS
else if (using_fpat)

             FPAT = FPAT
         RS = oldrs
         $0 = olddol0
     }

The BEGIN rule sets a private variable to the directory where pwcat is stored. Because it is used to help out an awk library routine, we have chosen to put it in /usr/local/libexec/awk; however, you might want it to be in a different directory on your system.

The function _pw_init() fills three copies of the user information into three associative arrays. The arrays are indexed by username (_pw_byname), by user ID number (_pw_byuid), and by order of occurrence (_pw_bycount). The variable _pw_inited is used for efficiency, as _pw_init() needs to be called only once.

Because this function uses getline to read information from pwcat, it first saves the values of FS, RS, and $0. It notes in the variable using_fw whether field splitting with FIELDWIDTHS is in effect or not. Doing so is necessary, as these functions could be called from anywhere within a user’s program, and the user may have his or her own way of splitting records and fields. This makes it possible to restore the correct field-splitting mechanism later. The test can only be true for gawk. It is false if using FS or FPAT, or on some other awk implementation.

The code that checks for using FPAT, using using_fpat and PROCINFO["FS"], is similar. The main part of the function uses a loop to read database lines, split the lines into

fields, and then store the lines into each array as necessary. When the loop is done, _pw_init() cleans up by closing the pipeline, setting _pw_inited to one, and restoring FS (and FIELDWIDTHS or FPAT if necessary), RS, and $0. The use of _pw_count is explained shortly.

The getpwnam() function takes a username as a string argument. If that user is in the database, it returns the appropriate line. Otherwise, it relies on the array reference to a nonexistent element to create the element with the null string as its value:

     function getpwnam(name)
     {
         _pw_init()
         return _pw_byname[name]
     }

Similarly, the getpwuid() function takes a user ID number argument. If that user number is in the database, it returns the appropriate line. Otherwise, it returns the null string:

     function getpwuid(uid)
     {
         _pw_init()
         return _pw_byuid[uid]
     }

The getpwent() function simply steps through the database, one entry at a time. It uses _pw_count to track its current position in the _pw_bycount array:

     function getpwent()

     {
         _pw_init()
         if (_pw_count < _pw_total)
             return _pw_bycount[++_pw_count]
         return ""
     }

The endpwent() function resets _pw_count to zero, so that subsequent calls to getpwent() start over again:

     function endpwent()
     {
         _pw_count = 0
     }

A conscious design decision in this suite is that each subroutine calls _pw_init() to initialize the database arrays. The overhead of running a separate process to generate the user database, and the I/O to scan it, are only incurred if the user’s main program actually calls one of these functions. If this library file is loaded along with a user’s program, but none of the routines are ever called, then there is no extra runtime overhead. (The alternative is move the body of _pw_init() into a BEGIN rule, which always runs pwcat. This simplifies the code but runs an extra process that may never be needed.)

In turn, calling _pw_init() is not too expensive, because the _pw_inited variable keeps the program from reading the data more than once. If you are worried about squeezing every last cycle out of your awk program, the check of _pw_inited could be moved out of _pw_init() and duplicated in all the other functions. In practice, this is not necessary, as most awk programs are I/O-bound, and such a change would clutter up the code.

The id program in Section 11.2.3 [Printing Out User Information], page 260, uses these functions.

10.6 Reading the Group Database

Much of the discussion presented in Section 10.5 [Reading the User Database], page 238, applies to the group database as well. Although there has traditionally been a well-known file (/etc/group) in a well-known format, the POSIX standard only provides a set of C library routines (<grp.h> and getgrent()) for accessing the information. Even though this file may exist, it may not have complete information. Therefore, as with the user database, it is necessary to have a small C program that generates the group database as its output. grcat, a C program that “cats” the group database, is as follows:

     /*
      * grcat.c
      *
      * Generate a printable version of the group database.
      */
     #include <stdio.h>
     #include <grp.h>
     int
     main(int argc, char **argv)

{
struct group *g;
int i;
while ((g = getgrent()) != NULL) {
   printf("%s:%s:%ld:", g->gr_name, g->gr_passwd,
                                (long) g->gr_gid);
   for (i = 0; g->gr_mem[i] != NULL; i++) {
       printf("%s", g->gr_mem[i]);
       if (g->gr_mem[i+1] != NULL)
           putchar(',');
   }
   putchar('\n');
}
endgrent();
return 0;
}

Each line in the group database represents one group. The fields are separated with colons and represent the following information:

Group Name The group’s name.

Group Password The group’s encrypted password. In practice, this field is never used; it is usually empty or set to ‘*’.

Group ID Number The group’s numeric group ID number; the association of name to number must be unique within the file. (On some systems it’s a C long, and not an int. Thus, we cast it to long for all cases.)

Group Member List A comma-separated list of usernames. These users are members of the group. Modern Unix systems allow users to be members of several groups simultaneously. If your system does, then there are elements "group1" through "groupN" in PROCINFO for those group ID numbers. (Note that PROCINFO is a gawk extension; see Section 7.5 [Predefined Variables], page 147.)

Here is what running grcat might produce:

$ grcat
⊣ wheel:*:0:arnold
⊣ nogroup:*:65534:
⊣ daemon:*:1:
⊣ kmem:*:2:
⊣ staff:*:10:arnold,miriam,andy
⊣ other:*:20:
...

Here are the functions for obtaining information from the group database. There are several, modeled after the C library functions of the same names:

# group.awk --- functions for dealing with the group file
BEGIN {
# Change to suit your system
_gr_awklib = "/usr/local/libexec/awk/"
}
function _gr_init(    oldfs, oldrs, olddol0, grcat,
                         using_fw, using_fpat, n, a, i)
{
if (_gr_inited)
    return
oldfs = FS
oldrs = RS
olddol0 = $0
using_fw = (PROCINFO["FS"] == "FIELDWIDTHS")
using_fpat = (PROCINFO["FS"] == "FPAT")
FS = ":"
RS = "\n"
grcat = _gr_awklib "grcat"
while ((grcat | getline) > 0) {
    if ($1 in _gr_byname)
        _gr_byname[$1] = _gr_byname[$1] "," $4
    else
        _gr_byname[$1] = $0
    if ($3 in _gr_bygid)
        _gr_bygid[$3] = _gr_bygid[$3] "," $4
    else
        _gr_bygid[$3] = $0
    n = split($4, a, "[ \t]*,[ \t]*")
    for (i = 1; i <= n; i++)
        if (a[i] in _gr_groupsbyuser)
            _gr_groupsbyuser[a[i]] = gr_groupsbyuser[a[i]] " " $1
        else
            _gr_groupsbyuser[a[i]] = $1
    _gr_bycount[++_gr_count] = $0
}
close(grcat)
_gr_count = 0
_gr_inited++
FS = oldfs
if (using_fw)
    FIELDWIDTHS = FIELDWIDTHS

         else if (using_fpat)
             FPAT = FPAT
         RS = oldrs
         $0 = olddol0
     }

The BEGIN rule sets a private variable to the directory where grcat is stored. Because it is used to help out an awk library routine, we have chosen to put it in /usr/local/libexec/awk. You might want it to be in a different directory on your system.

These routines follow the same general outline as the user database routines (see Section 10.5 [Reading the User Database], page 238). The _gr_inited variable is used to ensure that the database is scanned no more than once. The _gr_init() function first saves FS, RS, and $0, and then sets FS and RS to the correct values for scanning the group information. It also takes care to note whether FIELDWIDTHS or FPAT is being used, and to restore the appropriate field-splitting mechanism.

The group information is stored in several associative arrays. The arrays are indexed by group name (_gr_byname), by group ID number (_gr_bygid), and by position in the database (_gr_bycount). There is an additional array indexed by username (_gr_groupsbyuser), which is a space-separated list of groups to which each user belongs.

Unlike in the user database, it is possible to have multiple records in the database for the same group. This is common when a group has a large number of members. A pair of such entries might look like the following:

     tvpeople:*:101:johnny,jay,arsenio
     tvpeople:*:101:david,conan,tom,joan

For this reason, _gr_init() looks to see if a group name or group ID number is already seen. If so, the usernames are simply concatenated onto the previous list of users.^{11}

Finally, _gr_init() closes the pipeline to grcat, restores FS (and FIELDWIDTHS or FPAT, if necessary), RS, and $0, initializes _gr_count to zero (it is used later), and makes _gr_ inited nonzero.

The getgrnam() function takes a group name as its argument, and if that group exists, it is returned. Otherwise, it relies on the array reference to a nonexistent element to create the element with the null string as its value:

     function getgrnam(group)
     {
         _gr_init()
         return _gr_byname[group]
     }

The getgrgid() function is similar; it takes a numeric group ID and looks up the information associated with that group ID:

     function getgrgid(gid)
     {
         _gr_init()

11 There is a subtle problem with the code just presented. Suppose that the first time there were no names.

This code adds the names with a leading comma. It also doesn’t check that there is a $4.

         return _gr_bygid[gid]
     }

The getgruser() function does not have a C counterpart. It takes a username and returns the list of groups that have the user as a member:

     function getgruser(user)
     {
         _gr_init()
         return _gr_groupsbyuser[user]
     }

The getgrent() function steps through the database one entry at a time. It uses _gr_ count to track its position in the list:

     function getgrent()
     {
         _gr_init()
         if (++_gr_count in _gr_bycount)
             return _gr_bycount[_gr_count]
         return ""
     }

The endgrent() function resets _gr_count to zero so that getgrent() can start over again:

     function endgrent()
     {
         _gr_count = 0
     }

As with the user database routines, each function calls _gr_init() to initialize the arrays. Doing so only incurs the extra overhead of running grcat if these functions are used (as opposed to moving the body of _gr_init() into a BEGIN rule).

Most of the work is in scanning the database and building the various associative arrays. The functions that the user calls are themselves very simple, relying on awk’s associative arrays to do work.

The id program in Section 11.2.3 [Printing Out User Information], page 260, uses these functions.

10.7 Traversing Arrays of Arrays

Section 8.6 [Arrays of Arrays], page 171, described how gawk provides arrays of arrays. In particular, any element of an array may be either a scalar or another array. The isarray() function (see Section 9.1.7 [Getting Type Information], page 198) lets you distinguish an array from a scalar. The following function, walk_array(), recursively traverses an array, printing the element indices and values. You call it with the array and a string representing the name of the array:

     function walk_array(arr, name,      i)
     {
         for (i in arr) {
             if (isarray(arr[i]))

        walk_array(arr[i], (name "[" i "]"))
    else
        printf("%s[%s] = %s\n", name, i, arr[i])
}
}

It works by looping over each element of the array. If any given element is itself an array, the function calls itself recursively, passing the subarray and a new string representing the current index. Otherwise, the function simply prints the element’s name, index, and value. Here is a main program to demonstrate:

BEGIN {
a[1] = 1
a[2][1] = 21
a[2][2] = 22
a[3] = 3
a[4][1][1] = 411
a[4][2] = 42
walk_array(a, "a")
}

When run, the program produces the following output:

$ gawk -f walk_array.awk
⊣ a[1] = 1
⊣ a[2][1] = 21
⊣ a[2][2] = 22
⊣ a[3] = 3
⊣ a[4][1][1] = 411
⊣ a[4][2] = 42

The function just presented simply prints the name and value of each scalar array element. However, it is easy to generalize it, by passing in the name of a function to call when walking an array. The modified function looks like this:

function process_array(arr, name, process, do_arrays,   i, new_name)
{
for (i in arr) {
    new_name = (name "[" i "]")
    if (isarray(arr[i])) {
        if (do_arrays)
            @process(new_name, arr[i])
        process_array(arr[i], new_name, process, do_arrays)
    } else
        @process(new_name, arr[i])
}
}

The arguments are as follows:

arr The array.

name The name of the array (a string). process The name of the function to call.

do_arrays

If this is true, the function can handle elements that are subarrays.

If subarrays are to be processed, that is done before walking them further.

When run with the following scaffolding, the function produces the same results as does the earlier version of walk_array():

BEGIN {
a[1] = 1
a[2][1] = 21
a[2][2] = 22
a[3] = 3
a[4][1][1] = 411
a[4][2] = 42
process_array(a, "a", "do_print", 0)
}
function do_print(name, element)
{
printf "%s = %s\n", name, element
}

10.8 Summary

• Reading programs is an excellent way to learn Good Programming. The functions and programs provided in this chapter and the next are intended to serve that purpose.

• When writing general-purpose library functions, put some thought into how to name any global variables so that they won’t conflict with variables from a user’s program.

• The functions presented here fit into the following categories:

General problems Number-to-string conversion, testing assertions, rounding, random number generation, converting characters to numbers, joining strings, getting easily usable time-of-day information, and reading a whole file in one shot

Managing data files Noting data file boundaries, rereading the current file, checking for readable files, checking for zero-length files, and treating assignments as file names

Processing command-line options An awk version of the standard C getopt() function

Reading the user and group databases Two sets of routines that parallel the C library versions

Traversing arrays of arrays Two functions that traverse an array of arrays to any depth

10.9 Exercises

1. In Section 10.3.4 [Checking for Zero-Length Files], page 232, we presented the zerofile.awk program, which made use of gawk’s ARGIND variable. Can this problem be solved without relying on ARGIND? If so, how?

2. As a related challenge, revise that code to handle the case where an intervening value in ARGV is a variable assignment.

3. Section 10.7 [Traversing Arrays of Arrays], page 246, presented a function that walked a multidimensional array to print it out. However, walking an array and processing each element is a general-purpose operation. Generalize the walk_array() function by adding an additional parameter named process. Then, inside the loop, instead of printing the array element’s index and value, use the indirect function call syntax (see Section 9.3 [Indirect Function Calls], page 208) on process, passing it the index and the value. When calling walk_array(), you would pass the name of a user-defined function that expects to receive an index and a value, and then processes the element. Test your new version by printing the array; you should end up with output identical to that of the original version.

Chapter 11: Practical awk Programs 251

11 Practical awk Programs

Chapter 10 [A Library of awk Functions], page 217, presents the idea that reading programs in a language contributes to learning that language. This chapter continues that theme, presenting a potpourri of awk programs for your reading enjoyment. There are three sections. The first describes how to run the programs presented in this chapter.

The second presents awk versions of several common POSIX utilities. These are programs that you are hopefully already familiar with, and therefore whose problems are understood. By reimplementing these programs in awk, you can focus on the awk-related aspects of solving the programming problems.

The third is a grab bag of interesting programs. These solve a number of different data-manipulation and management problems. Many of the programs are short, which emphasizes awk’s ability to do a lot in just a few lines of code.

Many of these programs use library functions presented in Chapter 10 [A Library of awk Functions], page 217.

11.1 Running the Example Programs

To run a given program, you would typically do something like this:

     awk -f program -- options files

Here, program is the name of the awk program (such as cut.awk), options are any commandline options for the program that start with a ‘-’, and files are the actual data files.

If your system supports the ‘#!’ executable interpreter mechanism (see Section 1.1.4 [Executable awk Programs], page 19), you can instead run your program directly:

     cut.awk -c1-8 myfiles > results

If your awk is not gawk, you may instead need to use this:

     cut.awk -- -c1-8 myfiles > results

11.2 Reinventing Wheels for Fun and Profit

This section presents a number of POSIX utilities implemented in awk. Reinventing these programs in awk is often enjoyable, because the algorithms can be very clearly expressed, and the code is usually very concise and simple. This is true because awk does so much for you.

It should be noted that these programs are not necessarily intended to replace the installed versions on your system. Nor may all of these programs be fully compliant with the most recent POSIX standard. This is not a problem; their purpose is to illustrate awk language programming for “real-world” tasks.

The programs are presented in alphabetical order.

11.2.1 Cutting Out Fields and Columns

The cut utility selects, or “cuts,” characters or fields from its standard input and sends them to its standard output. Fields are separated by TABs by default, but you may supply a command-line option to change the field delimiter (i.e., the field-separator character). cut’s definition of fields is less general than awk’s.

A common use of cut might be to pull out just the login names of logged-on users from the output of who. For example, the following pipeline generates a sorted, unique list of the logged-on users:

     who | cut -c1-8 | sort | uniq

The options for cut are:

-c list Use list as the list of characters to cut out. Items within the list may be separated by commas, and ranges of characters can be separated with dashes. The list ‘1-8,15,22-35’ specifies characters 1 through 8, 15, and 22 through 35.

-f list Use list as the list of fields to cut out.

-d delim Use delim as the field-separator character instead of the TAB character.

-s Suppress printing of lines that do not contain the field delimiter.

The awk implementation of cut uses the getopt() library function (see Section 10.4 [Processing Command-Line Options], page 233) and the join() library function (see Section 10.2.6 [Merging an Array into a String], page 225).

The program begins with a comment describing the options, the library functions needed, and a usage() function that prints out a usage message and exits. usage() is called if invalid arguments are supplied:

     # cut.awk --- implement cut in awk
     # Options:
     #    -f list     Cut fields
     #    -d c        Field delimiter character
     #    -c list     Cut characters
     #
     #    -s          Suppress lines without the delimiter
     #
     # Requires getopt() and join() library functions
     function usage()
     {
         print("usage: cut [-f list] [-d c] [-s] [files...]") > "/dev/stderr"
         print("usage: cut [-c list] [files...]") > "/dev/stderr"
         exit 1
     }

Next comes a BEGIN rule that parses the command-line options. It sets FS to a single TAB character, because that is cut’s default field separator. The rule then sets the output field separator to be the same as the input field separator. A loop using getopt() steps through the command-line options. Exactly one of the variables by_fields or by_chars is set to true, to indicate that processing should be done by fields or by characters, respectively. When cutting by characters, the output field separator is set to the null string:

     BEGIN {
         FS = "\t"    # default

OFS = FS
while ((c = getopt(ARGC, ARGV, "sf:c:d:")) != -1) {
if (c == "f") {
by_fields = 1
fieldlist = Optarg
} else if (c == "c") {
by_chars = 1
fieldlist = Optarg
OFS = ""
} else if (c == "d") {
if (length(Optarg) > 1) {
    printf("cut: using first character of %s" \
           " for delimiter\n", Optarg) > "/dev/stderr"
    Optarg = substr(Optarg, 1, 1)
}
FS = Optarg
OFS = FS
if (FS == " ")    # defeat awk semantics
    FS = "[ ]"
} else if (c == "s")
suppress = 1
else
usage()
}
# Clear out options
for (i = 1; i < Optind; i++)
ARGV[i] = ""

The code must take special care when the field delimiter is a space. Using a single space (" ") for the value of FS is incorrect—awk would separate fields with runs of spaces, TABs, and/or newlines, and we want them to be separated with individual spaces. Also remember that after getopt() is through (as described in Section 10.4 [Processing Command-Line Options], page 233), we have to clear out all the elements of ARGV from 1 to Optind, so that awk does not try to process the command-line options as file names.

After dealing with the command-line options, the program verifies that the options make sense. Only one or the other of -c and -f should be used, and both require a field list. Then the program calls either set_fieldlist() or set_charlist() to pull apart the list of fields or characters:

if (by_fields && by_chars)
usage()
if (by_fields == 0 && by_chars == 0)
by_fields = 1    # default
if (fieldlist == "") {
print "cut: needs list for -c or -f" > "/dev/stderr"

             exit 1
         }
         if (by_fields)
             set_fieldlist()
         else
             set_charlist()
     }

set_fieldlist() splits the field list apart at the commas into an array. Then, for each element of the array, it looks to see if the element is actually a range, and if so, splits it apart. The function checks the range to make sure that the first number is smaller than the second. Each number in the list is added to the flist array, which simply lists the fields that will be printed. Normal field splitting is used. The program lets awk handle the job of doing the field splitting:

     function set_fieldlist(        n, m, i, j, k, f, g)
     {
         n = split(fieldlist, f, ",")
         j = 1    # index in flist
         for (i = 1; i <= n; i++) {
             if (index(f[i], "-") != 0) { # a range
                 m = split(f[i], g, "-")
                 if (m != 2 || g[1] >= g[2]) {
                     printf("cut: bad field list: %s\n",
                                       f[i]) > "/dev/stderr"
                     exit 1
                 }
                 for (k = g[1]; k <= g[2]; k++)
                     flist[j++] = k
             } else
                 flist[j++] = f[i]
         }
         nfields = j - 1
     }

The set_charlist() function is more complicated than set_fieldlist(). The idea here is to use gawk’s FIELDWIDTHS variable (see Section 4.6 [Reading Fixed-Width Data], page 72), which describes constant-width input. When using a character list, that is exactly what we have.

Setting up FIELDWIDTHS is more complicated than simply listing the fields that need to be printed. We have to keep track of the fields to print and also the intervening characters that have to be skipped. For example, suppose you wanted characters 1 through 8, 15, and 22 through 35. You would use ‘-c 1-8,15,22-35’. The necessary value for FIELDWIDTHS is "8 6 1 6 14". This yields five fields, and the fields to print are $1, $3, and $5. The intermediate fields are filler, which is stuff in between the desired data. flist lists the fields to print, and t tracks the complete field list, including filler fields:

     function set_charlist(    field, i, j, f, g, n, m, t,
                               filler, last, len)

{
field = 1   # count total fields
n = split(fieldlist, f, ",")
j = 1       # index in flist
for (i = 1; i <= n; i++) {
if (index(f[i], "-") != 0) { # range
m = split(f[i], g, "-")
if (m != 2 || g[1] >= g[2]) {
    printf("cut: bad character list: %s\n",
                   f[i]) > "/dev/stderr"
    exit 1
}
len = g[2] - g[1] + 1
if (g[1] > 1)  # compute length of filler
    filler = g[1] - last - 1
else
    filler = 0
if (filler)
    t[field++] = filler
t[field++] = len  # length of field
last = g[2]
flist[j++] = field - 1
} else {
if (f[i] > 1)
    filler = f[i] - last - 1
else
    filler = 0
if (filler)
    t[field++] = filler
t[field++] = 1
last = f[i]
flist[j++] = field - 1
}
}
FIELDWIDTHS = join(t, 1, field - 1)
nfields = j - 1
}

Next is the rule that processes the data. If the -s option is given, then suppress is true. The first if statement makes sure that the input record does have the field separator. If cut is processing fields, suppress is true, and the field separator character is not in the record, then the record is skipped.

If the record is valid, then gawk has split the data into fields, either using the character in FS or using fixed-length fields and FIELDWIDTHS. The loop goes through the list of fields that should be printed. The corresponding field is printed if it contains data. If the next field also has data, then the separator character is written out between the fields:

{

         if (by_fields && suppress && index($0, FS) == 0)
             next
         for (i = 1; i <= nfields; i++) {
             if ($flist[i] != "") {
                 printf "%s", $flist[i]
                 if (i < nfields && $flist[i+1] != "")
                     printf "%s", OFS
             }
         }
         print ""
     }

This version of cut relies on gawk’s FIELDWIDTHS variable to do the character-based cutting. It is possible in other awk implementations to use substr() (see Section 9.1.3 [String-Manipulation Functions], page 177), but it is also extremely painful. The FIELDWIDTHS variable supplies an elegant solution to the problem of picking the input line apart by characters.

11.2.2 Searching for Regular Expressions in Files

The egrep utility searches files for patterns. It uses regular expressions that are almost identical to those available in awk (see Chapter 3 [Regular Expressions], page 45). You invoke it as follows:

egrep [options] 'pattern' files … The pattern is a regular expression. In typical usage, the regular expression is quoted

to prevent the shell from expanding any of the special characters as file name wildcards. Normally, egrep prints the lines that matched. If multiple file names are provided on the command line, each output line is preceded by the name of the file and a colon.

The options to egrep are as follows:

-c Print out a count of the lines that matched the pattern, instead of the lines themselves.

-s Be silent. No output is produced and the exit value indicates whether the pattern was matched.

-v Invert the sense of the test. egrep prints the lines that do not match the pattern and exits successfully if the pattern is not matched.

-i Ignore case distinctions in both the pattern and the input data.

-l Only print (list) the names of the files that matched, not the lines that matched.

-e pattern

Use pattern as the regexp to match. The purpose of the -e option is to allow patterns that start with a ‘-’.

This version uses the getopt() library function (see Section 10.4 [Processing Command-Line Options], page 233) and the file transition library program (see Section 10.3.1 [Noting Data file Boundaries], page 229).

The program begins with a descriptive comment and then a BEGIN rule that processes the command-line arguments with getopt(). The -i (ignore case) option is particularly easy with gawk; we just use the IGNORECASE predefined variable (see Section 7.5 [Predefined Variables], page 147):

# egrep.awk --- simulate egrep in awk
#
# Options:
#    -c    count of lines
#    -s    silent - use exit value
#    -v    invert test, success if no match
#    -i    ignore case
#    -l    print filenames only
#    -e    argument is pattern
#
# Requires getopt and file transition library functions
BEGIN {
    while ((c = getopt(ARGC, ARGV, "ce:svil")) != -1) {
        if (c == "c")
            count_only++
        else if (c == "s")
            no_print++
        else if (c == "v")
            invert++
        else if (c == "i")
            IGNORECASE = 1
        else if (c == "l")
            filenames_only++
        else if (c == "e")
            pattern = Optarg
        else
            usage()
    }

Next comes the code that handles the egrep-specific behavior. If no pattern is supplied with -e, the first nonoption on the command line is used. The awk command-line arguments up to ARGV[Optind] are cleared, so that awk won’t try to process them as files. If no files are specified, the standard input is used, and if multiple files are specified, we make sure to note this so that the file names can precede the matched lines in the output:

    if (pattern == "")
        pattern = ARGV[Optind++]
    for (i = 1; i < Optind; i++)
        ARGV[i] = ""
    if (Optind >= ARGC) {
        ARGV[1] = "-"
        ARGC = 2

         } else if (ARGC - Optind > 1)
             do_filenames++
     #    if (IGNORECASE)
     #        pattern = tolower(pattern)
     }

The last two lines are commented out, as they are not needed in gawk. They should be uncommented if you have to use another version of awk.

The next set of lines should be uncommented if you are not using gawk. This rule translates all the characters in the input line into lowercase if the -i option is specified.^{1} The rule is commented out as it is not necessary with gawk:

     #{
     #    if (IGNORECASE)
     #        $0 = tolower($0)
     #}

The beginfile() function is called by the rule in ftrans.awk when each new file is processed. In this case, it is very simple; all it does is initialize a variable fcount to zero. fcount tracks how many lines in the current file matched the pattern. Naming the parameter junk shows we know that beginfile() is called with a parameter, but that we’re not interested in its value:

     function beginfile(junk)
     {
         fcount = 0
     }

The endfile() function is called after each file has been processed. It affects the output only when the user wants a count of the number of lines that matched. no_print is true only if the exit status is desired. count_only is true if line counts are desired. egrep therefore only prints line counts if printing and counting are enabled. The output format must be adjusted depending upon the number of files to process. Finally, fcount is added to total, so that we know the total number of lines that matched the pattern:

     function endfile(file)
     {
         if (! no_print && count_only) {
             if (do_filenames)
                 print file ":" fcount
             else
                 print fcount
         }
         total += fcount
     }

The BEGINFILE and ENDFILE special patterns (see Section 7.1.5 [The BEGINFILE and ENDFILE Special Patterns], page 135) could be used, but then the program would be gawk-specific. Additionally, this example was written before gawk acquired BEGINFILE and

1 It also introduces a subtle bug; if a match happens, we output the translated line, not the original.

ENDFILE.

The following rule does most of the work of matching lines. The variable matches is true if the line matched the pattern. If the user wants lines that did not match, the sense of matches is inverted using the ‘!’ operator. fcount is incremented with the value of matches, which is either one or zero, depending upon a successful or unsuccessful match. If the line does not match, the next statement just moves on to the next record.

A number of additional tests are made, but they are only done if we are not counting lines. First, if the user only wants the exit status (no_print is true), then it is enough to know that one line in this file matched, and we can skip on to the next file with nextfile. Similarly, if we are only printing file names, we can print the file name, and then skip to the next file with nextfile. Finally, each line is printed, with a leading file name and colon if necessary:

     {
         matches = ($0 ~ pattern)
         if (invert)
             matches = ! matches
         fcount += matches    # 1 or 0
         if (! matches)
             next
         if (! count_only) {
             if (no_print)
                 nextfile
             if (filenames_only) {
                 print FILENAME
                 nextfile
             }
             if (do_filenames)
                 print FILENAME ":" $0
             else
                 print
         }
     }

The END rule takes care of producing the correct exit status. If there are no matches, the exit status is one; otherwise, it is zero:

     END {
         exit (total == 0)
     }

The usage() function prints a usage message in case of invalid options, and then exits:

     function usage()

     {
         print("Usage: egrep [-csvil] [-e pat] [files ...]") > "/dev/stderr"
         print("\n\tegrep [-csvil] pat [files ...]") > "/dev/stderr"
         exit 1
     }

11.2.3 Printing Out User Information

The id utility lists a user’s real and effective user ID numbers, real and effective group ID numbers, and the user’s group set, if any. id only prints the effective user ID and group ID if they are different from the real ones. If possible, id also supplies the corresponding user and group names. The output might look like this:

     $ id
     ⊣ uid=1000(arnold) gid=1000(arnold) groups=1000(arnold),4(adm),7(lp),27(sudo)

This information is part of what is provided by gawk’s PROCINFO array (see Section 7.5 [Predefined Variables], page 147). However, the id utility provides a more palatable output than just individual numbers.

Here is a simple version of id written in awk. It uses the user database library functions (see Section 10.5 [Reading the User Database], page 238) and the group database library functions (see Section 10.6 [Reading the Group Database], page 242) from Chapter 10 [A Library of awk Functions], page 217.

The program is fairly straightforward. All the work is done in the BEGIN rule. The user and group ID numbers are obtained from PROCINFO. The code is repetitive. The entry in the user database for the real user ID number is split into parts at the ‘:’. The name is the first field. Similar code is used for the effective user ID number and the group numbers:

     # id.awk --- implement id in awk
     #
     # Requires user and group library functions
     # output is:
     # uid=12(foo) euid=34(bar) gid=3(baz) \
     #             egid=5(blat) groups=9(nine),2(two),1(one)
     BEGIN {
         uid = PROCINFO["uid"]
         euid = PROCINFO["euid"]
         gid = PROCINFO["gid"]
         egid = PROCINFO["egid"]
         printf("uid=%d", uid)
         pw = getpwuid(uid)
         pr_first_field(pw)
         if (euid != uid) {
             printf(" euid=%d", euid)
             pw = getpwuid(euid)
             pr_first_field(pw)
         }

printf(" gid=%d", gid)
pw = getgrgid(gid)
pr_first_field(pw)
if (egid != gid) {
printf(" egid=%d", egid)
pw = getgrgid(egid)
pr_first_field(pw)
}
for (i = 1; ("group" i) in PROCINFO; i++) {
if (i == 1)
    printf(" groups=")
group = PROCINFO["group" i]
printf("%d", group)
pw = getgrgid(group)
pr_first_field(pw)
if (("group" (i+1)) in PROCINFO)
    printf(",")
}
print ""
}
function pr_first_field(str,  a)
{
if (str != "") {
split(str, a, ":")
printf("(%s)", a[1])
}
}

The test in the for loop is worth noting. Any supplementary groups in the PROCINFO array have the indices "group1" through "groupN" for some N (i.e., the total number of supplementary groups). However, we don’t know in advance how many of these groups there are.

This loop works by starting at one, concatenating the value with "group", and then using in to see if that value is in the array (see Section 8.1.2 [Referring to an Array Element], page 161). Eventually, i is incremented past the last group in the array and the loop exits.

The loop is also correct if there are no supplementary groups; then the condition is false the first time it’s tested, and the loop body never executes.

The pr_first_field() function simply isolates out some code that is used repeatedly, making the whole program shorter and cleaner. In particular, moving the check for the empty string into this function saves several lines of code.

11.2.4 Splitting a Large File into Pieces

The split program splits large text files into smaller pieces. Usage is as follows:^{2}

split [-count] [file] [prefix] By default, the output files are named xaa, xab, and so on. Each file has 1,000 lines

in it, with the likely exception of the last file. To change the number of lines in each file, supply a number on the command line preceded with a minus sign (e.g., ‘-500’ for files with 500 lines in them instead of 1,000). To change the names of the output files to something like myfileaa, myfileab, and so on, supply an additional argument that specifies the file name prefix.

Here is a version of split in awk. It uses the ord() and chr() functions presented in Section 10.2.5 [Translating Between Characters and Numbers], page 223.

The program first sets its defaults, and then tests to make sure there are not too many arguments. It then looks at each argument in turn. The first argument could be a minus sign followed by a number. If it is, this happens to look like a negative number, so it is made positive, and that is the count of lines. The data file name is skipped over and the final argument is used as the prefix for the output file names:

     # split.awk --- do split in awk
     #
     # Requires ord() and chr() library functions
     # usage: split [-count] [file] [outname]
     BEGIN {
         outfile = "x"    # default
         count = 1000
         if (ARGC > 4)
             usage()
         i = 1
         if (i in ARGV && ARGV[i] ~ /^-[[:digit:]]+$/) {
             count = -ARGV[i]
             ARGV[i] = ""
             i++
         }
         # test argv in case reading from stdin instead of file
         if (i in ARGV)
             i++    # skip datafile name
         if (i in ARGV) {
             outfile = ARGV[i]
             ARGV[i] = ""
         }
         s1 = s2 = "a"
         out = (outfile s1 s2)

2 This is the traditional usage. The POSIX usage is different, but not relevant for what the program aims

to demonstrate.

     }

The next rule does most of the work. tcount (temporary count) tracks how many lines have been printed to the output file so far. If it is greater than count, it is time to close the current file and start a new one. s1 and s2 track the current suffixes for the file name. If they are both ‘z’, the file is just too big. Otherwise, s1 moves to the next letter in the alphabet and s2 starts over again at ‘a’:

     {
         if (++tcount > count) {
             close(out)
             if (s2 == "z") {
                 if (s1 == "z") {
                     printf("split: %s is too large to split\n",
                            FILENAME) > "/dev/stderr"
                     exit 1
                 }
                 s1 = chr(ord(s1) + 1)
                 s2 = "a"
             }
             else
                 s2 = chr(ord(s2) + 1)
             out = (outfile s1 s2)
             tcount = 1
         }
         print > out
     }

The usage() function simply prints an error message and exits:

     function usage()
     {
         print("usage: split [-num] [file] [outname]") > "/dev/stderr"
         exit 1
     }

This program is a bit sloppy; it relies on awk to automatically close the last file instead of doing it in an END rule. It also assumes that letters are contiguous in the character set, which isn’t true for EBCDIC systems.

11.2.5 Duplicating Output into Multiple Files

The tee program is known as a “pipe fitting.” tee copies its standard input to its standard output and also duplicates it to the files named on the command line. Its usage is as follows:

tee [-a] file …

The -a option tells tee to append to the named files, instead of truncating them and starting over.

The BEGIN rule first makes a copy of all the command-line arguments into an array named copy. ARGV[0] is not needed, so it is not copied. tee cannot use ARGV directly, because awk attempts to process each file name in ARGV as input data.

If the first argument is -a, then the flag variable append is set to true, and both ARGV[1] and copy[1] are deleted. If ARGC is less than two, then no file names were supplied and tee prints a usage message and exits. Finally, awk is forced to read the standard input by setting ARGV[1] to "-" and ARGC to two:

# tee.awk --- tee in awk
#
# Copy standard input to all named output files.
# Append content if -a option is supplied.
#
BEGIN {
for (i = 1; i < ARGC; i++)
    copy[i] = ARGV[i]
if (ARGV[1] == "-a") {
    append = 1
    delete ARGV[1]
    delete copy[1]
    ARGC--
}
if (ARGC < 2) {
    print "usage: tee [-a] file ..." > "/dev/stderr"
    exit 1
}
ARGV[1] = "-"
ARGC = 2
}

The following single rule does all the work. Because there is no pattern, it is executed for each line of input. The body of the rule simply prints the line into each file on the command line, and then to the standard output:

{
# moving the if outside the loop makes it run faster
if (append)
    for (i in copy)
        print >> copy[i]
else
    for (i in copy)
        print > copy[i]
print
}

It is also possible to write the loop this way:

for (i in copy)
if (append)
    print >> copy[i]
else
    print > copy[i]

This is more concise, but it is also less efficient. The ‘if’ is tested for each record and for each output file. By duplicating the loop body, the ‘if’ is only tested once for each input record. If there are N input records and M output files, the first method only executes N ‘if’ statements, while the second executes N*M ‘if’ statements.

Finally, the END rule cleans up by closing all the output files:

     END {
         for (i in copy)
             close(copy[i])
     }

11.2.6 Printing Nonduplicated Lines of Text

The uniq utility reads sorted lines of data on its standard input, and by default removes duplicate lines. In other words, it only prints unique lines—hence the name. uniq has a number of options. The usage is as follows:

uniq [-udc [-n]] [+n] [inputfile [outputfile]] The options for uniq are:

-d Print only repeated (duplicated) lines.

-u Print only nonrepeated (unique) lines.

-c Count lines. This option overrides -d and -u. Both repeated and nonrepeated lines are counted.

-n Skip n fields before comparing lines. The definition of fields is similar to awk’s default: nonwhitespace characters separated by runs of spaces and/or TABs.

+n Skip n characters before comparing lines. Any fields specified with ‘-n’ are skipped first.

inputfile

Data is read from the input file named on the command line, instead of from the standard input.

outputfile

The generated output is sent to the named output file, instead of to the standard output.

Normally uniq behaves as if both the -d and -u options are provided.

uniq uses the getopt() library function (see Section 10.4 [Processing Command-Line Options], page 233) and the join() library function (see Section 10.2.6 [Merging an Array into a String], page 225).

The program begins with a usage() function and then a brief outline of the options and their meanings in comments. The BEGIN rule deals with the command-line arguments and options. It uses a trick to get getopt() to handle options of the form ‘-25’, treating such an option as the option letter ‘2’ with an argument of ‘5’. If indeed two or more digits are supplied (Optarg looks like a number), Optarg is concatenated with the option digit and then the result is added to zero to make it into a number. If there is only one digit in the option, then Optarg is not needed. In this case, Optind must be decremented so that getopt() processes it next time. This code is admittedly a bit tricky.

If no options are supplied, then the default is taken, to print both repeated and nonrepeated lines. The output file, if provided, is assigned to outputfile. Early on, outputfile is initialized to the standard output, /dev/stdout:

# uniq.awk --- do uniq in awk
#
# Requires getopt() and join() library functions
function usage()
{
    print("Usage: uniq [-udc [-n]] [+n] [ in [ out ]]") > "/dev/stderr"
    exit 1
}
# -c    count lines. overrides -d and -u
# -d    only repeated lines
# -u    only nonrepeated lines
# -n    skip n fields
# +n    skip n characters, skip fields first
BEGIN {
    count = 1
    outputfile = "/dev/stdout"
    opts = "udc0:1:2:3:4:5:6:7:8:9:"
    while ((c = getopt(ARGC, ARGV, opts)) != -1) {
        if (c == "u")
            non_repeated_only++
        else if (c == "d")
            repeated_only++
        else if (c == "c")
            do_count++
        else if (index("0123456789", c) != 0) {
            # getopt() requires args to options
            # this messes us up for things like -5
            if (Optarg ~ /^[[:digit:]]+$/)
                fcount = (c Optarg) + 0
            else {
                fcount = c + 0
                Optind--
            }
        } else
            usage()
    }
    if (ARGV[Optind] ~ /^\+[[:digit:]]+$/) {
        charcount = substr(ARGV[Optind], 2) + 0
        Optind++

}
for (i = 1; i < Optind; i++)
ARGV[i] = ""
if (repeated_only == 0 && non_repeated_only == 0)
repeated_only = non_repeated_only = 1
if (ARGC - Optind == 2) {
outputfile = ARGV[ARGC - 1]
ARGV[ARGC - 1] = ""
}
}

The following function, are_equal(), compares the current line, $0, to the previous line, last. It handles skipping fields and characters. If no field count and no character count are specified, are_equal() returns one or zero depending upon the result of a simple string comparison of last and $0.

Otherwise, things get more complicated. If fields have to be skipped, each line is broken into an array using split() (see Section 9.1.3 [String-Manipulation Functions], page 177); the desired fields are then joined back into a line using join(). The joined lines are stored in clast and cline. If no fields are skipped, clast and cline are set to last and $0, respectively. Finally, if characters are skipped, substr() is used to strip off the leading charcount characters in clast and cline. The two strings are then compared and are_ equal() returns the result:

function are_equal(    n, m, clast, cline, alast, aline)
{
if (fcount == 0 && charcount == 0)
return (last == $0)
if (fcount > 0) {
n = split(last, alast)
m = split($0, aline)
clast = join(alast, fcount+1, n)
cline = join(aline, fcount+1, m)
} else {
clast = last
cline = $0
}
if (charcount) {
clast = substr(clast, charcount + 1)
cline = substr(cline, charcount + 1)
}
return (clast == cline)
}

The following two rules are the body of the program. The first one is executed only for the very first line of data. It sets last equal to $0, so that subsequent lines of text have something to be compared to.

The second rule does the work. The variable equal is one or zero, depending upon the results of are_equal()’s comparison. If uniq is counting repeated lines, and the lines are equal, then it increments the count variable. Otherwise, it prints the line and resets count, because the two lines are not equal.

If uniq is not counting, and if the lines are equal, count is incremented. Nothing is printed, as the point is to remove duplicates. Otherwise, if uniq is counting repeated lines and more than one line is seen, or if uniq is counting nonrepeated lines and only one line is seen, then the line is printed, and count is reset.

Finally, similar logic is used in the END rule to print the final line of input data:

NR == 1 {
last = $0
next
}
{
equal = are_equal()
if (do_count) {    # overrides -d and -u
    if (equal)
        count++
    else {
        printf("%4d %s\n", count, last) > outputfile
        last = $0
        count = 1    # reset
    }
    next
}
if (equal)
    count++
else {
    if ((repeated_only && count > 1) ||
        (non_repeated_only && count == 1))
            print last > outputfile
    last = $0
    count = 1
}
}
END {
if (do_count)
    printf("%4d %s\n", count, last) > outputfile
else if ((repeated_only && count > 1) ||

                 (non_repeated_only && count == 1))
             print last > outputfile
         close(outputfile)
     }

11.2.7 Counting Things

The wc (word count) utility counts lines, words, and characters in one or more input files. Its usage is as follows:

wc [-lwc] [files …] If no files are specified on the command line, wc reads its standard input. If there are

multiple files, it also prints total counts for all the files. The options and their meanings are as follows:

-l Count only lines.

-w Count only words. A “word” is a contiguous sequence of nonwhitespace characters, separated by spaces and/or TABs. Luckily, this is the normal way awk separates fields in its input data.

-c Count only characters.

Implementing wc in awk is particularly elegant, because awk does a lot of the work for us; it splits lines into words (i.e., fields) and counts them, it counts lines (i.e., records), and it can easily tell us how long a line is.

This program uses the getopt() library function (see Section 10.4 [Processing Command-Line Options], page 233) and the file-transition functions (see Section 10.3.1 [Noting Data file Boundaries], page 229).

This version has one notable difference from traditional versions of wc: it always prints the counts in the order lines, words, and characters. Traditional versions note the order of the -l, -w, and -c options on the command line, and print the counts in that order.

The BEGIN rule does the argument processing. The variable print_total is true if more than one file is named on the command line:

     # wc.awk --- count lines, words, characters
     # Options:
     #    -l    only count lines
     #    -w    only count words
     #    -c    only count characters
     #
     # Default is to count lines, words, characters
     #
     # Requires getopt() and file transition library functions
     BEGIN {
         # let getopt() print a message about
         # invalid options. we ignore them
         while ((c = getopt(ARGC, ARGV, "lwc")) != -1) {
             if (c == "l")

        do_lines = 1
    else if (c == "w")
        do_words = 1
    else if (c == "c")
        do_chars = 1
}
for (i = 1; i < Optind; i++)
    ARGV[i] = ""
# if no options, do all
if (! do_lines && ! do_words && ! do_chars)
    do_lines = do_words = do_chars = 1
print_total = (ARGC - i > 2)
}

The beginfile() function is simple; it just resets the counts of lines, words, and characters to zero, and saves the current file name in fname:

function beginfile(file)
{
lines = words = chars = 0
fname = FILENAME
}

The endfile() function adds the current file’s numbers to the running totals of lines, words, and characters. It then prints out those numbers for the file that was just read. It relies on beginfile() to reset the numbers for the following data file:

function endfile(file)
{
tlines += lines
twords += words
tchars += chars
if (do_lines)
    printf "\t%d", lines
if (do_words)
    printf "\t%d", words
if (do_chars)
    printf "\t%d", chars
printf "\t%s\n", fname
}

There is one rule that is executed for each line. It adds the length of the record, plus one, to chars.^{3} Adding one plus the record length is needed because the newline character separating records (the value of RS) is not part of the record itself, and thus not included in its length. Next, lines is incremented for each line read, and words is incremented by the value of NF, which is the number of “words” on this line:

# do per line

3 Because gawk understands multibyte locales, this code counts characters, not bytes.

     {
         chars += length($0) + 1    # get newline
         lines++
         words += NF
     }

Finally, the END rule simply prints the totals for all the files:

     END {
         if (print_total) {
             if (do_lines)
                 printf "\t%d", tlines
             if (do_words)
                 printf "\t%d", twords
             if (do_chars)
                 printf "\t%d", tchars
             print "\ttotal"
         }
     }

11.3 A Grab Bag of awk Programs

This section is a large “grab bag” of miscellaneous programs. We hope you find them both interesting and enjoyable.

11.3.1 Finding Duplicated Words in a Document

A common error when writing large amounts of prose is to accidentally duplicate words. Typically you will see this in text as something like “the the program does the following…” When the text is online, often the duplicated words occur at the end of one line and the the beginning of another, making them very difficult to spot.

This program, dupword.awk, scans through a file one line at a time and looks for adjacent occurrences of the same word. It also saves the last word on a line (in the variable prev) for comparison with the first word on the next line.

The first two statements make sure that the line is all lowercase, so that, for example, “The” and “the” compare equal to each other. The next statement replaces nonalphanumeric and nonwhitespace characters with spaces, so that punctuation does not affect the comparison either. The characters are replaced with spaces so that formatting controls don’t create nonsense words (e.g., the Texinfo ‘@code{NF}’ becomes ‘codeNF’ if punctuation is simply deleted). The record is then resplit into fields, yielding just the actual words on the line, and ensuring that there are no empty fields.

If there are no fields left after removing all the punctuation, the current record is skipped. Otherwise, the program loops through each word, comparing it to the previous one:

     # dupword.awk --- find duplicate words in text
     {
         $0 = tolower($0)
         gsub(/[^[:alnum:][:blank:]]/, " ");
         $0 = $0         # re-split
         if (NF == 0)

             next
         if ($1 == prev)
             printf("%s:%d: duplicate %s\n",
                 FILENAME, FNR, $1)
         for (i = 2; i <= NF; i++)
             if ($i == $(i-1))
                 printf("%s:%d: duplicate %s\n",
                     FILENAME, FNR, $i)
         prev = $NF
     }

11.3.2 An Alarm Clock Program

Nothing cures insomnia like a ringing alarm clock.

—Arnold Robbins Sleep is for web developers. —Erik Quanstrom

The following program is a simple “alarm clock” program. You give it a time of day and an optional message. At the specified time, it prints the message on the standard output. In addition, you can give it the number of times to repeat the message as well as a delay between repetitions.

This program uses the getlocaltime() function from Section 10.2.7 [Managing the Time of Day], page 225.

All the work is done in the BEGIN rule. The first part is argument checking and setting of defaults: the delay, the count, and the message to print. If the user supplied a message without the ASCII BEL character (known as the “alert” character, "\a"), then it is added to the message. (On many systems, printing the ASCII BEL generates an audible alert. Thus, when the alarm goes off, the system calls attention to itself in case the user is not looking at the computer.) Just for a change, this program uses a switch statement (see Section 7.4.5 [The switch Statement], page 141), but the processing could be done with a series of if-else statements instead. Here is the program:

     # alarm.awk --- set an alarm
     #
     # Requires getlocaltime() library function
     # usage: alarm time [ "message" [ count [ delay ] ] ]
     BEGIN {
         # Initial argument sanity checking
         usage1 = "usage: alarm time ['message' [count [delay]]]"
         usage2 = sprintf("\t(%s) time ::= hh:mm", ARGV[1])
         if (ARGC < 2) {
             print usage1 > "/dev/stderr"
             print usage2 > "/dev/stderr"
             exit 1
         }
         switch (ARGC) {

case 5:
    delay = ARGV[4] + 0
    # fall through
case 4:
    count = ARGV[3] + 0
    # fall through
case 3:
    message = ARGV[2]
    break
default:
    if (ARGV[1] !~ /[[:digit:]]?[[:digit:]]:[[:digit:]]{2}/) {
        print usage1 > "/dev/stderr"
        print usage2 > "/dev/stderr"
        exit 1
    }
    break
}
# set defaults for once we reach the desired time
if (delay == 0)
    delay = 180    # 3 minutes
if (count == 0)
    count = 5
if (message == "")
    message = sprintf("\aIt is now %s!\a", ARGV[1])
else if (index(message, "\a") == 0)
    message = "\a" message "\a"

The next section of code turns the alarm time into hours and minutes, converts it (if necessary) to a 24-hour clock, and then turns that time into a count of the seconds since midnight. Next it turns the current time into a count of seconds since midnight. The difference between the two is how long to wait before setting off the alarm:

# split up alarm time
split(ARGV[1], atime, ":")
hour = atime[1] + 0    # force numeric
minute = atime[2] + 0  # force numeric
# get current broken down time
getlocaltime(now)
# if time given is 12-hour hours and it's after that
# hour, e.g., `alarm 5:30' at 9 a.m. means 5:30 p.m.,
# then add 12 to real hour
if (hour < 12 && now["hour"] > hour)
    hour += 12
# set target time in seconds since midnight

target = (hour * 60 * 60) + (minute * 60)
# get current time in seconds since midnight
current = (now["hour"] * 60 * 60) + \
           (now["minute"] * 60) + now["second"]
# how long to sleep for
naptime = target - current
if (naptime <= 0) {
    print "alarm: time is in the past!" > "/dev/stderr"
    exit 1
}

Finally, the program uses the system() function (see Section 9.1.4 [Input/Output Functions], page 189) to call the sleep utility. The sleep utility simply pauses for the given number of seconds. If the exit status is not zero, the program assumes that sleep was interrupted and exits. If sleep exited with an OK status (zero), then the program prints the message in a loop, again using sleep to delay for however many seconds are necessary:

# zzzzzz..... go away if interrupted
if (system(sprintf("sleep %d", naptime)) != 0)
    exit 1
# time to notify!
command = sprintf("sleep %d", delay)
for (i = 1; i <= count; i++) {
    print message
    # if sleep command interrupted, go away
    if (system(command) != 0)
        break
}
exit 0
}

11.3.3 Transliterating Characters

The system tr utility transliterates characters. For example, it is often used to map uppercase letters into lowercase for further processing:

generate data | tr 'A-Z' 'a-z' | process data ...

tr requires two lists of characters.^{4} When processing the input, the first character in the first list is replaced with the first character in the second list, the second character in the first list is replaced with the second character in the second list, and so on. If there are more characters in the “from” list than in the “to” list, the last character of the “to” list is used for the remaining characters in the “from” list.

4 On some older systems, including Solaris, the system version of tr may require that the lists be written as

range expressions enclosed in square brackets (‘[a-z]’) and quoted, to prevent the shell from attempting a file name expansion. This is not a feature.

Once upon a time, a user proposed adding a transliteration function to gawk. The following program was written to prove that character transliteration could be done with a user-level function. This program is not as complete as the system tr utility, but it does most of the job.

The translate program was written long before gawk acquired the ability to split each character in a string into separate array elements. Thus, it makes repeated use of the substr(), index(), and gsub() built-in functions (see Section 9.1.3 [String-Manipulation Functions], page 177). There are two functions. The first, stranslate(), takes three arguments:

from A list of characters from which to translate

to A list of characters to which to translate

target The string on which to do the translation

Associative arrays make the translation part fairly easy. t_ar holds the “to” characters, indexed by the “from” characters. Then a simple loop goes through from, one character at a time. For each character in from, if the character appears in target, it is replaced with the corresponding to character.

The translate() function calls stranslate(), using $0 as the target. The main program sets two global variables, FROM and TO, from the command line, and then changes ARGV so that awk reads from the standard input.

Finally, the processing rule simply calls translate() for each record:

     # translate.awk --- do tr-like stuff
     # Bugs: does not handle things like tr A-Z a-z; it has
     # to be spelled out. However, if `to' is shorter than `from',
     # the last character in `to' is used for the rest of `from'.
     function stranslate(from, to, target,     lf, lt, ltarget, t_ar, i, c,
                                                                    result)
     {
         lf = length(from)
         lt = length(to)
         ltarget = length(target)
         for (i = 1; i <= lt; i++)
             t_ar[substr(from, i, 1)] = substr(to, i, 1)
         if (lt < lf)
             for (; i <= lf; i++)
                 t_ar[substr(from, i, 1)] = substr(to, lt, 1)
         for (i = 1; i <= ltarget; i++) {
             c = substr(target, i, 1)
             if (c in t_ar)
                 c = t_ar[c]
             result = result c
         }
         return result
     }

     function translate(from, to)
     {
         return $0 = stranslate(from, to, $0)
     }
     # main program
     BEGIN {
         if (ARGC < 3) {
             print "usage: translate from to" > "/dev/stderr"
             exit
         }
         FROM = ARGV[1]
         TO = ARGV[2]
         ARGC = 2
         ARGV[1] = "-"
     }
     {
         translate(FROM, TO)
         print
     }

It is possible to do character transliteration in a user-level function, but it is not necessarily efficient, and we (the gawk developers) started to consider adding a built-in function. However, shortly after writing this program, we learned that Brian Kernighan had added the toupper() and tolower() functions to his awk (see Section 9.1.3 [String-Manipulation Functions], page 177). These functions handle the vast majority of the cases where character transliteration is necessary, and so we chose to simply add those functions to gawk as well and then leave well enough alone.

An obvious improvement to this program would be to set up the t_ar array only once, in a BEGIN rule. However, this assumes that the “from” and “to” lists will never change throughout the lifetime of the program.

Another obvious improvement is to enable the use of ranges, such as ‘a-z’, as allowed by the tr utility. Look at the code for cut.awk (see Section 11.2.1 [Cutting Out Fields and Columns], page 251) for inspiration.

11.3.4 Printing Mailing Labels

Here is a “real-world”^{5} program. This script reads lists of names and addresses and generates mailing labels. Each page of labels has 20 labels on it, two across and 10 down. The addresses are guaranteed to be no more than five lines of data. Each address is separated from the next by a blank line.

The basic idea is to read 20 labels’ worth of data. Each line of each label is stored in the line array. The single rule takes care of filling the line array and printing the page when 20 labels have been read.

5 “Real world” is defined as “a program actually used to get something done.”

The BEGIN rule simply sets RS to the empty string, so that awk splits records at blank lines (see Section 4.1 [How Input Is Split into Records], page 59). It sets MAXLINES to 100, because 100 is the maximum number of lines on the page (20 · 5 = 100).

Most of the work is done in the printpage() function. The label lines are stored sequentially in the line array. But they have to print horizontally: line[1] next to line[6], line[2] next to line[7], and so on. Two loops accomplish this. The outer loop, controlled by i, steps through every 10 lines of data; this is each row of labels. The inner loop, controlled by j, goes through the lines within the row. As j goes from 0 to 4, ‘i+j’ is the jth line in the row, and ‘i+j+5’ is the entry next to it. The output ends up looking something like this:

     line 1          line 6
     line 2          line 7
     line 3          line 8
     line 4          line 9
     line 5          line 10
     ...

The printf format string ‘%-41s’ left-aligns the data and prints it within a fixed-width field.

As a final note, an extra blank line is printed at lines 21 and 61, to keep the output lined up on the labels. This is dependent on the particular brand of labels in use when the program was written. You will also note that there are two blank lines at the top and two blank lines at the bottom.

The END rule arranges to flush the final page of labels; there may not have been an even multiple of 20 labels in the data:

     # labels.awk --- print mailing labels
     # Each label is 5 lines of data that may have blank lines.
     # The label sheets have 2 blank lines at the top and 2 at
     # the bottom.
     BEGIN    { RS = "" ; MAXLINES = 100 }
     function printpage(    i, j)
     {
         if (Nlines <= 0)
             return
         printf "\n\n"        # header
         for (i = 1; i <= Nlines; i += 10) {
             if (i == 21 || i == 61)
                 print ""
             for (j = 0; j < 5; j++) {
                 if (i + j > MAXLINES)
                     break

        printf "   %-41s %s\n", line[i+j], line[i+j+5]
    }
    print ""
}
printf "\n\n"        # footer
delete line
}
# main rule
{
if (Count >= 20) {
    printpage()
    Count = 0
    Nlines = 0
}
n = split($0, a, "\n")
for (i = 1; i <= n; i++)
    line[++Nlines] = a[i]
for (; i <= 5; i++)
    line[++Nlines] = ""
Count++
}
END {
printpage()
}

11.3.5 Generating Word-Usage Counts

When working with large amounts of text, it can be interesting to know how often different words appear. For example, an author may overuse certain words, in which case he or she might wish to find synonyms to substitute for words that appear too often. This subsection develops a program for counting words and presenting the frequency information in a useful format.

At first glance, a program like this would seem to do the job:

# wordfreq-first-try.awk --- print list of word frequencies
{
for (i = 1; i <= NF; i++)
    freq[$i]++
}
END {
for (word in freq)
    printf "%s\t%d\n", word, freq[word]

     }

The program relies on awk’s default field-splitting mechanism to break each line up into “words” and uses an associative array named freq, indexed by each word, to count the number of times the word occurs. In the END rule, it prints the counts.

This program has several problems that prevent it from being useful on real text files:

• The awk language considers upper- and lowercase characters to be distinct. Therefore, “bartender” and “Bartender” are not treated as the same word. This is undesirable, because words are capitalized if they begin sentences in normal text, and a frequency analyzer should not be sensitive to capitalization.

• Words are detected using the awk convention that fields are separated just by whitespace. Other characters in the input (except newlines) don’t have any special meaning to awk. This means that punctuation characters count as part of words.

• The output does not come out in any useful order. You’re more likely to be interested in which words occur most frequently or in having an alphabetized table of how frequently each word occurs.

The first problem can be solved by using tolower() to remove case distinctions. The second problem can be solved by using gsub() to remove punctuation characters. Finally, we solve the third problem by using the system sort utility to process the output of the awk script. Here is the new version of the program:

     # wordfreq.awk --- print list of word frequencies
     {
         $0 = tolower($0)    # remove case distinctions
         # remove punctuation
         gsub(/[^[:alnum:]_[:blank:]]/, "", $0)
         for (i = 1; i <= NF; i++)
             freq[$i]++
     }
     END {
         for (word in freq)
             printf "%s\t%d\n", word, freq[word]
     }
   The regexp /[^[:alnum:]_[:blank:]]/ might have been written /[[:punct:]]/, but

then underscores would also be removed, and we want to keep them.

Assuming we have saved this program in a file named wordfreq.awk, and that the data is in file1, the following pipeline:

     awk -f wordfreq.awk file1 | sort -k 2nr

produces a table of the words appearing in file1 in order of decreasing frequency.

The awk program suitably massages the data and produces a word frequency table, which is not ordered. The awk script’s output is then sorted by the sort utility and printed on the screen.

The options given to sort specify a sort that uses the second field of each input line (skipping one field), that the sort keys should be treated as numeric quantities (otherwise ‘15’ would come before ‘5’), and that the sorting should be done in descending (reverse) order.

The sort could even be done from within the program, by changing the END action to:

     END {
         sort = "sort -k 2nr"
         for (word in freq)
             printf "%s\t%d\n", word, freq[word] | sort
         close(sort)
     }

This way of sorting must be used on systems that do not have true pipes at the commandline (or batch-file) level. See the general operating system documentation for more information on how to use the sort program.

11.3.6 Removing Duplicates from Unsorted Text

The uniq program (see Section 11.2.6 [Printing Nonduplicated Lines of Text], page 265) removes duplicate lines from sorted data.

Suppose, however, you need to remove duplicate lines from a data file but that you want to preserve the order the lines are in. A good example of this might be a shell history file. The history file keeps a copy of all the commands you have entered, and it is not unusual to repeat a command several times in a row. Occasionally you might want to compact the history by removing duplicate entries. Yet it is desirable to maintain the order of the original commands.

This simple program does the job. It uses two arrays. The data array is indexed by the text of each line. For each line, data[$0] is incremented. If a particular line has not been seen before, then data[$0] is zero. In this case, the text of the line is stored in lines[count]. Each element of lines is a unique command, and the indices of lines indicate the order in which those lines are encountered. The END rule simply prints out the lines, in order:

     # histsort.awk --- compact a shell history file
     # Thanks to Byron Rakitzis for the general idea
     {
         if (data[$0]++ == 0)
             lines[++count] = $0
     }
     END {
         for (i = 1; i <= count; i++)
             print lines[i]
     }

This program also provides a foundation for generating other useful information. For example, using the following print statement in the END rule indicates how often a particular command is used:

     print data[lines[i]], lines[i]

This works because data[$0] is incremented each time a line is seen.

11.3.7 Extracting Programs from Texinfo Source Files

Both this chapter and the previous chapter (Chapter 10 [A Library of awk Functions], page 217) present a large number of awk programs. If you want to experiment with these programs, it is tedious to type them in by hand. Here we present a program that can extract parts of a Texinfo input file into separate files.

   This book is written in Texinfo (http://www.gnu.org/software/texinfo/), the

GNU Project’s document formatting language. A single Texinfo source file can be used to produce both printed documentation, with TEX, and online documentation. (Texinfo is fully documented in the book Texinfo—The GNU Documentation Format, available from the Free Software Foundation, and also available online (http://www.gnu.org/software/

texinfo/manual/texinfo/).)

For our purposes, it is enough to know three things about Texinfo input files:

• The “at” symbol (‘@’) is special in Texinfo, much as the backslash (‘\’) is in C or awk. Literal ‘@’ symbols are represented in Texinfo source files as ‘@@’.

• Comments start with either ‘@c’ or ‘@comment’. The file-extraction program works by using special comments that start at the beginning of a line.

• Lines containing ‘@group’ and ‘@end group’ commands bracket example text that should not be split across a page boundary. (Unfortunately, TEX isn’t always smart enough to do things exactly right, so we have to give it some help.)

The following program, extract.awk, reads through a Texinfo source file and does two things, based on the special comments. Upon seeing ‘@c system ...’, it runs a command, by extracting the command text from the control line and passing it on to the system() function (see Section 9.1.4 [Input/Output Functions], page 189). Upon seeing ‘@c file filename’, each subsequent line is sent to the file filename, until ‘@c endfile’ is encountered. The rules in extract.awk match either ‘@c’ or ‘@comment’ by letting the ‘omment’ part be optional. Lines containing ‘@group’ and ‘@end group’ are simply removed. extract.awk uses the join() library function (see Section 10.2.6 [Merging an Array into a String], page 225).

The example programs in the online Texinfo source for GAWK: Effective AWK Programming (gawktexi.in) have all been bracketed inside ‘file’ and ‘endfile’ lines. The gawk distribution uses a copy of extract.awk to extract the sample programs and install many of them in a standard directory where gawk can find them. The Texinfo file looks something like this:

     ...
     This program has a @code{BEGIN} rule
     that prints a nice message:
     @example
     @c file examples/messages.awk
     BEGIN @{ print "Don't panic!" @}
     @c end file
     @end example
     It also prints some final advice:

     @example
     @c file examples/messages.awk
     END @{ print "Always avoid bored archaeologists!" @}
     @c end file
     @end example
     ...

extract.awk begins by setting IGNORECASE to one, so that mixed upper- and lowercase letters in the directives won’t matter.

The first rule handles calling system(), checking that a command is given (NF is at least three) and also checking that the command exits with a zero exit status, signifying OK:

     # extract.awk --- extract files and run programs from Texinfo files
     BEGIN    { IGNORECASE = 1 }
     /^@c(omment)?[ \t]+system/ {
         if (NF < 3) {
             e = ("extract: " FILENAME ":" FNR)
             e = (e  ": badly formed `system' line")
             print e > "/dev/stderr"
             next
         }
         $1 = ""
         $2 = ""
         stat = system($0)
         if (stat != 0) {
             e = ("extract: " FILENAME ":" FNR)
             e = (e ": warning: system returned " stat)
             print e > "/dev/stderr"
         }
     }

The variable e is used so that the rule fits nicely on the page.

The second rule handles moving data into files. It verifies that a file name is given in the directive. If the file named is not the current file, then the current file is closed. Keeping the current file open until a new file is encountered allows the use of the ‘>’ redirection for printing the contents, keeping open-file management simple.

The for loop does the work. It reads lines using getline (see Section 4.9 [Explicit Input with getline], page 78). For an unexpected end-of-file, it calls the unexpected_eof() function. If the line is an “endfile” line, then it breaks out of the loop. If the line is an ‘@group’ or ‘@end group’ line, then it ignores it and goes on to the next line. Similarly, comments within examples are also ignored.

Most of the work is in the following few lines. If the line has no ‘@’ symbols, the program can print it directly. Otherwise, each leading ‘@’ must be stripped off. To remove the ‘@’ symbols, the line is split into separate elements of the array a, using the split() function (see Section 9.1.3 [String-Manipulation Functions], page 177). The ‘@’ symbol is used as the separator character. Each element of a that is empty indicates two successive ‘@’ symbols in the original line. For each two empty elements (‘@@’ in the original file), we have to add a single ‘@’ symbol back in.

When the processing of the array is finished, join() is called with the value of SUBSEP (see Section 8.5 [Multidimensional Arrays], page 169), to rejoin the pieces back into a single line. That line is then printed to the output file:

/^@c(omment)?[ \t]+file/ {
if (NF != 3) {
e = ("extract: " FILENAME ":" FNR ": badly formed `file' line")
print e > "/dev/stderr"
next
}
if ($3 != curfile) {
if (curfile != "")
    close(curfile)
curfile = $3
}
for (;;) {
if ((getline line) <= 0)
    unexpected_eof()
if (line ~ /^@c(omment)?[ \t]+endfile/)
    break
else if (line ~ /^@(end[ \t]+)?group/)
    continue
else if (line ~ /^@c(omment+)?[ \t]+/)
    continue
if (index(line, "@") == 0) {
    print line > curfile
    continue
}
n = split(line, a, "@")
# if a[1] == "", means leading @,
# don't add one back in.
for (i = 2; i <= n; i++) {
    if (a[i] == "") { # was an @@
        a[i] = "@"
        if (a[i+1] == "")
            i++
    }
}
print join(a, 1, n, SUBSEP) > curfile
}
}

An important thing to note is the use of the ‘>’ redirection. Output done with ‘>’ only opens the file once; it stays open and subsequent output is appended to the file (see Section 5.6 [Redirecting Output of print and printf], page 98). This makes it easy to mix program text and explanatory prose for the same sample source file (as has been done here!) without any hassle. The file is only closed when a new data file name is encountered or at the end of the input file.

Finally, the function unexpected_eof() prints an appropriate error message and then exits. The END rule handles the final cleanup, closing the open file:

     function unexpected_eof()
     {
         printf("extract: %s:%d: unexpected EOF or error\n",
                          FILENAME, FNR) > "/dev/stderr"
         exit 1
     }
     END {
         if (curfile)
             close(curfile)
     }

11.3.8 A Simple Stream Editor

The sed utility is a stream editor, a program that reads a stream of data, makes changes to it, and passes it on. It is often used to make global changes to a large file or to a stream of data generated by a pipeline of commands. Although sed is a complicated program in its own right, its most common use is to perform global substitutions in the middle of a pipeline:

     command1 < orig.data | sed 's/old/new/g' | command2 > result

Here, ‘s/old/new/g’ tells sed to look for the regexp ‘old’ on each input line and globally replace it with the text ‘new’ (i.e., all the occurrences on a line). This is similar to awk’s gsub() function (see Section 9.1.3 [String-Manipulation Functions], page 177).

The following program, awksed.awk, accepts at least two command-line arguments: the pattern to look for and the text to replace it with. Any additional arguments are treated as data file names to process. If none are provided, the standard input is used:

     # awksed.awk --- do s/foo/bar/g using just print
     #    Thanks to Michael Brennan for the idea
     function usage()
     {
         print "usage: awksed pat repl [files...]" > "/dev/stderr"
         exit 1
     }
     BEGIN {
         # validate arguments
         if (ARGC < 3)
             usage()

         RS = ARGV[1]
         ORS = ARGV[2]
         # don't use arguments as files
         ARGV[1] = ARGV[2] = ""
     }
     # look ma, no hands!
     {
         if (RT == "")
             printf "%s", $0
         else
             print
     }

The program relies on gawk’s ability to have RS be a regexp, as well as on the setting of RT to the actual text that terminates the record (see Section 4.1 [How Input Is Split into Records], page 59).

The idea is to have RS be the pattern to look for. gawk automatically sets $0 to the text between matches of the pattern. This is text that we want to keep, unmodified. Then, by setting ORS to the replacement text, a simple print statement outputs the text we want to keep, followed by the replacement text.

There is one wrinkle to this scheme, which is what to do if the last record doesn’t end with text that matches RS. Using a print statement unconditionally prints the replacement text, which is not correct. However, if the file did not end in text that matches RS, RT is set to the null string. In this case, we can print $0 using printf (see Section 5.5 [Using printf Statements for Fancier Printing], page 92).

The BEGIN rule handles the setup, checking for the right number of arguments and calling usage() if there is a problem. Then it sets RS and ORS from the command-line arguments and sets ARGV[1] and ARGV[2] to the null string, so that they are not treated as file names (see Section 7.5.3 [Using ARGC and ARGV], page 155).

The usage() function prints an error message and exits. Finally, the single rule handles the printing scheme outlined earlier, using print or printf as appropriate, depending upon the value of RT.

11.3.9 An Easy Way to Use Library Functions

In Section 2.7 [Including Other Files into Your Program], page 42, we saw how gawk provides a built-in file-inclusion capability. However, this is a gawk extension. This section provides the motivation for making file inclusion available for standard awk, and shows how to do it using a combination of shell and awk programming.

Using library functions in awk can be very beneficial. It encourages code reuse and the writing of general functions. Programs are smaller and therefore clearer. However, using library functions is only easy when writing awk programs; it is painful when running them, requiring multiple -f options. If gawk is unavailable, then so too is the AWKPATH environment variable and the ability to put awk functions into a library directory (see Section 2.2 [Command-Line Options], page 31). It would be nice to be able to write programs in the following manner:

     # library functions
     @include getopt.awk
     @include join.awk
     ...
     # main program
     BEGIN {
         while ((c = getopt(ARGC, ARGV, "a:b:cde")) != -1)
             ...
         ...
     }

The following program, igawk.sh, provides this service. It simulates gawk’s searching of the AWKPATH variable and also allows nested includes (i.e., a file that is included with @include can contain further @include statements). igawk makes an effort to only include files once, so that nested includes don’t accidentally include a library function twice.

igawk should behave just like gawk externally. This means it should accept all of gawk’s command-line arguments, including the ability to have multiple source files specified via -f and the ability to mix command-line and library source files.

The program is written using the POSIX Shell (sh) command language.^{6} It works as follows:

1. Loop through the arguments, saving anything that doesn’t represent awk source code for later, when the expanded program is run.

2. For any arguments that do represent awk text, put the arguments into a shell variable that will be expanded. There are two cases: a. Literal text, provided with -e or --source. This text is just appended directly. b. Source file names, provided with -f. We use a neat trick and append ‘@include filename’ to the shell variable’s contents. Because the file-inclusion program works the way gawk does, this gets the text of the file included in the program at the correct point.

3. Run an awk program (naturally) over the shell variable’s contents to expand @include statements. The expanded program is placed in a second shell variable.

4. Run the expanded program with gawk and any other original command-line arguments that the user supplied (such as the data file names).

This program uses shell variables extensively: for storing command-line arguments and the text of the awk program that will expand the user’s program, for the user’s original program, and for the expanded program. Doing so removes some potential problems that might arise were we to use temporary files instead, at the cost of making the script somewhat more complicated.

The initial part of the program turns on shell tracing if the first argument is ‘debug’. The next part loops through all the command-line arguments. There are several cases

of interest:

6 Fully explaining the sh language is beyond the scope of this book. We provide some minimal explanations,

but see a good shell programming book if you wish to understand things in more depth.

-- This ends the arguments to igawk. Anything else should be passed on to the user’s awk program without being evaluated.

-W This indicates that the next option is specific to gawk. To make argument processing easier, the -W is appended to the front of the remaining arguments and the loop continues. (This is an sh programming trick. Don’t worry about it if you are not familiar with sh.)

-v, -F These are saved and passed on to gawk.

-f, --file, --file=, -Wfile=

The file name is appended to the shell variable program with an @include statement. The expr utility is used to remove the leading option part of the argument (e.g., ‘--file=’). (Typical sh usage would be to use the echo and sed utilities to do this work. Unfortunately, some versions of echo evaluate escape sequences in their arguments, possibly mangling the program text. Using expr avoids this problem.)

--source, --source=, -Wsource=

The source text is appended to program.

--version, -Wversion

igawk prints its version number, runs ‘gawk --version’ to get the gawk version information, and then exits.

If none of the -f, --file, -Wfile, --source, or -Wsource arguments are supplied, then the first nonoption argument should be the awk program. If there are no command-line arguments left, igawk prints an error message and exits. Otherwise, the first argument is appended to program. In any case, after the arguments have been processed, the shell variable program contains the complete text of the original awk program.

The program is as follows:

#! /bin/sh
# igawk --- like gawk but do @include processing
if [ "$1" = debug ]
then
set -x
shift
fi
# A literal newline, so that program text is formatted correctly
n='
'
# Initialize variables to empty
program=
opts=
while [ $# -ne 0 ] # loop over arguments
do

case $1 in
--)     shift
break ;;
-W)     shift
# The ${x?'message here'} construct prints a
# diagnostic if $x is the null string
set -- -W"${@?'missing operand'}"
continue ;;
-[vF])  opts="$opts $1 '${2?'missing operand'}'"
shift ;;
-[vF]*) opts="$opts '$1'" ;;
-f)     program="$program$n@include ${2?'missing operand'}"
shift ;;
-f*)    f=$(expr "$1" : '-f\(.*\)')
program="$program$n@include $f" ;;
-[W-]file=*)
f=$(expr "$1" : '-.file=\(.*\)')
program="$program$n@include $f" ;;
-[W-]file)
program="$program$n@include ${2?'missing operand'}"
shift ;;
-[W-]source=*)
t=$(expr "$1" : '-.source=\(.*\)')
program="$program$n$t" ;;
-[W-]source)
program="$program$n${2?'missing operand'}"
shift ;;
-[W-]version)
echo igawk: version 3.0 1>&2
gawk --version
exit 0 ;;
-[W-]*) opts="$opts '$1'" ;;
*)      break ;;
esac
shift

     done
     if [ -z "$program" ]
     then
          program=${1?'missing program'}
          shift
     fi
     # At this point, `program' has the program.

The awk program to process @include directives is stored in the shell variable expand_ prog. Doing this keeps the shell script readable. The awk program reads through the user’s program, one line at a time, using getline (see Section 4.9 [Explicit Input with getline], page 78). The input file names and @include statements are managed using a stack. As each @include is encountered, the current file name is “pushed” onto the stack and the file named in the @include directive becomes the current file name. As each file is finished, the stack is “popped,” and the previous input file becomes the current input file again. The process is started by making the original file the first one on the stack.

The pathto() function does the work of finding the full path to a file. It simulates gawk’s behavior when searching the AWKPATH environment variable (see Section 2.5.1 [The AWKPATH Environment Variable], page 39). If a file name has a ‘/’ in it, no path search is done. Similarly, if the file name is "-", then that string is used as-is. Otherwise, the file name is concatenated with the name of each directory in the path, and an attempt is made to open the generated file name. The only way to test if a file can be read in awk is to go ahead and try to read it with getline; this is what pathto() does.^{7} If the file can be read, it is closed and the file name is returned:

     expand_prog='
     function pathto(file,    i, t, junk)
     {
         if (index(file, "/") != 0)
             return file
         if (file == "-")
             return file
         for (i = 1; i <= ndirs; i++) {
             t = (pathlist[i] "/" file)
             if ((getline junk < t) > 0) {
                 # found it
                 close(t)
                 return t
             }
         }
         return ""

7 On some very old versions of awk, the test ‘getline junk < t’ can loop forever if the file exists but is

empty.

     }

The main program is contained inside one BEGIN rule. The first thing it does is set up the pathlist array that pathto() uses. After splitting the path on ‘:’, null elements are replaced with ".", which represents the current directory:

     BEGIN {
         path = ENVIRON["AWKPATH"]
         ndirs = split(path, pathlist, ":")
         for (i = 1; i <= ndirs; i++) {
             if (pathlist[i] == "")
                 pathlist[i] = "."
         }

The stack is initialized with ARGV[1], which will be "/dev/stdin". The main loop comes next. Input lines are read in succession. Lines that do not start with @include are printed verbatim. If the line does start with @include, the file name is in $2. pathto() is called to generate the full path. If it cannot, then the program prints an error message and continues.

The next thing to check is if the file is included already. The processed array is indexed by the full file name of each included file and it tracks this information for us. If the file is seen again, a warning message is printed. Otherwise, the new file name is pushed onto the stack and processing continues.

Finally, when getline encounters the end of the input file, the file is closed and the stack is popped. When stackptr is less than zero, the program is done:

         stackptr = 0
         input[stackptr] = ARGV[1] # ARGV[1] is first file
         for (; stackptr >= 0; stackptr--) {
             while ((getline < input[stackptr]) > 0) {
                 if (tolower($1) != "@include") {
                     print
                     continue
                 }
                 fpath = pathto($2)
                 if (fpath == "") {
                     printf("igawk: %s:%d: cannot find %s\n",
                         input[stackptr], FNR, $2) > "/dev/stderr"
                     continue
                 }
                 if (! (fpath in processed)) {
                     processed[fpath] = input[stackptr]
                     input[++stackptr] = fpath  # push onto stack
                 } else
                     print $2, "included in", input[stackptr],
                         "already included in",
                         processed[fpath] > "/dev/stderr"
             }

             close(input[stackptr])
         }
     }'  # close quote ends `expand_prog' variable
     processed_program=$(gawk -- "$expand_prog" /dev/stdin << EOF
     $program
     EOF
     )

The shell construct ‘command << marker’ is called a here document. Everything in the shell script up to the marker is fed to command as input. The shell processes the contents of the here document for variable and command substitution (and possibly other things as well, depending upon the shell).

The shell construct ‘$(...)’ is called command substitution. The output of the command inside the parentheses is substituted into the command line. Because the result is used in a variable assignment, it is saved as a single string, even if the results contain whitespace.

The expanded program is saved in the variable processed_program. It’s done in these steps:

1. Run gawk with the @include-processing program (the value of the expand_prog shell variable) reading standard input.

2. Standard input is the contents of the user’s program, from the shell variable program. Feed its contents to gawk via a here document.

3. Save the results of this processing in the shell variable processed_program by using command substitution.

The last step is to call gawk with the expanded program, along with the original options and command-line arguments that the user supplied:

     eval gawk $opts -- '"$processed_program"' '"$@"'

The eval command is a shell construct that reruns the shell’s parsing process. This keeps things properly quoted.

This version of igawk represents the fifth version of this program. There are four key simplifications that make the program work better:

• Using @include even for the files named with -f makes building the initial collected awk program much simpler; all the @include processing can be done once.

• Not trying to save the line read with getline in the pathto() function when testing for the file’s accessibility for use with the main program simplifies things considerably.

• Using a getline loop in the BEGIN rule does it all in one place. It is not necessary to call out to a separate loop for processing nested @include statements.

• Instead of saving the expanded program in a temporary file, putting it in a shell variable avoids some potential security problems. This has the disadvantage that the script relies upon more features of the sh language, making it harder to follow for those who aren’t familiar with sh.

Also, this program illustrates that it is often worthwhile to combine sh and awk programming together. You can usually accomplish quite a lot, without having to resort to low-level programming in C or C++, and it is frequently easier to do certain kinds of string and argument manipulation using the shell than it is in awk.

Finally, igawk shows that it is not always necessary to add new features to a program; they can often be layered on top.^{8}

11.3.10 Finding Anagrams from a Dictionary

An interesting programming challenge is to search for anagrams in a word list (such as /usr/share/dict/words on many GNU/Linux systems). One word is an anagram of another if both words contain the same letters (e.g., “babbling” and “blabbing”).

Column 2, Problem C, of Jon Bentley’s Programming Pearls, Second Edition, presents an elegant algorithm. The idea is to give words that are anagrams a common signature, sort all the words together by their signatures, and then print them. Dr. Bentley observes that taking the letters in each word and sorting them produces those common signatures.

The following program uses arrays of arrays to bring together words with the same signature and array sorting to print the words in sorted order:

     # anagram.awk --- An implementation of the anagram-finding algorithm
     #                 from Jon Bentley's "Programming Pearls," 2nd edition.
     #                 Addison Wesley, 2000, ISBN 0-201-65788-0.
     #                 Column 2, Problem C, section 2.8, pp 18-20.
     /'s$/   { next }        # Skip possessives

The program starts with a header, and then a rule to skip possessives in the dictionary file. The next rule builds up the data structure. The first dimension of the array is indexed by the signature; the second dimension is the word itself:

     {
         key = word2key($1)  # Build signature
         data[key][$1] = $1  # Store word with signature
     }

The word2key() function creates the signature. It splits the word apart into individual letters, sorts the letters, and then joins them back together:

     # word2key --- split word apart into letters, sort, and join back together
     function word2key(word,     a, i, n, result)
     {
         n = split(word, a, "")
         asort(a)
         for (i = 1; i <= n; i++)
             result = result a[i]
         return result
     }

8 gawk does @include processing itself in order to support the use of awk programs as Web CGI scripts.

Finally, the END rule traverses the array and prints out the anagram lists. It sends the output to the system sort command because otherwise the anagrams would appear in arbitrary order:

END {
    sort = "sort"
    for (key in data) {
        # Sort words with same key
        nwords = asorti(data[key], words)
        if (nwords == 1)
            continue
        # And print. Minor glitch: trailing space at end of each line
        for (j = 1; j <= nwords; j++)
            printf("%s ", words[j]) | sort
        print "" | sort
    }
    close(sort)
}

Here is some partial output when the program is run:

$ gawk -f anagram.awk /usr/share/dict/words | grep '^b'
...
babbled blabbed
babbler blabber brabble
babblers blabbers brabbles
babbling blabbing
babbly blabby
babel bable
babels beslab
babery yabber
...

11.3.11 And Now for Something Completely Different

The following program was written by Davide Brini and is published on his website (http://

backreference.org/2011/02/03/obfuscated-awk/). It serves as his signature in the

Usenet group comp.lang.awk. He supplies the following copyright terms:

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2008 Davide Brini

Copying and distribution of the code published in this page, with or without modification, are permitted in any medium without royalty provided the copyright notice and this notice are preserved.

Here is the program:

awk 'BEGIN{O="~"~"~";o="=="=="==";o+=+o;x=O""O;while(X++<=x+o+o)c=c"%c";
printf c,(x-O)*(x-O),x*(x-o)-o,x*(x-O)+x-O-o,+x*(x-O)-x+o,X*(o*o+O)+x-O,
X*(X-x)-o*o,(x+X)*o*o+o,x*(X-x)-O-O,x-O+(O+o+X+x)*(o+O),X*X-X*(x-O)-x+O,
O+X*(o*(o+O)+O),+x+O+X*o,x*(x-o),(o+X+x)*o*o-(x-O-O),O+(X-x)*(X+O),x-O}'

We leave it to you to determine what the program does. (If you are truly desperate to understand it, see Chris Johansen’s explanation, which is embedded in the Texinfo source file for this book.)

11.4 Summary

• The programs provided in this chapter continue on the theme that reading programs is an excellent way to learn Good Programming.

• Using ‘#!’ to make awk programs directly runnable makes them easier to use. Otherwise, invoke the program using ‘awk -f ...’.

• Reimplementing standard POSIX programs in awk is a pleasant exercise; awk’s expressive power lets you write such programs in relatively few lines of code, yet they are functionally complete and usable.

• One of standard awk’s weaknesses is working with individual characters. The ability to use split() with the empty string as the separator can considerably simplify such tasks.

• The examples here demonstrate the usefulness of the library functions from Chapter 10 [A Library of awk Functions], page 217, for a number of real (if small) programs.

• Besides reinventing POSIX wheels, other programs solved a selection of interesting problems, such as finding duplicate words in text, printing mailing labels, and finding anagrams.

11.5 Exercises

1. Rewrite cut.awk (see Section 11.2.1 [Cutting Out Fields and Columns], page 251) using split() with "" as the separator.

2. In Section 11.2.2 [Searching for Regular Expressions in Files], page 256, we mentioned that ‘egrep -i’ could be simulated in versions of awk without IGNORECASE by using tolower() on the line and the pattern. In a footnote there, we also mentioned that this solution has a bug: the translated line is output, and not the original one. Fix this problem.

3. The POSIX version of id takes options that control which information is printed. Modify the awk version (see Section 11.2.3 [Printing Out User Information], page 260) to accept the same arguments and perform in the same way.

4. The split.awk program (see Section 11.2.4 [Splitting a Large File into Pieces], page 262) assumes that letters are contiguous in the character set, which isn’t true for EBCDIC systems. Fix this problem. (Hint: Consider a different way to work through the alphabet, without relying on ord() and chr().)

5. In uniq.awk (see Section 11.2.6 [Printing Nonduplicated Lines of Text], page 265, the logic for choosing which lines to print represents a state machine, which is “a device that can be in one of a set number of stable conditions depending on its previous condition and on the present values of its inputs.”^{9} Brian Kernighan suggests that “an alternative approach to state machines is to just read the input into an array, then use indexing. It’s almost always easier code, and for most inputs where you would use this, just as fast.” Rewrite the logic to follow this suggestion.

9 This is the definition returned from entering define: state machine into Google.

6. Why can’t the wc.awk program (see Section 11.2.7 [Counting Things], page 269) just use the value of FNR in endfile()? Hint: Examine the code in Section 10.3.1 [Noting Data file Boundaries], page 229.

7. Manipulation of individual characters in the translate program (see Section 11.3.3 [Transliterating Characters], page 274) is painful using standard awk functions. Given that gawk can split strings into individual characters using "" as the separator, how might you use this feature to simplify the program?

8. The extract.awk program (see Section 11.3.7 [Extracting Programs from Texinfo Source Files], page 281) was written before gawk had the gensub() function. Use it to simplify the code.

9. Compare the performance of the awksed.awk program (see Section 11.3.8 [A Simple

Stream Editor], page 284) with the more straightforward:

     BEGIN {
         pat = ARGV[1]
         repl = ARGV[2]
         ARGV[1] = ARGV[2] = ""
     }
     { gsub(pat, repl); print }

10. What are the advantages and disadvantages of awksed.awk versus the real sed utility?

11. In Section 11.3.9 [An Easy Way to Use Library Functions], page 285, we mentioned that not trying to save the line read with getline in the pathto() function when testing for the file’s accessibility for use with the main program simplifies things considerably. What problem does this engender though?

12. As an additional example of the idea that it is not always necessary to add new features to a program, consider the idea of having two files in a directory in the search path:

default.awk

This file contains a set of default library functions, such as getopt() and

          assert().

site.awk This file contains library functions that are specific to a site or installation; i.e., locally developed functions. Having a separate file allows default.awk to change with new gawk releases, without requiring the system administrator to update it each time by adding the local functions.

One user suggested that gawk be modified to automatically read these files upon startup. Instead, it would be very simple to modify igawk to do this. Since igawk can process nested @include directives, default.awk could simply contain @include statements for the desired library functions. Make this change.

13. Modify anagram.awk (see Section 11.3.10 [Finding Anagrams from a Dictionary], page 292), to avoid the use of the external sort utility.

Part III: Moving Beyond Standard awk with gawk

Chapter 12: Advanced Features of gawk 299

12 Advanced Features of gawk

Write documentation as if whoever reads it is a violent psychopath who knows where you live.

—Steve English, as quoted by Peter Langston This chapter discusses advanced features in gawk. It’s a bit of a “grab bag” of items

that are otherwise unrelated to each other. First, we look at a command-line option that allows gawk to recognize nondecimal numbers in input data, not just in awk programs. Then, gawk’s special features for sorting arrays are presented. Next, two-way I/O, discussed briefly in earlier parts of this book, is described in full detail, along with the basics of TCP/IP networking. Finally, we see how gawk can profile an awk program, making it possible to tune it for performance.

Additional advanced features are discussed in separate chapters of their own:

• Chapter 13 [Internationalization with gawk], page 315, discusses how to internationalize your awk programs, so that they can speak multiple national languages.

• Chapter 14 [Debugging awk Programs], page 325, describes gawk’s built-in commandline debugger for debugging awk programs.

• Chapter 15 [Arithmetic and Arbitrary-Precision Arithmetic with gawk], page 341, describes how you can use gawk to perform arbitrary-precision arithmetic.

• Chapter 16 [Writing Extensions for gawk], page 355, discusses the ability to dynamically add new built-in functions to gawk.

12.1 Allowing Nondecimal Input Data

If you run gawk with the --non-decimal-data option, you can have nondecimal values in your input data:

     $ echo 0123 123 0x123 |
     > gawk --non-decimal-data '{ printf "%d, %d, %d\n", $1, $2, $3 }'
     ⊣ 83, 123, 291

For this feature to work, write your program so that gawk treats your data as numeric:

     $ echo 0123 123 0x123 | gawk '{ print $1, $2, $3 }'
     ⊣ 0123 123 0x123

The print statement treats its expressions as strings. Although the fields can act as numbers when necessary, they are still strings, so print does not try to treat them numerically. You need to add zero to a field to force it to be treated as a number. For example:

     $ echo 0123 123 0x123 | gawk --non-decimal-data '
     > { print $1, $2, $3
     >   print $1 + 0, $2 + 0, $3 + 0 }'
     ⊣ 0123 123 0x123
     ⊣ 83 123 291

Because it is common to have decimal data with leading zeros, and because using this facility could lead to surprising results, the default is to leave it disabled. If you want it, you must explicitly request it.

CAUTION: Use of this option is not recommended. It can break old programs very badly. Instead, use the strtonum() function to convert your data (see Section 9.1.3 [String-Manipulation Functions], page 177). This makes your programs easier to write and easier to read, and leads to less surprising results. This option may disappear in a future version of gawk.

12.2 Controlling Array Traversal and Array Sorting

gawk lets you control the order in which a ‘for (indx in array)’ loop traverses an array.

In addition, two built-in functions, asort() and asorti(), let you sort arrays based on the array values and indices, respectively. These two functions also provide control over the sorting criteria used to order the elements during sorting.

12.2.1 Controlling Array Traversal

By default, the order in which a ‘for (indx in array)’ loop scans an array is not defined; it is generally based upon the internal implementation of arrays inside awk.

Often, though, it is desirable to be able to loop over the elements in a particular order that you, the programmer, choose. gawk lets you do this.

Section 8.1.6 [Using Predefined Array Scanning Orders with gawk], page 164, describes how you can assign special, predefined values to PROCINFO["sorted_in"] in order to control the order in which gawk traverses an array during a for loop.

In addition, the value of PROCINFO["sorted_in"] can be a function name.^{1} This lets you traverse an array based on any custom criterion. The array elements are ordered according to the return value of this function. The comparison function should be defined with at least four arguments:

     function comp_func(i1, v1, i2, v2)
     {
         compare elements 1 and 2 in some fashion
         return < 0; 0; or > 0
     }

Here, i1 and i2 are the indices, and v1 and v2 are the corresponding values of the two elements being compared. Either v1 or v2, or both, can be arrays if the array being traversed contains subarrays as values. (See Section 8.6 [Arrays of Arrays], page 171, for more information about subarrays.) The three possible return values are interpreted as follows:

comp_func(i1, v1, i2, v2) < 0

Index i1 comes before index i2 during loop traversal.

comp_func(i1, v1, i2, v2) == 0

Indices i1 and i2 come together, but the relative order with respect to each other is undefined.

comp_func(i1, v1, i2, v2) > 0

Index i1 comes after index i2 during loop traversal.

Our first comparison function can be used to scan an array in numerical order of the indices:

1 This is why the predefined sorting orders start with an ‘@’ character, which cannot be part of an identifier.

function cmp_num_idx(i1, v1, i2, v2)
{
 # numerical index comparison, ascending order
 return (i1 - i2)
}

Our second function traverses an array based on the string order of the element values rather than by indices:

function cmp_str_val(i1, v1, i2, v2)
{
# string value comparison, ascending order
v1 = v1 ""
v2 = v2 ""
if (v1 < v2)
    return -1
return (v1 != v2)
}

The third comparison function makes all numbers, and numeric strings without any leading or trailing spaces, come out first during loop traversal:

function cmp_num_str_val(i1, v1, i2, v2,   n1, n2)
{
 # numbers before string value comparison, ascending order
 n1 = v1 + 0
 n2 = v2 + 0
 if (n1 == v1)
     return (n2 == v2) ? (n1 - n2) : -1
 else if (n2 == v2)
     return 1
 return (v1 < v2) ? -1 : (v1 != v2)
}

Here is a main program to demonstrate how gawk behaves using each of the previous functions:

BEGIN {
data["one"] = 10
data["two"] = 20
data[10] = "one"
data[100] = 100
data[20] = "two"
f[1] = "cmp_num_idx"
f[2] = "cmp_str_val"
f[3] = "cmp_num_str_val"
for (i = 1; i <= 3; i++) {
    printf("Sort function: %s\n", f[i])
    PROCINFO["sorted_in"] = f[i]
    for (j in data)

            printf("\tdata[%s] = %s\n", j, data[j])
        print ""
    }
}

Here are the results when the program is run:

$ gawk -f compdemo.awk
⊣ Sort function: cmp_num_idx  Sort by numeric index
⊣  data[two] = 20

⊣  data[one] = 10  Both strings are numerically zero

⊣  data[10] = one
⊣  data[20] = two
⊣  data[100] = 100

⊣ ⊣ Sort function: cmp_str_val Sort by element values as strings

⊣  data[one] = 10

⊣  data[100] = 100  String 100 is less than string 20

⊣  data[two] = 20
⊣  data[10] = one
⊣  data[20] = two

⊣ ⊣ Sort function: cmp_num_str_val Sort all numeric values before all strings

⊣  data[one] = 10
⊣  data[two] = 20
⊣  data[100] = 100
⊣  data[10] = one
⊣  data[20] = two

Consider sorting the entries of a GNU/Linux system password file according to login name. The following program sorts records by a specific field position and can be used for this purpose:

# passwd-sort.awk --- simple program to sort by field position
# field position is specified by the global variable POS
function cmp_field(i1, v1, i2, v2)
{
    # comparison by value, as string, and ascending order
    return v1[POS] < v2[POS] ? -1 : (v1[POS] != v2[POS])
}
{
    for (i = 1; i <= NF; i++)
        a[NR][i] = $i
}
END {
    PROCINFO["sorted_in"] = "cmp_field"
    if (POS < 1 || POS > NF)

             POS = 1
         for (i in a) {
             for (j = 1; j <= NF; j++)
                 printf("%s%c", a[i][j], j < NF ? ":" : "")
             print ""
         }
     }

The first field in each entry of the password file is the user’s login name, and the fields are separated by colons. Each record defines a subarray, with each field as an element in the subarray. Running the program produces the following output:

     $ gawk -v POS=1 -F: -f sort.awk /etc/passwd
     ⊣ adm:x:3:4:adm:/var/adm:/sbin/nologin
     ⊣ apache:x:48:48:Apache:/var/www:/sbin/nologin
     ⊣ avahi:x:70:70:Avahi daemon:/:/sbin/nologin
     ...

The comparison should normally always return the same value when given a specific pair of array elements as its arguments. If inconsistent results are returned, then the order is undefined. This behavior can be exploited to introduce random order into otherwise seemingly ordered data:

     function cmp_randomize(i1, v1, i2, v2)
     {
         # random order (caution: this may never terminate!)
         return (2 - 4 * rand())
     }

As already mentioned, the order of the indices is arbitrary if two elements compare equal. This is usually not a problem, but letting the tied elements come out in arbitrary order can be an issue, especially when comparing item values. The partial ordering of the equal elements may change the next time the array is traversed, if other elements are added to or removed from the array. One way to resolve ties when comparing elements with otherwise equal values is to include the indices in the comparison rules. Note that doing this may make the loop traversal less efficient, so consider it only if necessary. The following comparison functions force a deterministic order, and are based on the fact that the (string) indices of two elements are never equal:

     function cmp_numeric(i1, v1, i2, v2)
     {
         # numerical value (and index) comparison, descending order
         return (v1 != v2) ? (v2 - v1) : (i2 - i1)
     }
     function cmp_string(i1, v1, i2, v2)
     {
         # string value (and index) comparison, descending order
         v1 = v1 i1
         v2 = v2 i2
         return (v1 > v2) ? -1 : (v1 != v2)

     }

A custom comparison function can often simplify ordered loop traversal, and the sky is really the limit when it comes to designing such a function.

When string comparisons are made during a sort, either for element values where one or both aren’t numbers, or for element indices handled as strings, the value of IGNORECASE (see Section 7.5 [Predefined Variables], page 147) controls whether the comparisons treat corresponding upper- and lowercase letters as equivalent or distinct.

Another point to keep in mind is that in the case of subarrays, the element values can themselves be arrays; a production comparison function should use the isarray() function (see Section 9.1.7 [Getting Type Information], page 198) to check for this, and choose a defined sorting order for subarrays.

All sorting based on PROCINFO["sorted_in"] is disabled in POSIX mode, because the PROCINFO array is not special in that case.

As a side note, sorting the array indices before traversing the array has been reported to add a 15% to 20% overhead to the execution time of awk programs. For this reason, sorted array traversal is not the default.

12.2.2 Sorting Array Values and Indices with gawk

In most awk implementations, sorting an array requires writing a sort() function. This can be educational for exploring different sorting algorithms, but usually that’s not the point of the program. gawk provides the built-in asort() and asorti() functions (see Section 9.1.3 [String-Manipulation Functions], page 177) for sorting arrays. For example:

     populate the array data
     n = asort(data)
     for (i = 1; i <= n; i++)
         do something with data[i]

After the call to asort(), the array data is indexed from 1 to some number n, the total number of elements in data. (This count is asort()’s return value.) data[1] ≤ data[2] ≤ data[3], and so on. The default comparison is based on the type of the elements (see Section 6.3.2 [Variable Typing and Comparison Expressions], page 120). All numeric values come before all string values, which in turn come before all subarrays.

An important side effect of calling asort() is that the array’s original indices are irrevocably lost. As this isn’t always desirable, asort() accepts a second argument:

     populate the array source
     n = asort(source, dest)
     for (i = 1; i <= n; i++)
         do something with dest[i]

In this case, gawk copies the source array into the dest array and then sorts dest, destroying its indices. However, the source array is not affected.

Often, what’s needed is to sort on the values of the indices instead of the values of the elements. To do that, use the asorti() function. The interface and behavior are identical to that of asort(), except that the index values are used for sorting and become the values of the result array:

     { source[$0] = some_func($0) }
     END {
         n = asorti(source, dest)
         for (i = 1; i <= n; i++) {

Work with sorted indices directly:

             do something with dest[i]
             ...

Access original array via sorted indices:

             do something with source[dest[i]]
         }
     }

So far, so good. Now it starts to get interesting. Both asort() and asorti() accept a third string argument to control comparison of array elements. When we introduced asort() and asorti() in Section 9.1.3 [String-Manipulation Functions], page 177, we ignored this third argument; however, now is the time to describe how this argument affects these two functions.

Basically, the third argument specifies how the array is to be sorted. There are two possibilities. As with PROCINFO["sorted_in"], this argument may be one of the predefined names that gawk provides (see Section 8.1.6 [Using Predefined Array Scanning Orders with gawk], page 164), or it may be the name of a user-defined function (see Section 12.2.1 [Controlling Array Traversal], page 300).

In the latter case, the function can compare elements in any way it chooses, taking into account just the indices, just the values, or both. This is extremely powerful.

Once the array is sorted, asort() takes the values in their final order and uses them to fill in the result array, whereas asorti() takes the indices in their final order and uses them to fill in the result array.

NOTE: Copying array indices and elements isn’t expensive in terms of memory. Internally, gawk maintains reference counts to data. For example, when asort() copies the first array to the second one, there is only one copy of the original array elements’ data, even though both arrays use the values.

Because IGNORECASE affects string comparisons, the value of IGNORECASE also affects sorting for both asort() and asorti(). Note also that the locale’s sorting order does not come into play; comparisons are based on character values only.^{2}

12.3 Two-Way Communications with Another Process

It is often useful to be able to send data to a separate program for processing and then read the result. This can always be done with temporary files:

     # Write the data for processing
     tempfile = ("mydata." PROCINFO["pid"])
     while (not done with data)
         print data | ("subprogram > " tempfile)

2 This is true because locale-based comparison occurs only when in POSIX-compatibility mode, and be-

cause asort() and asorti() are gawk extensions, they are not available in that case.

     close("subprogram > " tempfile)
     # Read the results, remove tempfile when done
     while ((getline newdata < tempfile) > 0)
         process newdata appropriately
     close(tempfile)
     system("rm " tempfile)

This works, but not elegantly. Among other things, it requires that the program be run in a directory that cannot be shared among users; for example, /tmp will not do, as another user might happen to be using a temporary file with the same name.^{3}

However, with gawk, it is possible to open a two-way pipe to another process. The second process is termed a coprocess, as it runs in parallel with gawk. The two-way connection is created using the ‘|&’ operator (borrowed from the Korn shell, ksh):^{4}

     do {
         print data |& "subprogram"
         "subprogram" |& getline results
     } while (data left to process)
     close("subprogram")

The first time an I/O operation is executed using the ‘|&’ operator, gawk creates a twoway pipeline to a child process that runs the other program. Output created with print or printf is written to the program’s standard input, and output from the program’s standard output can be read by the gawk program using getline. As is the case with processes started by ‘|’, the subprogram can be any program, or pipeline of programs, that can be started by the shell.

There are some cautionary items to be aware of:

• As the code inside gawk currently stands, the coprocess’s standard error goes to the same place that the parent gawk’s standard error goes. It is not possible to read the child’s standard error separately.

• I/O buffering may be a problem. gawk automatically flushes all output down the pipe to the coprocess. However, if the coprocess does not flush its output, gawk may hang when doing a getline in order to read the coprocess’s results. This could lead to a situation known as deadlock, where each process is waiting for the other one to do something.

It is possible to close just one end of the two-way pipe to a coprocess, by supplying a second argument to the close() function of either "to" or "from" (see Section 5.9 [Closing Input and Output Redirections], page 102). These strings tell gawk to close the end of the pipe that sends data to the coprocess or the end that reads from it, respectively.

This is particularly necessary in order to use the system sort utility as part of a coprocess; sort must read all of its input data before it can produce any output. The sort program does not receive an end-of-file indication until gawk closes the write end of the pipe.

3 Michael Brennan suggests the use of rand() to generate unique file names. This is a valid point;

nevertheless, temporary files remain more difficult to use than two-way pipes.

4 This is very different from the same operator in the C shell and in Bash.

When you have finished writing data to the sort utility, you can close the "to" end of the pipe, and then start reading sorted data via getline. For example:

     BEGIN {
         command = "LC_ALL=C sort"
         n = split("abcdefghijklmnopqrstuvwxyz", a, "")
         for (i = n; i > 0; i--)
             print a[i] |& command
         close(command, "to")
         while ((command |& getline line) > 0)
             print "got", line
         close(command)
     }

This program writes the letters of the alphabet in reverse order, one per line, down the two-way pipe to sort. It then closes the write end of the pipe, so that sort receives an end-of-file indication. This causes sort to sort the data and write the sorted data back to the gawk program. Once all of the data has been read, gawk terminates the coprocess and exits.

As a side note, the assignment ‘LC_ALL=C’ in the sort command ensures traditional Unix (ASCII) sorting from sort. This is not strictly necessary here, but it’s good to know how to do this.

You may also use pseudo-ttys (ptys) for two-way communication instead of pipes, if your system supports them. This is done on a per-command basis, by setting a special element in the PROCINFO array (see Section 7.5.2 [Built-in Variables That Convey Information], page 149), like so:

     command = "sort -nr"           # command, save in convenience variable
     PROCINFO[command, "pty"] = 1   # update PROCINFO
     print ... |& command           # start two-way pipe
     ...

Using ptys usually avoids the buffer deadlock issues described earlier, at some loss in performance. If your system does not have ptys, or if all the system’s ptys are in use, gawk automatically falls back to using regular pipes.

12.4 Using gawk for Network Programming

     EMRED:

A host is a host from coast to coast, and nobody talks to a host that’s close, unless the host that isn’t close is busy, hung, or dead.

—Mike O’Brien (aka Mr. Protocol) In addition to being able to open a two-way pipeline to a coprocess on the same system

(see Section 12.3 [Two-Way Communications with Another Process], page 305), it is possible to make a two-way connection to another process on another system across an IP network connection.

You can think of this as just a very long two-way pipeline to a coprocess. The way gawk decides that you want to use TCP/IP networking is by recognizing special file names that begin with one of ‘/inet/’, ‘/inet4/’, or ‘/inet6/’.

   The full syntax of the special file name is /net-type/protocol/local-port/remote-

host/remote-port. The components are:

net-type Specifies the kind of Internet connection to make. Use ‘/inet4/’ to force IPv4, and ‘/inet6/’ to force IPv6. Plain ‘/inet/’ (which used to be the only option) uses the system default, most likely IPv4.

protocol The protocol to use over IP. This must be either ‘tcp’, or ‘udp’, for a TCP or UDP IP connection, respectively. TCP should be used for most applications.

local-port The local TCP or UDP port number to use. Use a port number of ‘0’ when you want the system to pick a port. This is what you should do when writing a TCP or UDP client. You may also use a well-known service name, such as ‘smtp’ or ‘http’, in which case gawk attempts to determine the predefined port number using the C getaddrinfo() function.

remote-host The IP address or fully qualified domain name of the Internet host to which you want to connect.

remote-port The TCP or UDP port number to use on the given remote-host. Again, use ‘0’ if you don’t care, or else a well-known service name.

NOTE: Failure in opening a two-way socket will result in a nonfatal error being returned to the calling code. The value of ERRNO indicates the error (see Section 7.5.2 [Built-in Variables That Convey Information], page 149).

Consider the following very simple example:

     BEGIN {
         Service = "/inet/tcp/0/localhost/daytime"
         Service |& getline
         print $0
         close(Service)
     }

This program reads the current date and time from the local system’s TCP daytime server. It then prints the results and closes the connection.

Because this topic is extensive, the use of gawk for TCP/IP programming is documented separately. See TCP/IP Internetworking with gawk (http://www.gnu.org/software/ gawk/manual/gawkinet/), which comes as part of the gawk distribution, for a much more complete introduction and discussion, as well as extensive examples.

12.5 Profiling Your awk Programs

You may produce execution traces of your awk programs. This is done by passing the option --profile to gawk. When gawk has finished running, it creates a profile of your program in a file named awkprof.out. Because it is profiling, it also executes up to 45% slower than gawk normally does.

As shown in the following example, the --profile option can be used to change the name of the file where gawk will write the profile:

gawk --profile=myprog.prof -f myprog.awk data1 data2

In the preceding example, gawk places the profile in myprog.prof instead of in awkprof.out.

Here is a sample session showing a simple awk program, its input data, and the results from running gawk with the --profile option. First, the awk program:

BEGIN { print "First BEGIN rule" }
END { print "First END rule" }
/foo/ {
    print "matched /foo/, gosh"
    for (i = 1; i <= 3; i++)
        sing()
}
{
    if (/foo/)
        print "if is true"
    else
        print "else is true"
}
BEGIN { print "Second BEGIN rule" }
END { print "Second END rule" }
function sing(    dummy)
{
    print "I gotta be me!"
}

Following is the input data:

foo
bar
baz
foo
junk

Here is the awkprof.out that results from running the gawk profiler on this program and data (this example also illustrates that awk programmers sometimes get up very early in the morning to work):

    # gawk profile, created Mon Sep 29 05:16:21 2014
    # BEGIN rule(s)
    BEGIN {

1          print "First BEGIN rule"
   }
   BEGIN {
1          print "Second BEGIN rule"
   }
   # Rule(s)
5  /foo/ { # 2
2          print "matched /foo/, gosh"
6          for (i = 1; i <= 3; i++) {
6                  sing()
           }
   }
5  {
5          if (/foo/) { # 2
2                  print "if is true"
3          } else {
3                  print "else is true"
           }
   }
   # END rule(s)
   END {
1          print "First END rule"
   }
   END {
1          print "Second END rule"
   }

   # Functions, listed alphabetically
6  function sing(dummy)
   {
6          print "I gotta be me!"
   }

This example illustrates many of the basic features of profiling output. They are as follows:

• The program is printed in the order BEGIN rules, BEGINFILE rules, pattern–action rules, ENDFILE rules, END rules, and functions, listed alphabetically. Multiple BEGIN and END rules retain their separate identities, as do multiple BEGINFILE and ENDFILE rules.

• Pattern–action rules have two counts. The first count, to the left of the rule, shows how many times the rule’s pattern was tested. The second count, to the right of the rule’s opening left brace in a comment, shows how many times the rule’s action was executed. The difference between the two indicates how many times the rule’s pattern evaluated to false.

• Similarly, the count for an if-else statement shows how many times the condition was tested. To the right of the opening left brace for the if’s body is a count showing how many times the condition was true. The count for the else indicates how many times the test failed.

• The count for a loop header (such as for or while) shows how many times the loop test was executed. (Because of this, you can’t just look at the count on the first statement in a rule to determine how many times the rule was executed. If the first statement is a loop, the count is misleading.)

• For user-defined functions, the count next to the function keyword indicates how many times the function was called. The counts next to the statements in the body show how many times those statements were executed.

• The layout uses “K&R” style with TABs. Braces are used everywhere, even when the body of an if, else, or loop is only a single statement.

• Parentheses are used only where needed, as indicated by the structure of the program and the precedence rules. For example, ‘(3 + 5) * 4’ means add three and five, then multiply the total by four. However, ‘3 + 5 * 4’ has no parentheses, and means ‘3 + (5

* 4)’.

• Parentheses are used around the arguments to print and printf only when the print or printf statement is followed by a redirection. Similarly, if the target of a redirection isn’t a scalar, it gets parenthesized.

• gawk supplies leading comments in front of the BEGIN and END rules, the BEGINFILE and ENDFILE rules, the pattern–action rules, and the functions.

The profiled version of your program may not look exactly like what you typed when you wrote it. This is because gawk creates the profiled version by “pretty-printing” its internal representation of the program. The advantage to this is that gawk can produce a standard representation. The disadvantage is that all source code comments are lost. Also, things such as:

 /foo/

come out as:

 /foo/   {
     print $0
 }

which is correct, but possibly unexpected.

Besides creating profiles when a program has completed, gawk can produce a profile while it is running. This is useful if your awk program goes into an infinite loop and you want to see what has been executed. To use this feature, run gawk with the --profile option in the background:

 $ gawk --profile -f myprog &

     [1] 13992

The shell prints a job number and process ID number; in this case, 13992. Use the kill command to send the USR1 signal to gawk:

     $ kill -USR1 13992

As usual, the profiled version of the program is written to awkprof.out, or to a different file if one was specified with the --profile option.

Along with the regular profile, as shown earlier, the profile file includes a trace of any active functions:

     # Function Call Stack:
     #   3. baz
     #   2. bar
     #   1. foo
     # -- main --

You may send gawk the USR1 signal as many times as you like. Each time, the profile and function call trace are appended to the output profile file.

If you use the HUP signal instead of the USR1 signal, gawk produces the profile and the function call trace and then exits.

When gawk runs on MS-Windows systems, it uses the INT and QUIT signals for producing the profile, and in the case of the INT signal, gawk exits. This is because these systems don’t support the kill command, so the only signals you can deliver to a program are those generated by the keyboard. The INT signal is generated by the Ctrl-c or Ctrl-BREAK key, while the QUIT signal is generated by the Ctrl-\ key.

Finally, gawk also accepts another option, --pretty-print. When called this way, gawk “pretty-prints” the program into awkprof.out, without any execution counts.

NOTE: The --pretty-print option still runs your program. This will change in the next major release.

12.6 Summary

• The --non-decimal-data option causes gawk to treat octal- and hexadecimal-looking input data as octal and hexadecimal. This option should be used with caution or not at all; use of strtonum() is preferable. Note that this option may disappear in a future version of gawk.

• You can take over complete control of sorting in ‘for (indx in array)’ array traversal by setting PROCINFO["sorted_in"] to the name of a user-defined function that does the comparison of array elements based on index and value.

• Similarly, you can supply the name of a user-defined comparison function as the third argument to either asort() or asorti() to control how those functions sort arrays. Or you may provide one of the predefined control strings that work for PROCINFO["sorted_

    in"].

• You can use the ‘|&’ operator to create a two-way pipe to a coprocess. You read from the coprocess with getline and write to it with print or printf. Use close() to close off the coprocess completely, or optionally, close off one side of the two-way communications.

• By using special file names with the ‘|&’ operator, you can open a TCP/IP (or UDP/IP) connection to remote hosts on the Internet. gawk supports both IPv4 and IPv6.

• You can generate statement count profiles of your program. This can help you determine which parts of your program may be taking the most time and let you tune them more easily. Sending the USR1 signal while profiling causes gawk to dump the profile and keep going, including a function call stack.

• You can also just “pretty-print” the program. This currently also runs the program, but that will change in the next major release.

Chapter 13: Internationalization with gawk 315

13 Internationalization with gawk

Once upon a time, computer makers wrote software that worked only in English. Eventually, hardware and software vendors noticed that if their systems worked in the native languages of non-English-speaking countries, they were able to sell more systems. As a result, internationalization and localization of programs and software systems became a common practice.

For many years, the ability to provide internationalization was largely restricted to programs written in C and C++. This chapter describes the underlying library gawk uses for internationalization, as well as how gawk makes internationalization features available at the awk program level. Having internationalization available at the awk level gives software developers additional flexibility—they are no longer forced to write in C or C++ when internationalization is a requirement.

13.1 Internationalization and Localization

Internationalization means writing (or modifying) a program once, in such a way that it can use multiple languages without requiring further source code changes. Localization means providing the data necessary for an internationalized program to work in a particular language. Most typically, these terms refer to features such as the language used for printing error messages, the language used to read responses, and information related to how numerical and monetary values are printed and read.

13.2 GNU gettext

gawk uses GNU gettext to provide its internationalization features. The facilities in GNU gettext focus on messages: strings printed by a program, either directly or via formatting

with printf or sprintf().^{1}

When using GNU gettext, each application has its own text domain. This is a unique name, such as ‘kpilot’ or ‘gawk’, that identifies the application. A complete application may have multiple components—programs written in C or C++, as well as scripts written in sh or awk. All of the components use the same text domain.

To make the discussion concrete, assume we’re writing an application named guide. Internationalization consists of the following steps, in this order:

1. The programmer reviews the source for all of guide’s components and marks each string that is a candidate for translation. For example, "`-F': option required" is a good candidate for translation. A table with strings of option names is not (e.g., gawk’s --profile option should remain the same, no matter what the local language).

2. The programmer indicates the application’s text domain ("guide") to the gettext library, by calling the textdomain() function.

3. Messages from the application are extracted from the source code and collected into a portable object template file (guide.pot), which lists the strings and their translations. The translations are initially empty. The original (usually English) messages serve as the key for lookup of the translations.

1 For some operating systems, the gawk port doesn’t support GNU gettext. Therefore, these features are

not available if you are using one of those operating systems. Sorry.

4. For each language with a translator, guide.pot is copied to a portable object file (.po) and translations are created and shipped with the application. For example, there might be a fr.po for a French translation.

5. Each language’s .po file is converted into a binary message object (.gmo) file. A message object file contains the original messages and their translations in a binary format that allows fast lookup of translations at runtime.

6. When guide is built and installed, the binary translation files are installed in a standard place.

7. For testing and development, it is possible to tell gettext to use .gmo files in a different directory than the standard one by using the bindtextdomain() function.

8. At runtime, guide looks up each string via a call to gettext(). The returned string is the translated string if available, or the original string if not.

9. If necessary, it is possible to access messages from a different text domain than the one belonging to the application, without having to switch the application’s default text domain back and forth.

In C (or C++), the string marking and dynamic translation lookup are accomplished by wrapping each string in a call to gettext():

     printf("%s", gettext("Don't Panic!\n"));

The tools that extract messages from source code pull out all strings enclosed in calls to

gettext().

The GNU gettext developers, recognizing that typing ‘gettext(...)’ over and over again is both painful and ugly to look at, use the macro ‘_’ (an underscore) to make things easier:

     /* In the standard header file: */
     #define _(str) gettext(str)
     /* In the program text: */
     printf("%s", _("Don't Panic!\n"));

This reduces the typing overhead to just three extra characters per string and is considerably easier to read as well.

There are locale categories for different types of locale-related information. The defined locale categories that gettext knows about are:

LC_MESSAGES

Text messages. This is the default category for gettext operations, but it is possible to supply a different one explicitly, if necessary. (It is almost never necessary to supply a different category.)

LC_COLLATE

Text-collation information (i.e., how different characters and/or groups of characters sort in a given language).

LC_CTYPE Character-type information (alphabetic, digit, upper- or lowercase, and so on) as well as character encoding. This information is accessed via the POSIX character classes in regular expressions, such as /[[:alnum:]]/ (see Section 3.4 [Using Bracket Expressions], page 50).

LC_MONETARY

Monetary information, such as the currency symbol, and whether the symbol goes before or after a number.

LC_NUMERIC

Numeric information, such as which characters to use for the decimal point and the thousands separator.^{2}

LC_TIME Time- and date-related information, such as 12- or 24-hour clock, month printed before or after the day in a date, local month abbreviations, and so on.

LC_ALL All of the above. (Not too useful in the context of gettext.)

13.3 Internationalizing awk Programs

gawk provides the following variables for internationalization:

TEXTDOMAIN

This variable indicates the application’s text domain. For compatibility with GNU gettext, the default value is "messages".

_"your message here"

String constants marked with a leading underscore are candidates for translation at runtime. String constants without a leading underscore are not translated.

gawk provides the following functions for internationalization:

dcgettext(string [, domain [, category]])

Return the translation of string in text domain domain for locale category category. The default value for domain is the current value of TEXTDOMAIN. The default value for category is "LC_MESSAGES". If you supply a value for category, it must be a string equal to one of the known locale categories described in the previous section. You must also supply a text domain. Use TEXTDOMAIN if you want to use the current domain.

CAUTION: The order of arguments to the awk version of the dcgettext() function is purposely different from the order for the C version. The awk version’s order was chosen to be simple and to allow for reasonable awk-style default arguments.

dcngettext(string1, string2, number [, domain [, category]])

Return the plural form used for number of the translation of string1 and string2 in text domain domain for locale category category. string1 is the English singular variant of a message, and string2 is the English plural variant of the same message. The default value for domain is the current value of TEXTDOMAIN. The default value for category is "LC_MESSAGES". The same remarks about argument order as for the dcgettext() function apply.

2 Americans use a comma every three decimal places and a period for the decimal point, while many

Europeans do exactly the opposite: 1,234.56 versus 1.234,56.

bindtextdomain(directory [, domain ])

Change the directory in which gettext looks for .gmo files, in case they will not or cannot be placed in the standard locations (e.g., during testing). Return the directory in which domain is “bound.” The default domain is the value of TEXTDOMAIN. If directory is the null string (""), then bindtextdomain() returns the current binding for the given domain.

To use these facilities in your awk program, follow these steps:

1. Set the variable TEXTDOMAIN to the text domain of your program. This is best done in a BEGIN rule (see Section 7.1.4 [The BEGIN and END Special Patterns], page 134), or it can also be done via the -v command-line option (see Section 2.2 [Command-Line Options], page 31):

BEGIN {
    TEXTDOMAIN = "guide"
    ...
}

2. Mark all translatable strings with a leading underscore (‘_’) character. It must be

adjacent to the opening quote of the string. For example:

print _"hello, world"
x = _"you goofed"
printf(_"Number of users is %d\n", nusers)

3. If you are creating strings dynamically, you can still translate them, using the

dcgettext() built-in function:^{3}

if (groggy)
    message = dcgettext("%d customers disturbing me\n", "adminprog")
else
    message = dcgettext("enjoying %d customers\n", "adminprog")
printf(message, ncustomers)

Here, the call to dcgettext() supplies a different text domain ("adminprog") in which to find the message, but it uses the default "LC_MESSAGES" category. The previous example only works if ncustomers is greater than one. This example would be better done with dcngettext():

if (groggy)
    message = dcngettext("%d customer disturbing me\n",
                         "%d customers disturbing me\n", "adminprog")
else
    message = dcngettext("enjoying %d customer\n",
                         "enjoying %d customers\n", "adminprog")
printf(message, ncustomers)

4. During development, you might want to put the .gmo file in a private directory for

testing. This is done with the bindtextdomain() built-in function:

BEGIN {
   TEXTDOMAIN = "guide"   # our text domain

3 Thanks to Bruno Haible for this example.

            if (Testing) {
                # where to find our files
                bindtextdomain("testdir")
                # joe is in charge of adminprog
                bindtextdomain("../joe/testdir", "adminprog")
            }
            ...
         }

See Section 13.5 [A Simple Internationalization Example], page 321, for an example program showing the steps to create and use translations from awk.

13.4 Translating awk Programs

Once a program’s translatable strings have been marked, they must be extracted to create the initial .pot file. As part of translation, it is often helpful to rearrange the order in which arguments to printf are output.

gawk’s --gen-pot command-line option extracts the messages and is discussed next. After that, printf’s ability to rearrange the order for printf arguments at runtime is covered.

13.4.1 Extracting Marked Strings

Once your awk program is working, and all the strings have been marked and you’ve set (and perhaps bound) the text domain, it is time to produce translations. First, use the --gen-pot command-line option to create the initial .pot file:

     gawk --gen-pot -f guide.awk > guide.pot

When run with --gen-pot, gawk does not execute your program. Instead, it parses it as usual and prints all marked strings to standard output in the format of a GNU gettext Portable Object file. Also included in the output are any constant strings that appear as the first argument to dcgettext() or as the first and second argument to dcngettext().^{4} You should distribute the generated .pot file with your awk program; translators will eventually use it to provide you translations that you can also then distribute. See Section 13.5 [A Simple Internationalization Example], page 321, for the full list of steps to go through to create and test translations for guide.

13.4.2 Rearranging printf Arguments

Format strings for printf and sprintf() (see Section 5.5 [Using printf Statements for Fancier Printing], page 92) present a special problem for translation. Consider the following:^{5}

     printf(_"String `%s' has %d characters\n",
               string, length(string)))

A possible German translation for this might be:

     "%d Zeichen lang ist die Zeichenkette `%s'\n"

4 The xgettext utility that comes with GNU gettext can handle .awk files.

5 This example is borrowed from the GNU gettext manual.

The problem should be obvious: the order of the format specifications is different from the original! Even though gettext() can return the translated string at runtime, it cannot change the argument order in the call to printf.

To solve this problem, printf format specifiers may have an additional optional element, which we call a positional specifier. For example:

"%2$d Zeichen lang ist die Zeichenkette `%1$s'\n"

Here, the positional specifier consists of an integer count, which indicates which argument to use, and a ‘$’. Counts are one-based, and the format string itself is not included. Thus, in the following example, ‘string’ is the first argument and ‘length(string)’ is the second:

$ gawk 'BEGIN {
>     string = "Don\47t Panic"
>     printf "%2$d characters live in \"%1$s\"\n",
>                         string, length(string)
> }'
⊣ 11 characters live in "Don't Panic"

If present, positional specifiers come first in the format specification, before the flags, the field width, and/or the precision.

Positional specifiers can be used with the dynamic field width and precision capability:

$ gawk 'BEGIN {
>    printf("%*.*s\n", 10, 20, "hello")
>    printf("%3$*2$.*1$s\n", 20, 10, "hello")
> }'
⊣  hello
⊣  hello

NOTE: When using ‘*’ with a positional specifier, the ‘*’ comes first, then the integer position, and then the ‘$’. This is somewhat counterintuitive.

gawk does not allow you to mix regular format specifiers and those with positional specifiers in the same string:

$ gawk 'BEGIN { printf "%d %3$s\n", 1, 2, "hi" }'
 error gawk: cmd. line:1: fatal: must use `count$' on all formats or none

NOTE: There are some pathological cases that gawk may fail to diagnose. In such cases, the output may not be what you expect. It’s still a bad idea to try mixing them, even if gawk doesn’t detect it.

Although positional specifiers can be used directly in awk programs, their primary purpose is to help in producing correct translations of format strings into languages different from the one in which the program is first written.

13.4.3 awk Portability Issues

gawk’s internationalization features were purposely chosen to have as little impact as possible on the portability of awk programs that use them to other versions of awk. Consider this program:

BEGIN {
    TEXTDOMAIN = "guide"
    if (Test_Guide)   # set with -v

         bindtextdomain("/test/guide/messages")
     print _"don't panic!"
 }

As written, it won’t work on other versions of awk. However, it is actually almost portable, requiring very little change:

• Assignments to TEXTDOMAIN won’t have any effect, because TEXTDOMAIN is not special in other awk implementations.

• Non-GNU versions of awk treat marked strings as the concatenation of a variable named _with the string following it.^{6} Typically, the variable _has the null string ("") as its value, leaving the original string constant as the result.

• By defining “dummy” functions to replace dcgettext(), dcngettext(), and bindtextdomain(), the awk program can be made to run, but all the messages are output in the original language. For example:

     function bindtextdomain(dir, domain)
     {
         return dir
     }
     function dcgettext(string, domain, category)
     {
         return string
     }
     function dcngettext(string1, string2, number, domain, category)
     {
         return (number == 1 ? string1 : string2)
     }

• The use of positional specifications in printf or sprintf() is not portable. To support gettext() at the C level, many systems’ C versions of sprintf() do support positional specifiers. But it works only if enough arguments are supplied in the function call. Many versions of awk pass printf formats and arguments unchanged to the underlying C library version of sprintf(), but only one format and argument at a time. What happens if a positional specification is used is anybody’s guess. However, because the positional specifications are primarily for use in translated format strings, and because non-GNU awks never retrieve the translated string, this should not be a problem in practice.

13.5 A Simple Internationalization Example

Now let’s look at a step-by-step example of how to internationalize and localize a simple awk program, using guide.awk as our original source:

 BEGIN {
     TEXTDOMAIN = "guide"
     bindtextdomain(".")  # for testing

6 This is good fodder for an “Obfuscated awk” contest.

    print _"Don't Panic"
    print _"The Answer Is", 42
    print "Pardon me, Zaphod who?"
}

Run ‘gawk --gen-pot’ to create the .pot file:

$ gawk --gen-pot -f guide.awk > guide.pot

This produces:

#: guide.awk:4
msgid "Don't Panic"
msgstr ""
#: guide.awk:5
msgid "The Answer Is"
msgstr ""

This original portable object template file is saved and reused for each language into which the application is translated. The msgid is the original string and the msgstr is the translation.

NOTE: Strings not marked with a leading underscore do not appear in the

guide.pot file.

Next, the messages must be translated. Here is a translation to a hypothetical dialect of English, called “Mellow”:^{7}

$ cp guide.pot guide-mellow.po
Add translations to guide-mellow.po ...

Following are the translations:

#: guide.awk:4
msgid "Don't Panic"
msgstr "Hey man, relax!"
#: guide.awk:5
msgid "The Answer Is"
msgstr "Like, the scoop is"

The next step is to make the directory to hold the binary message object file and then to create the guide.mo file. We pretend that our file is to be used in the en_US.UTF-8 locale, because we have to use a locale name known to the C gettext routines. The directory layout shown here is standard for GNU gettext on GNU/Linux systems. Other versions of gettext may use a different layout:

$ mkdir en_US.UTF-8 en_US.UTF-8/LC_MESSAGES

The msgfmt utility does the conversion from human-readable .po file to machine-readable .mo file. By default, msgfmt creates a file named messages. This file must be renamed and placed in the proper directory (using the -o option) so that gawk can find it:

7 Perhaps it would be better if it were called “Hippy.” Ah, well.

 $ msgfmt guide-mellow.po -o en_US.UTF-8/LC_MESSAGES/guide.mo

Finally, we run the program to test it:

 $ gawk -f guide.awk
  ⊣ Hey man, relax!
  ⊣ Like, the scoop is 42
  ⊣ Pardon me, Zaphod who?

If the three replacement functions for dcgettext(), dcngettext(), and bindtextdomain() (see Section 13.4.3 [awk Portability Issues], page 320) are in a file named libintl.awk, then we can run guide.awk unchanged as follows:

 $ gawk --posix -f guide.awk -f libintl.awk
  ⊣ Don't Panic
  ⊣ The Answer Is 42
  ⊣ Pardon me, Zaphod who?

13.6 gawk Can Speak Your Language

gawk itself has been internationalized using the GNU gettext package. (GNU gettext is described in complete detail in GNU gettext utilities (http://www.gnu.org/software/ gettext/manual/).) As of this writing, the latest version of GNU gettext is version 0.19.4

(ftp://ftp.gnu.org/gnu/gettext/gettext-0.19.4.tar.gz).

If a translation of gawk’s messages exists, then gawk produces usage messages, warnings, and fatal errors in the local language.

13.7 Summary

• Internationalization means writing a program such that it can use multiple languages without requiring source code changes. Localization means providing the data necessary for an internationalized program to work in a particular language.

• gawk uses GNU gettext to let you internationalize and localize awk programs. A program’s text domain identifies the program for grouping all messages and other data together.

• You mark a program’s strings for translation by preceding them with an underscore. Once that is done, the strings are extracted into a .pot file. This file is copied for each language into a .po file, and the .po files are compiled into .gmo files for use at runtime.

• You can use positional specifications with sprintf() and printf to rearrange the placement of argument values in formatted strings and output. This is useful for the translation of format control strings.

• The internationalization features have been designed so that they can be easily worked around in a standard awk.

• gawk itself has been internationalized and ships with a number of translations for its messages.

Chapter 14: Debugging awk Programs 325

14 Debugging awk Programs

It would be nice if computer programs worked perfectly the first time they were run, but in real life, this rarely happens for programs of any complexity. Thus, most programming languages have facilities available for “debugging” programs, and now awk is no exception.

The gawk debugger is purposely modeled after the GNU Debugger (GDB) (http://www. gnu.org/software/gdb/) command-line debugger. If you are familiar with GDB, learning how to use gawk for debugging your program is easy.

14.1 Introduction to the gawk Debugger

This section introduces debugging in general and begins the discussion of debugging in

gawk.

14.1.1 Debugging in General

(If you have used debuggers in other languages, you may want to skip ahead to Section 14.1.3 [awk Debugging], page 326.)

Of course, a debugging program cannot remove bugs for you, because it has no way of knowing what you or your users consider a “bug” versus a “feature.” (Sometimes, we humans have a hard time with this ourselves.) In that case, what can you expect from such a tool? The answer to that depends on the language being debugged, but in general, you can expect at least the following:

• The ability to watch a program execute its instructions one by one, giving you, the programmer, the opportunity to think about what is happening on a time scale of seconds, minutes, or hours, rather than the nanosecond time scale at which the code usually runs.

• The opportunity to not only passively observe the operation of your program, but to control it and try different paths of execution, without having to change your source files.

• The chance to see the values of data in the program at any point in execution, and also to change that data on the fly, to see how that affects what happens afterward. (This often includes the ability to look at internal data structures besides the variables you actually defined in your code.)

• The ability to obtain additional information about your program’s state or even its internal structure.

All of these tools provide a great amount of help in using your own skills and understanding of the goals of your program to find where it is going wrong (or, for that matter, to better comprehend a perfectly functional program that you or someone else wrote).

14.1.2 Debugging Concepts

Before diving in to the details, we need to introduce several important concepts that apply to just about all debuggers. The following list defines terms used throughout the rest of this chapter:

Stack frame Programs generally call functions during the course of their execution. One function can call another, or a function can call itself (recursion). You can view the chain of called functions (main program calls A, which calls B, which calls C), as a stack of executing functions: the currently running function is the topmost one on the stack, and when it finishes (returns), the next one down then becomes the active function. Such a stack is termed a call stack. For each function on the call stack, the system maintains a data area that contains the function’s parameters, local variables, and return value, as well as any other “bookkeeping” information needed to manage the call stack. This data area is termed a stack frame. gawk also follows this model, and gives you access to the call stack and to each stack frame. You can see the call stack, as well as from where each function on the stack was invoked. Commands that print the call stack print information about each stack frame (as detailed later on).

Breakpoint During debugging, you often wish to let the program run until it reaches a certain point, and then continue execution from there one statement (or instruction) at a time. The way to do this is to set a breakpoint within the program. A breakpoint is where the execution of the program should break off (stop), so that you can take over control of the program’s execution. You can add and remove as many breakpoints as you like.

Watchpoint A watchpoint is similar to a breakpoint. The difference is that breakpoints are oriented around the code: stop when a certain point in the code is reached. A watchpoint, however, specifies that program execution should stop when a data value is changed. This is useful, as sometimes it happens that a variable receives an erroneous value, and it’s hard to track down where this happens just by looking at the code. By using a watchpoint, you can stop whenever a variable is assigned to, and usually find the errant code quite quickly.

14.1.3 awk Debugging

Debugging an awk program has some specific aspects that are not shared with programs written in other languages.

First of all, the fact that awk programs usually take input line by line from a file or files and operate on those lines using specific rules makes it especially useful to organize viewing the execution of the program in terms of these rules. As we will see, each awk rule is treated almost like a function call, with its own specific block of instructions.

In addition, because awk is by design a very concise language, it is easy to lose sight of everything that is going on “inside” each line of awk code. The debugger provides the opportunity to look at the individual primitive instructions carried out by the higher-level awk commands.

14.2 Sample gawk Debugging Session

In order to illustrate the use of gawk as a debugger, let’s look at a sample debugging session. We will use the awk implementation of the POSIX uniq command described earlier (see Section 11.2.6 [Printing Nonduplicated Lines of Text], page 265) as our example.

14.2.1 How to Start the Debugger

Starting the debugger is almost exactly like running gawk normally, except you have to pass an additional option, --debug, or the corresponding short option, -D. The file(s) containing the program and any supporting code are given on the command line as arguments to one or more -f options. (gawk is not designed to debug command-line programs, only programs contained in files.) In our case, we invoke the debugger like this:

     $ gawk -D -f getopt.awk -f join.awk -f uniq.awk -1 inputfile

where both getopt.awk and uniq.awk are in $AWKPATH. (Experienced users of GDB or similar debuggers should note that this syntax is slightly different from what you are used to. With the gawk debugger, you give the arguments for running the program in the command line to the debugger rather than as part of the run command at the debugger prompt.) The -1 is an option to uniq.awk.

Instead of immediately running the program on inputfile, as gawk would ordinarily do, the debugger merely loads all the program source files, compiles them internally, and then gives us a prompt:

     gawk>

from which we can issue commands to the debugger. At this point, no code has been executed.

14.2.2 Finding the Bug

Let’s say that we are having a problem using (a faulty version of) uniq.awk in the “fieldskipping” mode, and it doesn’t seem to be catching lines which should be identical when skipping the first field, such as:

     awk is a wonderful program!
     gawk is a wonderful program!

This could happen if we were thinking (C-like) of the fields in a record as being numbered in a zero-based fashion, so instead of the lines:

     clast = join(alast, fcount+1, n)
     cline = join(aline, fcount+1, m)

we wrote:

     clast = join(alast, fcount, n)
     cline = join(aline, fcount, m)

The first thing we usually want to do when trying to investigate a problem like this is to put a breakpoint in the program so that we can watch it at work and catch what it is doing wrong. A reasonable spot for a breakpoint in uniq.awk is at the beginning of the function are_equal(), which compares the current line with the previous one. To set the breakpoint, use the b (breakpoint) command:

     gawk> b are_equal
     ⊣ Breakpoint 1 set at file `awklib/eg/prog/uniq.awk', line 63

The debugger tells us the file and line number where the breakpoint is. Now type ‘r’ or ‘run’ and the program runs until it hits the breakpoint for the first time:

     gawk> r
     ⊣ Starting program:

     ⊣ Stopping in Rule ...
     ⊣ Breakpoint 1, are_equal(n, m, clast, cline, alast, aline)
              at `awklib/eg/prog/uniq.awk':63
     ⊣ 63          if (fcount == 0 && charcount == 0)
     gawk>

Now we can look at what’s going on inside our program. First of all, let’s see how we got to where we are. At the prompt, we type ‘bt’ (short for “backtrace”), and the debugger responds with a listing of the current stack frames:

     gawk> bt
     ⊣ #0  are_equal(n, m, clast, cline, alast, aline)
              at `awklib/eg/prog/uniq.awk':68
     ⊣ #1  in main() at `awklib/eg/prog/uniq.awk':88

This tells us that are_equal() was called by the main program at line 88 of uniq.awk. (This is not a big surprise, because this is the only call to are_equal() in the program, but in more complex programs, knowing who called a function and with what parameters can be the key to finding the source of the problem.)

Now that we’re in are_equal(), we can start looking at the values of some variables. Let’s say we type ‘p n’ (p is short for “print”). We would expect to see the value of n, a parameter to are_equal(). Actually, the debugger gives us:

     gawk> p n
     ⊣ n = untyped variable

In this case, n is an uninitialized local variable, because the function was called without arguments (see Section 6.4 [Function Calls], page 126).

A more useful variable to display might be the current record:

     gawk> p $0
     ⊣ $0 = "gawk is a wonderful program!"

This might be a bit puzzling at first, as this is the second line of our test input. Let’s look at NR:

     gawk> p NR
     ⊣ NR = 2

So we can see that are_equal() was only called for the second record of the file. Of course, this is because our program contains a rule for ‘NR == 1’:

     NR == 1 {
         last = $0
         next
     }

OK, let’s just check that that rule worked correctly:

     gawk> p last
     ⊣ last = "awk is a wonderful program!"

Everything we have done so far has verified that the program has worked as planned, up to and including the call to are_equal(), so the problem must be inside this function. To investigate further, we must begin “stepping through” the lines of are_equal(). We start by typing ‘n’ (for “next”):

gawk> n
⊣ 66          if (fcount > 0) {

This tells us that gawk is now ready to execute line 66, which decides whether to give the lines the special “field-skipping” treatment indicated by the -1 command-line option. (Notice that we skipped from where we were before, at line 63, to here, because the condition

in line 63, ‘if (fcount == 0 && charcount == 0)’, was false.)

Continuing to step, we now get to the splitting of the current and last records:

gawk> n
⊣ 67              n = split(last, alast)
gawk> n
⊣ 68              m = split($0, aline)

At this point, we should be curious to see what our records were split into, so we try to look:

gawk> p n m alast aline
⊣ n = 5
⊣ m = untyped variable
⊣ alast = array, 5 elements
⊣ aline = untyped variable

(The p command can take more than one argument, similar to awk’s print statement.)

This is kind of disappointing, though. All we found out is that there are five elements in alast; m and aline don’t have values because we are at line 68 but haven’t executed it yet. This information is useful enough (we now know that none of the words were accidentally left out), but what if we want to see inside the array?

The first choice would be to use subscripts:

gawk> p alast[0]
⊣ "0" not in array `alast'

Oops!

gawk> p alast[1]
⊣ alast["1"] = "awk"

This would be kind of slow for a 100-member array, though, so gawk provides a shortcut (reminiscent of another language not to be mentioned):

gawk> p @alast
⊣ alast["1"] = "awk"
⊣ alast["2"] = "is"
⊣ alast["3"] = "a"
⊣ alast["4"] = "wonderful"
⊣ alast["5"] = "program!"

It looks like we got this far OK. Let’s take another step or two:

gawk> n
⊣ 69              clast = join(alast, fcount, n)
gawk> n
⊣ 70              cline = join(aline, fcount, m)

Well, here we are at our error (sorry to spoil the suspense). What we had in mind was to join the fields starting from the second one to make the virtual record to compare, and if the first field were numbered zero, this would work. Let’s look at what we’ve got:

     gawk> p cline clast
     ⊣ cline = "gawk is a wonderful program!"
     ⊣ clast = "awk is a wonderful program!"

Hey, those look pretty familiar! They’re just our original, unaltered input records. A little thinking (the human brain is still the best debugging tool), and we realize that we were off by one!

We get out of the debugger:

     gawk> q
     ⊣ The program is running. Exit anyway (y/n)? y

Then we get into an editor:

     clast = join(alast, fcount+1, n)
     cline = join(aline, fcount+1, m)

and problem solved!

14.3 Main Debugger Commands

The gawk debugger command set can be divided into the following categories: • Breakpoint control

• Execution control

• Viewing and changing data

• Working with the stack

• Getting information

• Miscellaneous

Each of these are discussed in the following subsections. In the following descriptions, commands that may be abbreviated show the abbreviation on a second description line. A debugger command name may also be truncated if that partial name is unambiguous. The debugger has the built-in capability to automatically repeat the previous command just by hitting Enter. This works for the commands list, next, nexti, step, stepi, and continue executed without any argument.

14.3.1 Control of Breakpoints

As we saw earlier, the first thing you probably want to do in a debugging session is to get your breakpoints set up, because your program will otherwise just run as if it was not under the debugger. The commands for controlling breakpoints are:

break [[filename:]n | function] ["expression"] b [[filename:]n | function] ["expression"]

Without any argument, set a breakpoint at the next instruction to be executed in the selected stack frame. Arguments can be one of the following:

n Set a breakpoint at line number n in the current source file. filename:n

Set a breakpoint at line number n in source file filename.

function Set a breakpoint at entry to (the first instruction of) function function.

Each breakpoint is assigned a number that can be used to delete it from the breakpoint list using the delete command. With a breakpoint, you may also supply a condition. This is an awk expression (enclosed in double quotes) that the debugger evaluates whenever the breakpoint is reached. If the condition is true, then the debugger stops execution and prompts for a command. Otherwise, it continues executing the program.

clear [[filename:]n | function] Without any argument, delete any breakpoint at the next instruction to be executed in the selected stack frame. If the program stops at a breakpoint, this deletes that breakpoint so that the program does not stop at that location again. Arguments can be one of the following:

n Delete breakpoint(s) set at line number n in the current source file.

filename:n Delete breakpoint(s) set at line number n in source file filename.

function Delete breakpoint(s) set at entry to function function.

condition n "expression"

Add a condition to existing breakpoint or watchpoint n. The condition is an awk expression enclosed in double quotes that the debugger evaluates whenever the breakpoint or watchpoint is reached. If the condition is true, then the debugger stops execution and prompts for a command. Otherwise, the debugger continues executing the program. If the condition expression is not specified, any existing condition is removed (i.e., the breakpoint or watchpoint is made unconditional).

delete [n1 n2 …] [n–m] d [n1 n2 …] [n–m]

Delete specified breakpoints or a range of breakpoints. Delete all defined breakpoints if no argument is supplied.

disable [n1 n2 … | n–m] Disable specified breakpoints or a range of breakpoints. Without any argument, disable all breakpoints.

enable [del | once] [n1 n2 …] [n–m] e [del | once] [n1 n2 …] [n–m]

Enable specified breakpoints or a range of breakpoints. Without any argument, enable all breakpoints. Optionally, you can specify how to enable the breakpoints:

del Enable the breakpoints temporarily, then delete each one when the program stops at it.

once Enable the breakpoints temporarily, then disable each one when the program stops at it.

ignore n count

Ignore breakpoint number n the next count times it is hit.

tbreak [[filename:]n | function] t [[filename:]n | function]

Set a temporary breakpoint (enabled for only one stop). The arguments are the same as for break.

14.3.2 Control of Execution

Now that your breakpoints are ready, you can start running the program and observing its behavior. There are more commands for controlling execution of the program than we saw in our earlier example:

commands [n]
silent

… end Set a list of commands to be executed upon stopping at a breakpoint or watch-

point. n is the breakpoint or watchpoint number. Without a number, the last one set is used. The actual commands follow, starting on the next line, and terminated by the end command. If the command silent is in the list, the usual messages about stopping at a breakpoint and the source line are not printed. Any command in the list that resumes execution (e.g., continue) terminates the list (an implicit end), and subsequent commands are ignored. For example:

               gawk> commands
               > silent
               > printf "A silent breakpoint; i = %d\n", i
               > info locals
               > set i = 10
               > continue
               > end
               gawk>
continue [count]

c [count] Resume program execution. If continued from a breakpoint and count is specified, ignore the breakpoint at that location the next count times before stopping.

finish Execute until the selected stack frame returns. Print the returned value.

next [count] n [count] Continue execution to the next source line, stepping over function calls. The

argument count controls how many times to repeat the action, as in step.

nexti [count] ni [count] Execute one (or count) instruction(s), stepping over function calls.

return [value] Cancel execution of a function call. If value (either a string or a number) is specified, it is used as the function’s return value. If used in a frame other than the innermost one (the currently executing function; i.e., frame number 0), discard all inner frames in addition to the selected one, and the caller of that frame becomes the innermost frame.

run

r Start/restart execution of the program. When restarting, the debugger retains the current breakpoints, watchpoints, command history, automatic display variables, and debugger options.

step [count] s [count] Continue execution until control reaches a different source line in the current

stack frame, stepping inside any function called within the line. If the argument count is supplied, steps that many times before stopping, unless it encounters a breakpoint or watchpoint.

stepi [count] si [count] Execute one (or count) instruction(s), stepping inside function calls. (For il-

lustration of what is meant by an “instruction” in gawk, see the output shown under dump in Section 14.3.6 [Miscellaneous Commands], page 337.)

until [[filename:]n | function] u [[filename:]n | function]

Without any argument, continue execution until a line past the current line in the current stack frame is reached. With an argument, continue execution until the specified location is reached, or the current stack frame returns.

14.3.3 Viewing and Changing Data

The commands for viewing and changing variables inside of gawk are:

display [var | $n]

Add variable var (or field $n) to the display list. The value of the variable or field is displayed each time the program stops. Each variable added to the list is identified by a unique number:

     gawk> display x
     ⊣ 10: x = 1

This displays the assigned item number, the variable name, and its current value. If the display variable refers to a function parameter, it is silently deleted from the list as soon as the execution reaches a context where no such variable of the given name exists. Without argument, display displays the current values of items on the list.

eval "awk statements"

Evaluate awk statements in the context of the running program. You can do anything that an awk program would do: assign values to variables, call functions, and so on.

eval param, … awk statements end This form of eval is similar, but it allows you to define “local variables” that

exist in the context of the awk statements, instead of using variables or function parameters defined by the program. print var1[, var2 …] p var1[, var2 …]

Print the value of a gawk variable or field. Fields must be referenced by constants:

     gawk> print $3

This prints the third field in the input record (if the specified field does not exist, it prints ‘Null field’). A variable can be an array element, with the subscripts being constant string values. To print the contents of an array, prefix the name of the array with the ‘@’ symbol:

     gawk> print @a

This prints the indices and the corresponding values for all elements in the array

a.

printf format [, arg …] Print formatted text. The format may include escape sequences, such as ‘\n’ (see Section 3.2 [Escape Sequences], page 46). No newline is printed unless one is specified.

set var=value Assign a constant (number or string) value to an awk variable or field. String values must be enclosed between double quotes ("…").

You can also set special awk variables, such as FS, NF, NR, and so on.

watch var | $n ["expression"]
w var | $n ["expression"]

Add variable var (or field $n) to the watch list. The debugger then stops whenever the value of the variable or field changes. Each watched item is assigned a number that can be used to delete it from the watch list using the unwatch command.

With a watchpoint, you may also supply a condition. This is an awk expression (enclosed in double quotes) that the debugger evaluates whenever the watchpoint is reached. If the condition is true, then the debugger stops execution and prompts for a command. Otherwise, gawk continues executing the program.

undisplay [n]

Remove item number n (or all items, if no argument) from the automatic display list.

unwatch [n]

Remove item number n (or all items, if no argument) from the watch list.

14.3.4 Working with the Stack

Whenever you run a program that contains any function calls, gawk maintains a stack of all of the function calls leading up to where the program is right now. You can see how you got to where you are, and also move around in the stack to see what the state of things was in the functions that called the one you are in. The commands for doing this are:

backtrace [count]

bt [count] where [count]

Print a backtrace of all function calls (stack frames), or innermost count frames if count > 0. Print the outermost count frames if count < 0. The backtrace displays the name and arguments to each function, the source file name, and the line number. The alias where for backtrace is provided for longtime GDB users who may be used to that command.

down [count] Move count (default 1) frames down the stack toward the innermost frame. Then select and print the frame.

frame [n]

f [n] Select and print stack frame n. Frame 0 is the currently executing, or innermost, frame (function call); frame 1 is the frame that called the innermost one. The highest-numbered frame is the one for the main program. The printed information consists of the frame number, function and argument names, source file, and the source line.

up [count] Move count (default 1) frames up the stack toward the outermost frame. Then select and print the frame.

14.3.5 Obtaining Information About the Program and the Debugger State

Besides looking at the values of variables, there is often a need to get other sorts of information about the state of your program and of the debugging environment itself. The gawk debugger has one command that provides this information, appropriately called info. info is used with one of a number of arguments that tell it exactly what you want to know:

info what

i what The value for what should be one of the following:

args List arguments of the selected frame.

break List all currently set breakpoints.

display List all items in the automatic display list.

frame Give a description of the selected stack frame.

functions

List all function definitions including source file names and line numbers.

locals List local variables of the selected frame.

source Print the name of the current source file. Each time the program stops, the current source file is the file containing the current instruction. When the debugger first starts, the current source file is the first file included via the -f option. The ‘list filename:lineno’ command can be used at any time to change the current source. sources List all program sources.

variables

List all global variables.

watch List all items in the watch list.

Additional commands give you control over the debugger, the ability to save the debugger’s state, and the ability to run debugger commands from a file. The commands are:

option [name[=value]] o [name[=value]]

Without an argument, display the available debugger options and their current values. ‘option name’ shows the current value of the named option. ‘option name=value’ assigns a new value to the named option. The available options are:

history_size

Set the maximum number of lines to keep in the history file ./.gawk_history. The default is 100.

listsize Specify the number of lines that list prints. The default is 15.

outfile Send gawk output to a file; debugger output still goes to standard output. An empty string ("") resets output to standard output.

prompt Change the debugger prompt. The default is ‘gawk> ’.

save_history [on | off]

Save command history to file ./.gawk_history. The default is on.

save_options [on | off]

Save current options to file ./.gawkrc upon exit. The default is on. Options are read back into the next session upon startup.

trace [on | off]

Turn instruction tracing on or off. The default is off.

save filename Save the commands from the current session to the given file name, so that they can be replayed using the source command.

source filename Run command(s) from a file; an error in any command does not terminate execution of subsequent commands. Comments (lines starting with ‘#’) are allowed in a command file. Empty lines are ignored; they do not repeat the last command. You can’t restart the program by having more than one run command in the file. Also, the list of commands may include additional source commands; however, the gawk debugger will not source the same file more than once in order to avoid infinite recursion. In addition to, or instead of, the source command, you can use the -D file or --debug=file command-line options to execute commands from a file noninteractively (see Section 2.2 [Command-Line Options], page 31).

14.3.6 Miscellaneous Commands

There are a few more commands that do not fit into the previous categories, as follows:

dump [filename] Dump byte code of the program to standard output or to the file named in filename. This prints a representation of the internal instructions that gawk executes to implement the awk commands in a program. This can be very enlightening, as the following partial dump of Davide Brini’s obfuscated code (see Section 11.3.11 [And Now for Something Completely Different], page 293) demonstrates:

gawk> dump
⊣  # BEGIN

⊣

⊣ [  1:0xfcd340] Op_rule           : [in_rule = BEGIN] [source_file = brini.awk]
⊣ [  1:0xfcc240] Op_push_i         : "~" [MALLOC|STRING|STRCUR]
⊣ [  1:0xfcc2a0] Op_push_i         : "~" [MALLOC|STRING|STRCUR]
⊣ [  1:0xfcc280] Op_match          :
⊣ [  1:0xfcc1e0] Op_store_var      : O
⊣ [  1:0xfcc2e0] Op_push_i         : "==" [MALLOC|STRING|STRCUR]
⊣ [  1:0xfcc340] Op_push_i         : "==" [MALLOC|STRING|STRCUR]
⊣ [  1:0xfcc320] Op_equal          :
⊣ [  1:0xfcc200] Op_store_var      : o
⊣ [  1:0xfcc380] Op_push           : o
⊣ [  1:0xfcc360] Op_plus_i         : 0 [MALLOC|NUMCUR|NUMBER]
⊣ [  1:0xfcc220] Op_push_lhs       : o [do_reference = true]
⊣ [  1:0xfcc300] Op_assign_plus    :
⊣ [   :0xfcc2c0] Op_pop            :
⊣ [  1:0xfcc400] Op_push           : O
⊣ [  1:0xfcc420] Op_push_i         : "" [MALLOC|STRING|STRCUR]
⊣ [   :0xfcc4a0] Op_no_op          :
⊣ [  1:0xfcc480] Op_push           : O
⊣ [   :0xfcc4c0] Op_concat         : [expr_count = 3] [concat_flag = 0]
⊣ [  1:0xfcc3c0] Op_store_var      : x
⊣ [  1:0xfcc440] Op_push_lhs       : X [do_reference = true]
⊣ [  1:0xfcc3a0] Op_postincrement  :
⊣ [  1:0xfcc4e0] Op_push           : x
⊣ [  1:0xfcc540] Op_push           : o
⊣ [  1:0xfcc500] Op_plus           :
⊣ [  1:0xfcc580] Op_push           : o
⊣ [  1:0xfcc560] Op_plus           :
⊣ [  1:0xfcc460] Op_leq            :
⊣ [   :0xfcc5c0] Op_jmp_false      : [target_jmp = 0xfcc5e0]
⊣ [  1:0xfcc600] Op_push_i         : "%c" [MALLOC|STRING|STRCUR]
⊣ [   :0xfcc660] Op_no_op          :
⊣ [  1:0xfcc520] Op_assign_concat  : c
⊣ [   :0xfcc620] Op_jmp            : [target_jmp = 0xfcc440]

⊣

...

⊣

⊣ [     2:0xfcc5a0] Op_K_printf         : [expr_count = 17] [redir_type = ""]
⊣ [      :0xfcc140] Op_no_op            :
⊣ [      :0xfcc1c0] Op_atexit           :
⊣ [      :0xfcc640] Op_stop             :
⊣ [      :0xfcc180] Op_no_op            :
⊣ [      :0xfcd150] Op_after_beginfile  :
⊣ [      :0xfcc160] Op_no_op            :

      ⊣ [      :0xfcc1a0] Op_after_endfile    :
      gawk>
help

h Print a list of all of the gawk debugger commands with a short summary of their usage. ‘help command’ prints the information about the command command.

list [- | + | n | filename:n | n–m | function] l [- | + | n | filename:n | n–m | function]

Print the specified lines (default 15) from the current source file or the file named filename. The possible arguments to list are as follows:

- (Minus) Print lines before the lines last printed.

+ Print lines after the lines last printed. list without any argument does the same thing.

n Print lines centered around line number n.

n–m Print lines from n to m.

filename:n Print lines centered around line number n in source file filename. This command may change the current source file.

function Print lines centered around the beginning of the function function. This command may change the current source file.

quit

q Exit the debugger. Debugging is great fun, but sometimes we all have to tend to other obligations in life, and sometimes we find the bug and are free to go on to the next one! As we saw earlier, if you are running a program, the debugger warns you when you type ‘q’ or ‘quit’, to make sure you really want to quit.

trace [on | off]

Turn on or off continuous printing of the instructions that are about to be executed, along with the awk lines they implement. The default is off. It is to be hoped that most of the “opcodes” in these instructions are fairly self-explanatory, and using stepi and nexti while trace is on will make them into familiar friends.

14.4 Readline Support

If gawk is compiled with the GNU Readline library (http://cnswww.cns.cwru.edu/ php/chet/readline/readline.html), you can take advantage of that library’s command completion and history expansion features. The following types of completion are available:

Command completion Command names.

Source file name completion Source file names. Relevant commands are break, clear, list, tbreak, and

until.

Argument completion Non-numeric arguments to a command. Relevant commands are enable and

      info.

Variable name completion Global variable names, and function arguments in the current context if the program is running. Relevant commands are display, print, set, and watch.

14.5 Limitations

We hope you find the gawk debugger useful and enjoyable to work with, but as with any program, especially in its early releases, it still has some limitations. A few that it’s worth being aware of are:

• At this point, the debugger does not give a detailed explanation of what you did wrong when you type in something it doesn’t like. Rather, it just responds ‘syntax error’. When you do figure out what your mistake was, though, you’ll feel like a real guru.

• If you perused the dump of opcodes in Section 14.3.6 [Miscellaneous Commands], page 337 (or if you are already familiar with gawk internals), you will realize that much of the internal manipulation of data in gawk, as in many interpreters, is done on a stack. Op_push, Op_pop, and the like are the “bread and butter” of most gawk code. Unfortunately, as of now, the gawk debugger does not allow you to examine the stack’s contents. That is, the intermediate results of expression evaluation are on the stack, but cannot be printed. Rather, only variables that are defined in the program can be printed. Of course, a workaround for this is to use more explicit variables at the debugging stage and then change back to obscure, perhaps more optimal code later.

• There is no way to look “inside” the process of compiling regular expressions to see if you got it right. As an awk programmer, you are expected to know the meaning of

/[^[:alnum:][:blank:]]/.

• The gawk debugger is designed to be used by running a program (with all its parameters) on the command line, as described in Section 14.2.1 [How to Start the Debugger], page 327. There is no way (as of now) to attach or “break into” a running program. This seems reasonable for a language that is used mainly for quickly executing, short programs.

• The gawk debugger only accepts source code supplied with the -f option.

14.6 Summary

• Programs rarely work correctly the first time. Finding bugs is called debugging, and a program that helps you find bugs is a debugger. gawk has a built-in debugger that works very similarly to the GNU Debugger, GDB.

• Debuggers let you step through your program one statement at a time, examine and change variable and array values, and do a number of other things that let you understand what your program is actually doing (as opposed to what it is supposed to do).

• Like most debuggers, the gawk debugger works in terms of stack frames, and lets you set both breakpoints (stop at a point in the code) and watchpoints (stop when a data value changes).

• The debugger command set is fairly complete, providing control over breakpoints, execution, viewing and changing data, working with the stack, getting information, and other tasks.

• If the GNU Readline library is available when gawk is compiled, it is used by the debugger to provide command-line history and editing.

Chapter 15: Arithmetic and Arbitrary-Precision Arithmetic with gawk 341

15 Arithmetic and Arbitrary-Precision Arithmetic

with gawk

This chapter introduces some basic concepts relating to how computers do arithmetic and defines some important terms. It then proceeds to describe floating-point arithmetic, which is what awk uses for all its computations, including a discussion of arbitrary-precision floating-point arithmetic, which is a feature available only in gawk. It continues on to present arbitrary-precision integers, and concludes with a description of some points where gawk and the POSIX standard are not quite in agreement.

NOTE: Most users of gawk can safely skip this chapter. But if you want to do scientific calculations with gawk, this is the place to be.

15.1 A General Description of Computer Arithmetic

Until now, we have worked with data as either numbers or strings. Ultimately, however, computers represent everything in terms of binary digits, or bits. A decimal digit can take on any of 10 values: zero through nine. A binary digit can take on any of two values, zero or one. Using binary, computers (and computer software) can represent and manipulate numerical and character data. In general, the more bits you can use to represent a particular thing, the greater the range of possible values it can take on.

Modern computers support at least two, and often more, ways to do arithmetic. Each kind of arithmetic uses a different representation (organization of the bits) for the numbers. The kinds of arithmetic that interest us are:

Decimal arithmetic This is the kind of arithmetic you learned in elementary school, using paper and pencil (and/or a calculator). In theory, numbers can have an arbitrary number of digits on either side (or both sides) of the decimal point, and the results of a computation are always exact. Some modern systems can do decimal arithmetic in hardware, but usually you need a special software library to provide access to these instructions. There are also libraries that do decimal arithmetic entirely in software. Despite the fact that some users expect gawk to be performing decimal arithmetic,^{1} it does not do so.

Integer arithmetic In school, integer values were referred to as “whole” numbers—that is, numbers without any fractional part, such as 1, 42, or −17. The advantage to integer numbers is that they represent values exactly. The disadvantage is that their range is limited. In computers, integer values come in two flavors: signed and unsigned. Signed values may be negative or positive, whereas unsigned values are always greater than or equal to zero. In computer systems, integer arithmetic is exact, but the possible range of values is limited. Integer arithmetic is generally faster than floating-point arithmetic.

1 We don’t know why they expect this, but they do.

Floating-point arithmetic Floating-point numbers represent what were called in school “real” numbers (i.e., those that have a fractional part, such as 3.1415927). The advantage to floating-point numbers is that they can represent a much larger range of values than can integers. The disadvantage is that there are numbers that they cannot represent exactly.

Modern systems support floating-point arithmetic in hardware, with a limited range of values. There are software libraries that allow the use of arbitraryprecision floating-point calculations.

POSIX awk uses double-precision floating-point numbers, which can hold more digits than single-precision floating-point numbers. gawk has facilities for performing arbitrary-precision floating-point arithmetic, which we describe in more detail shortly.

Computers work with integer and floating-point values of different ranges. Integer values are usually either 32 or 64 bits in size. Single-precision floating-point values occupy 32 bits, whereas double-precision floating-point values occupy 64 bits. Floating-point values are always signed. The possible ranges of values are shown in Table 15.1.

Numeric representation  Minimum value  Maximum value
32-bit signed integer  −2,147,483,648  2,147,483,647
32-bit unsigned integer  0  4,294,967,295
64-bit signed integer  −9,223,372,036,854,775,808 9,223,372,036,854,775,807
64-bit unsigned integer 0 18,446,744,073,709,551,615

Single-precision floating point (approximate)

1.175494^{−38} 3.402823^{38}

Double-precision floating point (approximate)

2.225074^{−308} 1.797693^{308}

Table 15.1: Value ranges for different numeric representations

15.2 Other Stuff to Know

The rest of this chapter uses a number of terms. Here are some informal definitions that should help you work your way through the material here:

Accuracy A floating-point calculation’s accuracy is how close it comes to the real (paper and pencil) value.

Error The difference between what the result of a computation “should be” and what it actually is. It is best to minimize error as much as possible.

Exponent The order of magnitude of a value; some number of bits in a floating-point value store the exponent.

Inf A special value representing infinity. Operations involving another number and infinity produce infinity.

NaN “Not a number.”^{2} A special value that results from attempting a calculation that has no answer as a real number. In such a case, programs can either receive a floating-point exception, or get NaN back as the result. The IEEE 754 standard recommends that systems return NaN. Some examples:

sqrt(-1) This makes sense in the range of complex numbers, but not in the range of real numbers, so the result is NaN.

log(-8) −8 is out of the domain of log(), so the result is NaN.

Normalized How the significand (see later in this list) is usually stored. The value is adjusted so that the first bit is one, and then that leading one is assumed instead of physically stored. This provides one extra bit of precision.

Precision The number of bits used to represent a floating-point number. The more bits, the more digits you can represent. Binary and decimal precisions are related approximately, according to the formula:

prec = 3.322 · dps

Here, prec denotes the binary precision (measured in bits) and dps (short for decimal places) is the decimal digits.

Rounding mode How numbers are rounded up or down when necessary. More details are provided later.

Significand A floating-point value consists of the significand multiplied by 10 to the power of the exponent. For example, in 1.2345e67, the significand is 1.2345.

Stability From the Wikipedia article on numerical stability (http://en.wikipedia. org/wiki/Numerical_stability): “Calculations that can be proven not to magnify approximation errors are called numerically stable.”

See the Wikipedia article on accuracy and precision (http://en.wikipedia.org/wiki/ Accuracy_and_precision) for more information on some of those terms.

On modern systems, floating-point hardware uses the representation and operations defined by the IEEE 754 standard. Three of the standard IEEE 754 types are 32-bit single precision, 64-bit double precision, and 128-bit quadruple precision. The standard also specifies extended precision formats to allow greater precisions and larger exponent ranges. (awk uses only the 64-bit double-precision format.)

Table 15.2 lists the precision and exponent field values for the basic IEEE 754 binary formats.

2 Thanks to Michael Brennan for this description, which we have paraphrased, and for the examples.

Name  Total bits  Precision  Minimum exponent  Maximum exponent
Single  32  24  −126  +127
Double  64  53  −1022  +1023
Quadruple  128  113  −16382  +16383

Table 15.2: Basic IEEE format values

NOTE: The precision numbers include the implied leading one that gives them one extra bit of significand.

15.3 Arbitrary-Precision Arithmetic Features in gawk

By default, gawk uses the double-precision floating-point values supplied by the hardware of the system it runs on. However, if it was compiled to do so, gawk uses the GNU

MPFR (http://www.mpfr.org) and GNU MP (http://gmplib.org) (GMP) libraries

for arbitrary-precision arithmetic on numbers. You can see if MPFR support is available like so:

     $ gawk --version
     ⊣ GNU Awk 4.1.2, API: 1.1 (GNU MPFR 3.1.0-p3, GNU MP 5.0.2)
     ⊣ Copyright (C) 1989, 1991-2015 Free Software Foundation.
     ...

(You may see different version numbers than what’s shown here. That’s OK; what’s important is to see that GNU MPFR and GNU MP are listed in the output.)

Additionally, there are a few elements available in the PROCINFO array to provide information about the MPFR and GMP libraries (see Section 7.5.2 [Built-in Variables That Convey Information], page 149).

The MPFR library provides precise control over precisions and rounding modes, and gives correctly rounded, reproducible, platform-independent results. With the -M commandline option, all floating-point arithmetic operators and numeric functions can yield results to any desired precision level supported by MPFR.

Two predefined variables, PREC and ROUNDMODE, provide control over the working precision and the rounding mode. The precision and the rounding mode are set globally for every operation to follow. See Section 15.4.4 [Setting the Precision], page 348, and Section 15.4.5 [Setting the Rounding Mode], page 349, for more information.

15.4 Floating-Point Arithmetic: Caveat Emptor!

Math class is tough!

—Teen Talk Barbie, July 1992

This section provides a high-level overview of the issues involved when doing lots of floating-point arithmetic.^{3} The discussion applies to both hardware and arbitrary-precision floating-point arithmetic.

CAUTION: The material here is purposely general. If you need to do serious computer arithmetic, you should do some research first, and not rely just on what we tell you.

15.4.1 Floating-Point Arithmetic Is Not Exact

Binary floating-point representations and arithmetic are inexact. Simple values like 0.1 cannot be precisely represented using binary floating-point numbers, and the limited precision of floating-point numbers means that slight changes in the order of operations or the precision of intermediate storage can change the result. To make matters worse, with arbitrary-precision floating-point arithmetic, you can set the precision before starting a computation, but then you cannot be sure of the number of significant decimal places in the final result.

15.4.1.1 Many Numbers Cannot Be Represented Exactly

So, before you start to write any code, you should think about what you really want and what’s really happening. Consider the two numbers in the following example:

     x = 0.875             # 1/2 + 1/4 + 1/8
     y = 0.425

Unlike the number in y, the number stored in x is exactly representable in binary because it can be written as a finite sum of one or more fractions whose denominators are all powers of two. When gawk reads a floating-point number from program source, it automatically rounds that number to whatever precision your machine supports. If you try to print the numeric content of a variable using an output format string of "%.17g", it may not produce the same number as you assigned to it:

     $ gawk 'BEGIN { x = 0.875; y = 0.425
     >               printf("%0.17g, %0.17g\n", x, y) }'
     ⊣ 0.875, 0.42499999999999999

Often the error is so small you do not even notice it, and if you do, you can always specify how much precision you would like in your output. Usually this is a format string like "%.15g", which, when used in the previous example, produces an output identical to the input.

15.4.1.2 Be Careful Comparing Values

Because the underlying representation can be a little bit off from the exact value, comparing floating-point values to see if they are exactly equal is generally a bad idea. Here is an example where it does not work like you would expect:

     $ gawk 'BEGIN { print (0.1 + 12.2 == 12.3) }'
     ⊣ 0

3 There is a very nice paper on floating-point arithmetic (http://www.validlab.com/goldberg/paper.

pdf) by David Goldberg, “What Every Computer Scientist Should Know About Floating-Point Arith-

metic,” ACM Computing Surveys 23, 1 (1991-03): 5-48. This is worth reading if you are interested in

the details, but it does require a background in computer science.

The general wisdom when comparing floating-point values is to see if they are within some small range of each other (called a delta, or tolerance). You have to decide how small a delta is important to you. Code to do this looks something like the following:

delta = 0.00001                 # for example
difference = abs(a) - abs(b)    # subtract the two values
if (difference < delta)
    # all ok
else
    # not ok

(We assume that you have a simple absolute value function named abs() defined elsewhere in your program.)

15.4.1.3 Errors Accumulate

The loss of accuracy during a single computation with floating-point numbers usually isn’t enough to worry about. However, if you compute a value that is the result of a sequence of floating-point operations, the error can accumulate and greatly affect the computation itself. Here is an attempt to compute the value of π using one of its many series representations:

BEGIN {
    x = 1.0 / sqrt(3.0)
    n = 6
    for (i = 1; i < 30; i++) {
        n = n * 2.0
        x = (sqrt(x * x + 1) - 1) / x
        printf("%.15f\n", n * x)
    }
}

When run, the early errors propagate through later computations, causing the loop to terminate prematurely after attempting to divide by zero:

$ gawk -f pi.awk
⊣ 3.215390309173475
⊣ 3.159659942097510
⊣ 3.146086215131467
⊣ 3.142714599645573
...
⊣ 3.224515243534819
⊣ 2.791117213058638
⊣ 0.000000000000000
 error gawk: pi.awk:6: fatal: division by zero attempted

Here is an additional example where the inaccuracies in internal representations yield an unexpected result:

$ gawk 'BEGIN {
>   for (d = 1.1; d <= 1.5; d += 0.1)    # loop five times (?)
>       i++
>   print i
> }'

     ⊣ 4

15.4.2 Getting the Accuracy You Need

Can arbitrary-precision arithmetic give exact results? There are no easy answers. The standard rules of algebra often do not apply when using floating-point arithmetic. Among other things, the distributive and associative laws do not hold completely, and order of operation may be important for your computation. Rounding error, cumulative precision loss, and underflow are often troublesome.

When gawk tests the expressions ‘0.1 + 12.2’ and ‘12.3’ for equality using the machine double-precision arithmetic, it decides that they are not equal! (See Section 15.4.1.2 [Be Careful Comparing Values], page 345.) You can get the result you want by increasing the precision; 56 bits in this case does the job:

     $ gawk -M -v PREC=56 'BEGIN { print (0.1 + 12.2 == 12.3) }'
     ⊣ 1

If adding more bits is good, perhaps adding even more bits of precision is better? Here is what happens if we use an even larger value of PREC:

     $ gawk -M -v PREC=201 'BEGIN { print (0.1 + 12.2 == 12.3) }'
     ⊣ 0

This is not a bug in gawk or in the MPFR library. It is easy to forget that the finite number of bits used to store the value is often just an approximation after proper rounding. The test for equality succeeds if and only if all bits in the two operands are exactly the same. Because this is not necessarily true after floating-point computations with a particular precision and effective rounding mode, a straight test for equality may not work. Instead, compare the two numbers to see if they are within the desirable delta of each other.

In applications where 15 or fewer decimal places suffice, hardware double-precision arithmetic can be adequate, and is usually much faster. But you need to keep in mind that every floating-point operation can suffer a new rounding error with catastrophic consequences, as illustrated by our earlier attempt to compute the value of π. Extra precision can greatly enhance the stability and the accuracy of your computation in such cases.

Additionally, you should understand that repeated addition is not necessarily equivalent to multiplication in floating-point arithmetic. In the example in Section 15.4.1.3 [Errors Accumulate], page 346:

     $ gawk 'BEGIN {
     >   for (d = 1.1; d <= 1.5; d += 0.1)    # loop five times (?)
     >       i++
     >   print i
     > }'
     ⊣ 4

you may or may not succeed in getting the correct result by choosing an arbitrarily large value for PREC. Reformulation of the problem at hand is often the correct approach in such situations.

15.4.3 Try a Few Extra Bits of Precision and Rounding

Instead of arbitrary-precision floating-point arithmetic, often all you need is an adjustment of your logic or a different order for the operations in your calculation. The stability and the accuracy of the computation of π in the earlier example can be enhanced by using the following simple algebraic transformation:

     (sqrt(x * x + 1) - 1) / x ≡ x / (sqrt(x * x + 1) + 1)

After making this change, the program converges to π in under 30 iterations:

     $ gawk -f pi2.awk
     ⊣ 3.215390309173473
     ⊣ 3.159659942097501
     ⊣ 3.146086215131436
     ⊣ 3.142714599645370
     ⊣ 3.141873049979825
     ...
     ⊣ 3.141592653589797
     ⊣ 3.141592653589797

15.4.4 Setting the Precision

gawk uses a global working precision; it does not keep track of the precision or accuracy of individual numbers. Performing an arithmetic operation or calling a built-in function rounds the result to the current working precision. The default working precision is 53 bits, which you can modify using the predefined variable PREC. You can also set the value to one of the predefined case-insensitive strings shown in Table 15.3, to emulate an IEEE 754 binary format.

PREC  IEEE 754 binary format
"half"  16-bit half-precision
"single" Basic 32-bit single precision
"double" Basic 64-bit double precision
"quad"  Basic 128-bit quadruple precision
"oct"  256-bit octuple precision

Table 15.3: Predefined precision strings for PREC

The following example illustrates the effects of changing precision on arithmetic operations:

     $ gawk -M -v PREC=100 'BEGIN { x = 1.0e-400; print x + 0
     >   PREC = "double"; print x + 0 }'
     ⊣ 1e-400
     ⊣ 0

CAUTION: Be wary of floating-point constants! When reading a floating-point constant from program source code, gawk uses the default precision (that of a C double), unless overridden by an assignment to the special variable PREC on the command line, to store it internally as an MPFR number. Changing the precision using PREC in the program text does not change the precision of a constant. If you need to represent a floating-point constant at a higher precision than the default and cannot use a command-line assignment to PREC, you should either specify the constant as a string, or as a rational number, whenever possible.

The following example illustrates the differences among various ways to print a floating-point constant:

          $ gawk -M 'BEGIN { PREC = 113; printf("%0.25f\n", 0.1) }'
          ⊣ 0.1000000000000000055511151
          $ gawk -M -v PREC=113 'BEGIN { printf("%0.25f\n", 0.1) }'
          ⊣ 0.1000000000000000000000000
          $ gawk -M 'BEGIN { PREC = 113; printf("%0.25f\n", "0.1") }'
          ⊣ 0.1000000000000000000000000
          $ gawk -M 'BEGIN { PREC = 113; printf("%0.25f\n", 1/10) }'
          ⊣ 0.1000000000000000000000000

15.4.5 Setting the Rounding Mode

The ROUNDMODE variable provides program-level control over the rounding mode. The correspondence between ROUNDMODE and the IEEE rounding modes is shown in Table 15.4.

Rounding mode  IEEE name  ROUNDMODE
Round to nearest, ties to even  roundTiesToEven        "N" or "n"
Round toward positive infinity  roundTowardPositive    "U" or "u"
Round toward negative infinity  roundTowardNegative    "D" or "d"

Round toward zero  roundTowardZero        "Z" or "z"

Round to nearest, ties away from zero  roundTiesToAway        "A" or "a"

Table 15.4: gawk rounding modes

ROUNDMODE has the default value "N", which selects the IEEE 754 rounding mode roundTiesToEven. In Table 15.4, the value "A" selects roundTiesToAway. This is only available if your version of the MPFR library supports it; otherwise, setting ROUNDMODE to "A" has no effect.

The default mode roundTiesToEven is the most preferred, but the least intuitive. This method does the obvious thing for most values, by rounding them up or down to the nearest digit. For example, rounding 1.132 to two digits yields 1.13, and rounding 1.157 yields 1.16.

However, when it comes to rounding a value that is exactly halfway between, things do not work the way you probably learned in school. In this case, the number is rounded to the nearest even digit. So rounding 0.125 to two digits rounds down to 0.12, but rounding 0.6875 to three digits rounds up to 0.688. You probably have already encountered this rounding mode when using printf to format floating-point numbers. For example:

     BEGIN {
         x = -4.5
         for (i = 1; i < 10; i++) {
             x += 1.0
             printf("%4.1f => %2.0f\n", x, x)
         }
     }

produces the following output when run on the author’s system:^{4}

4 It is possible for the output to be completely different if the C library in your system does not use the

IEEE 754 even-rounding rule to round halfway cases for printf.

     -3.5 => -4
     -2.5 => -2
     -1.5 => -2
     -0.5 => 0
      0.5 => 0
      1.5 => 2
      2.5 => 2
      3.5 => 4
      4.5 => 4

The theory behind roundTiesToEven is that it more or less evenly distributes upward and downward rounds of exact halves, which might cause any accumulating round-off error to cancel itself out. This is the default rounding mode for IEEE 754 computing functions and operators.

The other rounding modes are rarely used. Rounding toward positive infinity

(roundTowardPositive) and toward negative infinity (roundTowardNegative) are often

used to implement interval arithmetic, where you adjust the rounding mode to calculate upper and lower bounds for the range of output. The roundTowardZero mode can be used for converting floating-point numbers to integers. The rounding mode roundTiesToAway rounds the result to the nearest number and selects the number with the larger magnitude if a tie occurs.

Some numerical analysts will tell you that your choice of rounding style has tremendous impact on the final outcome, and advise you to wait until final output for any rounding. Instead, you can often avoid round-off error problems by setting the precision initially to some value sufficiently larger than the final desired precision, so that the accumulation of round-off error does not influence the outcome. If you suspect that results from your computation are sensitive to accumulation of round-off error, look for a significant difference in output when you change the rounding mode to be sure.

15.5 Arbitrary-Precision Integer Arithmetic with gawk

When given the -M option, gawk performs all integer arithmetic using GMP arbitraryprecision integers. Any number that looks like an integer in a source or data file is stored as an arbitrary-precision integer. The size of the integer is limited only by the available memory. For example, the following computes 5^{4^{3^{2}}} , the result of which is beyond the limits of ordinary hardware double-precision floating-point values:

     $ gawk -M 'BEGIN {
     >   x = 5^4^3^2
     >   print "number of digits =", length(x)
     >   print substr(x, 1, 20), "...", substr(x, length(x) - 19, 20)
     > }'
     ⊣ number of digits = 183231
     ⊣ 62060698786608744707 ... 92256259918212890625

If instead you were to compute the same value using arbitrary-precision floating-point values, the precision needed for correct output (using the formula prec = 3.322 · dps) would be 3.322 · 183231, or 608693.

The result from an arithmetic operation with an integer and a floating-point value is a floating-point value with a precision equal to the working precision. The following program calculates the eighth term in Sylvester’s sequence^{5} using a recurrence:

     $ gawk -M 'BEGIN {
     >   s = 2.0
     >   for (i = 1; i <= 7; i++)
     >       s = s * (s - 1) + 1
     >   print s
     > }'
     ⊣ 113423713055421845118910464

The output differs from the actual number, 113,423,713,055,421,844,361,000,443, because the default precision of 53 bits is not enough to represent the floating-point results exactly. You can either increase the precision (100 bits is enough in this case), or replace the floatingpoint constant ‘2.0’ with an integer, to perform all computations using integer arithmetic to get the correct output.

Sometimes gawk must implicitly convert an arbitrary-precision integer into an arbitraryprecision floating-point value. This is primarily because the MPFR library does not always provide the relevant interface to process arbitrary-precision integers or mixed-mode numbers as needed by an operation or function. In such a case, the precision is set to the minimum value necessary for exact conversion, and the working precision is not used for this purpose. If this is not what you need or want, you can employ a subterfuge and convert the integer to floating point first, like this:

     gawk -M 'BEGIN { n = 13; print (n + 0.0) % 2.0 }'

You can avoid this issue altogether by specifying the number as a floating-point value to begin with:

     gawk -M 'BEGIN { n = 13.0; print n % 2.0 }'

Note that for this particular example, it is likely best to just use the following:

     gawk -M 'BEGIN { n = 13; print n % 2 }'

15.6 Standards Versus Existing Practice

Historically, awk has converted any nonnumeric-looking string to the numeric value zero, when required. Furthermore, the original definition of the language and the original POSIX standards specified that awk only understands decimal numbers (base 10), and not octal (base 8) or hexadecimal numbers (base 16).

Changes in the language of the 2001 and 2004 POSIX standards can be interpreted to imply that awk should support additional features. These features are:

• Interpretation of floating-point data values specified in hexadecimal notation (e.g., 0xDEADBEEF). (Note: data values, not source code constants.)

• Support for the special IEEE 754 floating-point values “not a number” (NaN), positive infinity (“inf”), and negative infinity (“−inf”). In particular, the format for these values is as specified by the ISO 1999 C standard, which ignores case and can allow implementation-dependent additional characters after the ‘nan’ and allow either ‘inf’

5 Weisstein, Eric W. Sylvester’s Sequence. From MathWorld—A Wolfram Web Resource

(http://mathworld.wolfram.com/SylvestersSequence.html).

or ‘infinity’.

The first problem is that both of these are clear changes to historical practice:

• The gawk maintainer feels that supporting hexadecimal floating-point values, in particular, is ugly, and was never intended by the original designers to be part of the language.

• Allowing completely alphabetic strings to have valid numeric values is also a very severe departure from historical practice.

The second problem is that the gawk maintainer feels that this interpretation of the standard, which required a certain amount of “language lawyering” to arrive at in the first place, was not even intended by the standard developers. In other words, “We see how you got where you are, but we don’t think that that’s where you want to be.”

Recognizing these issues, but attempting to provide compatibility with the earlier versions of the standard, the 2008 POSIX standard added explicit wording to allow, but not require, that awk support hexadecimal floating-point values and special values for “not a number” and infinity.

Although the gawk maintainer continues to feel that providing those features is inadvisable, nevertheless, on systems that support IEEE floating point, it seems reasonable to provide some way to support NaN and infinity values. The solution implemented in gawk is as follows:

• With the --posix command-line option, gawk becomes “hands off.” String values are passed directly to the system library’s strtod() function, and if it successfully returns a numeric value, that is what’s used.^{6} By definition, the results are not portable across different systems. They are also a little surprising:

     $ echo nanny | gawk --posix '{ print $1 + 0 }'
     ⊣ nan
     $ echo 0xDeadBeef | gawk --posix '{ print $1 + 0 }'
     ⊣ 3735928559

• Without --posix, gawk interprets the four string values ‘+inf’, ‘-inf’, ‘+nan’, and ‘-nan’ specially, producing the corresponding special numeric values. The leading sign acts a signal to gawk (and the user) that the value is really numeric. Hexadecimal floating point is not supported (unless you also use --non-decimal-data, which is not recommended). For example:

     $ echo nanny | gawk '{ print $1 + 0 }'
     ⊣ 0
     $ echo +nan | gawk '{ print $1 + 0 }'
     ⊣ nan
     $ echo 0xDeadBeef | gawk '{ print $1 + 0 }'
     ⊣ 0

gawk ignores case in the four special values. Thus, ‘+nan’ and ‘+NaN’ are the same.

6 You asked for it, you got it.

15.7 Summary

• Most computer arithmetic is done using either integers or floating-point values. Standard awk uses double-precision floating-point values.

• In the early 1990s Barbie mistakenly said, “Math class is tough!” Although math isn’t tough, floating-point arithmetic isn’t the same as pencil-and-paper math, and care must be taken: − Not all numbers can be represented exactly. − Comparing values should use a delta, instead of being done directly with ‘==’ and

‘!=’. − Errors accumulate. − Operations are not always truly associative or distributive.

• Increasing the accuracy can help, but it is not a panacea.

• Often, increasing the accuracy and then rounding to the desired number of digits produces reasonable results.

• Use -M (or --bignum) to enable MPFR arithmetic. Use PREC to set the precision in bits, and ROUNDMODE to set the IEEE 754 rounding mode.

• With -M, gawk performs arbitrary-precision integer arithmetic using the GMP library. This is faster and more space-efficient than using MPFR for the same calculations.

• There are several areas with respect to floating-point numbers where gawk disagrees with the POSIX standard. It pays to be aware of them.

• Overall, there is no need to be unduly suspicious about the results from floating-point arithmetic. The lesson to remember is that floating-point arithmetic is always more complex than arithmetic using pencil and paper. In order to take advantage of the power of floating-point arithmetic, you need to know its limitations and work within them. For most casual use of floating-point arithmetic, you will often get the expected result if you simply round the display of your final results to the correct number of significant decimal digits.

• As general advice, avoid presenting numerical data in a manner that implies better precision than is actually the case.

Chapter 16: Writing Extensions for gawk 355

16 Writing Extensions for gawk

It is possible to add new functions written in C or C++ to gawk using dynamically loaded libraries. This facility is available on systems that support the C dlopen() and dlsym() functions. This chapter describes how to create extensions using code written in C or C++.

If you don’t know anything about C programming, you can safely skip this chapter, although you may wish to review the documentation on the extensions that come with gawk (see Section 16.7 [The Sample Extensions in the gawk Distribution], page 398), and the information on the gawkextlib project (see Section 16.8 [The gawkextlib Project], page 407). The sample extensions are automatically built and installed when gawk is.

NOTE: When --sandbox is specified, extensions are disabled (see Section 2.2 [Command-Line Options], page 31).

16.1 Introduction

An extension (sometimes called a plug-in) is a piece of external compiled code that gawk can load at runtime to provide additional functionality, over and above the built-in capabilities described in the rest of this book.

Extensions are useful because they allow you (of course) to extend gawk’s functionality. For example, they can provide access to system calls (such as chdir() to change directory) and to other C library routines that could be of use. As with most software, “the sky is the limit”; if you can imagine something that you might want to do and can write in C or C++, you can write an extension to do it!

Extensions are written in C or C++, using the application programming interface (API) defined for this purpose by the gawk developers. The rest of this chapter explains the facilities that the API provides and how to use them, and presents a small example extension. In addition, it documents the sample extensions included in the gawk distribution and describes the gawkextlib project. See Section C.5 [Extension API Design], page 453, for a discussion of the extension mechanism goals and design.

16.2 Extension Licensing

Every dynamic extension must be distributed under a license that is compatible with the GNU GPL (see [GNU General Public License], page 475).

In order for the extension to tell gawk that it is properly licensed, the extension must define the global symbol plugin_is_GPL_compatible. If this symbol does not exist, gawk emits a fatal error and exits when it tries to load your extension.

The declared type of the symbol should be int. It does not need to be in any allocated section, though. The code merely asserts that the symbol exists in the global scope. Something like this is enough:

     int plugin_is_GPL_compatible;

16.3 How It Works at a High Level

Communication between gawk and an extension is two-way. First, when an extension is loaded, gawk passes it a pointer to a struct whose fields are function pointers. This is shown in Figure 16.1.

Figure 16.1: Loading the extension

The extension can call functions inside gawk through these function pointers, at runtime, without needing (link-time) access to gawk’s symbols. One of these function pointers is to a function for “registering” new functions. This is shown in Figure 16.2.

Figure 16.2: Registering a new function

In the other direction, the extension registers its new functions with gawk by passing function pointers to the functions that provide the new feature (do_chdir(), for example). gawk associates the function pointer with a name and can then call it, using a defined calling convention. This is shown in Figure 16.3.

Figure 16.3: Calling the new function

The do_xxx() function, in turn, then uses the function pointers in the API struct to do its work, such as updating variables or arrays, printing messages, setting ERRNO, and so on.

Convenience macros make calling through the function pointers look like regular function calls so that extension code is quite readable and understandable.

Although all of this sounds somewhat complicated, the result is that extension code is quite straightforward to write and to read. You can see this in the sample extension filefuncs.c (see Section 16.6 [Example: Some File Functions], page 388) and also in the testext.c code for testing the APIs.

Some other bits and pieces:

• The API provides access to gawk’s do_xxx values, reflecting command-line options, like do_lint, do_profiling, and so on (see Section 16.4.12 [API Variables], page 385). These are informational: an extension cannot affect their values inside gawk. In addition, attempting to assign to them produces a compile-time error.

• The API also provides major and minor version numbers, so that an extension can check if the gawk it is loaded with supports the facilities it was compiled with. (Version mismatches “shouldn’t” happen, but we all know how that goes.) See Section 16.4.12.1 [API Version Constants and Variables], page 385, for details.

16.4 API Description

C or C++ code for an extension must include the header file gawkapi.h, which declares the functions and defines the data types used to communicate with gawk. This (rather large) section describes the API in detail.

16.4.1 Introduction

Access to facilities within gawk is achieved by calling through function pointers passed into your extension.

API function pointers are provided for the following kinds of operations:

• Allocating, reallocating, and releasing memory.

• Registration functions. You may register:

− Extension functions − Exit callbacks − A version string − Input parsers − Output wrappers − Two-way processors All of these are discussed in detail later in this chapter.

• Printing fatal, warning, and “lint” warning messages.

• Updating ERRNO, or unsetting it.

• Accessing parameters, including converting an undefined parameter into an array.

• Symbol table access: retrieving a global variable, creating one, or changing one.

• Creating and releasing cached values; this provides an efficient way to use values for multiple variables and can be a big performance win.

• Manipulating arrays:

− Retrieving, adding, deleting, and modifying elements − Getting the count of elements in an array − Creating a new array − Clearing an array − Flattening an array for easy C-style looping over all its indices and elements

Some points about using the API:

• The following types, macros, and/or functions are referenced in gawkapi.h. For correct use, you must therefore include the corresponding standard header file before including

gawkapi.h:

C entity  Header file

EOF                 <stdio.h>
Values for errno      <errno.h>
FILE                <stdio.h>
NULL                <stddef.h>
memcpy()            <string.h>
memset()            <string.h>
size_t              <sys/types.h>
struct stat         <sys/stat.h>

Due to portability concerns, especially to systems that are not fully standardscompliant, it is your responsibility to include the correct files in the correct way. This requirement is necessary in order to keep gawkapi.h clean, instead of becoming a portability hodge-podge as can be seen in some parts of the gawk source code.

• The gawkapi.h file may be included more than once without ill effect. Doing so, however, is poor coding practice.

• Although the API only uses ISO C 90 features, there is an exception; the “constructor” functions use the inline keyword. If your compiler does not support this keyword, you should either place ‘-Dinline=''’ on your command line or use the GNU Autotools and include a config.h file in your extensions.

• All pointers filled in by gawk point to memory managed by gawk and should be treated by the extension as read-only. Memory for all strings passed into gawk from the extension must come from calling one of gawk_malloc(), gawk_calloc(), or gawk_ realloc(), and is managed by gawk from then on.

• The API defines several simple structs that map values as seen from awk. A value can be a double, a string, or an array (as in multidimensional arrays, or when creating a new array). String values maintain both pointer and length, because embedded nul characters are allowed.

NOTE: By intent, strings are maintained using the current multibyte encoding (as defined by LC_xxx environment variables) and not using wide characters. This matches how gawk stores strings internally and also how characters are likely to be input into and output from files.

• When retrieving a value (such as a parameter or that of a global variable or array element), the extension requests a specific type (number, string, scalar, value cookie, array, or “undefined”). When the request is “undefined,” the returned value will have the real underlying type.

However, if the request and actual type don’t match, the access function returns “false” and fills in the type of the actual value that is there, so that the extension can, e.g., print an error message (such as “scalar passed where array expected”).

You may call the API functions by using the function pointers directly, but the interface is not so pretty. To make extension code look more like regular code, the gawkapi.h header file defines several macros that you should use in your code. This section presents the macros as if they were functions.

16.4.2 General-Purpose Data Types

I have a true love/hate relationship with unions.

—Arnold Robbins

That’s the thing about unions: the compiler will arrange things so they can accommodate both love and hate.

—Chet Ramey

The extension API defines a number of simple types and structures for general-purpose use. Additional, more specialized, data structures are introduced in subsequent sections, together with the functions that use them.

The general-purpose types and structures are as follows:

typedef void *awk_ext_id_t;

A value of this type is received from gawk when an extension is loaded. That value must then be passed back to gawk as the first parameter of each API function.

#define awk_const ...

This macro expands to ‘const’ when compiling an extension, and to nothing when compiling gawk itself. This makes certain fields in the API data structures unwritable from extension code, while allowing gawk to use them as it needs to.

typedef enum awk_bool {
awk_false = 0,
awk_true
} awk_bool_t;

A simple Boolean type.

typedef struct awk_string {
char *str;      /* data */
size_t len;     /* length thereof, in chars */
} awk_string_t;

This represents a mutable string. gawk owns the memory pointed to if it supplied the value. Otherwise, it takes ownership of the memory pointed to. Such memory must come from calling one of the gawk_malloc(), gawk_calloc(),

or gawk_realloc() functions!

As mentioned earlier, strings are maintained using the current multibyte encoding.

typedef enum {
AWK_UNDEFINED,
AWK_NUMBER,
AWK_STRING,
AWK_ARRAY,
AWK_SCALAR,         /* opaque access to a variable */
AWK_VALUE_COOKIE    /* for updating a previously created value */
} awk_valtype_t;

This enum indicates the type of a value. It is used in the following struct.

typedef struct awk_value {
awk_valtype_t val_type;
union {
awk_string_t       s;
double             d;
awk_array_t        a;
awk_scalar_t       scl;
awk_value_cookie_t vc;
} u;
} awk_value_t;

An “awk value.” The val_type member indicates what kind of value the union holds, and each member is of the appropriate type.

#define str_value      u.s
#define num_value      u.d
#define array_cookie   u.a
#define scalar_cookie  u.scl
#define value_cookie   u.vc

Using these macros makes accessing the fields of the awk_value_t more readable.

typedef void *awk_scalar_t;

Scalars can be represented as an opaque type. These values are obtained from gawk and then passed back into it. This is discussed in a general fashion in the text following this list, and in more detail in Section 16.4.10.2 [Variable Access and Update by Cookie], page 373.

typedef void *awk_value_cookie_t;

A “value cookie” is an opaque type representing a cached value. This is also discussed in a general fashion in the text following this list, and in more detail in Section 16.4.10.3 [Creating and Using Cached Values], page 375.

Scalar values in awk are either numbers or strings. The awk_value_t struct represents values. The val_type member indicates what is in the union.

Representing numbers is easy—the API uses a C double. Strings require more work. Because gawk allows embedded nul bytes in string values, a string must be represented as a pair containing a data pointer and length. This is the awk_string_t type.

Identifiers (i.e., the names of global variables) can be associated with either scalar values or with arrays. In addition, gawk provides true arrays of arrays, where any given array element can itself be an array. Discussion of arrays is delayed until Section 16.4.11 [Array Manipulation], page 377.

The various macros listed earlier make it easier to use the elements of the union as if they were fields in a struct; this is a common coding practice in C. Such code is easier to write and to read, but it remains your responsibility to make sure that the val_type member correctly reflects the type of the value in the awk_value_t struct.

Conceptually, the first three members of the union (number, string, and array) are all that is needed for working with awk values. However, because the API provides routines for accessing and changing the value of a global scalar variable only by using the variable’s name, there is a performance penalty: gawk must find the variable each time it is accessed and changed. This turns out to be a real issue, not just a theoretical one.

Thus, if you know that your extension will spend considerable time reading and/or changing the value of one or more scalar variables, you can obtain a scalar cookie^{1} object for that variable, and then use the cookie for getting the variable’s value or for changing the variable’s value. The awk_scalar_t type holds a scalar cookie, and the scalar_cookie macro provides access to the value of that type in the awk_value_t struct. Given a scalar cookie, gawk can directly retrieve or modify the value, as required, without having to find it first.

1 See the “cookie” entry in the Jargon file (http://catb.org/jargon/html/C/cookie.html) for a

definition of cookie, and the “magic cookie” entry in the Jargon file (http://catb.org/jargon/html/

M/magic-cookie.html) for a nice example. See also the entry for “Cookie” in the [Glossary], page 463.

The awk_value_cookie_t type and value_cookie macro are similar. If you know that you wish to use the same numeric or string value for one or more variables, you can create the value once, retaining a value cookie for it, and then pass in that value cookie whenever you wish to set the value of a variable. This saves storage space within the running gawk process and reduces the time needed to create the value.

16.4.3 Memory Allocation Functions and Convenience Macros

The API provides a number of memory allocation functions for allocating memory that can be passed to gawk, as well as a number of convenience macros. This subsection presents them all as function prototypes, in the way that extension code would use them:

void *gawk_malloc(size_t size);

Call the correct version of malloc() to allocate storage that may be passed to

          gawk.
void *gawk_calloc(size_t nmemb, size_t size);

Call the correct version of calloc() to allocate storage that may be passed to

          gawk.
void *gawk_realloc(void *ptr, size_t size);

Call the correct version of realloc() to allocate storage that may be passed

          to gawk.
void gawk_free(void *ptr);

Call the correct version of free() to release storage that was allocated with

          gawk_malloc(), gawk_calloc(), or gawk_realloc().

The API has to provide these functions because it is possible for an extension to be compiled and linked against a different version of the C library than was used for the gawk executable.^{2} If gawk were to use its version of free() when the memory came from an unrelated version of malloc(), unexpected behavior would likely result.

Two convenience macros may be used for allocating storage from gawk_malloc() and gawk_realloc(). If the allocation fails, they cause gawk to exit with a fatal error message. They should be used as if they were procedure calls that do not return a value:

#define emalloc(pointer, type, size, message) ...

The arguments to this macro are as follows:

pointer The pointer variable to point at the allocated storage.

type The type of the pointer variable. This is used to create a cast for

                    the call to gawk_malloc().

size The total number of bytes to be allocated.

message A message to be prefixed to the fatal error message. Typically this is the name of the function using the macro.

For example, you might allocate a string value like so:

2 This is more common on MS-Windows systems, but it can happen on Unix-like systems as well.

     awk_value_t result;
     char *message;
     const char greet[] = "Don't Panic!";
     emalloc(message, char *, sizeof(greet), "myfunc");
     strcpy(message, greet);
     make_malloced_string(message, strlen(message), & result);
#define erealloc(pointer, type, size, message) ...
This is like emalloc(), but it calls gawk_realloc() instead of gawk_malloc().

The arguments are the same as for the emalloc() macro.

16.4.4 Constructor Functions

The API provides a number of constructor functions for creating string and numeric values, as well as a number of convenience macros. This subsection presents them all as function prototypes, in the way that extension code would use them:

static inline awk_value_t *
make_const_string(const char *string, size_t length, awk_value_t *result);

This function creates a string value in the awk_value_t variable pointed to by result. It expects string to be a C string constant (or other string data), and automatically creates a copy of the data for storage in result. It returns

result.
static inline awk_value_t *
make_malloced_string(const char *string, size_t length, awk_value_t *result);

This function creates a string value in the awk_value_t variable pointed to by result. It expects string to be a ‘char *’ value pointing to data previously

obtained from gawk_malloc(), gawk_calloc(), or gawk_realloc(). The idea

here is that the data is passed directly to gawk, which assumes responsibility for it. It returns result.

static inline awk_value_t *
make_null_string(awk_value_t *result);

This specialized function creates a null string (the “undefined” value) in the awk_value_t variable pointed to by result. It returns result.

static inline awk_value_t *
make_number(double num, awk_value_t *result);

This function simply creates a numeric value in the awk_value_t variable pointed to by result.

16.4.5 Registration Functions

This section describes the API functions for registering parts of your extension with gawk.

16.4.5.1 Registering An Extension Function

Extension functions are described by the following record:

typedef struct awk_ext_func {
const char *name;

awk_value_t *(*function)(int num_actual_args, awk_value_t *result);
size_t num_expected_args;
} awk_ext_func_t;

The fields are:

const char *name;

The name of the new function. awk-level code calls the function by this name. This is a regular C string. Function names must obey the rules for awk identifiers. That is, they must begin with either an English letter or an underscore, which may be followed by any number of letters, digits, and underscores. Letter case in function names is significant.

awk_value_t *(*function)(int num_actual_args, awk_value_t *result);

This is a pointer to the C function that provides the extension’s functionality. The function must fill in *result with either a number or a string. gawk takes ownership of any string memory. As mentioned earlier, string memory must

come from one of gawk_malloc(), gawk_calloc(), or gawk_realloc().

The num_actual_args argument tells the C function how many actual parameters were passed from the calling awk code. The function must return the value of result. This is for the convenience of the calling code inside gawk.

size_t num_expected_args;

This is the number of arguments the function expects to receive. Each extension function may decide what to do if the number of arguments isn’t what it expected. As with real awk functions, it is likely OK to ignore extra arguments.

Once you have a record representing your extension function, you register it with gawk using this API function:

awk_bool_t add_ext_func(const char *namespace, const awk_ext_func_t *func);

This function returns true upon success, false otherwise. The namespace parameter is currently not used; you should pass in an empty string (""). The func pointer is the address of a struct representing your function, as just described.

16.4.5.2 Registering An Exit Callback Function

An exit callback function is a function that gawk calls before it exits. Such functions are useful if you have general “cleanup” tasks that should be performed in your extension (such as closing database connections or other resource deallocations). You can register such a function with gawk using the following function:

void awk_atexit(void (*funcp)(void *data, int exit_status),
      void *arg0);

The parameters are:

funcp A pointer to the function to be called before gawk exits. The data parameter will be the original value of arg0. The exit_status parameter is the exit status value that gawk intends to pass to the exit() system call. arg0 A pointer to private data that gawk saves in order to pass to the

function pointed to by funcp.

Exit callback functions are called in last-in, first-out (LIFO) order—that is, in the reverse order in which they are registered with gawk.

16.4.5.3 Registering An Extension Version String

You can register a version string that indicates the name and version of your extension with gawk, as follows:

void register_ext_version(const char *version);

Register the string pointed to by version with gawk. Note that gawk does not copy the version string, so it should not be changed.

gawk prints all registered extension version strings when it is invoked with the --version option.

16.4.5.4 Customized Input Parsers

By default, gawk reads text files as its input. It uses the value of RS to find the end of the record, and then uses FS (or FIELDWIDTHS or FPAT) to split it into fields (see Chapter 4 [Reading Input Files], page 59). Additionally, it sets the value of RT (see Section 7.5 [Predefined Variables], page 147).

If you want, you can provide your own custom input parser. An input parser’s job is to return a record to the gawk record-processing code, along with indicators for the value and length of the data to be used for RT, if any.

To provide an input parser, you must first provide two functions (where XXX is a prefix name for your extension):

awk_bool_t XXX_can_take_file(const awk_input_buf_t *iobuf);

This function examines the information available in iobuf (which we discuss shortly). Based on the information there, it decides if the input parser should be used for this file. If so, it should return true. Otherwise, it should return false. It should not change any state (variable values, etc.) within gawk.

awk_bool_t XXX_take_control_of(awk_input_buf_t *iobuf);

When gawk decides to hand control of the file over to the input parser, it calls this function. This function in turn must fill in certain fields in the awk_input_ buf_t structure and ensure that certain conditions are true. It should then return true. If an error of some kind occurs, it should not fill in any fields and should return false; then gawk will not use the input parser. The details are presented shortly.

Your extension should package these functions inside an awk_input_parser_t, which looks like this:

     typedef struct awk_input_parser {
         const char *name;   /* name of parser */
         awk_bool_t (*can_take_file)(const awk_input_buf_t *iobuf);
         awk_bool_t (*take_control_of)(awk_input_buf_t *iobuf);
         awk_const struct awk_input_parser *awk_const next;   /* for gawk */

     } awk_input_parser_t;

The fields are:

const char *name;

The name of the input parser. This is a regular C string.

awk_bool_t (*can_take_file)(const awk_input_buf_t *iobuf);

A pointer to your XXX_can_take_file() function.

awk_bool_t (*take_control_of)(awk_input_buf_t *iobuf);

A pointer to your XXX_take_control_of() function.

awk_const struct input_parser *awk_const next;

This is for use by gawk; therefore it is marked awk_const so that the extension cannot modify it.

The steps are as follows:

1. Create a static awk_input_parser_t variable and initialize it appropriately.

2. When your extension is loaded, register your input parser with gawk using the register_input_parser() API function (described next).

An awk_input_buf_t looks like this:

     typedef struct awk_input {
         const char *name;       /* filename */
         int fd;                 /* file descriptor */
     #define INVALID_HANDLE (-1)
         void *opaque;           /* private data for input parsers */
         int (*get_record)(char **out, struct awk_input *iobuf,
                           int *errcode, char **rt_start, size_t *rt_len);
         ssize_t (*read_func)();
         void (*close_func)(struct awk_input *iobuf);
         struct stat sbuf;       /* stat buf */
     } awk_input_buf_t;

The fields can be divided into two categories: those for use (initially, at least) by XXX_ can_take_file(), and those for use by XXX_take_control_of(). The first group of fields and their uses are as follows:

const char *name;

The name of the file.

int fd; A file descriptor for the file. If gawk was able to open the file, then fd will not be equal to INVALID_HANDLE. Otherwise, it will.

struct stat sbuf;

If the file descriptor is valid, then gawk will have filled in this structure via a call to the fstat() system call.

The XXX_can_take_file() function should examine these fields and decide if the input parser should be used for the file. The decision can be made based upon gawk state (the value of a variable defined previously by the extension and set by awk code), the name of the file, whether or not the file descriptor is valid, the information in the struct stat, or any combination of these factors.

Once XXX_can_take_file() has returned true, and gawk has decided to use your input parser, it calls XXX_take_control_of(). That function then fills either the get_record field or the read_func field in the awk_input_buf_t. It must also ensure that fd is not set to INVALID_HANDLE. The following list describes the fields that may be filled by XXX_take_

control_of():
void *opaque;

This is used to hold any state information needed by the input parser for this file. It is “opaque” to gawk. The input parser is not required to use this pointer.

int (*get_record)(char **out,
        struct awk_input *iobuf,
        int *errcode,
        char **rt_start,
        size_t *rt_len);

This function pointer should point to a function that creates the input records. Said function is the core of the input parser. Its behavior is described in the text following this list.

ssize_t (*read_func)();

This function pointer should point to a function that has the same behavior as the standard POSIX read() system call. It is an alternative to the get_record pointer. Its behavior is also described in the text following this list.

void (*close_func)(struct awk_input *iobuf);

This function pointer should point to a function that does the “teardown.” It should release any resources allocated by XXX_take_control_of(). It may also close the file. If it does so, it should set the fd field to INVALID_HANDLE.

If fd is still not INVALID_HANDLE after the call to this function, gawk calls the regular close() system call.

Having a “teardown” function is optional. If your input parser does not need it, do not set this field. Then, gawk calls the regular close() system call on the file descriptor, so it should be valid.

The XXX_get_record() function does the work of creating input records. The parameters are as follows:

char **out

This is a pointer to a char * variable that is set to point to the record. gawk makes its own copy of the data, so the extension must manage this storage.

struct awk_input *iobuf

This is the awk_input_buf_t for the file. The fields should be used for reading data (fd) and for managing private state (opaque), if any.

int *errcode

If an error occurs, *errcode should be set to an appropriate code from

<errno.h>.

char **rt_start
size_t *rt_len

If the concept of a “record terminator” makes sense, then *rt_start should be set to point to the data to be used for RT, and *rt_len should be set to the length of the data. Otherwise, *rt_len should be set to zero. gawk makes its own copy of this data, so the extension must manage this storage.

The return value is the length of the buffer pointed to by *out, or EOF if end-of-file was reached or an error occurred.

It is guaranteed that errcode is a valid pointer, so there is no need to test for a NULL value. gawk sets *errcode to zero, so there is no need to set it unless an error occurs.

If an error does occur, the function should return EOF and set *errcode to a value greater than zero. In that case, if *errcode does not equal zero, gawk automatically updates the ERRNO variable based on the value of *errcode. (In general, setting ‘*errcode = errno’ should do the right thing.)

As an alternative to supplying a function that returns an input record, you may instead supply a function that simply reads bytes, and let gawk parse the data into records. If you do so, the data should be returned in the multibyte encoding of the current locale. Such a function should follow the same behavior as the read() system call, and you fill in the read_func pointer with its address in the awk_input_buf_t structure.

By default, gawk sets the read_func pointer to point to the read() system call. So your extension need not set this field explicitly.

NOTE: You must choose one method or the other: either a function that returns a record, or one that returns raw data. In particular, if you supply a function to get a record, gawk will call it, and will never call the raw read function.

gawk ships with a sample extension that reads directories, returning records for each entry in a directory (see Section 16.7.6 [Reading Directories], page 404). You may wish to use that code as a guide for writing your own input parser.

When writing an input parser, you should think about (and document) how it is expected to interact with awk code. You may want it to always be called, and to take effect as appropriate (as the readdir extension does). Or you may want it to take effect based upon the value of an awk variable, as the XML extension from the gawkextlib project does (see Section 16.8 [The gawkextlib Project], page 407). In the latter case, code in a BEGINFILE section can look at FILENAME and ERRNO to decide whether or not to activate an input parser (see Section 7.1.5 [The BEGINFILE and ENDFILE Special Patterns], page 135).

You register your input parser with the following function:

void register_input_parser(awk_input_parser_t *input_parser);

Register the input parser pointed to by input_parser with gawk.

16.4.5.5 Customized Output Wrappers

An output wrapper is the mirror image of an input parser. It allows an extension to take over the output to a file opened with the ‘>’ or ‘>>’ I/O redirection operators (see Section 5.6 [Redirecting Output of print and printf], page 98).

The output wrapper is very similar to the input parser structure:

typedef struct awk_output_wrapper {
const char *name;   /* name of the wrapper */
awk_bool_t (*can_take_file)(const awk_output_buf_t *outbuf);
awk_bool_t (*take_control_of)(awk_output_buf_t *outbuf);
awk_const struct awk_output_wrapper *awk_const next;  /* for gawk */
} awk_output_wrapper_t;

The members are as follows:

const char *name;

This is the name of the output wrapper.

awk_bool_t (*can_take_file)(const awk_output_buf_t *outbuf);

This points to a function that examines the information in the awk_output_ buf_t structure pointed to by outbuf. It should return true if the output wrapper wants to take over the file, and false otherwise. It should not change any state (variable values, etc.) within gawk.

awk_bool_t (*take_control_of)(awk_output_buf_t *outbuf);

The function pointed to by this field is called when gawk decides to let the output wrapper take control of the file. It should fill in appropriate members of the awk_output_buf_t structure, as described next, and return true if successful, false otherwise.

awk_const struct output_wrapper *awk_const next;

This is for use by gawk; therefore it is marked awk_const so that the extension cannot modify it.

The awk_output_buf_t structure looks like this:

typedef struct awk_output_buf {
const char *name;   /* name of output file */
const char *mode;   /* mode argument to fopen */
FILE *fp;           /* stdio file pointer */
awk_bool_t redirected;  /* true if a wrapper is active */
void *opaque;       /* for use by output wrapper */
size_t (*gawk_fwrite)(const void *buf, size_t size, size_t count,
           FILE *fp, void *opaque);
int (*gawk_fflush)(FILE *fp, void *opaque);
int (*gawk_ferror)(FILE *fp, void *opaque);
int (*gawk_fclose)(FILE *fp, void *opaque);
} awk_output_buf_t;
Here too, your extension will define XXX_can_take_file() and XXX_take_control_

of() functions that examine and update data members in the awk_output_buf_t. The data members are as follows:

const char *name;

The name of the output file.

const char *mode;

The mode string (as would be used in the second argument to fopen()) with which the file was opened.

FILE *fp; The FILE pointer from <stdio.h>. gawk opens the file before attempting to find an output wrapper.

awk_bool_t redirected;

This field must be set to true by the XXX_take_control_of() function.

void *opaque;

This pointer is opaque to gawk. The extension should use it to store a pointer to any private data associated with the file.

size_t (*gawk_fwrite)(const void *buf, size_t size, size_t count,
                      FILE *fp, void *opaque);
int (*gawk_fflush)(FILE *fp, void *opaque);
int (*gawk_ferror)(FILE *fp, void *opaque);
int (*gawk_fclose)(FILE *fp, void *opaque);

These pointers should be set to point to functions that perform the equivalent function as the <stdio.h> functions do, if appropriate. gawk uses these function pointers for all output. gawk initializes the pointers to point to internal “pass-through” functions that just call the regular <stdio.h> functions, so an extension only needs to redefine those functions that are appropriate for what it does.

The XXX_can_take_file() function should make a decision based upon the name and mode fields, and any additional state (such as awk variable values) that is appropriate.

When gawk calls XXX_take_control_of(), that function should fill in the other fields as appropriate, except for fp, which it should just use normally.

You register your output wrapper with the following function:

void register_output_wrapper(awk_output_wrapper_t *output_wrapper);

Register the output wrapper pointed to by output_wrapper with gawk.

16.4.5.6 Customized Two-way Processors

A two-way processor combines an input parser and an output wrapper for two-way I/O with the ‘|&’ operator (see Section 5.6 [Redirecting Output of print and printf], page 98). It makes identical use of the awk_input_parser_t and awk_output_buf_t structures as described earlier.

A two-way processor is represented by the following structure:

     typedef struct awk_two_way_processor {
         const char *name;   /* name of the two-way processor */
         awk_bool_t (*can_take_two_way)(const char *name);
         awk_bool_t (*take_control_of)(const char *name,
                                       awk_input_buf_t *inbuf,
                                       awk_output_buf_t *outbuf);
         awk_const struct awk_two_way_processor *awk_const next;  /* for gawk */
     } awk_two_way_processor_t;

The fields are as follows:

const char *name;

The name of the two-way processor.

awk_bool_t (*can_take_two_way)(const char *name);

The function pointed to by this field should return true if it wants to take over two-way I/O for this file name. It should not change any state (variable values, etc.) within gawk.

awk_bool_t (*take_control_of)(const char *name,
                    awk_input_buf_t *inbuf,
                    awk_output_buf_t *outbuf);

The function pointed to by this field should fill in the awk_input_buf_t and awk_outut_buf_t structures pointed to by inbuf and outbuf, respectively. These structures were described earlier.

awk_const struct two_way_processor *awk_const next;

This is for use by gawk; therefore it is marked awk_const so that the extension cannot modify it.

As with the input parser and output processor, you provide “yes I can take this” and

“take over for this” functions, XXX_can_take_two_way() and XXX_take_control_of().

You register your two-way processor with the following function:

void register_two_way_processor(awk_two_way_processor_t *two_way_processor);

Register the two-way processor pointed to by two_way_processor with gawk.

16.4.6 Printing Messages

You can print different kinds of warning messages from your extension, as described here. Note that for these functions, you must pass in the extension ID received from gawk when the extension was loaded:^{3}

void fatal(awk_ext_id_t id, const char *format, ...);

Print a message and then cause gawk to exit immediately.

void warning(awk_ext_id_t id, const char *format, ...);

Print a warning message.

void lintwarn(awk_ext_id_t id, const char *format, ...);

Print a “lint warning.” Normally this is the same as printing a warning message, but if gawk was invoked with ‘--lint=fatal’, then lint warnings become fatal error messages.

All of these functions are otherwise like the C printf() family of functions, where the format parameter is a string with literal characters and formatting codes intermixed.

16.4.7 Updating ERRNO

The following functions allow you to update the ERRNO variable:

void update_ERRNO_int(int errno_val);

Set ERRNO to the string equivalent of the error code in errno_val. The value should be one of the defined error codes in <errno.h>, and gawk turns it into a (possibly translated) string using the C strerror() function.

3 Because the API uses only ISO C 90 features, it cannot make use of the ISO C 99 variadic macro feature

to hide that parameter. More’s the pity.

void update_ERRNO_string(const char *string);

Set ERRNO directly to the string value of ERRNO. gawk makes a copy of the value

          of string.
void unset_ERRNO(void);

Unset ERRNO.

16.4.8 Requesting Values

All of the functions that return values from gawk work in the same way. You pass in an awk_valtype_t value to indicate what kind of value you expect. If the actual value matches what you requested, the function returns true and fills in the awk_value_t result. Otherwise, the function returns false, and the val_type member indicates the type of the actual value. You may then print an error message or reissue the request for the actual value type, as appropriate. This behavior is summarized in Table 16.1.

Type of Actual Value

String  Number  Array  Undefined
String  String  String  False  False

Number Number if can be converted, else false Number False

False

Type  Array  False  False  Array  False
Requested  Scalar  Scalar  Scalar  False  False
Undefined  String  Number  Array  Undefined
Value cookie False  False  False  False

Table 16.1: API value types returned

16.4.9 Accessing and Updating Parameters

Two functions give you access to the arguments (parameters) passed to your extension function. They are:

awk_bool_t get_argument(size_t count,
                        awk_valtype_t wanted,
                        awk_value_t *result);

Fill in the awk_value_t structure pointed to by result with the countth argument. Return true if the actual type matches wanted, and false otherwise. In the latter case, result->val_type indicates the actual type (see Table 16.1). Counts are zero-based—the first argument is numbered zero, the second one, and so on. wanted indicates the type of value expected.

awk_bool_t set_argument(size_t count, awk_array_t array);

Convert a parameter that was undefined into an array; this provides call by reference for arrays. Return false if count is too big, or if the argument’s type is not undefined. See Section 16.4.11 [Array Manipulation], page 377, for more information on creating arrays.

16.4.10 Symbol Table Access

Two sets of routines provide access to global variables, and one set allows you to create and release cached values.

16.4.10.1 Variable Access and Update by Name

The following routines provide the ability to access and update global awk-level variables by name. In compiler terminology, identifiers of different kinds are termed symbols, thus the “sym” in the routines’ names. The data structure that stores information about symbols is termed a symbol table. The functions are as follows:

awk_bool_t sym_lookup(const char *name,
                      awk_valtype_t wanted,
                      awk_value_t *result);

Fill in the awk_value_t structure pointed to by result with the value of the variable named by the string name, which is a regular C string. wanted indicates the type of value expected. Return true if the actual type matches wanted, and false otherwise. In the latter case, result->val_type indicates the actual type (see Table 16.1).

awk_bool_t sym_update(const char *name, awk_value_t *value);

Update the variable named by the string name, which is a regular C string. The variable is added to gawk’s symbol table if it is not there. Return true if everything worked, and false otherwise.

Changing types (scalar to array or vice versa) of an existing variable is not allowed, nor may this routine be used to update an array. This routine cannot be used to update any of the predefined variables (such as ARGC or NF).

An extension can look up the value of gawk’s special variables. However, with the exception of the PROCINFO array, an extension cannot change any of those variables.

CAUTION: It is possible for the lookup of PROCINFO to fail. This happens if the awk program being run does not reference PROCINFO; in this case, gawk doesn’t bother to create the array and populate it.

16.4.10.2 Variable Access and Update by Cookie

A scalar cookie is an opaque handle that provides access to a global variable or array. It is an optimization that avoids looking up variables in gawk’s symbol table every time access is needed. This was discussed earlier, in Section 16.4.2 [General-Purpose Data Types], page 359.

The following functions let you work with scalar cookies:

awk_bool_t sym_lookup_scalar(awk_scalar_t cookie,
                             awk_valtype_t wanted,
                             awk_value_t *result);

Retrieve the current value of a scalar cookie. Once you have obtained a scalar cookie using sym_lookup(), you can use this function to get its value more efficiently. Return false if the value cannot be retrieved.

awk_bool_t sym_update_scalar(awk_scalar_t cookie, awk_value_t *value);

Update the value associated with a scalar cookie. Return false if the new value is not of type AWK_STRING or AWK_NUMBER. Here too, the predefined variables may not be updated.

It is not obvious at first glance how to work with scalar cookies or what their raison d’être really is. In theory, the sym_lookup() and sym_update() routines are all you really need to work with variables. For example, you might have code that looks up the value of a variable, evaluates a condition, and then possibly changes the value of the variable based on the result of that evaluation, like so:

     /*  do_magic --- do something really great */
     static awk_value_t *
     do_magic(int nargs, awk_value_t *result)
     {
         awk_value_t value;
         if (   sym_lookup("MAGIC_VAR", AWK_NUMBER, & value)
             && some_condition(value.num_value)) {
                 value.num_value += 42;
                 sym_update("MAGIC_VAR", & value);
         }
         return make_number(0.0, result);
     }

This code looks (and is) simple and straightforward. So what’s the problem?

Well, consider what happens if awk-level code associated with your extension calls the magic() function (implemented in C by do_magic()), once per record, while processing hundreds of thousands or millions of records. The MAGIC_VAR variable is looked up in the symbol table once or twice per function call!

The symbol table lookup is really pure overhead; it is considerably more efficient to get a cookie that represents the variable, and use that to get the variable’s value and update it as needed.^{4}

Thus, the way to use cookies is as follows. First, install your extension’s variable in gawk’s symbol table using sym_update(), as usual. Then get a scalar cookie for the variable using

sym_lookup():
     static awk_scalar_t magic_var_cookie;    /* cookie for MAGIC_VAR */
     static void
     my_extension_init()
     {
         awk_value_t value;
         /* install initial value */

4 The difference is measurable and quite real. Trust us.

sym_update("MAGIC_VAR", make_number(42.0, & value));
/* get the cookie */
sym_lookup("MAGIC_VAR", AWK_SCALAR, & value);
/* save the cookie */
magic_var_cookie = value.scalar_cookie;
...
}

Next, use the routines in this section for retrieving and updating the value through the cookie. Thus, do_magic() now becomes something like this:

/*  do_magic --- do something really great */
static awk_value_t *
do_magic(int nargs, awk_value_t *result)
{
awk_value_t value;
if (   sym_lookup_scalar(magic_var_cookie, AWK_NUMBER, & value)
    && some_condition(value.num_value)) {
        value.num_value += 42;
        sym_update_scalar(magic_var_cookie, & value);
}
...
return make_number(0.0, result);
}

NOTE: The previous code omitted error checking for presentation purposes. Your extension code should be more robust and carefully check the return values from the API functions.

16.4.10.3 Creating and Using Cached Values

The routines in this section allow you to create and release cached values. Like scalar cookies, in theory, cached values are not necessary. You can create numbers and strings using the functions in Section 16.4.4 [Constructor Functions], page 363. You can then assign those values to variables using sym_update() or sym_update_scalar(), as you like.

However, you can understand the point of cached values if you remember that every string value’s storage must come from gawk_malloc(), gawk_calloc(), or gawk_ realloc(). If you have 20 variables, all of which have the same string value, you must create 20 identical copies of the string.^{5}

It is clearly more efficient, if possible, to create a value once, and then tell gawk to reuse the value for multiple variables. That is what the routines in this section let you do. The functions are as follows:

5 Numeric values are clearly less problematic, requiring only a C double to store.

awk_bool_t create_value(awk_value_t *value, awk_value_cookie_t *result);

Create a cached string or numeric value from value for efficient later assignment. Only values of type AWK_NUMBER and AWK_STRING are allowed. Any other type is rejected. AWK_UNDEFINED could be allowed, but doing so would result in inferior performance.

awk_bool_t release_value(awk_value_cookie_t vc);

Release the memory associated with a value cookie obtained from create_

 value().

You use value cookies in a fashion similar to the way you use scalar cookies. In the extension initialization routine, you create the value cookie:

static awk_value_cookie_t answer_cookie;  /* static value cookie */
static void
my_extension_init()
{
awk_value_t value;
char *long_string;
size_t long_string_len;
/* code from earlier */
...
/* ... fill in long_string and long_string_len ... */
make_malloced_string(long_string, long_string_len, & value);
create_value(& value, & answer_cookie);    /* create cookie */
...
}

Once the value is created, you can use it as the value of any number of variables:

static awk_value_t *
do_magic(int nargs, awk_value_t *result)
{
awk_value_t new_value;
...    /* as earlier */
value.val_type = AWK_VALUE_COOKIE;
value.value_cookie = answer_cookie;
sym_update("VAR1", & value);
sym_update("VAR2", & value);
...
sym_update("VAR100", & value);
...
}

Using value cookies in this way saves considerable storage, as all of VAR1 through VAR100 share the same value.

You might be wondering, “Is this sharing problematic? What happens if awk code assigns a new value to VAR1; are all the others changed too?”

That’s a great question. The answer is that no, it’s not a problem. Internally, gawk uses reference-counted strings. This means that many variables can share the same string value, and gawk keeps track of the usage. When a variable’s value changes, gawk simply decrements the reference count on the old value and updates the variable to use the new value.

Finally, as part of your cleanup action (see Section 16.4.5.2 [Registering An Exit Callback Function], page 364) you should release any cached values that you created, using release_

value().

16.4.11 Array Manipulation

The primary data structure^{6} in awk is the associative array (see Chapter 8 [Arrays in awk], page 159). Extensions need to be able to manipulate awk arrays. The API provides a number of data structures for working with arrays, functions for working with individual elements, and functions for working with arrays as a whole. This includes the ability to “flatten” an array so that it is easy for C code to traverse every element in an array. The array data structures integrate nicely with the data structures for values to make it easy to both work with and create true arrays of arrays (see Section 16.4.2 [General-Purpose Data Types], page 359).

16.4.11.1 Array Data Types

The data types associated with arrays are as follows:

typedef void *awk_array_t;

If you request the value of an array variable, you get back an awk_array_t value. This value is opaque^{7} to the extension; it uniquely identifies the array but can only be used by passing it into API functions or receiving it from API functions. This is very similar to way ‘FILE *’ values are used with the <stdio.h> library routines.

typedef struct awk_element {
    /* convenience linked list pointer, not used by gawk */
    struct awk_element *next;
    enum {
        AWK_ELEMENT_DEFAULT = 0,  /* set by gawk */
        AWK_ELEMENT_DELETE = 1    /* set by extension */
    } flags;
    awk_value_t index;
    awk_value_t value;
} awk_element_t;

The awk_element_t is a “flattened” array element. awk produces an array of these inside the awk_flat_array_t (see the next item). Individual elements may be marked for deletion. New elements must be added individually, one at a time, using the separate API for that purpose. The fields are as follows:

6 OK, the only data structure.

7 It is also a “cookie,” but the gawk developers did not wish to overuse this term.

struct awk_element *next;

This pointer is for the convenience of extension writers. It allows an extension to create a linked list of new elements that can then be added to an array in a loop that traverses the list.

enum { ... } flags;

A set of flag values that convey information between the extension and gawk. Currently there is only one: AWK_ELEMENT_DELETE. Setting it causes gawk to delete the element from the original array upon release of the flattened array.

index

value The index and value of the element, respectively. All memory pointed to by index and value belongs to gawk.

typedef struct awk_flat_array {
awk_const void *awk_const opaque1;    /* for use by gawk */
awk_const void *awk_const opaque2;    /* for use by gawk */
awk_const size_t count;     /* how many elements */
awk_element_t elements[1];  /* will be extended */
} awk_flat_array_t;

This is a flattened array. When an extension gets one of these from gawk, the elements array is of actual size count. The opaque1 and opaque2 pointers are for use by gawk; therefore they are marked awk_const so that the extension cannot modify them.

16.4.11.2 Array Functions

The following functions relate to individual array elements:

awk_bool_t get_element_count(awk_array_t a_cookie, size_t *count);

For the array represented by a_cookie, place in *count the number of elements it contains. A subarray counts as a single element. Return false if there is an error.

awk_bool_t get_array_element(awk_array_t a_cookie,
                   const awk_value_t *const index,
                   awk_valtype_t wanted,
                   awk_value_t *result);

For the array represented by a_cookie, return in *result the value of the element whose index is index. wanted specifies the type of value you wish to retrieve. Return false if wanted does not match the actual type or if index is not in the array (see Table 16.1). The value for index can be numeric, in which case gawk converts it to a string. Using nonintegral values is possible, but requires that you understand how such values are converted to strings (see Section 6.1.4 [Conversion of Strings and Numbers], page 111); thus, using integral values is safest. As with all strings passed into gawk from an extension, the string value of

index must come from gawk_malloc(), gawk_calloc(), or gawk_realloc(),

and gawk releases the storage.

awk_bool_t set_array_element(awk_array_t a_cookie,
                   const awk_value_t *const index,
                   const awk_value_t *const value);

In the array represented by a_cookie, create or modify the element whose index is given by index. The ARGV and ENVIRON arrays may not be changed, although the PROCINFO array can be.

awk_bool_t set_array_element_by_elem(awk_array_t a_cookie,
                           awk_element_t element);
Like set_array_element(), but take the index and value from element. This

is a convenience macro.

awk_bool_t del_array_element(awk_array_t a_cookie,
                   const awk_value_t* const index);

Remove the element with the given index from the array represented by a_ cookie. Return true if the element was removed, or false if the element did not exist in the array.

The following functions relate to arrays as a whole:

awk_array_t create_array(void);

Create a new array to which elements may be added. See Section 16.4.11.4 [How To Create and Populate Arrays], page 383, for a discussion of how to create a new array and add elements to it.

awk_bool_t clear_array(awk_array_t a_cookie);

Clear the array represented by a_cookie. Return false if there was some kind of problem, true otherwise. The array remains an array, but after calling this function, it has no elements. This is equivalent to using the delete statement (see Section 8.4 [The delete Statement], page 168).

awk_bool_t flatten_array(awk_array_t a_cookie, awk_flat_array_t **data);

For the array represented by a_cookie, create an awk_flat_array_t structure and fill it in. Set the pointer whose address is passed as data to point to this structure. Return true upon success, or false otherwise. See Section 16.4.11.3 [Working With All The Elements of an Array], page 379, for a discussion of how to flatten an array and work with it.

awk_bool_t release_flattened_array(awk_array_t a_cookie,
                         awk_flat_array_t *data);

When done with a flattened array, release the storage using this function. You must pass in both the original array cookie and the address of the created awk_flat_array_t structure. The function returns true upon success, false otherwise.

16.4.11.3 Working With All The Elements of an Array

To flatten an array is to create a structure that represents the full array in a fashion that makes it easy for C code to traverse the entire array. Some of the code in extension/testext.c does this, and also serves as a nice example showing how to use the APIs.

We walk through that part of the code one step at a time. First, the gawk script that drives the test extension:

@load "testext"
BEGIN {
n = split("blacky rusty sophie raincloud lucky", pets)
printf("pets has %d elements\n", length(pets))
ret = dump_array_and_delete("pets", "3")
printf("dump_array_and_delete(pets) returned %d\n", ret)
if ("3" in pets)
    printf("dump_array_and_delete() did NOT remove index \"3\"!\n")
else
    printf("dump_array_and_delete() did remove index \"3\"!\n")
print ""
}

This code creates an array with split() (see Section 9.1.3 [String-Manipulation Functions], page 177) and then calls dump_array_and_delete(). That function looks up the array whose name is passed as the first argument, and deletes the element at the index passed in the second argument. The awk code then prints the return value and checks if the element was indeed deleted. Here is the C code that implements dump_array_and_delete(). It has been edited slightly for presentation.

The first part declares variables, sets up the default return value in result, and checks that the function was called with the correct number of arguments:

static awk_value_t *
dump_array_and_delete(int nargs, awk_value_t *result)
{
awk_value_t value, value2, value3;
awk_flat_array_t *flat_array;
size_t count;
char *name;
int i;
assert(result != NULL);
make_number(0.0, result);
if (nargs != 2) {
    printf("dump_array_and_delete: nargs not right "
           "(%d should be 2)\n", nargs);
    goto out;
}

The function then proceeds in steps, as follows. First, retrieve the name of the array, passed as the first argument, followed by the array itself. If either operation fails, print an error message and return:

/* get argument named array as flat array and print it */
if (get_argument(0, AWK_STRING, & value)) {
    name = value.str_value.str;

if (sym_lookup(name, AWK_ARRAY, & value2))
    printf("dump_array_and_delete: sym_lookup of %s passed\n",
           name);
else {
    printf("dump_array_and_delete: sym_lookup of %s failed\n",
           name);
    goto out;
}
} else {
printf("dump_array_and_delete: get_argument(0) failed\n");
goto out;
}

For testing purposes and to make sure that the C code sees the same number of elements as the awk code, the second step is to get the count of elements in the array and print it:

if (! get_element_count(value2.array_cookie, & count)) {
printf("dump_array_and_delete: get_element_count failed\n");
goto out;
}
printf("dump_array_and_delete: incoming size is %lu\n",
   (unsigned long) count);

The third step is to actually flatten the array, and then to double-check that the count in the awk_flat_array_t is the same as the count just retrieved:

if (! flatten_array(value2.array_cookie, & flat_array)) {
printf("dump_array_and_delete: could not flatten array\n");
goto out;
}
if (flat_array->count != count) {
printf("dump_array_and_delete: flat_array->count (%lu)"
       " != count (%lu)\n",
        (unsigned long) flat_array->count,
        (unsigned long) count);
goto out;
}

The fourth step is to retrieve the index of the element to be deleted, which was passed as the second argument. Remember that argument counts passed to get_argument() are zero-based, and thus the second argument is numbered one:

if (! get_argument(1, AWK_STRING, & value3)) {
printf("dump_array_and_delete: get_argument(1) failed\n");
goto out;
}

The fifth step is where the “real work” is done. The function loops over every element in the array, printing the index and element values. In addition, upon finding the element with the index that is supposed to be deleted, the function sets the AWK_ELEMENT_DELETE bit in the flags field of the element. When the array is released, gawk traverses the flattened array, and deletes any elements that have this flag bit set:

for (i = 0; i < flat_array->count; i++) {
printf("\t%s[\"%.*s\"] = %s\n",
    name,
    (int) flat_array->elements[i].index.str_value.len,
    flat_array->elements[i].index.str_value.str,
    valrep2str(& flat_array->elements[i].value));
if (strcmp(value3.str_value.str,
           flat_array->elements[i].index.str_value.str) == 0) {
    flat_array->elements[i].flags |= AWK_ELEMENT_DELETE;
    printf("dump_array_and_delete: marking element \"%s\" "
           "for deletion\n",
        flat_array->elements[i].index.str_value.str);
}
}

The sixth step is to release the flattened array. This tells gawk that the extension is no longer using the array, and that it should delete any elements marked for deletion. gawk also frees any storage that was allocated, so you should not use the pointer (flat_array in this code) once you have called release_flattened_array():

if (! release_flattened_array(value2.array_cookie, flat_array)) {
printf("dump_array_and_delete: could not release flattened array\n");
goto out;
}

Finally, because everything was successful, the function sets the return value to success, and returns:

make_number(1.0, result);
out:
return result;
}

Here is the output from running this part of the test:

pets has 5 elements
dump_array_and_delete: sym_lookup of pets passed
dump_array_and_delete: incoming size is 5
pets["1"] = "blacky"
pets["2"] = "rusty"
pets["3"] = "sophie"
dump_array_and_delete: marking element "3" for deletion
pets["4"] = "raincloud"
pets["5"] = "lucky"
dump_array_and_delete(pets) returned 1
dump_array_and_delete() did remove index "3"!

16.4.11.4 How To Create and Populate Arrays

Besides working with arrays created by awk code, you can create arrays and populate them as you see fit, and then awk code can access them and manipulate them.

There are two important points about creating arrays from extension code:

• You must install a new array into gawk’s symbol table immediately upon creating it. Once you have done so, you can then populate the array. Similarly, if installing a new array as a subarray of an existing array, you must add the new array to its parent before adding any elements to it. Thus, the correct way to build an array is to work “top down.” Create the array, and immediately install it in gawk’s symbol table using sym_update(), or install it as an element in a previously existing array using set_array_element(). We show example code shortly.

• Due to gawk internals, after using sym_update() to install an array into gawk, you have to retrieve the array cookie from the value passed in to sym_update() before doing anything else with it, like so:

     awk_value_t value;
     awk_array_t new_array;
     new_array = create_array();
     val.val_type = AWK_ARRAY;
     val.array_cookie = new_array;
     /* install array in the symbol table */
     sym_update("array", & val);
     new_array = val.array_cookie;    /* YOU MUST DO THIS */

If installing an array as a subarray, you must also retrieve the value of the array cookie after the call to set_element().

The following C code is a simple test extension to create an array with two regular elements and with a subarray. The leading #include directives and boilerplate variable declarations (see Section 16.4.13 [Boilerplate Code], page 386) are omitted for brevity. The first step is to create a new array and then install it in the symbol table:

 /* create_new_array --- create a named array */
 static void
 create_new_array()
 {
     awk_array_t a_cookie;
     awk_array_t subarray;
     awk_value_t index, value;
     a_cookie = create_array();
     value.val_type = AWK_ARRAY;
     value.array_cookie = a_cookie;

if (! sym_update("new_array", & value))
    printf("create_new_array: sym_update(\"new_array\") failed!\n");
a_cookie = value.array_cookie;

Note how a_cookie is reset from the array_cookie field in the value structure.

The second step is to install two regular values into new_array:

(void) make_const_string("hello", 5, & index);
(void) make_const_string("world", 5, & value);
if (! set_array_element(a_cookie, & index, & value)) {
    printf("fill_in_array: set_array_element failed\n");
    return;
}
(void) make_const_string("answer", 6, & index);
(void) make_number(42.0, & value);
if (! set_array_element(a_cookie, & index, & value)) {
    printf("fill_in_array: set_array_element failed\n");
    return;
}

The third step is to create the subarray and install it:

(void) make_const_string("subarray", 8, & index);
subarray = create_array();
value.val_type = AWK_ARRAY;
value.array_cookie = subarray;
if (! set_array_element(a_cookie, & index, & value)) {
    printf("fill_in_array: set_array_element failed\n");
    return;
}
subarray = value.array_cookie;

The final step is to populate the subarray with its own element:

(void) make_const_string("foo", 3, & index);
(void) make_const_string("bar", 3, & value);
if (! set_array_element(subarray, & index, & value)) {
    printf("fill_in_array: set_array_element failed\n");
    return;
}
}

Here is a sample script that loads the extension and then dumps the array:

@load "subarray"
function dumparray(name, array,     i)
{
for (i in array)
    if (isarray(array[i]))
        dumparray(name "[\"" i "\"]", array[i])

             else
                 printf("%s[\"%s\"] = %s\n", name, i, array[i])
     }
     BEGIN {
         dumparray("new_array", new_array);
     }

Here is the result of running the script:

     $ AWKLIBPATH=$PWD ./gawk -f subarray.awk
     ⊣ new_array["subarray"]["foo"] = bar
     ⊣ new_array["hello"] = world
     ⊣ new_array["answer"] = 42

(See Section 16.5 [How gawk Finds Extensions], page 388, for more information on the AWKLIBPATH environment variable.)

16.4.12 API Variables

The API provides two sets of variables. The first provides information about the version of the API (both with which the extension was compiled, and with which gawk was compiled). The second provides information about how gawk was invoked.

16.4.12.1 API Version Constants and Variables

The API provides both a “major” and a “minor” version number. The API versions are available at compile time as constants:

GAWK_API_MAJOR_VERSION

The major version of the API

GAWK_API_MINOR_VERSION

The minor version of the API

The minor version increases when new functions are added to the API. Such new functions are always added to the end of the API struct.

The major version increases (and the minor version is reset to zero) if any of the data types change size or member order, or if any of the existing functions change signature.

It could happen that an extension may be compiled against one version of the API but loaded by a version of gawk using a different version. For this reason, the major and minor API versions of the running gawk are included in the API struct as read-only constant integers:

api->major_version

The major version of the running gawk

api->minor_version

The minor version of the running gawk

It is up to the extension to decide if there are API incompatibilities. Typically, a check like this is enough:

     if (api->major_version != GAWK_API_MAJOR_VERSION
         || api->minor_version < GAWK_API_MINOR_VERSION) {

             fprintf(stderr, "foo_extension: version mismatch with gawk!\n");
             fprintf(stderr, "\tmy version (%d, %d), gawk version (%d, %d)\n",
                     GAWK_API_MAJOR_VERSION, GAWK_API_MINOR_VERSION,
                     api->major_version, api->minor_version);
             exit(1);
     }

Such code is included in the boilerplate dl_load_func() macro provided in gawkapi.h (discussed in Section 16.4.13 [Boilerplate Code], page 386).

16.4.12.2 Informational Variables

The API provides access to several variables that describe whether the corresponding command-line options were enabled when gawk was invoked. The variables are:

do_debug This variable is true if gawk was invoked with --debug option.

do_lint This variable is true if gawk was invoked with --lint option.

do_mpfr This variable is true if gawk was invoked with --bignum option.

do_profile

This variable is true if gawk was invoked with --profile option.

do_sandbox

This variable is true if gawk was invoked with --sandbox option.

do_traditional

This variable is true if gawk was invoked with --traditional option.

The value of do_lint can change if awk code modifies the LINT predefined variable (see Section 7.5 [Predefined Variables], page 147). The others should not change during execution.

16.4.13 Boilerplate Code

As mentioned earlier (see Section 16.3 [How It Works at a High Level], page 355), the function definitions as presented are really macros. To use these macros, your extension must provide a small amount of boilerplate code (variables and functions) toward the top of your source file, using predefined names as described here. The boilerplate needed is also provided in comments in the gawkapi.h header file:

     /* Boilerplate code: */
     int plugin_is_GPL_compatible;
     static gawk_api_t *const api;
     static awk_ext_id_t ext_id;
     static const char *ext_version = NULL; /* or ... = "some string" */
     static awk_ext_func_t func_table[] = {
         { "name", do_name, 1 },
         /* ... */
     };

/* EITHER: */
static awk_bool_t (*init_func)(void) = NULL;
/* OR: */
static awk_bool_t
init_my_extension(void)
{
...
}
static awk_bool_t (*init_func)(void) = init_my_extension;
dl_load_func(func_table, some_name, "name_space_in_quotes")

These variables and functions are as follows:

int plugin_is_GPL_compatible;

This asserts that the extension is compatible with the GNU GPL (see [GNU General Public License], page 475). If your extension does not have this, gawk will not load it (see Section 16.2 [Extension Licensing], page 355).

static gawk_api_t *const api;

This global static variable should be set to point to the gawk_api_t pointer that gawk passes to your dl_load() function. This variable is used by all of the macros.

static awk_ext_id_t ext_id;

This global static variable should be set to the awk_ext_id_t value that gawk passes to your dl_load() function. This variable is used by all of the macros.

static const char *ext_version = NULL; /* or ... = "some string" */

This global static variable should be set either to NULL, or to point to a string giving the name and version of your extension.

static awk_ext_func_t func_table[] = { ... };

This is an array of one or more awk_ext_func_t structures, as described earlier (see Section 16.4.5.1 [Registering An Extension Function], page 363). It can then be looped over for multiple calls to add_ext_func().

static awk_bool_t (*init_func)(void) = NULL;
         OR
static awk_bool_t init_my_extension(void) { ... }
static awk_bool_t (*init_func)(void) = init_my_extension;

If you need to do some initialization work, you should define a function that does it (creates variables, opens files, etc.) and then define the init_func pointer to point to your function. The function should return awk_false upon failure, or awk_true if everything goes well. If you don’t need to do any initialization, define the pointer and initialize it to

NULL.

dl_load_func(func_table, some_name, "name_space_in_quotes")

This macro expands to a dl_load() function that performs all the necessary initializations.

The point of all the variables and arrays is to let the dl_load() function (from the dl_load_func() macro) do all the standard work. It does the following:

1. Check the API versions. If the extension major version does not match gawk’s, or if the extension minor version is greater than gawk’s, it prints a fatal error message and exits.

2. Load the functions defined in func_table. If any of them fails to load, it prints a warning message but continues on.

3. If the init_func pointer is not NULL, call the function it points to. If it returns awk_ false, print a warning message.

4. If ext_version is not NULL, register the version string with gawk.

16.5 How gawk Finds Extensions

Compiled extensions have to be installed in a directory where gawk can find them. If gawk is configured and built in the default fashion, the directory in which to find extensions is /usr/local/lib/gawk. You can also specify a search path with a list of directories to search for compiled extensions. See Section 2.5.2 [The AWKLIBPATH Environment Variable], page 40, for more information.

16.6 Example: Some File Functions

No matter where you go, there you are.

—Buckaroo Banzai Two useful functions that are not in awk are chdir() (so that an awk program can

change its directory) and stat() (so that an awk program can gather information about a file). In order to illustrate the API in action, this section implements these functions for gawk in an extension.

16.6.1 Using chdir() and stat()

This section shows how to use the new functions at the awk level once they’ve been integrated into the running gawk interpreter. Using chdir() is very straightforward. It takes one argument, the new directory to change to:

     @load "filefuncs"
     ...
     newdir = "/home/arnold/funstuff"
     ret = chdir(newdir)
     if (ret < 0) {
         printf("could not change to %s: %s\n", newdir, ERRNO) > "/dev/stderr"
         exit 1
     }
     ...

The return value is negative if the chdir() failed, and ERRNO (see Section 7.5 [Predefined Variables], page 147) is set to a string indicating the error.

Using stat() is a bit more complicated. The C stat() function fills in a structure that has a fair amount of information. The right way to model this in awk is to fill in an associative array with the appropriate information:

     file = "/home/arnold/.profile"
     ret = stat(file, fdata)
     if (ret < 0) {
         printf("could not stat %s: %s\n",
                  file, ERRNO) > "/dev/stderr"
         exit 1
     }
     printf("size of %s is %d bytes\n", file, fdata["size"])

The stat() function always clears the data array, even if the stat() fails. It fills in the following elements:

"name" The name of the file that was stat()ed.

"dev"

"ino" The file’s device and inode numbers, respectively.

"mode" The file’s mode, as a numeric value. This includes both the file’s type and its permissions.

"nlink" The number of hard links (directory entries) the file has.

"uid"

"gid" The numeric user and group ID numbers of the file’s owner.

"size" The size in bytes of the file.

"blocks" The number of disk blocks the file actually occupies. This may not be a function of the file’s size if the file has holes.

"atime"
"mtime"

"ctime" The file’s last access, modification, and inode update times, respectively. These are numeric timestamps, suitable for formatting with strftime() (see Section 9.1.5 [Time Functions], page 192).

"pmode" The file’s “printable mode.” This is a string representation of the file’s type and permissions, such as is produced by ‘ls -l’—for example, "drwxr-xr-x".

"type" A printable string representation of the file’s type. The value is one of the following:

          "blockdev"
          "chardev"

The file is a block or character device (“special file”).

          "directory"

The file is a directory.

"fifo" The file is a named pipe (also known as a FIFO).

"file" The file is just a regular file.

"socket" The file is an AF_UNIX (“Unix domain”) socket in the filesystem.

          "symlink"

The file is a symbolic link.

"devbsize"

The size of a block for the element indexed by "blocks". This information is derived from either the DEV_BSIZE constant defined in <sys/param.h> on most systems, or the S_BLKSIZE constant in <sys/stat.h> on BSD systems. For some other systems, a priori knowledge is used to provide a value. Where no value can be determined, it defaults to 512.

Several additional elements may be present, depending upon the operating system and the type of the file. You can test for them in your awk program by using the in operator (see Section 8.1.2 [Referring to an Array Element], page 161):

"blksize"

The preferred block size for I/O to the file. This field is not present on all POSIX-like systems in the C stat structure.

"linkval"

If the file is a symbolic link, this element is the name of the file the link points to (i.e., the value of the link).

"rdev"
"major"

"minor" If the file is a block or character device file, then these values represent the numeric device number and the major and minor components of that number, respectively.

16.6.2 C Code for chdir() and stat()

Here is the C code for these extensions.^{8}

The file includes a number of standard header files, and then includes the gawkapi.h header file, which provides the API definitions. Those are followed by the necessary variable declarations to make use of the API macros and boilerplate code (see Section 16.4.13 [Boilerplate Code], page 386):

     #ifdef HAVE_CONFIG_H
     #include <config.h>
     #endif
     #include <stdio.h>
     #include <assert.h>
     #include <errno.h>
     #include <stdlib.h>
     #include <string.h>
     #include <unistd.h>

8 This version is edited slightly for presentation. See extension/filefuncs.c in the gawk distribution for

the complete version.

#include <sys/types.h>
#include <sys/stat.h>
#include "gawkapi.h"
#include "gettext.h"
#define _(msgid)  gettext(msgid)
#define N_(msgid) msgid
#include "gawkfts.h"
#include "stack.h"
static const gawk_api_t *api;    /* for convenience macros to work */
static awk_ext_id_t *ext_id;
static awk_bool_t init_filefuncs(void);
static awk_bool_t (*init_func)(void) = init_filefuncs;
static const char *ext_version = "filefuncs extension: version 1.0";
int plugin_is_GPL_compatible;

By convention, for an awk function foo(), the C function that implements it is called do_foo(). The function should have two arguments. The first is an int, usually called nargs, that represents the number of actual arguments for the function. The second is a pointer to an awk_value_t structure, usually named result:

/*  do_chdir --- provide dynamically loaded chdir() function for gawk */
static awk_value_t *
do_chdir(int nargs, awk_value_t *result)
{
    awk_value_t newdir;
    int ret = -1;
    assert(result != NULL);
    if (do_lint && nargs != 1)
        lintwarn(ext_id,
                 _("chdir: called with incorrect number of arguments, "
                   "expecting 1"));

The newdir variable represents the new directory to change to, which is retrieved with get_argument(). Note that the first argument is numbered zero.

If the argument is retrieved successfully, the function calls the chdir() system call. If the chdir() fails, ERRNO is updated:

    if (get_argument(0, AWK_STRING, & newdir)) {
        ret = chdir(newdir.str_value.str);
        if (ret < 0)
            update_ERRNO_int(errno);
    }

Finally, the function returns the return value to the awk level:

    return make_number(ret, result);
}

The stat() extension is more involved. First comes a function that turns a numeric mode into a printable representation (e.g., octal 0644 becomes ‘-rw-r--r--’). This is omitted here for brevity:

/* format_mode --- turn a stat mode field into something readable */
static char *
format_mode(unsigned long fmode)
{
    ...
}

Next comes a function for reading symbolic links, which is also omitted here for brevity:

/* read_symlink --- read a symbolic link into an allocated buffer.
   ... */
static char *
read_symlink(const char *fname, size_t bufsize, ssize_t *linksize)
{
    ...
}

Two helper functions simplify entering values in the array that will contain the result of

the stat():
/* array_set --- set an array element */
static void
array_set(awk_array_t array, const char *sub, awk_value_t *value)
{
    awk_value_t index;
    set_array_element(array,
                      make_const_string(sub, strlen(sub), & index),
                      value);
}
/* array_set_numeric --- set an array element with a number */
static void
array_set_numeric(awk_array_t array, const char *sub, double num)
{
    awk_value_t tmp;
    array_set(array, sub, make_number(num, & tmp));

}

The following function does most of the work to fill in the awk_array_t result array with values obtained from a valid struct stat. This work is done in a separate function to support the stat() function for gawk and also to support the fts() extension, which is included in the same file but whose code is not shown here (see Section 16.7.1 [File-Related Functions], page 398).

The first part of the function is variable declarations, including a table to map file types to strings:

/* fill_stat_array --- do the work to fill an array with stat info */
static int
fill_stat_array(const char *name, awk_array_t array, struct stat *sbuf)
{
    char *pmode;    /* printable mode */
    const char *type = "unknown";
    awk_value_t tmp;
    static struct ftype_map {
        unsigned int mask;
        const char *type;
    } ftype_map[] = {
        { S_IFREG, "file" },
        { S_IFBLK, "blockdev" },
        { S_IFCHR, "chardev" },
        { S_IFDIR, "directory" },
#ifdef S_IFSOCK
        { S_IFSOCK, "socket" },
#endif
#ifdef S_IFIFO
        { S_IFIFO, "fifo" },
#endif
#ifdef S_IFLNK
        { S_IFLNK, "symlink" },
#endif
#ifdef S_IFDOOR /* Solaris weirdness */
        { S_IFDOOR, "door" },
#endif /* S_IFDOOR */
    };
    int j, k;

The destination array is cleared, and then code fills in various elements based on values

in the struct stat:
    /* empty out the array */
    clear_array(array);
    /* fill in the array */
    array_set(array, "name", make_const_string(name, strlen(name),
                                               & tmp));

array_set_numeric(array, "dev", sbuf->st_dev);
array_set_numeric(array, "ino", sbuf->st_ino);
array_set_numeric(array, "mode", sbuf->st_mode);
array_set_numeric(array, "nlink", sbuf->st_nlink);
array_set_numeric(array, "uid", sbuf->st_uid);
array_set_numeric(array, "gid", sbuf->st_gid);
array_set_numeric(array, "size", sbuf->st_size);
array_set_numeric(array, "blocks", sbuf->st_blocks);
array_set_numeric(array, "atime", sbuf->st_atime);
array_set_numeric(array, "mtime", sbuf->st_mtime);
array_set_numeric(array, "ctime", sbuf->st_ctime);
/* for block and character devices, add rdev,
   major and minor numbers */
if (S_ISBLK(sbuf->st_mode) || S_ISCHR(sbuf->st_mode)) {
    array_set_numeric(array, "rdev", sbuf->st_rdev);
    array_set_numeric(array, "major", major(sbuf->st_rdev));
    array_set_numeric(array, "minor", minor(sbuf->st_rdev));
}

The latter part of the function makes selective additions to the destination array, depending upon the availability of certain members and/or the type of the file. It then returns zero, for success:

#ifdef HAVE_STRUCT_STAT_ST_BLKSIZE
array_set_numeric(array, "blksize", sbuf->st_blksize);
#endif /* HAVE_STRUCT_STAT_ST_BLKSIZE */
pmode = format_mode(sbuf->st_mode);
array_set(array, "pmode", make_const_string(pmode, strlen(pmode),
                                            & tmp));
/* for symbolic links, add a linkval field */
if (S_ISLNK(sbuf->st_mode)) {
    char *buf;
    ssize_t linksize;
    if ((buf = read_symlink(name, sbuf->st_size,
                & linksize)) != NULL)
        array_set(array, "linkval",
                  make_malloced_string(buf, linksize, & tmp));
    else
        warning(ext_id, _("stat: unable to read symbolic link `%s'"),
                name);
}
/* add a type field */
type = "unknown";   /* shouldn't happen */

for (j = 0, k = sizeof(ftype_map)/sizeof(ftype_map[0]); j < k; j++) {
    if ((sbuf->st_mode & S_IFMT) == ftype_map[j].mask) {
        type = ftype_map[j].type;
        break;
    }
}
array_set(array, "type", make_const_string(type, strlen(type), & tmp));
return 0;
}

The third argument to stat() was not discussed previously. This argument is optional. If present, it causes do_stat() to use the stat() system call instead of the lstat() system call. This is done by using a function pointer: statfunc. statfunc is initialized to point to lstat() (instead of stat()) to get the file information, in case the file is a symbolic link. However, if the third argument is included, statfunc is set to point to stat(), instead.

Here is the do_stat() function, which starts with variable declarations and argument checking:

/* do_stat --- provide a stat() function for gawk */
static awk_value_t *
do_stat(int nargs, awk_value_t *result)
{
awk_value_t file_param, array_param;
char *name;
awk_array_t array;
int ret;
struct stat sbuf;
/* default is lstat() */
int (*statfunc)(const char *path, struct stat *sbuf) = lstat;
assert(result != NULL);
if (nargs != 2 && nargs != 3) {
    if (do_lint)
        lintwarn(ext_id,
           _("stat: called with wrong number of arguments"));
    return make_number(-1, result);
}

Then comes the actual work. First, the function gets the arguments. Next, it gets the information for the file. If the called function (lstat() or stat()) returns an error, the code sets ERRNO and returns:

/* file is first arg, array to hold results is second */
if (   ! get_argument(0, AWK_STRING, & file_param)
    || ! get_argument(1, AWK_ARRAY, & array_param)) {
    warning(ext_id, _("stat: bad parameters"));

    return make_number(-1, result);
}
if (nargs == 3) {
    statfunc = stat;
}
name = file_param.str_value.str;
array = array_param.array_cookie;
/* always empty out the array */
clear_array(array);
/* stat the file; if error, set ERRNO and return */
ret = statfunc(name, & sbuf);
if (ret < 0) {
    update_ERRNO_int(errno);
    return make_number(ret, result);
}

The tedious work is done by fill_stat_array(), shown earlier. When done, the function returns the result from fill_stat_array():

ret = fill_stat_array(name, array, & sbuf);
return make_number(ret, result);
}

Finally, it’s necessary to provide the “glue” that loads the new function(s) into gawk. The filefuncs extension also provides an fts() function, which we omit here (see

Section 16.7.1 [File-Related Functions], page 398). For its sake, there is an initialization function:

/* init_filefuncs --- initialization routine */
static awk_bool_t
init_filefuncs(void)
{
...
}

We are almost done. We need an array of awk_ext_func_t structures for loading each function into gawk:

static awk_ext_func_t func_table[] = {
{ "chdir", do_chdir, 1 },
{ "stat",  do_stat, 2 },
#ifndef __MINGW32__
{ "fts",   do_fts, 3 },
#endif
};

Each extension must have a routine named dl_load() to load everything that needs to be loaded. It is simplest to use the dl_load_func() macro in gawkapi.h:

/* define the dl_load() function using the boilerplate macro */
dl_load_func(func_table, filefuncs, "")

And that’s it!

16.6.3 Integrating the Extensions

Now that the code is written, it must be possible to add it at runtime to the running gawk interpreter. First, the code must be compiled. Assuming that the functions are in a file named filefuncs.c, and idir is the location of the gawkapi.h header file, the following steps^{9} create a GNU/Linux shared library:

$ gcc -fPIC -shared -DHAVE_CONFIG_H -c -O -g -Iidir filefuncs.c
$ gcc -o filefuncs.so -shared filefuncs.o

Once the library exists, it is loaded by using the @load keyword:

# file testff.awk
@load "filefuncs"
BEGIN {
"pwd" | getline curdir  # save current directory
close("pwd")
chdir("/tmp")
system("pwd")   # test it
chdir(curdir)   # go back
print "Info for testff.awk"
ret = stat("testff.awk", data)
print "ret =", ret
for (i in data)
    printf "data[\"%s\"] = %s\n", i, data[i]
print "testff.awk modified:",
    strftime("%m %d %Y %H:%M:%S", data["mtime"])
print "\nInfo for JUNK"
ret = stat("JUNK", data)
print "ret =", ret
for (i in data)
    printf "data[\"%s\"] = %s\n", i, data[i]
print "JUNK modified:", strftime("%m %d %Y %H:%M:%S", data["mtime"])
}

9 In practice, you would probably want to use the GNU Autotools (Automake, Autoconf, Libtool, and

gettext) to configure and build your libraries. Instructions for doing so are beyond the scope of this book. See Section 16.8 [The gawkextlib Project], page 407, for Internet links to the tools.

The AWKLIBPATH environment variable tells gawk where to find extensions (see Section 16.5 [How gawk Finds Extensions], page 388). We set it to the current directory and run the program:

$ AWKLIBPATH=$PWD gawk -f testff.awk
⊣ /tmp
⊣ Info for testff.awk
⊣ ret = 0
⊣ data["blksize"] = 4096
⊣ data["devbsize"] = 512
⊣ data["mtime"] = 1412004710
⊣ data["mode"] = 33204
⊣ data["type"] = file
⊣ data["dev"] = 2053
⊣ data["gid"] = 1000
⊣ data["ino"] = 10358899
⊣ data["ctime"] = 1412004710
⊣ data["blocks"] = 8
⊣ data["nlink"] = 1
⊣ data["name"] = testff.awk
⊣ data["atime"] = 1412004716
⊣ data["pmode"] = -rw-rw-r--
⊣ data["size"] = 666
⊣ data["uid"] = 1000
⊣ testff.awk modified: 09 29 2014 18:31:50

⊣

⊣ Info for JUNK
⊣ ret = -1
⊣ JUNK modified: 01 01 1970 02:00:00

16.7 The Sample Extensions in the gawk Distribution

This section provides a brief overview of the sample extensions that come in the gawk distribution. Some of them are intended for production use (e.g., the filefuncs, readdir, and inplace extensions). Others mainly provide example code that shows how to use the extension API.

16.7.1 File-Related Functions

The filefuncs extension provides three different functions, as follows. The usage is:

@load "filefuncs"

This is how you load the extension.

result = chdir("/some/directory")

The chdir() function is a direct hook to the chdir() system call to change the current directory. It returns zero upon success or a value less than zero upon error. In the latter case, it updates ERRNO.

result = stat("/some/path", statdata [, follow])

The stat() function provides a hook into the stat() system call. It returns zero upon success or a value less than zero upon error. In the latter case, it updates ERRNO. By default, it uses the lstat() system call. However, if passed a third argument, it uses stat() instead. In all cases, it clears the statdata array. When the call is successful, stat() fills the statdata array with information retrieved from the filesystem, as follows:

Subscript  Field in struct stat  File type
"name"  The file name  All

"dev"        st_dev  All
"ino"        st_ino  All
"mode"       st_mode  All
"nlink"      st_nlink  All
"uid"        st_uid  All
"gid"        st_gid  All
"size"       st_size  All
"atime"      st_atime  All
"mtime"      st_mtime  All
"ctime"      st_ctime  All
"rdev"       st_rdev  Device files
"major"      st_major  Device files
"minor"      st_minor  Device files
"blksize"    st_blksize  All

"pmode"

A human-readable version of the mode value, like that printed by ls (for example,

             "-rwxr-xr-x")

All

"linkval"  The value of the symbolic link  Symbolic links

"type"

The type of the file as a string—one

             of "file", "blockdev", "chardev",
             "directory",  "socket",  "fifo",
             "symlink", "door", or "unknown" (not all

systems support all file types)

All

flags = or(FTS_PHYSICAL, ...)
result = fts(pathlist, flags, filedata)

Walk the file trees provided in pathlist and fill in the filedata array, as described next. flags is the bitwise OR of several predefined values, also described in a moment. Return zero if there were no errors, otherwise return −1.

The fts() function provides a hook to the C library fts() routines for traversing file hierarchies. Instead of returning data about one file at a time in a stream, it fills in a multidimensional array with data about each file and directory encountered in the requested hierarchies.

The arguments are as follows: pathlist An array of file names. The element values are used; the index values are

ignored.

flags This should be the bitwise OR of one or more of the following predefined constant flag values. At least one of FTS_LOGICAL or FTS_PHYSICAL must be provided; otherwise fts() returns an error value and sets ERRNO. The flags are:

FTS_LOGICAL

Do a “logical” file traversal, where the information returned for a symbolic link refers to the linked-to file, and not to the symbolic link itself. This flag is mutually exclusive with FTS_PHYSICAL.

FTS_PHYSICAL

Do a “physical” file traversal, where the information returned for a symbolic link refers to the symbolic link itself. This flag is mutually exclusive with FTS_LOGICAL.

FTS_NOCHDIR

As a performance optimization, the C library fts() routines change directory as they traverse a file hierarchy. This flag disables that optimization.

FTS_COMFOLLOW

Immediately follow a symbolic link named in pathlist, whether or

not FTS_LOGICAL is set.
FTS_SEEDOT

By default, the C library fts() routines do not return entries for . (dot) and .. (dot-dot). This option causes entries for dot-dot to also be included. (The extension always includes an entry for dot; more on this in a moment.)

FTS_XDEV During a traversal, do not cross onto a different mounted filesystem.

filedata The filedata array holds the results. fts() first clears it. Then it creates an element in filedata for every element in pathlist. The index is the name of the directory or file given in pathlist. The element for this index is itself an array. There are two cases:

The path is a file In this case, the array contains two or three elements:

"path" The full path to this file, starting from the “root” that was given in the pathlist array.

"stat" This element is itself an array, containing the same information as provided by the stat() function described earlier for its statdata argument. The element may not be present if the stat() system call for the file failed.

"error" If some kind of error was encountered, the array will also contain an element named "error", which is a string describing the error.

The path is a directory In this case, the array contains one element for each entry in the directory. If an entry is a file, that element is the same as for files, just described. If the entry is a directory, that element is (recursively) an array describing the subdirectory. If FTS_SEEDOT was provided in the flags, then there will also be an element named "..". This element will be an array containing the data as provided

                    by stat().

In addition, there will be an element whose index is ".". This element is an array containing the same two or three elements as

                    for a file: "path", "stat", and "error".

The fts() function returns zero if there were no errors. Otherwise, it returns −1. NOTE: The fts() extension does not exactly mimic the interface of the C library fts() routines, choosing instead to provide an interface that is based on associative arrays, which is more comfortable to use from an awk program. This includes the lack of a comparison function, because gawk already provides powerful array sorting facilities. Although an fts_read()-like interface could have been provided, this felt less natural than simply creating a multidimensional array to represent the file hierarchy and its information.

See test/fts.awk in the gawk distribution for an example use of the fts() extension function.

16.7.2 Interface to fnmatch()

This extension provides an interface to the C library fnmatch() function. The usage is:

@load "fnmatch"

This is how you load the extension.

result = fnmatch(pattern, string, flags)

The return value is zero on success, FNM_NOMATCH if the string did not match the pattern, or a different nonzero value if an error occurred.

In addition to the fnmatch() function, the fnmatch extension adds one constant (FNM_ NOMATCH), and an array of flag values named FNM.

The arguments to fnmatch() are:

pattern The file name wildcard to match

string The file name string

flag Either zero, or the bitwise OR of one or more of the flags in the FNM array

The flags are as follows:

Array element  Corresponding flag defined by fnmatch()

FNM["CASEFOLD"]      FNM_CASEFOLD
FNM["FILE_NAME"]     FNM_FILE_NAME
FNM["LEADING_DIR"]   FNM_LEADING_DIR
FNM["NOESCAPE"]      FNM_NOESCAPE

FNM["PATHNAME"]      FNM_PATHNAME
FNM["PERIOD"]        FNM_PERIOD

Here is an example:

     @load "fnmatch"
     ...
     flags = or(FNM["PERIOD"], FNM["NOESCAPE"])
     if (fnmatch("*.a", "foo.c", flags) == FNM_NOMATCH)
         print "no match"

16.7.3 Interface to fork(), wait(), and waitpid()

The fork extension adds three functions, as follows:

@load "fork"

This is how you load the extension.

pid = fork()

This function creates a new process. The return value is zero in the child and the process ID number of the child in the parent, or −1 upon error. In the latter case, ERRNO indicates the problem. In the child, PROCINFO["pid"] and PROCINFO["ppid"] are updated to reflect the correct values.

ret = waitpid(pid)

This function takes a numeric argument, which is the process ID to wait for. The return value is that of the waitpid() system call.

ret = wait()

This function waits for the first child to die. The return value is that of the wait() system call.

There is no corresponding exec() function.

Here is an example:

     @load "fork"
     ...
     if ((pid = fork()) == 0)
         print "hello from the child"
     else
         print "hello from the parent"

16.7.4 Enabling In-Place File Editing

The inplace extension emulates GNU sed’s -i option, which performs “in-place” editing of each input file. It uses the bundled inplace.awk include file to invoke the extension properly:

# inplace --- load and invoke the inplace extension.
@load "inplace"
# Please set INPLACE_SUFFIX to make a backup copy.  For example, you may
# want to set INPLACE_SUFFIX to .bak on the command line or in a BEGIN rule.
# N.B. We call inplace_end() in the BEGINFILE and END rules so that any
# actions in an ENDFILE rule will be redirected as expected.
BEGINFILE {
    if (_inplace_filename != "")
        inplace_end(_inplace_filename, INPLACE_SUFFIX)
    inplace_begin(_inplace_filename = FILENAME, INPLACE_SUFFIX)
}
END {
    inplace_end(FILENAME, INPLACE_SUFFIX)
}

For each regular file that is processed, the extension redirects standard output to a temporary file configured to have the same owner and permissions as the original. After the file has been processed, the extension restores standard output to its original destination. If INPLACE_SUFFIX is not an empty string, the original file is linked to a backup file name created by appending that suffix. Finally, the temporary file is renamed to the original file name.

The _inplace_filename variable serves to keep track of the current filename so as to not invoke inplace_end() before processing the first file.

If any error occurs, the extension issues a fatal error to terminate processing immediately without damaging the original file.

Here are some simple examples:

$ gawk -i inplace '{ gsub(/foo/, "bar") }; { print }' file1 file2 file3

To keep a backup copy of the original files, try this:

$ gawk -i inplace -v INPLACE_SUFFIX=.bak '{ gsub(/foo/, "bar") }
> { print }' file1 file2 file3

16.7.5 Character and Numeric values: ord() and chr()

The ordchr extension adds two functions, named ord() and chr(), as follows:

@load "ordchr"

This is how you load the extension.

number = ord(string)

Return the numeric value of the first character in string.

char = chr(number)

Return a string whose first character is that represented by number.

These functions are inspired by the Pascal language functions of the same name. Here is an example:

     @load "ordchr"
     ...
     printf("The numeric value of 'A' is %d\n", ord("A"))
     printf("The string value of 65 is %s\n", chr(65))

16.7.6 Reading Directories

The readdir extension adds an input parser for directories. The usage is as follows:

     @load "readdir"

When this extension is in use, instead of skipping directories named on the command line (or with getline), they are read, with each entry returned as a record.

The record consists of three fields. The first two are the inode number and the file name, separated by a forward slash character. On systems where the directory entry contains the file type, the record has a third field (also separated by a slash), which is a single letter indicating the type of the file. The letters and their corresponding file types are shown in Table 16.2.

Letter  File type
b  Block device
c  Character device
d  Directory
f  Regular file
l  Symbolic link
p  Named pipe (FIFO)
s  Socket
u  Anything else (unknown)

Table 16.2: File types returned by the readdir extension

On systems without the file type information, the third field is always ‘u’.

NOTE: On GNU/Linux systems, there are filesystems that don’t support the d_type entry (see the readdir(3) manual page), and so the file type is always ‘u’. You can use the filefuncs extension to call stat() in order to get correct type information.

Here is an example:

     @load "readdir"
     ...
     BEGIN { FS = "/" }
     { print "file name is", $2 }

16.7.7 Reversing Output

The revoutput extension adds a simple output wrapper that reverses the characters in each output line. Its main purpose is to show how to write an output wrapper, although it may be mildly amusing for the unwary. Here is an example:

     @load "revoutput"
     BEGIN {
         REVOUT = 1
         print "don't panic" > "/dev/stdout"
     }

The output from this program is ‘cinap t'nod’.

16.7.8 Two-Way I/O Example

The revtwoway extension adds a simple two-way processor that reverses the characters in each line sent to it for reading back by the awk program. Its main purpose is to show how to write a two-way processor, although it may also be mildly amusing. The following example shows how to use it:

     @load "revtwoway"
     BEGIN {
         cmd = "/magic/mirror"
         print "don't panic" |& cmd
         cmd |& getline result
         print result
         close(cmd)
     }

The output from this program also is: ‘cinap t'nod’.

16.7.9 Dumping and Restoring an Array

The rwarray extension adds two functions, named writea() and reada(), as follows:

@load "rwarray"

This is how you load the extension.

ret = writea(file, array)

This function takes a string argument, which is the name of the file to which to dump the array, and the array itself as the second argument. writea() understands arrays of arrays. It returns one on success, or zero upon failure.

ret = reada(file, array)

reada() is the inverse of writea(); it reads the file named as its first argument, filling in the array named as the second argument. It clears the array first. Here too, the return value is one on success, or zero upon failure.

The array created by reada() is identical to that written by writea() in the sense that the contents are the same. However, due to implementation issues, the array traversal order of the re-created array is likely to be different from that of the original array. As array traversal order in awk is by default undefined, this is (technically) not a problem. If you need to guarantee a particular traversal order, use the array sorting features in gawk to do so (see Section 12.2 [Controlling Array Traversal and Array Sorting], page 300).

The file contains binary data. All integral values are written in network byte order. However, double-precision floating-point values are written as native binary data. Thus, arrays containing only string data can theoretically be dumped on systems with one byte order and restored on systems with a different one, but this has not been tried.

Here is an example:

@load "rwarray"
...
ret = writea("arraydump.bin", array)
...
ret = reada("arraydump.bin", array)

16.7.10 Reading an Entire File

The readfile extension adds a single function named readfile(), and an input parser:

@load "readfile"

This is how you load the extension.

result = readfile("/some/path")

The argument is the name of the file to read. The return value is a string containing the entire contents of the requested file. Upon error, the function returns the empty string and sets ERRNO.

BEGIN { PROCINFO["readfile"] = 1 }

In addition, the extension adds an input parser that is activated if PROCINFO["readfile"] exists. When activated, each input file is returned in its entirety as $0. RT is set to the null string.

Here is an example:

@load "readfile"
...
contents = readfile("/path/to/file");
if (contents == "" && ERRNO != "") {
print("problem reading file", ERRNO) > "/dev/stderr"
...
}

16.7.11 Extension Time Functions

The time extension adds two functions, named gettimeofday() and sleep(), as follows:

@load "time"

This is how you load the extension.

the_time = gettimeofday()

Return the time in seconds that has elapsed since 1970-01-01 UTC as a floatingpoint value. If the time is unavailable on this platform, return −1 and set ERRNO. The returned time should have sub-second precision, but the actual precision may vary based on the platform. If the standard C gettimeofday() system call is available on this platform, then it simply returns the value. Otherwise, if on MS-Windows, it tries to use GetSystemTimeAsFileTime().

result = sleep(seconds)

Attempt to sleep for seconds seconds. If seconds is negative, or the attempt to sleep fails, return −1 and set ERRNO. Otherwise, return zero after sleeping for the indicated amount of time. Note that seconds may be a floating-point (nonintegral) value. Implementation details: depending on platform availability, this function tries to use nanosleep() or select() to implement the delay.

16.7.12 API Tests

The testext extension exercises parts of the extension API that are not tested by the other samples. The extension/testext.c file contains both the C code for the extension and awk test code inside C comments that run the tests. The testing framework extracts the awk code and runs the tests. See the source file for more information.

16.8 The gawkextlib Project

The gawkextlib (http://sourceforge.net/projects/gawkextlib/) project provides a

number of gawk extensions, including one for processing XML files. This is the evolution of the original xgawk (XML gawk) project.

As of this writing, there are seven extensions:

• errno extension

• GD graphics library extension

• PDF extension

• PostgreSQL extension

• MPFR library extension (this provides access to a number of MPFR functions that gawk’s native MPFR support does not)

• Redis extension

• XML parser extension, using the Expat (http://expat.sourceforge.net) XML parsing library

You can check out the code for the gawkextlib project using the Git (http://git-scm. com) distributed source code control system. The command is as follows:

     git clone git://git.code.sf.net/p/gawkextlib/code gawkextlib-code

You will need to have the Expat (http://expat.sourceforge.net) XML parser library installed in order to build and use the XML extension.

In addition, you must have the GNU Autotools installed (Autoconf (http://www.

gnu.org/software/autoconf), Automake (http://www.gnu.org/software/automake),
Libtool (http://www.gnu.org/software/libtool), and GNU gettext (http://www.
gnu.org/software/gettext)).

The simple recipe for building and testing gawkextlib is as follows. First, build and install gawk:

     cd .../path/to/gawk/code
     ./configure --prefix=/tmp/newgawk  Install in /tmp/newgawk for now

make && make check Build and check that all is OK

     make install  Install gawk
   Next, go to http://sourceforge.net/projects/gawkextlib/files to download

gawkextlib and any extensions that you would like to build. The README file at that site explains how to build the code. If you installed gawk in a non-standard location, you will

need to specify ./configure --with-gawk=/path/to/gawk to find it. You may need to use

the sudo utility to install both gawk and gawkextlib, depending upon how your system works.

If you write an extension that you wish to share with other gawk users, consider doing so through the gawkextlib project. See the project’s website for more information.

16.9 Summary

• You can write extensions (sometimes called plug-ins) for gawk in C or C++ using the application programming interface (API) defined by the gawk developers.

• Extensions must have a license compatible with the GNU General Public License (GPL), and they must assert that fact by declaring a variable named plugin_is_

GPL_compatible.

• Communication between gawk and an extension is two-way. gawk passes a struct to the extension that contains various data fields and function pointers. The extension can then call into gawk via the supplied function pointers to accomplish certain tasks.

• One of these tasks is to “register” the name and implementation of new awk-level functions with gawk. The implementation takes the form of a C function pointer with a defined signature. By convention, implementation functions are named do_XXXX() for some awk-level function XXXX().

• The API is defined in a header file named gawkapi.h. You must include a number of standard header files before including it in your source file.

• API function pointers are provided for the following kinds of operations:

• Allocating, reallocating, and releasing memory

• Registration functions (you may register extension functions, exit callbacks, a version string, input parsers, output wrappers, and two-way processors)

• Printing fatal, warning, and “lint” warning messages

• Updating ERRNO, or unsetting it

• Accessing parameters, including converting an undefined parameter into an array • Symbol table access (retrieving a global variable, creating one, or changing one)

• Creating and releasing cached values; this provides an efficient way to use values for multiple variables and can be a big performance win

• Manipulating arrays (retrieving, adding, deleting, and modifying elements; getting the count of elements in an array; creating a new array; clearing an array; and flattening an array for easy C-style looping over all its indices and elements)

• The API defines a number of standard data types for representing awk values, array elements, and arrays.

• The API provides convenience functions for constructing values. It also provides memory management functions to ensure compatibility between memory allocated by gawk and memory allocated by an extension.

• All memory passed from gawk to an extension must be treated as read-only by the extension.

• All memory passed from an extension to gawk must come from the API’s memory allocation functions. gawk takes responsibility for the memory and releases it when appropriate.

• The API provides information about the running version of gawk so that an extension can make sure it is compatible with the gawk that loaded it.

• It is easiest to start a new extension by copying the boilerplate code described in this chapter. Macros in the gawkapi.h header file make this easier to do.

• The gawk distribution includes a number of small but useful sample extensions. The gawkextlib project includes several more (larger) extensions. If you wish to write an extension and contribute it to the community of gawk users, the gawkextlib project is the place to do so.

16.10 Exercises

1. Add functions to implement system calls such as chown(), chmod(), and umask() to the file operations extension presented in Section 16.6.2 [C Code for chdir() and stat()], page 390.

2. Write an input parser that prints a prompt if the input is a from a “terminal” device. You can use the isatty() function to tell if the input file is a terminal. (Hint: this function is usually expensive to call; try to call it just once.) The content of the prompt should come from a variable settable by awk-level code. You can write the prompt to stanard error. However, for best results, open a new file descriptor (or file pointer) on /dev/tty and print the prompt there, in case standard error has been redirected. Why is standard error a better choice than standard output for writing the prompt? Which reading mechanism should you replace, the one to get a record, or the one to read raw bytes?

3. (Hard.) How would you provide namespaces in gawk, so that the names of functions in different extensions don’t conflict with each other? If you come up with a really good scheme, contact the gawk maintainer to tell him about it.

4. Write a wrapper script that provides an interface similar to ‘sed -i’ for the “inplace” extension presented in Section 16.7.4 [Enabling In-Place File Editing], page 402.

Part IV: Appendices

Appendix A: The Evolution of the awk Language 413

Appendix A The Evolution of the awk Language

This book describes the GNU implementation of awk, which follows the POSIX specification. Many longtime awk users learned awk programming with the original awk implementation in Version 7 Unix. (This implementation was the basis for awk in Berkeley Unix, through 4.3-Reno. Subsequent versions of Berkeley Unix, and, for a while, some systems derived from 4.4BSD-Lite, used various versions of gawk for their awk.) This chapter briefly describes the evolution of the awk language, with cross-references to other parts of the book where you can find more information.

A.1 Major Changes Between V7 and SVR3.1

The awk language evolved considerably between the release of Version 7 Unix (1978) and the new version that was first made generally available in System V Release 3.1 (1987). This section summarizes the changes, with cross-references to further details:

• The requirement for ‘;’ to separate rules on a line (see Section 1.6 [awk Statements Versus Lines], page 27)

• User-defined functions and the return statement (see Section 9.2 [User-Defined Functions], page 199)

• The delete statement (see Section 8.4 [The delete Statement], page 168).

• The do-while statement (see Section 7.4.3 [The do-while Statement], page 140)

• The built-in functions atan2(), cos(), sin(), rand(), and srand() (see Section 9.1.2 [Numeric Functions], page 176)

• The built-in functions gsub(), sub(), and match() (see Section 9.1.3 [String-Manipulation Functions], page 177)

• The built-in functions close() and system() (see Section 9.1.4 [Input/Output Functions], page 189)

• The ARGC, ARGV, FNR, RLENGTH, RSTART, and SUBSEP predefined variables (see Section 7.5 [Predefined Variables], page 147)

• Assignable $0 (see Section 4.4 [Changing the Contents of a Field], page 65)

• The conditional expression using the ternary operator ‘?:’ (see Section 6.3.4 [Conditional Expressions], page 125)

• The expression ‘indx in array’ outside of for statements (see Section 8.1.2 [Referring to an Array Element], page 161)

• The exponentiation operator ‘^’ (see Section 6.2.1 [Arithmetic Operators], page 114) and its assignment operator form ‘^=’ (see Section 6.2.3 [Assignment Expressions], page 116)

• C-compatible operator precedence, which breaks some old awk programs (see Section 6.5 [Operator Precedence (How Operators Nest)], page 128)

• Regexps as the value of FS (see Section 4.5 [Specifying How Fields Are Separated], page 67) and as the third argument to the split() function (see Section 9.1.3 [String-Manipulation Functions], page 177), rather than using only the first character of FS

• Dynamic regexps as operands of the ‘~’ and ‘!~’ operators (see Section 3.6 [Using Dynamic Regexps], page 52)

• The escape sequences ‘\b’, ‘\f’, and ‘\r’ (see Section 3.2 [Escape Sequences], page 46)

• Redirection of input for the getline function (see Section 4.9 [Explicit Input with getline], page 78)

• Multiple BEGIN and END rules (see Section 7.1.4 [The BEGIN and END Special Patterns], page 134)

• Multidimensional arrays (see Section 8.5 [Multidimensional Arrays], page 169)

A.2 Changes Between SVR3.1 and SVR4

The System V Release 4 (1989) version of Unix awk added these features (some of which originated in gawk):

• The ENVIRON array (see Section 7.5 [Predefined Variables], page 147)

• Multiple -f options on the command line (see Section 2.2 [Command-Line Options], page 31)

• The -v option for assigning variables before program execution begins (see Section 2.2 [Command-Line Options], page 31)

• The -- signal for terminating command-line options

• The ‘\a’, ‘\v’, and ‘\x’ escape sequences (see Section 3.2 [Escape Sequences], page 46)

• A defined return value for the srand() built-in function (see Section 9.1.2 [Numeric Functions], page 176)

• The toupper() and tolower() built-in string functions for case translation (see Section 9.1.3 [String-Manipulation Functions], page 177)

• A cleaner specification for the ‘%c’ format-control letter in the printf function (see Section 5.5.2 [Format-Control Letters], page 93)

• The ability to dynamically pass the field width and precision ("%*.*d") in the argument list of printf and sprintf() (see Section 5.5.2 [Format-Control Letters], page 93)

• The use of regexp constants, such as /foo/, as expressions, where they are equivalent to using the matching operator, as in ‘$0 ~ /foo/’ (see Section 6.1.2 [Using Regular Expression Constants], page 109)

• Processing of escape sequences inside command-line variable assignments (see Section 6.1.3.2 [Assigning Variables on the Command Line], page 111)

A.3 Changes Between SVR4 and POSIX awk

The POSIX Command Language and Utilities standard for awk (1992) introduced the following changes into the language:

• The use of -W for implementation-specific options (see Section 2.2 [Command-Line Options], page 31)

• The use of CONVFMT for controlling the conversion of numbers to strings (see Section 6.1.4 [Conversion of Strings and Numbers], page 111)

• The concept of a numeric string and tighter comparison rules to go with it (see Section 6.3.2 [Variable Typing and Comparison Expressions], page 120)

• The use of predefined variables as function parameter names is forbidden (see Section 9.2.1 [Function Definition Syntax], page 199)

• More complete documentation of many of the previously undocumented features of the language

In 2012, a number of extensions that had been commonly available for many years were finally added to POSIX. They are:

• The fflush() built-in function for flushing buffered output (see Section 9.1.4 [Input/Output Functions], page 189)

• The nextfile statement (see Section 7.4.9 [The nextfile Statement], page 145)

• The ability to delete all of an array at once with ‘delete array’ (see Section 8.4 [The delete Statement], page 168)

See Section A.7 [Common Extensions Summary], page 423, for a list of common extensions not permitted by the POSIX standard.

The 2008 POSIX standard can be found online at http://www.opengroup.org/

onlinepubs/9699919799/.

A.4 Extensions in Brian Kernighan’s awk

Brian Kernighan has made his version available via his home page (see Section B.5 [Other Freely Available awk Implementations], page 443).

This section describes common extensions that originally appeared in his version of awk:

• The ‘**’ and ‘**=’ operators (see Section 6.2.1 [Arithmetic Operators], page 114 and Section 6.2.3 [Assignment Expressions], page 116)

• The use of func as an abbreviation for function (see Section 9.2.1 [Function Definition Syntax], page 199)

• The fflush() built-in function for flushing buffered output (see Section 9.1.4 [Input/Output Functions], page 189)

See Section A.7 [Common Extensions Summary], page 423, for a full list of the extensions available in his awk.

A.5 Extensions in gawk Not in POSIX awk

The GNU implementation, gawk, adds a large number of features. They can all be disabled with either the --traditional or --posix options (see Section 2.2 [Command-Line Options], page 31).

A number of features have come and gone over the years. This section summarizes the additional features over POSIX awk that are in the current version of gawk.

• Additional predefined variables:

     − The ARGIND, BINMODE, ERRNO, FIELDWIDTHS, FPAT, IGNORECASE, LINT, PROCINFO,

RT, and TEXTDOMAIN variables (see Section 7.5 [Predefined Variables], page 147)

• Special files in I/O redirections:

     − The /dev/stdin, /dev/stdout, /dev/stderr, and /dev/fd/N special file names

(see Section 5.8 [Special File names in gawk], page 101) − The /inet, /inet4, and ‘/inet6’ special files for TCP/IP networking using ‘|&’ to specify which version of the IP protocol to use (see Section 12.4 [Using gawk for Network Programming], page 307)

• Changes and/or additions to the language: − The ‘\x’ escape sequence (see Section 3.2 [Escape Sequences], page 46) − Full support for both POSIX and GNU regexps (see Chapter 3 [Regular Expres-

sions], page 45) − The ability for FS and for the third argument to split() to be null strings (see

Section 4.5.3 [Making Each Character a Separate Field], page 69) − The ability for RS to be a regexp (see Section 4.1 [How Input Is Split into Records],

page 59) − The ability to use octal and hexadecimal constants in awk program source code

(see Section 6.1.1.2 [Octal and Hexadecimal Numbers], page 107) − The ‘|&’ operator for two-way I/O to a coprocess (see Section 12.3 [Two-Way

Communications with Another Process], page 305) − Indirect function calls (see Section 9.3 [Indirect Function Calls], page 208) − Directories on the command line produce a warning and are skipped (see Sec-

tion 4.11 [Directories on the Command Line], page 86)

• New keywords:

− The BEGINFILE and ENDFILE special patterns (see Section 7.1.5 [The BEGINFILE

and ENDFILE Special Patterns], page 135) − The switch statement (see Section 7.4.5 [The switch Statement], page 141)

• Changes to standard awk functions:

− The optional second argument to close() that allows closing one end of a twoway pipe to a coprocess (see Section 12.3 [Two-Way Communications with Another Process], page 305)

− POSIX compliance for gsub() and sub() with --posix − The length() function accepts an array argument and returns the number of

elements in the array (see Section 9.1.3 [String-Manipulation Functions], page 177) − The optional third argument to the match() function for capturing text-matching

subexpressions within a regexp (see Section 9.1.3 [String-Manipulation Functions], page 177)

− Positional specifiers in printf formats for making translations easier (see Sec-

tion 13.4.2 [Rearranging printf Arguments], page 319) − The split() function’s additional optional fourth argument, which is an array

to hold the text of the field separators (see Section 9.1.3 [String-Manipulation Functions], page 177)

• Additional functions only in gawk:

− The gensub(), patsplit(), and strtonum() functions for more powerful text

manipulation (see Section 9.1.3 [String-Manipulation Functions], page 177) − The asort() and asorti() functions for sorting arrays (see Section 12.2 [Con-

trolling Array Traversal and Array Sorting], page 300) − The mktime(), systime(), and strftime() functions for working with time-

stamps (see Section 9.1.5 [Time Functions], page 192)

− The and(), compl(), lshift(), or(), rshift(), and xor() functions for bit

manipulation (see Section 9.1.6 [Bit-Manipulation Functions], page 196) − The isarray() function to check if a variable is an array or not (see Section 9.1.7

[Getting Type Information], page 198)

− The bindtextdomain(), dcgettext(), and dcngettext() functions for interna-

tionalization (see Section 13.3 [Internationalizing awk Programs], page 317)

• Changes and/or additions in the command-line options:

− The AWKPATH environment variable for specifying a path search for the -f

command-line option (see Section 2.2 [Command-Line Options], page 31) − The AWKLIBPATH environment variable for specifying a path search for the -l

command-line option (see Section 2.2 [Command-Line Options], page 31)

− The -b, -c, -C, -d, -D, -e, -E, -g, -h, -i, -l, -L, -M, -n, -N, -o, -O, -p, -P, -r, -S,

-t, and -V short options. Also, the ability to use GNU-style long-named options

   that start with --, and the --assign, --bignum, --characters-as-bytes,
   --copyright, --debug, --dump-variables, --exec, --field-separator,
   --file, --gen-pot, --help, --include, --lint, --lint-old, --load,
   --non-decimal-data, --optimize, --posix, --pretty-print, --profile,
   --re-interval, --sandbox, --source, --traditional, --use-lc-numeric, and

--version long options (see Section 2.2 [Command-Line Options], page 31).

• Support for the following obsolete systems was removed from the code and the documentation for gawk version 4.0: − Amiga − Atari − BeOS − Cray − MIPS RiscOS − MS-DOS with the Microsoft Compiler − MS-Windows with the Microsoft Compiler − NeXT − SunOS 3.x, Sun 386 (Road Runner) − Tandem (non-POSIX) − Prestandard VAX C compiler for VAX/VMS − GCC for VAX and Alpha has not been tested for a while.

• Support for the following obsolete system was removed from the code for gawk version 4.1: − Ultrix

A.6 History of gawk Features

This section describes the features in gawk over and above those in POSIX awk, in the order they were added to gawk.

Version 2.10 of gawk introduced the following features:

• The AWKPATH environment variable for specifying a path search for the -f command-line option (see Section 2.2 [Command-Line Options], page 31).

• The IGNORECASE variable and its effects (see Section 3.8 [Case Sensitivity in Matching], page 55).

• The /dev/stdin, /dev/stdout, /dev/stderr and /dev/fd/N special file names (see

Section 5.8 [Special File names in gawk], page 101). Version 2.13 of gawk introduced the following features:

• The FIELDWIDTHS variable and its effects (see Section 4.6 [Reading Fixed-Width Data], page 72).

• The systime() and strftime() built-in functions for obtaining and printing timestamps (see Section 9.1.5 [Time Functions], page 192).

• Additional command-line options (see Section 2.2 [Command-Line Options], page 31):

− The -W lint option to provide error and portability checking for both the source

code and at runtime. − The -W compat option to turn off the GNU extensions. − The -W posix option for full POSIX compliance.

Version 2.14 of gawk introduced the following feature:

• The next file statement for skipping to the next data file (see Section 7.4.9 [The nextfile Statement], page 145). Version 2.15 of gawk introduced the following features:

• New variables (see Section 7.5 [Predefined Variables], page 147):

− ARGIND, which tracks the movement of FILENAME through ARGV.

− ERRNO, which contains the system error message when getline returns −1 or

      close() fails.
• The /dev/pid, /dev/ppid, /dev/pgrpid, and /dev/user special file names. These

have since been removed.

• The ability to delete all of an array at once with ‘delete array’ (see Section 8.4 [The delete Statement], page 168).

• Command-line option changes (see Section 2.2 [Command-Line Options], page 31): − The ability to use GNU-style long-named options that start with --. − The --source option for mixing command-line and library-file source code.

Version 3.0 of gawk introduced the following features:

• New or changed variables:

− IGNORECASE changed, now applying to string comparison as well as regexp opera-

tions (see Section 3.8 [Case Sensitivity in Matching], page 55). − RT, which contains the input text that matched RS (see Section 4.1 [How Input Is

Split into Records], page 59).

• Full support for both POSIX and GNU regexps (see Chapter 3 [Regular Expressions], page 45).

• The gensub() function for more powerful text manipulation (see Section 9.1.3 [String-Manipulation Functions], page 177).

• The strftime() function acquired a default time format, allowing it to be called with no arguments (see Section 9.1.5 [Time Functions], page 192).

• The ability for FS and for the third argument to split() to be null strings (see Section 4.5.3 [Making Each Character a Separate Field], page 69).

• The ability for RS to be a regexp (see Section 4.1 [How Input Is Split into Records], page 59).

• The next file statement became nextfile (see Section 7.4.9 [The nextfile Statement], page 145).

• The fflush() function from BWK awk (then at Bell Laboratories; see Section 9.1.4 [Input/Output Functions], page 189).

• New command-line options:

− The --lint-old option to warn about constructs that are not available in the original Version 7 Unix version of awk (see Section A.1 [Major Changes Between V7 and SVR3.1], page 413).

− The -m option from BWK awk. (Brian was still at Bell Laboratories at the time.)

This was later removed from both his awk and from gawk. − The --re-interval option to provide interval expressions in regexps (see Sec-

tion 3.3 [Regular Expression Operators], page 48). − The --traditional option was added as a better name for --compat (see Sec-

tion 2.2 [Command-Line Options], page 31).

• The use of GNU Autoconf to control the configuration process (see Section B.2.1 [Compiling gawk for Unix-Like Systems], page 433).

• Amiga support. This has since been removed. Version 3.1 of gawk introduced the following features:

• New variables (see Section 7.5 [Predefined Variables], page 147):

− BINMODE, for non-POSIX systems, which allows binary I/O for input and/or output files (see Section B.3.1.4 [Using gawk on PC Operating Systems], page 437).

− LINT, which dynamically controls lint warnings.

− PROCINFO, an array for providing process-related information.

− TEXTDOMAIN, for setting an application’s internationalization text domain (see

Chapter 13 [Internationalization with gawk], page 315).

• The ability to use octal and hexadecimal constants in awk program source code (see Section 6.1.1.2 [Octal and Hexadecimal Numbers], page 107).

• The ‘|&’ operator for two-way I/O to a coprocess (see Section 12.3 [Two-Way Communications with Another Process], page 305).

• The /inet special files for TCP/IP networking using ‘|&’ (see Section 12.4 [Using gawk for Network Programming], page 307).

• The optional second argument to close() that allows closing one end of a two-way pipe to a coprocess (see Section 12.3 [Two-Way Communications with Another Process], page 305).

• The optional third argument to the match() function for capturing text-matching subexpressions within a regexp (see Section 9.1.3 [String-Manipulation Functions], page 177).

• Positional specifiers in printf formats for making translations easier (see Section 13.4.2 [Rearranging printf Arguments], page 319).

• A number of new built-in functions:

− The asort() and asorti() functions for sorting arrays (see Section 12.2 [Con-

trolling Array Traversal and Array Sorting], page 300).

   − The bindtextdomain(), dcgettext() and dcngettext() functions for interna-

tionalization (see Section 13.3 [Internationalizing awk Programs], page 317). − The extension() function and the ability to add new built-in functions dynami-

cally (see Chapter 16 [Writing Extensions for gawk], page 355). − The mktime() function for creating timestamps (see Section 9.1.5 [Time Func-

tions], page 192).

   − The and(), or(), xor(), compl(), lshift(), rshift(), and strtonum() func-

tions (see Section 9.1.6 [Bit-Manipulation Functions], page 196).

• The support for ‘next file’ as two words was removed completely (see Section 7.4.9 [The nextfile Statement], page 145).

• Additional command-line options (see Section 2.2 [Command-Line Options], page 31): − The --dump-variables option to print a list of all global variables. − The --exec option, for use in CGI scripts. − The --gen-po command-line option and the use of a leading underscore to mark strings that should be translated (see Section 13.4.1 [Extracting Marked Strings], page 319).

− The --non-decimal-data option to allow non-decimal input data (see Section 12.1

[Allowing Nondecimal Input Data], page 299). − The --profile option and pgawk, the profiling version of gawk, for producing execution profiles of awk programs (see Section 12.5 [Profiling Your awk Programs], page 308).

− The --use-lc-numeric option to force gawk to use the locale’s decimal point for parsing input data (see Section 6.1.4 [Conversion of Strings and Numbers], page 111).

• The use of GNU Automake to help in standardizing the configuration process (see Section B.2.1 [Compiling gawk for Unix-Like Systems], page 433).

• The use of GNU gettext for gawk’s own message output (see Section 13.6 [gawk Can Speak Your Language], page 323).

• BeOS support. This was later removed.

• Tandem support. This was later removed.

• The Atari port became officially unsupported and was later removed entirely.

• The source code changed to use ISO C standard-style function definitions.

• POSIX compliance for sub() and gsub() (see Section 9.1.3.1 [More about ‘\’ and ‘&’ with sub(), gsub(), and gensub()], page 186).

• The length() function was extended to accept an array argument and return the number of elements in the array (see Section 9.1.3 [String-Manipulation Functions], page 177).

• The strftime() function acquired a third argument to enable printing times as UTC (see Section 9.1.5 [Time Functions], page 192).

Version 4.0 of gawk introduced the following features:

• Variable additions:

− FPAT, which allows you to specify a regexp that matches the fields, instead of

matching the field separator (see Section 4.7 [Defining Fields by Content], page 74).

 − If PROCINFO["sorted_in"] exists, ‘for(iggy in foo)’ loops sort the indices be-

fore looping over them. The value of this element provides control over how the indices are sorted before the loop traversal starts (see Section 8.1.6 [Using Predefined Array Scanning Orders with gawk], page 164).

− PROCINFO["strftime"], which holds the default format for strftime() (see Sec-

tion 9.1.5 [Time Functions], page 192).

• The special files /dev/pid, /dev/ppid, /dev/pgrpid and /dev/user were removed.

• Support for IPv6 was added via the /inet6 special file. /inet4 forces IPv4 and /inet chooses the system default, which is probably IPv4 (see Section 12.4 [Using gawk for Network Programming], page 307).

• The use of ‘\s’ and ‘\S’ escape sequences in regular expressions (see Section 3.7 [gawk-Specific Regexp Operators], page 54).

• Interval expressions became part of default regular expressions (see Section 3.3 [Regular Expression Operators], page 48).

• POSIX character classes work even with --traditional (see Section 3.3 [Regular Expression Operators], page 48).

• break and continue became invalid outside a loop, even with --traditional (see Section 7.4.6 [The break Statement], page 142, and also see Section 7.4.7 [The continue Statement], page 143).

• fflush(), nextfile, and ‘delete array’ are allowed if --posix or --traditional,

since they are all now part of POSIX.

• An optional third argument to asort() and asorti(), specifying how to sort (see Section 9.1.3 [String-Manipulation Functions], page 177).

• The behavior of fflush() changed to match BWK awk and for POSIX; now both ‘fflush()’ and ‘fflush("")’ flush all open output redirections (see Section 9.1.4 [Input/Output Functions], page 189).

• The isarray() function which distinguishes if an item is an array or not, to make it possible to traverse arrays of arrays (see Section 9.1.7 [Getting Type Information], page 198).

• The patsplit() function which gives the same capability as FPAT, for splitting (see Section 9.1.3 [String-Manipulation Functions], page 177).

• An optional fourth argument to the split() function, which is an array to hold the values of the separators (see Section 9.1.3 [String-Manipulation Functions], page 177).

• Arrays of arrays (see Section 8.6 [Arrays of Arrays], page 171).

• The BEGINFILE and ENDFILE special patterns (see Section 7.1.5 [The BEGINFILE and ENDFILE Special Patterns], page 135).

• Indirect function calls (see Section 9.3 [Indirect Function Calls], page 208).

• switch / case are enabled by default (see Section 7.4.5 [The switch Statement], page 141).

• Command-line option changes (see Section 2.2 [Command-Line Options], page 31):

− The -b and --characters-as-bytes options which prevent gawk from treating

input as a multibyte string. − The redundant --compat, --copyleft, and --usage long options were removed. − The --gen-po option was finally renamed to the correct --gen-pot. − The --sandbox option which disables certain features. − All long options acquired corresponding short options, for use in ‘#!’ scripts.

• Directories named on the command line now produce a warning, not a fatal error, unless --posix or --traditional are used (see Section 4.11 [Directories on the Command Line], page 86).

• The gawk internals were rewritten, bringing the dgawk debugger and possibly improved performance (see Chapter 14 [Debugging awk Programs], page 325).

• Per the GNU Coding Standards, dynamic extensions must now define a global symbol indicating that they are GPL-compatible (see Section 16.2 [Extension Licensing], page 355).

• In POSIX mode, string comparisons use strcoll() / wcscoll() (see Section 6.3.2.3 [String Comparison with POSIX Rules], page 124).

• The option for raw sockets was removed, since it was never implemented (see Section 12.4 [Using gawk for Network Programming], page 307).

• Ranges of the form ‘[d-h]’ are treated as if they were in the C locale, no matter what kind of regexp is being used, and even if --posix (see Section A.8 [Regexp Ranges and Locales: A Long Sad Story], page 424).

• Support was removed for the following systems:

− Atari − Amiga − BeOS − Cray − MIPS RiscOS − MS-DOS with Microsoft Compiler − MS-Windows with Microsoft Compiler − NeXT − SunOS 3.x, Sun 386 (Road Runner) − Tandem (non-POSIX) − Prestandard VAX C compiler for VAX/VMS

Version 4.1 of gawk introduced the following features:

• Three new arrays: SYMTAB, FUNCTAB, and PROCINFO["identifiers"] (see Section 7.5.2 [Built-in Variables That Convey Information], page 149).

• The three executables gawk, pgawk, and dgawk, were merged into one, named just gawk. As a result the command-line options changed.

• Command-line option changes (see Section 2.2 [Command-Line Options], page 31): − The -D option invokes the debugger. − The -i and --include options load awk library files. − The -l and --load options load compiled dynamic extensions. − The -M and --bignum options enable MPFR. − The -o option only does pretty-printing. − The -p option is used for profiling. − The -R option was removed.

• Support for high precision arithmetic with MPFR. (see Chapter 15 [Arithmetic and Arbitrary-Precision Arithmetic with gawk], page 341).

• The and(), or() and xor() functions changed to allow any number of arguments, with a minimum of two (see Section 9.1.6 [Bit-Manipulation Functions], page 196).

• The dynamic extension interface was completely redone (see Chapter 16 [Writing Extensions for gawk], page 355).

• Redirected getline became allowed inside BEGINFILE and ENDFILE (see Section 7.1.5 [The BEGINFILE and ENDFILE Special Patterns], page 135).

• The where command was added to the debugger (see Section 14.3.4 [Working with the Stack], page 334).

A.7 Common Extensions Summary

The following table summarizes the common extensions supported by gawk, Brian Kernighan’s awk, and mawk, the three most widely used freely available versions of awk (see Section B.5 [Other Freely Available awk Implementations], page 443).

Feature  BWK awk   mawk   gawk Now standard
‘\x’ escape sequence  X  X  X
FS as null string  X  X  X
/dev/stdin special file X  X  X
/dev/stdout special file X  X  X
/dev/stderr special file X  X  X
delete without subscript X  X  X  X
fflush() function  X  X  X  X
length() of an array  X  X  X
nextfile statement  X  X  X  X
** and **= operators  X  X
func keyword  X  X
BINMODE variable  X  X
RS as regexp  X  X
Time-related functions  X  X

A.8 Regexp Ranges and Locales: A Long Sad Story

This section describes the confusing history of ranges within regular expressions and their interactions with locales, and how this affected different versions of gawk.

The original Unix tools that worked with regular expressions defined character ranges (such as ‘[a-z]’) to match any character between the first character in the range and the last character in the range, inclusive. Ordering was based on the numeric value of each character in the machine’s native character set. Thus, on ASCII-based systems, ‘[a-z]’ matched all the lowercase letters, and only the lowercase letters, as the numeric values for the letters from ‘a’ through ‘z’ were contiguous. (On an EBCDIC system, the range ‘[a-z]’ includes additional nonalphabetic characters as well.)

Almost all introductory Unix literature explained range expressions as working in this fashion, and in particular, would teach that the “correct” way to match lowercase letters was with ‘[a-z]’, and that ‘[A-Z]’ was the “correct” way to match uppercase letters. And indeed, this was true.^{1}

The 1992 POSIX standard introduced the idea of locales (see Section 6.6 [Where You Are Makes a Difference], page 129). Because many locales include other letters besides the plain 26 letters of the English alphabet, the POSIX standard added character classes (see Section 3.4 [Using Bracket Expressions], page 50) as a way to match different kinds of characters besides the traditional ones in the ASCII character set.

However, the standard changed the interpretation of range expressions. In the "C" and "POSIX" locales, a range expression like ‘[a-dx-z]’ is still equivalent to ‘[abcdxyz]’, as in ASCII. But outside those locales, the ordering was defined to be based on collation order.

What does that mean? In many locales, ‘A’ and ‘a’ are both less than ‘B’. In other words, these locales sort characters in dictionary order, and ‘[a-dx-z]’ is typically not equivalent to ‘[abcdxyz]’; instead, it might be equivalent to ‘[ABCXYabcdxyz]’, for example.

This point needs to be emphasized: much literature teaches that you should use ‘[a-z]’ to match a lowercase character. But on systems with non-ASCII locales, this also matches all of the uppercase characters except ‘A’ or ‘Z’! This was a continuous cause of confusion, even well into the twenty-first century.

To demonstrate these issues, the following example uses the sub() function, which does text replacement (see Section 9.1.3 [String-Manipulation Functions], page 177). Here, the intent is to remove trailing uppercase characters:

     $ echo something1234abc | gawk-3.1.8 '{ sub("[A-Z]*$", ""); print }'
     ⊣ something1234a

This output is unexpected, as the ‘bc’ at the end of ‘something1234abc’ should not normally match ‘[A-Z]*’. This result is due to the locale setting (and thus you may not see it on your system).

Similar considerations apply to other ranges. For example, ‘["-/]’ is perfectly valid in ASCII, but is not valid in many Unicode locales, such as en_US.UTF-8.

Early versions of gawk used regexp matching code that was not locale-aware, so ranges had their traditional interpretation.

1 And Life was good.

When gawk switched to using locale-aware regexp matchers, the problems began; especially as both GNU/Linux and commercial Unix vendors started implementing non-ASCII locales, and making them the default. Perhaps the most frequently asked question became something like, “Why does ‘[A-Z]’ match lowercase letters?!?”

This situation existed for close to 10 years, if not more, and the gawk maintainer grew weary of trying to explain that gawk was being nicely standards-compliant, and that the issue was in the user’s locale. During the development of version 4.0, he modified gawk to always treat ranges in the original, pre-POSIX fashion, unless --posix was used (see Section 2.2 [Command-Line Options], page 31).^{2}

Fortunately, shortly before the final release of gawk 4.0, the maintainer learned that the 2008 standard had changed the definition of ranges, such that outside the "C" and "POSIX" locales, the meaning of range expressions was undefined.^{3}

By using this lovely technical term, the standard gives license to implementors to implement ranges in whatever way they choose. The gawk maintainer chose to apply the pre-POSIX meaning both with the default regexp matching and when --traditional or --posix are used. In all cases gawk remains POSIX-compliant.

A.9 Major Contributors to gawk

Always give credit where credit is due.

—Anonymous This section names the major contributors to gawk and/or this book, in approximate

chronological order:

• Dr. Alfred V. Aho, Dr. Peter J. Weinberger, and Dr. Brian W. Kernighan, all of Bell Laboratories, designed and implemented Unix awk, from which gawk gets the majority of its feature set.

• Paul Rubin did the initial design and implementation in 1986, and wrote the first draft (around 40 pages) of this book.

• Jay Fenlason finished the initial implementation.

• Diane Close revised the first draft of this book, bringing it to around 90 pages.

• Richard Stallman helped finish the implementation and the initial draft of this book. He is also the founder of the FSF and the GNU Project.

• John Woods contributed parts of the code (mostly fixes) in the initial version of gawk.

• In 1988, David Trueman took over primary maintenance of gawk, making it compatible with “new” awk, and greatly improving its performance.

• Conrad Kwok, Scott Garfinkle, and Kent Williams did the initial ports to MS-DOS with various versions of MSC.

• Pat Rankin provided the VMS port and its documentation.

2 And thus was born the Campaign for Rational Range Interpretation (or RRI). A number of GNU

tools have already implemented this change, or will soon. Thanks to Karl Berry for coining the phrase

“Rational Range Interpretation.”

3 See the standard (http://pubs.opengroup.org/onlinepubs/9699919799/basedefs/V1_chap09.

html#tag_09_03_05) and its rationale (http://pubs.opengroup.org/onlinepubs/9699919799/xrat/

V4_xbd_chap09.html#tag_21_09_03_05).

• Hal Peterson provided help in porting gawk to Cray systems. (This is no longer supported.)

• Kai Uwe Rommel provided the initial port to OS/2 and its documentation.

• Michal Jaegermann provided the port to Atari systems and its documentation. (This port is no longer supported.) He continues to provide portability checking, and has done a lot of work to make sure gawk works on non-32-bit systems.

• Fred Fish provided the port to Amiga systems and its documentation. (With Fred’s sad passing, this is no longer supported.)

• Scott Deifik currently maintains the MS-DOS port using DJGPP.

• Eli Zaretskii currently maintains the MS-Windows port using MinGW.

• Juan Grigera provided a port to Windows32 systems. (This is no longer supported.)

• For many years, Dr. Darrel Hankerson acted as coordinator for the various ports to different PC platforms and created binary distributions for various PC operating systems. He was also instrumental in keeping the documentation up to date for the various PC platforms.

• Christos Zoulas provided the extension() built-in function for dynamically adding new functions. (This was obsoleted at gawk 4.1.)

• Jürgen Kahrs contributed the initial version of the TCP/IP networking code and documentation, and motivated the inclusion of the ‘|&’ operator.

• Stephen Davies provided the initial port to Tandem systems and its documentation. (However, this is no longer supported.) He was also instrumental in the initial work to integrate the byte-code internals into the gawk code base.

• Matthew Woehlke provided improvements for Tandem’s POSIX-compliant systems.

• Martin Brown provided the port to BeOS and its documentation. (This is no longer supported.)

• Arno Peters did the initial work to convert gawk to use GNU Automake and GNU

  gettext.

• Alan J. Broder provided the initial version of the asort() function as well as the code for the optional third argument to the match() function.

• Andreas Buening updated the gawk port for OS/2.

• Isamu Hasegawa, of IBM in Japan, contributed support for multibyte characters.

• Michael Benzinger contributed the initial code for switch statements.

• Patrick T.J. McPhee contributed the code for dynamic loading in Windows32 environments. (This is no longer supported.)

• Anders Wallin helped keep the VMS port going for several years.

• Assaf Gordon contributed the code to implement the --sandbox option.

• John Haque made the following contributions:

− The modifications to convert gawk into a byte-code interpreter, including the de-

bugger − The addition of true arrays of arrays − The additional modifications for support of arbitrary-precision arithmetic − The initial text of Chapter 15 [Arithmetic and Arbitrary-Precision Arithmetic with

gawk], page 341 − The work to merge the three versions of gawk into one, for the 4.1 release − Improved array internals for arrays indexed by integers − The improved array sorting features were also driven by John, together with Pat

Rankin

• Panos Papadopoulos contributed the original text for Section 2.7 [Including Other Files into Your Program], page 42.

• Efraim Yawitz contributed the original text for Chapter 14 [Debugging awk Programs], page 325.

• The development of the extension API first released with gawk 4.1 was driven primarily by Arnold Robbins and Andrew Schorr, with notable contributions from the rest of the development team.

• John Malmberg contributed significant improvements to the OpenVMS port and the related documentation.

• Antonio Giovanni Colombo rewrote a number of examples in the early chapters that were severely dated, for which I am incredibly grateful.

• Arnold Robbins has been working on gawk since 1988, at first helping David Trueman, and as the primary maintainer since around 1994.

A.10 Summary

• The awk language has evolved over time. The first release was with V7 Unix, circa 1978. In 1987, for System V Release 3.1, major additions, including user-defined functions, were made to the language. Additional changes were made for System V Release 4, in 1989. Since then, further minor changes have happened under the auspices of the POSIX standard.

• Brian Kernighan’s awk provides a small number of extensions that are implemented in common with other versions of awk.

• gawk provides a large number of extensions over POSIX awk. They can be disabled with either the --traditional or --posix options.

• The interaction of POSIX locales and regexp matching in gawk has been confusing over the years. Today, gawk implements Rational Range Interpretation, where ranges of the form ‘[a-z]’ match only the characters numerically between ‘a’ through ‘z’ in the machine’s native character set. Usually this is ASCII, but it can be EBCDIC on IBM S/390 systems.

• Many people have contributed to gawk development over the years. We hope that the list provided in this chapter is complete and gives the appropriate credit where credit is due.

Appendix B: Installing gawk 429

Appendix B Installing gawk

This appendix provides instructions for installing gawk on the various platforms that are supported by the developers. The primary developer supports GNU/Linux (and Unix), whereas the other ports are contributed. See Section B.4 [Reporting Problems and Bugs], page 442, for the email addresses of the people who maintain the respective ports.

B.1 The gawk Distribution

This section describes how to get the gawk distribution, how to extract it, and then what is in the various files and subdirectories.

B.1.1 Getting the gawk Distribution

There are two ways to get GNU software:

• Copy it from someone else who already has it.

• Retrieve gawk from the Internet host ftp.gnu.org, in the directory /gnu/gawk. Both anonymous ftp and http access are supported. If you have the wget program, you can use a command like the following:

         wget http://ftp.gnu.org/gnu/gawk/gawk-4.1.2.tar.gz

The GNU software archive is mirrored around the world. The up-to-date list of mirror sites is available from the main FSF website (http://www.gnu.org/order/ftp.html). Try to use one of the mirrors; they will be less busy, and you can usually find one closer to your site.

B.1.2 Extracting the Distribution

gawk is distributed as several tar files compressed with different compression programs: gzip, bzip2, and xz. For simplicity, the rest of these instructions assume you are using the one compressed with the GNU Gzip program (gzip).

Once you have the distribution (e.g., gawk-4.1.2.tar.gz), use gzip to expand the file and then use tar to extract it. You can use the following pipeline to produce the gawk distribution:

     gzip -d -c gawk-4.1.2.tar.gz | tar -xvpf -

On a system with GNU tar, you can let tar do the decompression for you:

     tar -xvpzf gawk-4.1.2.tar.gz

Extracting the archive creates a directory named gawk-4.1.2 in the current directory.

The distribution file name is of the form gawk-V.R.P.tar.gz. The V represents the major version of gawk, the R represents the current release of version V, and the P represents a patch level, meaning that minor bugs have been fixed in the release. The current patch level is 2, but when retrieving distributions, you should get the version with the highest version, release, and patch level. (Note, however, that patch levels greater than or equal to 70 denote “beta” or nonproduction software; you might not want to retrieve such a version unless you don’t mind experimenting.) If you are not on a Unix or GNU/Linux system, you need to make other arrangements for getting and extracting the gawk distribution. You should consult a local expert.

B.1.3 Contents of the gawk Distribution

The gawk distribution has a number of C source files, documentation files, subdirectories, and files related to the configuration process (see Section B.2 [Compiling and Installing gawk on Unix-Like Systems], page 433), as well as several subdirectories related to different non-Unix operating systems:

Various ‘.c’, ‘.y’, and ‘.h’ files These files contain the actual gawk source code.

ABOUT-NLS

A file containing information about GNU gettext and translations.

AUTHORS A file with some information about the authorship of gawk. It exists only to satisfy the pedants at the Free Software Foundation.

README
README_d/README.*

Descriptive files: README for gawk under Unix and the rest for the various hardware and software combinations.

INSTALL A file providing an overview of the configuration and installation process.

ChangeLog

A detailed list of source code changes as bugs are fixed or improvements made.

ChangeLog.0

An older list of source code changes.

NEWS A list of changes to gawk since the last release or patch.

NEWS.0 An older list of changes to gawk.

COPYING The GNU General Public License.

POSIX.STD

A description of behaviors in the POSIX standard for awk that are left undefined, or where gawk may not comply fully, as well as a list of things that the POSIX standard should describe but does not.

doc/awkforai.txt

Pointers to the original draft of a short article describing why gawk is a good language for artificial intelligence (AI) programming.

doc/bc_notes

A brief description of gawk’s “byte code” internals.

doc/README.card
doc/ad.block
doc/awkcard.in
doc/cardfonts
doc/colors
doc/macros
doc/no.colors
doc/setter.outline

The troff source for a five-color awk reference card. A modern version of troff such as GNU troff (groff) is needed to produce the color version. See the file README.card for instructions if you have an older troff.

doc/gawk.1

The troff source for a manual page describing gawk. This is distributed for the convenience of Unix users.

doc/gawktexi.in
doc/sidebar.awk

The Texinfo source file for this book. It should be processed by

doc/sidebar.awk before processing with texi2dvi or texi2pdf to produce

a printed document, and with makeinfo to produce an Info or HTML file. The Makefile takes care of this processing and produces printable output via

texi2dvi or texi2pdf.
doc/gawk.texi

The file produced after processing gawktexi.in with sidebar.awk.

doc/gawk.info

The generated Info file for this book.

doc/gawkinet.texi

The Texinfo source file for TCP/IP Internetworking with gawk. It should be processed with TEX (via texi2dvi or texi2pdf) to produce a printed document and with makeinfo to produce an Info or HTML file.

doc/gawkinet.info

The generated Info file for TCP/IP Internetworking with gawk.

doc/igawk.1

The troff source for a manual page describing the igawk program presented in Section 11.3.9 [An Easy Way to Use Library Functions], page 285.

doc/Makefile.in

The input file used during the configuration process to generate the actual Makefile for creating the documentation.

Makefile.am
*/Makefile.am

Files used by the GNU Automake software for generating the Makefile.in files used by Autoconf and configure.

Makefile.in
aclocal.m4
bisonfix.awk
config.guess
configh.in
configure.ac
configure
custom.h
depcomp
install-sh
missing_d/*
mkinstalldirs

m4/* These files and subdirectories are used when configuring and compiling gawk for various Unix systems. Most of them are explained in Section B.2 [Compiling and Installing gawk on Unix-Like Systems], page 433. The rest are there to support the main infrastructure.

po/* The po library contains message translations.

awklib/extract.awk
awklib/Makefile.am
awklib/Makefile.in
awklib/eg/*

The awklib directory contains a copy of extract.awk (see Section 11.3.7 [Extracting Programs from Texinfo Source Files], page 281), which can be used to extract the sample programs from the Texinfo source file for this book. It also contains a Makefile.in file, which configure uses to generate a Makefile. Makefile.am is used by GNU Automake to create Makefile.in. The library functions from Chapter 10 [A Library of awk Functions], page 217, and the igawk program from Section 11.3.9 [An Easy Way to Use Library Functions], page 285, are included as ready-to-use files in the gawk distribution. They are installed as part of the installation process. The rest of the programs in this book are available in appropriate subdirectories of awklib/eg.

extension/*

The source code, manual pages, and infrastructure files for the sample extensions included with gawk. See Chapter 16 [Writing Extensions for gawk], page 355, for more information.

posix/* Files needed for building gawk on POSIX-compliant systems.

pc/* Files needed for building gawk under MS-Windows and OS/2 (see Section B.3.1 [Installation on PC Operating Systems], page 435, for details).

vms/* Files needed for building gawk under Vax/VMS and OpenVMS (see Section B.3.2 [Compiling and Installing gawk on Vax/VMS and OpenVMS], page 439, for details).

test/* A test suite for gawk. You can use ‘make check’ from the top-level gawk directory to run your version of gawk against the test suite. If gawk successfully passes ‘make check’, then you can be confident of a successful port.

B.2 Compiling and Installing gawk on Unix-Like Systems

Usually, you can compile and install gawk by typing only two commands. However, if you use an unusual system, you may need to configure gawk for your system yourself.

B.2.1 Compiling gawk for Unix-Like Systems

The normal installation steps should work on all modern commercial Unix-derived systems, GNU/Linux, BSD-based systems, and the Cygwin environment for MS-Windows.

After you have extracted the gawk distribution, cd to gawk-4.1.2. As with most GNU software, you configure gawk for your system by running the configure program. This program is a Bourne shell script that is generated automatically using GNU Autoconf. (The Autoconf software is described fully in Autoconf—Generating Automatic Configuration Scripts, which can be found online at the Free Software Foundation’s website (http://

www.gnu.org/software/autoconf/manual/index.html).)

To configure gawk, simply run configure:

     sh ./configure

This produces a Makefile and config.h tailored to your system. The config.h file describes various facts about your system. You might want to edit the Makefile to change the CFLAGS variable, which controls the command-line options that are passed to the C compiler (such as optimization levels or compiling for debugging).

Alternatively, you can add your own values for most make variables on the command line, such as CC and CFLAGS, when running configure:

     CC=cc CFLAGS=-g sh ./configure

See the file INSTALL in the gawk distribution for all the details.

After you have run configure and possibly edited the Makefile, type:

     make

Shortly thereafter, you should have an executable version of gawk. That’s all there is to it! To verify that gawk is working properly, run ‘make check’. All of the tests should succeed. If these steps do not work, or if any of the tests fail, check the files in the README_d directory to see if you’ve found a known problem. If the failure is not described there, send in a bug report (see Section B.4 [Reporting Problems and Bugs], page 442).

Of course, once you’ve built gawk, it is likely that you will wish to install it. To do so, you need to run the command ‘make install’, as a user with the appropriate permissions. How to do this varies by system, but on many systems you can use the sudo command to do so. The command then becomes ‘sudo make install’. It is likely that you will be asked for your password, and you will have to have been set up previously as a user who is allowed to run the sudo command.

B.2.2 Additional Configuration Options

There are several additional options you may use on the configure command line when compiling gawk from scratch, including:

--disable-extensions

Disable configuring and building the sample extensions in the extension directory. This is useful for cross-compiling. The default action is to dynamically check if the extensions can be configured and compiled.

--disable-lint

Disable all lint checking within gawk. The --lint and --lint-old options (see Section 2.2 [Command-Line Options], page 31) are accepted, but silently do nothing. Similarly, setting the LINT variable (see Section 7.5.1 [Built-in Variables That Control awk], page 147) has no effect on the running awk program. When used with the GNU Compiler Collection’s (GCC’s) automatic dead-code-elimination, this option cuts almost 23K bytes off the size of the gawk executable on GNU/Linux x86 64 systems. Results on other systems and with other compilers are likely to vary. Using this option may bring you some slight performance improvement.

CAUTION: Using this option will cause some of the tests in the test suite to fail. This option may be removed at a later date.

--disable-nls

Disable all message-translation facilities. This is usually not desirable, but it may bring you some slight performance improvement.

--with-whiny-user-strftime

Force use of the included version of the C strftime() function for deficient systems.

Use the command ‘./configure --help’ to see the full list of options supplied by

configure.

B.2.3 The Configuration Process

This section is of interest only if you know something about using the C language and Unix-like operating systems.

The source code for gawk generally attempts to adhere to formal standards wherever possible. This means that gawk uses library routines that are specified by the ISO C standard and by the POSIX operating system interface standard. The gawk source code requires using an ISO C compiler (the 1990 standard).

Many Unix systems do not support all of either the ISO or the POSIX standards. The missing_d subdirectory in the gawk distribution contains replacement versions of those functions that are most likely to be missing.

The config.h file that configure creates contains definitions that describe features of the particular operating system where you are attempting to compile gawk. The three things described by this file are: what header files are available, so that they can be correctly included, what (supposedly) standard functions are actually available in your C libraries, and various miscellaneous facts about your operating system. For example, there may not be an

st_blksize element in the stat structure. In this case, ‘HAVE_STRUCT_STAT_ST_BLKSIZE’

is undefined.

It is possible for your C compiler to lie to configure. It may do so by not exiting with an error when a library function is not available. To get around this, edit the custom.h file. Use an ‘#ifdef’ that is appropriate for your system, and either #define any constants that configure should have defined but didn’t, or #undef any constants that configure defined and should not have. The custom.h file is automatically included by the config.h file.

It is also possible that the configure program generated by Autoconf will not work on your system in some other fashion. If you do have a problem, the configure.ac file is the input for Autoconf. You may be able to change this file and generate a new version of configure that works on your system (see Section B.4 [Reporting Problems and Bugs], page 442, for information on how to report problems in configuring gawk). The same mechanism may be used to send in updates to configure.ac and/or custom.h.

B.3 Installation on Other Operating Systems

This section describes how to install gawk on various non-Unix systems.

B.3.1 Installation on PC Operating Systems

This section covers installation and usage of gawk on Intel architecture machines running MS-DOS, any version of MS-Windows, or OS/2. In this section, the term “Windows32” refers to any of Microsoft Windows 95/98/ME/NT/2000/XP/Vista/7/8.

The limitations of MS-DOS (and MS-DOS shells under the other operating systems) have meant that various “DOS extenders” are often used with programs such as gawk. The varying capabilities of Microsoft Windows 3.1 and Windows32 can add to the confusion. For an overview of the considerations, refer to README_d/README.pc in the distribution.

B.3.1.1 Installing a Prepared Distribution for PC Systems

If you have received a binary distribution prepared by the MS-DOS maintainers, then gawk and the necessary support files appear under the gnu directory, with executables in gnu/bin, libraries in gnu/lib/awk, and manual pages under gnu/man. This is designed for easy installation to a /gnu directory on your drive—however, the files can be installed anywhere provided AWKPATH is set properly. Regardless of the installation directory, the first line of igawk.cmd and igawk.bat (in gnu/bin) may need to be edited.

The binary distribution contains a separate file describing the contents. In particular, it may include more than one version of the gawk executable.

OS/2 (32 bit, EMX) binary distributions are prepared for the /usr directory of your preferred drive. Set UNIXROOT to your installation drive (e.g., ‘e:’) if you want to install gawk onto another drive than the hardcoded default ‘c:’. Executables appear in /usr/bin, libraries under /usr/share/awk, manual pages under /usr/man, Texinfo documentation under /usr/info, and NLS files under /usr/share/locale. Note that the files can be installed anywhere provided AWKPATH is set properly.

If you already have a file /usr/info/dir from another package do not overwrite it! Instead enter the following commands at your prompt (replace ‘x:’ by your installation drive):

     install-info --info-dir=x:/usr/info x:/usr/info/gawk.info
     install-info --info-dir=x:/usr/info x:/usr/info/gawkinet.info

The binary distribution may contain a separate file containing additional or more detailed installation instructions.

B.3.1.2 Compiling gawk for PC Operating Systems

gawk can be compiled for MS-DOS, Windows32, and OS/2 using the GNU development tools from DJ Delorie (DJGPP: MS-DOS only), MinGW (Windows32) or Eberhard Mattes (EMX: MS-DOS, Windows32 and OS/2). The file README_d/README.pc in the gawk distribution contains additional notes, and pc/Makefile contains important information on compilation options.

To build gawk for MS-DOS and Windows32, copy the files in the pc directory (except for ChangeLog) to the directory with the rest of the gawk sources, then invoke make with the appropriate target name as an argument to build gawk. The Makefile copied from the pc directory contains a configuration section with comments and may need to be edited in order to work with your make utility.

The Makefile supports a number of targets for building various MS-DOS and Windows32 versions. A list of targets is printed if the make command is given without a target. As an example, to build gawk using the DJGPP tools, enter ‘make djgpp’. (The DJGPP

tools needed for the build may be found at ftp://ftp.delorie.com/pub/djgpp/current/

v2gnu/.) To build a native MS-Windows binary of gawk using the MinGW tools, type ‘make

mingw32’.

The 32 bit EMX version of gawk works “out of the box” under OS/2. However, it is highly recommended to use GCC 2.95.3 for the compilation. In principle, it is possible to compile gawk the following way:

$ ./configure
$ make

This is not recommended, though. To get an OMF executable you should use the following commands at your sh prompt:

$ CFLAGS="-O2 -Zomf -Zmt"
$ export CFLAGS
$ LDFLAGS="-s -Zcrtdll -Zlinker /exepack:2 -Zlinker /pm:vio -Zstack 0x6000"
$ export LDFLAGS
$ RANLIB="echo"
$ export RANLIB
$ ./configure --prefix=c:/usr
$ make AR=emxomfar

These are just suggestions for use with GCC 2.x. You may use any other set of (selfconsistent) environment variables and compiler flags.

If you use GCC 2.95 it is recommended to use also:

$ LIBS="-lgcc"
$ export LIBS

You can also get an a.out executable if you prefer:

$ CFLAGS="-O2 -Zmt"
$ export CFLAGS
$ LDFLAGS="-s -Zstack 0x6000"
$ LIBS="-lgcc"
$ unset RANLIB
$ ./configure --prefix=c:/usr
$ make

NOTE: Compilation of a.out executables also works with GCC 3.2. Versions later than GCC 3.2 have not been tested successfully.

‘make install’ works as expected with the EMX build.

NOTE: Ancient OS/2 ports of GNU make are not able to handle the Makefiles of this package. If you encounter any problems with make, try GNU Make 3.79.1 or later versions. You should find the latest version on ftp://hobbes.nmsu.

     edu/pub/os2/.^{1}

B.3.1.3 Testing gawk on PC Operating Systems

Using make to run the standard tests and to install gawk requires additional Unix-like tools, including sh, sed, and cp. In order to run the tests, the test/*.ok files may need to be converted so that they have the usual MS-DOS-style end-of-line markers. Alternatively, run make check CMP="diff -a" to use GNU diff in text mode instead of cmp to compare the resulting files.

Most of the tests work properly with Stewartson’s shell along with the companion utilities or appropriate GNU utilities. However, some editing of test/Makefile is required. It is recommended that you copy the file pc/Makefile.tst over the file test/Makefile as a replacement. Details can be found in README_d/README.pc and in the file pc/Makefile.tst.

On OS/2 the pid test fails because spawnl() is used instead of fork()/execl() to start child processes. Also the mbfw1 and mbprintf1 tests fail because the needed multibyte functionality is not available.

B.3.1.4 Using gawk on PC Operating Systems

Under MS-DOS and MS-Windows, the Cygwin and MinGW environments support both the ‘|&’ operator and TCP/IP networking (see Section 12.4 [Using gawk for Network Programming], page 307). EMX (OS/2 only) supports at least the ‘|&’ operator.

The MS-DOS and MS-Windows versions of gawk search for program files as described in Section 2.5.1 [The AWKPATH Environment Variable], page 39. However, semicolons (rather than colons) separate elements in the AWKPATH variable. If AWKPATH is not set or is empty, then the default search path is ‘.;c:/lib/awk;c:/gnu/lib/awk’.

The search path for OS/2 (32 bit, EMX) is determined by the prefix directory (most likely /usr or c:/usr) that has been specified as an option of the configure script as is the case for the Unix versions. If c:/usr is the prefix directory then the default search path contains . and c:/usr/share/awk. Additionally, to support binary distributions of gawk for OS/2 systems whose drive ‘c:’ might not support long file names or might not exist at all, there is a special environment variable. If UNIXROOT specifies a drive then this specific drive is also searched for program files. E.g., if UNIXROOT is set to e: the complete default

search path is ‘.;c:/usr/share/awk;e:/usr/share/awk’.

An sh-like shell (as opposed to command.com under MS-DOS or cmd.exe under MS-Windows or OS/2) may be useful for awk programming. The DJGPP collection of tools includes an MS-DOS port of Bash, and several shells are available for OS/2, including ksh.

Under MS-Windows, OS/2 and MS-DOS, gawk (and many other text programs) silently translates end-of-line ‘\r\n’ to ‘\n’ on input and ‘\n’ to ‘\r\n’ on output. A special BINMODE variable (c.e.) allows control over these translations and is interpreted as follows:

1 As of November 2014, this site is still there, but the author could not find a package for GNU Make.

• If BINMODE is "r" or one, then binary mode is set on read (i.e., no translations on reads).

• If BINMODE is "w" or two, then binary mode is set on write (i.e., no translations on writes).

• If BINMODE is "rw" or "wr" or three, binary mode is set for both read and write.

• BINMODE=non-null-string is the same as ‘BINMODE=3’ (i.e., no translations on reads or writes). However, gawk issues a warning message if the string is not one of "rw" or

    "wr".

The modes for standard input and standard output are set one time only (after the command line is read, but before processing any of the awk program). Setting BINMODE for standard input or standard output is accomplished by using an appropriate ‘-v BINMODE=N’ option on the command line. BINMODE is set at the time a file or pipe is opened and cannot be changed midstream.

The name BINMODE was chosen to match mawk (see Section B.5 [Other Freely Available awk Implementations], page 443). mawk and gawk handle BINMODE similarly; however, mawk adds a ‘-W BINMODE=N’ option and an environment variable that can set BINMODE, RS, and ORS. The files binmode[1-3].awk (under gnu/lib/awk in some of the prepared binary distributions) have been chosen to match mawk’s ‘-W BINMODE=N’ option. These can be changed or discarded; in particular, the setting of RS giving the fewest “surprises” is open to debate. mawk uses ‘RS = "\r\n"’ if binary mode is set on read, which is appropriate for files with the MS-DOS-style end-of-line.

To illustrate, the following examples set binary mode on writes for standard output and other files, and set ORS as the “usual” MS-DOS-style end-of-line:

     gawk -v BINMODE=2 -v ORS="\r\n" ...

or:

     gawk -v BINMODE=w -f binmode2.awk ...

These give the same result as the ‘-W BINMODE=2’ option in mawk. The following changes the record separator to "\r\n" and sets binary mode on reads, but does not affect the mode on standard input:

     gawk -v RS="\r\n" -e "BEGIN { BINMODE = 1 }" ...

or:

     gawk -f binmode1.awk ...

With proper quoting, in the first example the setting of RS can be moved into the BEGIN rule.

B.3.1.5 Using gawk In The Cygwin Environment

gawk can be built and used “out of the box” under MS-Windows if you are using the Cygwin environment (http://www.cygwin.com). This environment provides an excellent simulation of GNU/Linux, using Bash, GCC, GNU Make, and other GNU programs. Compilation and installation for Cygwin is the same as for a Unix system:

     tar -xvpzf gawk-4.1.2.tar.gz
     cd gawk-4.1.2
     ./configure

     make && make check

When compared to GNU/Linux on the same system, the ‘configure’ step on Cygwin takes considerably longer. However, it does finish, and then the ‘make’ proceeds as usual.

B.3.1.6 Using gawk In The MSYS Environment

In the MSYS environment under MS-Windows, gawk automatically uses binary mode for reading and writing files. Thus, there is no need to use the BINMODE variable.

This can cause problems with other Unix-like components that have been ported to MS-Windows that expect gawk to do automatic translation of "\r\n", because it won’t.

B.3.2 Compiling and Installing gawk on Vax/VMS and OpenVMS

This subsection describes how to compile and install gawk under VMS. The older designation “VMS” is used throughout to refer to OpenVMS.

B.3.2.1 Compiling gawk on VMS

To compile gawk under VMS, there is a DCL command procedure that issues all the necessary CC and LINK commands. There is also a Makefile for use with the MMS and MMK utilities. From the source directory, use either:

     $ @[.vms]vmsbuild.com

or:

     $ MMS/DESCRIPTION=[.vms]descrip.mms gawk

or:

     $ MMK/DESCRIPTION=[.vms]descrip.mms gawk

MMK is an open source, free, near-clone of MMS and can better handle ODS-5 volumes with upper- and lowercase file names. MMK is available from https://github.com/

endlesssoftware/mmk.

With ODS-5 volumes and extended parsing enabled, the case of the target parameter may need to be exact.

gawk has been tested under VAX/VMS 7.3 and Alpha/VMS 7.3-1 using Compaq C V6.4, and under Alpha/VMS 7.3, Alpha/VMS 7.3-2, and IA64/VMS 8.3. The most recent builds used HP C V7.3 on Alpha VMS 8.3 and both Alpha and IA64 VMS 8.4 used HP C 7.3.^{2}

See Section B.3.2.5 [The VMS GNV Project], page 442, for information on building gawk as a PCSI kit that is compatible with the GNV product.

B.3.2.2 Compiling gawk Dynamic Extensions on VMS

The extensions that have been ported to VMS can be built using one of the following commands:

     $ MMS/DESCRIPTION=[.vms]descrip.mms extensions

or:

     $ MMK/DESCRIPTION=[.vms]descrip.mms extensions

gawk uses AWKLIBPATH as either an environment variable or a logical name to find the dynamic extensions.

2 The IA64 architecture is also known as “Itanium.”

Dynamic extensions need to be compiled with the same compiler options for floatingpoint, pointer size, and symbol name handling as were used to compile gawk itself. Alpha and Itanium should use IEEE floating point. The pointer size is 32 bits, and the symbol name handling should be exact case with CRC shortening for symbols longer than 32 bits.

For Alpha and Itanium:

     /name=(as_is,short)
     /float=ieee/ieee_mode=denorm_results

For VAX:

     /name=(as_is,short)

Compile-time macros need to be defined before the first VMS-supplied header file is included, as follows:

     #if (__CRTL_VER >= 70200000) && !defined (__VAX)
     #define _LARGEFILE 1
     #endif
     #ifndef __VAX
     #ifdef __CRTL_VER
     #if __CRTL_VER >= 80200000
     #define _USE_STD_STAT 1
     #endif
     #endif
     #endif

If you are writing your own extensions to run on VMS, you must supply these definitions yourself. The config.h file created when building gawk on VMS does this for you; if instead you use that file or a similar one, then you must remember to include it before any VMSsupplied header files.

B.3.2.3 Installing gawk on VMS

To use gawk, all you need is a “foreign” command, which is a DCL symbol whose value begins with a dollar sign. For example:

     $ GAWK :== $disk1:[gnubin]gawk

Substitute the actual location of gawk.exe for ‘$disk1:[gnubin]’. The symbol should be placed in the login.com of any user who wants to run gawk, so that it is defined every time the user logs on. Alternatively, the symbol may be placed in the system-wide sylogin.com procedure, which allows all users to run gawk.

If your gawk was installed by a PCSI kit into the GNV$GNU: directory tree, the program

will be known as GNV$GNU:[bin]gnv$gawk.exe and the help file will be GNV$GNU:[vms_
help]gawk.hlp.

The PCSI kit also installs a GNV$GNU:[vms_bin]gawk_verb.cld file that can be used to add gawk and awk as DCL commands.

For just the current process you can use:

     $ set command gnv$gnu:[vms_bin]gawk_verb.cld
   Or the system manager can use GNV$GNU:[vms_bin]gawk_verb.cld to add the gawk

and awk to the system-wide ‘DCLTABLES’.

The DCL syntax is documented in the gawk.hlp file.

Optionally, the gawk.hlp entry can be loaded into a VMS help library:

     $ LIBRARY/HELP sys$help:helplib [.vms]gawk.hlp

(You may want to substitute a site-specific help library rather than the standard VMS library ‘HELPLIB’.) After loading the help text, the command:

     $ HELP GAWK

provides information about both the gawk implementation and the awk programming language.

The logical name ‘AWK_LIBRARY’ can designate a default location for awk program files. For the -f option, if the specified file name has no device or directory path information in it, gawk looks in the current directory first, then in the directory specified by the translation of ‘AWK_LIBRARY’ if the file is not found. If, after searching in both directories, the file still is not found, gawk appends the suffix ‘.awk’ to the file name and retries the file search. If ‘AWK_LIBRARY’ has no definition, a default value of ‘SYS$LIBRARY:’ is used for it.

B.3.2.4 Running gawk on VMS

Command-line parsing and quoting conventions are significantly different on VMS, so examples in this book or from other sources often need minor changes. They are minor though, and all awk programs should run correctly.

Here are a couple of trivial tests:

     $ gawk -- "BEGIN {print ""Hello, World!""}"
     $ gawk -"W" version
     ! could also be -"W version" or "-W version"

Note that uppercase and mixed-case text must be quoted.

The VMS port of gawk includes a DCL-style interface in addition to the original shell-style interface (see the help entry for details). One side effect of dual command-line parsing is that if there is only a single parameter (as in the quoted string program), the command becomes ambiguous. To work around this, the normally optional -- flag is required to force Unix-style parsing rather than DCL parsing. If any other dash-type options (or multiple parameters such as data files to process) are present, there is no ambiguity and -- can be omitted.

The exit value is a Unix-style value and is encoded into a VMS exit status value when the program exits.

The VMS severity bits will be set based on the exit value. A failure is indicated by 1, and VMS sets the ERROR status. A fatal error is indicated by 2, and VMS sets the FATAL status. All other values will have the SUCCESS status. The exit value is encoded to comply with VMS coding standards and will have the C_FACILITY_NO of 0x350000 with the constant 0xA000 added to the number shifted over by 3 bits to make room for the severity codes.

To extract the actual gawk exit code from the VMS status, use:

     unix_status = (vms_status .and. &x7f8) / 8

A C program that uses exec() to call gawk will get the original Unix-style exit value.

Older versions of gawk for VMS treated a Unix exit code 0 as 1, a failure as 2, a fatal error as 4, and passed all the other numbers through. This violated the VMS exit status coding requirements.

VAX/VMS floating point uses unbiased rounding. See Section 10.2.3 [Rounding Numbers], page 222.

VMS reports time values in GMT unless one of the SYS$TIMEZONE_RULE or TZ logical names is set. Older versions of VMS, such as VAX/VMS 7.3, do not set these logical names.

The default search path, when looking for awk program files specified by the -f option, is "SYS$DISK:[],AWK_LIBRARY:". The logical name AWKPATH can be used to override this default. The format of AWKPATH is a comma-separated list of directory specifications. When defining it, the value should be quoted so that it retains a single translation and not a multitranslation RMS searchlist.

B.3.2.5 The VMS GNV Project

The VMS GNV package provides a build environment similar to POSIX with ports of a collection of open source tools. The gawk found in the GNV base kit is an older port. Currently, the GNV project is being reorganized to supply individual PCSI packages for

each component. See https://sourceforge.net/p/gnv/wiki/InstallingGNVPackages/.

The normal build procedure for gawk produces a program that is suitable for use with GNV.

The file vms/gawk_build_steps.txt in the distribution documents the procedure for building a VMS PCSI kit that is compatible with GNV.

B.3.2.6 Some VMS Systems Have An Old Version of gawk

Some versions of VMS have an old version of gawk. To access it, define a symbol, as follows:

     $ gawk :== $sys$common:[syshlp.examples.tcpip.snmp]gawk.exe

This is apparently version 2.15.6, which is extremely old. We recommend compiling and using the current version.

B.4 Reporting Problems and Bugs

There is nothing more dangerous than a bored archaeologist.

—Douglas Adams, The Hitchhiker’s Guide to the Galaxy If you have problems with gawk or think that you have found a bug, report it to the

developers; we cannot promise to do anything, but we might well want to fix it.

Before reporting a bug, make sure you have really found a genuine bug. Carefully reread the documentation and see if it says you can do what you’re trying to do. If it’s not clear whether you should be able to do something or not, report that too; it’s a bug in the documentation!

Before reporting a bug or trying to fix it yourself, try to isolate it to the smallest possible awk program and input data file that reproduce the problem. Then send us the program and data file, some idea of what kind of Unix system you’re using, the compiler you used to compile gawk, and the exact results gawk gave you. Also say what you expected to occur; this helps us decide whether the problem is really in the documentation.

Make sure to include the version number of gawk you are using. You can get this information with the command ‘gawk --version’.

Once you have a precise problem description, send email to bug-gawk@gnu.org.

The gawk maintainers subscribe to this address, and thus they will receive your bug report. Although you can send mail to the maintainers directly, the bug reporting address is preferred because the email list is archived at the GNU Project. All email must be in English. This is the only language understood in common by all the maintainers.

CAUTION: Do not try to report bugs in gawk by posting to the Usenet/Internet newsgroup comp.lang.awk. The gawk developers do occasionally read this newsgroup, but there is no guarantee that we will see your posting. The steps described here are the only officially recognized way for reporting bugs. Really. NOTE: Many distributions of GNU/Linux and the various BSD-based operating systems have their own bug reporting systems. If you report a bug using your distribution’s bug reporting system, you should also send a copy to

     bug-gawk@gnu.org.

This is for two reasons. First, although some distributions forward bug reports “upstream” to the GNU mailing list, many don’t, so there is a good chance that the gawk maintainers won’t even see the bug report! Second, mail to the GNU list is archived, and having everything at the GNU Project keeps things self-contained and not dependent on other organizations.

Non-bug suggestions are always welcome as well. If you have questions about things that are unclear in the documentation or are just obscure features, ask on the bug list; we will try to help you out if we can.

If you find bugs in one of the non-Unix ports of gawk, send an email to the bug list, with a copy to the person who maintains that port. The maintainers are named in the following list, as well as in the README file in the gawk distribution. Information in the README file should be considered authoritative if it conflicts with this book.

The people maintaining the various gawk ports are:

Unix and POSIX systems  Arnold Robbins, arnold@skeeve.com
MS-DOS with DJGPP  Scott Deifik, scottd.mail@sbcglobal.net
MS-Windows with MinGW Eli Zaretskii, eliz@gnu.org

OS/2  Andreas Buening, andreas.buening@nexgo.de

VMS  John Malmberg, wb8tyw@qsl.net
z/OS (OS/390)  Dave Pitts, dpitts@cozx.com

If your bug is also reproducible under Unix, send a copy of your report to the

bug-gawk@gnu.org email list as well.

B.5 Other Freely Available awk Implementations

It’s kind of fun to put comments like this in your awk code:

         // Do C++ comments work? answer: yes! of course

—Michael Brennan There are a number of other freely available awk implementations. This section briefly

describes where to get them:

Unix awk Brian Kernighan, one of the original designers of Unix awk, has made his implementation of awk freely available. You can retrieve this version via his home page (http://www.cs.princeton.edu/~bwk). It is available in several archive formats:

Shell archive

          http://www.cs.princeton.edu/~bwk/btl.mirror/awk.shar

Compressed tar file

          http://www.cs.princeton.edu/~bwk/btl.mirror/awk.tar.gz
Zip file  http://www.cs.princeton.edu/~bwk/btl.mirror/awk.zip

You can also retrieve it from GitHub:

     git clone git://github.com/onetrueawk/awk bwkawk

This command creates a copy of the Git (http://git-scm.com) repository in a directory named bwkawk. If you leave that argument off the git command line, the repository copy is created in a directory named awk. This version requires an ISO C (1990 standard) compiler; the C compiler from GCC (the GNU Compiler Collection) works quite nicely. See Section A.7 [Common Extensions Summary], page 423, for a list of extensions in this awk that are not in POSIX awk. As a side note, Dan Bornstein has created a Git repository tracking all the versions of BWK awk that he could find. It’s available at git://github.com/

danfuzz/one-true-awk.

mawk Michael Brennan wrote an independent implementation of awk, called mawk. It is available under the GPL (see [GNU General Public License], page 475), just as gawk is. The original distribution site for the mawk source code no longer has it. A copy

is available at http://www.skeeve.com/gawk/mawk1.3.3.tar.gz.

In 2009, Thomas Dickey took on mawk maintenance. Basic information is avail-

able on the project’s web page (http://www.invisible-island.net/mawk).
The download URL is http://invisible-island.net/datafiles/release/
mawk.tar.gz.

Once you have it, gunzip may be used to decompress this file. Installation is similar to gawk’s (see Section B.2 [Compiling and Installing gawk on Unix-Like Systems], page 433). See Section A.7 [Common Extensions Summary], page 423, for a list of extensions in mawk that are not in POSIX awk.

awka Written by Andrew Sumner, awka translates awk programs into C, compiles them, and links them with a library of functions that provide the core awk functionality. It also has a number of extensions.

The awk translator is released under the GPL, and the library is under the LGPL.

To get awka, go to http://sourceforge.net/projects/awka.

The project seems to be frozen; no new code changes have been made since approximately 2001.

pawk Nelson H.F. Beebe at the University of Utah has modified BWK awk to provide timing and profiling information. It is different from gawk with the --profile option (see Section 12.5 [Profiling Your awk Programs], page 308) in that it uses CPU-based profiling, not line-count profiling. You may find it at either

ftp://ftp.math.utah.edu/pub/pawk/pawk-20030606.tar.gz or http://
www.math.utah.edu/pub/pawk/pawk-20030606.tar.gz.

BusyBox awk BusyBox is a GPL-licensed program providing small versions of many applications within a single executable. It is aimed at embedded systems. It includes a full implementation of POSIX awk. When building it, be careful not to do ‘make install’ as it will overwrite copies of other applications in your /usr/local/bin. For more information, see the project’s home page (http://

busybox.net).

The OpenSolaris POSIX awk The versions of awk in /usr/xpg4/bin and /usr/xpg6/bin on Solaris are more or less POSIX-compliant. They are based on the awk from Mortice Kern Systems for PCs. We were able to make this code compile and work under GNU/Linux with 1–2 hours of work. Making it more generally portable (using GNU Autoconf and/or Automake) would take more work, and this has not been done, at least to our knowledge. The source code used to be available from the OpenSolaris website. However, that project was ended and the website shut down. Fortunately, the Illu-

mos project (http://wiki.illumos.org/display/illumos/illumos+Home)

makes this implementation available. You can view the files one at a time from

https://github.com/joyent/illumos-joyent/blob/master/usr/src/cmd/
awk_xpg4.

jawk This is an interpreter for awk written in Java. It claims to be a full interpreter, although because it uses Java facilities for I/O and for regexp matching, the language it supports is different from POSIX awk. More information is available

on the project’s home page (http://jawk.sourceforge.net).

Libmawk This is an embeddable awk interpreter derived from mawk. For more information,

see http://repo.hu/projects/libmawk/.

pawk This is a Python module that claims to bring awk-like features to Python. See

https://github.com/alecthomas/pawk for more information. (This is not

related to Nelson Beebe’s modified version of BWK awk, described earlier.)

QSE awk This is an embeddable awk interpreter. For more information, see http://

code.google.com/p/qse/ and http://awk.info/?tools/qse.

QTawk This is an independent implementation of awk distributed under the GPL. It has a large number of extensions over standard awk and may not be 100%

syntactically compatible with it. See http://www.quiktrim.org/QTawk.html

for more information, including the manual and a download link. The project may also be frozen; no new code changes have been made since approximately 2008.

Other versions See also the “Versions and implementations” section of the

Wikipedia article (http://en.wikipedia.org/wiki/Awk_language#
Versions_and_implementations) on awk for information on additional

versions.

B.6 Summary

• The gawk distribution is available from the GNU Project’s main distribution site,

ftp.gnu.org. The canonical build recipe is:

wget http://ftp.gnu.org/gnu/gawk/gawk-4.1.2.tar.gz
tar -xvpzf gawk-4.1.2.tar.gz
cd gawk-4.1.2
./configure && make && make check

• gawk may be built on non-POSIX systems as well. The currently supported systems are MS-Windows using DJGPP, MSYS, MinGW, and Cygwin, OS/2 using EMX, and both Vax/VMS and OpenVMS. Instructions for each system are included in this appendix.

• Bug reports should be sent via email to bug-gawk@gnu.org. Bug reports should be in English and should include the version of gawk, how it was compiled, and a short program and data file that demonstrate the problem.

• There are a number of other freely available awk implementations. Many are POSIXcompliant; others are less so.

Appendix C: Implementation Notes 447

Appendix C Implementation Notes

This appendix contains information mainly of interest to implementers and maintainers of gawk. Everything in it applies specifically to gawk and not to other implementations.

C.1 Downward Compatibility and Debugging

See Section A.5 [Extensions in gawk Not in POSIX awk], page 415, for a summary of the GNU extensions to the awk language and program. All of these features can be turned off by invoking gawk with the --traditional option or with the --posix option.

If gawk is compiled for debugging with ‘-DDEBUG’, then there is one more option available on the command line:

-Y
--parsedebug

Print out the parse stack information as the program is being parsed.

This option is intended only for serious gawk developers and not for the casual user. It probably has not even been compiled into your version of gawk, since it slows down execution.

C.2 Making Additions to gawk

If you find that you want to enhance gawk in a significant fashion, you are perfectly free to do so. That is the point of having free software; the source code is available and you are free to change it as you want (see [GNU General Public License], page 475).

This section discusses the ways you might want to change gawk as well as any considerations you should bear in mind.

C.2.1 Accessing The gawk Git Repository

As gawk is Free Software, the source code is always available. Section B.1 [The gawk Distribution], page 429, describes how to get and build the formal, released versions of

gawk.

However, if you want to modify gawk and contribute back your changes, you will probably wish to work with the development version. To do so, you will need to access the gawk source code repository. The code is maintained using the Git distributed version control system (http://git-scm.com). You will need to install it if your system doesn’t have it. Once you have done so, use the command:

     git clone git://git.savannah.gnu.org/gawk.git

This clones the gawk repository. If you are behind a firewall that does not allow you to use the Git native protocol, you can still access the repository using:

     git clone http://git.savannah.gnu.org/r/gawk.git

Once you have made changes, you can use ‘git diff’ to produce a patch, and send that to the gawk maintainer; see Section B.4 [Reporting Problems and Bugs], page 442, for how to do that.

Once upon a time there was Git–CVS gateway for use by people who could not install Git. However, this gateway no longer works, so you may have better luck using a more modern version control system like Bazaar, that has a Git plug-in for working with Git repositories.

C.2.2 Adding New Features

You are free to add any new features you like to gawk. However, if you want your changes to be incorporated into the gawk distribution, there are several steps that you need to take in order to make it possible to include them:

1. Before building the new feature into gawk itself, consider writing it as an extension (see Chapter 16 [Writing Extensions for gawk], page 355). If that’s not possible, continue with the rest of the steps in this list.

2. Be prepared to sign the appropriate paperwork. In order for the FSF to distribute your changes, you must either place those changes in the public domain and submit a signed statement to that effect, or assign the copyright in your changes to the FSF. Both of these actions are easy to do and many people have done so already. If you have questions, please contact me (see Section B.4 [Reporting Problems and Bugs],

page 442), or assign@gnu.org.

3. Get the latest version. It is much easier for me to integrate changes if they are relative to the most recent distributed version of gawk, or better yet, relative to the latest code in the Git repository. If your version of gawk is very old, I may not be able to integrate your changes at all. (See Section B.1.1 [Getting the gawk Distribution], page 429, for information on getting the latest version of gawk.)

4. Follow the GNU Coding Standards. This document describes how GNU software should be written. If you haven’t read it, please do so, preferably before starting to modify gawk. (The GNU Coding Standards are available from the GNU Project’s website (http://www.gnu.org/prep/standards/). Texinfo, Info, and DVI versions are also available.)

5. Use the gawk coding style. The C code for gawk follows the instructions in the GNU Coding Standards, with minor exceptions. The code is formatted using the traditional “K&R” style, particularly as regards to the placement of braces and the use of TABs. In brief, the coding rules for gawk are as follows:

• Use ANSI/ISO style (prototype) function headers when defining functions.

• Put the name of the function at the beginning of its own line.

• Put the return type of the function, even if it is int, on the line above the line with the name and arguments of the function.

• Put spaces around parentheses used in control structures (if, while, for, do,

    switch, and return).

• Do not put spaces in front of parentheses used in function calls.

• Put spaces around all C operators and after commas in function calls.

• Do not use the comma operator to produce multiple side effects, except in for loop initialization and increment parts, and in macro bodies.

• Use real TABs for indenting, not spaces.

• Use the “K&R” brace layout style.

• Use comparisons against NULL and '\0' in the conditions of if, while, and for statements, as well as in the cases of switch statements, instead of just the plain pointer or character value.

• Use true and false for bool values, the NULL symbolic constant for pointer values, and the character constant '\0' where appropriate, instead of 1 and 0.

• Provide one-line descriptive comments for each function.

• Do not use the alloca() function for allocating memory off the stack. Its use causes more portability trouble than is worth the minor benefit of not having to free the storage. Instead, use malloc() and free().

• Do not use comparisons of the form ‘! strcmp(a, b)’ or similar. As Henry Spencer once said, “strcmp() is not a boolean!” Instead, use ‘strcmp(a, b) == 0’.

• If adding new bit flag values, use explicit hexadecimal constants (0x001, 0x002, 0x004, and son on) instead of shifting one left by successive amounts (‘(1<<0)’, ‘(1<<1)’, and so on).

NOTE: If I have to reformat your code to follow the coding style used in gawk, I may not bother to integrate your changes at all.

6. Update the documentation. Along with your new code, please supply new sections and/or chapters for this book. If at all possible, please use real Texinfo, instead of just supplying unformatted ASCII text (although even that is better than no documentation at all). Conventions to be followed in GAWK: Effective AWK Programming are provided after the ‘@bye’ at the end of the Texinfo source file. If possible, please update the man page as well. You will also have to sign paperwork for your documentation changes.

7. Submit changes as unified diffs. Use ‘diff -u -r -N’ to compare the original gawk source tree with your version. I recommend using the GNU version of diff, or best of all, ‘git diff’ or ‘git format-patch’. Send the output produced by diff to me when you submit your changes. (See Section B.4 [Reporting Problems and Bugs], page 442, for the electronic mail information.) Using this format makes it easy for me to apply your changes to the master version of the gawk source code (using patch). If I have to apply the changes manually, using a text editor, I may not do so, particularly if there are lots of changes.

8. Include an entry for the ChangeLog file with your submission. This helps further minimize the amount of work I have to do, making it easier for me to accept patches. It is simplest if you just make this part of your diff.

Although this sounds like a lot of work, please remember that while you may write the new code, I have to maintain it and support it. If it isn’t possible for me to do that with a minimum of extra work, then I probably will not.

C.2.3 Porting gawk to a New Operating System

If you want to port gawk to a new operating system, there are several steps:

1. Follow the guidelines in the previous section concerning coding style, submission of diffs, and so on.

2. Be prepared to sign the appropriate paperwork. In order for the FSF to distribute your code, you must either place your code in the public domain and submit a signed statement to that effect, or assign the copyright in your code to the FSF. Both of these actions are easy to do and many people have done so already. If you have questions, please contact me, or gnu@gnu.org.

3. When doing a port, bear in mind that your code must coexist peacefully with the rest of gawk and the other ports. Avoid gratuitous changes to the system-independent parts of the code. If at all possible, avoid sprinkling ‘#ifdef’s just for your port throughout the code.

If the changes needed for a particular system affect too much of the code, I probably will not accept them. In such a case, you can, of course, distribute your changes on your own, as long as you comply with the GPL (see [GNU General Public License], page 475).

4. A number of the files that come with gawk are maintained by other people. Thus, you should not change them unless it is for a very good reason; i.e., changes are not out of the question, but changes to these files are scrutinized extra carefully.

The files are dfa.c, dfa.h, getopt.c, getopt.h, getopt1.c, getopt_int.h,
gettext.h, regcomp.c, regex.c, regex.h, regex_internal.c, regex_internal.h,
and regexec.c.

5. A number of other files are provided by the GNU Autotools (Autoconf, Automake, and GNU gettext). You should not change them either, unless it is for a very good

reason. The files are ABOUT-NLS, config.guess, config.rpath, config.sub, depcomp,
INSTALL, install-sh, missing, mkinstalldirs, xalloc.h, and ylwrap.

6. Be willing to continue to maintain the port. Non-Unix operating systems are supported by volunteers who maintain the code needed to compile and run gawk on their systems. If no-one volunteers to maintain a port, it becomes unsupported and it may be necessary to remove it from the distribution.

7. Supply an appropriate gawkmisc.??? file. Each port has its own gawkmisc.??? that implements certain operating system specific functions. This is cleaner than a plethora of ‘#ifdef’s scattered throughout the code. The gawkmisc.c in the main source directory includes the appropriate gawkmisc.??? file from each subdirectory. Be sure to update it as well.

Each port’s gawkmisc.??? file has a suffix reminiscent of the machine or operating system for the port—for example, pc/gawkmisc.pc and vms/gawkmisc.vms. The use of separate suffixes, instead of plain gawkmisc.c, makes it possible to move files from a port’s subdirectory into the main subdirectory, without accidentally destroying the real gawkmisc.c file. (Currently, this is only an issue for the PC operating system ports.)

8. Supply a Makefile as well as any other C source and header files that are necessary for your operating system. All your code should be in a separate subdirectory, with a name that is the same as, or reminiscent of, either your operating system or the computer system. If possible, try to structure things so that it is not necessary to move files out of the subdirectory into the main source directory. If that is not possible, then be sure to avoid using names for your files that duplicate the names of files in the main source directory.

9. Update the documentation. Please write a section (or sections) for this book describing the installation and compilation steps needed to compile and/or install gawk for your system.

Following these steps makes it much easier to integrate your changes into gawk and have them coexist happily with other operating systems’ code that is already there.

In the code that you supply and maintain, feel free to use a coding style and brace layout that suits your taste.

C.2.4 Why Generated Files Are Kept In Git

If you look at the gawk source in the Git repository, you will notice that it includes files that are automatically generated by GNU infrastructure tools, such as Makefile.in from Automake and even configure from Autoconf.

This is different from many Free Software projects that do not store the derived files, because that keeps the repository less cluttered, and it is easier to see the substantive changes when comparing versions and trying to understand what changed between commits.

However, there are several reasons why the gawk maintainer likes to have everything in the repository.

First, because it is then easy to reproduce any given version completely, without relying upon the availability of (older, likely obsolete, and maybe even impossible to find) other tools.

As an extreme example, if you ever even think about trying to compile, oh, say, the V7 awk, you will discover that not only do you have to bootstrap the V7 yacc to do so, but you also need the V7 lex. And the latter is pretty much impossible to bring up on a modern GNU/Linux system.^{1}

(Or, let’s say gawk 1.2 required bison whatever-it-was in 1989 and that there was no awkgram.c file in the repository. Is there a guarantee that we could find that bison version? Or that it would build?)

If the repository has all the generated files, then it’s easy to just check them out and build. (Or easier, depending upon how far back we go.)

And that brings us to the second (and stronger) reason why all the files really need to be in Git. It boils down to who do you cater to—the gawk developer(s), or the user who just wants to check out a version and try it out?

The gawk maintainer wants it to be possible for any interested awk user in the world to just clone the repository, check out the branch of interest and build it. Without their having to have the correct version(s) of the autotools.^{2} That is the point of the bootstrap.sh file. It touches the various other files in the right order such that

     # The canonical incantation for building GNU software:
     ./bootstrap.sh && ./configure && make

1 We tried. It was painful.

2 There is one GNU program that is (in our opinion) severely difficult to bootstrap from the Git repository.

For example, on the author’s old (but still working) PowerPC Macintosh with Mac OS X 10.5, it was

necessary to bootstrap a ton of software, starting with Git itself, in order to try to work with the latest

code. It’s not pleasant, and especially on older systems, it’s a big waste of time.

Starting with the latest tarball was no picnic either. The maintainers had dropped .gz and .bz2 files

and only distribute .tar.xz files. It was necessary to bootstrap xz first!

will just work.

This is extremely important for the master and gawk-X.Y-stable branches.

Further, the gawk maintainer would argue that it’s also important for the gawk developers. When he tried to check out the xgawk branch^{3} to build it, he couldn’t. (No ltmain.sh file, and he had no idea how to create it, and that was not the only problem.)

He felt extremely frustrated. With respect to that branch, the maintainer is no different than Jane User who wants to try to build gawk-4.1-stable or master from the repository.

Thus, the maintainer thinks that it’s not just important, but critical, that for any given branch, the above incantation just works.

A third reason to have all the files is that without them, using ‘git bisect’ to try to find the commit that introduced a bug is exceedingly difficult. The maintainer tried to do that on another project that requires running bootstrapping scripts just to create configure and so on; it was really painful. When the repository is self-contained, using git bisect in it is very easy.

What are some of the consequences and/or actions to take?

1. We don’t mind that there are differing files in the different branches as a result of different versions of the autotools. A. It’s the maintainer’s job to merge them and he will deal with it.

    B. He is really good at ‘git diff x y > /tmp/diff1 ; gvim /tmp/diff1’ to remove

the diffs that aren’t of interest in order to review code.

2. It would certainly help if everyone used the same versions of the GNU tools as he does, which in general are the latest released versions of Automake, Autoconf, bison, and

    GNU gettext.

Installing from source is quite easy. It’s how the maintainer worked for years (and still works). He had /usr/local/bin at the front of his PATH and just did:

         wget http://ftp.gnu.org/gnu/package/package-x.y.z.tar.gz
         tar -xpzvf package-x.y.z.tar.gz
         cd package-x.y.z
         ./configure && make && make check
         make install    # as root

Most of the above was originally written by the maintainer to other gawk developers. It raised the objection from one of the developers “… that anybody pulling down the source from Git is not an end user.”

However, this is not true. There are “power awk users” who can build gawk (using the magic incantation shown previously) but who can’t program in C. Thus, the major branches should be kept buildable all the time.

It was then suggested that there be a cron job to create nightly tarballs of “the source.” Here, the problem is that there are source trees, corresponding to the various branches! So, nightly tarballs aren’t the answer, especially as the repository can go for weeks without significant change being introduced.

Fortunately, the Git server can meet this need. For any given branch named branchname, use:

3 A branch (since removed) created by one of the other developers that did not include the generated files.

     wget http://git.savannah.gnu.org/cgit/gawk.git/snapshot/gawk-branchname.tar.gz

to retrieve a snapshot of the given branch.

C.3 Probable Future Extensions

AWK is a language similar to PERL, only considerably more elegant.

—Arnold Robbins Hey!

—Larry Wall The TODO file in the master branch of the gawk Git repository lists possible future

enhancements. Some of these relate to the source code, and others to possible new features. Please see that file for the list. See Section C.2 [Making Additions to gawk], page 447, if you are interested in tackling any of the projects listed there.

C.4 Some Limitations of the Implementation

This following table describes limits of gawk on a Unix-like system (although it is variable even then). Other systems may have different limits.

Item  Limit
Characters in a character class  2^(number of bits per byte)
Length of input record  MAX_INT
Length of output record  Unlimited
Length of source line  Unlimited
Number of fields in a record  MAX_LONG
Number of file redirections  Unlimited
Number of input records in one file MAX_LONG
Number of input records total  MAX_LONG

Number of pipe redirections min(number of processes per user, number of open files)

Numeric values  Double-precision floating point (if not using MPFR)
Size of a field  MAX_INT
Size of a literal string  MAX_INT
Size of a printf string  MAX_INT

C.5 Extension API Design

This section documents the design of the extension API, including a discussion of some of the history and problems that needed to be solved.

The first version of extensions for gawk was developed in the mid-1990s and released with gawk 3.1 in the late 1990s. The basic mechanisms and design remained unchanged for close to 15 years, until 2012.

The old extension mechanism used data types and functions from gawk itself, with a “clever hack” to install extension functions.

gawk included some sample extensions, of which a few were really useful. However, it was clear from the outset that the extension mechanism was bolted onto the side and was not really well thought out.

C.5.1 Problems With The Old Mechanism

The old extension mechanism had several problems:

• It depended heavily upon gawk internals. Any time the NODE structure^{4} changed, an extension would have to be recompiled. Furthermore, to really write extensions required understanding something about gawk’s internal functions. There was some documentation in this book, but it was quite minimal.

• Being able to call into gawk from an extension required linker facilities that are common on Unix-derived systems but that did not work on MS-Windows systems; users wanting extensions on MS-Windows had to statically link them into gawk, even though MS-Windows supports dynamic loading of shared objects.

• The API would change occasionally as gawk changed; no compatibility between versions was ever offered or planned for.

Despite the drawbacks, the xgawk project developers forked gawk and developed several significant extensions. They also enhanced gawk’s facilities relating to file inclusion and shared object access.

A new API was desired for a long time, but only in 2012 did the gawk maintainer and the xgawk developers finally start working on it together. More information about the xgawk project is provided in Section 16.8 [The gawkextlib Project], page 407.

C.5.2 Goals For A New Mechanism

Some goals for the new API were:

• The API should be independent of gawk internals. Changes in gawk internals should not be visible to the writer of an extension function.

• The API should provide binary compatibility across gawk releases as long as the API itself does not change.

• The API should enable extensions written in C or C++ to have roughly the same “appearance” to awk-level code as awk functions do. This means that extensions should have: − The ability to access function parameters. − The ability to turn an undefined parameter into an array (call by reference). − The ability to create, access and update global variables. − Easy access to all the elements of an array at once (“array flattening”) in order to

loop over all the element in an easy fashion for C code. − The ability to create arrays (including gawk’s true arrays of arrays).

Some additional important goals were:

• The API should use only features in ISO C 90, so that extensions can be written using the widest range of C and C++ compilers. The header should include the appropriate ‘#ifdef __cplusplus’ and ‘extern "C"’ magic so that a C++ compiler could be used. (If using C++, the runtime system has to be smart enough to call any constructors and destructors, as gawk is a C program. As of this writing, this has not been tested.)

4 A critical central data structure inside gawk.

• The API mechanism should not require access to gawk’s symbols^{5} by the compiletime or dynamic linker, in order to enable creation of extensions that also work on MS-Windows.

During development, it became clear that there were other features that should be available to extensions, which were also subsequently provided:

• Extensions should have the ability to hook into gawk’s I/O redirection mechanism. In particular, the xgawk developers provided a so-called “open hook” to take over reading records. During development, this was generalized to allow extensions to hook into input processing, output processing, and two-way I/O.

• An extension should be able to provide a “call back” function to perform cleanup actions when gawk exits.

• An extension should be able to provide a version string so that gawk’s --version option can provide information about extensions as well.

The requirement to avoid access to gawk’s symbols is, at first glance, a difficult one to meet.

One design, apparently used by Perl and Ruby and maybe others, would be to make the mainline gawk code into a library, with the gawk utility a small C main() function linked against the library.

This seemed like the tail wagging the dog, complicating build and installation and making a simple copy of the gawk executable from one system to another (or one place to another on the same system!) into a chancy operation.

Pat Rankin suggested the solution that was adopted. See Section 16.3 [How It Works at a High Level], page 355, for the details.

C.5.3 Other Design Decisions

As an arbitrary design decision, extensions can read the values of predefined variables and arrays (such as ARGV and FS), but cannot change them, with the exception of PROCINFO.

The reason for this is to prevent an extension function from affecting the flow of an awk program outside its control. While a real awk function can do what it likes, that is at the discretion of the programmer. An extension function should provide a service or make a C API available for use within awk, and not mess with FS or ARGC and ARGV.

In addition, it becomes easy to start down a slippery slope. How much access to gawk facilities do extensions need? Do they need getline? What about calling gsub() or compiling regular expressions? What about calling into awk functions? (That would be messy.)

In order to avoid these issues, the gawk developers chose to start with the simplest, most basic features that are still truly useful.

Another decision is that although gawk provides nice things like MPFR, and arrays indexed internally by integers, these features are not being brought out to the API in order to keep things simple and close to traditional awk semantics. (In fact, arrays indexed internally by integers are so transparent that they aren’t even documented!)

5 The symbols are the variables and functions defined inside gawk. Access to these symbols by code

external to gawk loaded dynamically at runtime is problematic on MS-Windows.

Additionally, all functions in the API check that their pointer input parameters are not NULL. If they are, they return an error. (It is a good idea for extension code to verify that pointers received from gawk are not NULL. Such a thing should not happen, but the gawk developers are only human, and they have been known to occasionally make mistakes.)

With time, the API will undoubtedly evolve; the gawk developers expect this to be driven by user needs. For now, the current API seems to provide a minimal yet powerful set of features for creating extensions.

C.5.4 Room For Future Growth

The API can later be expanded, in two ways:

• gawk passes an “extension id” into the extension when it first loads the extension. The extension then passes this id back to gawk with each function call. This mechanism allows gawk to identify the extension calling into it, should it need to know.

• Similarly, the extension passes a “name space” into gawk when it registers each extension function. This accommodates a possible future mechanism for grouping extension functions and possibly avoiding name conflicts.

Of course, as of this writing, no decisions have been made with respect to any of the above.

C.6 Compatibility For Old Extensions

Chapter 16 [Writing Extensions for gawk], page 355, describes the supported API and mechanisms for writing extensions for gawk. This API was introduced in version 4.1. However, for many years gawk provided an extension mechanism that required knowledge of gawk internals and that was not as well designed.

In order to provide a transition period, gawk version 4.1 continues to support the original extension mechanism. This will be true for the life of exactly one major release. This support will be withdrawn, and removed from the source code, at the next major release.

Briefly, original-style extensions should be compiled by including the awk.h header file in the extension source code. Additionally, you must define the identifier ‘GAWK’ when building (use ‘-DGAWK’ with Unix-style compilers). Otherwise, the definitions in gawkapi.h will cause conflicts with those in awk.h and your extension will not compile.

Just as in previous versions, you load an old-style extension with the extension() builtin function (which is not otherwise documented). This function in turn finds and loads the shared object file containing the extension and calls its dl_load() C routine.

Because original-style and new-style extensions use different initialization routines (dl_ load() versus dlload()), they may safely be installed in the same directory (to be found by AWKLIBPATH) without conflict.

The gawk development team strongly recommends that you convert any old extensions that you may have to use the new API described in Chapter 16 [Writing Extensions for gawk], page 355.

C.7 Summary

• gawk’s extensions can be disabled with either the --traditional option or with the --posix option. The --parsedebug option is available if gawk is compiled with

‘-DDEBUG’.

• The source code for gawk is maintained in a publicly accessible Git repository. Anyone may check it out and view the source.

• Contributions to gawk are welcome. Following the steps outlined in this chapter will make it easier to integrate your contributions into the code base. This applies both to new feature contributions and to ports to additional operating systems.

• gawk has some limits—generally those that are imposed by the machine architecture.

• The extension API design was intended to solve a number of problems with the previous extension mechanism, enable features needed by the xgawk project, and provide binary compatibility going forward.

• The previous extension mechanism is still supported in version 4.1 of gawk, but it will be removed in the next major release.

Appendix D: Basic Programming Concepts 459

Appendix D Basic Programming Concepts

This appendix attempts to define some of the basic concepts and terms that are used throughout the rest of this book. As this book is specifically about awk, and not about computer programming in general, the coverage here is by necessity fairly cursory and simplistic. (If you need more background, there are many other introductory texts that you should refer to instead.)

D.1 What a Program Does

At the most basic level, the job of a program is to process some input data and produce results. See Figure D.1.

Figure D.1: General Program Flow

The “program” in the figure can be either a compiled program^{1} (such as ls), or it may be interpreted. In the latter case, a machine-executable program such as awk reads your program, and then uses the instructions in your program to process the data.

When you write a program, it usually consists of the following, very basic set of steps, as shown in Figure D.2:

Figure D.2: Basic Program Steps Initialization

1 Compiled programs are typically written in lower-level languages such as C, C++, or Ada, and then

translated, or compiled, into a form that the computer can execute directly.

These are the things you do before actually starting to process data, such as checking arguments, initializing any data you need to work with, and so on. This step corresponds to awk’s BEGIN rule (see Section 7.1.4 [The BEGIN and END Special Patterns], page 134). If you were baking a cake, this might consist of laying out all the mixing bowls and the baking pan, and making sure you have all the ingredients that you need.

Processing This is where the actual work is done. Your program reads data, one logical chunk at a time, and processes it as appropriate. In most programming languages, you have to manually manage the reading of data, checking to see if there is more each time you read a chunk. awk’s pattern-action paradigm (see Chapter 1 [Getting Started with awk], page 17) handles the mechanics of this for you. In baking a cake, the processing corresponds to the actual labor: breaking eggs, mixing the flour, water, and other ingredients, and then putting the cake into the oven.

Clean Up Once you’ve processed all the data, you may have things you need to do before exiting. This step corresponds to awk’s END rule (see Section 7.1.4 [The BEGIN and END Special Patterns], page 134). After the cake comes out of the oven, you still have to wrap it in plastic wrap to keep anyone from tasting it, as well as wash the mixing bowls and utensils.

An algorithm is a detailed set of instructions necessary to accomplish a task, or process data. It is much the same as a recipe for baking a cake. Programs implement algorithms. Often, it is up to you to design the algorithm and implement it, simultaneously.

The “logical chunks” we talked about previously are called records, similar to the records a company keeps on employees, a school keeps for students, or a doctor keeps for patients. Each record has many component parts, such as first and last names, date of birth, address, and so on. The component parts are referred to as the fields of the record.

The act of reading data is termed input, and that of generating results, not too surprisingly, is termed output. They are often referred to together as “input/output,” and even more often, as “I/O” for short. (You will also see “input” and “output” used as verbs.)

awk manages the reading of data for you, as well as the breaking it up into records and fields. Your program’s job is to tell awk what to do with the data. You do this by describing patterns in the data to look for, and actions to execute when those patterns are seen. This data-driven nature of awk programs usually makes them both easier to write and easier to read.

D.2 Data Values in a Computer

In a program, you keep track of information and values in things called variables. A variable is just a name for a given value, such as first_name, last_name, address, and so on. awk has several predefined variables, and it has special names to refer to the current input record and the fields of the record. You may also group multiple associated values under one name, as an array.

Data, particularly in awk, consists of either numeric values, such as 42 or 3.1415927, or string values. String values are essentially anything that’s not a number, such as a name. Strings are sometimes referred to as character data, since they store the individual characters that comprise them. Individual variables, as well as numeric and string variables, are referred to as scalar values. Groups of values, such as arrays, are not scalars.

Section 15.1 [A General Description of Computer Arithmetic], page 341, provided a basic introduction to numeric types (integer and floating-point) and how they are used in a computer. Please review that information, including a number of caveats that were presented.

While you are probably used to the idea of a number without a value (i.e., zero), it takes a bit more getting used to the idea of zero-length character data. Nevertheless, such a thing exists. It is called the null string. The null string is character data that has no value. In other words, it is empty. It is written in awk programs like this: "".

Humans are used to working in decimal; i.e., base 10. In base 10, numbers go from 0 to 9, and then “roll over” into the next column. (Remember grade school? 42 = 4 x 10 + 2.)

There are other number bases though. Computers commonly use base 2 or binary, base 8 or octal, and base 16 or hexadecimal. In binary, each column represents two times the value in the column to its right. Each column may contain either a 0 or a 1. Thus, binary 1010 represents (1 x 8) + (0 x 4) + (1 x 2) + (0 x 1), or decimal 10. Octal and hexadecimal are discussed more in Section 6.1.1.2 [Octal and Hexadecimal Numbers], page 107.

At the very lowest level, computers store values as groups of binary digits, or bits. Modern computers group bits into groups of eight, called bytes. Advanced applications sometimes have to manipulate bits directly, and gawk provides functions for doing so.

Programs are written in programming languages. Hundreds, if not thousands, of programming languages exist. One of the most popular is the C programming language. The C language had a very strong influence on the design of the awk language.

There have been several versions of C. The first is often referred to as “K&R” C, after the initials of Brian Kernighan and Dennis Ritchie, the authors of the first book on C. (Dennis Ritchie created the language, and Brian Kernighan was one of the creators of awk.)

In the mid-1980s, an effort began to produce an international standard for C. This work culminated in 1989, with the production of the ANSI standard for C. This standard became an ISO standard in 1990. In 1999, a revised ISO C standard was approved and released. Where it makes sense, POSIX awk is compatible with 1999 ISO C.

Glossary 463

Glossary

Action A series of awk statements attached to a rule. If the rule’s pattern matches an input record, awk executes the rule’s action. Actions are always enclosed in braces. (See Section 7.3 [Actions], page 137.)

Ada A programming language originally defined by the U.S. Department of Defense for embedded programming. It was designed to enforce good Software Engineering practices.

Amazing awk Assembler Henry Spencer at the University of Toronto wrote a retargetable assembler completely as sed and awk scripts. It is thousands of lines long, including machine descriptions for several eight-bit microcomputers. It is a good example of a program that would have been better written in another language. You

can get it from http://awk.info/?awk100/aaa.

Amazingly Workable Formatter (awf)

Henry Spencer at the University of Toronto wrote a formatter that accepts a large subset of the ‘nroff -ms’ and ‘nroff -man’ formatting commands, using

awk and sh. It is available from http://awk.info/?tools/awf.

Anchor The regexp metacharacters ‘^’ and ‘$’, which force the match to the beginning or end of the string, respectively.

ANSI The American National Standards Institute. This organization produces many standards, among them the standards for the C and C++ programming languages. These standards often become international standards as well. See also “ISO.”

Argument An argument can be two different things. It can be an option or a file name passed to a command while invoking it from the command line, or it can be something passed to a function inside a program, e.g. inside awk. In the latter case, an argument can be passed to a function in two ways. Either it is given to the called function by value, i.e., a copy of the value of the variable is made available to the called function, but the original variable cannot be modified by the function itself; or it is given by reference, i.e., a pointer to the interested variable is passed to the function, which can then directly modify it. In awk scalars are passed by value, and arrays are passed by reference. See “Pass By Value/Reference.”

Array A grouping of multiple values under the same name. Most languages just provide sequential arrays. awk provides associative arrays.

Assertion A statement in a program that a condition is true at this point in the program. Useful for reasoning about how a program is supposed to behave.

Assignment An awk expression that changes the value of some awk variable or data object. An object that you can assign to is called an lvalue. The assigned values are called rvalues. See Section 6.2.3 [Assignment Expressions], page 116.

Associative Array Arrays in which the indices may be numbers or strings, not just sequential integers in a fixed range.

awk Language The language in which awk programs are written.

awk Program An awk program consists of a series of patterns and actions, collectively known as rules. For each input record given to the program, the program’s rules are all processed in turn. awk programs may also contain function definitions.

awk Script Another name for an awk program.

Bash The GNU version of the standard shell (the Bourne-Again SHell). See also “Bourne Shell.”

Binary Base-two notation, where the digits are 0–1. Since electronic circuitry works “naturally” in base 2 (just think of Off/On), everything inside a computer is calculated using base 2. Each digit represents the presence (or absence) of a power of 2 and is called a bit. So, for example, the base-two number 10101 is the same as decimal 21, ((1 x 16) + (1 x 4) + (1 x 1)). Since base-two numbers quickly become very long to read and write, they are usually grouped by 3 (i.e., they are read as octal numbers), or by 4 (i.e., they are read as hexadecimal numbers). There is no direct way to insert base 2 numbers in a C program. If need arises, such numbers are usually inserted as octal or hexadecimal numbers. The number of base-two digits that fit into registers used for representing integer numbers in computers is a rough indication of the computing power of the computer itself. Most computers nowadays use 64 bits for representing integer numbers in their registers, but 32-bit, 16-bit and 8-bit registers have been widely used in the past. See Section 6.1.1.2 [Octal and Hexadecimal Numbers], page 107.

Bit Short for “Binary Digit.” All values in computer memory ultimately reduce to binary digits: values that are either zero or one. Groups of bits may be interpreted differently—as integers, floating-point numbers, character data, addresses of other memory objects, or other data. awk lets you work with floatingpoint numbers and strings. gawk lets you manipulate bit values with the built-in functions described in Section 9.1.6 [Bit-Manipulation Functions], page 196. Computers are often defined by how many bits they use to represent integer values. Typical systems are 32-bit systems, but 64-bit systems are becoming increasingly popular, and 16-bit systems have essentially disappeared.

Boolean Expression Named after the English mathematician Boole. See also “Logical Expression.”

Bourne Shell The standard shell (/bin/sh) on Unix and Unix-like systems, originally written by Steven R. Bourne at Bell Laboratories. Many shells (Bash, ksh, pdksh, zsh) are generally upwardly compatible with the Bourne shell.

Braces The characters ‘{’ and ‘}’. Braces are used in awk for delimiting actions, compound statements, and function bodies.

Bracket Expression Inside a regular expression, an expression included in square brackets, meant to designate a single character as belonging to a specified character class. A bracket expression can contain a list of one or more characters, like ‘[abc]’, a range of characters, like ‘[A-Z]’, or a name, delimited by ‘:’, that designates a known set of characters, like ‘[:digit:]’. The form of bracket expression enclosed between ‘:’ is independent of the underlying representation of the character themselves, which could utilize the ASCII, ECBDIC, or Unicode codesets, depending on the architecture of the computer system, and on localization. See also “Regular Expression.”

Built-in Function The awk language provides built-in functions that perform various numerical, I/O-related, and string computations. Examples are sqrt() (for the square root of a number) and substr() (for a substring of a string). gawk provides functions for timestamp management, bit manipulation, array sorting, type checking, and runtime string translation. (See Section 9.1 [Built-in Functions], page 175.)

Built-in Variable

ARGC, ARGV, CONVFMT, ENVIRON, FILENAME, FNR, FS, NF, NR, OFMT, OFS, ORS,

RLENGTH, RSTART, RS, and SUBSEP are the variables that have special meaning

to awk. In addition, ARGIND, BINMODE, ERRNO, FIELDWIDTHS, FPAT, IGNORECASE,

LINT, PROCINFO, RT, and TEXTDOMAIN are the variables that have special meaning to gawk. Changing some of them affects awk’s running environment. (See Section 7.5 [Predefined Variables], page 147.)

C The system programming language that most GNU software is written in. The awk programming language has C-like syntax, and this book points out similarities between awk and C when appropriate. In general, gawk attempts to be as similar to the 1990 version of ISO C as makes sense.

C Shell The C Shell (csh or its improved version, tcsh) is a Unix shell that was created by Bill Joy in the late 1970s. The C shell was differentiated from other shells by its interactive features and overall style, which looks more like C. The C Shell is not backward compatible with the Bourne Shell, so special attention is required when converting scripts written for other Unix shells to the C shell, especially with regard to the management of shell variables. See also “Bourne Shell.”

C++ A popular object-oriented programming language derived from C.

Character Class See “Bracket Expression.”

Character List See “Bracket Expression.”

Character Set The set of numeric codes used by a computer system to represent the characters (letters, numbers, punctuation, etc.) of a particular country or place. The most common character set in use today is ASCII (American Standard Code for Information Interchange). Many European countries use an extension of ASCII known as ISO-8859-1 (ISO Latin-1). The Unicode character set (http://www. unicode.org) is increasingly popular and standard, and is particularly widely used on GNU/Linux systems.

CHEM A preprocessor for pic that reads descriptions of molecules and produces pic input for drawing them. It was written in awk by Brian Kernighan and Jon

Bentley, and is available from http://netlib.org/typesetting/chem.

Comparison Expression A relation that is either true or false, such as ‘a < b’. Comparison expressions are used in if, while, do, and for statements, and in patterns to select which input records to process. (See Section 6.3.2 [Variable Typing and Comparison Expressions], page 120.)

Compiler A program that translates human-readable source code into machine-executable object code. The object code is then executed directly by the computer. See also “Interpreter.”

Complemented Bracket Expression The negation of a bracket expression. All that is not described by a given bracket expression. The symbol ‘^’ precedes the negated bracket expression. E.g.: ‘[[^:digit:]’ designates whatever character is not a digit. ‘[^bad]’ designates whatever character is not one of the letters ‘b’, ‘a’, or ‘d’. See “Bracket Expression.”

Compound Statement A series of awk statements, enclosed in curly braces. Compound statements may be nested. (See Section 7.4 [Control Statements in Actions], page 138.)

Computed Regexps See “Dynamic Regular Expressions.”

Concatenation Concatenating two strings means sticking them together, one after another, producing a new string. For example, the string ‘foo’ concatenated with the string ‘bar’ gives the string ‘foobar’. (See Section 6.2.2 [String Concatenation], page 115.)

Conditional Expression An expression using the ‘?:’ ternary operator, such as ‘expr1 ? expr2 : expr3’. The expression expr1 is evaluated; if the result is true, the value of the whole expression is the value of expr2; otherwise the value is expr3. In either case, only one of expr2 and expr3 is evaluated. (See Section 6.3.4 [Conditional Expressions], page 125.)

Control Statement A control statement is an instruction to perform a given operation or a set of operations inside an awk program, if a given condition is true. Control statements are: if, for, while, and do (see Section 7.4 [Control Statements in Actions], page 138).

Cookie A peculiar goodie, token, saying or remembrance produced by or presented to a program. (With thanks to Professor Doug McIlroy.)

Coprocess A subordinate program with which two-way communications is possible.

Curly Braces See “Braces.”

Dark Corner An area in the language where specifications often were (or still are) not clear, leading to unexpected or undesirable behavior. Such areas are marked in this book with the picture of a flashlight in the margin and are indexed under the heading “dark corner.”

Data Driven A description of awk programs, where you specify the data you are interested in processing, and what to do when that data is seen.

Data Objects These are numbers and strings of characters. Numbers are converted into strings and vice versa, as needed. (See Section 6.1.4 [Conversion of Strings and Numbers], page 111.)

Deadlock The situation in which two communicating processes are each waiting for the other to perform an action.

Debugger A program used to help developers remove “bugs” from (de-bug) their programs.

Double Precision An internal representation of numbers that can have fractional parts. Double precision numbers keep track of more digits than do single precision numbers, but operations on them are sometimes more expensive. This is the way awk stores numeric values. It is the C type double.

Dynamic Regular Expression A dynamic regular expression is a regular expression written as an ordinary expression. It could be a string constant, such as "foo", but it may also be an expression whose value can vary. (See Section 3.6 [Using Dynamic Regexps], page 52.)

Empty String See “Null String.”

Environment A collection of strings, of the form ‘name=val’, that each program has available to it. Users generally place values into the environment in order to provide information to various programs. Typical examples are the environment variables

HOME and PATH.

Epoch The date used as the “beginning of time” for timestamps. Time values in most systems are represented as seconds since the epoch, with library functions available for converting these values into standard date and time formats. The epoch on Unix and POSIX systems is 1970-01-01 00:00:00 UTC. See also “GMT” and “UTC.”

Escape Sequences A special sequence of characters used for describing nonprinting characters, such as ‘\n’ for newline or ‘\033’ for the ASCII ESC (Escape) character. (See Section 3.2 [Escape Sequences], page 46.)

Extension An additional feature or change to a programming language or utility not defined by that language’s or utility’s standard. gawk has (too) many extensions over POSIX awk.

FDL See “Free Documentation License.”

Field When awk reads an input record, it splits the record into pieces separated by whitespace (or by a separator regexp that you can change by setting the predefined variable FS). Such pieces are called fields. If the pieces are of fixed length, you can use the built-in variable FIELDWIDTHS to describe their lengths. If you wish to specify the contents of fields instead of the field separator, you can use the predefined variable FPAT to do so. (See Section 4.5 [Specifying How Fields Are Separated], page 67, Section 4.6 [Reading Fixed-Width Data], page 72, and Section 4.7 [Defining Fields by Content], page 74.)

Flag A variable whose truth value indicates the existence or nonexistence of some condition.

Floating-Point Number Often referred to in mathematical terms as a “rational” or real number, this is just a number that can have a fractional part. See also “Double Precision” and “Single Precision.”

Format Format strings control the appearance of output in the strftime() and sprintf() functions, and in the printf statement as well. Also, data conversions from numbers to strings are controlled by the format strings contained in the predefined variables CONVFMT and OFMT. (See Section 5.5.2 [Format-Control Letters], page 93.)

Fortran Shorthand for FORmula TRANslator, one of the first programming languages available for scientific calculations. It was created by John Backus, and has been available since 1957. It is still in use today.

Free Documentation License This document describes the terms under which this book is published and may be copied. (See [GNU Free Documentation License], page 487.)

Free Software Foundation A nonprofit organization dedicated to the production and distribution of freely distributable software. It was founded by Richard M. Stallman, the author of the original Emacs editor. GNU Emacs is the most widely used version of Emacs today.

FSF See “Free Software Foundation.”

Function A part of an awk program that can be invoked from every point of the program, to perform a task. awk has several built-in functions. Users can define their own functions in every part of the program. Function can be recursive, i.e., they may invoke themselves. See Chapter 9 [Functions], page 175. In gawk it is also possible to have functions shared among different programs, and included where required using the @include directive (see Section 2.7 [Including Other Files into Your Program], page 42). In gawk the name of the function that should be invoked can be generated at run time, i.e., dynamically. The gawk extension API provides constructor functions (see Section 16.4.4 [Constructor Functions], page 363).

gawk The GNU implementation of awk.

General Public License This document describes the terms under which gawk and its source code may be distributed. (See [GNU General Public License], page 475.)

GMT “Greenwich Mean Time.” This is the old term for UTC. It is the time of day used internally for Unix and POSIX systems. See also “Epoch” and “UTC.”

GNU “GNU’s not Unix”. An on-going project of the Free Software Foundation to create a complete, freely distributable, POSIX-compliant computing environment.

GNU/Linux A variant of the GNU system using the Linux kernel, instead of the Free Software Foundation’s Hurd kernel. The Linux kernel is a stable, efficient, fullfeatured clone of Unix that has been ported to a variety of architectures. It is most popular on PC-class systems, but runs well on a variety of other systems too. The Linux kernel source code is available under the terms of the GNU General Public License, which is perhaps its most important aspect.

GPL See “General Public License.”

Hexadecimal Base 16 notation, where the digits are 0–9 and A–F, with ‘A’ representing 10, ‘B’ representing 11, and so on, up to ‘F’ for 15. Hexadecimal numbers are written in C using a leading ‘0x’, to indicate their base. Thus, 0x12 is 18 ((1 x 16) + 2). See Section 6.1.1.2 [Octal and Hexadecimal Numbers], page 107.

I/O Abbreviation for “Input/Output,” the act of moving data into and/or out of a running program.

Input Record A single chunk of data that is read in by awk. Usually, an awk input record consists of one line of text. (See Section 4.1 [How Input Is Split into Records], page 59.)

Integer A whole number, i.e., a number that does not have a fractional part.

Internationalization The process of writing or modifying a program so that it can use multiple languages without requiring further source code changes.

Interpreter A program that reads human-readable source code directly, and uses the instructions in it to process data and produce results. awk is typically (but not always) implemented as an interpreter. See also “Compiler.”

Interval Expression A component of a regular expression that lets you specify repeated matches of some part of the regexp. Interval expressions were not originally available in awk programs.

ISO The International Organization for Standardization. This organization produces international standards for many things, including programming languages, such as C and C++. In the computer arena, important standards like those for C, C++, and POSIX become both American national and ISO international standards simultaneously. This book refers to Standard C as “ISO C”

throughout. See the ISO website (http://www.iso.org/iso/home/about.

htm) for more information about the name of the organization and its languageindependent three-letter acronym.

Java A modern programming language originally developed by Sun Microsystems (now Oracle) supporting Object-Oriented programming. Although usually implemented by compiling to the instructions for a standard virtual machine (the JVM), the language can be compiled to native code.

Keyword In the awk language, a keyword is a word that has special meaning. Keywords are reserved and may not be used as variable names.

gawk’s keywords are: BEGIN, BEGINFILE, END, ENDFILE, break, case, continue,
defaultdelete, do...while, else, exit, for...in, for, function, func, if,
next, nextfile, switch, and while.

Korn Shell The Korn Shell (ksh) is a Unix shell which was developed by David Korn at Bell Laboratories in the early 1980s. The Korn Shell is backward-compatible with the Bourne shell and includes many features of the C shell. See also “Bourne Shell.”

Lesser General Public License This document describes the terms under which binary library archives or shared objects, and their source code may be distributed.

LGPL See “Lesser General Public License.”

Linux See “GNU/Linux.”

Localization The process of providing the data necessary for an internationalized program to work in a particular language.

Logical Expression An expression using the operators for logic, AND, OR, and NOT, written ‘&&’, ‘||’, and ‘!’ in awk. Often called Boolean expressions, after the mathematician who pioneered this kind of mathematical logic.

Lvalue An expression that can appear on the left side of an assignment operator. In most languages, lvalues can be variables or array elements. In awk, a field designator can also be used as an lvalue.

Matching The act of testing a string against a regular expression. If the regexp describes the contents of the string, it is said to match it.

Metacharacters Characters used within a regexp that do not stand for themselves. Instead, they denote regular expression operations, such as repetition, grouping, or alternation.

Nesting Nesting is where information is organized in layers, or where objects contain other similar objects. In gawk the @include directive can be nested. The “natural” nesting of arithmetic and logical operations can be changed using parentheses (see Section 6.5 [Operator Precedence (How Operators Nest)], page 128).

No-op An operation that does nothing.

Null String A string with no characters in it. It is represented explicitly in awk programs by placing two double quote characters next to each other (""). It can appear in input data by having two successive occurrences of the field separator appear next to each other.

Number A numeric-valued data object. Modern awk implementations use double precision floating-point to represent numbers. Ancient awk implementations used single precision floating-point.

Octal Base-eight notation, where the digits are 0–7. Octal numbers are written in C using a leading ‘0’, to indicate their base. Thus, 013 is 11 ((1 x 8) + 3). See Section 6.1.1.2 [Octal and Hexadecimal Numbers], page 107.

Output Record A single chunk of data that is written out by awk. Usually, an awk output record consists of one or more lines of text. See Section 4.1 [How Input Is Split into Records], page 59.

Pattern Patterns tell awk which input records are interesting to which rules. A pattern is an arbitrary conditional expression against which input is tested. If the condition is satisfied, the pattern is said to match the input record. A typical pattern might compare the input record against a regular expression. (See Section 7.1 [Pattern Elements], page 131.)

PEBKAC An acronym describing what is possibly the most frequent source of computer usage problems. (Problem Exists Between Keyboard And Chair.)

Plug-in See “Extensions.”

POSIX The name for a series of standards that specify a Portable Operating System interface. The “IX” denotes the Unix heritage of these standards. The main standard of interest for awk users is IEEE Standard for Information Technology, Standard 1003.1-2008. The 2008 POSIX standard can be found online at

http://www.opengroup.org/onlinepubs/9699919799/.

Precedence The order in which operations are performed when operators are used without explicit parentheses.

Private Variables and/or functions that are meant for use exclusively by library functions and not for the main awk program. Special care must be taken when naming such variables and functions. (See Section 10.1 [Naming Library Function Global Variables], page 218.)

Range (of input lines)

A sequence of consecutive lines from the input file(s). A pattern can specify ranges of input lines for awk to process or it can specify single lines. (See Section 7.1 [Pattern Elements], page 131.)

Record See “Input record” and “Output record.”

Recursion When a function calls itself, either directly or indirectly. If this is clear, stop, and proceed to the next entry. Otherwise, refer to the entry for “recursion.”

Redirection Redirection means performing input from something other than the standard input stream, or performing output to something other than the standard output stream. You can redirect input to the getline statement using the ‘<’, ‘|’, and ‘|&’ operators. You can redirect the output of the print and printf statements to a file or a system command, using the ‘>’, ‘>>’, ‘|’, and ‘|&’ operators. (See Section 4.9 [Explicit Input with getline], page 78, and Section 5.6 [Redirecting Output of print and printf], page 98.)

Reference Counts An internal mechanism in gawk to minimize the amount of memory needed to store the value of string variables. If the value assumed by a variable is used in more than one place, only one copy of the value itself is kept, and the associated reference count is increased when the same value is used by an additional variable, and decresed when the related variable is no longer in use. When the reference count goes to zero, the memory space used to store the value of the variable is freed.

Regexp See “Regular Expression.”

Regular Expression A regular expression (“regexp” for short) is a pattern that denotes a set of strings, possibly an infinite set. For example, the regular expression ‘R.*xp’ matches any string starting with the letter ‘R’ and ending with the letters ‘xp’. In awk, regular expressions are used in patterns and in conditional expressions. Regular expressions may contain escape sequences. (See Chapter 3 [Regular Expressions], page 45.)

Regular Expression Constant A regular expression constant is a regular expression written within slashes, such as /foo/. This regular expression is chosen when you write the awk program and cannot be changed during its execution. (See Section 3.1 [How to Use Regular Expressions], page 45.)

Regular Expression Operators See “Metacharacters.”

Rounding Rounding the result of an arithmetic operation can be tricky. More than one way of rounding exists, and in gawk it is possible to choose which method should be used in a program. See Section 15.4.5 [Setting the Rounding Mode], page 349.

Rule A segment of an awk program that specifies how to process single input records. A rule consists of a pattern and an action. awk reads an input record; then, for each rule, if the input record satisfies the rule’s pattern, awk executes the rule’s action. Otherwise, the rule does nothing for that input record.

Rvalue A value that can appear on the right side of an assignment operator. In awk, essentially every expression has a value. These values are rvalues.

Scalar A single value, be it a number or a string. Regular variables are scalars; arrays and functions are not.

Search Path In gawk, a list of directories to search for awk program source files. In the shell, a list of directories to search for executable programs.

sed See “Stream Editor.”

Seed The initial value, or starting point, for a sequence of random numbers.

Shell The command interpreter for Unix and POSIX-compliant systems. The shell works both interactively, and as a programming language for batch files, or shell scripts.

Short-Circuit The nature of the awk logical operators ‘&&’ and ‘||’. If the value of the entire expression is determinable from evaluating just the lefthand side of these operators, the righthand side is not evaluated. (See Section 6.3.3 [Boolean Expressions], page 124.)

Side Effect A side effect occurs when an expression has an effect aside from merely producing a value. Assignment expressions, increment and decrement expressions, and function calls have side effects. (See Section 6.2.3 [Assignment Expressions], page 116.)

Single Precision An internal representation of numbers that can have fractional parts. Single precision numbers keep track of fewer digits than do double precision numbers, but operations on them are sometimes less expensive in terms of CPU time. This is the type used by some ancient versions of awk to store numeric values. It is the C type float.

Space The character generated by hitting the space bar on the keyboard.

Special File A file name interpreted internally by gawk, instead of being handed directly to the underlying operating system—for example, /dev/stderr. (See Section 5.8 [Special File names in gawk], page 101.)

Statement An expression inside an awk program in the action part of a pattern–action rule, or inside an awk function. A statement can be a variable assignment, an array operation, a loop, etc.

Stream Editor A program that reads records from an input stream and processes them one or more at a time. This is in contrast with batch programs, which may expect to read their input files in entirety before starting to do anything, as well as with interactive programs which require input from the user.

String A datum consisting of a sequence of characters, such as ‘I am a string’. Constant strings are written with double quotes in the awk language and may contain escape sequences. (See Section 3.2 [Escape Sequences], page 46.)

Tab The character generated by hitting the TAB key on the keyboard. It usually expands to up to eight spaces upon output.

Text Domain A unique name that identifies an application. Used for grouping messages that are translated at runtime into the local language.

Timestamp A value in the “seconds since the epoch” format used by Unix and POSIX systems. Used for the gawk functions mktime(), strftime(), and systime(). See also “Epoch,” “GMT,” and “UTC.”

Unix A computer operating system originally developed in the early 1970’s at AT&T Bell Laboratories. It initially became popular in universities around the world and later moved into commercial environments as a software development system and network server system. There are many commercial versions of Unix, as well as several work-alike systems whose source code is freely available (such

as GNU/Linux, NetBSD (http://www.netbsd.org), FreeBSD (http://www.
freebsd.org), and OpenBSD (http://www.openbsd.org)).

UTC The accepted abbreviation for “Universal Coordinated Time.” This is standard time in Greenwich, England, which is used as a reference time for day and date calculations. See also “Epoch” and “GMT.”

Variable A name for a value. In awk, variables may be either scalars or arrays.

Whitespace A sequence of space, TAB, or newline characters occurring inside an input record or a string.

GNU General Public License 475

GNU General Public License

Version 3, 29 June 2007

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2007 Free Software Foundation, Inc. http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

Preamble

The GNU General Public License is a free, copyleft license for software and other kinds of works.

The licenses for most software and other practical works are designed to take away your freedom to share and change the works. By contrast, the GNU General Public License is intended to guarantee your freedom to share and change all versions of a program—to make sure it remains free software for all its users. We, the Free Software Foundation, use the GNU General Public License for most of our software; it applies also to any other work released this way by its authors. You can apply it to your programs, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for them if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs, and that you know you can do these things.

To protect your rights, we need to prevent others from denying you these rights or asking you to surrender the rights. Therefore, you have certain responsibilities if you distribute copies of the software, or if you modify it: responsibilities to respect the freedom of others.

For example, if you distribute copies of such a program, whether gratis or for a fee, you must pass on to the recipients the same freedoms that you received. You must make sure that they, too, receive or can get the source code. And you must show them these terms so they know their rights.

Developers that use the GNU GPL protect your rights with two steps: (1) assert copyright on the software, and (2) offer you this License giving you legal permission to copy, distribute and/or modify it.

For the developers’ and authors’ protection, the GPL clearly explains that there is no warranty for this free software. For both users’ and authors’ sake, the GPL requires that modified versions be marked as changed, so that their problems will not be attributed erroneously to authors of previous versions.

Some devices are designed to deny users access to install or run modified versions of the software inside them, although the manufacturer can do so. This is fundamentally incompatible with the aim of protecting users’ freedom to change the software. The systematic pattern of such abuse occurs in the area of products for individuals to use, which is precisely where it is most unacceptable. Therefore, we have designed this version of the GPL to prohibit the practice for those products. If such problems arise substantially in other domains, we stand ready to extend this provision to those domains in future versions of the GPL, as needed to protect the freedom of users.

Finally, every program is threatened constantly by software patents. States should not allow patents to restrict development and use of software on general-purpose computers, but in those that do, we wish to avoid the special danger that patents applied to a free program could make it effectively proprietary. To prevent this, the GPL assures that patents cannot be used to render the program non-free.

The precise terms and conditions for copying, distribution and modification follow.

TERMS AND CONDITIONS

0. Definitions. “This License” refers to version 3 of the GNU General Public License. “Copyright” also means copyright-like laws that apply to other kinds of works, such as semiconductor masks. “The Program” refers to any copyrightable work licensed under this License. Each licensee is addressed as “you”. “Licensees” and “recipients” may be individuals or organizations. To “modify” a work means to copy from or adapt all or part of the work in a fashion requiring copyright permission, other than the making of an exact copy. The resulting work is called a “modified version” of the earlier work or a work “based on” the earlier work. A “covered work” means either the unmodified Program or a work based on the Program. To “propagate” a work means to do anything with it that, without permission, would make you directly or secondarily liable for infringement under applicable copyright law, except executing it on a computer or modifying a private copy. Propagation includes copying, distribution (with or without modification), making available to the public, and in some countries other activities as well. To “convey” a work means any kind of propagation that enables other parties to make or receive copies. Mere interaction with a user through a computer network, with no transfer of a copy, is not conveying. An interactive user interface displays “Appropriate Legal Notices” to the extent that it includes a convenient and prominently visible feature that (1) displays an appropriate copyright notice, and (2) tells the user that there is no warranty for the work (except to the extent that warranties are provided), that licensees may convey the work under this License, and how to view a copy of this License. If the interface presents a list of user commands or options, such as a menu, a prominent item in the list meets this criterion.

1. Source Code. The “source code” for a work means the preferred form of the work for making modifications to it. “Object code” means any non-source form of a work. A “Standard Interface” means an interface that either is an official standard defined by a recognized standards body, or, in the case of interfaces specified for a particular programming language, one that is widely used among developers working in that language.

The “System Libraries” of an executable work include anything, other than the work as a whole, that (a) is included in the normal form of packaging a Major Component, but which is not part of that Major Component, and (b) serves only to enable use of the work with that Major Component, or to implement a Standard Interface for which an implementation is available to the public in source code form. A “Major Component”, in this context, means a major essential component (kernel, window system, and so on) of the specific operating system (if any) on which the executable work runs, or a compiler used to produce the work, or an object code interpreter used to run it.

The “Corresponding Source” for a work in object code form means all the source code needed to generate, install, and (for an executable work) run the object code and to modify the work, including scripts to control those activities. However, it does not include the work’s System Libraries, or general-purpose tools or generally available free programs which are used unmodified in performing those activities but which are not part of the work. For example, Corresponding Source includes interface definition files associated with source files for the work, and the source code for shared libraries and dynamically linked subprograms that the work is specifically designed to require, such as by intimate data communication or control flow between those subprograms and other parts of the work.

The Corresponding Source need not include anything that users can regenerate automatically from other parts of the Corresponding Source.

The Corresponding Source for a work in source code form is that same work.

2. Basic Permissions.

All rights granted under this License are granted for the term of copyright on the Program, and are irrevocable provided the stated conditions are met. This License explicitly affirms your unlimited permission to run the unmodified Program. The output from running a covered work is covered by this License only if the output, given its content, constitutes a covered work. This License acknowledges your rights of fair use or other equivalent, as provided by copyright law.

You may make, run and propagate covered works that you do not convey, without conditions so long as your license otherwise remains in force. You may convey covered works to others for the sole purpose of having them make modifications exclusively for you, or provide you with facilities for running those works, provided that you comply with the terms of this License in conveying all material for which you do not control copyright. Those thus making or running the covered works for you must do so exclusively on your behalf, under your direction and control, on terms that prohibit them from making any copies of your copyrighted material outside their relationship with you.

Conveying under any other circumstances is permitted solely under the conditions stated below. Sublicensing is not allowed; section 10 makes it unnecessary.

3. Protecting Users’ Legal Rights From Anti-Circumvention Law.

No covered work shall be deemed part of an effective technological measure under any applicable law fulfilling obligations under article 11 of the WIPO copyright treaty adopted on 20 December 1996, or similar laws prohibiting or restricting circumvention of such measures.

When you convey a covered work, you waive any legal power to forbid circumvention of technological measures to the extent such circumvention is effected by exercising rights under this License with respect to the covered work, and you disclaim any intention to limit operation or modification of the work as a means of enforcing, against the work’s users, your or third parties’ legal rights to forbid circumvention of technological measures.

4. Conveying Verbatim Copies.

You may convey verbatim copies of the Program’s source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice; keep intact all notices stating that this License and any non-permissive terms added in accord with section 7 apply to the code; keep intact all notices of the absence of any warranty; and give all recipients a copy of this License along with the Program. You may charge any price or no price for each copy that you convey, and you may offer support or warranty protection for a fee.

5. Conveying Modified Source Versions. You may convey a work based on the Program, or the modifications to produce it from the Program, in the form of source code under the terms of section 4, provided that you also meet all of these conditions: a. The work must carry prominent notices stating that you modified it, and giving a

relevant date. b. The work must carry prominent notices stating that it is released under this Li-

cense and any conditions added under section 7. This requirement modifies the requirement in section 4 to “keep intact all notices”.

c. You must license the entire work, as a whole, under this License to anyone who comes into possession of a copy. This License will therefore apply, along with any applicable section 7 additional terms, to the whole of the work, and all its parts, regardless of how they are packaged. This License gives no permission to license the work in any other way, but it does not invalidate such permission if you have separately received it.

d. If the work has interactive user interfaces, each must display Appropriate Legal Notices; however, if the Program has interactive interfaces that do not display Appropriate Legal Notices, your work need not make them do so.

A compilation of a covered work with other separate and independent works, which are not by their nature extensions of the covered work, and which are not combined with it such as to form a larger program, in or on a volume of a storage or distribution medium, is called an “aggregate” if the compilation and its resulting copyright are not used to limit the access or legal rights of the compilation’s users beyond what the individual works permit. Inclusion of a covered work in an aggregate does not cause this License to apply to the other parts of the aggregate.

6. Conveying Non-Source Forms.

You may convey a covered work in object code form under the terms of sections 4 and 5, provided that you also convey the machine-readable Corresponding Source under the terms of this License, in one of these ways: a. Convey the object code in, or embodied in, a physical product (including a phys-

ical distribution medium), accompanied by the Corresponding Source fixed on a durable physical medium customarily used for software interchange.

b. Convey the object code in, or embodied in, a physical product (including a physical distribution medium), accompanied by a written offer, valid for at least three years and valid for as long as you offer spare parts or customer support for that product model, to give anyone who possesses the object code either (1) a copy of the Corresponding Source for all the software in the product that is covered by this License, on a durable physical medium customarily used for software interchange, for a price no more than your reasonable cost of physically performing this conveying of source, or (2) access to copy the Corresponding Source from a network server at no charge.

c. Convey individual copies of the object code with a copy of the written offer to provide the Corresponding Source. This alternative is allowed only occasionally and noncommercially, and only if you received the object code with such an offer, in accord with subsection 6b.

d. Convey the object code by offering access from a designated place (gratis or for a charge), and offer equivalent access to the Corresponding Source in the same way through the same place at no further charge. You need not require recipients to copy the Corresponding Source along with the object code. If the place to copy the object code is a network server, the Corresponding Source may be on a different server (operated by you or a third party) that supports equivalent copying facilities, provided you maintain clear directions next to the object code saying where to find the Corresponding Source. Regardless of what server hosts the Corresponding Source, you remain obligated to ensure that it is available for as long as needed to satisfy these requirements.

e. Convey the object code using peer-to-peer transmission, provided you inform other peers where the object code and Corresponding Source of the work are being offered to the general public at no charge under subsection 6d.

A separable portion of the object code, whose source code is excluded from the Corresponding Source as a System Library, need not be included in conveying the object code work. A “User Product” is either (1) a “consumer product”, which means any tangible personal property which is normally used for personal, family, or household purposes, or (2) anything designed or sold for incorporation into a dwelling. In determining whether a product is a consumer product, doubtful cases shall be resolved in favor of coverage. For a particular product received by a particular user, “normally used” refers to a typical or common use of that class of product, regardless of the status of the particular user or of the way in which the particular user actually uses, or expects or is expected to use, the product. A product is a consumer product regardless of whether the product has substantial commercial, industrial or non-consumer uses, unless such uses represent the only significant mode of use of the product. “Installation Information” for a User Product means any methods, procedures, authorization keys, or other information required to install and execute modified versions of a covered work in that User Product from a modified version of its Corresponding Source.

The information must suffice to ensure that the continued functioning of the modified object code is in no case prevented or interfered with solely because modification has been made. If you convey an object code work under this section in, or with, or specifically for use in, a User Product, and the conveying occurs as part of a transaction in which the right of possession and use of the User Product is transferred to the recipient in perpetuity or for a fixed term (regardless of how the transaction is characterized), the Corresponding Source conveyed under this section must be accompanied by the Installation Information. But this requirement does not apply if neither you nor any third party retains the ability to install modified object code on the User Product (for example, the work has been installed in ROM). The requirement to provide Installation Information does not include a requirement to continue to provide support service, warranty, or updates for a work that has been modified or installed by the recipient, or for the User Product in which it has been modified or installed. Access to a network may be denied when the modification itself materially and adversely affects the operation of the network or violates the rules and protocols for communication across the network. Corresponding Source conveyed, and Installation Information provided, in accord with this section must be in a format that is publicly documented (and with an implementation available to the public in source code form), and must require no special password or key for unpacking, reading or copying.

7. Additional Terms. “Additional permissions” are terms that supplement the terms of this License by making exceptions from one or more of its conditions. Additional permissions that are applicable to the entire Program shall be treated as though they were included in this License, to the extent that they are valid under applicable law. If additional permissions apply only to part of the Program, that part may be used separately under those permissions, but the entire Program remains governed by this License without regard to the additional permissions. When you convey a copy of a covered work, you may at your option remove any additional permissions from that copy, or from any part of it. (Additional permissions may be written to require their own removal in certain cases when you modify the work.) You may place additional permissions on material, added by you to a covered work, for which you have or can give appropriate copyright permission. Notwithstanding any other provision of this License, for material you add to a covered work, you may (if authorized by the copyright holders of that material) supplement the terms of this License with terms: a. Disclaiming warranty or limiting liability differently from the terms of sections 15

and 16 of this License; or b. Requiring preservation of specified reasonable legal notices or author attributions

in that material or in the Appropriate Legal Notices displayed by works containing it; or

c. Prohibiting misrepresentation of the origin of that material, or requiring that modified versions of such material be marked in reasonable ways as different from the original version; or d. Limiting the use for publicity purposes of names of licensors or authors of the

material; or e. Declining to grant rights under trademark law for use of some trade names, trade-

marks, or service marks; or f. Requiring indemnification of licensors and authors of that material by anyone who

conveys the material (or modified versions of it) with contractual assumptions of liability to the recipient, for any liability that these contractual assumptions directly impose on those licensors and authors.

All other non-permissive additional terms are considered “further restrictions” within the meaning of section 10. If the Program as you received it, or any part of it, contains a notice stating that it is governed by this License along with a term that is a further restriction, you may remove that term. If a license document contains a further restriction but permits relicensing or conveying under this License, you may add to a covered work material governed by the terms of that license document, provided that the further restriction does not survive such relicensing or conveying. If you add terms to a covered work in accord with this section, you must place, in the relevant source files, a statement of the additional terms that apply to those files, or a notice indicating where to find the applicable terms. Additional terms, permissive or non-permissive, may be stated in the form of a separately written license, or stated as exceptions; the above requirements apply either way.

8. Termination.

You may not propagate or modify a covered work except as expressly provided under this License. Any attempt otherwise to propagate or modify it is void, and will automatically terminate your rights under this License (including any patent licenses granted under the third paragraph of section 11). However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, you do not qualify to receive new licenses for the same material under section 10.

9. Acceptance Not Required for Having Copies. You are not required to accept this License in order to receive or run a copy of the Program. Ancillary propagation of a covered work occurring solely as a consequence of using peer-to-peer transmission to receive a copy likewise does not require acceptance.

However, nothing other than this License grants you permission to propagate or modify any covered work. These actions infringe copyright if you do not accept this License. Therefore, by modifying or propagating a covered work, you indicate your acceptance of this License to do so.

10. Automatic Licensing of Downstream Recipients. Each time you convey a covered work, the recipient automatically receives a license from the original licensors, to run, modify and propagate that work, subject to this License. You are not responsible for enforcing compliance by third parties with this License. An “entity transaction” is a transaction transferring control of an organization, or substantially all assets of one, or subdividing an organization, or merging organizations. If propagation of a covered work results from an entity transaction, each party to that transaction who receives a copy of the work also receives whatever licenses to the work the party’s predecessor in interest had or could give under the previous paragraph, plus a right to possession of the Corresponding Source of the work from the predecessor in interest, if the predecessor has it or can get it with reasonable efforts. You may not impose any further restrictions on the exercise of the rights granted or affirmed under this License. For example, you may not impose a license fee, royalty, or other charge for exercise of rights granted under this License, and you may not initiate litigation (including a cross-claim or counterclaim in a lawsuit) alleging that any patent claim is infringed by making, using, selling, offering for sale, or importing the Program or any portion of it.

11. Patents.

A “contributor” is a copyright holder who authorizes use under this License of the Program or a work on which the Program is based. The work thus licensed is called the contributor’s “contributor version”. A contributor’s “essential patent claims” are all patent claims owned or controlled by the contributor, whether already acquired or hereafter acquired, that would be infringed by some manner, permitted by this License, of making, using, or selling its contributor version, but do not include claims that would be infringed only as a consequence of further modification of the contributor version. For purposes of this definition, “control” includes the right to grant patent sublicenses in a manner consistent with the requirements of this License. Each contributor grants you a non-exclusive, worldwide, royalty-free patent license under the contributor’s essential patent claims, to make, use, sell, offer for sale, import and otherwise run, modify and propagate the contents of its contributor version. In the following three paragraphs, a “patent license” is any express agreement or commitment, however denominated, not to enforce a patent (such as an express permission to practice a patent or covenant not to sue for patent infringement). To “grant” such a patent license to a party means to make such an agreement or commitment not to enforce a patent against the party. If you convey a covered work, knowingly relying on a patent license, and the Corresponding Source of the work is not available for anyone to copy, free of charge and under the terms of this License, through a publicly available network server or other readily accessible means, then you must either (1) cause the Corresponding Source to be so available, or (2) arrange to deprive yourself of the benefit of the patent license for this particular work, or (3) arrange, in a manner consistent with the requirements of this License, to extend the patent license to downstream recipients. “Knowingly relying” means you have actual knowledge that, but for the patent license, your conveying the covered work in a country, or your recipient’s use of the covered work in a country, would infringe one or more identifiable patents in that country that you have reason to believe are valid.

If, pursuant to or in connection with a single transaction or arrangement, you convey, or propagate by procuring conveyance of, a covered work, and grant a patent license to some of the parties receiving the covered work authorizing them to use, propagate, modify or convey a specific copy of the covered work, then the patent license you grant is automatically extended to all recipients of the covered work and works based on it.

A patent license is “discriminatory” if it does not include within the scope of its coverage, prohibits the exercise of, or is conditioned on the non-exercise of one or more of the rights that are specifically granted under this License. You may not convey a covered work if you are a party to an arrangement with a third party that is in the business of distributing software, under which you make payment to the third party based on the extent of your activity of conveying the work, and under which the third party grants, to any of the parties who would receive the covered work from you, a discriminatory patent license (a) in connection with copies of the covered work conveyed by you (or copies made from those copies), or (b) primarily for and in connection with specific products or compilations that contain the covered work, unless you entered into that arrangement, or that patent license was granted, prior to 28 March 2007.

Nothing in this License shall be construed as excluding or limiting any implied license or other defenses to infringement that may otherwise be available to you under applicable patent law.

12. No Surrender of Others’ Freedom.

If conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot convey a covered work so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not convey it at all. For example, if you agree to terms that obligate you to collect a royalty for further conveying from those to whom you convey the Program, the only way you could satisfy both those terms and this License would be to refrain entirely from conveying the Program.

13. Use with the GNU Affero General Public License.

Notwithstanding any other provision of this License, you have permission to link or combine any covered work with a work licensed under version 3 of the GNU Affero General Public License into a single combined work, and to convey the resulting work. The terms of this License will continue to apply to the part which is the covered work, but the special requirements of the GNU Affero General Public License, section 13, concerning interaction through a network will apply to the combination as such.

14. Revised Versions of this License.

The Free Software Foundation may publish revised and/or new versions of the GNU General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

Each version is given a distinguishing version number. If the Program specifies that a certain numbered version of the GNU General Public License “or any later version” applies to it, you have the option of following the terms and conditions either of that numbered version or of any later version published by the Free Software Foundation. If the Program does not specify a version number of the GNU General Public License, you may choose any version ever published by the Free Software Foundation.

If the Program specifies that a proxy can decide which future versions of the GNU General Public License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Program.

Later license versions may give you additional or different permissions. However, no additional obligations are imposed on any author or copyright holder as a result of your choosing to follow a later version.

15. Disclaimer of Warranty.

THERE IS NO WARRANTY FOR THE PROGRAM, TO THE EXTENT PER-MITTED BY APPLICABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPYRIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE PROGRAM “AS IS” WITHOUT WARRANTY OF ANY KIND, EITHER EX-PRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUALITY AND PERFORMANCE OF THE PROGRAM IS WITH YOU. SHOULD THE PROGRAM PROVE DEFEC-TIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

16. Limitation of Liability.

IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRITING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MODIFIES AND/OR CONVEYS THE PROGRAM AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, IN-CIDENTAL OR CONSEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE PROGRAM (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUS-TAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE PROGRAM TO OPERATE WITH ANY OTHER PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAM-AGES.

17. Interpretation of Sections 15 and 16.

If the disclaimer of warranty and limitation of liability provided above cannot be given local legal effect according to their terms, reviewing courts shall apply local law that most closely approximates an absolute waiver of all civil liability in connection with the Program, unless a warranty or assumption of liability accompanies a copy of the Program in return for a fee.

END OF TERMS AND CONDITIONS

How to Apply These Terms to Your New Programs

If you develop a new program, and you want it to be of the greatest possible use to the public, the best way to achieve this is to make it free software which everyone can redistribute and change under these terms.

To do so, attach the following notices to the program. It is safest to attach them to the start of each source file to most effectively state the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

      one line to give the program's name and a brief idea of what it does.
      Copyright (C) year name of author
      This program is free software: you can redistribute it and/or modify
      it under the terms of the GNU General Public License as published by
      the Free Software Foundation, either version 3 of the License, or (at
      your option) any later version.
      This program is distributed in the hope that it will be useful, but
      WITHOUT ANY WARRANTY; without even the implied warranty of
      MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE.  See the GNU
      General Public License for more details.
      You should have received a copy of the GNU General Public License
      along with this program.  If not, see http://www.gnu.org/licenses/.

Also add information on how to contact you by electronic and paper mail.

If the program does terminal interaction, make it output a short notice like this when it starts in an interactive mode:

      program Copyright (C) year name of author
      This program comes with ABSOLUTELY NO WARRANTY; for details type `show w'.
      This is free software, and you are welcome to redistribute it
      under certain conditions; type `show c' for details.

The hypothetical commands ‘show w’ and ‘show c’ should show the appropriate parts of the General Public License. Of course, your program’s commands might be different; for a GUI interface, you would use an “about box”.

You should also get your employer (if you work as a programmer) or school, if any, to sign a “copyright disclaimer” for the program, if necessary. For more information on this, and how to apply and follow the GNU GPL, see http://www.gnu.org/licenses/.

The GNU General Public License does not permit incorporating your program into proprietary programs. If your program is a subroutine library, you may consider it more useful to permit linking proprietary applications with the library. If this is what you want to do, use the GNU Lesser General Public License instead of this License. But first, please

read http://www.gnu.org/philosophy/why-not-lgpl.html.

GNU Free Documentation License 487

GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

 http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

The “publisher” means any person or entity that distributes copies of the Document to the public.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition.

The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

    Copyright (C)  year  your name.
    Permission is granted to copy, distribute and/or modify this document
    under the terms of the GNU Free Documentation License, Version 1.3
    or any later version published by the Free Software Foundation;
    with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
    Texts.  A copy of the license is included in the section entitled ``GNU
    Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Index 495

Index

!

! (exclamation point), ! operator … 125, 128, 133,

259 ! (exclamation point), != operator . . . . . . . 122, 128 ! (exclamation point), !~ operator . . . . . 45, 52, 55,

108, 122, 123, 129, 132

"

" (double quote), in regexp constants . . . . . . . . . . 53 " (double quote), in shell commands . . . . . . . . . . . 22

#

# (number sign), #! (executable scripts). . . . . . . . 19 # (number sign), commenting . . . . . . . . . . . . . . . . . . 20

$

$ (dollar sign), $ field operator . . . . . . . . . . . . 63, 128 $ (dollar sign), incrementing fields and arrays

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 $ (dollar sign), regexp operator . . . . . . . . . . . . . . . . 48

%

% (percent sign), % operator . . . . . . . . . . . . . . . . . . . 128 % (percent sign), %= operator . . . . . . . . . . . . . 118, 129

&

& (ampersand), && operator . . . . . . . . . . . . . . 125, 129 & (ampersand), gsub()/gensub()/sub() functions

and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186

’

' (single quote) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ' (single quote) in gawk command lines . . . . . . . . 19 ' (single quote), in shell commands . . . . . . . . . . . . 21 ' (single quote), vs. apostrophe . . . . . . . . . . . . . . . . 20 ' (single quote), with double quotes. . . . . . . . . . . . 22

(

() (parentheses), in a profile . . . . . . . . . . . . . . . . . . 311 () (parentheses), regexp operator . . . . . . . . . . . . . . 49

*

* (asterisk), * operator, as multiplication operator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

* (asterisk), * operator, as regexp operator. . . . . 49 * (asterisk), * operator, null strings, matching

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 * (asterisk), ** operator . . . . . . . . . . . . . . . . . 115, 128 * (asterisk), **= operator . . . . . . . . . . . . . . . . 118, 129 * (asterisk), *= operator . . . . . . . . . . . . . . . . . 118, 129

+

+ (plus sign), + operator . . . . . . . . . . . . . . . . . . . . . . 128 + (plus sign), ++ operator . . . . . . . . . . . . . . . . 119, 128 + (plus sign), += operator . . . . . . . . . . . . . . . . 117, 129 + (plus sign), regexp operator . . . . . . . . . . . . . . . . . . 49

,

, (comma), in range patterns . . . . . . . . . . . . . . . . . 133

-

- (hyphen), - operator. . . . . . . . . . . . . . . . . . . . . . . . 128 - (hyphen), -- operator. . . . . . . . . . . . . . . . . . 119, 128 - (hyphen), -= operator. . . . . . . . . . . . . . . . . . 118, 129 - (hyphen), filenames beginning with . . . . . . . . . . 32 - (hyphen), in bracket expressions . . . . . . . . . . . . . 51 --assign option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 --bignum option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 --characters-as-bytes option . . . . . . . . . . . . . . . . 32 --copyright option . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 --debug option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 --disable-extensions configuration option . . 433 --disable-lint configuration option . . . . . . . . . 433 --disable-nls configuration option . . . . . . . . . . 434 --dump-variables option. . . . . . . . . . . . . . . . . . . . . . 33 --dump-variables option, using for library

functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 --exec option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 --field-separator option. . . . . . . . . . . . . . . . . . . . . 31 --file option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 --gen-pot option . . . . . . . . . . . . . . . . . . . . . . . . . 34, 319 --help option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 --include option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 --lint option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31, 34 --lint-old option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 --load option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 --non-decimal-data option. . . . . . . . . . . . . . . 35, 299

--non-decimal-data option, strtonum() function

and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 --optimize option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 --posix option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 --posix option, --traditional option and …. 36 --pretty-print option . . . . . . . . . . . . . . . . . . . . . . . . 35 --profile option . . . . . . . . . . . . . . . . . . . . . . . . . 35, 308 --re-interval option . . . . . . . . . . . . . . . . . . . . . . . . . 36 --sandbox option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 --sandbox option, disabling system() function

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 --sandbox option, input redirection with getline

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 --sandbox option, output redirection with print,

printf. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 --source option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 --traditional option . . . . . . . . . . . . . . . . . . . . . . . . . 32 --traditional option, --posix option and …. 36 --use-lc-numeric option. . . . . . . . . . . . . . . . . . . . . . 35 --version option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

--with-whiny-user-strftime configuration option

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 434 -b option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 -c option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 -C option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 -d option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 -D option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 -e option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33, 37 -E option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 -f option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 31 -F option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 -F option, -Ft sets FS to TAB . . . . . . . . . . . . . . . . . 36 -F option, command-line. . . . . . . . . . . . . . . . . . . . . . . 69 -f option, multiple uses. . . . . . . . . . . . . . . . . . . . . . . . 36 -g option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 -h option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 -i option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 -l option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 -L option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 -M option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 -n option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 -N option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 -o option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 -O option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 -p option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 -P option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 -r option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 -S option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 -v option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32, 111 -V option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 -W option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

.

. (period), regexp operator . . . . . . . . . . . . . . . . . . . . 48 .gmo files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316 .gmo files, specifying directory of . . . . . . . . . 316, 317 .mo files, converting from .po . . . . . . . . . . . . . . . . . 322 .po files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315, 319 .po files, converting to .mo . . . . . . . . . . . . . . . . . . . 322 .pot files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315

/

/ (forward slash) to enclose regular expressions

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 / (forward slash), / operator. . . . . . . . . . . . . . . . . . 128 / (forward slash), /= operator . . . . . . . . . . . . 118, 129 / (forward slash), /= operator, vs. /=.../ regexp

constant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 / (forward slash), patterns and . . . . . . . . . . . . . . . 132 /= operator vs. /=.../ regexp constant . . . . . . . 118 /dev/... special files . . . . . . . . . . . . . . . . . . . . . . . . . 101 /dev/fd/N special files (gawk). . . . . . . . . . . . . . . . . 101 /inet/... special files (gawk). . . . . . . . . . . . . . . . . 307 /inet4/... special files (gawk) . . . . . . . . . . . . . . . 307 /inet6/... special files (gawk) . . . . . . . . . . . . . . . 307

;

; (semicolon), AWKPATH variable and. . . . . . . . . . . 437 ; (semicolon), separating statements in actions

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29, 137, 138

                              <

< (left angle bracket), < operator . . . . . . . . . 122, 128 < (left angle bracket), < operator (I/O). . . . . . . . . 80 < (left angle bracket), <= operator . . . . . . . . 122, 128

=

= (equals sign), = operator . . . . . . . . . . . . . . . . . . . . 116 = (equals sign), == operator . . . . . . . . . . . . . . 122, 128

                              >

> (right angle bracket), > operator . . . . . . . 122, 128 > (right angle bracket), > operator (I/O) . . . . . . . 98 > (right angle bracket), >= operator . . . . . . 122, 128 > (right angle bracket), >> operator (I/O) . . 98, 128

?

? (question mark), ?: operator. . . . . . . . . . . . . . . . 129 ? (question mark), regexp operator . . . . . . . . . 49, 55

@

@-notation for indirect function calls . . . . . . . . . . 209 @include directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 @load directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

[

[] (square brackets), regexp operator . . . . . . . . . . 49

                              ^

^(caret), ^operator . . . . . . . . . . . . . . . . . . . . . . . . . . 128 ^(caret), ^= operator . . . . . . . . . . . . . . . . . . . . 118, 129 ^(caret), in bracket expressions. . . . . . . . . . . . . . . . 51 ^(caret), in FS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 ^(caret), regexp operator . . . . . . . . . . . . . . . . . . 48, 55 _(underscore), C macro . . . . . . . . . . . . . . . . . . . . . . 316 _(underscore), in names of private variables . . 218 _(underscore), translatable string . . . . . . . . . . . . 318 _gr_init() user-defined function . . . . . . . . . . . . . 243 _ord_init() user-defined function . . . . . . . . . . . . 223 _pw_init() user-defined function . . . . . . . . . . . . . 240

\

\ (backslash) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 \ (backslash), \" escape sequence . . . . . . . . . . . . . . 47 \ (backslash), \' operator (gawk) . . . . . . . . . . . . . . 55 \ (backslash), \/ escape sequence . . . . . . . . . . . . . . 47 \ (backslash), \< operator (gawk) . . . . . . . . . . . . . . 54 \ (backslash), \> operator (gawk) . . . . . . . . . . . . . . 54 \ (backslash), \` operator (gawk) . . . . . . . . . . . . . . 55 \ (backslash), \a escape sequence . . . . . . . . . . . . . . 46 \ (backslash), \b escape sequence . . . . . . . . . . . . . . 46 \ (backslash), \B operator (gawk) . . . . . . . . . . . . . . 54 \ (backslash), \f escape sequence . . . . . . . . . . . . . . 46 \ (backslash), \n escape sequence . . . . . . . . . . . . . . 46 \ (backslash), \nnn escape sequence . . . . . . . . . . . 46 \ (backslash), \r escape sequence . . . . . . . . . . . . . . 46 \ (backslash), \s operator (gawk) . . . . . . . . . . . . . . 54 \ (backslash), \S operator (gawk) . . . . . . . . . . . . . . 54 \ (backslash), \t escape sequence . . . . . . . . . . . . . . 46 \ (backslash), \v escape sequence . . . . . . . . . . . . . . 46 \ (backslash), \w operator (gawk) . . . . . . . . . . . . . . 54 \ (backslash), \W operator (gawk) . . . . . . . . . . . . . . 54 \ (backslash), \x escape sequence . . . . . . . . . . . . . . 46 \ (backslash), \y operator (gawk) . . . . . . . . . . . . . . 54 \ (backslash), as field separator . . . . . . . . . . . . . . . . 70 \ (backslash), continuing lines and . . . . . . . . . . . . . 27 \ (backslash), continuing lines and, comments and

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 \ (backslash), continuing lines and, in csh. . . . . . 28 \ (backslash), gsub()/gensub()/sub() functions

and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 \ (backslash), in bracket expressions . . . . . . . . . . . 51 \ (backslash), in escape sequences . . . . . . . . . . 46, 47 \ (backslash), in escape sequences, POSIX and

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \ (backslash), in regexp constants . . . . . . . . . . . . . . 53 \ (backslash), in shell commands. . . . . . . . . . . . . . . 21 \ (backslash), regexp operator . . . . . . . . . . . . . . . . . 48

|

| (vertical bar) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 | (vertical bar), | operator (I/O) . . . . . . 81, 98, 128 | (vertical bar), |& operator (I/O) …. 82, 99, 128,

306 | (vertical bar), |& operator (I/O), pipes, closing

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 | (vertical bar), || operator. . . . . . . . . . . . . . 125, 129 {} (braces) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311 {} (braces), actions and . . . . . . . . . . . . . . . . . . . . . . 137 {} (braces), statements, grouping . . . . . . . . . . . . . 138

~

~ (tilde), ~ operator . . . . . 45, 52, 55, 108, 122, 123, 129, 132

A

accessing fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 accessing global variables from extensions . . . . . 373 account information . . . . . . . . . . . . . . . . . . . . . 238, 242 actions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 actions, control statements in . . . . . . . . . . . . . . . . . 138 actions, default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 actions, empty. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 Ada programming language. . . . . . . . . . . . . . . . . . . 463 adding, features to gawk . . . . . . . . . . . . . . . . . . . . . . 448 adding, fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 advanced features, fixed-width data . . . . . . . . . . . . 72 advanced features, gawk . . . . . . . . . . . . . . . . . . . . . . 299 advanced features, network programming . . . . . 307 advanced features, nondecimal input data . . . . . 299 advanced features, processes, communicating with

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 advanced features, specifying field content. . . . . . 74 Aho, Alfred . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 425 alarm clock example program . . . . . . . . . . . . . . . . . 272 alarm.awk program. . . . . . . . . . . . . . . . . . . . . . . . . . . 272 algorithms. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 allocating memory for extensions . . . . . . . . . . . . . 362 amazing awk assembler (aaa) . . . . . . . . . . . . . . . . . 463 amazingly workable formatter (awf). . . . . . . . . . . 463 ambiguity, syntactic: /= operator vs. /=.../ regexp

constant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 ampersand (&), && operator . . . . . . . . . . . . . . 125, 129 ampersand (&), gsub()/gensub()/sub() functions

and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 anagram.awk program . . . . . . . . . . . . . . . . . . . . . . . . 292 anagrams, finding . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 AND bitwise operation . . . . . . . . . . . . . . . . . . . . . . . 196 and Boolean-logic operator . . . . . . . . . . . . . . . . . . . 124 and() function (gawk) . . . . . . . . . . . . . . . . . . . . . . . . 196 ANSI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 463 API informational variables. . . . . . . . . . . . . . . . . . . 386 API version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 385 arbitrary precision . . . . . . . . . . . . . . . . . . . . . . . . . . . . 341 arbitrary precision integers . . . . . . . . . . . . . . . . . . . 350 archaeologists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 442 arctangent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 ARGC/ARGV variables . . . . . . . . . . . . . . . . . . . . . . . . . . 149 ARGC/ARGV variables, command-line arguments

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 ARGC/ARGV variables, how to use . . . . . . . . . . . . . . 155 ARGC/ARGV variables, portability and . . . . . . . . . . . 20 ARGIND variable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 ARGIND variable, command-line arguments. . . . . . 37 arguments, command-line . . . . . . . . . . . . 37, 149, 155 arguments, command-line, invoking awk . . . . . . . . 31 arguments, in function calls. . . . . . . . . . . . . . . . . . . 126 arguments, processing . . . . . . . . . . . . . . . . . . . . . . . . 233 ARGV array, indexing into. . . . . . . . . . . . . . . . . . . . . . . 37 arithmetic operators . . . . . . . . . . . . . . . . . . . . . . . . . . 114 array manipulation in extensions. . . . . . . . . . . . . . 377 array members . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 array scanning order, controlling . . . . . . . . . . . . . . 164 array, number of elements . . . . . . . . . . . . . . . . . . . . 180 arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 arrays of arrays. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 arrays, an example of using . . . . . . . . . . . . . . . . . . . 162 arrays, and IGNORECASE variable . . . . . . . . . . . . . . 161 arrays, as parameters to functions . . . . . . . . . . . . 205 arrays, associative . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 arrays, associative, library functions and . . . . . . 218 arrays, deleting entire contents. . . . . . . . . . . . . . . . 169 arrays, elements that don’t exist . . . . . . . . . . . . . . 161 arrays, elements, assigning values . . . . . . . . . . . . . 162 arrays, elements, deleting . . . . . . . . . . . . . . . . . . . . . 168 arrays, elements, order of access by in operator

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 arrays, elements, retrieving number of . . . . . . . . 178 arrays, for statement and . . . . . . . . . . . . . . . . . . . . 163 arrays, indexing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 arrays, merging into strings . . . . . . . . . . . . . . . . . . . 225 arrays, multidimensional . . . . . . . . . . . . . . . . . . . . . . 169 arrays, multidimensional, scanning . . . . . . . . . . . . 171 arrays, numeric subscripts . . . . . . . . . . . . . . . . . . . . 167 arrays, referencing elements. . . . . . . . . . . . . . . . . . . 161 arrays, scanning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 arrays, sorting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 arrays, sorting, and IGNORECASE variable . . . . . . 305 arrays, sparse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 arrays, subscripts, uninitialized variables as … 167 arrays, unassigned elements. . . . . . . . . . . . . . . . . . . 161 artificial intelligence, gawk and . . . . . . . . . . . . . . . . 430 ASCII . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224, 465 asort() function (gawk) . . . . . . . . . . . . . . . . . 178, 304 asort() function (gawk), arrays, sorting . . . . . . 304 asorti() function (gawk) . . . . . . . . . . . . . . . . 178, 304 asorti() function (gawk), arrays, sorting . . . . . 304 assert() function (C library) . . . . . . . . . . . . . . . . 220 assert() user-defined function . . . . . . . . . . . . . . . 221 assertions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 assign values to variables, in debugger . . . . . . . . 334 assignment operators . . . . . . . . . . . . . . . . . . . . . . . . . 116 assignment operators, evaluation order . . . . . . . . 118 assignment operators, lvalues/rvalues . . . . . . . . . 116 assignments as filenames. . . . . . . . . . . . . . . . . . . . . . 232 associative arrays. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 asterisk (*), * operator, as multiplication operator

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 asterisk (*), * operator, as regexp operator. . . . . 49 asterisk (*), * operator, null strings, matching

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 asterisk (*), ** operator . . . . . . . . . . . . . . . . . 115, 128 asterisk (*), **= operator . . . . . . . . . . . . . . . . 118, 129 asterisk (*), *= operator . . . . . . . . . . . . . . . . . 118, 129 atan2() function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 automatic displays, in debugger. . . . . . . . . . . . . . . 335 awf (amazingly workable formatter) program . . 463 awk debugging, enabling . . . . . . . . . . . . . . . . . . . . . . . 33 awk language, POSIX version . . . . . . . . . . . . . . . . . 118 awk profiling, enabling . . . . . . . . . . . . . . . . . . . . . . . . . 35 awk programs . . . . . . . . . . . . . . . . . . . . . . . . . . 17, 19, 26 awk programs, complex . . . . . . . . . . . . . . . . . . . . . . . . 29 awk programs, documenting . . . . . . . . . . . . . . . 20, 218 awk programs, examples of. . . . . . . . . . . . . . . . . . . . 251 awk programs, execution of . . . . . . . . . . . . . . . . . . . 144 awk programs, internationalizing . . . . . . . . . 198, 317 awk programs, lengthy . . . . . . . . . . . . . . . . . . . . . . . . . 18 awk programs, lengthy, assertions . . . . . . . . . . . . . 220 awk programs, location of . . . . . . . . . . . . . . . 31, 33, 34 awk programs, one-line examples . . . . . . . . . . . . . . . 25 awk programs, profiling . . . . . . . . . . . . . . . . . . . . . . . 308 awk programs, running . . . . . . . . . . . . . . . . . . . . . 17, 18 awk programs, running, from shell scripts . . . . . . 17 awk programs, running, without input files . . . . . 18 awk programs, shell variables in . . . . . . . . . . . . . . . 136 awk, function of. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 awk, gawk and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5, 7 awk, history of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 awk, implementation issues, pipes . . . . . . . . . . . . . 100 awk, implementations . . . . . . . . . . . . . . . . . . . . . . . . . 443 awk, implementations, limits . . . . . . . . . . . . . . . . . . . 83 awk, invoking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 awk, new vs. old . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 awk, new vs. old, OFMT variable. . . . . . . . . . . . . . . . 112 awk, POSIX and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 awk, POSIX and, See Also POSIX awk . . . . . . . . . . 5 awk, regexp constants and . . . . . . . . . . . . . . . . . . . . 123 awk, See Also gawk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 awk, terms describing . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 awk, uses for. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5, 17, 29 awk, versions of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 413 awk, versions of, changes between SVR3.1 and

SVR4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 414 awk, versions of, changes between SVR4 and

POSIX awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 414 awk, versions of, changes between V7 and SVR3.1

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 413 awk, versions of, See Also Brian Kernighan’s awk

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 415, 444 awka compiler for awk. . . . . . . . . . . . . . . . . . . . . . . . . 444 AWKLIBPATH environment variable . . . . . . . . . . . . . . 40 AWKPATH environment variable . . . . . . . . . . . . . 39, 437 awkprof.out file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 awksed.awk program . . . . . . . . . . . . . . . . . . . . . . . . . 284 awkvars.out file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

B

b debugger command (alias for break) . . . . . . . . 330 backslash (\) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 backslash (\), \" escape sequence . . . . . . . . . . . . . . 47 backslash (\), \' operator (gawk) . . . . . . . . . . . . . . 55 backslash (\), \/ escape sequence . . . . . . . . . . . . . . 47 backslash (\), \< operator (gawk) . . . . . . . . . . . . . . 54 backslash (\), \> operator (gawk) . . . . . . . . . . . . . . 54 backslash (\), \` operator (gawk) . . . . . . . . . . . . . . 55 backslash (\), \a escape sequence . . . . . . . . . . . . . . 46 backslash (\), \b escape sequence . . . . . . . . . . . . . . 46 backslash (\), \B operator (gawk) . . . . . . . . . . . . . . 54 backslash (\), \f escape sequence . . . . . . . . . . . . . . 46 backslash (\), \n escape sequence . . . . . . . . . . . . . . 46 backslash (\), \nnn escape sequence . . . . . . . . . . . 46 backslash (\), \r escape sequence . . . . . . . . . . . . . . 46 backslash (\), \s operator (gawk) . . . . . . . . . . . . . . 54 backslash (\), \S operator (gawk) . . . . . . . . . . . . . . 54 backslash (\), \t escape sequence . . . . . . . . . . . . . . 46 backslash (\), \v escape sequence . . . . . . . . . . . . . . 46 backslash (\), \w operator (gawk) . . . . . . . . . . . . . . 54 backslash (\), \W operator (gawk) . . . . . . . . . . . . . . 54 backslash (\), \x escape sequence . . . . . . . . . . . . . . 46 backslash (\), \y operator (gawk) . . . . . . . . . . . . . . 54 backslash (\), as field separator . . . . . . . . . . . . . . . . 70 backslash (\), continuing lines and . . . . . . . . . . . . . 27 backslash (\), continuing lines and, comments and

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 backslash (\), continuing lines and, in csh. . . . . . 28 backslash (\), gsub()/gensub()/sub() functions

and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 backslash (\), in bracket expressions . . . . . . . . . . . 51 backslash (\), in escape sequences . . . . . . . . . . 46, 47 backslash (\), in escape sequences, POSIX and

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 backslash (\), in regexp constants . . . . . . . . . . . . . . 53 backslash (\), in shell commands. . . . . . . . . . . . . . . 21 backslash (\), regexp operator . . . . . . . . . . . . . . . . . 48 backtrace debugger command . . . . . . . . . . . . . . . . 335 Beebe, Nelson H.F. . . . . . . . . . . . . . . . . . . . . . . . 12, 445 BEGIN pattern. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67, 134 BEGIN pattern, and profiling . . . . . . . . . . . . . . . . . . 309 BEGIN pattern, assert() user-defined function and

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 BEGIN pattern, Boolean patterns and. . . . . . . . . . 133 BEGIN pattern, exit statement and . . . . . . . . . . . 146 BEGIN pattern, getline and . . . . . . . . . . . . . . . . . . . 83 BEGIN pattern, headings, adding . . . . . . . . . . . . . . . 90 BEGIN pattern, next/nextfile statements and

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135, 145 BEGIN pattern, OFS/ORS variables, assigning values

to. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 BEGIN pattern, operators and . . . . . . . . . . . . . . . . . 134 BEGIN pattern, print statement and . . . . . . . . . . 135 BEGIN pattern, pwcat program . . . . . . . . . . . . . . . . 241 BEGIN pattern, running awk programs and . . . . . 252 BEGIN pattern, TEXTDOMAIN variable and . . . . . . 318 BEGINFILE pattern. . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 BEGINFILE pattern, Boolean patterns and . . . . . 133 beginfile() user-defined function . . . . . . . . . . . . 230 Bentley, Jon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 466 Benzinger, Michael . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 Berry, Karl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 425 binary input/output . . . . . . . . . . . . . . . . . . . . . . . . . . 147 bindtextdomain() function (C library) . . . . . . . 316 bindtextdomain() function (gawk) . . . . . . . 199, 317 bindtextdomain() function (gawk), portability and

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 BINMODE variable. . . . . . . . . . . . . . . . . . . . . . . . . 147, 437 bit-manipulation functions . . . . . . . . . . . . . . . . . . . . 196 bits2str() user-defined function . . . . . . . . . . . . . 197 bitwise AND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 bitwise complement. . . . . . . . . . . . . . . . . . . . . . . . . . . 196 bitwise OR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 bitwise XOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 bitwise, complement . . . . . . . . . . . . . . . . . . . . . . . . . . 196 bitwise, operations. . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 bitwise, shift . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 body, in actions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 body, in loops . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 Boolean expressions . . . . . . . . . . . . . . . . . . . . . . . . . . 124 Boolean expressions, as patterns . . . . . . . . . . . . . . 132 Boolean operators, See Boolean expressions … 124 Bourne shell, quoting rules for . . . . . . . . . . . . . . . . . 21 braces ({}) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311 braces ({}), actions and . . . . . . . . . . . . . . . . . . . . . . 137 braces ({}), statements, grouping . . . . . . . . . . . . . 138 bracket expressions . . . . . . . . . . . . . . . . . . . . . . . . 49, 50 bracket expressions, character classes. . . . . . . . . . . 51 bracket expressions, collating elements . . . . . . . . . 52 bracket expressions, collating symbols . . . . . . . . . . 52 bracket expressions, complemented. . . . . . . . . . . . . 49 bracket expressions, equivalence classes . . . . . . . . 52 bracket expressions, non-ASCII . . . . . . . . . . . . . . . . 52 bracket expressions, range expressions. . . . . . . . . . 50 break debugger command . . . . . . . . . . . . . . . . . . . . 330 break statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 breakpoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 breakpoint at location, how to delete . . . . . . . . . 331 breakpoint commands . . . . . . . . . . . . . . . . . . . . . . . . 332 breakpoint condition . . . . . . . . . . . . . . . . . . . . . . . . . 331 breakpoint, delete by number . . . . . . . . . . . . . . . . . 331 breakpoint, how to disable or enable . . . . . . . . . . 331 breakpoint, setting . . . . . . . . . . . . . . . . . . . . . . . . . . . 330 Brennan, Michael . . . . . . 2, 3, 12, 169, 284, 443, 444 Brian Kernighan’s awk . . . . . 29, 47, 55, 69, 82, 115,

135, 143, 144, 146, 169, 185, 186, 189 Brian Kernighan’s awk, extensions . . . . . . . . . . . . 415 Brian Kernighan’s awk, source code . . . . . . . . . . . 444 Brini, Davide . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293 Brink, Jeroen. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 Broder, Alan J.. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 Brown, Martin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 BSD-based operating systems. . . . . . . . . . . . . . . . . 474 bt debugger command (alias for backtrace) . . 335 Buening, Andreas . . . . . . . . . . . . . . . . . . . . 12, 426, 443 buffering, input/output . . . . . . . . . . . . . . . . . . 191, 306 buffering, interactive vs. noninteractive . . . . . . . 190 buffers, flushing. . . . . . . . . . . . . . . . . . . . . . . . . . 189, 191 buffers, operators for . . . . . . . . . . . . . . . . . . . . . . . . . . 54 bug reports, email address, bug-gawk@gnu.org

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 443 bug-gawk@gnu.org bug reporting address . . . . . 443 built-in functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 built-in functions, evaluation order. . . . . . . . . . . . 175 BusyBox Awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 445

C

c.e., See common extensions . . . . . . . . . . . . . . . . . . . 10 call by reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 call by value. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 call stack, display in debugger . . . . . . . . . . . . . . . . 335 caret (^), ^operator . . . . . . . . . . . . . . . . . . . . . . . . . . 128 caret (^), ^= operator . . . . . . . . . . . . . . . . . . . . 118, 129 caret (^), in bracket expressions. . . . . . . . . . . . . . . . 51 caret (^), regexp operator . . . . . . . . . . . . . . . . . . 48, 55 case keyword . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 case sensitivity, and regexps . . . . . . . . . . . . . . . . . . 148 case sensitivity, and string comparisons . . . . . . . 148 case sensitivity, array indices and . . . . . . . . . . . . . 161 case sensitivity, converting case . . . . . . . . . . . . . . . 186 case sensitivity, example programs . . . . . . . . . . . . 217 case sensitivity, gawk. . . . . . . . . . . . . . . . . . . . . . . . . . . 55 case sensitivity, regexps and . . . . . . . . . . . . . . . . . . . 55 CGI, awk scripts for . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 character classes, See bracket expressions. . . . . . . 49 character lists in regular expression . . . . . . . . . . . . 50 character lists, See bracket expressions . . . . . . . . . 49 character sets (machine character encodings)

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224, 465 character sets, See Also bracket expressions …. 49 characters, counting . . . . . . . . . . . . . . . . . . . . . . . . . . 269 characters, transliterating. . . . . . . . . . . . . . . . . . . . . 274 characters, values of as numbers . . . . . . . . . . . . . . 223 Chassell, Robert J. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 chdir() extension function . . . . . . . . . . . . . . . . . . . 398 chem utility. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 466 chr() extension function . . . . . . . . . . . . . . . . . . . . . 403 chr() user-defined function . . . . . . . . . . . . . . . . . . . 223 clear debugger command . . . . . . . . . . . . . . . . . . . . 331 Cliff random numbers . . . . . . . . . . . . . . . . . . . . . . . . 223 cliff_rand() user-defined function . . . . . . . . . . . 223 close file or coprocess . . . . . . . . . . . . . . . . . . . . . . . . . 189 close() function . . . . . . . . . . . . . . . . . . . . . . . . 102, 189 close() function, portability . . . . . . . . . . . . . . . . . 103 close() function, return value . . . . . . . . . . . . . . . . 104 close() function, two-way pipes and . . . . . . . . . 306 Close, Diane . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 425 Collado, Manuel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 collating elements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 collating symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Colombo, Antonio . . . . . . . . . . . . . . . . . . . . . . . . 12, 427 columns, aligning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 columns, cutting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 comma (,), in range patterns . . . . . . . . . . . . . . . . . 133 command completion, in debugger . . . . . . . . . . . . 338 command line, arguments . . . . . . . . . . . . 37, 149, 155 command line, directories on. . . . . . . . . . . . . . . . . . . 86 command line, formats . . . . . . . . . . . . . . . . . . . . . . . . 17 command line, FS on, setting . . . . . . . . . . . . . . . . . . 69 command line, invoking awk from . . . . . . . . . . . . . . 31 command line, option -f. . . . . . . . . . . . . . . . . . . . . . . 18 command line, options . . . . . . . . . . . . . . . . . . . . . . . . . 31 command line, options, end of . . . . . . . . . . . . . . . . . 32 command line, variables, assigning on . . . . . . . . . 111 command-line options, processing . . . . . . . . . . . . . 233 command-line options, string extraction. . . . . . . 319 commands debugger command . . . . . . . . . . . . . . . . . 332 commands to execute at breakpoint. . . . . . . . . . . 332 commenting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 commenting, backslash continuation and . . . . . . . 28 common extensions, ** operator . . . . . . . . . . . . . . 114 common extensions, **= operator . . . . . . . . . . . . . 118 common extensions, /dev/stderr special file . . 101 common extensions, /dev/stdin special file … 101 common extensions, /dev/stdout special file . . 101 common extensions, \x escape sequence . . . . . . . . 46 common extensions, BINMODE variable . . . . . . . . . 437 common extensions, delete to delete entire arrays

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 common extensions, func keyword . . . . . . . . . . . . 200 common extensions, length() applied to an array

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 common extensions, RS as a regexp . . . . . . . . . . . . 61 common extensions, single character fields . . . . . 69 comp.lang.awk newsgroup . . . . . . . . . . . . . . . . . . . . 443 comparison expressions . . . . . . . . . . . . . . . . . . . . . . . 120 comparison expressions, as patterns . . . . . . . . . . . 132 comparison expressions, string vs. regexp . . . . . 123 compatibility mode (gawk), extensions . . . . . . . . 415 compatibility mode (gawk), file names. . . . . . . . . 102 compatibility mode (gawk), hexadecimal numbers

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 compatibility mode (gawk), octal numbers …. 108 compatibility mode (gawk), specifying . . . . . . . . . . 32 compiled programs. . . . . . . . . . . . . . . . . . . . . . . 459, 466 compiling gawk for Cygwin . . . . . . . . . . . . . . . . . . . 438 compiling gawk for MS-DOS and MS-Windows

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 compiling gawk for VMS . . . . . . . . . . . . . . . . . . . . . . 439 compiling gawk with EMX for OS/2 . . . . . . . . . . 436 compl() function (gawk) . . . . . . . . . . . . . . . . . . . . . . 196 complement, bitwise . . . . . . . . . . . . . . . . . . . . . . . . . . 196 compound statements, control statements and

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 concatenating . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 condition debugger command . . . . . . . . . . . . . . . . 331 conditional expressions . . . . . . . . . . . . . . . . . . . . . . . 125 configuration option, --disable-extensions . . 433 configuration option, --disable-lint. . . . . . . . . 433 configuration option, --disable-nls . . . . . . . . . . 434 configuration option,

                                                    --with-whiny-user-strftime. . . . . . . . . . . . 434

configuration options, gawk . . . . . . . . . . . . . . . . . . . 433 constant regexps. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 constants, nondecimal . . . . . . . . . . . . . . . . . . . . . . . . 299 constants, numeric. . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 constants, types of. . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 continue program, in debugger . . . . . . . . . . . . . . . . 332 continue statement . . . . . . . . . . . . . . . . . . . . . . . . . . 143 control statements . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 controlling array scanning order . . . . . . . . . . . . . . 164 convert string to lower case . . . . . . . . . . . . . . . . . . . 186 convert string to number . . . . . . . . . . . . . . . . . . . . . 183 convert string to upper case . . . . . . . . . . . . . . . . . . 186 converting integer array subscripts . . . . . . . . . . . . 167 converting, dates to timestamps . . . . . . . . . . . . . . 193 converting, numbers to strings . . . . . . . . . . . 111, 198 converting, strings to numbers . . . . . . . . . . . 111, 198 CONVFMT variable. . . . . . . . . . . . . . . . . . . . . . . . . 112, 147 CONVFMT variable, and array subscripts . . . . . . . . 167 cookie . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 467 coprocesses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99, 306 coprocesses, closing . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 coprocesses, getline from . . . . . . . . . . . . . . . . . . . . . 82 cos() function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 cosine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 counting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 csh utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 csh utility, |& operator, comparison with. . . . . . 306 csh utility, POSIXLY_CORRECT environment variable

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 ctime() user-defined function. . . . . . . . . . . . . . . . . 202 currency symbols, localization . . . . . . . . . . . . . . . . 316 current system time . . . . . . . . . . . . . . . . . . . . . . . . . . 193 custom.h file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 434 customized input parser . . . . . . . . . . . . . . . . . . . . . . 365 customized output wrapper . . . . . . . . . . . . . . . . . . . 368 customized two-way processor . . . . . . . . . . . . . . . . 370 cut utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 cut.awk program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252

D

d debugger command (alias for delete) . . . . . . . 331 d.c., See dark corner . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 dark corner . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 467 dark corner, "0" is actually true . . . . . . . . . . . . . . 120 dark corner, /= operator vs. /=.../ regexp

constant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 dark corner, ^, in FS . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 dark corner, array subscripts. . . . . . . . . . . . . . . . . . 168 dark corner, break statement . . . . . . . . . . . . . . . . . 143 dark corner, close() function . . . . . . . . . . . . . . . . 104 dark corner, command-line arguments. . . . . . . . . 111 dark corner, continue statement. . . . . . . . . . . . . . 144 dark corner, CONVFMT variable . . . . . . . . . . . . . . . . . 112 dark corner, escape sequences . . . . . . . . . . . . . . . . . . 38 dark corner, escape sequences, for metacharacters

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 dark corner, exit statement . . . . . . . . . . . . . . . . . . 146 dark corner, field separators . . . . . . . . . . . . . . . . . . . 71 dark corner, FILENAME variable . . . . . . . . . . . . 83, 150 dark corner, FNR/NR variables . . . . . . . . . . . . . . . . . 155 dark corner, format-control characters . . . . . . 93, 94 dark corner, FS as null string . . . . . . . . . . . . . . . . . . 69 dark corner, input files. . . . . . . . . . . . . . . . . . . . . . . . . 61 dark corner, invoking awk . . . . . . . . . . . . . . . . . . . . . . 31 dark corner, length() function . . . . . . . . . . . . . . . 180 dark corner, locale’s decimal point character . . 112 dark corner, multiline records . . . . . . . . . . . . . . . . . . 76 dark corner, NF variable, decrementing . . . . . . . . . 66 dark corner, OFMT variable . . . . . . . . . . . . . . . . . . . . . 92 dark corner, regexp as second argument to index()

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 dark corner, regexp constants . . . . . . . . . . . . . . . . . 109 dark corner, regexp constants, /= operator and

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 dark corner, regexp constants, as arguments to

user-defined functions . . . . . . . . . . . . . . . . . . . . 109 dark corner, split() function . . . . . . . . . . . . . . . . 183 dark corner, strings, storing. . . . . . . . . . . . . . . . . . . . 63 dark corner, value of ARGV[0] . . . . . . . . . . . . . . . . . 150 data, fixed-width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 data-driven languages . . . . . . . . . . . . . . . . . . . . . . . . 460 database, group, reading. . . . . . . . . . . . . . . . . . . . . . 242 database, users, reading . . . . . . . . . . . . . . . . . . . . . . 238 date utility, GNU . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 date utility, POSIX . . . . . . . . . . . . . . . . . . . . . . . . . . 195 dates, converting to timestamps . . . . . . . . . . . . . . 193 dates, information related to, localization . . . . . 317 Davies, Stephen. . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 426 Day, Robert P.J.. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 dcgettext() function (gawk). . . . . . . . . . . . . 199, 317 dcgettext() function (gawk), portability and

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 dcngettext() function (gawk) . . . . . . . . . . . 199, 317 dcngettext() function (gawk), portability and

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 deadlocks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 debugger commands, b (break) . . . . . . . . . . . . . . . 330 debugger commands, backtrace . . . . . . . . . . . . . . 335 debugger commands, break . . . . . . . . . . . . . . . . . . . 330 debugger commands, bt (backtrace) . . . . . . . . . 335 debugger commands, c (continue). . . . . . . . . . . . 332 debugger commands, clear . . . . . . . . . . . . . . . . . . . 331 debugger commands, commands . . . . . . . . . . . . . . . 332 debugger commands, condition . . . . . . . . . . . . . . 331 debugger commands, continue . . . . . . . . . . . . . . . 332 debugger commands, d (delete) . . . . . . . . . . . . . . 331 debugger commands, delete. . . . . . . . . . . . . . . . . . 331 debugger commands, disable . . . . . . . . . . . . . . . . 331 debugger commands, display . . . . . . . . . . . . . . . . 333 debugger commands, down . . . . . . . . . . . . . . . . . . . . 335 debugger commands, dump . . . . . . . . . . . . . . . . . . . . 337 debugger commands, e (enable) . . . . . . . . . . . . . . 331 debugger commands, enable. . . . . . . . . . . . . . . . . . 331 debugger commands, end . . . . . . . . . . . . . . . . . . . . . 332 debugger commands, eval . . . . . . . . . . . . . . . . . . . . 333 debugger commands, f (frame) . . . . . . . . . . . . . . . 335 debugger commands, finish. . . . . . . . . . . . . . . . . . 332 debugger commands, frame . . . . . . . . . . . . . . . . . . . 335 debugger commands, h (help) . . . . . . . . . . . . . . . . 338 debugger commands, help . . . . . . . . . . . . . . . . . . . . 338 debugger commands, i (info) . . . . . . . . . . . . . . . . 335 debugger commands, ignore. . . . . . . . . . . . . . . . . . 332 debugger commands, info . . . . . . . . . . . . . . . . . . . . 335 debugger commands, l (list) . . . . . . . . . . . . . . . . 338 debugger commands, list . . . . . . . . . . . . . . . . . . . . 338 debugger commands, n (next) . . . . . . . . . . . . . . . . 332 debugger commands, next . . . . . . . . . . . . . . . . . . . . 332 debugger commands, nexti . . . . . . . . . . . . . . . . . . . 332 debugger commands, ni (nexti) . . . . . . . . . . . . . . 332 debugger commands, o (option) . . . . . . . . . . . . . . 336 debugger commands, option. . . . . . . . . . . . . . . . . . 336 debugger commands, p (print) . . . . . . . . . . . . . . . 333 debugger commands, print . . . . . . . . . . . . . . . . . . . 333 debugger commands, printf. . . . . . . . . . . . . . . . . . 334 debugger commands, q (quit) . . . . . . . . . . . . . . . . 338 debugger commands, quit . . . . . . . . . . . . . . . . . . . . 338 debugger commands, r (run) . . . . . . . . . . . . . . . . . 332 debugger commands, return. . . . . . . . . . . . . . . . . . 332 debugger commands, run . . . . . . . . . . . . . . . . . . . . . 332 debugger commands, s (step) . . . . . . . . . . . . . . . . 333 debugger commands, set . . . . . . . . . . . . . . . . . . . . . 334 debugger commands, si (stepi) . . . . . . . . . . . . . . 333 debugger commands, silent. . . . . . . . . . . . . . . . . . 332 debugger commands, step . . . . . . . . . . . . . . . . . . . . 333 debugger commands, stepi . . . . . . . . . . . . . . . . . . . 333 debugger commands, t (tbreak) . . . . . . . . . . . . . . 332 debugger commands, tbreak. . . . . . . . . . . . . . . . . . 332 debugger commands, trace . . . . . . . . . . . . . . . . . . . 338 debugger commands, u (until) . . . . . . . . . . . . . . . 333 debugger commands, undisplay . . . . . . . . . . . . . . 334 debugger commands, until . . . . . . . . . . . . . . . . . . . 333 debugger commands, unwatch . . . . . . . . . . . . . . . . 334 debugger commands, up . . . . . . . . . . . . . . . . . . . . . . 335 debugger commands, w (watch) . . . . . . . . . . . . . . . 334 debugger commands, watch . . . . . . . . . . . . . . . . . . . 334 debugger commands, where (backtrace) . . . . . . 335 debugger default list amount. . . . . . . . . . . . . . . . . . 336 debugger history file . . . . . . . . . . . . . . . . . . . . . . . . . . 336 debugger history size . . . . . . . . . . . . . . . . . . . . . . . . . 336 debugger options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 336 debugger prompt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 336 debugger, how to start. . . . . . . . . . . . . . . . . . . . . . . . 327 debugger, read commands from a file . . . . . . . . . 336 debugging awk programs . . . . . . . . . . . . . . . . . . . . . . 325 debugging gawk, bug reports . . . . . . . . . . . . . . . . . . 442 decimal point character, locale specific . . . . . . . . . 36 decrement operators . . . . . . . . . . . . . . . . . . . . . . . . . . 119 default keyword . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Deifik, Scott . . . . . . . . . . . . . . . . . . . . . . . . . 12, 426, 443 delete array . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 delete breakpoint at location. . . . . . . . . . . . . . . . . . 331 delete breakpoint by number. . . . . . . . . . . . . . . . . . 331 delete debugger command . . . . . . . . . . . . . . . . . . . 331 delete statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 delete watchpoint. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 334 deleting elements in arrays. . . . . . . . . . . . . . . . . . . . 168 deleting entire arrays . . . . . . . . . . . . . . . . . . . . . . . . . 169 Demaille, Akim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 describe call stack frame, in debugger . . . . . . . . . 335 differences between gawk and awk . . . . . . . . . . . . . 180 differences in awk and gawk, ARGC/ARGV variables

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 differences in awk and gawk, ARGIND variable . . 150 differences in awk and gawk, array elements,

deleting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169

                                                differences in awk and gawk, AWKLIBPATH

environment variable . . . . . . . . . . . . . . . . . . . . . . 40 differences in awk and gawk, AWKPATH environment

variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 differences in awk and gawk, BEGIN/END patterns

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

                                                differences in awk and gawk, BEGINFILE/ENDFILE

patterns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 differences in awk and gawk, BINMODE variable

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147, 437 differences in awk and gawk, close() function

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103, 104 differences in awk and gawk, command-line

directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 differences in awk and gawk, ERRNO variable. … 150 differences in awk and gawk, error messages … 100 differences in awk and gawk, FIELDWIDTHS variable

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 differences in awk and gawk, FPAT variable . . . . . 147 differences in awk and gawk, FUNCTAB variable . . 151 differences in awk and gawk, function arguments

(gawk) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 differences in awk and gawk, getline command

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 differences in awk and gawk, IGNORECASE variable

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 differences in awk and gawk, implementation

limitations . . . . . . . . . . . . . . . . . . . . . . . . . . . 83, 100 differences in awk and gawk, indirect function calls

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 differences in awk and gawk, input/output

operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82, 99 differences in awk and gawk, line continuations

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 differences in awk and gawk, LINT variable . . . . . 148 differences in awk and gawk, match() function

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181

                                                differences in awk and gawk, print/printf

statements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 differences in awk and gawk, PROCINFO array … 151 differences in awk and gawk, read timeouts . . . . . 84 differences in awk and gawk, record separators . . 61 differences in awk and gawk, regexp constants . . 109 differences in awk and gawk, regular expressions

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 differences in awk and gawk, RS/RT variables …. 62 differences in awk and gawk, RT variable . . . . . . . 153 differences in awk and gawk, single-character fields

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 differences in awk and gawk, split() function

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 differences in awk and gawk, strings . . . . . . . . . . . 107 differences in awk and gawk, strings, storing …. 63 differences in awk and gawk, SYMTAB variable . . 153 differences in awk and gawk, TEXTDOMAIN variable

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 differences in awk and gawk, trunc-mod operation

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 directories, command-line . . . . . . . . . . . . . . . . . . . . . . 86 directories, searching . . . . . . . . . . . . . . . . . . . . . . . . . 295 directories, searching for loadable extensions … 40 directories, searching for source files. . . . . . . . . . . . 39 disable breakpoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . 331 disable debugger command . . . . . . . . . . . . . . . . . . 331 display debugger command . . . . . . . . . . . . . . . . . . 333 display debugger options. . . . . . . . . . . . . . . . . . . . . . 336 division . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 do-while statement. . . . . . . . . . . . . . . . . . . . . . . . . . . 140 do-while statement, use of regexps in . . . . . . . . . . 45 documentation, of awk programs . . . . . . . . . . . . . . 218 documentation, online . . . . . . . . . . . . . . . . . . . . . . . . . 10 documents, searching . . . . . . . . . . . . . . . . . . . . . . . . . 271 dollar sign ($), $ field operator . . . . . . . . . . . . 63, 128 dollar sign ($), incrementing fields and arrays

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 dollar sign ($), regexp operator . . . . . . . . . . . . . . . . 48 double quote ("), in regexp constants . . . . . . . . . . 53 double quote ("), in shell commands . . . . . . . . . . . 22 down debugger command . . . . . . . . . . . . . . . . . . . . . 335 Drepper, Ulrich. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Duman, Patrice . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 dump all variables of a program. . . . . . . . . . . . . . . . 33 dump debugger command . . . . . . . . . . . . . . . . . . . . . 337 dupword.awk program . . . . . . . . . . . . . . . . . . . . . . . . 271 dynamic profiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311 dynamically loaded extensions . . . . . . . . . . . . . . . . 355

E

e debugger command (alias for enable) . . . . . . . 331 EBCDIC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 effective group ID of gawk user. . . . . . . . . . . . . . . . 151 effective user ID of gawk user . . . . . . . . . . . . . . . . . 151 egrep utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51, 256 egrep.awk program. . . . . . . . . . . . . . . . . . . . . . . . . . . 257 elements in arrays, assigning values . . . . . . . . . . . 162 elements in arrays, deleting . . . . . . . . . . . . . . . . . . . 168 elements in arrays, order of access by in operator

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 elements in arrays, scanning . . . . . . . . . . . . . . . . . . 163 elements of arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 email address for bug reports, bug-gawk@gnu.org

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 443 EMISTERED. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 empty array elements . . . . . . . . . . . . . . . . . . . . . . . . . 161 empty pattern . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 empty strings. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 empty strings, See null strings . . . . . . . . . . . . . . . . . 68 enable breakpoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . 331 enable debugger command . . . . . . . . . . . . . . . . . . . 331 end debugger command. . . . . . . . . . . . . . . . . . . . . . . 332 END pattern . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 END pattern, and profiling. . . . . . . . . . . . . . . . . . . . . 309 END pattern, assert() user-defined function and

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 END pattern, Boolean patterns and . . . . . . . . . . . . 133 END pattern, exit statement and. . . . . . . . . . . . . . 146 END pattern, next/nextfile statements and . . 135,

145 END pattern, operators and. . . . . . . . . . . . . . . . . . . . 134 END pattern, print statement and. . . . . . . . . . . . . 135 ENDFILE pattern . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 ENDFILE pattern, Boolean patterns and . . . . . . . 133 endfile() user-defined function . . . . . . . . . . . . . . 230 endgrent() function (C library) . . . . . . . . . . . . . . 246 endgrent() user-defined function . . . . . . . . . . . . . 246 endpwent() function (C library) . . . . . . . . . . . . . . 242 endpwent() user-defined function . . . . . . . . . . . . . 242 English, Steve . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 ENVIRON array . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 environment variables used by gawk . . . . . . . . . . . . 39 environment variables, in ENVIRON array . . . . . . 150 epoch, definition of . . . . . . . . . . . . . . . . . . . . . . . . . . . 467 equals sign (=), = operator . . . . . . . . . . . . . . . . . . . . 116 equals sign (=), == operator . . . . . . . . . . . . . . 122, 128 EREs (Extended Regular Expressions) . . . . . . . . . 51 ERRNO variable . . . . . . . . . . . . . . . . . . . . . . . . . . . 150, 308 ERRNO variable, with BEGINFILE pattern . . . . . . . 136 ERRNO variable, with close() function . . . . . . . . 104 ERRNO variable, with getline command . . . . . . . . 78 error handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 error handling, ERRNO variable and . . . . . . . . . . . . 150 error output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

escape processing, gsub()/gensub()/sub()

functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 escape sequences, in strings . . . . . . . . . . . . . . . . . . . . 46 eval debugger command . . . . . . . . . . . . . . . . . . . . . 333 evaluate expressions, in debugger . . . . . . . . . . . . . 333 evaluation order . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 evaluation order, concatenation . . . . . . . . . . . . . . . 115 evaluation order, functions. . . . . . . . . . . . . . . . . . . . 175 examining fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 exclamation point (!), ! operator … 125, 128, 259 exclamation point (!), != operator . . . . . . . 122, 128 exclamation point (!), !~ operator . . . . . 45, 52, 55,

108, 122, 123, 129, 132 exit statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 exit status, of gawk . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 exit status, of VMS. . . . . . . . . . . . . . . . . . . . . . . . . . . 441 exit the debugger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 338 exp() function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 expand utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Expat XML parser library . . . . . . . . . . . . . . . . . . . . 407 exponent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 expressions, as patterns. . . . . . . . . . . . . . . . . . . . . . . 131 expressions, assignment . . . . . . . . . . . . . . . . . . . . . . . 116 expressions, Boolean. . . . . . . . . . . . . . . . . . . . . . . . . . 124 expressions, comparison . . . . . . . . . . . . . . . . . . . . . . 120 expressions, conditional. . . . . . . . . . . . . . . . . . . . . . . 125 expressions, matching, See comparison expressions

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 expressions, selecting . . . . . . . . . . . . . . . . . . . . . . . . . 125 Extended Regular Expressions (EREs) . . . . . . . . . 51 extension API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 357 extension API informational variables. . . . . . . . . 386 extension API version . . . . . . . . . . . . . . . . . . . . . . . . 385 extension API, version number . . . . . . . . . . . . . . . 153 extension example . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 extension registration . . . . . . . . . . . . . . . . . . . . . . . . . 363 extension search path. . . . . . . . . . . . . . . . . . . . . . . . . 388 extensions distributed with gawk . . . . . . . . . . . . . . 398 extensions, allocating memory . . . . . . . . . . . . . . . . 362 extensions, Brian Kernighan’s awk. . . . . . . . 415, 423 extensions, common, ** operator . . . . . . . . . . . . . 114 extensions, common, **= operator . . . . . . . . . . . . 118 extensions, common, /dev/stderr special file

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 extensions, common, /dev/stdin special file . . 101 extensions, common, /dev/stdout special file

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 extensions, common, \x escape sequence . . . . . . . 46 extensions, common, BINMODE variable . . . . . . . . 437 extensions, common, delete to delete entire arrays

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 extensions, common, fflush() function . . . . . . . 189 extensions, common, func keyword . . . . . . . . . . . 200 extensions, common, length() applied to an array

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 extensions, common, RS as a regexp . . . . . . . . . . . . 61 extensions, common, single character fields . . . . . 69 extensions, in gawk, not in POSIX awk . . . . . . . . 415 extensions, loading, @load directive . . . . . . . . . . . . 43 extensions, mawk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 423 extensions, where to find . . . . . . . . . . . . . . . . . . . . . 407 extract.awk program . . . . . . . . . . . . . . . . . . . . . . . . 282 extraction, of marked strings (internationalization)

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319

F

f debugger command (alias for frame) . . . . . . . . 335 false, logical . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 FDL (Free Documentation License) . . . . . . . . . . . 487 features, adding to gawk . . . . . . . . . . . . . . . . . . . . . . 448 features, deprecated . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 features, undocumented. . . . . . . . . . . . . . . . . . . . . . . . 44 Fenlason, Jay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 425 fflush() function . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 field numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 field operator $ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 field operators, dollar sign as. . . . . . . . . . . . . . . . . . . 63 field separator, in multiline records. . . . . . . . . . . . . 76 field separator, on command line . . . . . . . . . . . . . . . 69 field separator, POSIX and . . . . . . . . . . . . . . . . . . . . 71 field separators. . . . . . . . . . . . . . . . . . . . . . . 67, 147, 148 field separators, choice of . . . . . . . . . . . . . . . . . . . . . . 67 field separators, FIELDWIDTHS variable and …. 147 field separators, FPAT variable and . . . . . . . . . . . . 147 field separators, POSIX and . . . . . . . . . . . . . . . . . . . 63 field separators, regular expressions as . . . . . . 67, 68 field separators, See Also OFS . . . . . . . . . . . . . . . . . . 66 field separators, spaces as . . . . . . . . . . . . . . . . . . . . . 253 fields. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59, 63, 460 fields, adding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 fields, changing contents of. . . . . . . . . . . . . . . . . . . . . 65 fields, cutting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 fields, examining. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 fields, number of. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 fields, numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 fields, printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 fields, separating. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 fields, single-character . . . . . . . . . . . . . . . . . . . . . . . . . 69 FIELDWIDTHS variable . . . . . . . . . . . . . . . . . . . . . 72, 147 file descriptors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 file inclusion, @include directive . . . . . . . . . . . . . . . 42 file names, distinguishing . . . . . . . . . . . . . . . . . . . . . 150 file names, in compatibility mode . . . . . . . . . . . . . 102 file names, standard streams in gawk . . . . . . . . . . 101 FILENAME variable. . . . . . . . . . . . . . . . . . . . . . . . . 59, 150 FILENAME variable, getline, setting with . . . . . . . 83 filenames, assignments as . . . . . . . . . . . . . . . . . . . . . 232 files, .gmo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316 files, .gmo, specifying directory of . . . . . . . . 316, 317 files, .mo, converting from .po . . . . . . . . . . . . . . . . 322 files, .po. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315, 319 files, .po, converting to .mo . . . . . . . . . . . . . . . . . . . 322 files, .pot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 files, /dev/... special files . . . . . . . . . . . . . . . . . . . . 101 files, /inet/... (gawk) . . . . . . . . . . . . . . . . . . . . . . . 307 files, /inet4/... (gawk) . . . . . . . . . . . . . . . . . . . . . . 307 files, /inet6/... (gawk) . . . . . . . . . . . . . . . . . . . . . . 307 files, awk programs in . . . . . . . . . . . . . . . . . . . . . . . . . . 18 files, awkprof.out . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 files, awkvars.out . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 files, closing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 files, descriptors, See file descriptors . . . . . . . . . . 100 files, group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 files, initialization and cleanup . . . . . . . . . . . . . . . . 229 files, input, See input files. . . . . . . . . . . . . . . . . . . . . . 18 files, log, timestamps in. . . . . . . . . . . . . . . . . . . . . . . 192 files, managing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 files, managing, data file boundaries . . . . . . . . . . 229 files, message object . . . . . . . . . . . . . . . . . . . . . . . . . . 316 files, message object, converting from portable

object files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 files, message object, specifying directory of . . 316,

317 files, multiple passes over . . . . . . . . . . . . . . . . . . . . . . 38 files, multiple, duplicating output into . . . . . . . . 263 files, output, See output files. . . . . . . . . . . . . . . . . . 102 files, password . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 files, portable object . . . . . . . . . . . . . . . . . . . . . 315, 319 files, portable object template . . . . . . . . . . . . . . . . 315 files, portable object, converting to message object

files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 files, portable object, generating . . . . . . . . . . . . . . . 34 files, processing, ARGIND variable and. . . . . . . . . . 150 files, reading. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 files, reading, multiline records . . . . . . . . . . . . . . . . . 76 files, searching for regular expressions . . . . . . . . . 256 files, skipping. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 files, source, search path for . . . . . . . . . . . . . . . . . . 295 files, splitting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 files, Texinfo, extracting programs from . . . . . . . 281 find substring in string . . . . . . . . . . . . . . . . . . . . . . . 180 finding extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 finish debugger command . . . . . . . . . . . . . . . . . . . 332 Fish, Fred . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 fixed-width data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 flag variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125, 263 floating-point, numbers, arbitrary precision … 341 floating-point, VAX/VMS . . . . . . . . . . . . . . . . . . . . 442 flush buffered output . . . . . . . . . . . . . . . . . . . . . . . . . 189 fnmatch() extension function . . . . . . . . . . . . . . . . . 401 FNR variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59, 151 FNR variable, changing . . . . . . . . . . . . . . . . . . . . . . . . 155 for statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 for statement, looping over arrays . . . . . . . . . . . . 163 fork() extension function . . . . . . . . . . . . . . . . . . . . 402 format specifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 format specifiers, mixing regular with positional

specifiers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320 format specifiers, printf statement . . . . . . . . . . . . 93 format specifiers, strftime() function (gawk)

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 format time string . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 formats, numeric output . . . . . . . . . . . . . . . . . . . . . . . 92 formatting output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 formatting strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 forward slash (/) to enclose regular expressions

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 forward slash (/), / operator. . . . . . . . . . . . . . . . . . 128 forward slash (/), /= operator . . . . . . . . . . . . 118, 129 forward slash (/), /= operator, vs. /=.../ regexp

constant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 forward slash (/), patterns and . . . . . . . . . . . . . . . 132 FPAT variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74, 147 frame debugger command . . . . . . . . . . . . . . . . . . . . 335 Free Documentation License (FDL) . . . . . . . . . . . 487 Free Software Foundation (FSF) . . . . . 10, 429, 468,

469 FreeBSD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 FS variable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67, 147 FS variable, --field-separator option and …. 31 FS variable, as null string . . . . . . . . . . . . . . . . . . . . . . 69 FS variable, as TAB character. . . . . . . . . . . . . . . . . . 36 FS variable, changing value of . . . . . . . . . . . . . . . . . . 67 FS variable, running awk programs and . . . . . . . . 252 FS variable, setting from command line . . . . . . . . 69 FS, containing ^. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 FS, in multiline records . . . . . . . . . . . . . . . . . . . . . . . . 76 FSF (Free Software Foundation) . . . . . 10, 429, 468,

469 fts() extension function . . . . . . . . . . . . . . . . . . . . . 399 FUNCTAB array . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 function calls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 function calls, indirect . . . . . . . . . . . . . . . . . . . . . . . . 208 function calls, indirect, @-notation for . . . . . . . . . 209 function definition example . . . . . . . . . . . . . . . . . . . 201 function pointers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 functions, arrays as parameters to . . . . . . . . . . . . 205 functions, built-in. . . . . . . . . . . . . . . . . . . . . . . . 126, 175 functions, built-in, evaluation order . . . . . . . . . . . 175 functions, defining . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 functions, library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 functions, library, assertions . . . . . . . . . . . . . . . . . . 220 functions, library, associative arrays and . . . . . . 218 functions, library, C library . . . . . . . . . . . . . . . . . . . 233 functions, library, character values as numbers

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 functions, library, Cliff random numbers . . . . . . 223 functions, library, command-line options . . . . . . 233 functions, library, example program for using . . 285 functions, library, group database, reading …. 242 functions, library, managing data files . . . . . . . . . 229 functions, library, managing time . . . . . . . . . . . . . 225 functions, library, merging arrays into strings . . 225 functions, library, rounding numbers . . . . . . . . . . 222 functions, library, user database, reading . . . . . . 238 functions, names of . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 functions, recursive . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 functions, string-translation. . . . . . . . . . . . . . . . . . . 198 functions, undefined . . . . . . . . . . . . . . . . . . . . . . . . . . 206 functions, user-defined . . . . . . . . . . . . . . . . . . . . . . . . 199 functions, user-defined, calling . . . . . . . . . . . . . . . . 202 functions, user-defined, counts, in a profile …. 311 functions, user-defined, library of . . . . . . . . . . . . . 217 functions, user-defined, next/nextfile statements

and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145, 146

G

G-d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Garfinkle, Scott . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 425 gawk program, dynamic profiling . . . . . . . . . . . . . . 311 gawk version. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 gawk, ARGIND variable in . . . . . . . . . . . . . . . . . . . . . . . 37 gawk, awk and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5, 7 gawk, bitwise operations in. . . . . . . . . . . . . . . . . . . . 196 gawk, break statement in . . . . . . . . . . . . . . . . . . . . . 143 gawk, character classes and . . . . . . . . . . . . . . . . . . . . 52 gawk, coding style in. . . . . . . . . . . . . . . . . . . . . . . . . . 448 gawk, command-line options, and regular

expressions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 gawk, configuring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 434 gawk, configuring, options. . . . . . . . . . . . . . . . . . . . . 433 gawk, continue statement in. . . . . . . . . . . . . . . . . . 144 gawk, distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . 430 gawk, ERRNO variable in . . . . . 78, 104, 136, 150, 308 gawk, escape sequences. . . . . . . . . . . . . . . . . . . . . . . . . 47 gawk, extensions, disabling . . . . . . . . . . . . . . . . . . . . . 35 gawk, features, adding . . . . . . . . . . . . . . . . . . . . . . . . 448 gawk, features, advanced . . . . . . . . . . . . . . . . . . . . . . 299 gawk, field separators and. . . . . . . . . . . . . . . . . . . . . 148 gawk, FIELDWIDTHS variable in. . . . . . . . . . . . . 72, 147 gawk, file names in. . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 gawk, format-control characters. . . . . . . . . . . . . 93, 94 gawk, FPAT variable in. . . . . . . . . . . . . . . . . . . . . 74, 147 gawk, FUNCTAB array in . . . . . . . . . . . . . . . . . . . . . . . 151 gawk, function arguments and. . . . . . . . . . . . . . . . . 175 gawk, hexadecimal numbers and . . . . . . . . . . . . . . 108 gawk, IGNORECASE variable in. … 55, 148, 161, 178,

305 gawk, implementation issues . . . . . . . . . . . . . . . . . . 447 gawk, implementation issues, debugging . . . . . . . 447 gawk, implementation issues, downward

compatibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . 447 gawk, implementation issues, limits. . . . . . . . . . . . . 83 gawk, implementation issues, pipes . . . . . . . . . . . . 100 gawk, installing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 gawk, internationalization and, See

internationalization . . . . . . . . . . . . . . . . . . . . . . 315 gawk, interpreter, adding code to. . . . . . . . . . . . . . 397 gawk, interval expressions and. . . . . . . . . . . . . . . . . . 50 gawk, line continuation in . . . . . . . . . . . . . . . . . . . . . 126 gawk, LINT variable in . . . . . . . . . . . . . . . . . . . . . . . . 148 gawk, list of contributors to . . . . . . . . . . . . . . . . . . . 425 gawk, MS-DOS version of . . . . . . . . . . . . . . . . . . . . . 437 gawk, MS-Windows version of. . . . . . . . . . . . . . . . . 437 gawk, newlines in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 gawk, octal numbers and. . . . . . . . . . . . . . . . . . . . . . 108 gawk, OS/2 version of . . . . . . . . . . . . . . . . . . . . . . . . 437 gawk, predefined variables and . . . . . . . . . . . . . . . . 147 gawk, PROCINFO array in . . . . . . . . 151, 153, 192, 307 gawk, regexp constants and . . . . . . . . . . . . . . . . . . . 109 gawk, regular expressions, case sensitivity . . . . . . 55 gawk, regular expressions, operators . . . . . . . . . . . . 54 gawk, regular expressions, precedence . . . . . . . . . . 50 gawk, RT variable in . . . . . . . . . . . . . . . . . . . 61, 78, 153 gawk, See Also awk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 gawk, source code, obtaining . . . . . . . . . . . . . . . . . . 429 gawk, splitting fields and . . . . . . . . . . . . . . . . . . . . . . . 73 gawk, string-translation functions . . . . . . . . . . . . . 198 gawk, SYMTAB array in. . . . . . . . . . . . . . . . . . . . . . . . . 153 gawk, TEXTDOMAIN variable in . . . . . . . . . . . . . . . . . 149 gawk, timestamps. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 gawk, uses for . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 gawk, versions of, information about, printing . . 36 gawk, VMS version of . . . . . . . . . . . . . . . . . . . . . . . . . 439 gawk, word-boundary operator . . . . . . . . . . . . . . . . . 55 gawkextlib . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 gawkextlib project . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 General Public License (GPL) . . . . . . . . . . . . . . . . 469 General Public License, See GPL . . . . . . . . . . . . . . 10 generate time values . . . . . . . . . . . . . . . . . . . . . . . . . . 192 gensub() function (gawk) . . . . . . . . . . . . . . . . 109, 178 gensub() function (gawk), escape processing . . 186 getaddrinfo() function (C library) . . . . . . . . . . . 308 getgrent() function (C library) . . . . . . . . . 242, 246 getgrent() user-defined function . . . . . . . . 242, 246 getgrgid() function (C library) . . . . . . . . . . . . . . 245 getgrgid() user-defined function . . . . . . . . . . . . . 245 getgrnam() function (C library) . . . . . . . . . . . . . . 245 getgrnam() user-defined function . . . . . . . . . . . . . 245 getgruser() function (C library) . . . . . . . . . . . . . 246 getgruser() function, user-defined . . . . . . . . . . . 246 getline command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 getline command, _gr_init() user-defined

function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 getline command, _pw_init() function . . . . . . 241 getline command, coprocesses, using from …. 82,

102 getline command, deadlock and . . . . . . . . . . . . . 306 getline command, explicit input with . . . . . . . . . 78 getline command, FILENAME variable and . . . . . 83 getline command, return values. . . . . . . . . . . . . . . 78 getline command, variants. . . . . . . . . . . . . . . . . . . . 84 getline from a file . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 getline into a variable . . . . . . . . . . . . . . . . . . . . . . . . 79

                                                getline statement, BEGINFILE/ENDFILE patterns

and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 getlocaltime() user-defined function . . . . . . . . 225 getopt() function (C library) . . . . . . . . . . . . . . . . 233 getopt() user-defined function . . . . . . . . . . . . . . . 235 getpwent() function (C library) . . . . . . . . . 238, 241 getpwent() user-defined function . . . . . . . . 238, 241 getpwnam() function (C library) . . . . . . . . . . . . . . 241 getpwnam() user-defined function . . . . . . . . . . . . . 241 getpwuid() function (C library) . . . . . . . . . . . . . . 241 getpwuid() user-defined function . . . . . . . . . . . . . 241 gettext library. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 gettext library, locale categories . . . . . . . . . . . . . 316 gettext() function (C library) . . . . . . . . . . . . . . . 316 gettimeofday() extension function . . . . . . . . . . . 406 git utility . . . . . . . . . . . . . . . . . . . . . 407, 444, 447, 449 Git, use of for gawk source code . . . . . . . . . . . . . . . 451 GNITS mailing list . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 GNU awk, See gawk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 GNU Free Documentation License . . . . . . . . . . . . 487 GNU General Public License . . . . . . . . . . . . . . . . . 469 GNU Lesser General Public License. . . . . . . . . . . 470 GNU long options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 GNU long options, printing list of. . . . . . . . . . . . . . 34 GNU Project . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 469 GNU/Linux. . . . . . . . . . . . . . . . . . . . . . . . . . 10, 322, 474 Gordon, Assaf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 GPL (General Public License) . . . . . . . . . . . . 10, 469 GPL (General Public License), printing . . . . . . . . 32 grcat program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 Grigera, Juan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 group database, reading . . . . . . . . . . . . . . . . . . . . . . 242 group file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 group ID of gawk user . . . . . . . . . . . . . . . . . . . . . . . . 152 groups, information about . . . . . . . . . . . . . . . . . . . . 242 gsub() function . . . . . . . . . . . . . . . . . . . . . . . . . 109, 179 gsub() function, arguments of . . . . . . . . . . . . . . . . 185 gsub() function, escape processing. . . . . . . . . . . . 186

H

h debugger command (alias for help) . . . . . . . . . 338 Hankerson, Darrel . . . . . . . . . . . . . . . . . . . . . . . . 12, 426 Haque, John. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 Hartholz, Elaine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Hartholz, Marshall. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Hasegawa, Isamu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 help debugger command . . . . . . . . . . . . . . . . . . . . . 338 hexadecimal numbers . . . . . . . . . . . . . . . . . . . . . . . . . 107 hexadecimal values, enabling interpretation of . . 35 history expansion, in debugger . . . . . . . . . . . . . . . . 338 histsort.awk program . . . . . . . . . . . . . . . . . . . . . . . 280 Hughes, Phil . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 HUP signal, for dynamic profiling . . . . . . . . . . . . . . 312 hyphen (-), - operator. . . . . . . . . . . . . . . . . . . . . . . . 128 hyphen (-), -- operator. . . . . . . . . . . . . . . . . . 119, 128 hyphen (-), -= operator. . . . . . . . . . . . . . . . . . 118, 129 hyphen (-), filenames beginning with . . . . . . . . . . 32 hyphen (-), in bracket expressions . . . . . . . . . . . . . 51

I

i debugger command (alias for info) . . . . . . . . . 335 id utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 id.awk program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 if statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 if statement, actions, changing . . . . . . . . . . . . . . . 133 if statement, use of regexps in. . . . . . . . . . . . . . . . . 45 igawk.sh program . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 ignore breakpoint. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 332 ignore debugger command . . . . . . . . . . . . . . . . . . . 332 IGNORECASE variable . . . . . . . . . . . . . . . . . . . . . . . . . . 148 IGNORECASE variable, and array indices . . . . . . . . 161 IGNORECASE variable, and array sorting functions

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 IGNORECASE variable, in example programs …. 217 IGNORECASE variable, with ~ and !~ operators . . 55 Illumos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 445 Illumos, POSIX-compliant awk. . . . . . . . . . . . . . . . 445 implementation issues, gawk . . . . . . . . . . . . . . . . . . 447 implementation issues, gawk, limits . . . . . . . . 83, 100 implementation issues, gawk, debugging . . . . . . . 447 in operator . . . . . . . . . . . . . . . . . . . . . . . . . 122, 129, 141 in operator, index existence in multidimensional

arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 in operator, order of array access . . . . . . . . . . . . . 164 in operator, testing if array element exists …. 161 in operator, use in loops. . . . . . . . . . . . . . . . . . . . . . 163 including files, @include directive . . . . . . . . . . . . . . 42 increment operators . . . . . . . . . . . . . . . . . . . . . . . . . . 119 index() function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 indexing arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 indirect function calls. . . . . . . . . . . . . . . . . . . . . . . . . 208 indirect function calls, @-notation . . . . . . . . . . . . . 209 infinite precision. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 341 info debugger command . . . . . . . . . . . . . . . . . . . . . 335 initialization, automatic . . . . . . . . . . . . . . . . . . . . . . . 27 inplace extension . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 input files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 input files, closing . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 input files, counting elements in. . . . . . . . . . . . . . . 269 input files, examples . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 input files, reading. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 input files, running awk without . . . . . . . . . . . . . . . . 18 input files, variable assignments and . . . . . . . . . . . 38 input pipeline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 input record, length of . . . . . . . . . . . . . . . . . . . . . . . . 180 input redirection. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 input, data, nondecimal . . . . . . . . . . . . . . . . . . . . . . 299 input, explicit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 input, files, See input files. . . . . . . . . . . . . . . . . . . . . . 76 input, multiline records . . . . . . . . . . . . . . . . . . . . . . . . 76 input, splitting into records . . . . . . . . . . . . . . . . . . . . 59 input, standard . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 100 input/output functions . . . . . . . . . . . . . . . . . . . . . . . 189 input/output, binary . . . . . . . . . . . . . . . . . . . . . . . . . 147 input/output, from BEGIN and END . . . . . . . . . . . . 135 input/output, two-way . . . . . . . . . . . . . . . . . . . . . . . 306 insomnia, cure for . . . . . . . . . . . . . . . . . . . . . . . . . . . . 272 installation, VMS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 439 installing gawk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 instruction tracing, in debugger . . . . . . . . . . . . . . . 336 INT signal (MS-Windows). . . . . . . . . . . . . . . . . . . . . 312 int() function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 integer array indices . . . . . . . . . . . . . . . . . . . . . . . . . . 167 integers, arbitrary precision . . . . . . . . . . . . . . . . . . . 350 integers, unsigned . . . . . . . . . . . . . . . . . . . . . . . . . . . . 341 interacting with other programs . . . . . . . . . . . . . . 190 internationalization . . . . . . . . . . . . . . . . . . . . . . 198, 315 internationalization, localization . . . . . . . . . 149, 315 internationalization, localization, character classes

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 internationalization, localization, gawk and …. 315 internationalization, localization, locale categories

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316 internationalization, localization, marked strings

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 internationalization, localization, portability and

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320 internationalizing a program . . . . . . . . . . . . . . . . . . 315 interpreted programs . . . . . . . . . . . . . . . . . . . . 459, 469 interval expressions, regexp operator . . . . . . . . . . . 50 inventory-shipped file . . . . . . . . . . . . . . . . . . . . . . . . 23 invoke shell command . . . . . . . . . . . . . . . . . . . . . . . . 190 isarray() function (gawk). . . . . . . . . . . . . . . . . . . . 198 ISO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 470 ISO 8859-1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 465 ISO Latin-1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 465

J

Jacobs, Andrew . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 Jaegermann, Michal. . . . . . . . . . . . . . . . . . . . . . . 12, 426 Java implementation of awk. . . . . . . . . . . . . . . . . . . 445 Java programming language . . . . . . . . . . . . . . . . . . 470 jawk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 445 Jedi knights . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Johansen, Chris . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293 join() user-defined function. . . . . . . . . . . . . . . . . . 225

K

Kahrs, Jürgen . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 426 Kasal, Stepan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Kenobi, Obi-Wan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Kernighan, Brian . . . . . 6, 10, 12, 81, 115, 217, 415,

425, 444, 461, 466 kill command, dynamic profiling. . . . . . . . . . . . . 312 Knights, jedi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Kwok, Conrad. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 425

L

l debugger command (alias for list) . . . . . . . . . 338 labels.awk program . . . . . . . . . . . . . . . . . . . . . . . . . 277 Langston, Peter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 languages, data-driven . . . . . . . . . . . . . . . . . . . . . . . . 460 LC_ALL locale category. . . . . . . . . . . . . . . . . . . . . . . . 317 LC_COLLATE locale category . . . . . . . . . . . . . . . . . . . 316 LC_CTYPE locale category . . . . . . . . . . . . . . . . . . . . . 316 LC_MESSAGES locale category . . . . . . . . . . . . . . . . . . 316

LC_MESSAGES locale category, bindtextdomain()

function (gawk) . . . . . . . . . . . . . . . . . . . . . . . . . . 318 LC_MONETARY locale category . . . . . . . . . . . . . . . . . . 316 LC_NUMERIC locale category . . . . . . . . . . . . . . . . . . . 317 LC_TIME locale category . . . . . . . . . . . . . . . . . . . . . . 317 left angle bracket (<), < operator . . . . . . . . . 122, 128 left angle bracket (<), < operator (I/O). . . . . . . . . 80 left angle bracket (<), <= operator . . . . . . . . 122, 128 left shift. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 left shift, bitwise . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 leftmost longest match. . . . . . . . . . . . . . . . . . . . . . . . . 76 length of input record . . . . . . . . . . . . . . . . . . . . . . . . 180 length of string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 length() function . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 Lesser General Public License (LGPL) . . . . . . . . 470 LGPL (Lesser General Public License) . . . . . . . . 470 libmawk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 445 libraries of awk functions. . . . . . . . . . . . . . . . . . . . . . 217 libraries of awk functions, assertions. . . . . . . . . . . 220 libraries of awk functions, associative arrays and

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 libraries of awk functions, character values as

numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 libraries of awk functions, command-line options

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 libraries of awk functions, example program for

using . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 libraries of awk functions, group database, reading

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 libraries of awk functions, managing, data files

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 libraries of awk functions, managing, time . . . . . 225 libraries of awk functions, merging arrays into

strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 libraries of awk functions, rounding numbers . . 222 libraries of awk functions, user database, reading

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 line breaks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 line continuations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 line continuations, gawk. . . . . . . . . . . . . . . . . . . . . . . 126 line continuations, in print statement . . . . . . . . . 90 line continuations, with C shell . . . . . . . . . . . . . . . . 27 lines, blank, printing. . . . . . . . . . . . . . . . . . . . . . . . . . . 89 lines, counting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 lines, duplicate, removing . . . . . . . . . . . . . . . . . . . . . 280 lines, matching ranges of. . . . . . . . . . . . . . . . . . . . . . 133 lines, skipping between markers . . . . . . . . . . . . . . . 133 lint checking. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 lint checking, array elements . . . . . . . . . . . . . . . . . . 169 lint checking, array subscripts . . . . . . . . . . . . . . . . 168 lint checking, empty programs . . . . . . . . . . . . . . . . . 31 lint checking, issuing warnings . . . . . . . . . . . . . . . . . 34 lint checking, POSIXLY_CORRECT environment

variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 lint checking, undefined functions . . . . . . . . . . . . . 206 LINT variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 322, 474 list all global variables, in debugger . . . . . . . . . . . 336 list debugger command . . . . . . . . . . . . . . . . . . . . . 338 list function definitions, in debugger . . . . . . . . . . 335 loading extensions, @load directive . . . . . . . . . . . . . 43 loading, extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 local variables, in a function . . . . . . . . . . . . . . . . . . 203 locale categories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316 locale decimal point character. . . . . . . . . . . . . . . . . . 36 locale, definition of . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 localization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 localization, See internationalization, localization

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 log files, timestamps in . . . . . . . . . . . . . . . . . . . . . . . 192 log() function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 logarithm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 logical false/true . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 logical operators, See Boolean expressions . . . . . 124 login information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 long options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 loops . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 loops, break statement and . . . . . . . . . . . . . . . . . . . 142 loops, continue statements and. . . . . . . . . . . . . . . 141 loops, count for header, in a profile . . . . . . . . . . . 311 loops, do-while. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 loops, exiting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 loops, for, array scanning . . . . . . . . . . . . . . . . . . . . 163 loops, for, iterative. . . . . . . . . . . . . . . . . . . . . . . . . . . 140 loops, See Also while statement . . . . . . . . . . . . . . 139 loops, while . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 ls utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 lshift() function (gawk) . . . . . . . . . . . . . . . . . . . . . 197 lvalues/rvalues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116

M

mail-list file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 mailing labels, printing . . . . . . . . . . . . . . . . . . . . . . . 276 mailing list, GNITS. . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Malmberg, John . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 443 Malmberg, John E. . . . . . . . . . . . . . . . . . . . . . . . . . . . 427 mark parity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 marked string extraction (internationalization)

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 marked strings, extracting . . . . . . . . . . . . . . . . . . . . 319 Marx, Groucho . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 match regexp in string. . . . . . . . . . . . . . . . . . . . . . . . 180 match() function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 match() function, RSTART/RLENGTH variables . . 181 matching, expressions, See comparison expressions

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 matching, leftmost longest . . . . . . . . . . . . . . . . . . . . . 76 matching, null strings. . . . . . . . . . . . . . . . . . . . . . . . . 186 mawk utility . . . . . . . . . . . . . . . . . . 47, 82, 115, 146, 444 maximum precision supported by MPFR library

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 McIlroy, Doug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 467 McPhee, Patrick. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 message object files. . . . . . . . . . . . . . . . . . . . . . . . . . . 316 message object files, converting from portable

object files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 message object files, specifying directory of … 316,

317 messages from extensions . . . . . . . . . . . . . . . . . . . . . 371 metacharacters in regular expressions . . . . . . . . . . 48 metacharacters, escape sequences for . . . . . . . . . . . 48 minimum precision required by MPFR library

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 mktime() function (gawk) . . . . . . . . . . . . . . . . . . . . . 192 modifiers, in format specifiers . . . . . . . . . . . . . . . . . . 94 monetary information, localization . . . . . . . . . . . . 316 Moore, Duncan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 msgfmt utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 multiple precision . . . . . . . . . . . . . . . . . . . . . . . . . . . . 341 multiple-line records . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

N

n debugger command (alias for next) . . . . . . . . . 332 names, arrays/variables. . . . . . . . . . . . . . . . . . . . . . . 218 names, functions . . . . . . . . . . . . . . . . . . . . . . . . . 199, 218 namespace issues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 namespace issues, functions . . . . . . . . . . . . . . . . . . . 199 NetBSD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 networks, programming . . . . . . . . . . . . . . . . . . . . . . . 307 networks, support for . . . . . . . . . . . . . . . . . . . . . . . . . 102 newlines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27, 36, 125 newlines, as field separators . . . . . . . . . . . . . . . . . . . . 68 newlines, as record separators . . . . . . . . . . . . . . . . . . 59 newlines, in dynamic regexps . . . . . . . . . . . . . . . . . . 54 newlines, in regexp constants . . . . . . . . . . . . . . . . . . 54 newlines, printing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 newlines, separating statements in actions …. 137,

138 next debugger command . . . . . . . . . . . . . . . . . . . . . 332 next file statement. . . . . . . . . . . . . . . . . . . . . . . . . . 420 next statement . . . . . . . . . . . . . . . . . . . . . . . . . . 125, 144 next statement, BEGIN/END patterns and . . . . . . 135 next statement, BEGINFILE/ENDFILE patterns and

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 next statement, user-defined functions and. … 145 nextfile statement . . . . . . . . . . . . . . . . . . . . . . . . . . 145 nextfile statement, BEGIN/END patterns and . . 135

nextfile statement, BEGINFILE/ENDFILE patterns

and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 nextfile statement, user-defined functions and

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 nexti debugger command . . . . . . . . . . . . . . . . . . . . 332 NF variable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63, 151 NF variable, decrementing . . . . . . . . . . . . . . . . . . . . . . 66 ni debugger command (alias for nexti) . . . . . . . 332 noassign.awk program . . . . . . . . . . . . . . . . . . . . . . . 233 non-existent array elements . . . . . . . . . . . . . . . . . . . 161 not Boolean-logic operator. . . . . . . . . . . . . . . . . . . . 124 NR variable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59, 151 NR variable, changing . . . . . . . . . . . . . . . . . . . . . . . . . 155 null strings . . . . . . . . . . . . . . . . . . . . . . . 61, 68, 120, 461 null strings in gawk arguments, quoting and …. 22 null strings, and deleting array elements . . . . . . 168 null strings, as array subscripts . . . . . . . . . . . . . . . 168 null strings, converting numbers to strings …. 112 null strings, matching. . . . . . . . . . . . . . . . . . . . . . . . . 186 number as string of bits . . . . . . . . . . . . . . . . . . . . . . 198 number of array elements . . . . . . . . . . . . . . . . . . . . . 180 number sign (#), #! (executable scripts). . . . . . . . 19 number sign (#), commenting . . . . . . . . . . . . . . . . . . 20 numbers, as array subscripts . . . . . . . . . . . . . . . . . . 167 numbers, as values of characters . . . . . . . . . . . . . . 223 numbers, Cliff random . . . . . . . . . . . . . . . . . . . . . . . . 223 numbers, converting . . . . . . . . . . . . . . . . . . . . . 111, 198 numbers, converting, to strings. . . . . . . . . . . 147, 148 numbers, hexadecimal . . . . . . . . . . . . . . . . . . . . . . . . 107 numbers, octal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 numbers, rounding . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 numeric constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 numeric functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 numeric, output format . . . . . . . . . . . . . . . . . . . . . . . . 92 numeric, strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121

O

o debugger command (alias for option) . . . . . . . 336 obsolete features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 octal numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 octal values, enabling interpretation of . . . . . . . . . 35 OFMT variable. . . . . . . . . . . . . . . . . . . . . . . . . 92, 112, 148 OFMT variable, POSIX awk and . . . . . . . . . . . . . . . . . 92 OFS variable . . . . . . . . . . . . . . . . . . . . . . . . . . . 66, 91, 148 OpenBSD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 OpenSolaris . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 445 operating systems, BSD-based . . . . . . . . . . . . . . . . . 10 operating systems, PC, gawk on. . . . . . . . . . . . . . . 437 operating systems, PC, gawk on, installing …. 435 operating systems, porting gawk to. . . . . . . . . . . . 449 operating systems, See Also GNU/Linux, PC

operating systems, Unix. . . . . . . . . . . . . . . . . . 429 operations, bitwise. . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 operators, arithmetic . . . . . . . . . . . . . . . . . . . . . . . . . 114 operators, assignment . . . . . . . . . . . . . . . . . . . . . . . . 116 operators, assignment, evaluation order . . . . . . . 118 operators, Boolean, See Boolean expressions . . 124 operators, decrement/increment . . . . . . . . . . . . . . 119 operators, GNU-specific . . . . . . . . . . . . . . . . . . . . . . . 54 operators, input/output …. 80, 81, 82, 98, 99, 128 operators, logical, See Boolean expressions …. 124 operators, precedence . . . . . . . . . . . . . . . . . . . . 120, 128 operators, relational, See operators, comparison

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 operators, short-circuit . . . . . . . . . . . . . . . . . . . . . . . 125 operators, string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 operators, string-matching . . . . . . . . . . . . . . . . . . . . . 45 operators, string-matching, for buffers. . . . . . . . . . 54 operators, word-boundary (gawk) . . . . . . . . . . . . . . 55 option debugger command . . . . . . . . . . . . . . . . . . . 336 options, command-line . . . . . . . . . . . . . . . . . . . . . . . . . 31 options, command-line, end of . . . . . . . . . . . . . . . . . 32 options, command-line, invoking awk . . . . . . . . . . . 31 options, command-line, processing . . . . . . . . . . . . 233 options, deprecated . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 options, long . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 options, printing list of . . . . . . . . . . . . . . . . . . . . . . . . 34 OR bitwise operation . . . . . . . . . . . . . . . . . . . . . . . . . 196 or Boolean-logic operator . . . . . . . . . . . . . . . . . . . . . 124 or() function (gawk) . . . . . . . . . . . . . . . . . . . . . . . . . 197 ord() extension function . . . . . . . . . . . . . . . . . . . . . 403 ord() user-defined function . . . . . . . . . . . . . . . . . . . 223 order of evaluation, concatenation . . . . . . . . . . . . 115 ORS variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91, 148 output field separator, See OFS variable . . . . . . . . 66 output record separator, See ORS variable . . . . . . 91 output redirection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 output wrapper . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 368 output, buffering . . . . . . . . . . . . . . . . . . . . . . . . 189, 191 output, duplicating into files . . . . . . . . . . . . . . . . . . 263 output, files, closing . . . . . . . . . . . . . . . . . . . . . . . . . . 102 output, format specifier, OFMT . . . . . . . . . . . . . . . . . . 92 output, formatted . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 output, pipes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 output, printing, See printing . . . . . . . . . . . . . . . . . . 89 output, records . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 output, standard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

P

p debugger command (alias for print) . . . . . . . . 333 Papadopoulos, Panos . . . . . . . . . . . . . . . . . . . . . . . . . 427 parent process ID of gawk process . . . . . . . . . . . . . 152 parentheses (), in a profile. . . . . . . . . . . . . . . . . . . . 311 parentheses (), regexp operator . . . . . . . . . . . . . . . . 49 password file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 patsplit() function (gawk). . . . . . . . . . . . . . . . . . . 182 patterns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 patterns, comparison expressions as . . . . . . . . . . . 132 patterns, counts, in a profile . . . . . . . . . . . . . . . . . . 310 patterns, default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 patterns, empty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 patterns, expressions as. . . . . . . . . . . . . . . . . . . . . . . 131 patterns, ranges in. . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 patterns, regexp constants as . . . . . . . . . . . . . . . . . 132 patterns, types of. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 pawk (profiling version of Brian Kernighan’s awk)

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 445 pawk, awk-like facilities for Python . . . . . . . . . . . . 445 PC operating systems, gawk on . . . . . . . . . . . . . . . 437 PC operating systems, gawk on, installing . . . . . 435 percent sign (%), % operator . . . . . . . . . . . . . . . . . . . 128 percent sign (%), %= operator . . . . . . . . . . . . . 118, 129 period (.), regexp operator . . . . . . . . . . . . . . . . . . . . 48 Perl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 453 Peters, Arno . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 Peterson, Hal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 pipe, closing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 pipe, input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 pipe, output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 Pitts, Dave . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 443 Plauger, P.J. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 plug-in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 355 plus sign (+), + operator . . . . . . . . . . . . . . . . . . . . . . 128 plus sign (+), ++ operator . . . . . . . . . . . . . . . . 119, 128 plus sign (+), += operator . . . . . . . . . . . . . . . . 117, 129 plus sign (+), regexp operator . . . . . . . . . . . . . . . . . . 49 pointers to functions. . . . . . . . . . . . . . . . . . . . . . . . . . 208 portability. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 portability, #! (executable scripts) . . . . . . . . . . . . . 20 portability, ** operator and. . . . . . . . . . . . . . . . . . . 115 portability, **= operator and . . . . . . . . . . . . . . . . . 118 portability, ARGV variable . . . . . . . . . . . . . . . . . . . . . . 20 portability, backslash continuation and . . . . . . . . . 28 portability, backslash in escape sequences . . . . . . 47 portability, close() function and . . . . . . . . . . . . . 103 portability, data files as single record. . . . . . . . . . . 63 portability, deleting array elements . . . . . . . . . . . 169 portability, example programs . . . . . . . . . . . . . . . . 217 portability, functions, defining . . . . . . . . . . . . . . . . 201 portability, gawk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 449 portability, gettext library and. . . . . . . . . . . . . . . 315 portability, internationalization and . . . . . . . . . . . 320 portability, length() function . . . . . . . . . . . . . . . . 180 portability, new awk vs. old awk . . . . . . . . . . . . . . . 112 portability, next statement in user-defined

functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 portability, NF variable, decrementing . . . . . . . . . . 66 portability, operators . . . . . . . . . . . . . . . . . . . . . . . . . 120 portability, operators, not in POSIX awk . . . . . . 129 portability, POSIXLY_CORRECT environment variable

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 portability, substr() function . . . . . . . . . . . . . . . . 185 portable object files. . . . . . . . . . . . . . . . . . . . . . 315, 319 portable object files, converting to message object

files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 portable object files, generating . . . . . . . . . . . . . . . . 34 portable object template files . . . . . . . . . . . . . . . . . 315 porting gawk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 449 positional specifiers, printf statement …. 94, 319 positional specifiers, printf statement, mixing

with regular formats . . . . . . . . . . . . . . . . . . . . . 320 POSIX awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 118 POSIX awk, ** operator and. . . . . . . . . . . . . . . . . . 129 POSIX awk, **= operator and . . . . . . . . . . . . . . . . 118 POSIX awk, < operator and . . . . . . . . . . . . . . . . . . . . 80 POSIX awk, | I/O operator and . . . . . . . . . . . . . . . 82 POSIX awk, arithmetic operators and . . . . . . . . . 114 POSIX awk, backslashes in string constants …. 47 POSIX awk, BEGIN/END patterns . . . . . . . . . . . . . . 135 POSIX awk, bracket expressions and . . . . . . . . . . . 51 POSIX awk, bracket expressions and, character

classes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51, 52 POSIX awk, break statement and . . . . . . . . . . . . . 143 POSIX awk, changes in awk versions . . . . . . . . . . 414 POSIX awk, continue statement and . . . . . . . . . 144 POSIX awk, CONVFMT variable and . . . . . . . . . . . . 147 POSIX awk, date utility and. . . . . . . . . . . . . . . . . . 195 POSIX awk, field separators and. . . . . . . . . . . . 63, 71 POSIX awk, FS variable and . . . . . . . . . . . . . . . . . . 148 POSIX awk, function keyword in. . . . . . . . . . . . . 200 POSIX awk, functions and, gsub()/sub() . . . . . 188 POSIX awk, functions and, length() . . . . . . . . . 180 POSIX awk, GNU long options and . . . . . . . . . . . . 31 POSIX awk, interval expressions in . . . . . . . . . . . . . 50 POSIX awk, next/nextfile statements and… 145 POSIX awk, numeric strings and . . . . . . . . . . . . . . 121 POSIX awk, OFMT variable and . . . . . . . . . . . . 92, 112 POSIX awk, period (.), using . . . . . . . . . . . . . . . . . . 49 POSIX awk, printf format strings and. . . . . . . . . 96 POSIX awk, regular expressions and . . . . . . . . . . . 50 POSIX awk, timestamps and. . . . . . . . . . . . . . . . . . 192 POSIX mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35, 37 POSIX, awk and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 POSIX, gawk extensions not included in . . . . . . 415 POSIX, programs, implementing in awk . . . . . . . 251 POSIXLY_CORRECT environment variable . . . . . . . . 37 PREC variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 precedence. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120, 128 precedence, regexp operators. . . . . . . . . . . . . . . . . . . 50 predefined variables . . . . . . . . . . . . . . . . . . . . . . . . . . 147 predefined variables, -v option, setting with …. 32 predefined variables, conveying information … 149 predefined variables, user-modifiable . . . . . . . . . . 147 print debugger command . . . . . . . . . . . . . . . . . . . . 333 print statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 print statement, BEGIN/END patterns and . . . . . 135 print statement, commas, omitting . . . . . . . . . . . . 90 print statement, I/O operators in . . . . . . . . . . . . 128 print statement, line continuations and . . . . . . . . 90 print statement, OFMT variable and . . . . . . . . . . . 148 print statement, See Also redirection, of output

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 print statement, sprintf() function and . . . . . 222 print variables, in debugger . . . . . . . . . . . . . . . . . . . 333 printf debugger command . . . . . . . . . . . . . . . . . . . 334 printf statement . . . . . . . . . . . . . . . . . . . . . . . . . . 89, 92 printf statement, columns, aligning . . . . . . . . . . . 90 printf statement, format-control characters … 93 printf statement, I/O operators in . . . . . . . . . . . 128 printf statement, modifiers . . . . . . . . . . . . . . . . . . . 94 printf statement, positional specifiers …. 94, 319 printf statement, positional specifiers, mixing

with regular formats . . . . . . . . . . . . . . . . . . . . . 320 printf statement, See Also redirection, of output

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 printf statement, sprintf() function and. … 222 printf statement, syntax of . . . . . . . . . . . . . . . . . . . 92 printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 printing messages from extensions . . . . . . . . . . . . 371 printing, list of options . . . . . . . . . . . . . . . . . . . . . . . . 34 printing, mailing labels . . . . . . . . . . . . . . . . . . . . . . . 276 printing, unduplicated lines of text. . . . . . . . . . . . 265 printing, user information. . . . . . . . . . . . . . . . . . . . . 260 private variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 process group ID of gawk process . . . . . . . . . . . . . 152 process ID of gawk process . . . . . . . . . . . . . . . . . . . . 152 processes, two-way communications with . . . . . . 305 processing data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 459 PROCINFO array . . . . . . . . . . . . . . . . . . . . . 151, 192, 238 PROCINFO array, and communications via ptys

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 PROCINFO array, and group membership . . . . . . . 242 PROCINFO array, and user and group ID numbers

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 PROCINFO array, testing the field splitting . . . . . 241 PROCINFO array, uses. . . . . . . . . . . . . . . . . . . . . . . . . . 153 PROCINFO, values of sorted_in . . . . . . . . . . . . . . . . 165 profiling awk programs . . . . . . . . . . . . . . . . . . . . . . . . 308 profiling awk programs, dynamically . . . . . . . . . . 311 program identifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 program, definition of. . . . . . . . . . . . . . . . . . . . . . . . . . 17 programming conventions, --non-decimal-data

option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 programming conventions, ARGC/ARGV variables

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 programming conventions, exit statement …. 146 programming conventions, function parameters

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 programming conventions, functions, calling… 175 programming conventions, functions, writing . . 200 programming conventions, gawk extensions …. 391 programming conventions, private variable names

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 programming language, recipe for . . . . . . . . . . . . . . . 6 programming languages, Ada . . . . . . . . . . . . . . . . . 463 programming languages, Java . . . . . . . . . . . . . . . . . 470 programming languages, data-driven vs. procedural

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 programming, basic steps . . . . . . . . . . . . . . . . . . . . . 459 programming, concepts . . . . . . . . . . . . . . . . . . . . . . . 459 pwcat program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238

Q

q debugger command (alias for quit) . . . . . . . . . 338 QSE awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 445 Quanstrom, Erik . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 272 question mark (?), ?: operator. . . . . . . . . . . . . . . . 129 question mark (?), regexp operator . . . . . . . . . 49, 55 QuikTrim Awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 446 quit debugger command . . . . . . . . . . . . . . . . . . . . . 338 QUIT signal (MS-Windows) . . . . . . . . . . . . . . . . . . . 312 quoting in gawk command lines . . . . . . . . . . . . . . . . 19 quoting in gawk command lines, tricks for . . . . . . 22 quoting, for small awk programs . . . . . . . . . . . . . . . 20

R

r debugger command (alias for run) . . . . . . . . . . 332 Rakitzis, Byron. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 Ramey, Chet. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 359 rand() function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 random numbers, Cliff . . . . . . . . . . . . . . . . . . . . . . . . 223 random numbers, rand()/srand() functions . . 176 random numbers, seed of . . . . . . . . . . . . . . . . . . . . . 176 range expressions (regexps) . . . . . . . . . . . . . . . . . . . . 50 range patterns. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 range patterns, line continuation and . . . . . . . . . 134 Rankin, Pat. . . . . . . . . . . . . . . . . . . . . . . . . . 12, 117, 425 reada() extension function . . . . . . . . . . . . . . . . . . . 405 readable data files, checking . . . . . . . . . . . . . . . . . . 231 readable.awk program . . . . . . . . . . . . . . . . . . . . . . . 231 readdir extension . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 readfile() extension function. . . . . . . . . . . . . . . . 406 readfile() user-defined function . . . . . . . . . . . . . 227 reading input files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 recipe for a programming language . . . . . . . . . . . . . . 6 record separators . . . . . . . . . . . . . . . . . . . . . . . . . 59, 149 record separators, changing . . . . . . . . . . . . . . . . . . . . 61 record separators, regular expressions as . . . . . . . 61 record separators, with multiline records . . . . . . . 76 records . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59, 460 records, multiline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 records, printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 records, splitting input into . . . . . . . . . . . . . . . . . . . . 59 records, terminating . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 records, treating files as. . . . . . . . . . . . . . . . . . . . . . . . 63 recursive functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 redirect gawk output, in debugger . . . . . . . . . . . . . 336 redirection of input . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 redirection of output. . . . . . . . . . . . . . . . . . . . . . . . . . . 98 reference counting, sorting arrays . . . . . . . . . . . . . 305 regexp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 regexp constants . . . . . . . . . . . . . . . . . . . . . 46, 108, 123 regexp constants, /=.../, /= operator and …. 118 regexp constants, as patterns . . . . . . . . . . . . . . . . . 132 regexp constants, in gawk . . . . . . . . . . . . . . . . . . . . . 109 regexp constants, slashes vs. quotes . . . . . . . . . . . . 53 regexp constants, vs. string constants . . . . . . . . . . 53 register extension . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 regular expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 regular expressions as field separators . . . . . . . . . . 67 regular expressions, anchors in . . . . . . . . . . . . . . . . . 48 regular expressions, as field separators . . . . . . . . . 68 regular expressions, as patterns . . . . . . . . . . . 45, 131 regular expressions, as record separators . . . . . . . 61 regular expressions, case sensitivity. . . . . . . . 55, 148 regular expressions, computed . . . . . . . . . . . . . . . . . 52 regular expressions, constants, See regexp

constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 regular expressions, dynamic . . . . . . . . . . . . . . . . . . . 52 regular expressions, dynamic, with embedded

newlines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 regular expressions, gawk, command-line options

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 regular expressions, interval expressions and. … 36 regular expressions, leftmost longest match. . . . . 52 regular expressions, operators . . . . . . . . . . . . . . 45, 48 regular expressions, operators, for buffers . . . . . . 54 regular expressions, operators, for words . . . . . . . 54 regular expressions, operators, gawk . . . . . . . . . . . . 54 regular expressions, operators, precedence of … 50 regular expressions, searching for . . . . . . . . . . . . . 256 relational operators, See comparison operators

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 replace in string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 return debugger command . . . . . . . . . . . . . . . . . . . 332 return statement, user-defined functions . . . . . . 206 return value, close() function . . . . . . . . . . . . . . . . 104 rev() user-defined function . . . . . . . . . . . . . . . . . . . 202 revoutput extension. . . . . . . . . . . . . . . . . . . . . . . . . . 404 revtwoway extension. . . . . . . . . . . . . . . . . . . . . . . . . . 405 rewind() user-defined function . . . . . . . . . . . . . . . 231 right angle bracket (>), > operator . . . . . . . 122, 128 right angle bracket (>), > operator (I/O) . . . . . . . 98 right angle bracket (>), >= operator . . . . . . 122, 128 right angle bracket (>), >> operator (I/O) . . 98, 128 right shift . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 right shift, bitwise . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 Ritchie, Dennis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 461 RLENGTH variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 RLENGTH variable, match() function and . . . . . . . 181 Robbins, Arnold … 70, 81, 240, 272, 359, 427, 443,

453 Robbins, Bill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 Robbins, Harry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Robbins, Jean . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Robbins, Miriam . . . . . . . . . . . . . . . . . . . . . . 12, 81, 240 Rommel, Kai Uwe . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 round to nearest integer . . . . . . . . . . . . . . . . . . . . . . 176 round() user-defined function. . . . . . . . . . . . . . . . . 222 rounding numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 ROUNDMODE variable . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 RS variable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59, 149 RS variable, multiline records and . . . . . . . . . . . . . . 76 rshift() function (gawk) . . . . . . . . . . . . . . . . . . . . . 197 RSTART variable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 RSTART variable, match() function and . . . . . . . . 181 RT variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61, 78, 153 Rubin, Paul. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 425 rule, definition of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 run debugger command. . . . . . . . . . . . . . . . . . . . . . . 332 rvalues/lvalues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116

S

s debugger command (alias for step) . . . . . . . . . 333 sample debugging session . . . . . . . . . . . . . . . . . . . . . 326 sandbox mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 save debugger options . . . . . . . . . . . . . . . . . . . . . . . . 336 scalar or array. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 scalar values. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 scanning arrays. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 scanning multidimensional arrays . . . . . . . . . . . . . 171 Schorr, Andrew . . . . . . . . . . . . . . . . . . . . . . 12, 154, 427 Schreiber, Bert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Schreiber, Rita . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 search and replace in strings . . . . . . . . . . . . . . . . . . 178 search in string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 search paths . . . . . . . . . . . . . . . . . . . . . . . . 295, 437, 442 search paths, for loadable extensions . . . . . . . . . . . 40 search paths, for source files . . . . . 39, 295, 437, 442 searching, files for regular expressions . . . . . . . . . 256 searching, for words . . . . . . . . . . . . . . . . . . . . . . . . . . 271 sed utility . . . . . . . . . . . . . . . . . . . . . . . . . . . 71, 284, 463 seeding random number generator . . . . . . . . . . . . 176 semicolon (;), AWKPATH variable and. . . . . . . . . . . 437 semicolon (;), separating statements in actions

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29, 137, 138 separators, field. . . . . . . . . . . . . . . . . . . . . . . . . . 147, 148 separators, field, FIELDWIDTHS variable and. … 147 separators, field, FPAT variable and. . . . . . . . . . . . 147 separators, field, POSIX and . . . . . . . . . . . . . . . . . . . 63 separators, for records . . . . . . . . . . . . . . . . . 59, 61, 149 separators, for records, regular expressions as . . 61 separators, for statements in actions . . . . . . . . . . 137 separators, subscript. . . . . . . . . . . . . . . . . . . . . . . . . . 149 set breakpoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 330 set debugger command. . . . . . . . . . . . . . . . . . . . . . . 334 set directory of message catalogs . . . . . . . . . . . . . . 199 set watchpoint. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 334 shadowing of variable values . . . . . . . . . . . . . . . . . . 200 shell quoting, rules for . . . . . . . . . . . . . . . . . . . . . . . . . 21 shells, piping commands into . . . . . . . . . . . . . . . . . 100 shells, quoting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 shells, quoting, rules for . . . . . . . . . . . . . . . . . . . . . . . 21 shells, scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 shells, sea . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 shells, variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 shift, bitwise . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 short-circuit operators . . . . . . . . . . . . . . . . . . . . . . . . 125 show all source files, in debugger. . . . . . . . . . . . . . 336 show breakpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 335 show function arguments, in debugger . . . . . . . . 335 show local variables, in debugger. . . . . . . . . . . . . . 335 show name of current source file, in debugger . . 335 show watchpoints. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 336 si debugger command (alias for stepi) . . . . . . . 333 side effects . . . . . . . . . . . . . . . . . . . . . . . . . . 115, 119, 120 side effects, array indexing . . . . . . . . . . . . . . . . . . . . 161 side effects, asort() function . . . . . . . . . . . . . . . . . 304 side effects, assignment expressions . . . . . . . . . . . 116 side effects, Boolean operators . . . . . . . . . . . . . . . . 124 side effects, conditional expressions . . . . . . . . . . . 126 side effects, decrement/increment operators … 119 side effects, FILENAME variable . . . . . . . . . . . . . . . . . 83 side effects, function calls . . . . . . . . . . . . . . . . . . . . . 127 side effects, statements . . . . . . . . . . . . . . . . . . . . . . . 138 sidebar, A Constant’s Base Does Not Affect Its

Value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 sidebar, Backslash Before Regular Characters . . 47 sidebar, Changing FS Does Not Affect the Fields

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 sidebar, Changing NR and FNR. . . . . . . . . . . . . . . . . 155 sidebar, Controlling Output Buffering with

system() . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 sidebar, Escape Sequences for Metacharacters . . 48 sidebar, FS and IGNORECASE . . . . . . . . . . . . . . . . . . . . 72 sidebar, Interactive Versus Noninteractive

Buffering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 sidebar, Matching the Null String . . . . . . . . . . . . . 186 sidebar, Operator Evaluation Order . . . . . . . . . . . 120 sidebar, Piping into sh. . . . . . . . . . . . . . . . . . . . . . . . 100 sidebar, Pre-POSIX awk Used OFMT for String

Conversion. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 sidebar, Recipe for a Programming Language. … 6 sidebar, RS = "\0" Is Not Portable . . . . . . . . . . . . . 62 sidebar, So Why Does gawk Have BEGINFILE and

ENDFILE? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 sidebar, Syntactic Ambiguities Between ‘/=’ and

Regular Expressions . . . . . . . . . . . . . . . . . . . . . 118 sidebar, Understanding ‘#!’ . . . . . . . . . . . . . . . . . . . . 19 sidebar, Understanding $0 . . . . . . . . . . . . . . . . . . . . . 67 sidebar, Using \n in Bracket Expressions of

Dynamic Regexps . . . . . . . . . . . . . . . . . . . . . . . . . 54 sidebar, Using close()’s Return Value . . . . . . . . 104 SIGHUP signal, for dynamic profiling. . . . . . . . . . . 312 SIGINT signal (MS-Windows) . . . . . . . . . . . . . . . . . 312 signals, HUP/SIGHUP, for profiling. . . . . . . . . . . . . . 312 signals, INT/SIGINT (MS-Windows) . . . . . . . . . . . 312 signals, QUIT/SIGQUIT (MS-Windows). . . . . . . . . 312 signals, USR1/SIGUSR1, for profiling . . . . . . . . . . . 312 signature program . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293 SIGQUIT signal (MS-Windows) . . . . . . . . . . . . . . . . 312 SIGUSR1 signal, for dynamic profiling . . . . . . . . . 312 silent debugger command . . . . . . . . . . . . . . . . . . . 332 sin() function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 sine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 single quote (') . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 single quote (') in gawk command lines . . . . . . . . 19 single quote ('), in shell commands . . . . . . . . . . . . 21 single quote ('), vs. apostrophe . . . . . . . . . . . . . . . . 20 single quote ('), with double quotes. . . . . . . . . . . . 22 single-character fields . . . . . . . . . . . . . . . . . . . . . . . . . . 69 single-step execution, in the debugger . . . . . . . . . 332 Skywalker, Luke . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 sleep utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 sleep() extension function . . . . . . . . . . . . . . . . . . . 406 Solaris, POSIX-compliant awk . . . . . . . . . . . . . . . . 445 sort array . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 sort array indices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 sort function, arrays, sorting . . . . . . . . . . . . . . . . . . 304 sort utility. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 sort utility, coprocesses and . . . . . . . . . . . . . . . . . . 306 sorting characters in different languages. . . . . . . 316 source code, awka. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 444 source code, Brian Kernighan’s awk . . . . . . . . . . . 444 source code, BusyBox Awk . . . . . . . . . . . . . . . . . . . 445 source code, gawk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 source code, Illumos awk . . . . . . . . . . . . . . . . . . . . . . 445 source code, jawk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 445 source code, libmawk . . . . . . . . . . . . . . . . . . . . . . . . . 445 source code, mawk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 444 source code, mixing. . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 source code, pawk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 445 source code, pawk (Python version) . . . . . . . . . . . 445 source code, QSE awk . . . . . . . . . . . . . . . . . . . . . . . . 445 source code, QuikTrim Awk . . . . . . . . . . . . . . . . . . 446 source code, Solaris awk. . . . . . . . . . . . . . . . . . . . . . . 445 source files, search path for . . . . . . . . . . . . . . . . . . . 295 sparse arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 Spencer, Henry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 463 split string into array . . . . . . . . . . . . . . . . . . . . . . . . . 182 split utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 split() function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 split() function, array elements, deleting …. 169 split.awk program. . . . . . . . . . . . . . . . . . . . . . . . . . . 262 sprintf() function . . . . . . . . . . . . . . . . . . . . . . . 92, 183 sprintf() function, OFMT variable and . . . . . . . . 148 sprintf() function, print/printf statements and

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 sqrt() function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 square brackets ([]), regexp operator . . . . . . . . . . 49 square root . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 srand() function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 stack frame. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 Stallman, Richard . . . . . . . . . . . . . . . . 10, 11, 425, 468 standard error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 standard input. . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 100 standard output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 starting the debugger . . . . . . . . . . . . . . . . . . . . . . . . . 327 stat() extension function . . . . . . . . . . . . . . . . . . . . 398 statements, compound, control statements and

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 statements, control, in actions . . . . . . . . . . . . . . . . 138 statements, multiple . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 step debugger command . . . . . . . . . . . . . . . . . . . . . 333 stepi debugger command . . . . . . . . . . . . . . . . . . . . 333 stop automatic display, in debugger . . . . . . . . . . . 334 stream editors . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71, 284 strftime() function (gawk). . . . . . . . . . . . . . . . . . . 192 string constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 string constants, vs. regexp constants . . . . . . . . . . 53 string extraction (internationalization) . . . . . . . . 319 string length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 string operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 string, regular expression match . . . . . . . . . . . . . . 180 string-manipulation functions . . . . . . . . . . . . . . . . . 177 string-matching operators. . . . . . . . . . . . . . . . . . . . . . 45 string-translation functions . . . . . . . . . . . . . . . . . . . 198 strings splitting, example . . . . . . . . . . . . . . . . . . . . . 183 strings, converting . . . . . . . . . . . . . . . . . . . . . . . 111, 198 strings, converting letter case . . . . . . . . . . . . . . . . . 186 strings, converting, numbers to. . . . . . . . . . . 147, 148 strings, empty, See null strings . . . . . . . . . . . . . . . . . 61 strings, extracting . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 strings, for localization . . . . . . . . . . . . . . . . . . . . . . . 317 strings, length limitations. . . . . . . . . . . . . . . . . . . . . 107 strings, merging arrays into . . . . . . . . . . . . . . . . . . . 225 strings, null . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 strings, numeric . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 strtonum() function (gawk). . . . . . . . . . . . . . . . . . . 183

                                                strtonum() function (gawk), --non-decimal-data

option and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 sub() function. . . . . . . . . . . . . . . . . . . . . . . . . . . 109, 184 sub() function, arguments of . . . . . . . . . . . . . . . . . 185 sub() function, escape processing . . . . . . . . . . . . . 186 subscript separators . . . . . . . . . . . . . . . . . . . . . . . . . . 149 subscripts in arrays, multidimensional. . . . . . . . . 169 subscripts in arrays, multidimensional, scanning

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 subscripts in arrays, numbers as . . . . . . . . . . . . . . 167 subscripts in arrays, uninitialized variables as

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 SUBSEP variable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 SUBSEP variable, and multidimensional arrays

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 substitute in string . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 substr() function . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 substring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 Sumner, Andrew . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 444 supplementary groups of gawk process. . . . . . . . . 153 switch statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 SYMTAB array . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 syntactic ambiguity: /= operator vs. /=.../ regexp

constant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 system() function . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 systime() function (gawk). . . . . . . . . . . . . . . . . . . . 193

T

t debugger command (alias for tbreak) . . . . . . . 332 tbreak debugger command . . . . . . . . . . . . . . . . . . . 332 Tcl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 TCP/IP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 TCP/IP, support for . . . . . . . . . . . . . . . . . . . . . . . . . 102 tee utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 tee.awk program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 temporary breakpoint . . . . . . . . . . . . . . . . . . . . . . . . 332 terminating records . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 testbits.awk program . . . . . . . . . . . . . . . . . . . . . . . 197 testext extension . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 Texinfo . . . . . . . . . . . . . . . . . 9, 217, 271, 281, 431, 449 Texinfo, chapter beginnings in files . . . . . . . . . . . . . 48 Texinfo, extracting programs from source files

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 text, printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 text, printing, unduplicated lines of . . . . . . . . . . . 265 TEXTDOMAIN variable . . . . . . . . . . . . . . . . . . . . . 149, 317 TEXTDOMAIN variable, BEGIN pattern and . . . . . . 318 TEXTDOMAIN variable, portability and . . . . . . . . . . 321 textdomain() function (C library) . . . . . . . . . . . . 315 tilde (~), ~ operator . . . . . 45, 52, 55, 108, 122, 123,

129, 132 time functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 time, alarm clock example program . . . . . . . . . . . 272 time, localization and. . . . . . . . . . . . . . . . . . . . . . . . . 317 time, managing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 time, retrieving . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 timeout, reading input . . . . . . . . . . . . . . . . . . . . . . . . . 84 timestamps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192, 193 timestamps, converting dates to . . . . . . . . . . . . . . 193 timestamps, formatted. . . . . . . . . . . . . . . . . . . . . . . . 225 tolower() function . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 toupper() function . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 tr utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 trace debugger command . . . . . . . . . . . . . . . . . . . . 338 traceback, display in debugger . . . . . . . . . . . . . . . . 335 translate string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 translate.awk program . . . . . . . . . . . . . . . . . . . . . . 275 treating files, as single records . . . . . . . . . . . . . . . . . 63 troubleshooting, --non-decimal-data option… 35 troubleshooting, == operator . . . . . . . . . . . . . . . . . . 122 troubleshooting, awk uses FS not IFS . . . . . . . . . . . 67 troubleshooting, backslash before nonspecial

character . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 troubleshooting, division. . . . . . . . . . . . . . . . . . . . . . 114 troubleshooting, fatal errors, field widths,

specifying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 troubleshooting, fatal errors, printf format strings

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 troubleshooting, fflush() function . . . . . . . . . . . 190 troubleshooting, function call syntax . . . . . . . . . . 126 troubleshooting, gawk. . . . . . . . . . . . . . . . . . . . . . . . . 447 troubleshooting, gawk, bug reports . . . . . . . . . . . . 442 troubleshooting, gawk, fatal errors, function

arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 troubleshooting, getline function . . . . . . . . . . . . 231 troubleshooting, gsub()/sub() functions . . . . . . 185 troubleshooting, match() function . . . . . . . . . . . . 182 troubleshooting, print statement, omitting

commas. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 troubleshooting, printing. . . . . . . . . . . . . . . . . . . . . . . 99 troubleshooting, quotes with file names . . . . . . . 101 troubleshooting, readable data files . . . . . . . . . . . 231 troubleshooting, regexp constants vs. string

constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 troubleshooting, string concatenation . . . . . . . . . 115 troubleshooting, substr() function . . . . . . . . . . . 185 troubleshooting, system() function . . . . . . . . . . . 191 troubleshooting, typographical errors, global

variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 true, logical . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 Trueman, David . . . . . . . . . . . . . . . . . . . . . . . . 6, 12, 425 trunc-mod operation. . . . . . . . . . . . . . . . . . . . . . . . . . 114 truth values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 type conversion. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112

U

u debugger command (alias for until) . . . . . . . . 333 unassigned array elements . . . . . . . . . . . . . . . . . . . . 161 undefined functions . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 underscore (_), C macro . . . . . . . . . . . . . . . . . . . . . . 316 underscore (_), in names of private variables . . 218 underscore (_), translatable string . . . . . . . . . . . . 318 undisplay debugger command . . . . . . . . . . . . . . . . 334 undocumented features . . . . . . . . . . . . . . . . . . . . . . . . 44 Unicode . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224, 424, 465 uninitialized variables, as array subscripts. . . . . 167 uniq utility. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 uniq.awk program . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 Unix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 Unix awk, backslashes in escape sequences. . . . . . 47 Unix awk, close() function and . . . . . . . . . . . . . . 104 Unix awk, password files, field separators and… 70 Unix, awk scripts and . . . . . . . . . . . . . . . . . . . . . . . . . . 19 UNIXROOT variable, on OS/2 systems . . . . . . . . . . 437 unsigned integers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 341 until debugger command . . . . . . . . . . . . . . . . . . . . 333 unwatch debugger command . . . . . . . . . . . . . . . . . . 334 up debugger command . . . . . . . . . . . . . . . . . . . . . . . . 335 user database, reading . . . . . . . . . . . . . . . . . . . . . . . . 238 user-defined functions . . . . . . . . . . . . . . . . . . . . . . . . 199 user-defined, functions, counts, in a profile …. 311 user-defined, variables . . . . . . . . . . . . . . . . . . . . . . . . 110 user-modifiable variables. . . . . . . . . . . . . . . . . . . . . . 147 users, information about, printing. . . . . . . . . . . . . 260 users, information about, retrieving . . . . . . . . . . . 238 USR1 signal, for dynamic profiling . . . . . . . . . . . . . 312

V

values, numeric . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 values, string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 variable assignments and input files . . . . . . . . . . . . 38 variable typing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29, 460 variables, assigning on command line . . . . . . . . . 111 variables, built-in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 variables, flag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 variables, getline command into, using … 79, 80,

82, 83 variables, global, for library functions . . . . . . . . . 218 variables, global, printing list of . . . . . . . . . . . . . . . . 33 variables, initializing. . . . . . . . . . . . . . . . . . . . . . . . . . 110 variables, local to a function . . . . . . . . . . . . . . . . . . 203 variables, predefined . . . . . . . . . . . . . . . . . . . . . . . . . . 147 variables, predefined -v option, setting with …. 32 variables, predefined conveying information … 149 variables, private . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 variables, setting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 variables, shadowing . . . . . . . . . . . . . . . . . . . . . . . . . . 200 variables, types of . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 variables, types of, comparison expressions and

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 variables, uninitialized, as array subscripts …. 167 variables, user-defined . . . . . . . . . . . . . . . . . . . . . . . . 110 version of gawk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 version of gawk extension API . . . . . . . . . . . . . . . . 153 version of GNU MP library . . . . . . . . . . . . . . . . . . . 153 version of GNU MPFR library . . . . . . . . . . . . . . . . 153 vertical bar (|) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 vertical bar (|), | operator (I/O). . . . . . . . . . 81, 128 vertical bar (|), |& operator (I/O) …. 82, 128, 306 vertical bar (|), || operator. . . . . . . . . . . . . . 125, 129 Vinschen, Corinna . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

W

w debugger command (alias for watch) . . . . . . . . 334 w utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 wait() extension function . . . . . . . . . . . . . . . . . . . . 402 waitpid() extension function . . . . . . . . . . . . . . . . . 402 walk_array() user-defined function . . . . . . . . . . . 246 Wall, Larry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159, 453 Wallin, Anders . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 warnings, issuing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 watch debugger command . . . . . . . . . . . . . . . . . . . . 334 watchpoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 wc utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 wc.awk program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 Weinberger, Peter. . . . . . . . . . . . . . . . . . . . . . . . . . 6, 425 where debugger command . . . . . . . . . . . . . . . . . . . . 335 where debugger command (alias for backtrace)

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 335 while statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 while statement, use of regexps in . . . . . . . . . . . . . 45 whitespace, as field separators . . . . . . . . . . . . . . . . . 68 whitespace, functions, calling . . . . . . . . . . . . . . . . . 175 whitespace, newlines as . . . . . . . . . . . . . . . . . . . . . . . . 36 Williams, Kent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 425 Woehlke, Matthew . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 Woods, John . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 425 word boundaries, matching . . . . . . . . . . . . . . . . . . . . 54 word, regexp definition of . . . . . . . . . . . . . . . . . . . . . . 54 word-boundary operator (gawk) . . . . . . . . . . . . . . . . 55 wordfreq.awk program . . . . . . . . . . . . . . . . . . . . . . . 279 words, counting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 words, duplicate, searching for . . . . . . . . . . . . . . . . 271 words, usage counts, generating . . . . . . . . . . . . . . . 278 writea() extension function . . . . . . . . . . . . . . . . . . 405

X

xgettext utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 XOR bitwise operation . . . . . . . . . . . . . . . . . . . . . . . 196 xor() function (gawk) . . . . . . . . . . . . . . . . . . . . . . . . 197

Y

Yawitz, Efraim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 427

Z

Zaretskii, Eli . . . . . . . . . . . . . . . . . . . . . . . . . 12, 426, 443 zerofile.awk program . . . . . . . . . . . . . . . . . . . . . . . 232 Zoulas, Christos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426